A Bibliography of Publications in Scientific Programming

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

15 January 2025 Version 1.17

Title word cross-reference

+ [BDV03]. 1 [TPKP13]. 2 [LZZ^{+}23b, PPJ^{+}15, fTPsLZ23]. 3 [JR10, MRLF12]. ^{3} [FP00a]. ^{C} [SAE^{+}23]. ^{H} [Che93a, Che93b]. ^{SM} [Ano07a]. _{2} [WCZ^{+}16]. α [CZB24]. C [JFD20]. d [TOF^{+}14]. γ [DCH^{+}23]. K [Ano23-31, RLC04, HKVVNT18, Hon23, LZSC23, SAZ^{+}24, WYY21]. N [AAS^{+}23, HJ96, SJD20, War14, WHG93]. O(N) [HJ96]. Z [DZL^{+}20].

-body [HJ96, War14]. -D [TPKP13]. -dimensional [WHG93]. -Expansion [CZB24]. -Gram [SJD20]. -Grams [AAS^{+}23]. -Mean [SAZ^{+}24]. -Means [Hon23, LZSC23, Ano23-31, JFD20, RLC04]. -Medoids [HKVVNT18]. -Number [DZL^{+}20]. -spline-based [TOF^{+}14].

1 [HR12a, KBRS95]. 19 [AKA^{+}23, DD23a, Guo23b, JAED23, NHHW23, TMHM23, WW23a, ZZDY24, ZWZ^{+}23b]. 1999 [Met99a, Met99b].

2 [GBJ94, HR12b]. 2.0 [HMSM08]. 20 [Met99a]. 200 [GBJ94, MJ95]. 2001 [Ger02]. 2003 [BCCP05, Mor15, RMX05]. 2003/2008 [Mor15]. 2008 [Mor15]. 2012 [VR13]. 2048-processor [HMCH07]. 2D [HLKY19, HJO^{+}20, TZH20]. 2HOT [War14]. 2nd [Nag05b].

3 [GBH14]. 3-Dimensional [WZ23d]. 3D [Ano23f, Ano23h, Ano23g, Ano23r, CTZC24, Chu21, Lia23b, Liu23e, LDZZ23, QWZP20, TLH17, WWW23, YYL17, Yan23g, ZWS19, ZYY^{+}23, ZZDY24]. 3D-HEVC [ZWS19]. 3Es [SZC16]. 3p [ZMM^{+}23].

4M [DTV00].

5-Step [SY20]. 590 [Was95]. 5G [Ano23-28, LGZ23, RLU^{+}20, XWL^{+}23]. 5G-V2X [XWL^{+}23].

61850 [CSWW23]. 64-Line [WLCC18]. 6G [JGDK23].

77 [OPE^{+}95, Zim07].

860 [CBCM93]. 8i [KGBB09].

90 [McC96]. 95 [RMX05, Sch03a]. 95/2003 [RMX05]. 9th [SO11].

A-Share [HS23c]. A. [SQJG23]. ABC [JFPL16]. Abdominal [DBWY24, He24, LCL^{+}23a, ZSZ^{+}23b, ZWW^{+}23b]. Abdominis [WLLY23]. Ability [Jua23b]. Ablation [CWMZ23, HZHW23]. Abnormal [Ano23i, FYC^{+}20, Ma23c]. Abstract [AB18, FXYY17, Vol97, Rou08b]. Abstracted [SSSJ20]. abstractions [Hav00b]. Accelerated [TKZ^{+}24]. Accelerating [BF17]. Acceleration [Jes10, ZSS^{+}10]. Accelerator [TLC15, CIN^{+}96]. accelerators [KK11]. Acceptance [Liu22, QQWJWYQZ23]. Access [GBH20, SJD20, SCD21, XWL^{+}23, CGL08, GL04, GBH14, LOHA01, TKS02]. Accessing [TC96]. Accident [XCW18]. According [ZXDY19]. Account [LFS^{+}20, YLW21]. Accounting [QHJ23, XZ23c]. Accuracy [AWB^{+}23, GMCT15, PDROFRM16, BBDN11]. Accurate [LZ23c, TZZ23, ZSD22]. Achieve [HS20]. Achieving [BAN02, HLPY16, KFFZ05]. Acid [LXL^{+}23]. Acoustic [MLL23]. Acquisition [OKHS23, ZGS16]. across [GR93, IPST18]. ACT [MCB^{+}20]. ACT-SVM [MCB^{+}20]. Action [Ano24, JZS21, JWPW24, KLuH^{+}20, LLL23c, PZ24, QS23, QH23, Tan23, WW23b, Wan23e]. Active [HWJ18, LWZ^{+}23c, YWGZ20, ZCL^{+}24a, ZSZ^{+}23a, RLC04]. Activities [SDM^{+}19, SCW^{+}23]. Activity [JYLL23, ZG15]. Actor [CSM15, KPS16]. Actor-Fact [CSM15]. Acute [CL23b, LLY^{+}23b, RJA^{+}23, XNL^{+}22, Xue23, XZLZ24, ZCL^{+}24a, ZLZ^{+}21]. Ad [JRS^{+}21, BDM^{+}04]. Adaboost [ZFX17]. Adam [JHS^{+}20, SK21]. Adaptability [Ano23-28]. Adaptation [LX23a, SRO^{+}15, IRSD99]. Adaptation-Based [LX23a]. Adaptive [ASZF23, AZGJ24, Ano23-40, CLDX20, CZKY24, CZCL23a, CLG^{+}23, GL23, GPP^{+}17, HKH^{+}17, HJO^{+}20, HHZ^{+}23, HHCC23, LLW23a, NTNTTK^{+}19, OQZ23, PVL^{+}04, QWP23, Shu94, SKHZ24, TS23a, WZG^{+}17, WZW^{+}20, WZF24, WMZM23, YWX23, ZDMY23, ZWZ20, ZZWL24, ZQL23, ZPA19, BCC^{+}93, BAN02, CW93, EAS^{+}97, GRN99, OHS00, VCT05]. Addiction [PP23, PZL19]. address [HUN08]. Addressing [TFZ^{+}20]. Adenocarcinoma [Ano23-29]. Adenomyosis [ZXZZ23]. Adhesion [LLQ^{+}23, WZ23d]. Adhesions [HSS23]. ADIFOR [BCC^{+}92]. Adipocytes [DCH^{+}23]. Adipose [WWH^{+}23]. Adjacency [ZSS^{+}10]. Adjacency-based [ZSS^{+}10]. adjoints [FC01, HUN08]. Adjustment [LBY^{+}23, QWP23, ZPA19]. Administration [Hu23a, Liu23b]. Adopted [ZWW^{+}23b]. Adoption [AUA^{+}20, AUA^{+}23, CLF^{+}23, HHZ^{+}23, JUA^{+}23a, LCW22, LX23d, NDSG07]. Advance [ZYZL23]. Advance-Selling [ZYZL23]. Advanced [AAH^{+}23, ATZ^{+}15, GMS^{+}23, RWY^{+}17, TY23]. Advances [GDNVSMM19, ZLX^{+}23, Spo12, HPS^{+}13]. Advantages [DF14]. Adversarial [BC23, HZFH23, ML23a, Xia23a, ZH23]. Adverse [Wan23f]. Advertisement [ZFX17]. Advertising [Zha23i, ZYZL23].

AEP [YWGL20]. Aerial [YLZ23, ZDMY23]. Aerobic [ZQ22]. Aerobics [Che23a, WW23b, XHX23, ZWF23, ZZSQ23]. AES [AHM23]. Affairs [HL23d]. Affecting [Wan21]. Affine [AZGJ24]. Affinity [LJZ18]. Affirmed [WCF^{+}21]. after [Ano23v, BLL^{+}23, CZFX23, CLW^{+}23, HWY^{+}23, HSY^{+}23, LHC^{+}23, LWZ^{+}23c, PP23, SWH23, Wan23f, WZZ23, ZRX^{+}24]. against [CGMRO^{+}19, Xu23e]. Age [SY23]. Agent [CSU^{+}17, RKA^{+}24, CJS^{+}02, CGK^{+}05, PVL^{+}04]. Agent-Based [CSU^{+}17, RKA^{+}24, CJS^{+}02]. Agents [YW24]. Aggregation [LLZW23, ZZWL24, IMB^{+}13]. Agile [AKA16, JUA^{+}23a]. AGNES [WYC^{+}20a]. AGNES-SMOTE [WYC^{+}20a]. Agonists [QYL23]. Agreement [Yi23]. Agricultural [SL21]. Agriculture [LYZY24, ZM23]. AGV [HQL^{+}23]. AGVs [ZCG23]. AHP [Wan23d]. AI [Ano23f, SGPCLG19, TMHM23]. AI-Based [SGPCLG19, TMHM23]. Aided [Che23a, Xia23b, IJL^{+}01]. AIoT [BDAH20]. AIoT-Based [BDAH20]. AIPS [WHG93]. Air [LBW^{+}24, OZI^{+}23]. Aircraft [CIU20]. Airport [KZZ^{+}21]. Airway [HSL^{+}23]. Alan [Ste97]. Alarm [CHL^{+}23, HZJ23]. ALE [MLYH17]. AlexNet [JHS^{+}20]. AlexNet-Based [JHS^{+}20]. algebra [ACIK97, HKL^{+}12]. Algebraic [YLLY23, ZRX^{+}24, DHH00, Hav00a]. Algorithm [ACLS16, AHM23, ALA20, AE19, Ano22b, Ano23k, Ano23p, Ano23s, Ano23n, Ano23-30, Ano23-31, Ano23-35, Ano23-47, AB18, BLZG23, BH23, BLL^{+}23, Bia23, BF17, CZKY24, CWMZ23, Cha23b, CLTX16, CCG17, CZD^{+}20, CLF^{+}23, CL23b, CD23, CZY^{+}23, Che23c, CTD23, CH23b, CPG^{+}23, CWS^{+}24, CTZC24, CZP^{+}24, CBE23, CLXZ23, DYYW24, DBWY24, DLP^{+}23, Don23, DFT^{+}23, FMA16, FZH23, FLR23, FWX23, FHC^{+}22, FLY^{+}23, FMG^{+}23, GWZ23b, GSL23, GZD23, HJ15, HI18, HWSL22, Han23, HSH23, HSS23, HZL^{+}23, HLHC23, He23, He24, Hon23, Hon24, HPZ17, HTZL23, HHZ^{+}23, HZHW23, HSY^{+}23, HS23b, JKR92, JLJ24, JYLL23, JZH^{+}23, JFD20, JL23b, JSY^{+}21, JFY23, JHL^{+}23, Jua23c, KS15, LGL23, LGJ18, LXWD18, LCW22, Li23c, LHMC23, LWZ^{+}23b, LZL^{+}23, LWL^{+}23c, Li23b, LSS^{+}23, LZZ23a, LBY^{+}23, LX23b, LLP23, LWL23b, LZSC23, LBW^{+}24, LZC^{+}24, LLY^{+}23a, LWW16, LLGL20, LHC^{+}23, eLA20, LLS^{+}20, LX23d, LLZW23, Liu23b, LLW23a, LWZ^{+}23c, LZWL23]. Algorithm [LJ23, LZLJ23, LLW23b, LZK^{+}23, LHZ^{+}23, Liu23c, Liu23g, LTL^{+}23, LMQ^{+}24, LLGH24, LQW23, LL24, LZZ19, MTK^{+}23, MUJN20, Men23, MWQ16, NHHW23, OWL^{+}23, OQZ23, PP23, PSW23, PCYZ18, PWY23, Qin23, QWP23, QFX^{+}23, RAZ^{+}23, STM^{+}23, SZW^{+}23, SAN^{+}19, SC23, She23, SLBZ16, SHCZ23, SL23a, SYW23, SKT^{+}24, SK21, SZC16, Su23, SWZL24, SCZ18, SQ21, TT16, TH18, TXD^{+}20, Ton20, Tun23, VVNHTHNT20, WS15, WWLG16, WQLR17, WLCC18, WYC^{+}20a, WH20, WCW^{+}21, WHL^{+}21, WCF^{+}21, WZS^{+}22, WW23b, WKY23, WZL23, Wan23h, Wan23d, WCZW23, Wan23a, Wan23c, WZZ23, Wan23i, WLZJ24, WZC^{+}23, WSL23, Wu23c, WJZC23, WDY23, WZ23d, Wu23b, WFZ^{+}23, XZZY20, XWG23, XYL^{+}23, yXHtJC23, XSG23, XHSL17, XSZ18, XY23b, Xu23c, XZ23b, Xu23f, XWLL23, XLLL23, gXySR^{+}17, XZLZ24, YY23a, Yan23b, YYL17, YLWW18, YXZ23, Yan23f, Yan23e, YYW^{+}23, YGZ^{+}24, Yan22b, Ye23, Yi23, YLZ23, YLLY23, YY23c, YW24, YPC^{+}23, ZYLZ15, tZhYsHlZ19, Zen22, ZDMY23]. Algorithm [ZCL^{+}24a, ZFX17, ZG18, ZLZ^{+}22, ZYZ23, ZLHZ23, ZXZZ23, ZZZ^{+}23, ZCLL23, Zha23c, ZLSW23, ZW23a, ZZD23a, ZZW^{+}23, ZFFP24, ZD23a, Zha23m, ZLZ^{+}23, ZSZ^{+}23b, ZQL23, ZZ23, ZHL16, ZL22, Zho22, ZTL^{+}23, ZWJN23, ZC23, ZZH22, ZHHX23, Zhu23a, KHSJ95, MJLM07, MSLS13, War14, ZSS^{+}10]. Algorithm-Based [Ano23-31, CWMZ23, CL23b, CTD23, CH23b, FZH23, FMG^{+}23, GSL23, HSS23, Hon24, HHZ^{+}23, HSY^{+}23, JLJ24, JFY23, LWZ^{+}23b, LZL^{+}23, LX23b, LBW^{+}24, LLY^{+}23a, LWZ^{+}23c, LZWL23, Liu23c, NHHW23, WHL^{+}21, WCF^{+}21, WZS^{+}22, WCZW23, WZZ23, WZC^{+}23, XYL^{+}23, XLLL23, ZCL^{+}24a, ZLHZ23, ZZZ^{+}23, ZCLL23]. Algorithm-Segmented [SWZL24]. Algorithmic [LZ23a]. Algorithms [AHMMÁR18, CS94, ÇÜ19, CBRRC^{+}16, DDB^{+}18, FLS17, GNKuR23, HL23a, Hua21, LTR^{+}23, MABU18, RuHMS24, RWY^{+}17, SAJ^{+}20, WWW^{+}17, XH20, tXChL^{+}24, YCWL17, YZWF23, ZY23, ZZW17, ZWXL21, ZW23c, ZG23, CW93, DRL13, GIKP95, GA96, HJ96, KTP05, KK09, Lin04, LKDB10, NKV^{+}02, OHS00, RLC04, SZ09]. Algorithms-Based [Hua21]. Alice [Nag04]. aligned [AGIS94]. Alignment [XYY^{+}20]. All-Subgraph [SJ15]. All-to-all [MJ95]. Alleviates [HSL^{+}23]. Alliance [BH23]. Allocation [BS01, HLPY16, JFPL16, JFD20, MPA16, STM^{+}23, SL23a, SLDL17, fXhCL17, YW24, ZCSL17, Mor94, RMX05]. along [WCW^{+}21, XZLZ24, ZY23]. Alternate [HEVD19]. Altimetry [LJSJ23]. Amazon [JMR^{+}11, TFN11]. Ambiguous [ABK^{+}23]. American [ZG15]. Amesos2 [BHRT12]. Amino [LXL^{+}23]. among [LZF16]. Amplicon [Ma22]. Amplicon-Based [Ma22]. Analgesic [Ano23-47, WLLY23]. Analog [SHW23]. Analysing [DYW^{+}23]. Analysis [Ano22a, Ano23f, Ano23h, Ano23j, Ano23e, Ano23c, Ano23d, Ano24, AZX^{+}17, BGZL23, BLZG23, BH23, CTCG15, CLHH20, CZY20, CWD23, CL23b, CCHJ23, CZCL23b, CLW^{+}23, CH23b, CZS^{+}23, Cui23, CFY23, DRMG^{+}18, DMV^{+}19, DLP^{+}23, DS23, FYY23, FC22, FHT23, FMG^{+}23, GWZ23b, GMM20, GPX^{+}21, GZD23, Han23, HPZ16, HLW^{+}18, HCC23, HBCM94, HEVD19, Hu23a, HW23b, HS23c, HXPZ24, JLJ24, JZS21, JFY23, KAS15, KZS^{+}20, KPS16, KOM94, KS19, LKS18, LPBSSE17, LQLX18, Li23c, LWZ^{+}23b, LZL^{+}23, LCK^{+}23, LBW^{+}24, LLLZ23, LW23, Lin23b, LZWL23, LMQ^{+}24, Lou23, LLGH24, LZ23d, LML^{+}23, MCZ23, MLZ23, MR18, MSZZ20, MLHC20, MAC^{+}20, NHHW23, OL23, Pan23b, PP23, PVCPdPSZVG17, PZL19, PRW23, PWY23, RBJ^{+}19, SO18, SLS17, SQJG23, SD23, SAZ^{+}24, SC16, SZL^{+}23, SYW23, SMD20, SK21, SW23, Son23b, SWZL24, SWY23, SWH23, TZH20, TTL^{+}23, TH23, VVDM18, WHL^{+}18, WFHC20, WQT^{+}20, WZ22]. Analysis [WZL23, WYHS23, WCZW23, WG23, WZC^{+}23, WmG21, WZ23d, yXHtJC23, tXChL^{+}24, XXSL23, Xu23a, XLLL23, XZLZ24, YZQ^{+}23, YG16, YYY23, YB23, YPM^{+}21, YT22, YHH^{+}23, YC23, YYZ^{+}17, YB24, YL20, ZCL^{+}24a, ZCZ21, ZSD22, Zha22, Zha23f, ZM23, ZCLL23, Zha23i, ZCZ^{+}23, ZLZ^{+}23, ZCT^{+}20, AE03, BCK07, BB08, BBF^{+}04, BCHL05, CO93, FO96, HMSM08, JMR^{+}11, LDV07, Lin04, PPS12a, PPS^{+}12b, PDGQ05, Rou08b, SCSJ09, SS00, SGM^{+}08, Sne95, TKS02, TF04, WGW08, GM14]. Analysis-Based [Han23, TTL^{+}23, ZCT^{+}20]. analysis-driven [Rou08b]. Analytic [LFQA20]. Analytical [AAA^{+}23, JUA^{+}23a]. Analytics [ABK^{+}23, CLL17, KGU^{+}20, KIU^{+}21, KGU^{+}21, NGMAA23, WLZ^{+}21, YSL^{+}18, ZGL18, hKJ^{+}20]. Analyze [BLL^{+}23]. Analyzed [LTL^{+}23]. Analyzing [HNSG23, LSB12, LLZW23, Ma22, WZY^{+}23, ZCXQ16]. Ancient [DYT23]. ANCS [HLPY16]. AND/ [yXHtJC23]. Android [Ano23a, AAF19, JMGH21]. Anesthesia [BGZL23, BLZG23, CZFX23, LWZ^{+}23b, LBW^{+}24, WCW^{+}21, YWGL20]. Anesthetics [GSL23]. Anesthetized [LLGH24, WHL^{+}21]. Aneurysm [BLL^{+}23, LHC^{+}23, ZZZ^{+}23, ZSZ^{+}23b]. Aneurysms [LZL^{+}23]. Angina [HL23a]. Angiography [DLP^{+}23, Hon24, SW24, WSL23, ZZW^{+}23, ZLZ^{+}21]. Angle [ZC23]. angular [KMB09]. Anisotropic [HLC^{+}24]. Ankle [Zha23f]. ANN [JAED23, JAED23, NLMT23]. ANN-Particle [JAED23]. ANN-PSOCoG [JAED23]. Annealing [BNH^{+}20]. Annotation [PMEH17, XXM^{+}23, ZHW^{+}21]. Annotation-Based [ZHW^{+}21]. Anomaly [JGF23, LZK^{+}23, WYY21]. ANOVA [LYZ^{+}23, YHZG23]. ANOVA-Based [YHZG23]. ANP [Wen23]. Answer [KIU^{+}21]. Answering [WLZ^{+}21]. Answers [SHM97]. Ant [FTS^{+}20, FMA16, QHAN23, WCLL17]. Anterior [LZWL23]. Antibodies [XYL^{+}23]. Antiphishing [FA24]. Antispam [PDROFRM16]. AntStar [FMA16]. Aortic [ZSZ^{+}23b]. Aperture [LL23a, tZhYsHlZ19, PTS^{+}13]. API [UJS^{+}20]. Apoptosis [ZMM^{+}23]. App [AAA^{+}23]. Apparatus [CLW^{+}23]. Apparel [Ano23g]. Appearance [ZLS^{+}17]. AppLeS [COB^{+}00]. Application [Ano23f, Ano23h, Ano23j, Ano23k, Ano23m, Ano23l, Ano23i, Ano23-41, ATZ^{+}15, BCS01, Cha23b, CZD^{+}20, Che22, CZCL23b, CH23a, CWS^{+}24, CSM^{+}18, DYYW24, DZL^{+}20, FTS^{+}20, Gog23, HI18, He24, HYK^{+}20, HY23, HMQS23, JZ22, JP23, JZY23, KJG^{+}08, KIU^{+}21, LWL^{+}23a, Li22, LHMC23, LLL^{+}23a, Li23e, LWZ24, Lin23a, LWQ^{+}23, Liu23e, LL23d, LLY23c, LML^{+}23, Mal01, NTNTTK^{+}19, PPS^{+}12b, SO18, She23, SXP23, SH19, VVNHTHNT20, WH23, WJS^{+}23, Wu23c, WG24, Xia23a, XY23b, YTD15, Yan22a, YL23, YWX23, ZXZZ23, ZFFP24, yZcR23, ZLF^{+}22, BBF^{+}04, BFGH14, BBK^{+}11, DFP96, HLM^{+}05, IJL^{+}01, KeKR^{+}11, Bor04, WG21].

Applications [AK18, Alt23, Ang93,

ACMT19, CHM^{+}97, CLPVPI20, KPW^{+}18, LYTL20, LWYS16, MLHC20, RRM^{+}15, RLU^{+}20, SMA^{+}23, Sne95, Wan23g, WW23c, Zha23d, BBBM12, BDV03, CHS^{+}99, CKS^{+}09, HZ99, HHML05, LTB02, LPHD04, Nor07, PVL^{+}04, RKR^{+}99, SJR14, SKU^{+}09, Slo12, SB01, SBJV11, WGW08]. Applied [BL23, LD23, SMD20, ZWW^{+}23a, ZQL23, WGF93]. Applying [AlS16, CTS23, LFS^{+}17, PDROFRM16, SCB02]. Approach [ASS23, AXA20, AlS16, AUA^{+}20, AKA16, AK23, ABK^{+}23, BKMS23, BC23, CLA^{+}15, CGMRO^{+}19, CZS^{+}23, DDMS15, DEŁ01, FH23, FOT92, HGMM23, HLY^{+}17, HFH^{+}23, HKQ^{+}23, JUA^{+}23a, JNK24, JGF23, Jua23b, KGU^{+}20, KIU^{+}21, KGU^{+}21, LGC^{+}16, LWYS16, LG18, LLW^{+}22, LFQA20, NTNTTK^{+}19, Ona19, OZI^{+}23, PLY^{+}23, PVCPdPSZVG17, PDRC18, PPB^{+}19, RLU^{+}20, REPRL^{+}19, STM^{+}23, SPPH15, SSN15, SZX^{+}20, lWwFAH23, YZ16, ZWZ16, ZW23c, FP95, GG95, KJG^{+}08, KLS^{+}96, LSB12, LB02, PSU08]. Approach-Based [BKMS23]. Approaches [CPD21, DDB^{+}18, OB96]. Approximate [AMU^{+}23, LLL^{+}23b, WX23, GG05]. Approximated [PY17]. Approximation [HEVD19, YTD15]. APPs [CYJ^{+}20, ZYX^{+}21]. Apriori [JNK24]. Aptitude [KJA^{+}23]. AQUACOAST [MVIA20]. Aquifer [BRPV24]. Arabic [AWB^{+}23]. arbitrary [CGL08]. Arc [NYML17]. Arc-Length [NYML17]. Architecting [IR02]. Architectural [GWH^{+}23]. Architecture [AKA16, Ano23y, ARAM19, CL23c, CZS^{+}23, CSU^{+}17, JGDK23, KV23, KGMN15, LD23, ML23a, Ma23d, NLPL^{+}20, WZW^{+}20, ZZS^{+}23, BHK^{+}13, PMTL12, RX04, Rou08b, WB95, KCO^{+}05]. Architectures [CCTY15, CKKK15, KGIL19, BFGH14, BBK^{+}11, HPD09, LKDB10, SH94, SJR14, WSP94]. Area [LF23b, fTPsLZ23, XZNL23, ZFFP24, JEM07]. Areas [NHHW23, WZY^{+}23]. ARIMA [ZZWZ23]. arising [ARvW03]. Aristotle [FSV14]. arithmetic [Sch03a]. Arm [Qin23]. ARMS [CJS^{+}02]. Array [HBCM94, McC96, ESP^{+}12, Ott93, WHG93, BCS01]. arrays [FO96, RN07, TC96]. Arrhythmia [MLY23]. Arrival [AMB23]. Arrow [AN19]. Arrow-Connected [AN19]. Art [Ano23-41, Bry96, CPK19, SY23, Tia23, WX23, Wu23c, XL23, ZJ23, Zha23k, BDH^{+}10]. Arthritis [GLD^{+}24, Li23d, YPY23]. Arthritis-Related [Li23d]. Arthroplasty [LWZ^{+}23b]. Artificial [BGZL23, BL23, BLZG23, CZFX23, Che23a, Chu21, Dai23, Du23a, Du24, HLHC23, JL23a, LCW22, LWZ^{+}23b, LJL23, LBW^{+}24, LZ23b, Liu23g, LLY23c, MABU18, QK21, SY23, SCZ18, TY23, Wan23g, XY21, XY22, ZJ23, ZGL23, ZSZ^{+}23b, ZCL^{+}24b, Gil09, ZFFP24]. Artistic [Xia23b]. Arts [LLY24]. ASKALON [Pro07]. ASM [FZ23]. Aspect [WLZ^{+}21, YSZ^{+}23]. Aspect-Based [YSZ^{+}23]. Aspects [KAS15]. Aspen [SV15]. Aspiration [JLJ24]. Assembly [JBLB15]. Assess [ÁRCPS15, HMQS23]. Assessing [LLW^{+}14, MVIAM20, VVNHTHNT20]. Assessment [AFB23, BBI23, CZWY23, HZJ23, Jua23b, LL23c, LCH^{+}23, TYCL23, Wan23f, WYC20b, YG16, ZJZ23, ZWLL23, MCvM10]. Assigning [HY15, HKQ^{+}23]. Assigning-Weight [HY15]. Assignment [QYW^{+}17]. Assistance [Cao23b, Wu23a]. Assistant [QYW^{+}17]. Assisted [CPG^{+}23, Dai23, DS23, Liu23c, LL24, Ma23d]. Association [DRMG^{+}18, HLY^{+}17, ZYLZ15]. Association-Oriented [HLY^{+}17]. Assurance [SNH24]. Asthma [HSL^{+}23]. Asymmetric [RQD^{+}20]. Asymptotic [AMU^{+}23]. AT-BiGRU [LZ23d]. Ataxia [CZY20]. Atherosclerosis [LXRC21].

Atherosclerotic [Ano23-31]. Athlete [HH23b, LH23b]. Athletes [JWPW24, XHX23, ZZSQ23]. Atlas [CWLW23]. Atmospheric [CHZY17, OZI^{+}23, Ros00, SAN^{+}19, WFZ^{+}23, SS00]. atomistic [ADLN08, NKV^{+}02]. atomistic-to-continuum [ADLN08]. Atopic [Xu23e]. Atrial [CTZC24, LLY^{+}23a]. attachment [MPP^{+}04]. Attack [JRS^{+}21, KAA^{+}20, LL23c, WDD^{+}23]. Attack-Resistant [JRS^{+}21]. Attacks [CGMRO^{+}19, Pan23b]. Attention [Ano23-38, CLDX20, DyLJ23, FDY^{+}19, LZWJ23, Pan23a, PRW23, QFYY20, SLC^{+}23, SW23, WM23a, Yua20, Zha23b, Zhu23b]. Attention-Based [FDY^{+}19]. Attentional [GCB^{+}23]. Attentive [KYR^{+}23]. Attitude [HZ24, WW22, WY23]. Attitudes [SXTK18]. Attribute [XWG23, ZH19, ZL22]. Attributes [ZY24]. Audio [Sun23a]. Audit [BH23]. Auditing [MKS^{+}24]. Augmented [SAE^{+}23]. Augmenting [BC23]. Authentication [HCGGSS17, TXD^{+}20]. Author [Ano97a, Ano99a, Ano00, Ano08a, Ano09, Ano11a, Ano12, Ano14]. Authorship [ABR17]. auto [DF14, FML^{+}14]. auto-tuning [DF14, FML^{+}14]. Autocontouring [GL23]. Autoencoder [FC22, GLH23, WG23]. Autoencoders [WCR^{+}21]. Automated [SV15]. Automatic [AIA^{+}21, CSM^{+}18, EOS17, FHT23, FB99, GA18, HZ99, HX23, HCQ17, HFH^{+}23, Hua23, JJY^{+}03, KYR^{+}23, LX23a, LCL^{+}23a, LWQJ24, Luo23, MqWp23, NL19, PST19, QK21, SDM^{+}19, SC23, SLD^{+}16, WYY21, WWW23, XLY18, YJ23, ZYZ23, ZL23b, Zha23m, AGG^{+}97, BFGH14, IJL^{+}01, Kes96, LP99, OPE^{+}95]. Automatically [CBRRC^{+}16, HPGC19]. Automating [CQF05, PPS12a, PPS^{+}12b]. Automation [MAA^{+}21, SCKW23, HMSM08]. Automotive [HLY20]. Autonomic [KeKR^{+}11]. Autonomous [Ano23w]. Autopsies [DRMG^{+}18]. Autoregressive [ZCW^{+}19]. Autoscaling [YPM^{+}21]. Auxiliary [Ano23-37, ZYL23]. Available [RRV09]. AVBH [RQD^{+}20]. Aversion [LQLX18, ZLL20]. Avionics [GLBB17]. Avoiding [XWLL23, YZX17]. Aware [BHP^{+}03, CGC16, CLYL16, CPK19, FZW^{+}18, HS20, LG18, MJB15, OCC^{+}16, PY17, PBY^{+}20, RKA^{+}24, SMB17, TT16, TLW^{+}23, Xie20, ZCSL17, HJYC10, IMB^{+}13, Zim07, KKK^{+}18, ZWL^{+}19]. Awareness [LZC^{+}24, LNH^{+}20, ZHL16]. AWGN [AE19]. AWS [JMR^{+}11]. AWSMOTE [WZF24]. Axis [MTK^{+}23, ZMM^{+}23]. Azimuthal [ML24]. Azure [HLM^{+}11].

B [LWZC21, LMQ^{+}24, Ano23y, PDRC18, PZC^{+}23, ZCZ^{+}23]. B-Mode [LWZC21, LMQ^{+}24]. B.E. [Ano08b, VvAC^{+}09]. B/S [Ano23y]. B2C [AA23]. Babb [Per08]. Background [Hu23a, HX23, LZ23b, Liu23d, Lou23, LQW23, fTPsLZ23, WWYR22, Yan23e]. Backprojection [DBWY24, PWY23, PTS^{+}13]. backprojection-based [PTS^{+}13]. Backpropagation [LWZ^{+}23b, Pen23, ZLHZ23]. Backstepping [HHCC23]. Backup [CTCG15]. Badminton [WY23]. Baguatian [ZJZ23]. Balance [CKKK15, HSL^{+}23, PY17, ZCCS23]. Balanced [BGZL23, SL23a, XSG23]. Balancing [AFU18, CGC16, VMRM16, Xie20, LTB02, PVL^{+}04]. Ballooning [PGP19]. BAM [Ma22]. Band [WZY^{+}23]. banded [MSSG11]. Bandwidth [WNL^{+}20]. Bank [LL23d]. Banking [SWZS17]. Banks [MSZZ20]. Barriers [AAA^{+}23]. bartering [Özt04]. Based [ASS23, AKA^{+}23, ACA19, AN19, AXWS23, Ano22b, Ano23f, Ano23g, Ano23m, Ano23e, Ano23l, Ano23o, Ano23p, Ano23s, Ano23q, Ano23a, Ano23t, Ano23r, Ano23n, Ano23v, Ano23w, Ano23x, Ano23u, Ano23y, Ano23z, Ano23c, Ano23d, Ano23-30, Ano23-28, Ano23-31, Ano23-32, Ano23-33, Ano23-46, Ano23b, Ano23-41, Ano23-36, Ano23-37, Ano23-42, Ano23-38, Ano23-40, Ano23-45, Ano23-48, Ano24, ABR17, ACBV20, ATZ^{+}15, AA23, AB18, AUA^{+}23, AAA^{+}23, AAF19, AZX^{+}17, BGZL23, Bai23, BDAH20, BBI23, BD16, Bia23, BKU^{+}23, BKMS23, Cao23a, CC23, CZKY24, CWMZ23, Cha23b, Cha23a, CLA^{+}15, CLTX16, CHZY17, CSZ^{+}18, CZD^{+}20, CZFX23, CWD23, CL23b, CCHJ23, CSWW23, CD23, CZ23, Che23c, CYLD23, CTD23, CLW^{+}23, CWSF23, CWH23, CH23a, Che23a, CT23, CHL^{+}23, CH23b, CCM^{+}23, CFL^{+}23, CYSZ23, CTZC24, CL23c, CHTZ23, CLYL16, CZWY23, CW23, Cui23]. Based [CFY23, CSU^{+}17, DYYW24, DBWY24, DyLJ23, DRQ^{+}18, DWZ^{+}20, DPMB16, DCF17, Don23, DM23, DZL^{+}24, DFT^{+}23, DYT23, Du23a, Du23b, DD23b, DW23, FTS^{+}20, FA24, FDY^{+}19, FYH^{+}19, FYY23, FX23, FZH23, FSLG17, FHS^{+}23, FZW23, FWX23, FYC^{+}20, FYZ^{+}20, FC22, FHT23, Fen23b, Fen23a, FGSCR^{+}17, FZW^{+}18, FLY^{+}23, FMG^{+}23, FZ23, GAH^{+}20, GWZ23a, GSL23, Ge23, GLD^{+}24, GLH23, GMM20, GPX^{+}21, GN22, GZD23, Guo23a, HJ15, HLKY19, HBR^{+}20, HLL23, Han23, HSH23, HLW^{+}18, He20, HCC23, HSS23, HWY^{+}23, HZL^{+}23, HLHC23, He23, He24, HLC^{+}24, Hon23, Hon24, HY15, HYK^{+}20, HT22, Hu23a, HL23c, HH23a, HHZ^{+}23, HZFH23, HX23, HY23, HHD^{+}16, HWJ18, Hua21, HFH^{+}23, HZHW23, HSY^{+}23, HLH23, HW23a, HZ23, Hua23, HXPZ24, HZ24, Hui22, Hui23, HKQ^{+}23, JK22, JLJ24, JYLL23, JFPL16, JHS^{+}20, JZL^{+}20, JZS21, JW23a, JSY^{+}21, JFY23, JZY23, JLX^{+}20, Jua23c, KS15, KL23, KSYC23, KLK^{+}20a, KLK^{+}20b, KZZ^{+}20, KAA^{+}20]. Based [KKA^{+}22, KZK^{+}23, KJA^{+}23, KV23, KPS16, KZZ^{+}21, LGL23, LKS18, LWZ^{+}23a, LPBSSE17, LPBSSEPL19, LGJ18, LXWD18, bLwHcY18, LY20, LPD20, LXRC21, LLW^{+}22, Li23c, LHMC23, LWZ^{+}23b, LZL^{+}23, LX23a, LLQ^{+}23, LD23, LSS^{+}23, LZZ23a, LBY^{+}23, LXL^{+}23, LX23b, Li23d, LLP23, LWL23b, Li23f, LMZP23, Li23a, LF23a, LL23b, LZSC23, LCK^{+}23, LGZG23, LBW^{+}24, LLY24, LCH^{+}24, LZC^{+}24, LLLZ23, Lia23b, LY23b, Lia23a, LLY^{+}23a, LW23, LCL^{+}23a, LZ23b, LJZ18, eLA20, LLY^{+}20, Liu20, LSY20, LLZW23, Liu23f, LWW23a, LDZZ23, LWZ^{+}23c, LYL^{+}23, LZWL23, Liu23d, LZLJ23, Liu23a, LLL23c, LGZ23, LLW23b, LZK^{+}23, LHZ^{+}23, Liu23c, LLC^{+}23, Liu23g, LDW23, LYZY24, LYL^{+}24, LLGH24, LH23b, LZ23d, LJG^{+}23, LL24, Luo23, Lv23, LLT23, MCB^{+}20, Ma22, Ma23c, Ma23a, ML23a, Ma23d, MCZ23, MLZ23, MFB^{+}15, MPA16, MAS^{+}23, Men23, MLY23, MLHC20, ML23b, ML24, MCF^{+}23, MABU18, MLAS21, MM16, NHHW23, Ona19, OWL^{+}23, PWP21, Pan23a, PLJ^{+}23, Pan23b, PP23, Pen23, PZC^{+}23]. Based [PZ24, PCYZ18, QK21, QYL23, QHJ23, QHAN23, QWZP20, Qin23, QH23, QFYY20, QWP23, QGS23, Qiu23, QL23, RKA^{+}24, RWN20, RAZ^{+}23, RAdASGR^{+}19, REPRL^{+}19, SKB^{+}24, SDM^{+}19, SZW^{+}23, SPPH15, SGK^{+}23, SJD20, SSN15, SD23, SX23, SLH23, SLBZ16, SZX^{+}20, SFZ^{+}20, SZL^{+}23, SL23a, SYW23, SW24, SLP18, SZXL16, SW20, Son20, tS23b, SW23, Son23b, SGPCLG19, SH19, SF23, Su23, SCD21, SHW23, Sun23a, Sun23b, SL23c, Sun23c, SCKW23, SL23b, SNH24, THS^{+}15, TZF^{+}18, TZH20, TZZ23, TZCZ23, TZDZ20, TY23, TLZ^{+}23, TS23a, TZL^{+}24, TMHM23, TH18, TL17, TMY^{+}23, TF23, TTL^{+}23, TZC23, TT23, TXD^{+}20, TYCL23, Tun23, UAU^{+}23, WS15, WB16, WZG^{+}17, WXZW17, WWZ17, WHL^{+}18, WLCC18, WPDZ19, WYC^{+}20a, WFHC20, WDZY20, WH20, WCR^{+}21, WYY21, WHL^{+}21, WLP^{+}21, WCF^{+}21, WZS^{+}22, WZ22, WLJ22, Wan22, WW23b, WKY23, Wan23g, WZL23, WY23, WWW23, WL23b, Wan23f, WZ23a, WWH^{+}23]. Based [WYHS23, WZML23, Wan23b, Wan23d, WZY^{+}23, WCZW23, Wan23a, Wan23c, WZZ23, WZ23b, WW23a, WL23a, WM23b, WLQ23, Wan23i, Wan23e, WC23a, WLL^{+}23b, WG23, WZF24, WDD^{+}23, WZ23c, WZC^{+}23, Wei23, WZF^{+}20, WLZ^{+}21, WG22, WDY23, WZ23d, WC23b, Wu23a, XCW18, XY23a, XH20, XY21, XY22, XNLJ23, XLB^{+}23, XL23, XYL^{+}23, XSG23, XSS^{+}23, XBG17, XWL19, XTCC15, XML^{+}16, fXhCL17, XLY18, XYY^{+}19, Xu20, XXSL23, Xu23a, XYCC23, Xu23c, XWL^{+}23, Xu23b, XLLL23, XZNL23, XXM^{+}23, XC23, YZL23, YX23, YY23a, Yan23b, YZZ15, YZ16, YG16, YYL17, YCWL17, YLWW18, YXW^{+}22, Yan23f, Yan23d, Yan23e, Yan23c, YY23b, YHZ^{+}23, YB23, YGZ^{+}24, YDC^{+}24, Yan22b, Yan23h, YWGZ20, YT22, YSZ^{+}23, Ye23, YC23, YLLL20, Yin23, YT23, YB24, YZX17, YL20, Yu22, YNK23, YLLY23, Yua20, YZ20, YHZG23, YPC^{+}23, YN23, Zen22, ZCL^{+}24a, ZHZF20, ZCXQ16, ZWW16, ZFX17, ZLS^{+}17, ZCW^{+}19, ZWS19, ZLLS20, ZLLZ20, ZCL^{+}20, Zha20b]. Based [ZWG^{+}20, ZCZ21, ZQ22, Zha22, ZHY22, ZZWZ23, ZCG23, ZL23b, ZLHZ23, ZHL^{+}23, Zha23f, ZXZZ23, ZCS23, Zha23e, ZGFL23, ZZZ^{+}23, ZJ23, ZCLL23, Zha23c, ZH23, Zha23j, Zha23i, ZLWW23, Zha23g, ZSZ^{+}23a, ZLSW23, ZJZ23, ZW23a, ZZD23a, ZRX^{+}24, ZWCY16, ZWZ20, Zha23k, ZWZ23a, ZL23c, ZD23a, ZY23, ZYL23, Zha23m, ZWF23, ZLZ^{+}23, ZS23, ZW18, ZGY^{+}20, ZCL^{+}24b, ZY24, ZHW^{+}21, ZL22, ZZW22, ZW23c, ZWJN23, Zho23, ZW23b, ZLX^{+}23, ZHHX23, ZWLL23, ZZSQ23, ZSZ23c, Zhu23b, Zhu23a, ZCT^{+}20, ZGT^{+}23, Abd02, ABB^{+}02, AMM05, BMY13, BB08, BGLR93, CJS^{+}02, CL04, GRN99, GGJ04, GG05, HW23b, KK11, Lan01, OPP11, PTS^{+}13, PPS12a, Sch03b, TOF^{+}14, TKS02, Tým99, WOC99, ZSS^{+}10, SAN^{+}19]. Basic [BBG^{+}93]. Basin [FYH^{+}19, WZY^{+}23]. Basing [CLHH20]. Basis [MCB^{+}20]. Basketball [DZL23, HX23, LLL23c, WH20, WWYR22, Yan23d]. Batching [FH23]. Battery [CHTZ23, DZ18]. Bay [LF23b]. Bayes [CCM^{+}23, XY23a, YHZ^{+}23]. Bayesian [YL20, DRMG^{+}18, LB02, SLS17, WXL19, XWG23]. Bayesian-Based [YL20]. BE [RRV09]. Bearing [Sah20]. Bedrooms [GWH^{+}23]. Bee [HLHC23, LBW^{+}24, Liu23g, SCZ18]. before [Ano23v, BLL^{+}23, LHC^{+}23]. Behavior [Ano23c, Ano23-43, Ano23-48, CD23, LLW^{+}22, LWX23b, PZL19, SJD20, TZCZ23, TLZ^{+}23, XHC^{+}23, tXChL^{+}24, XC23, YDC^{+}24, YC23, TKS02]. Behavioral [LQLX18]. Behaviors [lWwFAH23]. Behaviour [IPST18, JLX^{+}20, GGJ04]. Beijing [ZFFP24]. Beijing-Tianjin-Hebei [ZFFP24]. being [BKP99]. Belief [Lia23a]. Belos [BHRT12]. Benchmark [Was95, WG21, WG24, DEvdV01, KGBB09, Mül03, PHH95, RST02]. Benchmarking [Gog23, RST02]. Benchmarks [PAR94, AE03, HB94, SBM^{+}10]. Benefit [BH23, CWD23, HLY20]. benefits [NJ08]. Benign [LX23d, Yin23]. BERT [LWL23b]. Berth [LFS^{+}20]. besides [ZWZ^{+}23b]. Best [LWL^{+}23a, Old00]. Bethesda [YLL^{+}23]. better [BPC^{+}09, GA96]. between [AK18, KZK^{+}23, LYL^{+}24, SH94, Son23a, Sun23c, SWH23, WWH^{+}23, WYHS23, WZZ23, WLLY23, WZC^{+}23, WWYR22, YWGL20, Zha20a, ZZH22]. Beyond [SF03a, Lin04]. Bi [HSL^{+}23]. Bibliometric [XZ22]. Bibliometrics [CCHJ23]. biclusters [HB12]. Bidimensional [ML23b]. Bidirectional [JKA^{+}23, LXL^{+}23, ZLDX21]. Biermann [Ste97]. Bifida [CTD23]. Big [AK23, Ano19, Ano23e, AZX^{+}17, Bai23, CTC^{+}15, CLL17, Cha23a, CNW^{+}23, CFY23, DS23, FWX23, FAA^{+}20, FC22, FYH23, HYK^{+}20, HL23d, Hui23, JWPW24, KGU^{+}20, KIU^{+}21, KGU^{+}21, KPW^{+}18, KKK^{+}18, LJF20, LD23, LLY24, Lia23a, LFQA20, Lin23a, Lin23b, LJZ18, LLZW23, LYZY24, MLHC20, MABU18, NGMAA23, NLPL^{+}20, PCYZ18, RMS21, SLP18, Son23a, SL23b, Tia23, UAK18, WJCZ21, WG23, WWL^{+}23, Wei23, WLZ^{+}21, XCW18, XHX23, Xu23c, YCWL17, YYY23, YB23, ZG18, ZZD23a, ZGL18, ZL23c, ZX17, ZTLH18, Zhu22, ZZSQ23]. Big-Data-Driven [Zhu22]. BiGRU [YZQ^{+}23, LZ23d, QFYY20]. Bike [Son23a, TFZ^{+}20]. Bilateral [LLL^{+}23b, XLLL23]. Bile [CZFX23]. Bilevel [fXhCL17]. Bilingual [ABNJMB23, LY20]. BiLSTM [CFL^{+}23]. BIM [Du23a, Du23b, LY23b]. Bin [BDAH20]. Binary [XSG23, YDC^{+}16, LSB12]. binding [VRW^{+}03]. Biobjective [CMTAC^{+}16]. Biodiesel [ZJZ^{+}16]. Biogeography [ASS23, PCYZ18]. Biogeography-Based [ASS23, PCYZ18]. Bioinspired [KAA^{+}23]. Biological [AHCZ12, HH23b, ZRP08]. Biomedical [BZKA20, SZX^{+}20]. Biometric [Guo23b, YHZ^{+}23]. biomolecular [BPK00]. Bipartite [WXL19, XWL19]. Bird [KZZ^{+}21]. Birds [SCAP16]. Birthmark [GNKuR23]. BIS [YWGL20]. Bispectral [LLGH24]. Bispectral-Based [LLGH24]. Bit [RQD^{+}20]. BitTorrent [DRL13]. Black [YDC^{+}16]. Bladder [HHZ^{+}23, LJ23]. BLAS [ARvW03]. Blended [qCtWfPbZ23, Hui22, XHC^{+}23]. Blending [HAW23]. Blind [Hua23]. Blitz [AJJF14]. Block [Ano23-47, GSL23, HHZ^{+}23, WCW^{+}21, ZYY^{+}23]. Block-Matching [ZYY^{+}23]. Blockchain [CHTZ23, Ge23, HAA^{+}22, KL23, KV23, Lvp23, Ma23a, RuHMS24, SCD21, TKKZ23, Xu23d, Yi23, ZCCS23]. Blocks [WLLY23]. Blogosphere [ZG15]. Blood [BLL^{+}23, FMG^{+}23, GAH^{+}23, LHC^{+}23, ZCLL23, ZZH22]. Blueberry [QWZP20]. Bocca [ANEA08]. bodies [AGIS94]. Body [CFY23, LZLZ19, WBZL17, WZC^{+}23, HJ96, War14]. Bolting [KKA^{+}22]. Boltzmann [TLH17, XSZ18]. Bone [DCH^{+}23, LYZ^{+}23, YHZG23]. Book [Ano07a, Ano07b, Ano07c, Ano11b, Bai08, Bor04, Bro11a, Bro11b, Bro12, Bro09, Bus09, CZH24, Fil14, GB10, Les10, Man08, Nag04, Nag05a, Nag05b, Nag09a, Nag09b, Nag11a, Nag11b, Nag12]. books [Met99a, Met99b]. Boosting [KAA^{+}23, PDROFRM16, Wu23c]. Boot [HHD^{+}16]. Border [Ano23-30, Liu23f, YZL23]. Borrower [LWT18]. Borrowing [LWT18]. Both [ZHL16]. Boundary [XWLL23]. Bow [SLS17]. Bow-Tie [SLS17]. BP [HLHC23, MCZ23, TS23a, Wan23h, Ye23]. BPMN [LY23a]. BPSO [FZW^{+}18]. Brachial [GSL23]. Brain [KSYC23, MUJN20, WMZM23, Yan23g, YT22]. Brain-Computer [MUJN20]. Branching [Ser18]. breadth [BAP13]. breadth-first [BAP13]. Breakdown [WKW^{+}23]. Breaking [HL23b]. Breast [LTR^{+}23]. Breeding [JZH^{+}23]. Bridge [Che23c, OPB^{+}20, YGZ^{+}24]. Brightness [LCH^{+}23]. bringing [Pla04]. Broadband [GGP09, IB09, KD09, SKU^{+}09, VGC09, WSZK09, WC23b]. Broadcasting [WNL^{+}20]. broadcasts [DRL13]. Broker [ACB^{+}02]. Bronchial [HSL^{+}23]. Brownian [TXD^{+}20]. BSML [VRW^{+}03]. BSP [SHM97]. Buckling [NLMT23]. Budesonide [Zha23a]. Buffer [YJ23]. Bug [YZX17]. Build [XZ23c]. Building [Ano23r, CYZ^{+}20, CYJ^{+}20, FX23, VvAC^{+}09, YB24, ZYX^{+}21, RLL^{+}02]. Buildings [WG21, WG24]. Bulk [HNSG23, ZWCY16]. Buoy [XZNL23]. Burgers [AMU^{+}23, MRLF12]. burst [BCHL05]. Bus [MM16]. Business [ARGC19, Hu23a, HW23b, KF20, LZYJ20, LWX23b]. butterfly [Sch03b].

C [CZP^{+}24, HHZ^{+}23, SWY23, Ang93, CO93, Che93a, Che93b, CP03, DNS97, DDMS15, LH93, LX23d]. C-Means [CZP^{+}24, HHZ^{+}23, LX23d]. C-Reactive [SWY23]. Cache [AlS16, NsP16, TLC15, Bai95, WC96]. caching [Tým99]. CAF [Mar05]. CaKernel [BBK^{+}11]. Calculation [FCZ^{+}17, LZ23c]. Calculus [LCK^{+}23, RMX05, Rou08b]. Calibration [CSZ^{+}19, fTPsLZ23, XLF^{+}24]. Calisthenics [WH23]. Call [Ano97b]. Called [Tun23]. Calls [UJS^{+}20]. Camera [ACLS16, KLK^{+}20a, SHCZ23, fTPsLZ23, XLF^{+}24]. Camera-Based [KLK^{+}20a]. can [Ang93, BFH96]. Cancer [Ano23v, Ano23u, CWMZ23, CZD^{+}20, CLF^{+}23, DBWY24, HLL23, HZHW23, KAA^{+}23, LWL^{+}23c, LSS^{+}23, LWQ^{+}23, LZZ^{+}23b, LJ23, LLC^{+}23, LTR^{+}23, SSY^{+}23, TY23, WZC^{+}23, XYL^{+}23, XZ23b, YLLY23, YHZG23, ZYZ23, ZCL^{+}24b, ZTL^{+}23, AMN^{+}12]. Cannula [TY23]. Canonical [ML23a, TZH20]. CanonicalProducer [KCO^{+}05]. CAP [FA24]. Capabilities [LZYJ20, MRLF12, PPS12a, PPS^{+}12b]. Capability [FHT23, HLS^{+}24, WmG21]. Capital [YP24]. Capsule [NL23, ZC20]. Capsules [SYW23]. Capture [WW22]. Capturing [FSLG17]. Car [Ano23-44, LJZ18]. Car-Sharing [LJZ18]. Carbon [CWD23, ZFFP24]. Carcinoma [CLW^{+}23, FLY^{+}23, PTZL23, WZ23a, WLZJ24]. Cardiac [HCC23, LLY^{+}23a]. Cardinality [BF17]. Cardiomyocytes [ZMM^{+}23]. Cardiovascular [Hua21, Wan23f]. Care [Hon24, KJA^{+}23]. Career [LF23b]. Carlo [WXL19, WCZW23]. Carotid [Ano23-31]. Cartesian [Zha23k]. Cartilage [ZRX^{+}24]. Carving [Lia23b]. Cascade [WDZY20]. Cascaded [WCR^{+}21]. Case [DS97, FYH^{+}19, Hav00a, LQZ^{+}21, LZYJ20, McC96, QHJ23, RRM^{+}15, SBK^{+}24, SPH20, AJJF14, GW11, GB94, GL04, MS00, Rou08b, ZS99]. Cases [SZWZ23]. Cassandra [CTCG15]. Categories [LLZW23]. Categorization [OKHS23]. Category [WLZ^{+}21, YLL^{+}23]. Catheter [CH23b, YLLY23]. Catheterization [YLLY23]. Causal [LDV07]. Caused [JFY23, ZH19]. Cava [CZFX23]. CBO [ASS23]. CBO-IE [ASS23]. CCA [ANEA08]. CCFD [LLY^{+}20]. CDIO [Liu23d]. Celecoxib [SYW23]. Cell [SCAP16, WLZJ24, ZXX^{+}23, Ano08b, GGP09, IB09, KD09, RRV09, SKU^{+}09, VvAC^{+}09, VGC09, WSZK09]. Cell/B.E. [VvAC^{+}09]. Cells [DCH^{+}23, FMG^{+}23]. CellSs [BPC^{+}09]. Cellular [YJ23, ZDMY23]. Center [CTCG15, CGC16, JAED23, TZDZ20, ZFFP24, ZL22]. Centered [DDMS15]. CenterFace [XYY^{+}20]. Centers [JFPL16, LZF16, ZXHQ17, ZWL^{+}19, ZHL16]. Central [CH23b, QYL23, YLLY23, Zho22]. Centralized [OKHS23]. Centric [TLC15, ZLF18]. Centroids [DDB^{+}18]. Ceramic [LH23a, TH23]. Ceramide [LJ23]. Cerclage [LTL^{+}23]. Cerebral [BLZG23, BLL^{+}23, DM23, FZH23, LZL^{+}23, LHC^{+}23, SKT^{+}24, WCF^{+}21, ZCL^{+}24a, ZZZ^{+}23, ZCLL23]. Cerebrospinal [SWH23, WZZ23]. Cerebrovascular [DLP^{+}23]. Cervical [KAA^{+}23, LZWL23, LTL^{+}23, SSY^{+}23]. CFD [Jes10, JR10, LYTL20, LLY^{+}20, MYH19, SM17]. Chain [DPMB16, LTF23, LCH^{+}24, LSY20, Liu23g, LWX23b, MTK^{+}23, SX23, TKKZ23, YZL23, ZJZ^{+}16, ZFFP24]. chains [HS03]. Challenge [PCYZ18]. Challenges [AUA^{+}23, CPD21, OKHS23, NJ08]. Challenging [RAdASGR^{+}19]. Chance [SZC16, ZHZS19]. Chance-Constrained [SZC16]. Change [FCSL23, HLW^{+}18, LJSJ23, SCW^{+}23, WHL^{+}18, YHH^{+}23]. Changes [LY23a, LZWL23, SWH23, YLLY23]. Channel [AE19, CLDX20, LTF23, ZDMY23]. Chaos [GAH^{+}20]. Chaos-Based [GAH^{+}20]. Chaotic [ASS23, YXZ23]. Character [DYT23, LGJ18, ZHL^{+}23]. Characteristic [CC23]. Characteristics [BLL^{+}23, Cui23, LZ23a, LLL^{+}23b, LHC^{+}23, SZ20, TY23, VMRM16, WZG^{+}17, Wen23, ZCZ^{+}23]. Characterization [PCGLO14, MJLM07]. Characterizing [OdSSP13]. charge [CBCM93]. Chart [ZLLZ20]. Cheaper [HUN08]. Check [PZH23]. Check-In [PZH23]. Checker [SLP18]. Checking [AMS17, RAA23]. Checklist [WKW^{+}23]. Checkpoint [CSC16]. checkpointing [IMB^{+}13]. Chemical [LGZG23]. Chemistry [KJG^{+}08, SE95]. Chemotherapy [HTZL23]. Chest [LWL^{+}23c, LSS^{+}23, SZW^{+}23, ZZDY24, ZWZ^{+}23b]. Chi [Ano23e]. Children [Li23c, Li23d, WZL23, WJZC23]. Chilean [CSM^{+}18]. China [LJSJ23, FYH^{+}19, HS23c, HXPZ24, JZL^{+}20, JW23b, LCH^{+}24, LLZW23, LQW23, OL23, PFQ^{+}23, SPH20, SWZS17, TH23, WHH^{+}21, WWYR22, YHH^{+}23, YCLK23, ZCL^{+}20]. Chinese [CZY20, DYT23, JHS^{+}20, JZL^{+}20, LWL23b, LGZG23, MSZZ20, PWP21, QFYY20, SZ20, WSM^{+}21, Xu23b, ZHZF20, ZHL^{+}23]. Chip [Ano22a, JK10]. CHIPPER [MM16]. Chips [TLC15]. Choice [ZW23a, AJJF14, FB99]. Cholecystectomy [BGZL23, CZFX23]. Cholesky [CDO^{+}96]. Choreography [AB18]. Chorus [LZ23b]. Chou [SY20]. Chronic [BLZG23, HCC23, JFY23, TY23, XZLZ24, ZCZ^{+}23]. CIM [LD23]. CirBiTree [SW20]. Circuit [AN19, SHW23]. Circular [CSZ^{+}19]. circulation [Hem00, SS00]. Cities [Fen23a, PPB^{+}19]. Citrullination [SW20]. Citrus [CLL^{+}20]. City [Ano23j, CWD23, HCM23, OZI^{+}23]. Clark [NO18]. Class [CZW^{+}23, HTZL23, HZ24, LLC^{+}20, NCMF15, QM23, SDM^{+}19, SO93, WZF24, Ang93]. Class-Imbalance [WZF24]. Class-specific [SO93, Ang93]. Class-Structured [HTZL23]. Classes [SLL^{+}17, Ott93, VHBR93, WHG93]. Classical [PDROFRM16]. Classification [AMH23, AK23, Ano22a, Ano23s, Ano23v, AWB^{+}23, AAS^{+}23, CLHH20, CL23b, CCHJ23, CLW^{+}23, CLYL16, ÇÜ19, DD23b, EMTCA^{+}19, FLR23, FYZ^{+}20, GAH^{+}23, HNSG23, HS23b, KYR^{+}23, LLC^{+}20, eLA20, Liu23e, MqWp23, PRW23, QWZP20, QGS23, SO18, SZL^{+}23, WZ22, WW23b, WWQ23, WLZ^{+}21, Wu23c, XSG23, XHSL17, Xu20, Xu23b, YXL^{+}20, YWGZ20, YT22, YSZ^{+}23, YZWF23, ZHY22, Zha23d, Zha23c, ZH23, ZL23c, ZS23, ZQL23, ZLDX21, KG08]. Classification-Based [CLW^{+}23]. Classifier [AZX^{+}17, LCL^{+}23b, MWQ16, YHZ^{+}23]. Classifiers [LL23a, LJG^{+}23, PDROFRM16]. Classify [SY20]. Classifying [WZG^{+}17]. Classroom [Cha23a, DS23, LLW^{+}22, Liu20, LL24, SF23, WZ23b, XXSL23, YN23, ZLSW23, ZKA23]. Cleaning [LWL23b]. Cleft [DHW^{+}23]. Click [ZFX17]. Click-Through [ZFX17]. Climate [ARAM19, CLPVPI20, SCW^{+}23]. Clinical [CZY20, HWSL22, HSL^{+}23, LLL^{+}23a, Li23d, SSY^{+}23, WCF^{+}21, WL23c, YPY23, ZWZ^{+}23b]. CLO3D [WC23a]. Clone [YXZ23]. Closed [MLL23]. Closure [KZK^{+}23]. Cloth [JSY^{+}21]. Clothing [Jua23c, ZLS^{+}17]. Cloud [ALL^{+}16, AlS16, AUA^{+}20, Ano23z, Ano23-30, Ano23-28, Ano23-46, Ano23-45, AUA^{+}23, BR11, CTCG15, CTC^{+}15, CL23c, CBE23, CMTAC^{+}16, DCF17, FZW^{+}18, HPZ16, Hon23, HYK^{+}20, HS20, JW23a, JL23b, KV23, KPS16, KS19, bLwHcY18, Li23c, LWW23b, Liu23d, LZLJ23, LQW23, Ma22, MKS^{+}24, OCC^{+}16, PBY^{+}20, QYZ17, QHJ23, STM^{+}23, SGK^{+}23, SKA^{+}20a, SKA^{+}20b, SAE^{+}23, SLP18, Son20, SH19, TPSM17, UAK18, WLP^{+}21, WH23, WL23b, WTY23, Wu16, WDY23, XST^{+}16, Yan23e, YPM^{+}21, ZWL^{+}19, ZGFL23, ZJZ23, ZHL16, ZHW^{+}21, ZTLH18, JMR^{+}11, SBJV11, TSCT11]. Cloud-Based [AUA^{+}23]. Cloud-Enabled [ZTLH18]. Cloudlet [CBE23]. Clouds [CLYL16, HLPY16, MFB^{+}15, ZWZ16, OPP11]. Cluster [ACCTA^{+}15, BA18, CW23, FA24, FYY23, FZH23, HS20, LLLZ23, LX23d, SHHI01, SCD^{+}15, ZCLL23, HMCH07, JEM07, GWZ16]. Cluster-Based [FA24]. Cluster-enabled [SHHI01]. Clustering [Ano23-31, CZD^{+}20, CLG^{+}23, CZP^{+}24, DYYW24, DDB^{+}18, FZH23, GDA^{+}23, GMM20, GZD23, HKVVNT18, Hon23, HZD17, HW23b, LXWD18, LZSC23, LJZ18, LLZW23, Liu23b, NTNTTK^{+}19, Ona19, OL23, SGK^{+}23, SD23, SAZ^{+}24, SWZL24, SLL^{+}17, VVNHTHNT20, WYC^{+}20a, WM23b, WWW^{+}17, XH20, tXChL^{+}24, YXZ23, ZCLL23, ZLSW23, ZW23c, DRL13, GGJ04]. Clustering-Based [Ona19, HW23b]. Clustering-Evolutionary [FZH23, ZCLL23]. Clusters [LYTL20, LWW16, BCHL05, BDV03, JR10, KK11, Nag05a]. CM [GBJ94, MJ95]. CM-2 [GBJ94]. CM-200 [GBJ94, MJ95]. CNN [LLW^{+}22, AXA20, BKMS23, CFL^{+}23, QFYY20, WTY23, XLY18, YZQ^{+}23, ZLWW23]. CNN-BiGRU [QFYY20]. CNN-BiGRU-AT [YZQ^{+}23]. CNN-BiLSTM-ECA [CFL^{+}23]. CNN-LSTM [BKMS23]. Co [LWCZ21, LWX23b, RN07, ZGW08, WCZ^{+}16, BCS01]. Co-Array [BCS01]. Co-arrays [RN07]. Co-Creation [LWX23b]. Co-Occurrence [LWCZ21]. co-regulations [ZGW08]. Coaching [Ano07a]. Coal [LZ23c, WZY^{+}23, XCW18]. Coarray [SM17]. Coarrays [RRM^{+}15]. coarse [KB96]. coarse-grain [KB96]. Coastal [MVIA20]. Coastline [YHH^{+}23]. CODAS [Xu23d]. Code [ACBV20, BCC^{+}92, BKLS01, CTS23, CCBPGA15, EOS17, FHT23, FXYY17, GNKuR23, HNG^{+}23, IPST18, RWN20, SSSJ20, tXChL^{+}24, ZLLS20, ZLLZ20, yZcR23, ACIK97, BHK^{+}13, CP03, Jes10, PKE^{+}10]. Codes [O’K00, BGLR93, IJL^{+}01, Kok07, Mal01, NAA^{+}03]. Codesign [MOS16]. Coding [ATZ^{+}15, BKMS23, LX23c, YPC^{+}23, ZWS19, FC01]. Coefficient [PLAAH23, XWL19]. Coevolution [LZ23b]. Cognition [Ano23-48]. Cognitive [GKK^{+}20, RKA^{+}24, TYCL23, Wan21]. coherence [WC96]. Cold [LCL^{+}23a, Liu23g, ZFFP24]. Collaborating [CL04]. Collaboration [HS20, HXPZ24]. Collaboration-Aware [HS20]. Collaborative [CL23a, LZZ23a, LLP23, LLGL20, LLY23c, MBBDP17, TKZ^{+}24, WLP^{+}21, Wan23d, fXhCL17, YWH20, Zha23j, FML^{+}14, KJG^{+}08, XNQF04]. Collapse [AXWS23]. Collection [GLW^{+}23, KV23, Wei23, ZGT^{+}23, Hav00b]. Collections [BBC^{+}10]. Collective [FML^{+}14]. College [Ano23o, Ano23a, Ano23-27, Ano23-33, Ano23-48, Cha23a, Lou23, QQWJWYQZ23, SF23, Wan21, WH23, Wan23h, WL23a, WDY23, Yan23e, WLJ22, Zha22]. Colleges [Ano23-46, Han23, Hon23, Liu22, LGZ23, WZ23b, yZcR23]. Collision [JSY^{+}21]. Colon [ZCL^{+}24b]. Colonic [GGSL23]. Colonoscopy [GGSL23]. Colony [FTS^{+}20, HLHC23, LBW^{+}24, Liu23g, QHAN23, SCZ18, WCLL17]. Color [HCC23, He24, LLY^{+}23a, fTPsLZ23, YLLY23, ZSD22, ZLHZ23]. Colorectal [PTZL23, YLLY23]. coloring [BÇCD12]. Columns [NLMT23]. Combination [Ano23-41, MLZ23, YWX23]. Combinations [AZGJ24]. combinatorial [BÇCD12]. Combined [DRR12, FHC^{+}22, GSL23, GLD^{+}24, GGSL23, HWSL22, HY15, HS23c, JLJ24, LX23b, LLY^{+}23a, SYW23, XLB^{+}23, XYL^{+}23, Yan23a, ZLL^{+}15, Zha23a, ZZZ^{+}23, ZZW^{+}23, ZLZ^{+}21]. Combining [DZL23, Hu23b, LGC^{+}16, MSH99, WW22, MS00]. combustion [SE95]. Command [FWX23]. Commerce [Ano23-30, Ano23-45, FYH23, Li24, Liu23f, YZL23, YT23, ZCCS23]. Commercial [MSZZ20, SWZS17, ZXX^{+}23]. Common [BD16]. Communicating [SWH23, WZZ23]. Communication [Ano23-46, AAF19, CM21, GBK^{+}96, LLP^{+}19, MAS^{+}23, RLU^{+}20, SCKW23, Wan23g, YJ23, Yan23e, Zha23i, Zhu22, BDM^{+}04, Bry96, ESSL99, KB01, MJ95, PKE^{+}10, TPKP13]. Communications [GBJ94, SGK^{+}23, GG95]. Communities [WXL19]. Community [Ano23-42, CT23, KCS15, SSN15, SL21, XWL19, ZW18]. Community-Based [SSN15]. Compaction [BA18]. Companies [ARGC19]. Comparative [GPX^{+}21, LZL^{+}23, Sah20, SO18, SMB17, YPM^{+}21, ZS99, OB96]. Comparing [FC01, PSM^{+}15, SCR20]. Comparison [ÇÜ19, HZHW23, PLAAH23, PBK01, WLLY23, YWGL20, ZWW^{+}23b, ZG15]. Competencies [ZGL23]. Competition [Su23]. Compilation [Ano99b, CP03, ESSL99]. Compile [CGMRO^{+}19, MSH99]. compile-time [MSH99]. Compiler [BMN^{+}97, DS97, DZKS13, FHT23, IRSD99, OSS94, SKS01, Sin99, WC96, Ang93, BGH99, EAS^{+}97, HCJ08, Mül03, OB96, SHHI01, SO93, Ben99]. Compiler-directed [DZKS13]. Compiler-enforced [WC96]. Compilers [CCTY15, Met99a, Met99b]. compiling [DDS99, GG95]. Complementary [SLC^{+}23, Ton20]. Complete [KKD^{+}20, Nag05b]. Completion [LTZS18, NO18, RWN20]. Complex [DMV^{+}19, LCH^{+}24, ML23b, Zho20, Che93a, DSS^{+}05, ESSL99].

Complexity [ABK^{+}23, BKMS23, KGIL19, Rou08a, ZWS19, ANEA08, PPS12a, PPS^{+}12b]. Component [TTL^{+}23, WmG21, ZCT^{+}20, KCO^{+}05, KJG^{+}08, Sch03b]. component-based [Sch03b]. Components [ADLN08, TLZ^{+}23, TMY^{+}23, ZCXQ16, CL04]. Composition [CCM^{+}23, HL23c, Zha23m, ZSZ23c]. Comprehension [SLL^{+}17, SLDL17]. Comprehensive [Cha23b, RST02]. Compressed [HSY^{+}23, LWZ24, QGS23]. compressible [WB95]. Compression [Ano23-39, HZL^{+}23, LZK^{+}23, YCWL17, IMB^{+}13, LLW^{+}14]. Computation [KGIL19, LD23, Nag05b, SRO^{+}15, DSZ96, HST^{+}93, IB09, KB01, Lar93, OSS94, PMTL12, PTS^{+}13, PKE^{+}10, WGF93]. Computational [DMV^{+}19, KAS15, LL23b, PLY^{+}23, RLPI19, SRTCL16, GW11, GHW00, HLM^{+}05, MCvM10, WD07, SLJ^{+}00]. Computations [GA18, CLM05, CFR14, HUN08, HJYC10, PCS99, ZSS^{+}10]. Compute [Ser18, JIC^{+}14]. compute-local [JIC^{+}14]. Computed [Ano23n, BLL^{+}23, CWMZ23, CLF^{+}23, CPG^{+}23, GGSL23, LHC^{+}23, LJ23, NHHW23, SYW23, SWH23, TY23, WCF^{+}21, WWH^{+}23, WLZJ24, WJZC23, XZLZ24, ZZZ^{+}23, ZSZ^{+}23b, ZZ23, ZHHX23]. Computer [Ano23j, Cao23a, JZH^{+}23, MUJN20, PZ24, RVJ^{+}23, Yan23c, ZG23, ZZSQ23, BTS^{+}14, IJL^{+}01, JEM07, KMB09, LDV07, Ste97, WCKD07, WSP94]. Computerized [DM23, LWL^{+}23c, XZ23c, ZTL^{+}23]. Computers [PAR94, TH23, HB94, LH93, OS99, PBK01, PMCF94, Shu94]. Computing [ALL^{+}16, Ano23z, Ano23-30, Ano23-28, Ano23-46, Ano23-45, BR11, CCF^{+}17, CTC^{+}15, CBE23, CEMM17, CPK19, CPD21, DMAC23, DK02, DDMS15, FZW^{+}18, GM14, GGP09, GN22, HPZ16, Hon23, LG18, LXWD18, bLwHcY18, Li23c, Liu23d, LGZ23, LQW23, Ma22, Nag04, OPB^{+}20, PSM^{+}15, PT09a, PT09b, QHJ23, SGK^{+}23, SKA^{+}20a, SKA^{+}20b, SLP18, Son20, UAK18, VR13, WWLG16, WH23, WY23, WTY23, WW23a, WDY23, XHL^{+}23, YTD15, Yan23e, ZGFL23, ZJZ23, ZHW^{+}21, ACB^{+}02, Ano08b, AB96, BKK^{+}11, BH12, BDG^{+}94, BBK^{+}11, BC99, BÇCD12, BDH^{+}10, CJS^{+}02, CM02, Che93b, DDMS14, DRR12, FB99, FP00b, GW11, GGMS99, GHH^{+}02, Hav00a, JIC^{+}14, KeKR^{+}11, KLS^{+}96, LPHD04, LG03, MMG^{+}02, NJ08, Nor07, NDSG07, Rou08a, SKU^{+}09, VCT05, VF95, WGW08, Zim07, HHML05, JEM07, SO11, Mic97]. Concentration [SAN^{+}19]. Concept [WWYR22, ZH19, BMY13]. concept-based [BMY13]. Concepts [DNS97, PMEH17, NDSG07]. Concrete [Men23]. Concurrence [ZLDX21]. Concurrency [YZX17]. Concurrent [BBC^{+}10]. Condition [ZMM^{+}23]. Conditional [BCK07, KPHY18]. Conditions [LSY20, Zho20]. Condor [FYY23]. Conference [GM14]. Confidence [ÁRCPS15]. Config [MS00]. configuration [CQF05]. Confounder [HMQS23]. Confrontation [WLQ23]. Congenital [JP23, YZWF23]. Congestion [HPZ17, PSW23]. Conic [SO18]. conjugate [GG05]. CONLAB [JKR92]. Connected [AN19]. Connection [WC23b, GBJ94, MJ95]. connectivity [PSU08]. conquer [MGK^{+}13]. Conquering [HLZ^{+}21]. Consensus [Ma23a, Ona19, Yi23]. considerations [GHST12]. considering [CLL^{+}18, FH23, GLW^{+}23, LLC^{+}20, SQ21, ZWG^{+}20, ZYZL23, ZLL20]. Consistency [Bak23, Ma23a, ZLL^{+}15]. Consolidation [CLYL16]. Constant [QK21, CGL08]. constant-time [CGL08]. Constrained [LTF23, LZLZ19, LLS^{+}20, MFB^{+}15, SZC16, YZ16, ZHZS19, Zha20b, LBvBW12].

Constraint [ML24, SW16, WM23b, YPC^{+}23]. Constraints [BD16, LR05]. Construct [FLC^{+}20, SLDL17]. Constructing [BBBM12, GLH23]. Construction [Ano23o, Ano23p, Ano23s, Ano23q, Ano23t, Ano23r, Ano23-32, CT23, Du23a, HCM23, HCGGSS17, LWW23a, LQW23, LL24, MLZ23, PLJ^{+}23, Son20, tS23b, WL23b, Wu23b, XY23a, XZ23a, Ye23, YB24, ZKA23, MOT97]. constructs [CP95, RS94]. Consumer [CWD23]. Consumption [ACMT19, FX23, WG21, WG24]. Container [JW23a, LGJ18, ZZD23a]. Container-Based [JW23a]. Content [DRQ^{+}18, FHS^{+}23, LZZ23a]. Content-Based [DRQ^{+}18]. Context [Ano23l, Ano23w, Ano23b, HL23d, RKA^{+}24, SC23, SH19, Zha23l]. Context-Aware [RKA^{+}24]. Context-Based [Ano23l]. Contextual [LMZP23, LWCZ21]. Contingent [ZJZ23]. Continuous [HI18, HLZ^{+}21, LFS^{+}20, SFZ^{+}20, SNH24, WKW^{+}23, XNLJ23]. continuum [ADLN08]. contour [RLC04]. Contract [Ge23, HAA^{+}22, KL23, LSY20, TKKZ23, YLW21, YP24, ZL23a]. Contraction [ZSZ^{+}23a, PMCF94]. Contrast [Ano23-31, ZXZZ23, ZCL^{+}24b]. Contrast-Enhanced [Ano23-31, ZXZZ23, ZCL^{+}24b]. Control [GBH20, HPZ17, HT22, HXX16, HHCC23, Liu23f, Luo23, PSW23, PDRC18, RWY^{+}17, SZWZ23, SCD21, WNL^{+}20, lWwFAH23, WJS^{+}23, WZF^{+}20, YX23, Yan22b, ZLSW23, FP95, GR93]. Controlled [CGA17]. Conventional [Alk19]. convergence [HMCH07]. Conversion [WKY23]. Converter [VMRM16]. Convex [XZNL23]. Convolution [TZL^{+}24]. Convolutional [Ano23-47, CD23, CYLD23, CT23, CPG^{+}23, HTZL23, JFY23, LLW^{+}22, LXL^{+}23, Liu22, LLL23c, MJK23, QWZP20, TTL^{+}23, WCR^{+}21, WW23b, WWW23, XL23, Yan23b, ZYZ23, ZS23]. Convolutional-Neural-Network-Algorithm-Based [Yan23b]. Cooperation [Ano23f, LQLX18, XY21, XY22, YZL23, Zho20]. Cooperative [KF20, LHZ^{+}23, CR05]. Coordinate [Bjø00, GHW00]. Coordinated [PRM^{+}14]. Coordination [CHM^{+}97, HCM23, LSY20, MTK^{+}23, YZL23]. coping [Hil97]. Copper [LFS^{+}17]. Coprocessor [HYS^{+}19, SRO^{+}15]. Coprocessors [MYH19, TSP^{+}15]. CORBA [Lan01]. Cord [ZZH22]. Core [CKKK15, DGL^{+}17, TC96, TLC15, JR10, JIC^{+}14, PTS^{+}13, PT09a, PT09b, Zhe10, DGH^{+}15, Nag05b]. coregulation [HB12]. Coronary [HL23a, LZ23a, SW24, Wan23f, WWH^{+}23, XNL^{+}22]. Corporate [tS23b]. Corpus [HNMB19]. Corral [JDVM10]. Correcting [tXChL^{+}24]. Correction [CSZ^{+}18, DCF17, fTPsLZ23, WJZC23]. Correlation [CZCL23b, GDA^{+}23, KLK^{+}20b, KZK^{+}23, LYL^{+}24, SWY23, TZH20, WWH^{+}23, WYHS23, WZC^{+}23, KMB09]. Correlation-Based [KLK^{+}20b]. Correntropy [WPDZ19]. Corrigendum [KGU^{+}20, SKA^{+}20a, WG21]. Corruption [DWZ^{+}20]. COSA [CCF^{+}17]. cosmological [War14]. cosmology [BAN02]. Cost [Ano23-30, BA18, CWH23, Du23a, HLY20, HWJ18, MFB^{+}15, MKH^{+}19, QYZ17, SH19, Wu16, Ye23, BKP99, JMR^{+}11, PMCF94, SZ04]. Cost-Benefit [HLY20]. Cost-Effective [Wu16]. Cotton [SKB^{+}24]. Counseling [QQWJWYQZ23]. Counting [LZWJ23, TZL^{+}24]. Countries [HLS^{+}24]. Coupled [YW24]. Coupling [WBZL17, GRN99]. Course [Ano23e, Ano23d, Ano23-36, Cao23b, JL23a, Wan21, Zha23j, ZW23a, Zhu23b, ZGW08]. Courses [JL23b, Liu23d, SF23, tXChL^{+}24, Met99a, Met99b]. Cover [FCSL23, SCW^{+}23]. COVID [AKA^{+}23, DD23a, Guo23b, JAED23, NHHW23, TMHM23, WW23a, ZZDY24, ZWZ^{+}23b]. COVID-19 [AKA^{+}23, DD23a, Guo23b, JAED23, NHHW23, TMHM23, WW23a, ZZDY24, ZWZ^{+}23b]. CP [WWW^{+}17]. CPFRS [JZY23]. CPGAN [ML23a]. CPI [CC23]. CPU [CFR14, LPBSSE17]. CPU/GPU [CFR14]. CPUs [LNK^{+}15, YTD15]. Crack [Men23]. CRAFT [PMM94]. Crane [KLuH^{+}20]. Crane-Robot [KLuH^{+}20]. Craniectomy [SWH23, WZZ23]. Craniocerebral [SWH23, WZZ23]. Craniotomy [BLZG23]. cratering [HLM^{+}05]. CRAUL [IRSD99]. Creating [BKLS01]. Creation [LWX23b, ZG23]. Creative [Ano23-35, Lia23b]. Credit [Pen23, Yan22b]. CRIM [KPHY18]. Crisis [Wan23a]. Criteria [HKQ^{+}23]. Criterial [QHAN23]. Criterion [WPDZ19]. Critical [GXY^{+}20, NM16, NLMT23, ZA10]. Criticality [GPP^{+}17]. Cross [AMS17, Ano23-30, FYZ^{+}20, HNMB19, Liu23f, MCF^{+}23, NLMT23, SCD21, XZNL23, YZL23, ZLLS20, ZCL^{+}20, Zha23h, BHK^{+}13, GA96]. cross-architecture [BHK^{+}13]. Cross-Border [Ano23-30, Liu23f, YZL23]. Cross-Checking [AMS17]. Cross-Domain [FYZ^{+}20, SCD21]. Cross-Efficiency [ZCL^{+}20]. Cross-Individual [MCF^{+}23]. Cross-Language [ZLLS20]. Cross-Lingual [HNMB19]. Cross-Modal [Zha23h]. Cross-Section [NLMT23]. Crossover [YWGL20]. Crowd [BYZ^{+}23, LZWJ23, SMA^{+}23, TZL^{+}24, ZAW^{+}21]. Crowdsensing [LQLX18]. Crowdsourcing [HKQ^{+}23, SMA^{+}23]. Crude [Ano22b, Li24]. crunching [GB94]. Crypto [AHM23]. Crypto-Watermarking [AHM23]. Cryptographic [GMS^{+}23]. Cryptography [TXD^{+}20]. Cryptosystem [GAH^{+}20]. crystals [GIKP95]. CSE [CHC15]. CSMRI [ZCL^{+}24a]. CT [Ano23i, Ano23u, BGZL23, CZD^{+}20, CL23b, CWLW23, FLR23, HZL^{+}23, Hua21, HZHW23, JP23, LZL^{+}23, LSS^{+}23, LCL^{+}23a, LYZ^{+}23, PWY23, SWY23, WZL23, WZ23a, WSL23, WMZM23, Xu23f, ZZW^{+}23, ZZDY24, ZWLL23, ZWZ^{+}23b, ZLZ^{+}21]. CU [ZCS23, ZWZ20, ZWZ23a]. Cuckoo [WS15]. Cucumber [QZY^{+}20]. CUDA [LYTL20]. Cultivation [WLJ22]. Cultural [Ano23-35, Jua23b, LG18, Lia23b, Lvp23, WWYR22, Xu23a]. Culture [LZ23b, Liu22]. Curative [LZWL23]. Currency [YDC^{+}24]. Current [Ano23z, LH23a]. Cursive [HNSG23, MSA^{+}20]. custom [NAA^{+}03]. Customer [ZLF18]. Customer-Centric [ZLF18]. Customers [SQ21]. customizable [WOC99]. Customize [HPGC19]. Customized [XST^{+}16, ZPA19]. Cutting [BRSP18]. CVM [ZJZ23]. CX [CM02]. Cyber [JGDK23]. Cyber-Twin [JGDK23]. Cycle [ZCXQ16]. Cylindricity [YLWW18].

D [Nag05a, Qin23, JR10, LZZ^{+}23b, MRLF12, PPJ^{+}15, TPKP13, fTPsLZ23]. D-P [Qin23]. D2D [YJ23]. Daily [SDM^{+}19]. Damage [Che23c, ZC23]. Dance [PZ24, QS23]. DANet [YXW^{+}22]. Danjiang [HLHC23]. Data [AMS17, ASS23, AlS16, AMB23, AHMMÁR18, ÁRAHMM18, AK23, Ano19, Ano23e, AZX^{+}17, Bai23, BH23, Bak23, CLLZ23, CC23, CIU20, CTCG15, CTC^{+}15, CLL17, Cha23a, CSC16, CGC16, CHZY17, qCtWfPbZ23, CZ23, CWS^{+}24, CNW^{+}23, ÇÜ19, CFY23, DQQ17, DCF17, DS23, DD23b, DW23, FSS^{+}17, FYH^{+}19, FLC^{+}20, FWX23, FAA^{+}20, FC22, Fen23b, FCSL23, FYH23, GBH20, GKK^{+}20, HJ15, HMC97, HKH^{+}17, HBSZ20, HNSG23, HYK^{+}20, HL23c, HCQ17, HYS^{+}19, HL23d, Hui23, JGF23, JFPL16, JZL^{+}20, JZ22, JWPW24, KL23, KGU^{+}20, KIU^{+}21, KGU^{+}21, KPS16, KPW^{+}18, KKK^{+}18, LR05, LKS18, LZF16, LJF20, Li23c, LD23, LWL23b, LLY24, Lia23a, LFQA20, Lin23a, Lin23b, LJZ18, LFY^{+}18, LLC^{+}20, LWQ^{+}23, LLZW23, LJSJ23, LZLJ23, LZK^{+}23, LLC^{+}23, LYZY24, LL24, Lv23, MZLC21, Ma22, MLHC20, MABU18, MJB15, NGMAA23, NLPL^{+}20, OKHS23, PZH23, PY17, PZL19, PFQ^{+}23, PPB^{+}19, PCYZ18, RMS21, RBJ^{+}19, RLPI19]. Data [SAJ^{+}20, SLP18, SZXL16, Son23a, Son23b, SGPCLG19, SSM^{+}02, SL23b, THS^{+}15, TZDZ20, Tia23, UAK18, WZG^{+}17, WZW^{+}20, Wan21, WCR^{+}21, WHH^{+}21, WJCZ21, Wan22, WZY^{+}23, lWwFAH23, WG23, WWL^{+}23, Wei23, WCG95, Wu16, WLZ^{+}21, WG21, WG24, XCW18, XHC^{+}23, XHX23, XZ23c, Xu23c, YLL^{+}23, YCWL17, YL23, YYY23, YB23, YHH^{+}23, YC23, YT23, ZCSL17, ZG18, ZWL^{+}19, ZLCW20, Zha22, ZWW^{+}23a, ZZD23a, ZGL18, ZL23c, ZQL23, ZX17, ZHL16, ZLDX21, ZW23b, ZZX^{+}17, ZTLH18, Zhu22, ZLX^{+}23, ZZSQ23, ZWZ^{+}23b, ZGT^{+}23, hKJ^{+}20, AMN^{+}12, AHCZ12, AGG^{+}97, CDD^{+}05, CvHK97, CP95, DRT^{+}14, Gil09, Hav00b, HJ96, HS03, IB09, IMB^{+}13, KG08, KB96, KGV97, KVW^{+}07, LOHA01, LLW^{+}14, Lin04, NPP^{+}00, Özt04, PSU08, PDGQ05, Pla04, PMCF94, RR07, RKR^{+}99, Rou08b, SVR^{+}07, Sin99, Slo12, SG96, TFN11, VRW^{+}03, VRM02, WHRH07, War96, ZRP08, ZS99, ZGW08, ZSS^{+}10, vDKH01]. Data-Acquisition [OKHS23]. data-aware [IMB^{+}13]. Data-Driven [CWS^{+}24, JWPW24, Lv23, MZLC21, PPB^{+}19, RLPI19, Wan21, Zha22]. Data-Intensive [AHMMÁR18, ÁRAHMM18, DRT^{+}14, Slo12]. Data-parallel [HMC97, WCG95, CvHK97, Hav00b, HJ96, ZS99]. data-structure-neutral [SG96]. Database [AIA^{+}21, CLN^{+}24, CTC^{+}15, GBH20, HMSW92, LX23c, BDLL94, MSLS13, WSZK09]. Databases [JNK24, TH18, WWZ17, TSCT11]. Dataset [ASS23, ABNJMB23, SAZ^{+}24]. Datasets [BC23, CGMRO^{+}19, SZAG15, BCHL05, HB12]. DAuGAN [BC23]. David [Ano07c, Nag05b]. DDT [AGG^{+}97]. de-allocation [RMX05]. DE-LS [SQ21]. DEA [Ano22b, YCLK23, ZCL^{+}20, Zha20b]. DEA-Malmquist [Ano22b, YCLK23]. Deadline [MFB^{+}15, WLP^{+}21]. Deadline-Constrained [MFB^{+}15]. deadlock [HPS^{+}13]. Debridement [CLHH20]. Debt [SCR20]. Debts [OL23]. Debugging [MR18, FP00b]. Decay [QWZP20]. Decision [AUA^{+}20, BH23, Che22, DZL^{+}20, FWX23, Fen23b, Gog23, HNG^{+}23, KLuH^{+}20, LRX^{+}17, LJF20, LD23, LWX23a, LJG^{+}23, Lv23, LWCJ23, MZLC21, MWQ16, PFQ^{+}23, QHAN23, SZL^{+}23, tS23b, Son23b, SNH24, TMY^{+}23, UJS^{+}20, XNL^{+}22, XNLJ23, XHL^{+}23, XSG23, YZL23, YL20, ZCS23, ZWZ20, ZWZ23a, ZLL20]. Decision-Maker [ZLL20]. Decision-Making [AUA^{+}20, Che22, Fen23b, Gog23, LRX^{+}17, Lv23, QHAN23]. Decision-Support [MZLC21]. Decisions [CL23a, ZGS16]. Deck [YZSZ17]. Decoction [Xu23e]. Decomposition [LDW23, ML23a, ML23b, WZML23, WWW^{+}17, DBVF01, GRN99, JHNP14]. Decompressive [SWH23, WZZ23]. Decoupling [ZLDX21]. Decremental [WM23a]. Decryption [GAH^{+}20, HGMM23]. Deduplication [CSC16, HBSZ20]. Deep [AXA20, ALA20, Ano19, Ano23l, Ano23p, Ano23r, Ano23u, Ano23b, Ano23-35, Ano23-40, AZX^{+}17, Bia23, Cha23a, CLG^{+}23, FX23, FYZ^{+}20, FS24, FZ23, GLD^{+}24, GGSL23, GRB23, HLL23, HL23a, HCC23, Hu23a, HTZL23, HH23a, HZHW23, HLH23, HW23a, HZ23, Hui22, JGF23, JSY^{+}21, LLQ^{+}23, Lia23a, LCL^{+}23a, eLA20, LWZC21, LWW23b, LDW23, LLT23, Ma23c, MUJN20, Men23, MLY23, OPB^{+}20, PTZL23, PVCPdPSZVG17, PRW23, QYL23, QWZP20, SKB^{+}24, SLH23, SXP23, SL23a, SWW^{+}24, SW24, tS23b, Tan23, TMG^{+}19, TYCL23, Wan22, Wan23f, WZ23a, WWH^{+}23, WYHS23, WWQ23, WZML23, WM23a, Wan23c, WL23a, Wan23e, WDD^{+}23, WLZ^{+}21, Wu23a, XLB^{+}23, XHX23, Xu20, Xu23b, YY23a, YXL^{+}20, YYY23, YGZ^{+}24, Yan23h, YT22, YT23, YZ20, ZHY22, Zha23d, ZYZ23, Zha23f, ZGFL23, Zha23c, ZLWW23, ZWZ20, ZWZ23a, ZD23a, ZYL23, ZWF23, ZX17, Zho22, ZWLL23, ZWW^{+}23b]. Deep-Learning-Based [Zha23f]. Default [ÇÜ19, GPX^{+}21]. Defect [AFB23, CWSF23, CWS^{+}24, CTZC24, FDY^{+}19, LCW22, LZSC23, LWW23b, LML^{+}23, XLY18, YYF^{+}19, ZGY^{+}20]. Defects [SZL^{+}23]. Defensive [Che22]. Defined [HPZ17, KV23, PLW^{+}20]. Definition [LYL^{+}23, HS03]. definition-use [HS03]. Deformation [TZZ23, ZWW^{+}23a]. Deforming [MLYH17]. Degenerative [ZYY^{+}23]. Degradation [LW23]. Degree [CZY20, CFY23, LWCZ21, WZC^{+}23, Wen23, ZCXQ16, ZWW16]. Dehazing [WFZ^{+}23]. Delaunay [LZ23c]. Delay [AMB23]. Delays [HXX16]. Delegated [ITF^{+}08]. Delivery [JHL^{+}23]. Delta [HWY^{+}23, Xu23a]. Demand [PPB^{+}19, ZJZ^{+}16, ZXHQ17, Ang93]. Demands [SLBZ16]. Demarcating [XZNL23]. Dementia [ZLT^{+}21]. Demons [LWZ^{+}23c]. Denoising [AZX^{+}17, FHC^{+}22, HSS23, LZL^{+}23, LLW23a, WHL^{+}21, WZML23, WG23, WZ23c, WZ23d, YPC^{+}23, ZYY^{+}23, ZZZ^{+}23]. Dense [KYR^{+}23, YTD15]. DenseNet [Ano23v, Yin23]. DenseNet-Based [Ano23v, Yin23]. Density [AFB23, CZW^{+}23, HKVVNT18, LPD20, LLZW23, TZL^{+}24, YXZ23, ZGY^{+}20]. Density-Based [LPD20]. dependence [BH02, Lin04]. Dependency [TLW^{+}23, ZHZF20]. Dependent [yXHtJC23, ZCLL23]. Deployment [LJZ18]. Depots [LJZ18]. Depression [CZY20]. Depth [DZL23, LBY^{+}23].

Derivative [BCC^{+}92, Son23a]. derivatives [Met99a, Met99b]. derived [RMX05]. Dermatitis [Xu23e]. Descent [NTNTTK^{+}19, WWW^{+}17]. Description [LXW24]. descriptions [MPP^{+}04]. Descriptors [HY23, XXM^{+}23]. Desert [WG22]. Desertification [Cui23, MVIAM20]. Design [Ano23h, Ano23j, Ano23g, Ano23w, Ano23x, Ano23y, Ano23-41, Ano23-35, Ano23-42, ABR17, BTV^{+}20, Cao23a, Cao23b, CFR14, CHL^{+}23, CDO^{+}96, CIN^{+}96, CFY23, Dai23, DDJ20, FSS^{+}17, FGSCR^{+}17, FCSL23, GHST12, GSM03, GPP^{+}17, HZJ23, HNMB19, HMR^{+}15, HT22, HS23a, HL23d, KL23, KS19, Li22, LY23b, LL23d, Liu23b, Liu23a, LLY23c, Lv23, O’K00, PPD05, RX04, Sah20, SS00, SC23, SY23, SLP18, SG96, SV15, SF23, Sun23b, SL23c, Tia23, Tun23, WKW^{+}23, Wu23b, Xia23b, XL23, XSS^{+}23, Xu23a, XYCC23, YLW21, ZWW16, ZJ23, Zha23m, Zhu22, DRT^{+}14, DSZ96, FP95, PDA^{+}08]. Designed [Ano23a]. Designing [Ano23-44, HPZ16, ML23a, Lin04]. designs [OHS00]. Detecting [Ano23v, CLL^{+}20]. Detection [AMY^{+}21, AXWS23, Ano23m, Ano23n, Ano23-31, BD16, BKU^{+}23, CLA^{+}15, CZFX23, Che23c, CWSF23, CWS^{+}24, DC22, DYYW24, DBWY24, FYC^{+}20, FXYY17, FZ23, GAH^{+}23, GGSL23, HJ15, HAW23, HJO^{+}20, He22, Hu23b, JGF23, JMGH21, JSY^{+}21, KKD^{+}20, KIU^{+}21, KJA^{+}23, KCS15, Li23f, Li23a, LZSC23, LCL^{+}23a, LWZC21, LDZZ23, LLW23b, LZK^{+}23, LXW24, LML^{+}23, Ma23c, Men23, MLL23, MLY23, MWZ^{+}23, QWZP20, QFYY20, QWP23, RAZ^{+}23, RAdASGR^{+}19, She23, SHCZ23, SZL^{+}23, SSY^{+}23, TZDZ20, TMHM23, UJS^{+}20, WCR^{+}21, WYY21, WTY23, WJS^{+}23, WSL23, XML^{+}16, XYY^{+}20, XZ23b, YXL^{+}20, YXW^{+}22, Yan23c, YDC^{+}24, YHH^{+}23, Yua20, YZ20, ZLLS20, ZLLZ20, ZQ22, ZLHZ23, ZSZ^{+}23a, ZW18, ZZH22, ZSZ23c, HPS^{+}13]. Deterioration [HSY^{+}23]. Determine [NM16]. Determining [ADS95, ÇÜ19, Wen23, BH02]. Detonation [NYML17]. Develop [CNW^{+}23]. Developing [CYJ^{+}20, HLS^{+}24, PCYZ18, ZYX^{+}21, ZHW^{+}21]. Development [AKA16, Ano07a, Ano23g, Ano24, BBI23, CLL17, CW23, CZS^{+}23, FSLG17, FS01, GDNVSMM19, GZD23, HNMB19, Hon23, HLM^{+}05, JZL^{+}20, JZS21, Lia23b, LLY^{+}20, LWX23a, PT09a, PT09b, QHJ23, REPRL^{+}19, SB01, SW23, Sun23c, VHBR93, WZ22, Wan23g, ZAW^{+}21, ZGL23, dBMT20, ARvW03, Ano93, JKR92, KJG^{+}08, SDS00]. Deviation [WZL23]. Device [FSLG17, FYC^{+}20, LGZG23]. Devices [CGMRO^{+}19, RLU^{+}20, SAJ^{+}20, Wen23]. Dexmedetomidine [GSL23, ZY23]. Dezocine [CZFX23]. DGR [Alk19]. Diabetes [LWZC21, ZLHZ23]. Diabetic [Ano23-34]. Diagnosing [FZH23]. Diagnosis [Ano23i, Ano23n, CWMZ23, CLF^{+}23, CL23b, CZY^{+}23, CTD23, CPG^{+}23, CWLW23, DHW^{+}23, DM23, FTS^{+}20, FLR23, FHC^{+}22, HWSL22, HSS23, He24, KAA^{+}23, LZL^{+}23, LLQ^{+}23, LX23b, LLLZ23, LLY^{+}23a, LJ23, Liu23c, LTR^{+}23, LYZ^{+}23, LML^{+}23, NHHW23, OWL^{+}23, QYL23, SZL^{+}23, SW24, SHW23, SWY23, TY23, WHL^{+}21, WCF^{+}21, WLQ23, WMZM23, XLB^{+}23, XYL^{+}23, Yan23b, YPY23, YYW^{+}23, Yin23, YHZG23, ZYZ23, ZXZZ23, ZZZ^{+}23, ZY23, ZYL23, ZSZ^{+}23b, Zho22, ZTL^{+}23, ZLT^{+}21, ZHHX23]. Diagnostic [LSS^{+}23, WZ23d, ZLZ^{+}21]. Diagram [YYL17]. Diagrams [AN19, NCMF15, NL19]. Diarrhea [WHH^{+}21]. DiDi [LZYJ20]. Difference [Ano23v, LBY^{+}23, Vol97, VHBR93]. Differences [WWYR22]. Different [GMCT15, LBW^{+}24, NL23, SAJ^{+}20, WLLY23]. Differential [NTNTTK^{+}19, QTD^{+}20, TZCZ23, ZCT^{+}20, Gus93, PPS^{+}12b]. Differentially [DQQ17]. Differentiated [GLW^{+}23]. Differentiation [DCH^{+}23, Yin23]. Diffusion [HTZL23, KSYC23, LX23d, LWZ^{+}23c, LYL^{+}24]. Diffusion-Weighted [HTZL23, LYL^{+}24]. digit [BSF96, HLZ^{+}21]. digit-index [BSF96]. Digital [AN19, Ano23b, Cao23a, Che23b, Han23, Hon24, KLK^{+}20a, Ma23d, QH23, SY23, Sun23a, WX23, WL23b, XYCC23, YB23, YDC^{+}24, ZW23c]. Digitization [Li23e]. Dilated [KYR^{+}23, TZL^{+}24]. dilemma [BFH96]. Dimension [CZH24]. Dimensional [Alk19, CTD23, DHW^{+}23, DDB^{+}18, JHL^{+}23, LLL^{+}23a, LZWL23, QZY^{+}20, WZW19, WFHC20, WW23b, WL23c, WLZJ24, WZ23d, YB23, ZXZZ23, ZZW^{+}23, HSS23, Kok07, KBRS95, LLQ^{+}23, WHG93]. Dimensionality [SZAG15, TF23]. Dimensioned [XYY^{+}19]. Direct [BHRT12]. directed [DZKS13]. Direction [BAP13, KLK^{+}20b]. Direction-optimizing [BAP13]. Directions [SKA^{+}20a, SKA^{+}20b]. Directive [XTCC15]. Directive-Based [XTCC15]. Directives [CCBPGA15, LOHA01]. Dirichlet [SLDL17]. Disambiguation [CYLD23]. Disc [WYHS23]. Discover [ZGW08, HB12]. discoveries [Gil09]. Discovering [DPMB16]. Discovery [GB15, LHZ^{+}23, TLZ^{+}23, AHCZ12, MPP^{+}04, WD07]. Discrete [LGL23]. Discretization [LPD20]. Discriminant [TF23]. Discrimination [CZCL23b, SW23, ZWJN23]. Discriminative [XXM^{+}23]. Discussion [KIU^{+}21]. Disease [AKA^{+}23, CZY^{+}23, HL23a, HCC23, KFU^{+}20, LZ23a, LWCZ21, TY23, WHL^{+}21, Wan23f, WWH^{+}23, XZLZ24, YZWF23]. Diseases [CPG^{+}23, DLP^{+}23, Hua21, LYZ^{+}23, ZYL23]. Disjunctive [NO18]. Dislocation [GWZ23b, ZRX^{+}24]. Disorders [LLGH24].

Dispatching [FLC^{+}20]. Display [LH23a, WC23a, WYC20b]. Dissemination [LH23a]. Distance [Ano23-27, Ano23-36, DYYW24, LG18, RKA^{+}24, RAZ^{+}23]. Distance-Aware [LG18]. Distancing [TMHM23]. Distant [BZKA20]. Distension [ZWW^{+}23b]. Distillation [LGZG23]. Distortion [CSZ^{+}18, ZWZ23a]. Distress [CL23b, LMQ^{+}24, SZW^{+}23, Zha23a]. Distributed [Ano23o, Ano23-48, BL18, BBG^{+}93, BMN^{+}97, CHL^{+}23, CLXZ23, FC22, GBH20, HYS^{+}19, KPS16, LX23c, Ma23a, MKS^{+}24, PSW23, SXTK18, SO11, SO15, SCKW23, TVCB23, TH18, YW24, ZY24, BKK^{+}11, DSS^{+}05, FP00a, FS01, FB99, HMCH07, JKR92, KMR^{+}97, KGV97, KABW11, LTB02, MSLS13, OB96, RLL^{+}02, RS95, SHHI01, Shu94, XNQF04, ZA10]. distributed-object [FS01]. Distributing [ASCH^{+}07]. Distribution [AAH^{+}23, BH23, CSWW23, Cui23, DZL^{+}24, JZ22, LLGL20, MAC^{+}20, MJB15, SL21, Su23, XWLL23, ZWW16, ZXHQ17, ZFFP24, ZZWL24, ACIK97, AGG^{+}97, CP95, KK09, NPP^{+}00]. distributions [KTP05]. Divergence [OL23, QL23]. Diverse [Ma23d]. divide [MGK^{+}13]. DLGWO [CYSZ23]. DLGWO-SVR [CYSZ23]. DMGA [CLXZ23]. DMZ [DRT^{+}14]. DNA [KTP05]. Do [GWH^{+}23, Ang93, GA96]. Doc2Vec [QGS23]. Documents [PMEH17]. Domain [DZL23, FYZ^{+}20, JBLB15, LY20, LX23a, LMZP23, LDZQ24, MAS^{+}23, SCD21, VVDM18, GRN99, JHNP14]. Domain-Based [MAS^{+}23]. Domain-Specific [VVDM18]. Dongarra [Nag05b]. Doppler [He24, YLLY23, ZLHZ23, ZZH22]. Dose [FMG^{+}23, Hon24, WCF^{+}21]. Double [JHL^{+}23, LFS^{+}20, YP24, Zhu23b]. Double-Layer [Zhu23b]. Double-Line [LFS^{+}20]. Double-Regional [JHL^{+}23]. Downregulates [DCH^{+}23]. DQN [CBE23]. Drama [Don23]. Drawing [VVDM18].

Drawn [AN19]. Drifting [XZNL23].

Driven [BR11, CWS^{+}24, DMV^{+}19, FYH23, HLW^{+}18, HCGGSS17, JWPW24, LFY^{+}18, LZ23c, Lv23, MZLC21, PPB^{+}19, RLPI19, Wan21, Zha22, Zhu22, Kes96, KVW^{+}07, Rou08b]. Drives [KPHY18]. Driving [LQW23]. Drones [LLGL20]. Drug [PP23, WCW^{+}21, ZHP^{+}23]. Drug-Target [ZHP^{+}23]. DSMPI [SSC97]. DSP [YZZ15]. Dual [ACCTA^{+}15, DYT23, LTF23, TZCZ23, ZGS16, ZLL20, ZPA19, ZWW^{+}23b]. Dual-Channel [LTF23]. Dual-Source [ZGS16]. Dual-stream [TZCZ23]. Duct [CZFX23]. Duplicate [LWL23b]. Duplication [RAZ^{+}23]. Durations [WLLY23]. during [AKA16, LTL^{+}23, XWLL23]. Duty [ZXX^{+}23]. DWD [SDS00]. DWI [ZLZ^{+}23]. Dynamic [AIA^{+}21, Ano23i, CMM^{+}02, CCG17, CZ23, CP03, GWZ16, HPZ16, HLPY16, HZFH23, JYLL23, JLX^{+}20, KLN^{+}04, LTB02, Li23b, LTZS18, LZYJ20, LQW23, RMX05, SX23, TH18, WD07, WHRH07, Xie20, YXL^{+}20, YYZ^{+}17, ZCG23, ZCL^{+}24b, ZGT^{+}23, HS03, MSCS14, RKR^{+}99, Shu94, VCT05]. dynamical [CO93]. Dynamically [XYY^{+}19]. dynamics [BCS01, BPK00, GIKP95, GHW00]. Dysfunction [Ano23n, TYCL23].

E-Commerce [Ano23-30, FYH23, Li24, Liu23f, YZL23, YT23, Ano23-45, ZCCS23]. E-Value [HMQS23]. E-Vector [MZLC21]. Each [CL23a]. Early [HL23b, HLM^{+}11, HSY^{+}23, Li23f, LLY^{+}23b, NHHW23, PLJ^{+}23, QWZP20, ZCZ21]. East [YHH^{+}23]. EC2 [TFN11]. ECA [CFL^{+}23]. Echocardiographic [HL23a]. Echocardiography [CTZC24, YXL^{+}20]. ECMWF [DTV00]. Eco [LQW23]. Eco-Efficiency [LQW23]. Ecocompensation [WKY23]. Ecological [WG22]. Ecology [WCZ^{+}16]. Economic [GZD23, Hu23a, Sun23c, Zho23]. Economics [HWY^{+}23, LQLX18]. Economy [XTZ20, ZM23]. Ecosystem [LQZ^{+}21, ZCCS23]. ed [Nag05b]. Eddy [XSZ18]. Edge [Ano23n, KLK^{+}20b, LG18, LXWD18, TMG^{+}19, WLP^{+}21, WZS^{+}22, WY23, WW23a, WJZC23]. Edge-Cloud [WLP^{+}21]. edited [Bry96]. Editor [Sne95]. Editorial [Ger02, O’K00]. editors [Pre99]. Education [Ano23z, Ano23-28, Ano23-46, Ano23-37, Ano23-45, Bai23, Chu21, DDJ20, Fen23b, Hon23, HL23c, LZ23b, LWW23a, LWX23a, LGZ23, LWX23b, Ma23d, SAZ^{+}24, She23, SW23, Sun23c, WL23b, Wan23h, Wan23d, Wan23a, WW23a, WC23b, XY23b, Yan23e, Zha23j, ZKA23, DDMS14]. Educational [LNH^{+}20, Liu23b, MAS^{+}23, MLHC20]. EEG [ADS95, CZB24, KJA^{+}23]. EEWMP [UAU^{+}23]. Effect [AKA16, Ano23d, Ano23-47, BLZG23, CWMZ23, CTZC24, FH23, Guo23a, HCC23, HZ23, JAED23, LZWL23, LTL^{+}23, SH19, TY23, WHL^{+}21, WLJ22, WCZW23, WLLY23, Wei23, WG22, WZ23d, YL23, ZXZZ23, ZRX^{+}24, ZZ23, ZW23c]. Effective [Don23, HLKY19, TSP^{+}15, Wu16]. Effectiveness [SBK^{+}24, XXSL23]. Effects [Ano23-34, FMG^{+}23, HLZ^{+}21, HZHW23, JHL^{+}23, RAdASGR^{+}19, SYW23, SWZL24, XLLL23, ZCLL23, CBCM93, GIKP95, LLW^{+}14]. Efficacy [GLD^{+}24, JLJ24, XZLZ24, Zha23a, ZY23, ZCL^{+}24b, ZWW^{+}23b]. Efficiency [AAF19, BKMS23, CGK^{+}05, HSS23, HW23b, JZH^{+}23, JFPL16, LKS18, LQW23, Mat94, MKH^{+}19, PLW^{+}20, SSSJ20, WZW19, ZCL^{+}20, ZM23, STD^{+}14]. Efficient [CCF^{+}17, CTCG15, CSM15, DMAC23, DRL13, EOS17, FAA^{+}20, HJ15, HL23b, IB09, LDZQ24, ML23a, MSA^{+}20, MOS16, MABU18, NTNTTK^{+}19, PTS^{+}13, RJA^{+}23, RuHMS24, RLC04, SGK^{+}23, SCAP16, SCD21, TT16, UAU^{+}23, WYY21, WDD^{+}23, XY21, XY22, ZXS^{+}23, CGL08, GL04, KABW11, SZ04, TFN11]. Efficiently [RLL^{+}02, HJ96]. Egg [Lia23b]. Eigensolver [BH96]. Ejector [CCBPGA15]. Elastic [ZHP^{+}23, MSCS14]. Elasticity [ALL^{+}16, Kok07]. Elastography [XYL^{+}23, ZCZ^{+}23]. Elderly [AAA^{+}23, Fen23a, HZL^{+}23, QFX^{+}23, Wan23f]. Electric [WCKD07, WCR^{+}21, XWL^{+}23, YX23]. Electrical [ZWLL23]. Electrocardiogram [YZWF23]. Electrocardiograph [LZ23a, MLY23]. Electroconvulsive [ZY23]. Electromagnetic [Ano22a]. Electronic [ZXHQ17, ZGT^{+}23, CL04]. Element [Ano07b, JBLB15, SO15, WZ22, VHBR93, WGF93, ZSS^{+}10]. Elements [Jua23c, Hem00]. Elephant [TZDZ20]. Elimination [PPJ^{+}15]. Elitist [NTNTTK^{+}19]. elliptic [AGIS94, JHNP14]. Ellis [Bry96]. ELM [ZFX17]. Emails [BKU^{+}23]. Embedded [CLN^{+}24, CLG^{+}23, JBLB15, MJK23, WLJ22, WLL23a, XYY^{+}19, YZZ15, PPS12a, PPS^{+}12b]. Embedding [BQR^{+}20, LXAC21, LCL^{+}23b, TZF^{+}18, TF23, WTY23]. Embedding-Based [TF23]. Embolism [ZZW^{+}23]. Emergency [CT23, DZL^{+}20]. Emerging [CSM^{+}18]. Emission [CPG^{+}23, TYCL23, WZS^{+}23]. Emotion [HZ23, HS23b, HZ24, NL23, Yu22]. Emotional [SK21]. Empirical [AFB23, AUA^{+}23, CTCG15, HBCM94, JZL^{+}20, ML23b, Sun23c]. Employability [Ano23b]. Employment [Ano23-33, Son23b]. EMRs [ZHZF20]. EMT [ACA19]. Emulating [Mor15]. Enabled [Alt23, ZTLH18, CDD^{+}05, KVW^{+}07, SHHI01]. Enabling [BTS^{+}14, GBH14, HJYC10, Mic97, XHSL17, vDKH01, Nor07]. Encephalopathy [KSYC23]. Encoding [RQD^{+}20, TT23, LSB12]. Encryption [GAH^{+}20, HGMM23, SAJ^{+}20]. Endometrial [HWSL22, LX23b, XZ23b].

Endoscope [DBWY24]. Endothelial [ZZD^{+}23b]. Energy [ACMT19, AAF19, CLYL16, CPK19, DMAC23, FX23, FAA^{+}20, FZW^{+}18, JFPL16, LLL^{+}23b, PBY^{+}20, QZY^{+}20, RuHMS24, SMB17, TT16, UAU^{+}23, WG21, WG24, Xie20, YX23, YXZ23, ZWL^{+}19, ZM23, ZHL16, CKS^{+}09, PRM^{+}14, STD^{+}14]. Energy-Aware [CLYL16, CPK19, FZW^{+}18, PBY^{+}20, Xie20]. Energy-Awareness [ZHL16]. Energy-Efficiency [JFPL16]. Energy-Efficient [DMAC23, FAA^{+}20, TT16, UAU^{+}23]. ENet [Wan23i]. enforced [WC96]. Engagement [SK21, Wan21]. engine [KVW^{+}07, NPP^{+}00, GGP09, IB09, KD09, SKU^{+}09, VGC09, WSZK09]. engineer [Hil97]. Engineering [AHMMÁR18, ÁRAHMM18, Ano07a, CHC15, Du23b, HCGGSS17, LHMC23, MBBDP17, Nag05b, HP02, MS00]. English [HNMB19, Ano23o, Ano23c, Ano23-27, Ano23-32, Ano23-36, Ano23-37, Ano23-38, ABNJMB23, Cao23b, Cha23a, CH23a, DC22, DS23, Du24, He20, He22, Hui22, Jua23b, Li23a, Liu20, LL23d, Pan23a, WLJ22, WL23b, Wei23, Yan23e, Zha22, Zha23m, ZW23b, ZSZ23c]. Enhanced [Ano23-31, CZB24, ZXZZ23, ZSZ^{+}23b, ZCL^{+}24b]. Enhancement [Ano23b, HLS^{+}24, LBY^{+}23, LMQ^{+}24, MLAS21, ZH23]. Enhancing [Ano19, FMA16, LDZZ23, RLU^{+}20, ZX17]. Enlightenment [Ano23z]. Ensemble [ACA19, HAW23, PLY^{+}23, SMD20, WZG^{+}17, ZQL23, AMN^{+}12]. Ensure [Bak23]. Enterprise [BH23, BYZ^{+}23, Hu23a, LZYJ20, PLJ^{+}23, YL23, ZCZ21]. Enterprises [Li23e, LWX23b, Pen23, Wan23c]. Entire [YZZ15]. Entity [HWJ18, TLW^{+}23, WSM^{+}21, ZHL^{+}23]. Entity-Aware [TLW^{+}23]. Entrepreneurship [Ano23p, Ano23b, Ano23-43, Ano23-45, LGZ23, Lou23, ZKA23]. Entropy [ASS23, CSZ^{+}18, CZ23, FYZ^{+}20, GSL23, LLC^{+}23, Ma23a, WG22, ZQ22]. Entry [HWY^{+}23]. Enumeration [EBDB24, SJ15]. Environment [CTC^{+}15, CLL^{+}20, GBK^{+}96, HKH^{+}17, KV23, KS19, Lia23a, QYZ17, SZ20, TPSM17, WWL^{+}23, WG22, XSS^{+}23, Xu23d, YCWL17, ZGFL23, BBF^{+}04, BGLR93, BDG^{+}94, BBLR03, CB99, FS01, HMC97, Lan01, Lar93, MSCS14, PVKE01, Pro07, VF95]. Environmental [Ano23-41, JZH^{+}23, YB24]. Environments [AK18, AHMMÁR18, ÁRAHMM18, CPK19, KPW^{+}18, LWW23b, OCC^{+}16, PGP19, RAdASGR^{+}19, WZW^{+}20, EAS^{+}97, GGMS99, GRC03, RLL^{+}02, VCT05, VRW^{+}03, ZA10, Bry96]. Ephedrae [XZLZ24]. Epicardial [WWH^{+}23]. epidermal [GGJ04]. Epilepsy [TYCL23]. Episode [ZY23]. Epoch [HJ15]. Epoch-Based [HJ15]. ePRO [CKS^{+}09]. ePRO-MP [CKS^{+}09]. Equalization [VMRM16]. Equation [AMU^{+}23, MRLF12]. Equations [YTD15, Gus93, PPS^{+}12b]. Equipment [CIU20, LWZ^{+}23a, LML^{+}23, SFZ^{+}20, SZL^{+}23]. ER-Store [LX23c]. Era [Ano23b, GZD23, Tia23, Xu23c, ZGL23]. Erasure [LX23c]. ERBB2 [Ano23-29]. Erection [Ano23i]. Erlang [IPST18]. Erosion [FCSL23]. ERP [AUA^{+}23]. Error [He22, tXChL^{+}24, YLWW18, ZSZ23c, HPS^{+}13]. Escort [QFX^{+}23]. Esophageal [Ano23-29, CLW^{+}23, LCW22, WZ23a]. ESports [HLZ^{+}21]. Establishing [OKHS23]. Estate [GWH^{+}23, SWDZ24]. Estimate [NLMT23]. Estimating [FCSL23, WXL19]. Estimation [AAH^{+}23, BF17, KLK^{+}20b, LJSJ23, WLL23a, WZS^{+}23, XWL19, gXySR^{+}17, ZL22, TOF^{+}14]. estimator [FP00a]. Estuary [PFQ^{+}23, KBRS95]. Ethnic [Jua23c]. Ethnics [AXA20]. EuroExa [ARAM19]. Europe [ZCL^{+}20]. EV [CHTZ23]. Evaluate [LQZ^{+}21]. Evaluated [ZZ23]. Evaluating [HWY^{+}23, HY15, LFQA20, SAJ^{+}20, SZ09, WmG21]. Evaluation [Ano23f, Ano23t, Ano23d, Ano23-34, Ano23-48, AA23, Cha23b, CTD23, Che23a, CTZC24, CEMM17, DWZ^{+}20, FHT23, Fen23b, FHC^{+}22, GSL23, GLD^{+}24, Guo23a, HLL23, Han23, HMQS23, Hui22, JHL^{+}23, LNK^{+}15, LLW^{+}22, LZ23b, LLW23a, LWZ^{+}23c, LWCJ23, PWP21, Pen23, QM23, SCD^{+}15, SFZ^{+}20, SXP23, SKT^{+}24, Son23b, Wan23h, WG21, WG22, WG24, XY23a, Xu23c, Xu23d, YLWW18, YL23, Yan23f, YN23, ZLS^{+}17, ZLF18, ZCL^{+}20, ZLZ^{+}22, ZXZZ23, Zha23a, ZRX^{+}24, ZWF23, ZCL^{+}24b, ZZW22, ZLF^{+}22, BC99, ZS99]. Event [AMH23, HXX16, HLH23, Kuf15, Ma23c, THS^{+}15, ZLX^{+}23, OPP11, PDRC18]. Event-B [PDRC18]. Event-Based [ZLX^{+}23, OPP11]. Event-Triggered [HXX16]. Events [CZH24, DMV^{+}19, Wan23f, ZA10]. Everything [JGDK23]. Evidence [JZL^{+}20, KV23, ZCZ21]. Evolution [Bor04, CYJ^{+}17, CCHJ23, CW23, HZD17, LY23a, LCH^{+}24, LQW23, MCZ23, NTNTTK^{+}19, QTD^{+}20, ZCCS23, ZCT^{+}20]. Evolutionary [Bia23, FZH23, WLL^{+}23b, tXChL^{+}24, YWH20, YXZ23, ZCLL23]. Evolving [FSS^{+}17, WM23a]. Exacerbation [XZLZ24]. Examination [Son20]. Examining [SXTK18]. Example [Ano23h, CM21, Son23a, WG22, ZG15, Den96]. Excess [SPH20]. Exchange [HH23a]. Excitation [LZZ^{+}23b]. Executing [LFQA20]. Execution [AB18, CTS23, JZZW20, Mar05, SAE^{+}23, TLZ^{+}23, FS01, HLM^{+}05, Slo12, ZS99]. executions [RR07]. Exercise [GWZ23b, He24, SYW23, WCZW23, ZQ22, ZLZ^{+}23, ZZ23]. Exercise-Induced [He24]. Exercises [KKD^{+}20, ZW23b]. Exhibition [HS23a]. Existing [Wu23b]. Expandable [GBH20]. Expands [WCZ^{+}16]. Expansion [AMU^{+}23, CZB24, WWZ17]. Expected [YZ16]. Experience [ZZJ^{+}24]. Experiences [ARGC19, CCF^{+}17, CvHK97, JR10, JDVM10, SM17, Lar93]. Experiment [Ano23z, OKHS23, BCCP05]. Experimental [WDY23]. Experiments [CM21, YPM^{+}21, BKK^{+}11, PDA^{+}08]. Expert [LL23c, OB96]. Explaining [SH94]. Explicit [SF03b, SF03a]. exploit [BPC^{+}09]. Exploitability [DWZ^{+}20]. Exploitation [ROMFST23]. Exploiting [HPD09]. Exploration [Ano23-27, JL23a, LF23b, SV15, Zha23h]. Explore [MVIA20, WZZ23]. Exploring [CYJ^{+}17, JK10, JIC^{+}14, MRLF12, ZGL23]. Express [DNS97]. expressing [JK10]. Expression [DyLJ23, Li23b, SDM^{+}19, Wan23b, XHSL17, ZC20, PCS99, ZGW08]. Expressive [XXM^{+}23]. Expressway [ZGT^{+}23]. Extended [DZ18, NL19, TC96, YZSZ17]. Extending [BCC^{+}00, NO18]. Extensible [MJK23]. Extension [Zha23i, PPD05]. Extensions [QYW^{+}17]. Extra [QGS23]. Extracting [BNH^{+}20, NCMF15]. Extraction [ABNJMB23, BZKA20, CLLZ23, DZL23, EOS17, LLT23, SH19, TLW^{+}23, YHZ^{+}23, YGZ^{+}24, ZHZF20, Zha23c, ZHP^{+}23, Abd02]. Extreme [HSS23, KAA^{+}23, WZ23d, XBG17, YWC^{+}18, ZGY^{+}20, BAN02]. Eye [SWDZ24, TS23a]. Eye-Tracking [SWDZ24].

Face [AXA20, ALA20, LCH^{+}23, TMHM23, XYY^{+}20, Yua20]. FaceFilter [ALA20]. Facet [WYHS23]. Facial [DyLJ23, Li23b, Wan23b, YN23, ZC20]. Facilitating [DMV^{+}19]. Fact [CSM15]. Factor [YCLK23, GGJ04]. factorial [PDA^{+}08]. Factorization [DEŁ01, TKZ^{+}24, CDO^{+}96, KD09, VGC09]. Factors [AA23, CLHH20, CL23b, CH23b, GWZ23b, JW23b, PWY23, Wan21, WCZW23]. factory [JMR^{+}11]. Failure [CIU20, LLY^{+}23a, SLS17]. Failures [LLY^{+}23b, MR18]. fair [BTS^{+}14]. Fake [HAW23]. Farming [KKA^{+}22, LSY20, MVIA20]. Farms [KKA^{+}22]. Fashion [CFY23, ZLS^{+}17]. Fast [Ano23-47, BSF96, CEMM17, JFPL16, LLL^{+}23b, QK21, SJ15, SKHZ24, WZW19, ZCS23, ZWZ23a, HST^{+}93, TVCB23]. Faster [XLY18]. Fastscat [HST^{+}93]. Fatigue [ZQ22]. Fatiguing [FH23]. Fatty [LYL^{+}24]. Fault [SLP18, SHW23, WLQ23, YYW^{+}23, ZCSL17, ZA10]. Fault-Aware [ZCSL17]. fault-tolerance [ZA10]. Faults [GRB23, HXX16]. FCN [TZC23]. Feasibility [TVCB23]. Feature [AK23, CLTX16, FTS^{+}20, HZJ23, HJO^{+}20, KS15, LWZ^{+}23b, Li23d, LYWS23, Liu23c, LXW24, MABU18, MWZ^{+}23, SMD20, SLDL17, SL23b, TZCZ23, WZW19, XJH23, YT22, YN23, ZLS^{+}17, ZWG^{+}20, Zha23c, ZXS^{+}23, ZYL23, ZHP^{+}23, ZZWL24]. Feature-Preserving [WZW19]. Features [BLZG23, CWMZ23, DRQ^{+}18, DCF17, DZL23, HL23a, HKQ^{+}23, JMGH21, LBW^{+}24, LDZZ23, LTR^{+}23, LJG^{+}23, Ton20, WZL23, WM23a, XXM^{+}23, YWX23]. Fee [ZMLZ18]. Feed [HI18]. Feed-Forward [HI18]. Feedback [FHS^{+}23, HPGC19, MPA16]. Feedback-Based [MPA16]. Feeder [SCKW23]. FEM [TZZ23]. Female [HLZ^{+}21, ZLZ^{+}22]. Femur [KYR^{+}23]. Feng [HSL^{+}23]. Festival [Wan22]. Fetal [CTD23, DHW^{+}23, Zho22, ZZH22]. FFT [Sch03b, TPKP13]. Fibrillation [LLY^{+}23a]. Fibrosis [ZCZ^{+}23]. Field [Ano23h, BL23, BLL^{+}23, CSZ^{+}19, LHC^{+}23, MLL23, WX23, VHBR93]. Fields [ACCTA^{+}15]. file [DZKS13, Met99a, Met99b]. Files [Ma22, CGL08]. Film [LZWL23]. Filter [ALA20, AZGJ24, LLL^{+}23b, XLLL23].

Filtering [FGROL^{+}16, JHL^{+}23, LZZ23a, LLP23, TKZ^{+}24, Wan23d, XWL19, XZNL23, ZYY^{+}23, Zha23j]. Filtering-Based [XZNL23]. Filtration [DLP^{+}23, PWY23]. Final [ZGS16]. Finance [ZW23c]. Financial [Hu23a, HW23a, JZL^{+}20, Li23e, Lin23a, MSZZ20, PLJ^{+}23, Pen23, SXP23, Yan22b, ZCZ21, LPHD04]. Financially [LTF23]. Financing [BH23]. Findings [Xue23]. Fine [JMGH21, JW23a, LLP^{+}19, ZZWL24, HPD09]. fine-grain [HPD09]. Fine-Grained [JMGH21, JW23a, LLP^{+}19, ZZWL24]. Fingerprint [KLK^{+}20a]. Fingerspelling [JHS^{+}20]. Finite [Ano07b, JBLB15, Vol97, VHBR93, WGF93, ZSS^{+}10]. Fink [Hil97]. Fire [CHL^{+}23, HZJ23, ZZWZ23]. Firefly [KAA^{+}23]. Firmware [QGS23]. First [ARAM19, LPBSSEPL19, ZY23, BAP13]. Fish [LCW22, LZLZ19, ZFFP24]. Fitting [VRM02, XWLL23]. five [GA96]. Fixed [Zha23l]. Flexibility [YZSZ17]. Flexible [ESSL99, HS23a, JNK24, QFX^{+}23, RS94, SW20, WCLL17, YZSZ17, GHST12, MSLS13]. Flight [AMB23, YZSZ17]. Flipped [Cha23a, DS23, Liu20, SF23, XXSL23]. Flood [RAA23, MCvM10]. Floor [GWH^{+}23, ZLZ^{+}22]. Flow [AFU18, ASZF23, Ano23i, BLL^{+}23, CLL^{+}18, CCBPGA15, LGL23, LHC^{+}23, MLAS21, REPRL^{+}19, SQ21, TZDZ20, TY23, TLH17, WCF18, WZZ23, lWwFAH23, XBG17, Xu23a, ZLLZ20, ZLCW20, ZWG^{+}20, ADLN08, FP95, GR93, WB95]. Flow-Shop [SQ21]. Flowchart [ZLLS20]. Flowchart-Based [ZLLS20]. Fluid [QK21, SWH23, WBZL17, WZZ23, GHW00]. Fly [HJ15]. Flying [OQZ23]. FMS [XHX23]. FOA [PLJ^{+}23]. FOA-SVR [PLJ^{+}23]. Focal [PRW23, XLF^{+}24]. Fog [LGZ23]. Following [YNK23]. Foods [ZXHQ17]. Football [Che22, FWX23, Hu23b, HLH23].

Footballer [YYY23]. footprint [SVR^{+}07]. force [AJ94]. Forecast [CFL^{+}23, SAN^{+}19, Xia23a, YY23a, Ye23, ZWG^{+}20, Zho23, DTV00, Ros00]. Forecasting [LYWS23, MAC^{+}20, PPB^{+}19, XBG17, ZCW^{+}19, ZGFL23, WCG95]. Foreground [Hu23b]. Foreign [Ano23l]. Forest [CCB^{+}20, LZK^{+}23, STM^{+}23, WDZY20, XY23b, Yan22b]. Foreword [Ano05, CK08]. Forgery [RAZ^{+}23]. Formal [RAA23]. format [CGL08]. Formation [SCAP16, BAN02]. Formed [KZK^{+}23]. Formula [Ano23-44, AJJF14, HSL^{+}23]. formulation [KHSJ95]. Fortran [Ben99, Ano93, DDS99, Zim07, ACIK97, AJJF14, AGG^{+}97, BCS01, BCC^{+}92, BMN^{+}97, CMZ92, DS97, Hig93, HMR^{+}15, HZ99, KMR^{+}97, KOM94, KGV97, LP99, McC96, MH95, Met99a, Met99b, Mor15, NCMF15, NL19, NDSG07, OPE^{+}95, OPB^{+}20, PMM94, PCS99, RRM^{+}15, RN07, RMX05, Sch03a, SM17, Sny07, Szy07]. Fortran-like [KGV97]. Fortran-P [OPE^{+}95]. FORTRAN/journal [Ano93]. Fortran90 [DNS97]. Fortran95 [DN04]. ForTrilinos [MRLF12]. ForUML [NCMF15, NL19]. Forums [KIU^{+}21]. Forward [ACCTA^{+}15, HI18]. Foundation [LHMC23]. Foundations [ZGL18]. Founded [NO18]. Four [YZL23]. Fourier [TVCB23]. FP [JFPL16, ZYLZ15]. FP-ABC [JFPL16]. FP-Growth [ZYLZ15]. FPGA [HGMM23, OCC^{+}16]. FPGA-Aware [OCC^{+}16]. FPGAs [ARAM19, dBMT20]. FPPM [JNK24]. Fractal [LZZ19, TXD^{+}20]. Fractional [PDA^{+}08, SCW^{+}23, XC23]. Fracture [FHC^{+}22, KYR^{+}23, XLB^{+}23]. Fractures [HZL^{+}23, PWY23]. Fracturing [YLLL20]. Fragmentation [GWZ16]. Frame [HL23b, LLT23, RAZ^{+}23]. Framework [BGLR93, BL18, CKKK15, CSU^{+}17, DWZ^{+}20, FGROL^{+}16, FAA^{+}20, HAA^{+}22, LCL^{+}23b, MAS^{+}23, MBBDP17, PZC^{+}23, Pla04, SDM^{+}19, SZAG15, SW16, SZXL16, Ton20, YL20, ZHW^{+}21, ACB^{+}02, ACIK97, BDM^{+}04, BBK^{+}11, BHK^{+}13, CR05, DSS^{+}05, MCvM10, OPP11, TPKP13]. Framework-based [BGLR93]. Free [ZWW16, Bjø00, GHW00]. Freight [HWY^{+}23, ZCL^{+}20]. Frequency [Che23b, CSM^{+}18, PPB^{+}19, SZW^{+}23, XLLL23, YPY23]. Frequent [DQQ17]. Fresh [SL21, ZXHQ17, ZYZL23]. Fried [XZLZ24]. friendly [Hem00]. Fuel [ZXX^{+}23]. Fujitsu [DTV00]. Fulfilling [ZXHQ17]. Full [LSS^{+}23, LZLZ19, PGP19]. Full-Body [LZLZ19]. Function [eLA20, SLH23, WCZW23, ZCZ21, ZLZ^{+}23, ZWLL23, KMB09]. Functional [BH96, FZH23, Jua23b, ZCLL23, DFP96, Den96, WC96]. Functions [HKVVNT18, NL23, SO18]. Fusion [Ano19, Ano23k, DFT^{+}23, HSH23, HYK^{+}20, JZ22, LLL^{+}23b, Lin23b, LYWS23, LLC^{+}23, LJG^{+}23, LYZ^{+}23, MWQ16, PP23, SMD20, TTL^{+}23, Ton20, WG23, XZ23c, ZHZF20, ZWW^{+}23a, ZHL^{+}23, ZXZZ23, ZXS^{+}23, Zha23b, ZWZ20, ZYL23, ZX17, PKE^{+}10]. Future [CLPVPI20, SKA^{+}20a, SKA^{+}20b, JIC^{+}14]. Futuristic [JGDK23]. Fuzzy [AUA^{+}20, BBI23, Cha23b, CZP^{+}24, DYT23, HBR^{+}20, Han23, HH23a, HHZ^{+}23, JFD20, JZY23, LX23d, Liu23b, NM16, SLS17, SZC16, SW20, SWZL24, Wan23a, Wen23, WmG21, XNL^{+}22, Xu23d, ZLSW23, ZGY^{+}20, ZLL20]. FWA [ZFFP24]. FWA-Artificial [ZFFP24]. FWHT [PLY^{+}23]. FWHT-RF [PLY^{+}23].

G [ABB^{+}02, KVW^{+}07, TF04]. Game [Bai23, Che22, DZL23, YWH20, ZL23a]. gamma [BCHL05]. gamma-ray [BCHL05]. GAN [QWP23, TZC23, Wan23b]. gap [SH94]. Garden [Ano23j, He23]. Garment [ZPA19]. Gas [SLS17]. Gasket [CWSF23]. Gastric [CLF^{+}23, DBWY24, ZTL^{+}23].

Gastrocnemius [Xue23]. Gastrointestinal [CLW^{+}23]. Gated [LXL^{+}23]. Gateway [BTV^{+}20]. Gathering [Wei23]. Gaussian [CZCL23a, JFD20, PLAAH23, WZ23c, XWLL23]. GCR [CKKK15]. GDMP [SSM^{+}02]. GDP [QHJ23]. Geant4 [SCD^{+}15]. Gender [Alk19, HLZ^{+}21]. Gene [DCH^{+}23, MABU18, XHSL17, ZGW08]. General [LWZ^{+}23b, WCW^{+}21, YWGL20, Hem00, Mat94, RR07, SS00, SF03b, SF03a]. Generalized [CBRRC^{+}16]. Generate [CLPVPI20]. Generated [BRSP18, CTS23, FHS^{+}23]. Generating [SO15, BCC^{+}92]. Generation [Ano23-42, AMT20, CGA17, Ma22, NL19, SZWZ23, Xia23a, YWGZ20, ZLLZ20, Zha23h, BDM^{+}04, BHK^{+}13]. Generation-Based [ZLLZ20]. Generative [HZFH23, ML23a, WLQ23, Xia23a]. generator [SF03b, SF03a]. generic [BH12, PPS12a, VHBR93]. Genes [LWQ^{+}23]. Genetic [BRSP18, Cha23b, CLTX16, Don23, KS15, MABU18, NHHW23, Pen23, STM^{+}23, Wan23h, XWG23, Yan23f, Zen22, Zha23k]. Genome [Ma22]. Genres [Xu23b]. gentle [Ste97]. Geological [HCQ17, ZZX^{+}17, ZTLH18]. Geometric [CSZ^{+}19, DCF17, WZW19]. Geometry [Cer21]. Gestational [LWZC21, ZLHZ23]. Gesture [CZKY24, MCF^{+}23]. GFDL [Hem00]. ghosting [MSLS13]. Giga [GKL^{+}96]. GIS [BRPV24, Cui23, FCSL23, ZCW^{+}19]. GIS-Based [ZCW^{+}19]. GIS-SWIAS [BRPV24]. GitHub [LQZ^{+}21]. Glaucoma [ZC23]. Glioma [YT22]. Gliomas [YHJ^{+}23]. Global [AKA16, GSL23, KGMN15, Pan23b, Ros00, WB16, XYY^{+}19, YYF^{+}19, ZAW^{+}21, FC01, HMCH07, Wan02]. Glomerular [DLP^{+}23]. Glowworm [TH18]. Glucose [ZMM^{+}23]. Glucose-Induced [ZMM^{+}23]. Glycosylation [LJ23]. GM1 [WHL^{+}21].

GMA [KCO^{+}05]. GMRES [PBK01]. Goal [Yan23d]. Gockenbach [Ano07b]. Gold [ABNJMB23]. Golgi [CCB^{+}20]. Gonadotropin [QYL23]. Gonadotropin-Releasing [QYL23]. Good [Ano07c, Mat03]. Goods [HH23a]. GOOSE [SCKW23]. Gordon [Per08]. Governance [WG22, Zen22]. governed [BH02]. Government [HL23d, MCZ23, OL23, WG22]. GPS [ZWW^{+}23a]. GPU [CFR14, CCBPGA15, GW11, Jes10, LYTL20, LPBSSE17, PGP19, TKZ^{+}24, TLH17, XTCC15, gXySR^{+}17, YTD15, YSL^{+}18]. GPU-Accelerated [TKZ^{+}24]. GPUs [BBDN11, EBDB24, GA18, XSZ18]. Grade [qCtWfPbZ23, YHJ^{+}23]. Gradient [KLK^{+}20b, KAA^{+}23, WWW^{+}17, XC23, GG05]. Graham [XZNL23]. Grain [LWL^{+}23a, HPD09, KB96]. Grained [JMGH21, JW23a, LLP^{+}19, ZZWL24]. Gram [SJD20]. Grammar [Ano23-32, He22, LL23d, ZSZ23c]. Grams [AAS^{+}23]. Granularity [WS15]. Granulation [HH23a]. Graph [Ano23e, BQR^{+}20, CD23, CYLD23, CLXZ23, FLC^{+}20, HLY^{+}17, HYS^{+}19, LL23c, Lin04, PMCF94, SQW^{+}21, TZF^{+}18, TF23, VVDM18, YSL^{+}18, ZG18, ZZX^{+}17]. Graphical [CYZ^{+}20, PHH95]. Graphics [GLBB17, PGP19, ZG23, BB09]. Graphs [LXAC21, Ser18]. Gravitational [ACCTA^{+}15]. Gravity [JAED23]. Gray [Ano23-33]. Great [Ste97]. Greater [LF23b]. Green [Ano23j, HYK^{+}20, HXPZ24, PBY^{+}20, QHJ23, XZ22]. Greenhouse [QZY^{+}20]. Grey [Guo23a, WH20, WM23b, ZCL^{+}20, ZLL20]. Grid [DK02, KCO^{+}05, ZL23c, ACB^{+}02, CJS^{+}02, GRC03, MPP^{+}04, Pla04, HP02, ABB^{+}02, ACB^{+}02, AMM05, BBF^{+}04, BBŁ08, COB^{+}00, CQF05, HHML05, HLM^{+}05, KFFZ05, KVW^{+}07, KLN^{+}04, MAG^{+}07, PPD05, Pro07, Rom02, SZ04, SSM^{+}02, TF04, VCT05]. Grid-based [ABB^{+}02]. Grid-enabled [KVW^{+}07]. GRIDCC [MAG^{+}07]. GridLab [KLN^{+}04]. Grids [PPJ^{+}15, SN02, ITF^{+}08, Özt04, SLJ^{+}00]. GridWay [HHML05]. Grinding [WS15]. GROMOS96 [BCS01]. Groovy [PSM^{+}15]. GroovyLab [PSM^{+}15]. Groundwater [MVIA20]. Group [CD23, DZL^{+}20, DPMB16, JYLL23, QQWJWYQZ23, SL21, ZLL20]. Growing [WSL23]. Growth [WQLR17, ZYLZ15, ZZH22, GGJ04, Szy07]. GRU [GWZ23a]. Grunwald [Bry96]. GSSIM [BKK^{+}11]. GTS [PKE^{+}10]. Guangdong [LF23b]. Guangdong-Hong [LF23b]. Guarantee [LGJ18]. Guest [Ger02, O’K00, Sne95]. Guest-Editorial [Ger02, O’K00]. Guidance [Yua20, ZWLL23]. Guide [Ano07c, Hil97]. Guided [Ano23-47, CLW^{+}23, HZL^{+}23, HZHW23, HSY^{+}23, LZWJ23, LCL^{+}23a, MWZ^{+}23, PP23, WCW^{+}21, Ye23]. Guidelines [ACMT19, DDJ20, GB94]. Guides [HSL^{+}23]. Guillotine [BRSP18]. Guillotine-Cutting [BRSP18]. Gym [KKD^{+}20]. Gynecological [WLLY23, WSL23].

H [Hil97, Mic97]. H.266 [ZCS23, ZWZ20]. H.266/VVC [ZCS23, ZWZ20]. Hadoop [LKS18, SZXL16]. Hadoop-Based [LKS18]. Hall [HW23b]. Hand [AN19]. Hand-Drawn [AN19]. Handling [Che93a, ZZD23a]. Handwritten [LGJ18]. Hangzhou [ZJZ23]. Hardware [CM21, DMAC23, DGH^{+}15, GLBB17, MOS16, SLP18, vDKH01]. Hardware-in-the-Loop [CM21]. Hardware-Level [SLP18]. Hardware/Software [MOS16]. Harmonic [DZL^{+}24, PSW23, ZSD22]. Harmony [CCG17, YLWW18]. Harvesting [LWL^{+}23a]. Hash [Ma23a, RQD^{+}20, ZD23a]. Hashed [JNK24, War14]. Hashing [FS24]. Hazardous [WQT^{+}20]. Hazards [YLW21, YP24]. HBase [CTCG15]. HCRCaaS [LGJ18]. Head [DM23, SW23]. Health [Ano19, KAA^{+}20, KJA^{+}23, Li23c, LLLZ23, NHHW23, SL23b, WWL^{+}23, YYY23, ZCZ^{+}23, ZX17]. Health-Related [KAA^{+}20]. Healthcare [ASS23, FAA^{+}20, HBR^{+}20, HYK^{+}20, NGMAA23, RMS21, RLU^{+}20, Wen23, hKJ^{+}20]. Heart [HL23a, LZ23a, LLY^{+}23a, Wan23f, WWH^{+}23, WDD^{+}23, YZWF23, ZZH22]. Heavy [YY23c, ZXX^{+}23]. Heavy-Duty [ZXX^{+}23]. Hebei [ZFFP24]. Heckman [WKY23]. help [BFH96]. Hemangioma [ZZD^{+}23b]. Hematoma [BLZG23, Xu23f]. Hemodialysis [Hon24]. Hemodynamic [YLLY23]. Hemorrhage [DM23, WCF^{+}21]. HeNCE [BDG^{+}94]. Hepatitis [ZCZ^{+}23]. Herba [XZLZ24]. Heritage [Lvp23]. Herniation [WYHS23]. Heroes [FYY23]. Hesitant [DYT23, ZLL20]. heterogeneity [LR05]. Heterogeneous [BB09, CCTY15, HMSW92, LPBSSE17, LWW16, SRO^{+}15, TT16, WWLG16, WCF18, WQT^{+}20, WWW^{+}17, WFZ^{+}23, ZCSL17, BCHL05, BDG^{+}94, BBK^{+}11, BDH^{+}10, DRL13, JHNP14, KK11, PVL^{+}04, PRM^{+}14, RLL^{+}02]. Heuristic [FH23, LWW16, PBY^{+}20, ZWXL21]. Heuristics [KGMN15, SLBZ16, FB99]. Heuristics-Based [SLBZ16]. HEVC [gXySR^{+}17, ZWS19]. Hidden [KZZ^{+}20, XML^{+}16, BH02, LB02, VRM02]. Hierarchical [AAA^{+}23, LLZW23, SLDL17, TLZ^{+}23, WYC^{+}20a, BDV03, DZKS13]. Hierarchically [YTD15]. Hierarchy [JUA^{+}23a, BPC^{+}09]. High [Ano99b, Ano08b, Ano23i, BKMS23, CTCG15, CSM^{+}18, CPK19, CPD21, DD23b, GM14, GGP09, GBK^{+}96, HMR^{+}15, HKL^{+}12, LLC^{+}20, Liu23d, LBvBW12, Nag05a, Nor07, PPB^{+}19, SZW^{+}23, TY23, WSZK09, XLLL23, YSL^{+}18, YXZ23, YPY23, YGZ^{+}24, YWGZ20, ZMM^{+}23, ZLZ^{+}21, dBMT20, AB96, BTS^{+}14, CB99, FP00b, GGMS99, IR02, KMB09, Lan03, LG03, MMG^{+}02, NJ08, VvAC^{+}09, Zim07, ACIK97, Ano93, AGG^{+}97, BMN^{+}97, DS97, KMR^{+}97, KOM94, KGV97, MH95]. High-Density [YXZ23]. High-Efficiency [BKMS23]. High-Flow [Ano23i]. High-Frequency [CSM^{+}18, PPB^{+}19, SZW^{+}23, XLLL23, YPY23]. High-Level [YWGZ20, dBMT20, LBvBW12]. High-Performance [CPK19, CPD21, GBK^{+}96, HMR^{+}15, LLC^{+}20, YSL^{+}18, HKL^{+}12, AB96, FP00b, KMB09, LG03]. High-Resolution [YGZ^{+}24, ZLZ^{+}21, VvAC^{+}09]. Higher [Fen23b, HL23c, LWW23a, WL23b]. Highly [gXySR^{+}17, GBH14]. highly-scalable [GBH14]. Hip [LWZ^{+}23b, WCW^{+}21]. Hitting [Qin23]. HMM [Abd02]. HMM-based [Abd02]. Hoc [JRS^{+}21, BDM^{+}04]. Hole [YDC^{+}16]. Holiday [Wan22]. Holmium [LCK^{+}23]. Home [MAA^{+}21]. Homogeneity [ZWS19, KTP05]. Homogeneous [DGL^{+}17, SMD20]. Homophily [KZK^{+}23]. Honey [XZLZ24]. Honey-Fried [XZLZ24]. Hong [LF23b]. Horizontal [JW23a, YLLL20]. Hormone [QYL23]. Hospital [DRMG^{+}18, WL23c]. Hospitals [JW23b]. host [Din99]. Hotspot [CLLZ23, LWCZ21, Wei23]. Households [LLZW23]. Houses [LY23b]. Housing [Zha23e]. HOV [LWCJ23]. HOVA [JNK24]. HOVA-FPPM [JNK24]. HPC [BFGH14, DGH^{+}15, KPW^{+}18, LYTL20, STD^{+}14]. HPF [BDV03]. HPGraph [YSL^{+}18]. HPM [SJD20]. HSIP [WWLG16]. HSR [HWY^{+}23]. Huaihe [FYH^{+}19]. Huanglian [Xu23e]. Hull [XZNL23]. Human [Alk19, Ano23w, Cao23a, CFY23, JZ22, PZ24, QS23, RVJ^{+}23, SL23a, SCW^{+}23, Sun23b, WM23b, Zho20, Zhu23a].

Human-Computer

[Cao23a, PZ24, RVJ^{+}23]. Human-Machine [Ano23w, Zho20]. Humanistic [WLJ22]. Humphrey [Ano07a]. Huss [Nag05b]. Hybrid [ÁRCPS15, Ano23f, ATZ^{+}15, CLN^{+}24, CLLZ23, CIU20, CL23c, CZWY23, DFP96, FCZ^{+}17, KS15, LYTL20, LZZ23a, LX23c, LLGL20, LJ23, Ma23d, MABU18, NTNTTK^{+}19, PWP21, PRW23, PCYZ18, STM^{+}23, SJD20, SW16, SZC16, SQ21, TZZ23, Tia23, Ton20, WS15, WXZW17, XJH23, YX23, ZLCW20, CFR14, JR10, KeKR^{+}11]. Hydrocephalus [SWH23, WZZ23]. Hydrology [CLPVPI20]. Hydrolyzed [ZWW^{+}23b]. Hydrostatic [Sah20]. HyperLogLog [BF17]. Hyperspectral [PFQ^{+}23, QWZP20, ZWJN23]. Hypertension [DM23]. Hypervisor [OCC^{+}16]. Hypothesis [YWGZ20]. Hypoxic [KSYC23]. Hypoxic-Ischemic [KSYC23]. Hysteresis [WS15]. Hysteroscopy [HWSL22, LX23b].

I/O [OSLK12, LKS18, NsP16]. IaaS [TPSM17]. IBM [KGBB09, Was95]. Ice [Ano23-28]. ICS [SX23]. Ideas [BBG^{+}93, Ste97]. Identification [ALA20, Ano23u, ABR17, CSWW23, CCB^{+}20, DQQ17, Guo23b, HL23b, JHS^{+}20, KZZ^{+}20, LPBSSEPL19, LWQ^{+}23, LYL^{+}23, MZLC21, SZX^{+}20, SKT^{+}24, SCR20, XC23, YLLL20, ZWF23, ZHP^{+}23, ZGT^{+}23]. Identifying [AXA20, Ano23-29, LLZW23, ZCXQ16]. Ideological [Ano23e, Ano23x, Ano23-46, Hon23, JL23b, SF23, Wan21, Wan23d]. IDL [ESSL99]. IE [ASS23]. IEC [CSWW23]. IFS [DTV00]. Ignorant [SMB17]. II [PPS^{+}12b, Per08, YPM^{+}21]. III [YLL^{+}23, YPM^{+}21]. IIOT [LFQA20, CZS^{+}23]. IIOT-Industrial [LFQA20]. Illegal [YDC^{+}24]. Illumination [Li23b, LBY^{+}23, RAdASGR^{+}19]. Image [AHM23, AXWS23, Ano23d, Ano23-39, Ano23-47, Ano24, BLL^{+}23, CLDX20, CWMZ23, CZD^{+}20, CLW^{+}23, CPG^{+}23, CTZC24, DRQ^{+}18, DLP^{+}23, DYT23, FTS^{+}20, FLR23, FZW23, Fen23a, FS24, FHC^{+}22, FMG^{+}23, GAH^{+}20, HGMM23, HY23, Hua21, HH23b, JZS21, KSYC23, KLK^{+}20b, LPBSSEPL19, LWZ^{+}23b, LBY^{+}23, LLL^{+}23b, Li23d, LF23a, LBW^{+}24, LW23, eLA20, Liu23e, LLW23a, LTL^{+}23, LDW23, LMQ^{+}24, LXW24, LLGH24, LH23b, LYZ^{+}23, LML^{+}23, ML23a, NHHW23, PP23, PZ24, QS23, Qin23, SLC^{+}23, Sun23b, SL23c, TVCB23, TTL^{+}23, TZC23, WCW^{+}21, WHL^{+}21, WZL23, WL23c, Wan23i, WG23, WZ23c, WYC20b, Wu23c, WJZC23, WZ23d, WFZ^{+}23, Xia23b, XWLL23, XLLL23, XXM^{+}23, Yan23g, YGZ^{+}24, YT22, YB24, Yu22, YPC^{+}23, ZCL^{+}24a, ZSD22, ZHY22, ZYZ23, ZXZZ23, ZH23, Zha23k, ZD23a, ZLZ^{+}23, ZSZ^{+}23b, ZZDY24, ZZ23, Zho22, ZWJN23, ZHHX23, ZWW^{+}23b, GKL^{+}96, RLC04]. Image-Assisted [CPG^{+}23]. Image-Guided [Ano23-47, WCW^{+}21]. Imagery [SWZL24]. Images [Ano23v, Ano23u, Ano23-31, BGZL23, CLF^{+}23, CZFX23, CZCL23b, CH23b, CWLW23, Cui23, DBWY24, FLY^{+}23, GWH^{+}23, GWZ23b, GSL23, GLD^{+}24, GAH^{+}23, GGSL23, HLKY19, HZL^{+}23, HZHW23, JLJ24, JHL^{+}23, Li22, LZL^{+}23, LWL^{+}23c, Li23b, LSS^{+}23, LCH^{+}23, LCK^{+}23, LL23a, LHC^{+}23, LCL^{+}23a, LWZC21, ML23b, ML24, MWZ^{+}23, QWZP20, SLH23, SYW23, SKT^{+}24, SWZL24, SWH23, TZH20, TXD^{+}20, WZS^{+}22, Wan23f, WWH^{+}23, WYHS23, WZS^{+}23, WZC^{+}23, WSL23, WMZM23, XZZY20, Xia23a, Xu23f, XWLL23, XZLZ24, ZYY^{+}23, Zha23f, ZZZ^{+}23, ZCLL23, ZC23, ZWLL23, VvAC^{+}09]. Imaging [Ano23i, Ano23n, BLZG23, CL23b, DBWY24, DM23, FZH23, FHC^{+}22, HLL23, HSS23, He24, HTZL23, HHZ^{+}23, HSY^{+}23, JFY23, LWZ^{+}23b, LX23b, LX23d, LWZ^{+}23c, Liu23c, LTR^{+}23, LYL^{+}24, LYZ^{+}23, MLAS21, OWL^{+}23, PTZL23, PWY23, QYL23, SWY23, TY23, TYCL23, WCF^{+}21, WZ23a, WCZW23, WZZ23, WLZJ24, XYL^{+}23, Yan23b, YLL^{+}23, YT22, YHZG23, tZhYsHlZ19, ZZZ^{+}23, ZCLL23, ZY23, ZCL^{+}24b, ZTL^{+}23]. Imaging-Assisted [Liu23c]. Imaging-Magnetic [HTZL23]. Imbalance [CZW^{+}23, LLC^{+}20, WZF24]. Imbalanced [BC23, Ona19, WZG^{+}17, ZLDX21]. Immune [FMG^{+}23, HSL^{+}23]. Impact [AWB^{+}23, CLHH20, CZH24, Chu21, DYW^{+}23, FCSL23, Guo23b, GRB23, HCM23, HLW^{+}18, Hu23a, Li23e, LZ23b, Lou23, MSZZ20, PZH23, WHL^{+}18, YB24, FSV14, HLM^{+}05]. Impedance [ZWLL23]. imperative [DFP96]. Imperial [Su23]. Implantation [BLL^{+}23, LHC^{+}23]. Implement [CM21]. Implementation [ABR17, BHP^{+}03, Cao23a, Cao23b, CTC^{+}15, CSU^{+}17, HGMM23, HT22, KL23, KMB09, Liu23b, MYH19, NsP16, SKU^{+}09, SSC97, TOF^{+}14, TMHM23, XSS^{+}23, BS01, BDV03, CBCM93, CDO^{+}96, CIN^{+}96, CW93, FS01, Gus93, PPD05, PCS99]. implementations [ADS95, BH96, PBK01]. implemented [Hav00b]. Implementing [Ano07b, CS94, HJ96, Lan03, VGC09, YZZ15, DN04]. implications [DF14]. Implicit [FHS^{+}23, HPGC19, SCSJ09]. Implicit-Feedback [HPGC19]. Important [Wan21]. Improve [AlS16, Ano23-28, GB15, KPHY18, Sin99]. Improved [Ano23-38, AB18, CLL^{+}20, CZY^{+}23, CBE23, FYZ^{+}20, FS24, GAH^{+}20, GZD23, HGMM23, HLHC23, HFH^{+}23, KZZ^{+}21, LXWD18, LLP23, LZZ^{+}23b, LHZ^{+}23, Liu23g, LZ23d, LLT23, Ma23a, Ma23b, Pan23a, QWZP20, QL23, RAZ^{+}23, SSSJ20, Su23, TZCZ23, TMY^{+}23, Tun23, WCLL17, WYC^{+}20a, WH20, Wan23b, Wan23d, WLQ23, Wan23i, WZ23c, Wu23c, yXHtJC23, XSG23, XLY18, YLWW18, Yan23g, Yan23f, Yan23d, YY23b, Yin23, ZYLZ15, ZGFL23, ZC20, ZZDY24, War14, vDKH01]. Improvement [PVCPdPSZVG17, SNH24, XNLJ23, XHL^{+}23]. Improving [Ano23-42, BBDN11, CTS23, EBDB24, HKH^{+}17, HLS^{+}24, HSL^{+}23, HH23b, LKS18, LTZS18, RR04]. Impulse [CHS^{+}99]. ImReMuDF [LYL^{+}23]. IMRT [FLY^{+}23]. In-Depth [DZL23]. In-Domain [LY20]. Inastemp [Bra17]. Incidence [NHHW23]. Incident [CGMRO^{+}19]. Incision [TZZ23]. Inclusive [ZW23c]. Income [SWZS17]. Incontinence [ZLZ^{+}22]. Incorporates [GPP^{+}17]. Incorporating [BGH99, LWCZ21, SPH20]. Increase [MLAS21]. Incremental [KLL^{+}18, WM23a, XWG23, ZG18, TOF^{+}14]. Independent [ZCT^{+}20, GL04]. Index [Ano99a, Ano00, Ano08a, Ano09, Ano11a, Ano12, Ano14, CTC^{+}15, LXRC21, SCR20, WZC^{+}23, XML^{+}16, YCLK23, ZL23c, Ano97a, BSF96]. Indicator [AFB23, FSV14]. Indicators [ZZS^{+}23]. Individual [GGJ04, MCF^{+}23, WBZL17]. Individual-based [GGJ04]. Indoor [CHL^{+}23, LWYS16]. Induced [He24, Zha23f, ZMM^{+}23]. Industrial [HNG^{+}23, HL23b, JZ22, LJF20, LZK^{+}23, Nag04, OKHS23, OZI^{+}23, SKA^{+}20a, SKA^{+}20b, SZWZ23, Son23a, WJCZ21, Yan23c, ZCXQ16, LFQA20]. Industry [Ano24, CW23, GMS^{+}23, HSH23, JZH^{+}23, JZS21, LH23a, LGZG23, TH23, XZ23a, XNLJ23, Xu23d]. Inertial [FSLG17, WW22]. Inexact [ZHZS19]. Infantile [ZZD^{+}23b]. Infarction [LWZ^{+}23c, ZCL^{+}24a]. Infection [OWL^{+}23, PWY23, ZCZ^{+}23]. Infection-Related [PWY23]. Infectious [LLW23a, WHH^{+}21, ZYL23]. Inference [BBI23, SW20, WXL19]. Inferior [CZFX23]. Inferring [KZZ^{+}20, PSU08]. Inflammation [HSL^{+}23]. inflation [OdSSP13]. Influence [BYZ^{+}23, LBW^{+}24, LW23, LWZC21, LWX23b, SY23, WDY23, Wu23b, YYZ^{+}17, SCSJ09]. Influences [ZH19, RX04]. Influencing [GWZ23b, JW23b, WCZW23]. influenza [LSB12]. informatics [BFH96]. Information [ASS23, AlS16, Ano23k, Ano23-27, Ano23d, ARGC19, Bak23, CLLZ23, CL23c, CCB^{+}20, DHW^{+}23, DFT^{+}23, HSH23, HLW^{+}18, HLY20, HS23a, HTZL23, HS23c, JW23b, bLwHcY18, LY20, LWZ24, Lin23b, LDZZ23, LWX23b, Met99a, Met99b, MBBDP17, PY17, RWY^{+}17, ROMFST23, TLW^{+}23, WHL^{+}18, WW23c, WLL^{+}23b, WWL^{+}23, WWW^{+}17, XZ23c, YDC^{+}24, Zha23b, ZCZ^{+}23, ZGY^{+}20, ZLL20, ZTLH18, PHH95]. Information-Balance-Aware [PY17]. Infrared [HY23, LLL^{+}23b, SZL^{+}23]. Infrastructure [CMTAC^{+}16, KKK^{+}18, MCZ23, STM^{+}23, SPPH15, WTY23, KeKR^{+}11, Rom02, SGM^{+}08]. Infrastructure-Based [SPPH15]. Infrequent [lWwFAH23]. Infringements [LPBSSEPL19]. Inheritance [Mor15]. Inhibit [ZMM^{+}23]. Inhibition [DCH^{+}23, LZZ19]. Initial [FZW^{+}18, YLL^{+}23]. Injury [CZFX23, FHC^{+}22, Li23d, SWY23, SWH23, WZS^{+}22, WZZ23, WZS^{+}23, WMZM23, XLB^{+}23, XWLL23, Xue23, Zha23f, ZRX^{+}24, ZZSQ23]. Inner [WG22]. Innovation [Ano23p, Ano23-46, Ano23-43, Ano23-45, CWD23, Hon23, LCH^{+}24, LZYJ20, Lou23, TH23, WmG21, YWH20]. Innovative [Lia23b, SW23]. input [GB94]. INR [ZYX^{+}21]. InSAR [ZWW^{+}23a]. Insect [CYZ^{+}20, CYJ^{+}20, ZYX^{+}21]. Inserted [CH23b, YLLY23]. Insights [WCZ^{+}16]. Inspired [BL18]. Institutions [LNH^{+}20]. Instruction [GR93]. instrument [KCO^{+}05]. Instruments [Xu23b]. Insufficiency [LTL^{+}23]. Insurance [HAA^{+}22, YCLK23]. Intangible [Lvp23].

Integer [QYW^{+}17]. Integrated [CL23c, DGH^{+}15, LZZ^{+}23b, RKR^{+}99, SL21, WHL^{+}18, ZS23]. Integrating [FMA16, HLY20, JL23b, Li22, lWwFAH23]. Integration [Ano23x, DDJ20, MLL23, SC23, ZYZL23, ZZS^{+}23, IRSD99]. Integrity [HHD^{+}16, MKS^{+}24]. Intel [CBCM93, CSU^{+}17, DGH^{+}15, SCD^{+}15, SRO^{+}15, TSP^{+}15]. Intelligence [BGZL23, BL23, BLZG23, CZFX23, Che23a, Chu21, Dai23, DMV^{+}19, Du23a, Du24, FLY^{+}23, JL23a, KF20, LWZ^{+}23b, LJL23, LL23b, LZ23b, LLY23c, NM16, RLPI19, SY23, SRTCL16, TY23, WCZ^{+}16, Wan23g, WWYR22, XY21, XY22, ZJ23, ZW23a, ZGL23, ZSZ^{+}23b, ZCL^{+}24b, Gil09]. Intelligence-Aided [Che23a]. Intelligence-Based [BGZL23, CZFX23, TY23, ZCL^{+}24b]. Intelligent [Ano23k, Ano23d, Ano23-30, Ano23-38, ABK^{+}23, BL23, CWMZ23, CYZ^{+}20, CYJ^{+}20, CT23, CZWY23, DC22, FX23, GBK^{+}96, GN22, HBR^{+}20, HSH23, Hon24, HT22, Hua21, HS23b, JLX^{+}20, KV23, LF23a, LWZ24, LLLZ23, LL23d, Liu23b, LZWL23, LWX23a, LTR^{+}23, LH23b, Ma23d, MAA^{+}21, PZL19, QK21, QTD^{+}20, RWN20, RBJ^{+}19, SMA^{+}23, SZC16, WX23, WLCC18, WCR^{+}21, Wu23a, XZ23a, XZ23c, XWL^{+}23, Xu23f, ZLZ^{+}22, ZZD23a, ZYX^{+}21, ZL23c, ZZX^{+}17, ZHHX23]. Intelligibility [SZ20]. Intel(R) [BGH99]. Intensive [AHMMÁR18, ÁRAHMM18, DRT^{+}14, Slo12]. Intention [Ano23-43, DFT^{+}23, HLZ^{+}21, QFX^{+}23]. Intentional [ZW23a]. Inter [ITF^{+}08]. Inter-operating [ITF^{+}08]. Interacting [WLL^{+}23b]. Interaction [Ano23w, Cao23a, Cao23b, FYH^{+}19, LWZ24, Liu23a, LWX23b, PZ24, RVJ^{+}23, REPRL^{+}19, SL23c, ZHP^{+}23]. Interactions [MCB^{+}20, MVIA20, PLY^{+}23]. Interactive [CH23a, FYH23, XYCC23, ZJ23, BBLR03, KVW^{+}07]. Interblock [FS24]. interchange [VRW^{+}03]. Interest [DPMB16, LWT18, MSZZ20, WDY23, XZ23b, ZW18, MGK^{+}13]. Interface [Ano23w, MUJN20, SLD^{+}16, Spo12, WOC99]. Interfaces [BKLS01]. Interference [AK18, SCSJ09]. Interferometry [WZY^{+}23]. intermediate [Tým99]. International [GM14, HH23a, PAR94, QM23, SO11, XY21, XY22, HB94]. Internet [HZFH23, JZL^{+}20, JL23b, JGDK23, JRS^{+}21, LJF20, LWW23a, Lou23, PSW23, PZL19, SW23, Wen23, XNLJ23, YC23, CL23a, Che22, CH23a, CHL^{+}23, DZL^{+}24, GN22, HNG^{+}23, HL23b, HT22, KKA^{+}22, LWZ24, Lia23b, LFQA20, Liu23f, LWX23a, LZK^{+}23, Lou23, OKHS23, Ray17, SL23b, WCZ^{+}16, WJCZ21, Yan23c, ZZD23a, ZLF^{+}22]. Interpolation [KLK^{+}20b]. Interpretation [WWZ17]. Interpreting [PDGQ05]. Interprocedural [HBCM94, HS03]. intervals [CGL08]. Intervention [CH23b, JWPW24, QQWJWYQZ23, Wan23f, ZWW^{+}23b]. Intracerebral [LX23d]. Intracranial [SKT^{+}24, Xu23f]. Intrauterine [HSS23, LLQ^{+}23, WZ23d, ZZH22]. Intravascular [Wan23f]. Intravenous [HSY^{+}23]. Intrepid [BEK^{+}12]. Intrinsics [Bra17]. Intrinsics-as-Template [Bra17]. Introduction [Ano01, Ano02, Ano04a, MH95, PT09a, Sne95, VC93, WSP94, Ste97]. Intrusion [BRPV24, DYYW24]. Intuitionistic [Xu23d]. Invariant [KLK^{+}20a, Liu23c]. Invasion [ZMM^{+}23]. Inventory [ZGS16]. inverse [Wan02]. inverses [GG05]. Investigating [HMCH07, OPP11]. Investigation [AUA^{+}23, NL23, Sah20]. Investment [tS23b, Xu23d, YLW21]. Investments [YP24]. Investor [HS23c]. Inviscid [AMU^{+}23]. Invocation [QYZ17]. Involvement [ZG23]. IoT [LWX23a, RLU^{+}20, SKB^{+}24, ZZS^{+}23, ASS23, Alt23, CLN^{+}24, CNW^{+}23, CGMRO^{+}19, FYC^{+}20, FAA^{+}20, GKK^{+}20, HBR^{+}20, KZZ^{+}20, Ma23d, QTD^{+}20, SKA^{+}20a, SKA^{+}20b, UAU^{+}23]. IoT-Assisted [Ma23d]. IOT-Based [SKB^{+}24, UAU^{+}23]. IoT-Enabled [Alt23]. IOT-Related [RLU^{+}20]. IP [HPZ17]. iPSC [CBCM93]. iPSC/860 [CBCM93]. Iranian [DD23a]. irradiation [GIKP95]. Irregular [KB96, LLP^{+}19, PCS99, LOHA01, LP99]. Irrigation [MVIA20, UAU^{+}23]. Ischemia [SW24]. Ischemic [DLP^{+}23, FZH23, HSY^{+}23, KSYC23, ZCLL23]. Isolated [LZK^{+}23]. Isomap [WFHC20]. Isomap-Based [WFHC20]. Isomorphic [HZD17]. Isorropia [BÇCD12]. Issue [JK10, PT09a, PT09b, SO11, VR13, BFGH14, BR11, GM14, HR12a, HR12b]. Issues [Ano99b, GMS^{+}23, McC96, SKA^{+}20a, SKA^{+}20b, ARvW03]. Itanium(R) [GHH^{+}02]. ITDPM [HZFH23]. Item [LLP23]. Iterative [BLL^{+}23, CLF^{+}23, CWS^{+}24, He24, LWL^{+}23c, LSS^{+}23, LHC^{+}23, LJ23, Ser18, WCF^{+}21, XZLZ24, ZTL^{+}23, BHRT12, DEvdV01, SG96]. iUML [PZC^{+}23]. iUML-B [PZC^{+}23].

J [Hil97]. Jack [Nag05b]. Jacobi [BH96]. JADE [CGK^{+}05]. Janet [BKLS01]. Jason [PCGLO14]. Java [Ano99b, BC99, BKLS01, DDS99, GRN99, GGMS99, MR02, MMG^{+}02, PMTL14, Tým99, WOC99]. Java-based [GRN99, WOC99]. Java^{T M} [BGH99]. Jiedu [Xu23e]. JIST [ARFS05]. JIT [BGH99]. JLAPACK [DDS99, PMTL12]. Job [LF23b, WCLL17, YZSZ17, FS01, KLN^{+}04]. job-execution [FS01]. Join [AMS17, CZ23]. Joint [GSL23, LCL^{+}23b, Qiu23, WYHS23, WZF^{+}20, XYY^{+}20, YWX23, Zha23f]. Joseph [Nag05a]. journal [Ano93]. Jr. [Hil97]. Judgment [QH23]. Just [YYF^{+}19, ZWZ23a, ARFS05]. Just-In-Time [YYF^{+}19, ARFS05]. Justice [Zha23h].

K-Modes [DYYW24]. K-Nearest [SHW23]. k-Reciprocal [TT23]. K2 [XWG23]. Kaiser [Bry96]. Kemari [KMR^{+}97]. Kenmi [Liu23d]. Keras [OPB^{+}20]. Kernel [FXYY17, HLC^{+}24, JFD20, eLA20, PWP21, SWW^{+}24]. Kernel-Based [HLC^{+}24]. kernels [FSV14]. Ketamine [WHL^{+}21]. Key [CCHJ23, Hem00, LLT23, ZGL23, ZZS^{+}23, ZD23b]. Keypoint [LXW24]. Keyword [WWZ17, Zha23b]. Kidneys [GL23]. KL [QL23]. Knee [SYW23, Yan23b, ZZ23]. Knowledge [Ano23-32, BL23, BQR^{+}20, DZL23, FLC^{+}20, HMSM08, LXAC21, LGZG23, LLY23c, MBBDP17, SQW^{+}21, SH19, TZF^{+}18, TLW^{+}23, XZ22, YG16, ZXS^{+}23, ZZX^{+}17, AHCZ12, Gil09]. Kokkos [ESP^{+}12]. Komzsik [Bor04]. Kong [LF23b]. Kong-Macao [LF23b]. Koniges [Nag04]. KSR [KBRS95]. KSR-1 [KBRS95]. Kubuqi [WG22]. Kurtosis [KSYC23]. Kurtosis-Based [KSYC23]. Kutta [LPD20]. Kyphoplasty [HZL^{+}23].

L [WZY^{+}23]. L-Band [WZY^{+}23]. Label [Bak23, ZQ22, ZLDX21]. Labels [CLG^{+}23]. laboratory [ABB^{+}02, WCW^{+}21]. LabVIEW [CM21]. laminar [FP95]. Lanczos [Bor04]. Land [Cui23, FCSL23, WKY23, ZHY22]. Landau [Hil97]. Landscape [Ano23f, Ano23j, YB23, ZJZ23]. Lane [LWCJ23]. Language [Ano23l, AAS^{+}23, Cer21, CHM^{+}97, CYZ^{+}20, CYJ^{+}20, CGA17, CP95, Hig93, HNMB19, JKA^{+}23, JBLB15, JHS^{+}20, LGZG23, NLPL^{+}20, REPRL^{+}19, VVDM18, Yan23h, ZLLS20, AJ94, AMM05, AJJF14, BBG^{+}93, CLM05, Che93a, Che93b, DSZ96, GGMS99, LG03, Met99a, Met99b, RS94, Szy07, VRW^{+}03, VF95, WC96]. Languages [CCTY15, HNSG23, Ray17, ZLWW23, HJ96, JK10, KGV97, Mar05, Zim07]. LAPACK [ARvW03, DDS99]. Laparoscopic [BGZL23, CZFX23, PTZL23, WLLY23, ZWLL23]. Laplace [CZY^{+}23]. Large [DZL^{+}20, HNMB19, HL23b, LLC^{+}20, SZAG15, SZXL16, TSCT11, TF23, XHSL17, XSZ18, YW24, BHRT12, BPK00, CGL08, JEM07, MJLM07, MSCS14, RS94, WGW08]. large-area [JEM07]. Large-Scale [HL23b, LLC^{+}20, SZAG15, XHSL17, YW24, MJLM07, WGW08]. Laryngeal [WZS^{+}22, WZS^{+}23, XWLL23]. Laser [LCK^{+}23, WJS^{+}23]. Latency [Kuf15, RS95]. Latent [BC23, HEVD19, LPBSSE17, SLDL17]. Lateral [LZZ19]. Latin [JK22]. Latin-Based [JK22]. Lattice [TLH17, XSZ18, ZH19, IB09]. Law [LPBSSEPL19, AHM23]. Layer [Ano23-39, KLL^{+}18, LZZ^{+}23b, Zhu23b]. Layout [LWZ24, VVDM18, DZKS13]. LCVP [LLS^{+}20]. Leaching [LFS^{+}17]. Leader [QHAN23]. leadership [WGW08]. Leaf [fTPsLZ23]. Leakage [SLS17]. Lean [YL23]. Learner [SK21]. Learning [ALA20, AKA^{+}23, Ano19, Ano23l, Ano23o, Ano23p, Ano23s, Ano23q, Ano23r, Ano23u, Ano23b, Ano23-35, Ano23-40, Ano23-48, ABR17, BQR^{+}20, BKU^{+}23, BKMS23, BTV^{+}20, CZB24, CIU20, qCtWfPbZ23, Che23c, CWSF23, CWS^{+}24, CZW^{+}23, Chu21, DDJ20, DM23, FX23, FYZ^{+}20, FZ23, GDNVSMM19, GLD^{+}24, GAH^{+}23, GLBB17, GNKuR23, GMM20, GGSL23, HLL23, HL23a, HAW23, He20, HCC23, HSS23, Hu23a, HH23a, HZFH23, HWJ18, HZHW23, HLH23, HW23a, HZ23, HZ24, Hui22, JGF23, JLJ24, JHS^{+}20, JL23a, KLL^{+}18, KGU^{+}20, KFU^{+}20, KAA^{+}20, KIU^{+}21, KGU^{+}21, LXAC21, LLQ^{+}23, Li23d, LZC^{+}24, LCL^{+}23a, LLC^{+}20, eLA20, LWW23b, LYWS23, LCL^{+}23b, LDW23, LLT23, Ma23d, MTK^{+}23, MUJN20, MLY23, MABU18, Ona19, OPB^{+}20, PTZL23, PLY^{+}23, Pan23b, PVCPdPSZVG17, PDROFRM16, PRW23, PPB^{+}19, QYL23, RBJ^{+}19, RQD^{+}20, RAdASGR^{+}19, SKB^{+}24, SDM^{+}19, She23, SLH23, SXP23, SL23a, SWW^{+}24, SW24, SK21, tS23b]. Learning [SW23, Sun23a, Tan23, Ton20, TYCL23, Wan22, WW23c, Wan23f, WZ23a, WWH^{+}23, WYHS23, WZML23, WM23a, WCZW23, Wan23c, Wan23e, WZF24, WDD^{+}23, WLZ^{+}21, WDY23, WZ23d, Wu23a, XWG23, XLB^{+}23, XHC^{+}23, tXChL^{+}24, XBG17, XHX23, XYY^{+}19, Xu20, YYY23, YGZ^{+}24, Yan23h, YWGZ20, YT22, YSZ^{+}23, YT23, YNK23, YWC^{+}18, YZ20, Zha22, ZHY22, Zha23d, Zha23f, ZGFL23, Zha23c, ZSZ^{+}23a, ZWZ20, ZWZ23a, ZY23, Zha23m, ZWF23, ZGY^{+}20, ZX17, Zho22, Zho23, ZW23b, ZZX^{+}17, ZLT^{+}21, ZWLL23, ZSZ23c, Zhu23a, ZZJ^{+}24, KG08, LB02, PVL^{+}04]. Learning-Based [AKA^{+}23, Ano23u, ABR17, BKU^{+}23, Che23c, DM23, GLD^{+}24, HLL23, HCC23, HW23a, KAA^{+}20, LLQ^{+}23, LCL^{+}23a, MLY23, QYL23, RAdASGR^{+}19, SLH23, TYCL23, Wan23f, WZ23a, WWH^{+}23, WYHS23, WLZ^{+}21, XLB^{+}23, YT22, ZWLL23, ZSZ23c]. Learning-Kernel [eLA20]. Lederman [Nag05b]. Left [YXL^{+}20]. Legacy [RRM^{+}15]. Legal [LL24]. Legend [FYY23]. Lending [GPX^{+}21, PWP21]. Length [NYML17, SX23, XLF^{+}24]. Lens [CSZ^{+}18]. LEO [CCBPGA15]. Leone [AAH^{+}23]. Lesion [XZ23b, ZH23]. Lesions [DBWY24, ZYY^{+}23, ZLZ^{+}21]. Letter [Pre99]. Leukemia [GAH^{+}23, RJA^{+}23]. Level [JYLL23, LJSJ23, OCC^{+}16, SLP18, TPSM17, Vol97, XXM^{+}23, YWGZ20, ZCLL23, ZZH22, dBMT20, COB^{+}00, LBvBW12, RRV09, AK18]. Level-Dependent [ZCLL23]. Level-Wise [TPSM17]. Levels [SWY23]. Levenshtein [RAZ^{+}23]. Leveraging [DRQ^{+}18]. LFRic [ARAM19]. Liberalization [MSZZ20]. libraries [ARvW03, PMTL12, PMTL14, SG96]. Library [Bra17, Cao23a, DW23, VVDM18, WL23b, XNLJ23, BŻvA^{+}01, ESP^{+}12, GHST12, GL04, GSM03, Lan03, VGC09]. Library-Based [XNLJ23]. License [RAdASGR^{+}19]. Lidar [CHZY17, WLCC18]. Lidar-Based [WLCC18]. Life [YCLK23, KFFZ05]. Lifetime [KPHY18]. Lifting [Sch03b]. Ligature [ZLWW23]. Light [CSZ^{+}19, MWZ^{+}23, QZY^{+}20, WFZ^{+}23]. LightGBM [HY23, ZWG^{+}20]. Lighting [XYCC23]. Lightweight [LZLJ23, Ott93]. like [KGV97, PMTL14, ZLWW23]. Limit [LWT18, ZGT^{+}23]. Linda [Lar93, Mat94]. Line [Cer21, CSZ^{+}19, CWS^{+}24, LFS^{+}20, MSA^{+}20, WLCC18, BBF^{+}04, RLL^{+}02, SCB02]. Linear [ACIK97, QYW^{+}17, WB16, YTD15, Zha23e, ZCT^{+}20, BHRT12, HKL^{+}12, Kok07, MSSG11, RR07, SG96]. Linearization [WB16]. Linearly [YW24]. Lines [Cer21, FH23, YXW^{+}22]. Lingual [HNMB19]. Linguistic [DZL^{+}20]. Link [CSM15, GMCT15, SSN15]. Linkage [Che22]. linked [HS03]. Linpack [KGBB09]. Linux [Nag05a, HHD^{+}16, MR18]. Lip [DHW^{+}23]. LitCovid [DD23a]. Literature [ACBV20, JKA^{+}23, WSM^{+}21, XHL^{+}23, XTZ20]. Lithotripsy [LCK^{+}23]. Live [JFPL16, YZ20, ZZW22]. Liver [Ano23v, CWMZ23, HZHW23, LYL^{+}24, ZYZ23, ZCZ^{+}23, ZCL^{+}24b]. Living [SDM^{+}19]. LMS [HCGGSS17]. LncRNA [ZMM^{+}23]. Load [AFU18, CKKK15, DZL^{+}24, NLMT23, WCR^{+}21, Xie20, ZGFL23, Din99, IRSD99, LTB02, PVL^{+}04]. Lobar [WZL23]. Local [GSL23, LGC^{+}16, LLL^{+}23b, Ma23b, XWLL23, YYF^{+}19, FSV14, JIC^{+}14, ZGW08]. Locality [MJB15, TLC15, HJYC10, RR04, Sin99, Zim07]. Locality-Aware [MJB15, HJYC10, Zim07]. Locality-Centric [TLC15]. Localization [FZ23, HLC^{+}24, RAZ^{+}23, YYL17, ZDMY23]. Location [KZZ^{+}20, KZK^{+}23, LZC^{+}24, LLS^{+}20, PLW^{+}20, Su23, SCZ18, ZXHQ17, Zha20a, ZFFP24]. Location-Based [KZZ^{+}20, KZK^{+}23]. Location-Constrained [LLS^{+}20]. Location-Sharing [PLW^{+}20]. Locations [Kuf15]. Log [WYY21, XHC^{+}23]. Log-Based [WYY21]. Logging [ZLL^{+}15]. Logic [AN19, Han23, NO18, XNL^{+}22]. Logical [CSWW23, Jua23b]. Logistic [SKHZ24, SWH23]. Logistics [Ano23-30, BYZ^{+}23, Liu23g, Su23, Wan23g, fXhCL17, ZFFP24, Zha23l]. Logs [SJD20]. Long [LYWS23, MCF^{+}23, RKA^{+}24, SF23, KTP05]. Long-Distance [RKA^{+}24]. lookup [WSB11]. Loop [CM21, NAA^{+}03]. loops [LOHA01]. LoRa [CHL^{+}23]. Loss [KPHY18, LQLX18, NL23, PRW23, YDC^{+}24, ZZWL24, ZLL20]. Louis [Bor04]. Louvain [LHZ^{+}23]. Low [BA18, CWD23, FHC^{+}22, FMG^{+}23, GDA^{+}23, Hon24, HEVD19, HWJ18, KGIL19, LZL^{+}23, Li23b, NHHW23, RS95, WCF^{+}21, YTD15, YXZ23, YPC^{+}23, ZWS19, ZZZ^{+}23, ZFFP24, ZL23c, HB12, SZ04, TPKP13]. Low-Carbon [CWD23, ZFFP24]. low-communication [TPKP13]. Low-Complexity [KGIL19, ZWS19]. Low-Cost [BA18, HWJ18, SZ04]. Low-Dose [FMG^{+}23, Hon24, WCF^{+}21]. Low-Energy [YXZ23]. Low-Incidence [NHHW23]. Low-Rank [FHC^{+}22, HEVD19, LZL^{+}23, YTD15, ZZZ^{+}23]. low-variance [HB12]. LPARX [KB96]. LS [SQ21]. LSA [LPBSSEPL19]. LSA-Based [LPBSSEPL19]. LSTM [BKMS23, GWZ23a, JZH^{+}23]. LTS [PDRC18]. LU [CDO^{+}96, DEŁ01]. Lumbar [WCW^{+}21, WYHS23, ZYY^{+}23]. Lung [AMN^{+}12, Ano23u, CZD^{+}20, LWL^{+}23c, LSS^{+}23, Li23d, LBW^{+}24, SWY23, TY23, WZL23]. Luxury [Ano23s]. Lymphoblast [RJA^{+}23]. Lymphocytic [RJA^{+}23].

M [AE19]. M-QAM [AE19]. M5P [PLAAH23]. Macao [LF23b, PZH23]. Machine [AKA^{+}23, Ano23s, Ano23q, Ano23w, ABR17, AB18, AGIS94, BKU^{+}23, BTV^{+}20, CIU20, CGC16, CLTX16, CZW^{+}23, DBWY24, DM23, FZH23, GDNVSMM19, GAH^{+}23, GLBB17, GNKuR23, Hav00b, HSS23, Hui23, JK22, JLJ24, JL23a, JWPW24, KLL^{+}18, KGU^{+}20, KAA^{+}20, KIU^{+}21, KGU^{+}21, Li23d, LCK^{+}23, LLC^{+}20, LLS^{+}20, MTK^{+}23, MJ95, MqWp23, MABU18, PWP21, Pan23a, Pan23b, PZC^{+}23, PDROFRM16, PPB^{+}19, RBJ^{+}19, SGK^{+}23, SK21, SCAP16, Sun23a, WW23c, WZS^{+}23, WZ23d, XBG17, ZCLL23, ZY23, Zha23m, ZGY^{+}20, ZHL16, Zho20, ZZW22, Zho23, ZLT^{+}21, ZSZ23c, Zhu23a, GBJ94]. Machine-Based [DBWY24, JK22, LCK^{+}23]. Machine-Part [SCAP16]. Machinery [WLQ23]. Machines [BMN^{+}97, GL23, SJ15, SO15, YWC^{+}18, BCC^{+}93, BCC^{+}00, HMCH07]. MAGMA [DGH^{+}15]. Magnetic [DBWY24, DLP^{+}23, FZH23, FSLG17, FHC^{+}22, HLL23, HTZL23, HSY^{+}23, JFY23, LWZ^{+}23b, Liu23c, LTL^{+}23, LTR^{+}23, LYL^{+}24, LYZ^{+}23, OWL^{+}23, PTZL23, SKT^{+}24, WYHS23, WCZW23, WZZ23, Yan23b, YT22, YHZG23, ZYZ23, ZZZ^{+}23, ZCLL23, ZY23, ZCL^{+}24b, ZLZ^{+}21, PKE^{+}10]. Mainland [SPH20]. Maintainability [ACBV20, SCR20]. Maintenance [Hon24]. make [Sch94]. Maker [ZLL20]. Makerspace [WZ22]. Makespan [CBE23]. Making [AUA^{+}20, Che22, DZL^{+}20, Fen23b, Gog23, LRX^{+}17, Lv23, QHAN23, SNH24, XNLJ23]. Malformation [Zho22]. Malicious [GKK^{+}20]. Malignant [LX23d, LYZ^{+}23, WZC^{+}23, WSL23, Yin23]. Malmquist [Ano22b, YCLK23]. Malware [JMGH21, WTY23]. Management [AMVGC^{+}23, ÁRAHMM18, Ano23k, BDAH20, BBI23, CHTZ23, CMTAC^{+}16, CFY23, Du23b, FLS17, HLY20, HL23c, HS23a, HCQ17, JUA^{+}23a, JZY23, LL23b, Lin23a, LWW23a, Lvp23, MSZZ20, NGMAA23, ROMFST23, SMB17, SXP23, SL23b, SNH24, THS^{+}15, TKKZ23, UAK18, UAU^{+}23, WW23c, Wan23c, WWL^{+}23, WLZ^{+}21, YG16, YL23, YC23, YP24, ZWL^{+}19, ZHZS19, ZZD23a, ZZW17, ZTLH18, ZLF^{+}22, BCK07, CJS^{+}02, IB09, KeKR^{+}11, KABW11, KLN^{+}04, PRM^{+}14, vDKH01]. Managing [ANEA08, AKA16, CDD^{+}05, PPS12a, PPS^{+}12b]. Manchu [ZLWW23]. MANETs [RKA^{+}24]. Maneuvering [XC23]. Manifold [GMM20]. Manipulation [Liu23d]. Manipulator [Luo23]. Manual [FH23]. Manufacturing [LLY23c, XZ23a]. Many [CKKK15, GWH^{+}23, TLC15, PTS^{+}13, DGH^{+}15]. Many-Core [CKKK15, TLC15, PTS^{+}13]. Many-Integrated-Core [DGH^{+}15]. Manycore [CCTY15, ESP^{+}12, MJB15]. Map [Ano23-32, TZL^{+}24, ZWXL21]. MAPANet [TT23]. Maple [PPD05]. Mapping [AIA^{+}21, CZY^{+}23, EMTCA^{+}19, HCQ17, QFX^{+}23, SMA^{+}23, XZ22, DSS^{+}05, Den96, DRR12, PMCF94, SZ04]. MapReduce [AMS17, CZ23, MPA16, ZWXL21]. MapReduce-Based [CZ23, MPA16]. maps [KTP05]. Marc [Nag05b]. Margin [TF23]. Marine [ZCW^{+}19]. Maritime [LPBSSEPL19]. Mark [Ano07b]. Marker [SSY^{+}23]. Market [CSM^{+}18, HS23c, JAED23, SPH20, Xia23a, YCLK23]. Marketing [Ano23s, Bia23, FYH23]. Markov [BH02, Guo23a, LB02, VRM02]. markup [VRW^{+}03]. Marrow [DCH^{+}23]. Martial [LLY24]. Mary [Bry96]. Mask [HCC23, TMHM23]. Mass [WZC^{+}23]. Massive [HYS^{+}19, LKS18, JK10]. Massively [GA96, LLY^{+}20, Nag04, LH93, OPE^{+}95, SS00, Sch94, WMR^{+}94, WB95, WSP94].

Matches [FWX23]. Matching [CLL^{+}18, CFY23, DCF17, LL23b, LZZ19, TLC15, YWC^{+}18, ZG18, ZYY^{+}23]. MatchMaking [ITF^{+}08]. Material [ZXS^{+}23, ZWJN23]. Materials [HMSW92, SZX^{+}20, WQT^{+}20, NKV^{+}02]. Mathematical [Fen23a, LGC^{+}16, LRX^{+}17]. Mathematics [Lv23, Yan23f]. Matheuristic [LGC^{+}16]. MATLAB [PMTL14, CB99, Kok07, MLAS21]. Matlab-Based [MLAS21]. MATLAB-like [PMTL14]. Matrices [CSM15]. Matrix [DEŁ01, FHC^{+}22, HEVD19, LZL^{+}23, TKZ^{+}24, VMRM16, YTD15, ZZZ^{+}23, BBDN11, CFR14, KHSJ95, Lin04, VGC09, WCW^{+}21]. Matter [KSYC23, SAN^{+}19]. Maturity [HLS^{+}24]. Maximizing [HY15]. Maximum [WPDZ19, ZGT^{+}23]. May [Met99a, Met99b]. MC [CTZC24]. MC2 [DTV00]. MCAF [ZHW^{+}21]. MCDM [HKQ^{+}23]. McrEngine [IMB^{+}13]. MDE [HCGGSS17]. Mean [SAZ^{+}24, WZL23]. Means [Ano23-31, CZP^{+}24, Hon23, HHZ^{+}23, JFD20, LZSC23, LX23d, LPBSSEPL19, RLC04]. Measure [CSZ^{+}18, JL23a]. Measurement [ALL^{+}16, Cha23b, FSLG17, LZ23c, SKA^{+}20a, SKA^{+}20b, SX23, SFZ^{+}20, WWH^{+}23, WJS^{+}23, Xu23f, ZY24, WGW08]. Measures [KGIL19]. Measuring [AUA^{+}20, ACMT19, DPMB16, TMHM23, fTPsLZ23, YCLK23, STD^{+}14]. Mechanical [LWL^{+}23a]. mechanics [DFP96, GW11]. Mechanism [Ano23b, Ano23-38, CL23a, DyLJ23, Ge23, HCGGSS17, KAA^{+}20, Li23e, LX23c, LQW23, Ma23a, PRW23, QFYY20, SLC^{+}23, THS^{+}15, TZDZ20, YWH20, Yi23, Yua20, ZW23a, Zha23b, Zhu23b, ZCCS23]. Mechanisms [QWP23, Ott93]. Meconium [JLJ24]. Media [Bia23, LW23, SC23, Zha23i]. Mediated [CH23b, WZC^{+}23]. Mediation [LF23b, MBBDP17]. Medical [AHM23, DRMG^{+}18, DQQ17, FTS^{+}20, GAH^{+}20, GLW^{+}23, HGMM23, HQL^{+}23, LZLJ23, LMQ^{+}24, TTL^{+}23, TXD^{+}20, WSM^{+}21, WL23c, XZZY20, YPC^{+}23, ZHL^{+}23, ZH23]. Medicine [XZLZ24]. Medium [Fen23a, Li23e, Wan23c, JK10]. Medium-Sized [Fen23a, Li23e, Wan23c]. Medoids [HKVVNT18]. Mega [Zha22]. Meituan [Ano23q]. Mellitus [LWZC21, ZLHZ23]. Melodies [ZL23b]. Member [JYLL23]. Memorial [Per08]. Memory [BMN^{+}97, CHS^{+}99, GCB^{+}23, LYWS23, MCF^{+}23, PGP19, SO15, SZXL16, SF23, TKS02, WCZW23, YZX17, ZLZ^{+}23, Bai95, BPC^{+}09, DBVF01, FSV14, GL04, GBH14, HMCH07, JKR92, JIC^{+}14, KMR^{+}97, KHSJ95, LR05, NAA^{+}03, OB96, Per08, PBK01, RS95, RMX05, SHHI01, SZ09, Shu94, ZS99]. Memory-Based [SZXL16]. Memristor [LD23]. Memristor-Based [LD23]. MEMS [LLY24]. Menhir [CB99]. Meniscus [FHC^{+}22, XLB^{+}23]. Mental [Li23c, NHHW23, PP23, SWZL24, hKJ^{+}20, ADS95]. Merchandise [YT23]. Merchants [Ano23q]. Merit [CC23]. Mesenchymal [DCH^{+}23, ZZD^{+}23b]. Mesh [KK09, MLYH17, NYML17, WZW19, BAN02, CL04, CW93, MSLS13, KK09]. mesh-based [CL04]. Meshes [MYH19, SM17, SO15]. message [PBK01]. message-passing [PBK01]. messages [RS95]. MESSIAHS [CS94]. Messina [Mic97]. Meta [ACA19, LCL^{+}23b]. Meta-Based [ACA19]. Meta-Learning [LCL^{+}23b]. Metabolic [LYL^{+}24]. Metadata [DWZ^{+}20, MPP^{+}04]. Metaheuristic [SQ21, Tun23]. Metals [YY23c]. metaprogramming [BMY13]. Metastasis [Ano23u, HLL23, XZ23b, YHZG23, ZCL^{+}24b]. Method [ÁRCPS15, Ano23d, Ano23-42, Ano23-38, Bai23, BBI23, Bia23, Bor04, CIU20, CSZ^{+}19, Che23a, CCM^{+}23, DC22, DZ18, DZL^{+}20, DZL^{+}24, Du23b, Du24, DD23b, FX23, FHS^{+}23, FXYY17, GMS^{+}23, GLH23, GL23, HLKY19, HLW^{+}18, HKVVNT18, HY15, HZD17, HS20, Hu23b, HX23, HHD^{+}16, HW23b, JW23a, JZY23, KLK^{+}20a, LWL^{+}23a, LXL^{+}23, LY23a, LGZG23, LDZQ24, LZ23b, LFY^{+}18, LYL^{+}23, LHZ^{+}23, LDW23, LH23b, LZ23d, LJG^{+}23, MqWp23, MM16, NM16, PLY^{+}23, PBY^{+}20, PZ24, QYZ17, QK21, QM23, Qin23, QH23, SBK^{+}24, SO18, SD23, SX23, SZWZ23, SL23a, SCR20, Sun23a, TZZ23, Tan23, TLZ^{+}23, TC96, fTPsLZ23, WX23, WB16, WZG^{+}17, WXL19, WZW19, WNL^{+}20, WYY21, Wan23b, WL23a, WLQ23, WWW^{+}17, WLZ^{+}21, WG22, WC23b, XZ23a, XWG23, XJH23, Xie20, XWL19, XLY18, XSZ18, Xu23d, Yan23a, YDC^{+}16, YJ23, Yan23g, Yan23f, Yan23d, Yan23c, YYW^{+}23, Yu22, YCLK23, ZCL^{+}24a, ZLS^{+}17, ZLCW20, Zha20b]. Method [ZGFL23, Zha23j, Zha23g, ZJZ23, Zha23b, ZC20, Zha23k, ZWZ23a, ZL23c, ZHP^{+}23, ZY24, ZLL20, ZLDX21, ZSZ23c, ZGT^{+}23, DN04, GG05, PBK01, TOF^{+}14, Ano07b]. Methodological [ACMT19, DDJ20]. Methodologies [CPD21, Hav00a]. methodology [HCJ08]. Methods [AAH^{+}23, BRSP18, CKKK15, GMCT15, Gog23, Li23a, LBW^{+}24, LWW23b, Mor94, SMD20, Vol97, Zha23h, BCC^{+}93, RR07, STD^{+}14]. Metric [LWYS16, XML^{+}16]. Metrics [AFB23, ACBV20, GMCT15, SNH24, LLW^{+}14, RX04]. Metropolitan [ZFFP24]. Mexican [DRMG^{+}18]. MHDFS [SZXL16]. MIC [CCBPGA15, CSU^{+}17, MYH19, ZWG^{+}20]. micro [ADLN08]. micro- [ADLN08]. Microanimation [SC23]. Microblog [FYH^{+}19, QFYY20]. Microclass [Yan23a]. Microcomputer [Ano22a]. MicroGrid [SLJ^{+}00]. Microimmune [YZ16]. MicroRNA [SWW^{+}24]. Microservice [MLHC20]. Microservice-Based [MLHC20]. Microtia [JP23]. Middleware [OKHS23, COB^{+}00, VCT05]. Midline [LZLZ19]. Migrant [Li23c]. migratable [BŻvA^{+}01]. Migrating [SCAP16]. Migration [Che23c, CCBPGA15, GWZ16, LZF16, ZMM^{+}23, ZZW22, HZ99, KLN^{+}04]. Milk [ZWW^{+}23b]. millennium [Met99a, Met99b]. MIMD [PBK01]. Mind [Ano23-32, FML^{+}14]. Mine [XCW18]. Mineral [MLAS21]. Minimal [LDZQ24]. Minimization [ZMLZ18]. Minimize [ABK^{+}23]. Minimizing [KPHY18]. Minimum [CBRRC^{+}16, LWL^{+}23a, MKH^{+}19, WZC^{+}23]. Mining [ASS23, AlS16, Ano23q, ÇÜ19, DD23a, Don23, DW23, Fen23b, HB12, HL23c, JNK24, KS15, Li23c, LWQ^{+}23, LL24, LLY^{+}23b, RuHMS24, SLD^{+}16, Wan22, WZY^{+}23, lWwFAH23, WG21, WG24, YL23, YT23, ZYLZ15, ZLLZ20, ZXS^{+}23, ZW23b, AMN^{+}12, AHCZ12, PSU08, ZRP08]. Minority [Jua23c]. miR22 [ZMM^{+}23]. miR22-3p-TXNIP [ZMM^{+}23]. Mirroring [SSM^{+}02]. Missing [WCZ^{+}16, WCR^{+}21]. mission [Nor07]. mission-enabling [Nor07]. Mitigating [AK18, OdSSP13]. Mix [CWD23]. Mixed [GPP^{+}17, Hui22, LG03, QYW^{+}17, RR07, SB01]. Mixed-Criticality [GPP^{+}17]. Mixed-Integer [QYW^{+}17]. Mixed-language [LG03]. Mixture [WZ23c]. MM5 [Mic00]. MMX^{T M} [BGH99]. Mobile [Alt23, Ano23o, ATZ^{+}15, CYJ^{+}17, JL23b, LG18, LD23, MR02, PLW^{+}20, RLU^{+}20, SAE^{+}23, WY23, WC23b, XWG23, YC23, ZHW^{+}21, CKS^{+}09]. MobileNet [Yan23d]. mobility [BDM^{+}04]. Modal [WFHC20, Zha23h, ZCT^{+}20]. Mode [DZ18, DW23, HHCC23, JL23b, LWZC21, LMQ^{+}24, ML23b, SW23, Tia23, WH23, WDY23, ZYZL23, SB01]. Model [AKA^{+}23, ACA19, ASZF23, Ano22b, Ano23o, Ano23s, Ano23q, Ano23t, Ano23-46, Ano23-38, Ano23-39, Ano23-44, Ano23-40, Ano23-48, ABR17, ARAM19, AA23, Bak23, BD16, BYZ^{+}23, CL23a, CZB24, CTS23, Cha23a, CHZY17, CFL^{+}23, CL23c, CZWY23, CFY23, DMV^{+}19, DDMS15, DS23, FA24, FCZ^{+}17, FC22, Fen23a, FGSCR^{+}17, FCSL23, GLH23, GBH20, GXY^{+}20, Guo23a, HPZ16, HCGGSS17, Hon23, HLS^{+}24, Hu23a, HZFH23, HSL^{+}23, HZ23, Hui22, JAED23, JWPW24, JZY23, JRS^{+}21, KS15, KLuH^{+}20, KAA^{+}23, KPS16, KF20, KZZ^{+}21, KS19, LQLX18, LWL^{+}23c, LSS^{+}23, LXL^{+}23, LY23a, LMZP23, Li23a, LL23c, LL23b, LGZG23, LLY24, Li24, LCH^{+}24, LZYJ20, LLGL20, Lin23a, LZ23b, Liu20, LZZ^{+}23b, LCL^{+}23b, Liu23d, LZLJ23, LQW23, LZ23d, LL24, LWX23b, LWCJ23, MCB^{+}20, MTK^{+}23, MAC^{+}20, MKH^{+}19, PWP21, Pan23a, PLJ^{+}23, QZY^{+}20, RKA^{+}24, RWN20, RAA23, SJD20, SLS17, SQJG23, SZC16, tS23b, SH19, Sun23c, TZZ23]. Model [TFZ^{+}20, TZC23, TXD^{+}20, WHL^{+}18, Wan21, WZ22, Wan22, WKY23, Wan23g, Wan23c, WLL23a, WM23b, WWL^{+}23, WDD^{+}23, WZ23c, WG21, WmG21, WG24, XY23a, XY21, XY22, XZ23a, XHX23, XTCC15, fXhCL17, Xu20, Xu23a, XZ23c, Xu23c, XWL^{+}23, XZ23b, Xu23b, YZQ^{+}23, Yan23g, Yan23f, Yan23e, YLW21, YZSZ17, YNK23, Yua20, ZXHQ17, ZLF18, ZCW^{+}19, ZHZS19, Zha20a, ZCZ21, Zha22, ZZWZ23, ZHL^{+}23, ZH23, Zha23i, ZLWW23, ZLSW23, ZL23a, ZC20, ZYX^{+}21, ZYL23, ZQL23, ZZS^{+}23, ZWZ^{+}23b, BHK^{+}13, BH02, BDV03, CIN^{+}96, DTV00, Hem00, HMCH07, KBRS95, Mar05, PMM94, Ros00, SZ09, SS00, SDS00, SBM^{+}10, Wan02, War96]. Model-Based [BD16, SH19, ZZWZ23]. Model-Centered [DDMS15]. Model-Driven [DMV^{+}19, HCGGSS17]. Modelica [CM21, Cer21]. Modelica-LabVIEW [CM21]. Modeling [Che23a, DD23a, DDMS14, DGL^{+}17, JKA^{+}23, LLP^{+}19, RAA23, REPRL^{+}19, SW16, WZY^{+}23, ZZD23a, ZLX^{+}23, ADLN08, BDM^{+}04, GRR^{+}03, LDV07, SLJ^{+}00].

Modelling

[ACCTA^{+}15, Cer21, CSU^{+}17, LPHD04]. Models [CCTY15, FLS17, GWZ23a, GPX^{+}21, GRB23, HLY20, Li23c, PRW23, QYW^{+}17, RWY^{+}17, TLZ^{+}23, WS15, WZW19, WSM^{+}21, lWwFAH23, YYF^{+}19, YWH20, ZZW17, LB02, MSSG11, VRM02]. Modern [Ano23-27, HMR^{+}15, LNK^{+}15, ZJ23, AJJF14, NDSG07]. Modes [DYYW24, JLX^{+}20, YZL23]. Modification [SY20]. Modified [AMY^{+}21, CCG17, CLHH20, UJS^{+}20, WQLR17, ZY23]. Modulation [AE19, Hua23, LX23a, VMRM16]. Module [WYC20b, BCS01, Sch03a]. modules [HB12]. Molecular [TYCL23, BCS01, BPK00, DFP96, GIKP95]. Mongolia [WG22]. Monitoring [BBŁ08, BDAH20, KKA^{+}22, Kuf15, Lia23a, LFY^{+}18, LYZY24, RAA23, TMG^{+}19, WZY^{+}23, XCW18, XSS^{+}23, YLLY23, ZWW^{+}23a, ZCZ^{+}23, BBF^{+}04, KCO^{+}05, TF04]. Monkey [Tun23]. Monocular [LWYS16, XLF^{+}24]. Monosialotetrahexosylganglioside [WHL^{+}21]. Monte [WXL19, WCZW23]. MOOC [Ano23-37, Yan23a]. Mooring [LFS^{+}20]. MOPEX [JEM07]. Moral [Liu22, YLW21, YP24]. Morphological [TTL^{+}23]. Mosaic [FMG^{+}23]. mosaics [JEM07]. Motion [FSLG17, LZLZ19, LDW23, WW22, gXySR^{+}17, ZWS19]. Motivated [TTL^{+}23, LLW^{+}14]. Motivators [JUA^{+}23a]. Motor [Ano23n, HH23b, Liu23e]. Move [CZB24]. Movement [LH23b, TS23a]. Movements [ZWF23]. Movie [KGU^{+}20, KGU^{+}21]. Moving [Luo23, NYML17, ZCT^{+}20]. MP [CKS^{+}09]. MPDATA [SRO^{+}15]. MPI [Nag05b, GBH14, HZ99, HPS^{+}13, JR10, KB01, LYTL20, MAC^{+}20, RST02, SM17, SB01]. MPI-3 [GBH14]. MPI/OpenMP [JR10, SB01]. MQHD [Xie20]. MRI [LYZ^{+}23, ZY23, Ano23v, GWZ23b, JHL^{+}23, LZL^{+}23, SWZL24, XLB^{+}23, Xue23, Yin23, ZCL^{+}24a, ZYY^{+}23, Zha23f, ZRX^{+}24, ZC23]. Mucosolvan [JLJ24]. Mulching [LWL^{+}23a]. Multi [CKKK15, GGMS99, LYTL20, LPBSSE17, LZWJ23, LL23c, PT09a, PT09b, XTCC15, XSZ18, CIN^{+}96, DF14, JR10, PVL^{+}04, Zhe10]. Multi- [CKKK15]. multi-agent [PVL^{+}04]. Multi-core [PT09a, PT09b, JR10, Zhe10]. Multi-CPU [LPBSSE17]. Multi-GPU [LYTL20, XTCC15]. Multi-GPUs [XSZ18]. Multi-language [GGMS99]. multi-model [CIN^{+}96]. multi-objective [DF14]. Multi-Scale [LZWJ23]. Multi-Target [LL23c]. Multiagent [BL18, KF20, YWH20, ZZJ^{+}24, PCGLO14]. Multibillion [SO15]. MultiCache [NsP16]. Multichatbot [MAS^{+}23]. Multiclass [AMH23]. Multicloud [HKH^{+}17, WZW^{+}20]. multicomputers [JKR92]. Multicore [BD16, DGL^{+}17, KGMN15, LNK^{+}15, NJ08, SJ15, YTD15, BB09, HPD09, LKDB10, PCGLO14, SZ09, SJR14]. Multicriteria [AUA^{+}20, Gog23, HNG^{+}23, SBK^{+}24, Son23b, SNH24, XNLJ23]. Multicue [LZLZ19]. Multidepot [WQT^{+}20]. Multidimensional [Ano23w, Cao23b, LLP23, NYML17, SD23, Wan21, ZW23c, ZKA23, ESP^{+}12, FO96]. Multidisciplinary [CHM^{+}97]. multidisk [GKL^{+}96]. Multifactor [Xu23c]. Multifactor-Weighted [Xu23c]. Multifeature [HBSZ20, SLC^{+}23, YHZ^{+}23, ZWZ20]. Multifocus [WG23]. Multifunctional [CLW^{+}23]. Multigenomic [LLC^{+}23]. Multigrained [WDZY20]. multigrid [AGIS94, GHST12]. Multihead [QFYY20]. Multikernel [GL23, XBG17]. Multilabel [WW23b, ZLDX21]. Multilayer [AMB23, PCYZ18, Yan22a]. Multilayered [NsP16]. Multilevel [FZH23, Han23, MFB^{+}15, ZCLL23, JJY^{+}03].

Multimean [HI18]. Multimedia [Ano23x, LL24, Tia23]. Multimodal [Ano23-34, DFT^{+}23, HWY^{+}23, LLZW23, WCZW23, WG23, Yu22, Zha23b, Zha23l, ZYL23]. Multimode [Liu23a, Qiu23, SL23c, Zha20b]. MULTIMOORA [DZ18]. Multiobjective [BNH^{+}20, MKH^{+}19, SBK^{+}24, TH18, TPSM17, WM23b, WZF^{+}20, Zha23g, dBMT20]. Multioperator [HWSL22, LX23b]. Multiparameter [Ano23i]. Multipath [ZCSL17]. Multiperson [WW22]. Multiphase [AFU18]. multiphysics [PPS12a, PPS^{+}12b]. Multiple [AMS17, BNH^{+}20, CLL17, CLA^{+}15, HKQ^{+}23, JLX^{+}20, Kuf15, LLGL20, LZLZ19, LJG^{+}23, MYH19, MWQ16, Mor15, QWP23, SWW^{+}24, SQ21, WS15, XSG23, XSZ18, YWX23, ZMLZ18, Zha23e, ZY24, MGK^{+}13]. Multiple-Relaxation-Time [XSZ18]. Multiplexing [WNL^{+}20]. multiplication [KHSJ95]. multiplications [BBDN11, GA96]. Multiplicative [WB16]. Multipolicy [YNK23]. Multiprocessing [RuHMS24]. multiprocessor [CKS^{+}09, GKL^{+}96]. multiprocessors [LR05, RS95, BDV03, AGIS94]. Multiprogramming [BHP^{+}03]. Multiproxies [ZZWL24]. Multipurpose [XZZY20]. Multiresolution [Che23b, TZH20, TZL^{+}24]. Multiresource [CLL^{+}18]. Multiscale [LYWS23, ADLN08]. Multisensor [JZ22]. Multislice [JP23, LJ23, LYZ^{+}23]. Multisource [qCtWfPbZ23, HSH23, HS23c]. Multistage [CLXZ23, YLLL20, ZHZS19]. Multisupercomputer [LZF16]. Multiswarm [BNH^{+}20]. Multitarget [YXW^{+}22]. Multitask [LWZ^{+}23a, YSZ^{+}23]. Multitasking [ATZ^{+}15]. Multithreaded [SCD^{+}15, HMC97]. multithreading [Sin99]. Multivariate [DDB^{+}18, ZWZ^{+}23b]. Multiview [BQR^{+}20, CZCL23b, YXL^{+}20, YWGZ20]. Multiway [AMS17]. Murals [LLW23b]. Muscle [Xue23]. Museum [XYCC23]. Mushroom [WDZY20]. Music [Ano23a, Ano23y, Dai23, HZ23, HS23b, She23, Wan23a, XY23a, Xu23b, Yu22, ZL23b, Zha23c, ZGL23]. Musical [HZ23, Xu23b]. MUST [HPS^{+}13]. Mutant [ZXDY19]. Mutants [LYL^{+}23]. Mutation [CZKY24, CZCL23a, WWQ23]. Mutual [YDC^{+}24, Szy07]. Mycoplasma [CWLW23, WJZC23]. Myocardial [LWZ^{+}23c, SW24]. Myofascial [JFY23].

NaaS [XST^{+}16]. Naive [CCM^{+}23, YHZ^{+}23, XY23a]. Name [CYLD23]. Named [HWJ18, WSM^{+}21]. Names [LDZQ24]. naming [Sch94]. nanofluidic [ADLN08]. NAS [SBM^{+}10]. Nasal [TY23]. Nasopharyngeal [FLY^{+}23]. Native [BKLS01, EOS17]. Natural [DC22, NLPL^{+}20, SLS17]. Navier [AMU^{+}23]. Navigation [AMT20, DCF17, MJLM07]. Navy [Ros00]. Nearest [CZD^{+}20, SHW23]. nearly [CGL08]. NEC [DTV00]. Need [GWH^{+}23]. Neighbor [CZD^{+}20, SHW23, WYY21]. Neighborhood [GCB^{+}23]. Neonatal [LMQ^{+}24, SZW^{+}23, Zha23a, ZWW^{+}23b]. Nephropathy [Ano23-34]. Nerve [Ano23-47, HHZ^{+}23, Liu23e, WZS^{+}22, WZS^{+}23, XWLL23]. Nervous [Zho22]. Nested [BS01]. Net [HPZ16, ZHP^{+}23, FZW23, HFH^{+}23, LZZ^{+}23b, Yan23g, ZZDY24]. Net/SLE [WOC99]. Netherlands [MCvM10]. Nets [LY23a]. Network [AMY^{+}21, Alk19, AMB23, Ano23t, Ano23x, Ano23c, Ano23-46, Ano23-36, Ano23-39, Ano23-47, AZX^{+}17, Bia23, CLDX20, Cao23a, CSWW23, CYLD23, CTD23, CT23, CCM^{+}23, CPG^{+}23, DPMB16, Du24, FDY^{+}19, FYY23, FZW23, FWX23, FYC^{+}20, GMCT15, GWZ23b, GLW^{+}23, GRB23, HI18, HKH^{+}17, HLHC23, HLC^{+}24, HLPY16, HTZL23, Hua23, HXPZ24, JSY^{+}21, JFY23, KYR^{+}23, KAS15, KZS^{+}20, KKK^{+}18, Kuf15, LLW^{+}22, LWZ^{+}23b, LXL^{+}23, LZWJ23, Li23f, Li23a, LL23c, LCH^{+}24, Lia23a, LFQA20, LWZC21, Liu22, LLL23c, LLC^{+}23, LTL^{+}23, LYZY24, LWQJ24, LZ23d, MCZ23, MLZ23, MABU18, Pen23, QWZP20, QTD^{+}20, RWN20, RLU^{+}20, SLS17, SW20, SY20, TZCZ23, TS23a, TZC23, WQLR17, WH20, WW23b, WWW23, WWQ23, Wan23h, WM23a, WZZ23, WZ23b, WL23a, WLQ23, Wan23i, WG23, WWYR22, XST^{+}16, XWG23, XL23, fXhCL17, Xu23c, XC23, YX23, Yan23b, YYL17, Yan22a, YJ23, Ye23, YYZ^{+}17, YLLL20, YZWF23, ZCXQ16, ZCW^{+}19, ZCG23, ZYZ23]. Network [ZL23b, ZLHZ23, Zha23j, Zha23i, ZWCY16, ZC20, ZD23a, ZS23, Zhu23b, ZD23b, ZWW^{+}23b, ASCH^{+}07, BDG^{+}94, CM02, CDD^{+}05, CIN^{+}96, DRT^{+}14, DRL13, PSU08, WOC99]. Network-Based [Hua23, JSY^{+}21]. Networking [GM14, KV23]. Networking-Based [KV23]. Networks [ÁRCPS15, Ano23e, Cha23a, CGC16, CD23, CSM15, DRMG^{+}18, GMCT15, GCB^{+}23, GKK^{+}20, HBR^{+}20, HJO^{+}20, He22, HPZ17, Hu23b, JRS^{+}21, KLL^{+}18, KZZ^{+}20, KZK^{+}23, LJL23, Ma23c, ML23a, Men23, MCF^{+}23, MJK23, NL23, PLW^{+}20, SSN15, SLC^{+}23, SF23, TZDZ20, TMG^{+}19, WXZW17, WXL19, WWW^{+}17, Xia23a, XWL19, XXSL23, Xu23b, YY23a, YXZ23, ZDMY23, ZWW16, ZM23, ZY24, ADS95, BDM^{+}04, DRL13]. Neural [AMY^{+}21, Alk19, AMB23, Ano23e, Ano23t, Ano23-47, AZX^{+}17, Bia23, Cao23a, Cha23a, CTD23, CT23, CCM^{+}23, CPG^{+}23, FDY^{+}19, FWX23, GWZ23b, GRB23, HI18, HBR^{+}20, He22, HLHC23, HTZL23, Hu23b, JSY^{+}21, JFY23, KLL^{+}18, LLW^{+}22, LWZ^{+}23b, LXL^{+}23, Li23a, LWZC21, Liu22, LLL23c, LTL^{+}23, LWQJ24, Ma23c, MCZ23, MLZ23, Men23, MJK23, MABU18, Pen23, QWZP20, RWN20, SW20, SY20, TS23a, TMG^{+}19, WH20, WW23b, WWW23, WWQ23, Wan23h, WZZ23, WL23a, WG23, XL23, XXSL23, Xu23c, Xu23b, XC23, YX23, YY23a, Yan23b, Yan22a, Ye23, YLLL20, YZWF23, ZYZ23, ZL23b, ZLHZ23, Zha23i, ZS23, ZWW^{+}23b, ADS95, CIN^{+}96]. neural-network [CIN^{+}96]. Neurological [HSY^{+}23, LLGH24]. neuronal [PSU08]. Neuroprotective [WHL^{+}21]. Neurorehabilitation [SWZL24]. neutral [SG96]. Newborn [LLW23a]. Newborns [JLJ24]. Newly [CTS23]. News [HAW23, Zha23d]. Next [Ma22, RN07]. Next-Generation [Ma22]. Night [LF23a]. Nimrod [PDA^{+}08]. Nimrod/E [PDA^{+}08]. NINJA [MMG^{+}02]. NKILA [ZMM^{+}23]. NLPA [NLPL^{+}20]. No [LGL23, WYC20b]. No-Reference [WYC20b]. No-Wait [LGL23]. NoC [MM16]. Node [Ano19, HLL23, HLC^{+}24, XTCC15, YCWL17, ZX17, BTS^{+}14, JR10]. Nodes [CSWW23, YYZ^{+}17, ZY24]. Nodules [FLR23, WZS^{+}22, WZS^{+}23, XWLL23, YLL^{+}23, Yin23]. Noise [SLH23, SZ20]. Non [YCLK23, JIC^{+}14, NAA^{+}03]. Non-Life [YCLK23]. non-regular [NAA^{+}03]. non-volatile [JIC^{+}14]. Nonalcoholic [LYL^{+}24]. nonaligned [AGIS94]. Noncontrolled [RAdASGR^{+}19]. Nondeterministic [MR18]. Nondisturbing [HPGC19]. Nondominated [ZXDY19]. Noninfectious [ZWW^{+}23b]. Noninterest [SWZS17]. Noninvasive [RVJ^{+}23, SZW^{+}23]. Nonlinear [DZL^{+}24, HI18, BCHL05]. Nonlocal [YPC^{+}23]. Nonmetric [CSZ^{+}18]. Nonnegative [TKZ^{+}24]. Nonrealization [DRMG^{+}18]. Norm [HLH23]. Normalized [AZGJ24, GG05]. North [LFY^{+}18]. NoSQL [AIA^{+}21, CTC^{+}15]. Notation [McC96]. Noticeable [ZWZ23a]. Notification [HPGC19]. Novel [BRSP18, Bra17, CZB24, CZKY24, DZL^{+}20, DDB^{+}18, FCZ^{+}17, FGROL^{+}16, FC22, GDNVSMM19, GMS^{+}23, HI18, HPZ16, JGF23, LWYS16, LLP23, PLY^{+}23, PCYZ18, ROMFST23, SAN^{+}19, WWLG16, Yan23g, Yi23, YLZ23]. NSGA [YPM^{+}21]. NSGA-II [YPM^{+}21]. NSGA-III [YPM^{+}21]. Nucleus [FZW23]. NUMA [BCC^{+}00, KKK^{+}18, MJB15, TKS02]. NUMA-Aware [KKK^{+}18]. NUMA-based [TKS02]. Number [DZL^{+}20, GB94, NM16, WXL19]. Number-crunching [GB94]. numbers [Che93a]. Numerical [MLZ23, Wu23b, BC99, BAN02, DSZ96, DHH00, MMG^{+}02, VF95, WCG95]. numerics [Bjø00]. NumPy [AJJF14]. Nursing [Ano23n, CH23b, GLD^{+}24, Li23d, LWZ^{+}23c, Liu23c, LYZ^{+}23, SKT^{+}24, WZ23a, WLZJ24, YLLY23, ZY23, yZcR23, ZCL^{+}24b]. Nutrition [ZWW^{+}23b]. NVivo [WZ22]. NVRAM [ZLL^{+}15].

O [OSLK12, LKS18, NsP16]. Object [AMY^{+}21, BCC^{+}93, CW93, Gus93, HLKY19, HMSW92, HJO^{+}20, LDZZ23, MOT97, NCMF15, NL19, QWP23, SSSJ20, BGLR93, BBG^{+}93, BKP99, FS01, HST^{+}93, KZRR94, KK11, VF95, WGF93]. object-based [KK11]. Object-Oriented [HMSW92, NCMF15, NL19, BCC^{+}93, CW93, Gus93, MOT97, BGLR93, BKP99, HST^{+}93, KZRR94, VF95, WGF93]. objective [DF14]. ObjectMath [VF95]. Objects [DDJ20, MR02]. Observational [HMQS23]. observations [HLM^{+}11]. observed [BH02]. Observer [KLuH^{+}20]. Obstacle [WLCC18]. Obstructive [HCC23, TY23, XZLZ24]. Obturator [HHZ^{+}23]. Occlusion [SHCZ23]. Occupancy [ZL23c]. Occurrence [JNK24, LWCZ21]. ocean [HMCH07, Mal01]. OCR [MSA^{+}20, ZLWW23]. oct [War14]. oct-tree [War14]. Octagonal [NM16]. ODE [RR04]. OFDM [RLU^{+}20]. Offensive [Che22]. Offloading [ZHW^{+}21]. Oil [Ano22b, Li24]. Oligonucleotide [KTP05]. Oliveira [Ano07c]. Omics [ZTL^{+}23]. Omnimedia [Cao23b]. OMPM2001 [AE03]. on-chip [JK10]. on-line [BBF^{+}04, RLL^{+}02, SCB02]. On-the-Fly [HJ15]. On-Water [QK21]. One [Alk19, GBH14]. One-Dimensional [Alk19]. Onion [Alt23, KKA^{+}22]. Online [AN19, Ano23y, Ano23-48, AMT20, Cao23b, CLLZ23, CH23a, HKH^{+}17, HAW23, JW23b, KGU^{+}20, KGU^{+}21, LWYS16, LWW23a, LWX23b, MLHC20, PWP21, QM23, SK21, Son20, SW23, SWDZ24, WW23a, tXChL^{+}24, YN23, yZcR23]. onto [DSS^{+}05]. Ontology [CHZY17, LMZP23, SDM^{+}19]. Ontology-Based [SDM^{+}19]. OON [VC93]. OON-SKI [VC93]. Open [Ano23t, ZWXL21, ZC23, SGM^{+}08, SGM^{+}08]. Open-Map [ZWXL21]. OpenACC [CCBPGA15, MLYH17]. OpenCL [FSV14, LNK^{+}15, THS^{+}15]. OpenFOAM [LPD20]. OpenMP [ARvW03, BHP^{+}03, BCC^{+}00, BS01, BDV03, CCBPGA15, DBVF01, HJYC10, JJY^{+}03, JR10, KB01, LOHA01, Mal01, Mar05, Mat03, MJB15, Mül03, NPP^{+}00, PVKE01, PKE^{+}10, SHHI01, SKS01, SB01, Wan02]. Operating [WSP94, ITF^{+}08, KZRR94, SCSJ09, WMR^{+}94]. Operation [CL23a, MCZ23, RWY^{+}17, ZCL^{+}20]. Operational [BYZ^{+}23, Ros00, WFHC20, ZCT^{+}20]. Operations [FLS17, YZSZ17, ZZW17]. Operator [ZCZ^{+}23]. operators [GRC03]. Opinion [GMM20, KS15, LYZY24, QHAN23]. Opinions [DRMG^{+}18]. Opposition [XYY^{+}19]. Opposition-Based [XYY^{+}19]. Optical [ASZF23]. Optimal [ACLS16, CWLW23, KLL^{+}18, KFU^{+}20, LZF16, LL23b, MWZ^{+}23, PPJ^{+}15, QZY^{+}20, ZWW16, Zha23l]. Optimisation [CBE23, GLW^{+}23, WFZ^{+}23, WD07]. Optimisers [Sah20]. Optimization [ASS23, Ano23-43, Ano23-44, Bia23, BNH^{+}20, CC23, CLTX16, CZCL23a, CWH23, CZWY23, CW23, FTS^{+}20, FMA16, FLS17, FGSCR^{+}17, GKK^{+}20, HI18, HSH23, HQL^{+}23, HL23c, JAED23, JZZW20, JL23b, JFY23, KAA^{+}23, LLP^{+}19, LCW22, LHMC23, LZZ23a, LJL23, Li23e, LZSC23, LY23b, LZ23b, LZLZ19, LLY^{+}20, LSY20, LWW23a, LCL^{+}23b, LQW23, LFS^{+}20, Lv23, MYH19, MFB^{+}15, McC96, MLL23, NTNTTK^{+}19, PCYZ18, QYZ17, QHAN23, RWY^{+}17, SLH23, SW16, SK21, SZC16, SCAP16, SGPCLG19, TH18, TPSM17, Tia23, Ton20, TLH17, Tun23, WB16, WCLL17, WQLR17, Wan21, WH23, lWwFAH23, WZ23b, WW23a, WZS^{+}23, WZF^{+}20, WMZM23, fXhCL17, XYY^{+}19, YZ16, YCWL17, Ye23, ZJZ^{+}16, ZLZ^{+}22, Zha23g, ZZW17, ZQL23, ZL22, BBLR03, DHH00, DZKS13, GSM03, LBvBW12, SKS01, Sin99, WSB11, XNQF04]. Optimization-Based [FTS^{+}20, TH18]. Optimization-Guided [Ye23]. optimizations [Ang93]. Optimized [Ano23-47, BLL^{+}23, CZKY24, CGC16, GWZ23a, HLHC23, JZH^{+}23, JHL^{+}23, LHC^{+}23, LX23d, PP23, PWY23, SYW23, THS^{+}15, YZX17, YLLY23, ZCL^{+}24a, ZLL^{+}15, ZLZ^{+}23, ZZ23, ZC23]. Optimizer [JHS^{+}20]. Optimizing [BMN^{+}97, CSC16, GRB23, LWT18, Liu23d, SVR^{+}07, YZZ15, Zhe10, ZKA23, BAP13, BHK^{+}13, CKS^{+}09, SO93]. Opus [CHM^{+}97]. Oral [Ano23t, Hui22, LWQ^{+}23]. Orchard [CLL^{+}20]. Order [CZH24, FH23, XWG23, BH02]. ordering [BÇCD12]. Orderliness [CL23a]. Orders [ZXHQ17]. ordinary [Gus93]. Organ [LLY^{+}23b]. Organizational [CL23a, LF23b, SAZ^{+}24]. Organophosphorus [SWY23]. Orientation [SC16]. Oriented [CZS^{+}23, DPMB16, HMSW92, HLY^{+}17, NCMF15, NL19, XZ23c, ZZS^{+}23, ZD23b, BGLR93, BCC^{+}93, BKP99, CW93, Gus93, HST^{+}93, KZRR94, MOT97, VF95, WGF93]. Original [ZXDY19]. Orthopedic [LBW^{+}24]. OS-Level [AK18]. Osteoarthritis [SYW23, ZZ23]. Osteoarthropathy [Yan23b]. Osteomyelitis [OWL^{+}23]. Osteoporosis [GLD^{+}24]. Osteoporotic [HZL^{+}23]. Osteosarcoma [HTZL23]. OTO [WDY23]. Otto [Nag05b]. Ottoman [JK22]. Out-of-Core [TC96, JIC^{+}14]. Outcome [AKA^{+}23]. Outcomes [LWZC21, ZLHZ23]. Outdoor [He23]. Outlier [XML^{+}16]. Outline [HY23]. Output [tXChL^{+}24]. Ovarian [ZHHX23]. Overall [YHJ^{+}23]. overcome [BFH96]. Overflow [YJ23]. Overheating [SZL^{+}23]. Overlap [DYYW24]. Overlapped [MYH19]. Overlapping [KB01, KCS15, PKE^{+}10]. Oversampling [WYC^{+}20a]. Overview [AJ94, GN22, HNSG23, WJCZ21, XNLJ23, HW12]. Oxygenation [ZCLL23].

P [Qin23, FP00a, OPE^{+}95]. P2P [ASCH^{+}07, LWT18, PWP21]. P4 [Lar93]. Pack [LGL23]. Package [SSM^{+}02]. Packages [SLL^{+}17]. packaging [CL04]. Packing [FH23]. PADDA [Ger02]. Pain [FLY^{+}23, He24, JFY23]. Paint [QQWJWYQZ23]. Painting [WX23, XL23]. Pair [HNMB19, SXTK18, YG16]. Pairs [LY20]. Pakistani [AUA^{+}23]. Palate [DHW^{+}23]. Pallet [bLwHcY18]. Pancreatitis [LLY^{+}23b]. Pandemic [WW23a]. Panel [JZL^{+}20]. Panoramic [Ano23-41]. Papers [Old00, VR13, Ano97b]. Paradigm [Ano19, WWZ17, ZX17]. Parallel [AFU18, Cao23a, CBCM93, CLA^{+}15, CLTX16, CR05, CEMM17, CSU^{+}17, CPD21, DDMS15, FP95, FOT92, GIKP95, GG05, GBK^{+}96, HYS^{+}19, Hui23, JFPL16, KGIL19, LWL23b, LPHD04, LLY^{+}20, Nag04, NYML17, O’K00, Ott93, PAR94, PVKE01, SZAG15, SO11, SJ15, SAZ^{+}24, SE95, SO15, SCAP16, TZL^{+}24, TLC15, WGF93, XH20, XSZ18, gXySR^{+}17, YX23, YDC^{+}16, YXZ23, YZSZ17, ARvW03, ARFS05, AJ94, ADS95, AGIS94, BCC^{+}93, BBK^{+}11, BBG^{+}93, BB09, BÇCD12, BFH96, CLM05, CM02, CvHK97, DBVF01, DRR12, EAS^{+}97, FP00a, FB99, GG95, GL04, GRN99, GRR^{+}03, GA96, HMC97, Hav00b, HZ99, HB94, HJ96, KMR^{+}97, KZRR94, KLS^{+}96, Lar93, LH93, Lin04, Mar05, Mic00, MSLS13, OPE^{+}95, OdSSP13, OSS94, PVL^{+}04, PMCF94, RR07, RKR^{+}99, RS94, SCSJ09, SS00, SDS00, Sch94, SJR14, SGM^{+}08, SBM^{+}10, Shu94, SCB02]. parallel [SBJV11, VGC09, WCKD07, War96, War14, WMR^{+}94, WB95, WCG95, WSP94, WGW08, ZS99, JEM07, PDGQ05]. parallel/distributed [FB99]. parallelisation [IJL^{+}01]. parallelism [BS01, HPD09, JK10, KB96, RRV09]. Parallelization [BA18, HLKY19, JR10, KBRS95, LYTL20, MLYH17, Vol97, BCS01, JJY^{+}03, Kes96, LP99, MSH99]. Parallelize [RRM^{+}15]. Parallelizing [KCS15, XHSL17, OB96]. Parameter [ASZF23, COB^{+}00, CLTX16, GRB23, LWZ^{+}23c, WZS^{+}23, YPM^{+}21, YLLL20, ZL22, PDA^{+}08, TOF^{+}14]. Parameters [LZWL23]. Parenteral [ZWW^{+}23b]. Parkbench [HB94]. Parkinson [CZY^{+}23, WHL^{+}21]. PARMACS [HZ99]. Parsing [SJD20, ZHZF20]. Part [SCAP16, ZGS16, HR12a, HR12b, PPS12a, PPS^{+}12b]. Parthenogenetic [SLBZ16]. Partial [CCB^{+}20, KLuH^{+}20, PPS^{+}12b]. Participation [HLZ^{+}21]. Particle [BNH^{+}20, CZCL23a, CZWY23, CW23, DCH^{+}23, GKK^{+}20, HI18, LZSC23, WMZM23, Ye23, YLZ23, ZLZ^{+}22, ZQL23, CBCM93, JAED23]. Particulate [SAN^{+}19]. Partition [CZ23, ZCS23]. Partitioned [XZNL23]. Partitioning [BÇCD12, HLKY19, HLY^{+}17, MOS16, CP95, LR05]. Partnership [KJG^{+}08]. Passenger [Xu23a, ZWG^{+}20]. passing [PBK01]. Pastes [MLAS21]. Patellar [GWZ23b, ZRX^{+}24]. Patent [HXPZ24]. Patent-Based [HXPZ24]. Path [AMT20, CLXZ23, DS23, HSH23, Hon23, LZF16, LCH^{+}24, LW23, Liu23g, NM16, QK21, yXHtJC23, Yan23e, YLZ23, Zho20, ZC23]. Pathogenic [CL23b]. Pathological [YT22]. Patients [AKA^{+}23, Ano23n, BGZL23, BLZG23, CZY20, CH23b, GXY^{+}20, HL23a, Hon24, Hua21, HSY^{+}23, KSYC23, LZ23a, LWL^{+}23c, LSS^{+}23, LBW^{+}24, LLY^{+}23a, Liu23c, LLGH24, LLY^{+}23b, NHHW23, PP23, SWZL24, TY23, TYCL23, WHL^{+}21, Wan23f, WZ23a, WCZW23, YLLY23, YZWF23, ZCL^{+}24a, ZYZ23, ZZW^{+}23, ZLZ^{+}23, ZWLL23]. Pattern [Ano24, ARGC19, CD23, CLW^{+}23, DWZ^{+}20, GRC03, JNK24, JZS21, Kes96, LQW23, REPRL^{+}19, Xu23c, ZG18, ZPA19, DRT^{+}14]. Pattern-Based [DWZ^{+}20, REPRL^{+}19]. Pattern-driven [Kes96]. Pattern-Recognition [ARGC19]. Patterns [EOS17, HMR^{+}15, Zhu22, CFR14, ESSL99, GSM03, ZS99]. Paul [Hil97, Mic97]. Payment [Ge23]. PBD [TZZ23]. pC [BBG^{+}93]. PCA [Jua23c]. PCN [FOT92]. PCNN [TTL^{+}23]. PDDP [War96]. PDE [KK09, LBvBW12, SF03b, SF03a]. PDE-constrained [LBvBW12]. PDEs [BEK^{+}12]. PDS [BDLL94]. PE [WDY23]. Peak [ZLCW20, SH94]. Pectoris [HL23a]. Pedestrian [TZCZ23]. Pediatric [CWLW23]. Peer [GPX^{+}21, XNQF04]. Peer-to-Peer [GPX^{+}21, XNQF04]. Pegasus [DSS^{+}05]. Pelvic [PWY23, ZLZ^{+}22]. Pelvis [JHL^{+}23]. Penetration [JZL^{+}20]. Penile [Ano23i]. Peninsula [PZH23]. Pension [HSH23, LJL23]. penultimate [Met99a, Met99b]. People [AXA20]. Perception [LLZW23, WY23]. Perceptron [AMB23, PCYZ18]. Perceptual [Ano23w, SZ20, TTL^{+}23]. Percutaneous [HZL^{+}23, Wan23f]. PerfExplorer [HMSM08]. Performance [ACA19, ACIK97, Ano93, Ano99b, ACCTA^{+}15, AA23, AGG^{+}97, BYZ^{+}23, BMN^{+}97, CTS23, CCBPGA15, CPK19, CPD21, DS97, DTV00, GA18, GB15, GLBB17, GM14, GGP09, GBK^{+}96, HKH^{+}17, HMR^{+}15, HP02, HS23b, JMR^{+}11, KMR^{+}97, KOM94, KZRR94, KGV97, LLP^{+}19, LNK^{+}15, LLC^{+}20, MJLM07, MSSG11, MH95, Nag05a, PRW23, Sah20, SXTK18, SCD^{+}15, SAZ^{+}24, SSC97, SMD20, SWZS17, TLH17, WZ22, WGW08, XJH23, XXSL23, Xu23d, YSL^{+}18, ZZS^{+}23, Ano08b, AB96, AE03, BBF^{+}04, BDLL94, BTS^{+}14, BDV03, CLM05, CMM^{+}02, CB99, CKS^{+}09, ESP^{+}12, FP00a, FSV14, FP00b, GGMS99, GRR^{+}03, HCJ08, HLM^{+}11, HKL^{+}12, HMSM08, IR02, KMB09, Lan03, LDV07, LG03, MMG^{+}02, NJ08, Nor07, SS00, SH94, SGM^{+}08, SBM^{+}10, SE95, TOF^{+}14, TF04, WHRH07, WSZK09]. performance-portability [ESP^{+}12]. performance-prediction [BDV03]. Performances [LWX23b]. performing [PMTL12]. Perinatal [LWZC21, ZLHZ23]. Perineotomy [Ano23-47]. Period [WZ23a]. Periodic [JNK24]. Perioperative [WZ23a]. Peripherally [CH23b, YLLY23]. PERLEX [ABNJMB23]. Permeability [PLAAH23]. permutations [BSF96]. Persian [ABNJMB23]. Persian-English [ABNJMB23]. persistence [CDD^{+}05]. Personal [Ma22, TMG^{+}19]. Personality [FZ23]. Personalized [TL17, YC23, ZLF18, ZW18, ZW23b]. Perspective [ACBV20, CWD23, CZWY23, DPMB16, JUA^{+}23a, LRX^{+}17, LWT18, LNH^{+}20, Lin23b, LWX23b, SKA^{+}20a, SKA^{+}20b, UAK18, WG22, ZW23c, ZKA23]. Perturbation [CZCL23a, Tun23]. peta [Zim07]. peta-scale [Zim07]. petaflops [Mic97]. PetaShare [KABW11]. PETCT [PP23]. Petri [LY23a, HPZ16]. Pfortran [BCS01]. PGHPF [BMN^{+}97]. Pharmaceutical [HMQS23]. Phase [ACLS16, SAN^{+}19, TC96, VMRM16]. Phi [ACCTA^{+}15, DGH^{+}15, SCD^{+}15, SRO^{+}15, TSP^{+}15]. Phones [AAF19, FA24]. Photon [CPG^{+}23]. Photoplethysmograph [LXRC21]. Physical [Ano23z, Guo23a, Wan23h, WWL^{+}23, XY23b, Zha23j]. physically [LLW^{+}14]. physics [BHK^{+}13, RX04]. Piano [Li22]. PIC [BDV03]. Picking [FH23]. Pictorial [Bry96]. Picture [JZY23]. Piecewise [XYY^{+}19]. Pig [JZH^{+}23]. Pipeline [DDMS15]. Pipelined [CWH23]. Pipelines [SLS17]. Pipelining [ATZ^{+}15, WCF18]. PIR [KAA^{+}20]. Piracy [LNH^{+}20]. Place [SQW^{+}21, FO96]. Placement [ACLS16, CGC16, FZW^{+}18, HKH^{+}17, LLS^{+}20, WZW^{+}20, ZHL16]. Placements [Wu16]. Plagiarism [FXYY17, HNMB19]. Plan [GWH^{+}23]. Plane [WLLY23]. Planning [QK21, SL21, YLZ23, ZMLZ18, ZCG23, Zha23g, Zho20]. Plans [ARGC19]. Plant [Ano23j, PLY^{+}23, SQJG23, fTPsLZ23, WQLR17, YY23c]. Plaque [Ano23-31]. Plasma [ZRX^{+}24, LG03]. Plastic [CWSF23, WJS^{+}23]. Plate [RAdASGR^{+}19]. Plateau [FHC^{+}22, XLB^{+}23]. Platelet [ZRX^{+}24]. Platelet-Rich [ZRX^{+}24]. Platform [Ano23p, Ano23a, CT23, EOS17, GWZ16, HL23d, Hui23, KF20, bLwHcY18, LZYJ20, LWX23a, LZLJ23, MLHC20, PWP21, She23, TKZ^{+}24, UAU^{+}23, WL23b, XST^{+}16, XTZ20, CGK^{+}05, PCGLO14]. Platforms [CLL17, GPX^{+}21, HCGGSS17, OKHS23, YT23, CFR14]. Playa [HKL^{+}12]. Player [Guo23a, HLZ^{+}21, Hu23b, HLH23, LLL23c, Qin23]. Players [Tan23, WY23]. Plexus [GSL23, WCW^{+}21]. PLS [BYZ^{+}23]. plume [WCKD07]. Plus [Lou23]. PMVS [He23]. Pneumonectomy [LLL^{+}23a]. Pneumonia [CWLW23, LLW23a, WZL23, WJZC23].

POI [LTZS18]. Point [DCF17, JFY23, LDZZ23, QL23, XYY^{+}20, KMB09]. Point-Based [LDZZ23]. pointer [HS03]. pointer-linked [HS03]. Poisoning [SWY23]. Polar [ZSD22]. Police [WG23]. Policy [CWD23, JFPL16, KFU^{+}20, LNH^{+}20, LZLJ23, SZ04]. Policymaking [Son23a]. Polish [ZG15]. Political [Ano23e, Ano23x, Hon23, JL23b, SF23, Wan21, Wan23d]. Politician [KZS^{+}20]. Pollen [WWW23]. Pollution [Lia23a, OZI^{+}23]. Polyadic [ML23a]. Polyhedral [SO18]. polymorphism [DN04]. Polypectomy [LCL^{+}23a]. Polyps [GGSL23, HWSL22, LX23b, LCL^{+}23a]. Polyshift [GBJ94]. Pool [LZC^{+}24]. Pooling [bLwHcY18]. Popular [Jua23c, SC23]. Port [DGH^{+}15]. portability [ESP^{+}12]. Portable [Bra17, BŻvA^{+}01, GL04, KMR^{+}97]. Portal [He24, KBG^{+}02]. Portfolio [FGSCR^{+}17, YLW21]. Porting [ARAM19, HHML05]. Portrait [FYC^{+}20]. Ports [ZMLZ18]. Positioning [PRW23, WW22]. Positron [TYCL23, WZS^{+}23]. Possibilistic [CZP^{+}24]. Possibility [AUA^{+}20, CZWY23]. Post [CLLZ23, CC23]. Post-Based [CC23]. Postal [YSSL23]. Postepidemic [GZD23]. Poster [Sun23b]. Posterior [CLHH20]. Postoperative [BGZL23, PWY23, WLLY23, ZLZ^{+}22]. Posture [Hu23b, HLH23, SW23, Yua20]. Potential [CLPVPI20, LPBSSEPL19, SLBZ16]. Poverty [LLZW23, ZW23c]. Poverty-Stricken [LLZW23]. Powder [ZWW^{+}23b]. Power [Ano23-44, Cer21, CCF^{+}17, CHTZ23, DZ18, DZL^{+}24, DGL^{+}17, FLC^{+}20, FYC^{+}20, KGBB09, LWZ^{+}23a, SFZ^{+}20, TMY^{+}23]. Power-Efficient [CCF^{+}17]. PPA [KZZ^{+}21]. PPA-ResNet [KZZ^{+}21]. PPAR [DCH^{+}23]. PPG [YHZ^{+}23]. ppOpen [TOF^{+}14]. ppOpen-AT [TOF^{+}14]. Practical [ARGC19, CGMRO^{+}19, KOM94, FML^{+}14]. Practice [DS23, LGZ23, WL23c]. Precise [HBCM94, Mor94]. Precision [ZM23, Sch03a]. Precocious [QYL23]. preconditioned [GG05]. preconditioners [JHNP14, MOT97]. preconditioning [GHST12]. Precursors [SWW^{+}24]. Predict [AKA^{+}23, HS23c, PLY^{+}23]. Predicting [ACA19, GXY^{+}20, GRB23, KZS^{+}20, LHMC23, XJH23, XHC^{+}23, YWGL20, ZWZ^{+}23b]. Prediction [AMB23, Ano19, Ano23m, Ano23-30, Ano23-33, CIU20, qCtWfPbZ23, CYSZ23, CW23, DZL^{+}24, FDY^{+}19, FX23, GMCT15, GWZ23a, GLBB17, GPX^{+}21, Guo23a, HLHC23, HW23a, JAED23, JZH^{+}23, LXRC21, LLP23, LLY24, Lia23a, LWW23b, LLC^{+}23, Luo23, LLY^{+}23b, MCB^{+}20, Pan23b, PLAAH23, Ros00, SKB^{+}24, SJD20, SSN15, SQJG23, SWW^{+}24, SPH20, TS23a, WHH^{+}21, WCZW23, Wan23a, WLL23a, WL23a, WLL^{+}23b, WDD^{+}23, XHX23, XXSL23, Xu23a, XZ23b, XC23, YYF^{+}19, YHJ^{+}23, YT23, ZFX17, ZLCW20, ZZWZ23, Zha23e, ZGY^{+}20, ZX17, Zho22, Zho23, ZZSQ23, ZD23b, AMN^{+}12, BDV03, FP95]. Predictive [SAN^{+}19, YY23c, GRN99]. predictive-adaptive [GRN99]. Prefabricated [LY23b]. Preface [Ano04b]. Prefiltering [GB15]. Pregnancy [LTL^{+}23]. Preliminary [Ano23-34, ZLT^{+}21, BKP99]. Premature [ZSZ^{+}23a]. Prenatal [DHW^{+}23]. Preoperative [WLLY23, ZLZ^{+}22]. Preparation [CIU20]. Preprocessing [NLPL^{+}20]. Preprocessor [CGA17]. Preschool [Bai23]. Presence [GKK^{+}20]. Preservation [IPST18]. Preserving [JRS^{+}21, WZW19]. Pressure [CKKK15, WL23a]. Pretraining [WSM^{+}21]. Prevent [YP24]. Prevention [KFU^{+}20, TMHM23, WZS^{+}22]. Prewarning [LFY^{+}18, XCW18]. Price [CFL^{+}23, CZH24, Zha23e]. Prices [HS23c]. Pricing [SL21, ZGS16, BTS^{+}14]. Primary [ZGL23, ZC23]. Primers [Ma22]. Primiparas [JHL^{+}23]. primordial [BAN02]. Principal [WmG21]. Printing [Ano23h, Chu21, Lia23b]. Prior [YDC^{+}24]. Prioritization [AAA^{+}23, SBK^{+}24]. Priority [YZSZ17]. Privacy [JRS^{+}21, KAA^{+}20, LZLJ23, PLW^{+}20, XWL^{+}23]. Privacy-Preserving [JRS^{+}21]. Private [DQQ17, tXChL^{+}24]. Probabilistic [XBG17, YL20]. Probabilities [Ser18]. Probability [CZW^{+}23, HKVVNT18]. Problem [BRSP18, CCG17, CBRRC^{+}16, GWZ16, LGL23, OZI^{+}23, SQ21, TFZ^{+}20, WCLL17, WQT^{+}20, yXHtJC23, YSSL23, ZWXL21, GB94, Lan01, VRW^{+}03, WGF93]. Problems [BD16, FMA16, Gog23, HYS^{+}19, JZY23, MABU18, SW16, SCAP16, Tun23, YW24, GHW00, JHNP14, SDS00, VHBR93]. Procalcitonin [SWY23]. Process [Ano23-46, AAA^{+}23, Chu21, FLS17, HS20, JUA^{+}23a, KLuH^{+}20, LFS^{+}17, LFQA20, Liu23a, MOS16, MAC^{+}20, PLAAH23, SL23c, lWwFAH23, XHL^{+}23, ZLLZ20, ZZS^{+}23, BH02, Shu94]. Process-Based [AAA^{+}23]. Processes [ARGC19, HS20, Ser18, CR05]. Processing [AXWS23, Ano24, Che23b, FC22, Fen23a, JZS21, LF23a, MLAS21, SL23c, SL23b, TVCB23, YB24, tZhYsHlZ19, BB09, Den96, FP95, TFN11, VF95]. Processor [MJK23, GHH^{+}02, HMCH07, KGBB09, Sin99]. Processors [ATZ^{+}15, MJB15, YZZ15, ARFS05, Ano08b, KMR^{+}97, OPE^{+}95, PTS^{+}13, PRM^{+}14, PCGLO14, Nag04]. Procurement [Ano22b]. Produced [CBRRC^{+}16]. Product [Ano23-35, Liu23a, WQLR17, XZNL23, ZYZL23, GA96, KHSJ95]. Production [HLY20, JZH^{+}23, Yan23a, ZGS16]. Productive [FOT92, ZJZ23]. Productivity [LQZ^{+}21, YSL^{+}18, YCLK23, Zim07]. Products [Lia23b, LYZY24, PVCPdPSZVG17, SAZ^{+}24, SL21, SL23c].

Professional [ÁRCPS15, ZGL23]. Profile [SC16]. profiling [BB08, CKS^{+}09]. Profit [KS19]. Prognosis [HSY^{+}23, WCZW23]. Prognostic [HTZL23, LWQ^{+}23, MTK^{+}23]. Program [CTS23, PST19, RWN20, SLL^{+}17, SLDL17, WKY23, CMM^{+}02, HST^{+}93, RR04]. programmable [HKL^{+}12]. programmer [OB96]. Programming [AMVGC^{+}23, AHMMÁR18, ÁRAHMM18, BRSP18, CCTY15, CMZ92, CYZ^{+}20, CYJ^{+}20, CPD21, DMAC23, DS97, DDMS15, DDJ20, DGH^{+}15, FOT92, GDNVSMM19, GBK^{+}96, KPW^{+}18, KGBB09, KK11, LGC^{+}16, LRX^{+}17, LJF20, MKH^{+}19, Nag04, NGMAA23, PT09a, PT09b, QYW^{+}17, Ray17, RLPI19, SXTK18, SW16, SZC16, SRTCL16, SGPCLG19, TFZ^{+}20, WB16, WCF18, WLZ^{+}21, XTCC15, XHSL17, fXhCL17, YZ16, ZHZS19, ZCG23, ZLL^{+}17, ZGL18, ZYX^{+}21, Zha23k, BH12, BBK^{+}11, BB09, BFH96, CLM05, Che93a, Che93b, CvHK97, DHH00, EAS^{+}97, GL04, GBH14, GGMS99, GHW00, HMC97, Hav00b, Lar93, Mat94, NDSG07, PVKE01, PMM94, PPS12a, SBM^{+}10, Slo12, Sny07, Szy07, VCT05, Wan02, War96, Hil97]. Programs [BCC^{+}92, CGA17, LKS18, MR18, MJB15, NO18, Den96, DF14, FP00a, GG95, GRR^{+}03, LDV07, LP99, OPE^{+}95, OdSSP13, RS94, SKS01, TKS02, Zhe10]. Progress [ZWZ^{+}23b]. Progressive [ZXS^{+}23]. Project [ARGC19, CCF^{+}17, Du23b, FGSCR^{+}17, LFY^{+}18, Ye23, Zha20b, HW12, HR12a, HR12b]. Projection [AZGJ24, Li23b]. Projections [SHW23]. Projects [HLS^{+}24, MCZ23]. Proliferation [ZMM^{+}23]. Promotes [ZMM^{+}23]. Propagation [CZD^{+}20, HZFH23, LW23, LJZ18, MWZ^{+}23]. properties [Din99]. Property [TVCB23]. Propofol [BLZG23]. Proportionate [AZGJ24]. Propulsion [ZXX^{+}23, WCKD07]. Prospect [MCZ23].

Prostate [LZZ^{+}23b, WZC^{+}23, YHZG23]. Protect [CGMRO^{+}19]. Protection [XWL^{+}23]. Protective [BLZG23]. Protein [CLA^{+}15, CCB^{+}20, MCB^{+}20, PLY^{+}23, SQJG23, SY20, SWY23, ZWW^{+}23b, KG08, WSZK09]. Protein-Protein [MCB^{+}20, PLY^{+}23]. Proteins [WLL^{+}23b]. Protocol [SZWZ23, MPP^{+}04]. Protocols [KAA^{+}20]. prototyping [DSZ96]. Provenance [PY17]. Province [QHJ23]. Provincial [JZL^{+}20, LCH^{+}24, LQW23, YCLK23]. Proving [RAA23]. provisioning [JDVM10, OPP11, Slo12]. Proximal [WPDZ19]. Pruning [LWQJ24]. PSEs [HP02]. Pseudo [NYML17]. PSO [CZKY24, YLLL20]. PSO-RBF [YLLL20]. PSOCoG [JAED23]. Psoriatic [YPY23]. PSPNet [YY23b]. PSR [SAN^{+}19]. PSR- [SAN^{+}19]. PSSM [WLL^{+}23b]. Psychological [Cha23b, Li23f, Wan23a, WL23a, VRM02]. Psychology [BL18, Yan22a]. Psychology-Inspired [BL18]. Puberty [QYL23]. Public [LL23b, LYZY24, MKS^{+}24, PAR94, RuHMS24, Son23a, TPSM17, WW23c, WG22, ZG23, HB94, RST02]. Publications [DD23a]. Pudendal [Ano23-47]. Pulmonary [FLR23, HCC23, JLJ24, TY23, XZLZ24, Yin23, Zha23a, ZZW^{+}23, ZWLL23]. Pulse [Ano22a]. PUMA [WMR^{+}94]. Purchase [SL21]. Purchasing [Li24]. purpose [Mat94]. Pursuit [YWC^{+}18]. PVAR [Sun23c]. Pyramid [HJO^{+}20]. Python [CLM05, DDMS14, Spo12]. PyTrilinos [Spo12].

QAM [AE19]. QCD [IB09]. Qigesan [Ano23-29]. Qingdao [YHH^{+}23]. QoS [HLPY16, LGJ18, MAG^{+}07, Slo12]. QR [CDO^{+}96, KD09, yZcR23]. Qu [HSL^{+}23]. Quadratic [NTNTTK^{+}19]. Quadrotor [HHCC23]. Qualification [CLL^{+}18].

QUALIFLEX [DZL^{+}20]. Qualitative [SZX^{+}20]. Quality [ÁRCPS15, CEMM17, Han23, HLHC23, HLS^{+}24, JW23b, LCH^{+}23, Son23b, SH19, SNH24, VVNHTHNT20, WYC20b, XWL19, Yan23f, YN23, KFFZ05, PMCF94]. quality-cost [PMCF94]. Quantification [KAA^{+}20, GW11]. Quantile [CZY20]. Quantitative [Ano23b, AE03, CZY^{+}23, DMV^{+}19, Lou23, Pan23b, WWH^{+}23, ZCL^{+}24b]. Quantum [Don23, PLJ^{+}23, RX04, KJG^{+}08]. Quasi [PPJ^{+}15]. Quasi-Optimal [PPJ^{+}15]. Quaternion [ZSD22]. Query [CWH23, HLY^{+}17, JZZW20, WWZ17]. Question [LL23d, WLZ^{+}21, ZS23]. Question-Answering-Style [WLZ^{+}21]. Questions [SHM97, SN02]. queueing [DDMS14]. QuPiD [KAA^{+}20].

R [KCO^{+}05, XLY18]. R-CNN [XLY18]. R-GMA [KCO^{+}05]. Rabbit [DCH^{+}23]. Race [Ano23-44, HJ15]. Racing [YZ16]. Radar [LL23a, tZhYsHlZ19, PTS^{+}13]. Radiofrequency [CWMZ23, HZHW23]. Radiomics [YHJ^{+}23]. Radiotherapy [CZD^{+}20]. Rail [HWY^{+}23, Zha20a, ZWG^{+}20]. Rail-Based [HWY^{+}23]. Railway [XLY18]. Random [CCB^{+}20, FZH23, LLC^{+}23, STM^{+}23, Ser18, SCZ18, SHW23, Tun23, WXZW17, XWG23, XY23b, Yan22b, ZCLL23]. Randomization [ZWZ16]. Ranging [XLF^{+}24]. Rank [FHC^{+}22, GDA^{+}23, HEVD19, LZL^{+}23, SMD20, YTD15, YYL17, YPC^{+}23, ZZZ^{+}23]. Ranking [Ano23q, FGSCR^{+}17, Gog23, NM16, SMD20]. Ranks [LFQA20]. Ransomware [UJS^{+}20]. Rapid [CLL17, WLZJ24]. Rate [Ano23-33, DLP^{+}23, HH23a, LWT18, MSZZ20, ZFX17]. Rating [LLP23]. Ratio [VMRM16]. Ray [GLD^{+}24, LZWL23, SZW^{+}23, ZWW^{+}23b, BCHL05]. Rays [FMG^{+}23]. RBF [YLLL20]. Reactive [SWY23]. Reading [He20, WLJ22]. ready [TSCT11]. Real [BDAH20, BD16, GWH^{+}23, GMS^{+}23, PDROFRM16, She23, SWDZ24, TMHM23, TH18, WJS^{+}23, XYL^{+}23, YSSL23, ZCZ^{+}23, Bry96, JR10, MAG^{+}07, SH94]. Real-Estate [GWH^{+}23]. Real-Time [BDAH20, BD16, TMHM23, TH18, XYL^{+}23, ZCZ^{+}23, MAG^{+}07]. Real-World [YSSL23]. Realization [Hua21, WC23b, Yan23e]. Realize [Ano23-27]. Reasoning [LMZP23]. Receiver [ZCZ^{+}23]. receptors [GGJ04]. Reciprocal [TT23]. Reclustering [YYW^{+}23]. Recognition [AXA20, AE19, Alk19, AN19, Ano23-39, Ano24, ARGC19, CZKY24, CZCL23b, CD23, DyLJ23, HLKY19, HY23, HWJ18, HZ23, HS23b, Hua23, JZS21, Jua23c, KZZ^{+}21, LWZ^{+}23a, LGJ18, LLW^{+}22, LX23a, Li23b, LXL^{+}23, Li23d, Li23a, LCH^{+}23, LL23a, LWCZ21, LLL23c, LDW23, LH23b, LJG^{+}23, LML^{+}23, Ma23b, ML23b, ML24, MCF^{+}23, NL23, PZ24, QS23, Qiu23, SDM^{+}19, Sun23a, TZH20, TZCZ23, Tan23, WDZY20, WSM^{+}21, WW23b, WWW23, Wan23b, Wan23e, XLY18, Xu23b, Xu23f, Yan23d, YY23b, YHZ^{+}23, Yu22, YWX23, Yua20, YN23, ZHL^{+}23, ZXS^{+}23, ZC20]. Recommendation [Ano23c, CCM^{+}23, CSM15, EMTCA^{+}19, FHS^{+}23, FYH23, GCB^{+}23, JYLL23, LG18, LZZ23a, LF23a, LTZS18, RBJ^{+}19, SLD^{+}16, TKZ^{+}24, TL17, WWZ17, Wan23d, YC23, ZLS^{+}17, ZW23b, Zhu23b, Zhu23a]. Recommendations [ZW18, AB96]. Recommender [DRQ^{+}18, GWH^{+}23]. Recommending [Zha23j]. reconfigurable [KMB09]. Reconstructed [Ano23-47, JHL^{+}23]. Reconstruction [Ano23r, BLL^{+}23, CLF^{+}23, CTZC24, He23, He24, JP23, LLL^{+}23a, LWL^{+}23c, LSS^{+}23, LHC^{+}23, LZWL23, LJ23, Ma23b, ML24, SZW^{+}23, SAN^{+}19, SYW23, WCF^{+}21, WLZJ24, XZLZ24, YLLY23, ZZW^{+}23, ZRX^{+}24, ZZ23, ZTL^{+}23]. Reconstruction- [SAN^{+}19]. Recovery [BGZL23, LFS^{+}17, MLAS21, WXZW17, WCR^{+}21]. Recreational [ZJZ23]. Rectal [LCL^{+}23a]. Recurrent [CCM^{+}23, FDY^{+}19, He22, LXL^{+}23, TMG^{+}19, WZS^{+}22, WZS^{+}23, XWLL23, XC23, ZL23b]. Recursive [Ano23-39, DEŁ01]. Recycling [DZ18, GLW^{+}23]. Redesign [YSSL23]. Reduce [ZWXL21]. Reduction [AMU^{+}23, Ano23-39, BKMS23, KS15, KKA^{+}22, MABU18, PWY23, SZAG15, SLH23, TF23, ZHL16, ZW23c, KHSJ95]. Redundancy [AFB23]. Redundant [LYL^{+}23]. Reentrant [CLL^{+}18]. Reference [LYL^{+}23, Nag05b, WYC20b]. Refinement [PZC^{+}23, BAN02, CW93]. Reform [Ano23z, JL23a, LGZ23, Yan23f]. Refractory [WJZC23]. Region [HCC23, WSL23, XZ23b, ZCG23]. Region-Growing [WSL23]. Region-Segmentation [ZCG23]. Regional [HCM23, JZL^{+}20, JHL^{+}23, WFZ^{+}23, ZJZ^{+}16]. regions [MGK^{+}13]. Registration [LXW24, QL23, WZL23, ZLZ^{+}23, ZC23]. Regression [CLHH20, CZY20, SKHZ24, SWH23, WPDZ19, Xu23a, Zha23e]. Regular [Ano08c, NAA^{+}03]. Regularized [YPC^{+}23]. Regulation [JZH^{+}23, KJA^{+}23, MCZ23, MTK^{+}23]. regulations [ZGW08]. Regurgitation [CTZC24]. Rehabilitation [Ano23n, Liu23c, SYW23, WCZW23, WLZJ24, YHZG23, ZCL^{+}24a, ZZ23]. Reidentification [TT23]. Reinforcement [KFU^{+}20, YNK23, PVL^{+}04]. Related [CH23b, KAA^{+}20, Li23d, PWY23, RLU^{+}20]. Relation [ABNJMB23, BZKA20, TLW^{+}23, ZHZF20]. Relational [AIA^{+}21, WWZ17, WM23b, KCO^{+}05]. Relationship [Son23a, Sun23c, SWH23, TL17, WZZ23, Zha20a, ZZH22]. Relative [LLC^{+}23, ZGY^{+}20]. Relaxation [XSZ18]. Relearning [ZLZ^{+}23]. Releasing [QYL23].

Reliability

[AAH^{+}23, AFB23, KPHY18, SMB17, TT16]. Reliability-Aware [SMB17, TT16]. Reliability-Ignorant [SMB17]. reliable [DRL13, KABW11]. Remanufacturing [ZGS16]. Remapping [KPHY18]. Remodeling [KSYC23, LWZ^{+}23c]. Remote [CTCG15, Cui23, DD23b, FCSL23, GL04, PFQ^{+}23, TZC23, Wan23i, XXM^{+}23, ZHY22, GBH14]. Remote-Sensing [PFQ^{+}23, XXM^{+}23]. Removal [BLZG23]. Removing [Ma22]. Renal [HLL23, WKW^{+}23, WLZJ24]. Rendering [Xia23b, XL23]. Renegotiation [YP24]. Rental [SWDZ24]. reordering [ZSS^{+}10]. Repair [XLLL23, YJ23]. Replacement [WCW^{+}21, WKW^{+}23]. Replacing [YP24]. Replay [MR18]. Replicas [Wu16]. Replication [LX23c, TH18]. Report [PAR94, HB94]. Reporting [YLL^{+}23]. Reports [SPH20]. Repositioning [TFZ^{+}20]. Repositories [SLD^{+}16]. Representation [GDA^{+}23, LWZ^{+}23a, LXAC21, Ma23b, Qiu23, ROMFST23, SLC^{+}23, TZF^{+}18, YWX23, Tým99]. Representations [Qiu23, TZH20]. Reproducibility [SPPH15]. Requirements [AKA16, ABK^{+}23, AA23, SDS00, SQ21]. Resampling [ZLDX21]. rescheduling [KLN^{+}04, SZ04]. Rescue [QK21, RAA23, Wan23a, ZZWZ23]. Research [Alt23, Ano22a, Ano22b, Ano23m, Ano23-41, Ano23-35, Ano23-36, Ano23-37, Ano23-42, Ano23-38, Ano23-39, Ano23-43, Ano23-40, BL23, BH23, CH23a, Che23a, CT23, CHL^{+}23, Du23b, DW23, FLS17, Fen23b, Fen23a, HSH23, HBSZ20, HLHC23, He23, Hon23, HHD^{+}16, HCQ17, HWJ18, HL23d, HW23a, HZ23, HS23b, HZ24, Hui23, LZF16, LJL23, LMZP23, Li23e, Li23a, LL23c, LF23a, LL23b, Li24, LCH^{+}24, LLLZ23, LY23b, LZ23b, LFY^{+}18, Liu20, LWW23b, LLW23b, LHZ^{+}23, LQW23, LWCJ23, MAC^{+}20, Pen23, PZ24, QM23, QFX^{+}23, SPH20, SW23, SCKW23, TH23, WQLR17, WW22, WM23b, Wan23e, WC23a, Wei23, WDY23, XCW18, Xia23a, Xia23b, XSG23, Xu23e, Xu23d, Yan23d, Yan23e, Yan23c, Yan23h, Yu22, ZYLZ15, Zen22, ZWW^{+}23a, ZLSW23, ZYL23, Zha23m, ZWF23, ZS23, ZG23, ZW23b, Zhu23a, ZD23b, AGG^{+}97, NKV^{+}02]. Resection [CLW^{+}23, LX23d]. Reservation [GPP^{+}17]. Reserved [Li23b]. Reshaping [TFN11]. Residential [Ano23-42, LY23b, Zha20a]. Residual [CLDX20, Men23, QWZP20]. Resistant [JRS^{+}21]. ResNet [KZZ^{+}21]. Resolution [Ano23-47, CLDX20, DD23b, YGZ^{+}24, ZLZ^{+}21, BAN02, VvAC^{+}09]. resolving [BAN02]. Resonance [DBWY24, DLP^{+}23, FZH23, FHC^{+}22, HLL23, HTZL23, HSY^{+}23, JFY23, LWZ^{+}23b, Liu23c, LTL^{+}23, LTR^{+}23, LYL^{+}24, LYZ^{+}23, OWL^{+}23, PTZL23, SKT^{+}24, WYHS23, WCZW23, WZZ23, Yan23b, YT22, YHZG23, ZYZ23, ZZZ^{+}23, ZCLL23, ZY23, ZCL^{+}24b, ZLZ^{+}21]. Resource [ACB^{+}02, Ano23c, Ano23-45, Don23, GWZ16, HPZ16, JFD20, JZ22, LXWD18, LL23b, MPA16, Özt04, STM^{+}23, UAK18, WW23c, WL23b, WC23b, fXhCL17, YW24, ZCSL17, ZWL^{+}19, Zha20b, ZD23a, Zhu23a, CJS^{+}02, JDVM10, KLN^{+}04, OPP11]. Resource-Constrained [Zha20b]. Resources [Ano23x, BL23, HLPY16, QTD^{+}20, SL23a, WM23b, ZHZS19, Zha23j, KVW^{+}07]. Respiratory [CL23b, LMQ^{+}24, PWY23, SZW^{+}23, Zha23a, ZYL23]. Response [CLL17, DYW^{+}23]. Responses [SCW^{+}23]. Responsible [YWH20]. Restart [CSC16]. Restoration [BL18, LLW23b, LLGH24, YY23c, ZWCY16]. Restriction [ZZH22]. results [BHK^{+}13]. Retinal [HFH^{+}23]. RetinaNet [YXL^{+}20]. Retracted [Ano19, Ano22a, Ano22b, Ano23f, Ano23h, Ano23j, Ano23g, Ano23k, Ano23m, Ano23e, Ano23l, Ano23i, Ano23o, Ano23p, Ano23s, Ano23q, Ano23a, Ano23t, Ano23r, Ano23n, Ano23v, Ano23w, Ano23x, Ano23u, Ano23y, Ano23z, Ano23c, Ano23-27, Ano23d, Ano23-30, Ano23-28, Ano23-29, Ano23-31, Ano23-32, Ano23-33, Ano23-46, Ano23-34, Ano23b, Ano23-41, Ano23-35, Ano23-36, Ano23-37, Ano23-42, Ano23-38, Ano23-39, Ano23-43, Ano23-44, Ano23-40, Ano23-45, Ano23-47, Ano23-48, Ano24, JZS21, ZX17]. RETRACTION [Che22, FC22, HT22, JZ22, Li22, Liu22, WLJ22, Wan22, WWYR22, XY21, Yan22a, ZQ22, Zha22, ZLF^{+}22]. Retrieval [BL23, DYT23, FS24, HLW^{+}18, SQW^{+}21, WX23, WHL^{+}18, Zha23b, ZD23a, ZZWL24]. Retrieval-Based [HLW^{+}18]. Retrospection [Ano23-44]. Return [SPH20]. Returns [WQLR17]. reusing [NAA^{+}03]. Reverse [HHD^{+}16]. reversing [HUN08]. Review [Ano07a, Ano07b, Ano07c, Ano11b, ACBV20, Bai08, Bor04, Bro11a, Bro11b, Bro12, Bro09, Bus09, Fil14, GAH^{+}23, GB10, HNG^{+}23, Hil97, JKA^{+}23, Les10, LXW24, Man08, Mic97, Nag04, Nag05a, Nag05b, Nag09a, Nag09b, Nag11a, Nag11b, Nag12, RMS21, Ste97, Wes08, XHL^{+}23, XTZ20]. Reviewers [Ano03, PT09b]. Reviews [KGU^{+}20, KGU^{+}21, WLZ^{+}21]. Revised [Ano19, ZX17]. RF [PLY^{+}23]. RFID [HL23b]. Rheumatoid [GLD^{+}24, Li23d]. Rhythm [ZZH22]. Rich [ZRX^{+}24]. Rigid [WBZL17]. Rigid-Fluid [WBZL17]. RISC [Hil97, Was95]. Risk [Ano19, BBI23, CH23b, CZWY23, ÇÜ19, GPX^{+}21, HW23a, JZY23, MVIAM20, OL23, SLS17, SFZ^{+}20, TMG^{+}19, WQT^{+}20, XNL^{+}22, YG16, Yan22b, ZCZ21, ZX17, ZZW22, ZZSQ23, MCvM10]. River [FYH^{+}19, HWY^{+}23, PFQ^{+}23, Xu23a]. rmvPFBAM [Ma22]. RnkHEU [XJH23]. Road [WLCC18, ZCG23, Zho20]. Robert [Per08]. Robot [KLuH^{+}20, QK21, QFX^{+}23, YNK23]. Robot-Following [YNK23]. Robust [AZX^{+}17, SK21, TXD^{+}20, WPDZ19, YWC^{+}18, ZJZ^{+}16, CM02]. Robustness [ZWW16]. ROC [ZCZ^{+}23]. Role [LJF20, LFQA20, ZZD^{+}23b]. Roller [Ano23m]. Rotating [WLQ23]. Rotation [KLK^{+}20a]. Rotation-Invariant [KLK^{+}20a]. Rough [PDROFRM16]. Route [ZCG23, Zha23l]. Router [GWZ16]. routines [CDO^{+}96]. Routing [CCG17, GLW^{+}23, HBSZ20, WQT^{+}20, YCWL17, YJ23, ZCSL17]. ROV [WYC20b]. rover [MJLM07]. Rubin [Hil97]. Rule [BBI23, SY20]. Rule-Based [BBI23]. Rules [DRMG^{+}18, ZYLZ15]. Run [EAS^{+}97, DN04, IRSD99, MSH99]. Run-time [EAS^{+}97, DN04, IRSD99, MSH99]. Runge [LPD20]. Runtime [Ano99b, UJS^{+}20, HPS^{+}13, NPP^{+}00, Pro07]. Rural [Xu23c, Zho23].

S [Ano07a, Ano23y]. SaaS [DYW^{+}23]. Sacral [WCW^{+}21]. Safe [AHM23]. Safety [LFY^{+}18, ZWW^{+}23b]. Sagittal [LZWL23]. Salary [CC23]. Sale [LTF23]. Sales [YT23, ZYZL23]. Saliency [MWZ^{+}23]. same [Mic00]. same-source [Mic00]. Sample [MKH^{+}19]. Sampling [AAH^{+}23, BC23, SKHZ24, WBZL17, WXL19, YZ16]. SAMR [LTB02]. SAR [tZhYsHlZ19, CZCL23b, LJG^{+}23, Ma23b, ML23b, ML24, Qiu23, TZH20, WZY^{+}23, YWX23]. SAT [MVIAM20]. Satellite [LJSJ23, WNL^{+}20, YGZ^{+}24]. Satisfaction [SW16, AMM05]. satisfaction-based [AMM05]. Sawzall [PDGQ05]. SC13 [GM14]. SC2000 [Old00]. Scala [PSM^{+}15]. Scalability [LPD20, SBJV11]. Scalable [BPK00, FS24, GBH20, HPZ16, HYS^{+}19, JHNP14, KGIL19, KCS15, NKV^{+}02, OHS00, XSZ18, gXySR^{+}17, CM02, DRR12, GBH14, IMB^{+}13, OS99, Ros00, Rou08a]. ScalaLab [PMTL12, PMTL14, PSM^{+}15]. ScaLAPACK [CDO^{+}96]. Scale [BRPV24, HJO^{+}20, HL23b, LZWJ23, LLC^{+}20, Liu23c, SZAG15, XHSL17, YW24, ZWW16, MJLM07, MSCS14, WGW08, Zim07]. Scale-Free [ZWW16]. SCALEA [TF04]. SCALEA-G [TF04]. Scaling [DYW^{+}23, DGL^{+}17, JUA^{+}23a, JW23a, KS19, NAA^{+}03, SD23, MGK^{+}13, SCSJ09]. Scan [ZWZ^{+}23b]. Scanning [Ano23g, WSZK09]. Scar [XLLL23]. SCASH [SHHI01]. Scattering [ZL22, HST^{+}93]. Scenario [FCSL23, WLP^{+}21]. Scenarios [CLPVPI20, PDROFRM16]. Scene [Ano23r, He23, Sun23a, YWGZ20, YB24]. Scenes [DC22]. Scenic [CL23c]. SCG [LHMC23]. schedules [NAA^{+}03]. Scheduling [CS94, CBE23, CLL^{+}18, HPZ16, HQL^{+}23, KKK^{+}18, KGMN15, LGL23, LXWD18, LWW16, LKDB10, LFS^{+}20, MFB^{+}15, MJB15, OCC^{+}16, PBY^{+}20, SGK^{+}23, SQ21, TT16, WWLG16, WCLL17, WLP^{+}21, XWL^{+}23, YZSZ17, Zha20b, ZWZ16, ZWXL21, DRR12, FB99, GR93, KLN^{+}04, SJR14, Shu94, SCB02, WD07, ARFS05, BPC^{+}09]. Schema [AIA^{+}21, VRW^{+}03]. Scheme [Ano23-42, ATZ^{+}15, DPMB16, LPD20, LL23b, Lv23, MKS^{+}24, PLW^{+}20, QTD^{+}20, Ser18, Zha20b, SJR14]. Schemes [NYML17, SMB17, SF03b, SF03a]. Schizophrenia [ZY23]. School [Liu23d, ZLSW23]. Schools [ZGL23]. Science [BR11, HMSW92, KJG^{+}08, KBG^{+}02, SC23, SGPCLG19, DRT^{+}14, Ste97, TSCT11, DRT^{+}14]. Science-Driven [BR11]. Scientific [AK18, AMVGC^{+}23, AHMMÁR18, ÁRAHMM18, Ano07c, BTV^{+}20, Che93b, DMAC23, DS97, DDMS15, DDJ20, GDNVSMM19, GHH^{+}02, LJF20, MFB^{+}15, Nag05b, NGMAA23, OPB^{+}20, PSM^{+}15, PT09a, PT09b, RLPI19, SPPH15, Sny07, SRTCL16, SGPCLG19, TPSM17, TH23, WLZ^{+}21, ZLL^{+}17, ZGL18, AJ94, ANEA08, BH12, BBŁ08, BGLR93, BÇCD12, CLM05, CHS^{+}99, DSS^{+}05, DDMS14, GG95, Gil09, Hav00a, HHML05, IJL^{+}01, JDVM10, KJG^{+}08, Mat94, NJ08, NDSG07, OSS94, PMTL12, PMTL14, Rou08b, SKU^{+}09, SLJ^{+}00, SBJV11, Szy07, VF95]. Scientist [Hil97]. Score [XHX23]. Scoring [SBK^{+}24, Zha23m]. Script [JK22, MSA^{+}20, ZLWW23, BBBM12]. Script-Based [ZLWW23]. scripting [PMTL14]. Scrum [CZS^{+}23, ZZS^{+}23]. Sculpture [ZG23]. SDC [BA18]. SDN [TZDZ20, WCF18, ZWCY16]. SDN-Based [TZDZ20, ZWCY16]. Sea [LJSJ23, LJSJ23]. Seals [LZSC23]. Search [CCG17, KAA^{+}20, LGC^{+}16, LZZ19, OQZ23, PSW23, Tun23, WHH^{+}21, yXHtJC23, XYY^{+}19, YLWW18, YL20, BAP13]. Searching [HYS^{+}19, WS15, WYY21, GA96]. Seasonal [KS19]. Seawater [BRPV24]. Second [ZKA23]. Secondary [CTC^{+}15]. Section [Ano08c, NLMT23]. Sections [TC96]. Sector [HXPZ24, SAZ^{+}24]. Secure [MAA^{+}21, SLP18, Yan23c]. Secured [HAA^{+}22]. Security [AlS16, Bak23, CGMRO^{+}19, GMS^{+}23, GBH20, HGMM23, LL23c, LZLJ23, SKA^{+}20b, WW23c, ZD23b, CQF05, SKA^{+}20a]. Security-Oriented [ZD23b]. SEDC [SLP18]. SEDC-Based [SLP18]. Seed [CLA^{+}15, MWZ^{+}23]. Seed-Based [CLA^{+}15]. SEER [AMN^{+}12]. Segmentation [CZD^{+}20, CWLW23, CEMM17, DLP^{+}23, FLR23, FZW23, Hua21, HFH^{+}23, KSYC23, LDZQ24, LZZ^{+}23b, MSA^{+}20, RJA^{+}23, SKT^{+}24, TZC23, WZS^{+}22, WZZ23, Wan23i, WJZC23, XWLL23, Yan23g, YY23b, ZCG23, ZYZ23, ZZDY24, ZHHX23, RLC04]. Segmented [SWZL24]. segments [Abd02]. SEI [Ano07a]. Seismic [CZWY23].

Selected [HKQ^{+}23, VR13]. Selecting [ZXDY19]. Selection [AK23, CLTX16, DZ18, FTS^{+}20, HY23, LY20, Liu23g, MWZ^{+}23, MKH^{+}19, QHAN23, SMA^{+}23, WYY21, XJH23, ZLF18, Zha20b, ZWG^{+}20, ZW23a, AMM05, WHRH07]. Selection-Guided [MWZ^{+}23]. Selective [AMB23, ZZH22]. Selective-Data [AMB23]. Self [GLD^{+}24, Pan23a, QQWJWYQZ23, WLL^{+}23b, WMZM23]. Self-Acceptance [QQWJWYQZ23]. Self-Adaptive [WMZM23]. Self-Attention [Pan23a]. Self-Efficacy [GLD^{+}24]. Self-Interacting [WLL^{+}23b]. Selfish [YW24]. Selling [ZYZL23]. SEM [BYZ^{+}23]. SEM-PLS [BYZ^{+}23]. Semantic [AB18, DZL23, HEVD19, KZS^{+}20, LPBSSE17, LMZP23, PMEH17, SD23, TZC23, Wan23i, YY23b, YWGZ20, MPP^{+}04, BB08, GB15]. Semantic-Based [YWGZ20]. Semantics [GMM20, NO18]. sEMG [CZKY24]. sEMG-Based [CZKY24]. Semi [KG08, IJL^{+}01]. semi-automatic [IJL^{+}01]. Semi-supervised [KG08]. Semiclosed [QZY^{+}20]. Semiseparable [YTD15]. Semisupervised [CLG^{+}23, WLQ23]. Sending [HPGC19]. Sensing [Cui23, Don23, DD23b, FCSL23, HSY^{+}23, LWZ24, PFQ^{+}23, TZC23, WLCC18, WW22, Wan23i, XXM^{+}23, ZHY22]. Sensitive [SAJ^{+}20]. Sensitivity [CO93, YWGL20]. Sensor [Ano23k, CSZ^{+}19, HLC^{+}24, LJL23, WS15, WXZW17, WLJ22, YZZ15, YYL17, YXZ23, ZM23, ZY24]. Sensor-Based [YZZ15]. Sensors [KKD^{+}20, LLY24, LZK^{+}23]. Sensory [Liu23e]. Sentence [AAS^{+}23, LY20]. Sentiment [FYZ^{+}20, GMM20, HS23c, LZ23d, PVCPdPSZVG17, QFYY20, SPH20, YZQ^{+}23, YSZ^{+}23]. Sentimental [ZW18]. Separability [TVCB23]. SEPB [HZL^{+}23]. Sepsis [GXY^{+}20]. Septal [CTZC24, LCW22]. Sequence [CCB^{+}20, NL19, YYL17, YZSZ17, Yu22, KG08, WSZK09]. sequences [KTP05, LSB12]. Sequencing [FH23, Ma22, HB12]. Sequential [LLW23b]. serial [CLM05]. Series [Ano07a, BC23, CLN^{+}24, JNK24, JGF23, Nag05b, ZWXL21, BCHL05]. Series-Reduce [ZWXL21]. server [BDLL94, GKL^{+}96]. servers [CDD^{+}05]. Service [CT23, CCM^{+}23, CL23c, CMTAC^{+}16, CZS^{+}23, DW23, Fen23a, GB15, HPGC19, HS20, JFD20, JW23b, LG18, LGJ18, SD23, SH19, TL17, WmG21, YSSL23, ZLF18, ZZS^{+}23, KFFZ05, MPP^{+}04, PHH95, WHRH07]. Service-Oriented [CZS^{+}23, ZZS^{+}23]. Services [AB18, CYJ^{+}17, HKH^{+}17, HZD17, HLY20, HS23a, MSZZ20, PVCPdPSZVG17, QYZ17, RKA^{+}24, RLU^{+}20, ZZW17, ZTLH18, AMM05, TSCT11]. Set [KKD^{+}20, PDROFRM16, QL23, YG16]. Sets [DQQ17, DYT23, Wu16]. Setting [LWCJ23]. Settlement [LHMC23, MLZ23]. Severe [WMZM23]. Severity [AKA^{+}23, LLW23a]. Sevoflurane [LLGH24]. sFas [ZZH22]. sFas/sFasL [ZZH22]. sFasL [ZZH22]. shallow [KBRS95]. shallow-water [KBRS95]. Shannon [CZ23]. Shape [CFY23]. Shaped [NLMT23]. Share [HS23c]. shared [DBVF01, HMCH07, NAA^{+}03, OB96, PBK01, SHHI01, TKS02, ZS99]. shared-memory [NAA^{+}03, ZS99]. Sharing [Ano23-45, DD23b, KL23, LJZ18, PLW^{+}20, Son23a, TFZ^{+}20, WC23b, BTS^{+}14, Ott93]. Sharpness [LCH^{+}23]. Shear [XYL^{+}23, ZCZ^{+}23]. Shifting [ZMLZ18]. Ship [CYSZ23, HZJ23, LFS^{+}20, LPBSSEPL19]. Ship-SIBISCaS [LPBSSEPL19]. Shipping [ZCW^{+}19]. Shock [SZW^{+}23]. Shoe [Ano23h]. Shooting [HX23, HLH23, LLL23c, QH23]. Shop [CLL^{+}18, LGL23, SQ21, WCLL17, YZSZ17]. Short [LYWS23, MCF^{+}23, SF23, WHH^{+}21, ZWG^{+}20]. Short-Term [LYWS23, MCF^{+}23, SF23, WHH^{+}21, ZWG^{+}20]. Shortest [CLXZ23, EBDB24, Liu23g, yXHtJC23]. SIBISCaS [LPBSSEPL19]. sided [LKDB10, GBH14]. Sieve [KK09]. SIG [SCR20]. Sign [AZGJ24, JHS^{+}20]. Signal [AE19, Che23b, WZF^{+}20, tZhYsHlZ19, ZMM^{+}23, Den96]. Signals [CZB24, Hua23, LZ23a, ADS95]. Signed [SSN15]. Significance [SSY^{+}23]. Silico [Ano23-29]. SIMD [Bra17, TSP^{+}15]. SIMD-Vectorization [Bra17]. SIMDization [IB09]. Similar [LWL23b]. Similarities [CLA^{+}15]. Similarity [AMU^{+}23, FCZ^{+}17, KGIL19, LLW23b, LLC^{+}23, PMEH17, SD23, TZF^{+}18, ZLLS20, ZLLZ20]. Similarity-Based [TZF^{+}18]. Simple [SAZ^{+}24]. Simplification [WZW19]. simplified [DN04]. Simulated [BNH^{+}20]. simulating [BDM^{+}04]. Simulation [AFU18, Ano23j, Cer21, CCBPGA15, Du23a, LZZ23a, LLY^{+}20, MCZ23, MLZ23, MVIA20, MLL23, O’K00, WQLR17, WBZL17, WL23c, WB95, Wu23b, XSZ18, ZZD23a, BBLR03, GIKP95, GGJ04, MJLM07, NKV^{+}02, OPP11, PPS12a, PPS^{+}12b, SE95, War14]. Simulations [SCD^{+}15, YDC^{+}16, KLS^{+}96, LLW^{+}14, LG03, MGK^{+}13, RX04, WCKD07]. Single [AlS16, Ano22a, Bak23, CLDX20, CSZ^{+}19, CPG^{+}23, KLL^{+}18, VMRM16, WZY^{+}23, XTCC15, AGIS94, DF14]. single- [DF14]. Single-Chip [Ano22a]. Single-Label [Bak23]. Single-Layer [KLL^{+}18]. Single-Phase-to-Three-Phase [VMRM16]. Single-Photon [CPG^{+}23]. single/multigrid [AGIS94]. Singularities [PPJ^{+}15]. Sink [WCZ^{+}16]. Sinogram [WCF^{+}21]. Site [SW20]. Sites [SQJG23, SY20, IR02]. Situ [fTPsLZ23]. Situation [ZD23b]. Situations [TMG^{+}19]. Size [Ser18, ZWZ23a]. Sized [Fen23a, Li23e, Wan23c]. Sizing [ZXX^{+}23].

SKaMPI [RST02]. Skater [Ano23m]. Sketches [AN19]. SKETRACK [AN19]. Skewed [MSA^{+}20]. Skewness [CZ23]. SKI [VC93]. Skillrank [ÁRCPS15]. Skills [ÁRCPS15, HH23b]. Skin [XLLL23]. sky [VvAC^{+}09]. SKYHI [Hem00]. SLA [ZWL^{+}19, ZHL16]. SLA-Aware [ZWL^{+}19]. SLAM [LWYS16]. SLE [WOC99]. Slicers [PST19]. Sliding [HHCC23]. Sloan [Nag05a]. Slogger [BB08]. Slope [AXWS23, SC16]. Sloping [WKY23]. Slow [ZCT^{+}20]. SLPA [KCS15]. Small [Fen23a, Li23e, Wan23c, tXChL^{+}24]. Small- [Fen23a, Li23e, Wan23c]. Smart [Alt23, BDAH20, CWD23, CL23c, FA24, Fen23a, Ge23, HCM23, HAA^{+}22, KL23, KKA^{+}22, LJL23, MAA^{+}21, PPB^{+}19, SMA^{+}23, TKKZ23, UAU^{+}23, ZCG23, Zha23h, ZLL^{+}17]. Smartphone [AAA^{+}23, KKD^{+}20]. SMC [SMB17]. Smear [GAH^{+}23]. SMEs [AUA^{+}23]. Smith [Mic97]. Smoothness [ASZF23]. SMOTE [WYC^{+}20a, WZF24, WDD^{+}23]. SMOTE-Based [WDD^{+}23]. SMP [CCBPGA15]. Snare [LCL^{+}23a]. Snir [Nag05b]. Snow [Ano23-28]. Social [ÁRCPS15, Ano23-28, Bia23, CSM15, DPMB16, FYY23, GMCT15, HKH^{+}17, KAS15, KZZ^{+}20, KZS^{+}20, KZK^{+}23, PLW^{+}20, SSN15, SLC^{+}23, TMHM23, Wan23a, YYZ^{+}17, Zen22]. Socialization [LF23b]. Society [Ma23d]. sockets [BŻvA^{+}01]. Soft [LHMC23, SHW23, WS15, XHL^{+}23, Zha23l]. Soft-Sensor [WS15]. Softcomputing [LFS^{+}17]. Software [AUA^{+}20, AKA16, Ano07a, Ano07c, AB96, ACMT19, ACBV20, ARGC19, ABK^{+}23, BBI23, CHC15, CZS^{+}23, DMAC23, DWZ^{+}20, DDJ20, DYW^{+}23, FDY^{+}19, GBJ94, GMS^{+}23, GNKuR23, GRB23, HLW^{+}18, HLS^{+}24, HPZ17, KV23, KPW^{+}18, LNH^{+}20, LQZ^{+}21, LLY^{+}20, LWW23b, MVIAM20, MOS16, OS99, O’K00, PT09a, PT09b, SX23, SAZ^{+}24, SLD^{+}16, SNH24, TLZ^{+}23, TLW^{+}23, Wes08, XNLJ23, XHL^{+}23, YYF^{+}19, YZX17, tZhYsHlZ19, ZAW^{+}21, ZGY^{+}20, ANEA08, BFH96, DHH00, GB94, Gus93, Hav00a, KJG^{+}08, LBvBW12, MS00, MOT97, PMTL12, PPS12a, PPS^{+}12b, RX04, Rou08b, SHHI01, WSB11]. Software-Defined [HPZ17, KV23]. Soil [FCSL23, LHMC23, PLAAH23, WLL23a]. Solid [BDAH20, CLN^{+}24, KPHY18]. Solid-State [CLN^{+}24, KPHY18]. Solo [RuHMS24]. Solution [AMU^{+}23, CLLZ23, CZW^{+}23, FH23, FGSCR^{+}17, HMCH07, SG96]. Solutions [DMAC23, FAA^{+}20, GDNVSMM19, AGIS94]. Solve [GMS^{+}23]. Solver [CKKK15, LPD20, MLYH17, MYH19, SM17, TLH17, DEvdV01, JR10, MSSG11, SCSJ09, SF03b, SF03a]. Solvers [Gog23, BHRT12, RR04]. Solving [BEK^{+}12, CCG17, LGL23, OZI^{+}23, SW16, SCAP16, WB16, YTD15, YZ16, YSSL23, Gus93, Lan01, MRLF12, VRW^{+}03]. Somatic [WWQ23]. Some [McC96, Lar93]. Sonar [DCF17, HLKY19]. Sorting [HQL^{+}23, SCAP16]. Source [FHT23, FXYY17, GNKuR23, HNG^{+}23, ZLLS20, ZLLZ20, ZGS16, Mic00, SGM^{+}08]. Sources [AMS17, NLPL^{+}20, ZCSL17]. South [LFY^{+}18, LJSJ23]. South-to-North [LFY^{+}18]. Southern [Ano23-28]. Space [BC23, DPMB16, MLL23, PGP19, SAN^{+}19, SV15, XML^{+}16, CBCM93, Nor07]. Spain [MVIAM20]. Spam [BKU^{+}23, FGROL^{+}16]. Spanish [PVCPdPSZVG17]. Spanning [CBRRC^{+}16]. Spare [ZGS16]. Spare-Part [ZGS16]. Spark [Hui23, JZZW20, TKZ^{+}24, XH20]. Sparrow [OQZ23, YLZ23]. Sparse [DEŁ01, LWZ^{+}23a, Li23b, Ma23b, SHW23, WG23, YPC^{+}23, ZHP^{+}23, BHRT12, CFR14, LP99, Lin04, SG96]. sparse-matrix [CFR14]. Sparsity [TTL^{+}23]. Sparsity-Motivated [TTL^{+}23]. Spatial [Cui23, DD23b, FYH^{+}19, Hua23, LDZZ23, LQW23, WZG^{+}17, Xu23c, ZCW^{+}19]. Spatial-Temporal [ZCW^{+}19]. Spatio [CD23]. Spatio-Temporal [CD23]. Spatiotemporal [YHH^{+}23, ZQ22, ZLX^{+}23]. Speaker [MZLC21]. SPEC [AE03]. Special [BFGH14, BR11, GM14, HR12a, HR12b, JK10, PT09a, PT09b, SO11, VR13]. specialization [BMY13]. Specialty [JL23a]. Specific [JBLB15, VVDM18, ZZW22, Ang93, SO93, ZGW08]. Specification [AB18, Hig93, Pro07, CMM^{+}02]. Specifications [VVDM18]. Specificity [YWGL20]. Speckle [LLY^{+}23a]. Spectra [SZX^{+}20]. Spectral [Ano23u, LXWD18, Ros00]. Spectroscopy [SZX^{+}20]. Spectrum [DD23b, LXL^{+}23]. Speech [Alk19, NL23, She23, SZ20, WZML23, Abd02]. Speed [DGL^{+}17, ZGT^{+}23]. SpeedShop [SGM^{+}08]. Spherical [NTNTTK^{+}19]. spike [PSU08, MSSG11]. Spina [CTD23]. Spinal [Ano23u, CLHH20]. Spine [WYHS23, ZYY^{+}23, KLS^{+}96]. SPINET [KLS^{+}96]. Spinocerebellar [CZY20]. Spiral [JP23, LJ23, LYZ^{+}23, OQZ23]. Spirit [WLJ22]. Spitzer [JEM07]. spline [TOF^{+}14]. Split [ZWZ20]. SPOC [Hui22, Liu20]. Spoken [Yan23h]. Sport [Zha23f]. Sport-Induced [Zha23f]. Sports [Ano23-28, Ano24, CW23, HLZ^{+}21, HLH23, JZS21, JWPW24, LLLZ23, Liu23d, LLT23, MqWp23, SL23b, Wan23e, Wu23a, Xu20, XXSL23, Xue23, Yan23b, Yan22a]. Spot [ZYZL23]. Spot-Selling [ZYZL23]. Spots [CL23c]. Sprawl [HCM23]. Spring [Wan22]. SPSA [YLZ23]. SQALE [SCR20]. SQL [JZZW20]. Square [WZL23, GG05]. Squeeze [LZZ^{+}23b]. Squeeze-and-Excitation [LZZ^{+}23b]. ST [ZWG^{+}20]. ST-LightGBM [ZWG^{+}20]. Stability [SC16]. stack [Tým99]. stack-based [Tým99]. Staff [Zha23g]. Stage [HZ23, TFZ^{+}20, ZDMY23]. Staging [HLL23, LJ23]. Standard [BKMS23, FP00b, Met99a, Met99b, RN07]. Standardization [SZX^{+}20]. Standardized [QH23]. star [BAN02]. Starting [Cer21]. State [AB18, BDH^{+}10, CLN^{+}24, CPK19, HZ24, KPHY18, PP23, PZC^{+}23, ADS95]. State-of-the-Art [CPK19, BDH^{+}10]. Statements [ZXDY19]. States [WWYR22]. Static [Den96, GG95, ACIK97]. Statically [SSSJ20]. Station [ZWG^{+}20]. Statistical [CLPVPI20, KTP05, SMD20, WNL^{+}20, Din99, OHS00]. Status [BRPV24, LH23a, Liu22]. Steel [NLMT23]. Steepest [NTNTTK^{+}19]. Stem [DCH^{+}23]. Stemming [AWB^{+}23]. Stencil [CKKK15, GA18, SRO^{+}15]. Stent [BLL^{+}23, LHC^{+}23]. Step [LPBSSEPL19, SY20, WKY23]. Stepped [Che23b]. Stepping [WBZL17]. Steps [ARAM19]. Stereotype [HLZ^{+}21]. Sterling [Mic97]. Steve [Nag05b, Bry96]. Steven [Nag05b]. Stewart [Ano07c]. STM32 [Ano22a, XSS^{+}23]. Stochastic [HPZ16, HXX16, LZZ19, TFZ^{+}20, WWW^{+}17, yXHtJC23, ZHZS19, ZWZ16]. Stock [CFL^{+}23, CSM^{+}18, GWZ23a, HS23c, JAED23, SPH20, Xia23a, YY23a]. Stocks [LYWS23]. Stokes [AMU^{+}23]. Storage [CLN^{+}24, CHZY17, GM14, HYK^{+}20, LD23, LX23c, MKS^{+}24, SAJ^{+}20, Slo12, SZXL16, Xie20, DZKS13, KABW11]. Store [LX23c]. Stowage [ZMLZ18]. Strassen [KHSJ95]. Strategies [ARGC19, EBDB24, OCC^{+}16, DBVF01, FC01, PCS99, Sch94]. Strategy [ASZF23, Ano23-43, Ano23-45, CHTZ23, FZW^{+}18, GLW^{+}23, GA18, GB15, HBSZ20, HL23b, LJZ18, PSW23, RRM^{+}15, WS15, WXZW17, WLP^{+}21, Wu16, YX23, Yan22b, YL20, ZWL^{+}19, ZYZL23, MGK^{+}13]. Stratification [XNL^{+}22]. Straw [LWL^{+}23a]. Stream [Ano23m, BA18, TZCZ23]. Streaming [FTS^{+}20, HLY^{+}17]. streams [Pla04]. Strength [AAH^{+}23, Nag04]. Stress [AAH^{+}23, KJA^{+}23, Li23f, Wan23a, ZLZ^{+}22]. Stress-Strength [AAH^{+}23]. Stricken [LLZW23]. strides [Bai95]. String [TLC15]. Stroke [AN19, Ano23n, FZH23, HSY^{+}23, Liu23c, SKT^{+}24, SWZL24, WCZW23, ZCLL23, ZLZ^{+}23, ZLZ^{+}21]. Stroke-Based [AN19]. Strong [SCSJ09, ZCXQ16]. Structural [KSYC23, LB02]. Structure [CLA^{+}15, Che23c, FYZ^{+}20, HLW^{+}18, TLW^{+}23, XWG23, YB23, Zha23g, SG96]. Structure-Driven [HLW^{+}18]. Structured [BFH96, CSZ^{+}19, HTZL23]. Structures [ZCT^{+}20, HS03, KGV97]. Student [ACA19, Ano23k, Ano23-33, Ano23b, Ano23-44, Ano23-40, Ano23-48, HZ24, Lou23, QQWJWYQZ23, SXTK18, Wan21, WLJ22, WL23a, WWL^{+}23, WDY23, XJH23, YZQ^{+}23, ZLF^{+}22]. Students [Ano23-28]. Studies [HMQS23, ZS99]. Study [AAH^{+}23, Ano23-34, Ano23-45, ACCTA^{+}15, AAF19, Chu21, DS97, EMTCA^{+}19, FYH^{+}19, HBCM94, HX23, JLX^{+}20, bLwHcY18, LZ23a, LQZ^{+}21, LZYJ20, LSY20, LHZ^{+}23, LYL^{+}24, McC96, PWP21, QHJ23, RRM^{+}15, SMA^{+}23, SMB17, Son23a, Sun23c, YHZ^{+}23, YY23c, ZW23a, ZLZ^{+}23, ZW23c, Zhu22, AJJF14, BKP99, GW11, GB94, Hav00a, MS00, OB96, SBM^{+}10]. Style [WLZ^{+}21, XL23, DHH00, Ano07c]. Sub [CCB^{+}20]. Sub-Golgi [CCB^{+}20]. Subband [AZGJ24]. Subdural [BLZG23]. Subflow [ZLCW20]. Subgrade [XLY18]. Subgraph [SJ15]. submesh [Mor94]. SubRF Seq [CCB^{+}20]. Subsidence [WZY^{+}23]. Subspace [LZLZ19]. Subspaces [LXAC21]. Substation [LML^{+}23, SZL^{+}23]. Substitution [HS20]. Subsystem [CL23a]. Subtraction [Hon24, HX23]. Subtype [LLC^{+}23]. Subway [MLZ23]. Suely [Ano07c]. Sufficient [ZXS^{+}23]. Suitable [LWL^{+}23a]. Summarization [PY17]. Summarize [BRPV24]. Summarizing [KGU^{+}21, KGU^{+}20]. Sundance [LBvBW12]. Sunway [LLY^{+}20]. Super [Ano23-47, CLDX20, LWQJ24, VR13]. Super-Resolution [CLDX20]. Supercomputer [GBK^{+}96, SH94]. supercomputers [Hil97]. Supercomputing [BCCP05]. Supernova [JMR^{+}11]. SuperPruner [LWQJ24]. Supersonic [CCBPGA15]. Supervised [GLH23, ZZWL24, KG08]. Supervision [BZKA20]. Supply [LTF23, LSY20, LWX23b, MTK^{+}23, TKKZ23, YZL23, ZJZ^{+}16]. Support [CS94, CLTX16, DBWY24, FZH23, GL23, HNG^{+}23, Hui23, JGDK23, LJF20, LD23, LCK^{+}23, LWX23a, MCB^{+}20, MZLC21, MqWp23, OSLK12, PWP21, PFQ^{+}23, RWN20, tS23b, TMY^{+}23, WPDZ19, WZS^{+}23, XNL^{+}22, XHL^{+}23, XTCC15, ZCLL23, BCC^{+}93, CHS^{+}99, EAS^{+}97, HCJ08, HMSM08, Pro07, RKR^{+}99, WSB11, WSP94]. Supported [Du24, CMM^{+}02]. Supporter [KZS^{+}20]. Supporting [CYZ^{+}20, ZA10, BBLR03]. Supports [LY23a]. Surface [LZSC23, TZZ23]. Surfactant [JLJ24]. Surfactants [Zha23a]. Surgery [DM23, FLR23, LBW^{+}24, LZWL23, PTZL23, WYHS23, WLLY23, WJS^{+}23, ZRX^{+}24, ZWLL23]. Surgical [CTZC24, WLZJ24]. Surrogate [FGSCR^{+}17]. Surveillance [SHCZ23, TZCZ23]. Survey [ÁRAHMM18, CPK19, CPD21, FLS17, HCQ17, OKHS23, Ray17, XH20, AB96, BCK07]. surveys [JEM07]. Survival [GXY^{+}20, YHJ^{+}23, AMN^{+}12]. Susceptibility [CZY^{+}23]. Suspect [WG23]. Suspicious [KZZ^{+}20]. Sustainability [CNW^{+}23]. Sustainable [KS19, MCZ23, QHJ23]. SVM [CZKY24, MCB^{+}20, WZF24, XSG23, Yan23e]. SVM-Based [WZF24]. SVR [CYSZ23, PLJ^{+}23]. Swapping [HLZ^{+}21]. Swarm [BNH^{+}20, CZCL23a, CZWY23, CW23, FLY^{+}23, GKK^{+}20, HI18, JAED23, LCW22, LZSC23, Sah20, TH18, WMZM23, Ye23, YLZ23, ZLZ^{+}22, ZW23a, ZFFP24, ZQL23]. Swarms [AMT20]. Sweep [COB^{+}00, YPM^{+}21, PDA^{+}08]. SWIAS [BRPV24]. Switches [WCF18]. Switching [WS15]. SWOT [TH23]. SX [DTV00]. SX-4M [DTV00]. symbolic [SF03b, SF03a, VF95]. Symmetric [BDV03]. Symposium [SO11]. Symptom [CZY20, DQQ17]. Symptoms [LWCZ21]. Synchronization [MM16]. Syncope [ZZW^{+}23]. Syndrome [CL23b, JLJ24, LMQ^{+}24, LYL^{+}24, SZW^{+}23, WZL23, XNL^{+}22, Zha23a]. Syntactic [TLW^{+}23]. Syntax [FXYY17]. Synthesis [ZL23b, dBMT20]. Synthetic [CGA17, LL23a, tZhYsHlZ19, PTS^{+}13]. System [AlS16, Ano23f, Ano23a, Ano23r, Ano23x, Ano23y, Ano23c, Ano23-33, Ano23-36, Ano23-37, Ano23-40, BBI23, CLN^{+}24, Cao23a, Che22, CH23a, CHL^{+}23, DRQ^{+}18, Du24, FMA16, FYH^{+}19, FYH23, GWH^{+}23, HBR^{+}20, HZJ23, HNG^{+}23, He20, HQL^{+}23, HT22, HL23c, HS23a, HX23, HYS^{+}19, HS23b, JL23a, KL23, KKA^{+}22, KJA^{+}23, LPBSSE17, LWT18, LJF20, Li22, LTF23, LD23, LZZ23a, Li23f, LL23c, LLGL20, Liu23f, LWW23a, LL23d, Liu23b, LGZ23, LYZY24, MZLC21, MKS^{+}24, MAA^{+}21, MAS^{+}23, MBBDP17, PFQ^{+}23, RAA23, SZC16, Son20, Son23b, SCD21, Sun23b, TFZ^{+}20, TMHM23, TMY^{+}23, WLJ22, WW22, Wan23d, Was95, WJS^{+}23, Wu23a, XCW18, XNL^{+}22, XZ23a, XHL^{+}23, XL23, XSS^{+}23, XYCC23, YTD15, YLL^{+}23, YZZ15, YYW^{+}23, YSSL23, YN23, tZhYsHlZ19, ZXX^{+}23, ZGS16, Zho22, ZZW22, ZLF^{+}22, Zhu23b, ZKA23, BB08, CO93, CJS^{+}02, CHS^{+}99, FP95, IMB^{+}13, KMR^{+}97, KZRR94, KABW11, KLN^{+}04, Lar93, MSSG11]. system [MAG^{+}07, SCSJ09, SHHI01, TF04, WMR^{+}94, WSP94, XNQF04, Ros00].

System-of-Systems [RAA23].

System/6000 [Was95]. Systematic

[ACBV20, EMTCA^{+}19, GAH^{+}23, JKA^{+}23, RMS21, SMA^{+}23, XHL^{+}23]. Systems [BD16, BL18, CPD21, DGL^{+}17, EMTCA^{+}19, GBJ94, GCB^{+}23, GPP^{+}17, GBH20, Guo23b, HXX16, Kuf15, LX23c, MPA16, MJB15, PT09a, PT09b, RKA^{+}24, RBJ^{+}19, RuHMS24, RAA23, TT16, Xie20, YZZ15, ZCG23, ZG23, ZZJ^{+}24, ADLN08, Bai95, BHRT12, BTS^{+}14, BPK00, DSS^{+}05, DBVF01, DZKS13, DDMS14, LTB02, OS99, PVL^{+}04, SZ04, Sch94, STD^{+}14, SG96, WMR^{+}94, WGW08, Zhe10]. systolic [CIN^{+}96].

T [FP00a]. Table [WCF18, WSB11]. TACO [SJR14]. Tactical [FWX23]. Tag [HL23b]. Tai [Ano23e]. TaihuLight [LLY^{+}20]. Tailings [MLAS21]. Take [WG22]. Taking [Ano23h, LFS^{+}20, Son23a, YLW21]. Talent [Zen22]. Talents [Wei23]. TAME [SAE^{+}23]. TAP [WLLY23]. Target [CSZ^{+}19, CZCL23b, HY23, LL23c, LL23a, LJG^{+}23, Luo23, Ma23b, ML23b, ML24, MKH^{+}19, Qiu23, TZH20, WH20, WW22, XZ23b, YWX23, ZHP^{+}23, ZLL20]. Targeted [ZCL^{+}24b]. Targeting [ZMM^{+}23]. Targets [Ano23-29]. Task [CLYL16, HLKY19, HKQ^{+}23, LWW16, MJB15, SGK^{+}23, WWLG16, WLP^{+}21, ZWXL21, OdSSP13, RR07, RKR^{+}99, RRV09]. Task-Based [SGK^{+}23]. task-level [RRV09]. Tasks [ZWXL21, BS01, DRR12, PKE^{+}10]. Taxation [Lin23a]. Taxi [LZYJ20, YY23b]. Taxonomy [SKA^{+}20a, SKA^{+}20b]. TCM [WZL23]. Teachers [ZGL23]. Teaching [Ano23l, Ano23a, Ano23x, Ano23y, Ano23z, Ano23c, Ano23-27, Ano23d, Ano23-32, Ano23-36, Ano23-37, Ano23-43, BL23, Bai23, Cha23a, CH23a, Dai23, DDMS15, DS23, Du24, Han23, Hui22, JL23a, JL23b, LLW^{+}22, Li22, Liu20, LGZ23, LL24, Lv23, MqWp23, QM23, QYW^{+}17, Tia23, WH23, Wan23h, WZ23b, WWYR22, WDY23, XY23a, XY21, XY22, XHC^{+}23, XXSL23, Yan23f, YN23, Zha22, ZL23b, Zha23g, ZLSW23, ZD23a, yZcR23]. Team [LHZ^{+}23]. Teams [Ano07a]. Technical [FWX23, SCR20]. Technique [HJ15, LLL^{+}23a, MSA^{+}20, MOS16, SGK^{+}23, UJS^{+}20, VMRM16]. Techniques [AlS16, AHMMÁR18, ÁRAHMM18, BC23, CPK19, GMS^{+}23, GN22, HNSG23, SDM^{+}19, ZRX^{+}24, Abd02, BMY13, BPC^{+}09, KJG^{+}08, MS00, SKS01]. Technological [TH23]. Technologies [CCHJ23, RWY^{+}17, ZD23b, BB08, Mic97, vDKH01]. Technology [Ano23h, Ano23g, Ano23-27, Ano23-41, Ano23-36, Ano24, ARGC19, Bai23, CZD^{+}20, Dai23, DHW^{+}23, Don23, Du23b, Du24, FZW23, FC22, FYH23, HLY20, HL23c, HCQ17, HH23b, HXPZ24, JZS21, JGDK23, LH23a, LF23a, LWZ24, Lia23b, LY23b, Liu23e, LWW23a, LWX23a, Lou23, LML^{+}23, Lv23, Lvp23, Pan23a, QS23, QH23, RVJ^{+}23, SZL^{+}23, Son20, Sun23a, SWDZ24, WW22, Wan23g, WL23c, Wei23, WWYR22, XZ22, XWG23, XYCC23, XZ23c, YB23, ZQ22, ZL23b, ZXZZ23, ZJ23, Zha23j, ZZD23a, yZcR23, ZLF^{+}22, BGH99, OSS94, VCT05]. Temperature [WLL23a, WJS^{+}23]. Template [BMY13, Bra17, COB^{+}00, LZZ19, PPS12a, CP03]. Template-based [PPS12a]. Templates [MWQ16]. Temporal [CD23, Cui23, FSS^{+}17, LPD20, LZC^{+}24, ZCW^{+}19, PSU08]. ten [SN02]. Tensor [KHSJ95, LTZS18, LX23d, LWZ^{+}23c, WWW^{+}17, Lan03]. Terabits [KKK^{+}18]. TeraGyroid [BCCP05]. Terahertz [LXL^{+}23, SZX^{+}20]. Term [LYWS23, MCF^{+}23, SF23, WHH^{+}21, ZWG^{+}20]. Terminal [Ano23o, YX23, ZZD23a]. Terrain [WXZW17]. Terrorist [Pan23b]. Test [SBK^{+}24, SZWZ23, CIN^{+}96]. Testing [AUA^{+}20, DWZ^{+}20, PST19, yZcR23]. Tests [BNH^{+}20, Was95]. Text [AMH23, AK23, Ano23q, AWB^{+}23, AAS^{+}23, AZX^{+}17, DC22, DD23a, Li23a, LDZQ24, LZ23d, LLY^{+}23b, ML23a, NLPL^{+}20, PRW23, SO18, Zha23d, Zha23h, TFN11]. Text-to-Image [ML23a]. Texts [GMM20, YZQ^{+}23]. Textual [CLLZ23, CYJ^{+}20, ROMFST23]. Texture [Xia23b, ZWS19, ZCS23]. thaliana [SQJG23]. Their [GMCT15]. them [TSCT11]. Thematic [Ano23l]. theoretical [SH94]. Theory [Guo23a, Liu23f, MCZ23, PDROFRM16, Wan21, ZCZ21, ZZW22]. Therapeutic [LTL^{+}23, ZCLL23]. Therapy [QQWJWYQZ23, SYW23, WKW^{+}23, ZY23, ZCL^{+}24b, ZZ23]. Things [LJF20, CL23a, Che22, CH23a, CHL^{+}23, DZL^{+}24, GN22, HNG^{+}23, HL23b, HT22, KKA^{+}22, LWZ24, Lia23b, LFQA20, Liu23f, LWX23a, LZK^{+}23, OKHS23, Ray17, SL23b, WCZ^{+}16, WJCZ21, Yan23c, ZZD23a, ZLF^{+}22]. Things-Based [KKA^{+}22]. Thinking [JL23b, Jua23b, SW23]. Thomas [Mic97]. Thoracic [CLHH20, FLR23]. Thoracolumbar [HZL^{+}23]. Thoracoscopic [LLL^{+}23a]. Thread [KKK^{+}18, HPD09]. threads [BS01]. Threat [HLZ^{+}21]. Three [CTD23, DHW^{+}23, EBDB24, HSS23, JHL^{+}23, LLL^{+}23a, LLQ^{+}23, LZWL23, MTK^{+}23, QZY^{+}20, VMRM16, WZW19, WFHC20, WW23b, WL23c, WLZJ24, YB23, ZXZZ23, ZZW^{+}23, ZH19, KBRS95, Mar05]. Three-Axis [MTK^{+}23]. Three-Dimensional [CTD23, DHW^{+}23, JHL^{+}23, LLL^{+}23a, LZWL23, QZY^{+}20, WZW19, WFHC20, WW23b, WL23c, WLZJ24, YB23, ZXZZ23, ZZW^{+}23, HSS23, LLQ^{+}23, KBRS95]. Three-Way [ZH19]. Threshold [SLH23]. Thrombolysis [HSY^{+}23]. Thrombosis [CH23b]. Thrust [QK21, Sah20]. Thyroglobulin [XYL^{+}23]. Thyroid [CPG^{+}23, WZS^{+}22, WZS^{+}23, XYL^{+}23, XWLL23, YLL^{+}23]. Tianjin [ZFFP24]. Tibial [FHC^{+}22, OWL^{+}23, XLB^{+}23]. Tie [SLS17]. Ties [KZZ^{+}20]. Tile [XXM^{+}23, LKDB10]. Tile-Level [XXM^{+}23]. Tillage [LWL^{+}23a]. Time [BDAH20, BD16, BC23, CLN^{+}24, CCG17, CZH24, DYW^{+}23, GLW^{+}23, GMS^{+}23, GLH23, GXY^{+}20, HXX16, JNK24, JGF23, PZC^{+}23, She23, SLBZ16, TMHM23, TH18, WBZL17, WM23b, XYL^{+}23, yXHtJC23, XSZ18, YYF^{+}19, Zha23l, ZCZ^{+}23, ZCT^{+}20, BCHL05, CGL08, DN04, EAS^{+}97, IRSD99, MAG^{+}07, MSH99, OdSSP13, ZGW08, ZA10, ARFS05]. time-course [ZGW08]. Time-Critical [GXY^{+}20, ZA10]. Time-Dependent [yXHtJC23]. time-series [BCHL05]. Time-Varying [ZCT^{+}20]. Tiny [TMY^{+}23]. Tissue [CSU^{+}17, WWH^{+}23]. Tolerance [SLP18, ZA10]. Tolerant [HXX16]. Toll [ZGT^{+}23]. Tomb [LLW23b]. Tomography [Ano23n, BLL^{+}23, CWMZ23, CLF^{+}23, CPG^{+}23, DM23, GGSL23, LWL^{+}23c, LHC^{+}23, LJ23, NHHW23, SYW23, SWH23, TY23, TYCL23, WCF^{+}21, WWH^{+}23, WZS^{+}23, WLZJ24, WJZC23, XZLZ24, ZZZ^{+}23, ZSZ^{+}23b, ZZ23, ZTL^{+}23, ZHHX23, ZWLL23, DRL13, SCB02]. Tool [ACBV20, BRPV24, CYZ^{+}20, CLPVPI20, DMV^{+}19, MVIA20, MLAS21, WSB11, AGG^{+}97, BKK^{+}11, CKS^{+}09, GRR^{+}03, SLJ^{+}00]. Tool-Based [ACBV20]. Toolkit [YZZ15, IJL^{+}01]. toolkits [BÇCD12]. Tools [AMVGC^{+}23, CPK19, KJG^{+}08, ROMFST23, STD^{+}14, WJCZ21, XZ22, AB96, Met99a, Met99b, MOT97, RLL^{+}02]. top [SN02]. Topic [DD23a, DPMB16, GLH23, GXY^{+}20, LY20, SLC^{+}23]. Topology [CSWW23, HZFH23, WZ23b]. Topp [AAH^{+}23]. TOPSIS [JZY23]. Total [YCLK23]. Touch [CYJ^{+}20]. Tourism [CL23c, LF23a, LQW23, Wan22, Xu23a, Xu23c, Zho23]. Tourist [FS24]. Toxicity [WDZY20]. Tpetra [BH12]. trace [CGL08].

Traceability [CHTZ23, KL23]. Track [CYSZ23, XC23]. Tracking [Ano23m, LLY^{+}23a, LZLZ19, Luo23, Qin23, SWDZ24, WH20, CBCM93]. Trade [Ano22b, HH23a, Li24, Yi23]. tradeoff [PMCF94]. tradeoffs [AJJF14]. Trading [CSM^{+}18]. Traditional [Jua23c, Liu22, WG21, WG24, Xu23b]. Traffic [CGC16, FYC^{+}20, JLX^{+}20, LJZ18, WZF^{+}20, XBG17, Yan23c, ZCW^{+}19, ZL23a, BDM^{+}04]. Traffic-Aware [CGC16]. Train [Ano23-44, ZCL^{+}20]. Training [AMB23, Che23a, Guo23a, HI18, JWPW24, LGZG23, LLY24, LLT23, PCYZ18, QH23, SKHZ24, Wan23e, Xu20]. Trait [FZ23]. Trajectory [Ano23m, Luo23, Qin23]. Trampoline [LDW23]. Transaction [HH23a, Yi23]. Transactional [YZX17]. Transesophageal [CTZC24]. Transfacet [CLHH20]. Transfer [AHM23, CZB24, CWSF23, JHS^{+}20, LFY^{+}18, She23, ZWG^{+}20]. Transfers [KKK^{+}18, THS^{+}15]. Transform [ARGC19, CTS23, ZSD22, TVCB23]. Transformation [AIA^{+}21, AMU^{+}23, GNKuR23, HNG^{+}23, Liu23c, ZZD^{+}23b, Lin04]. transformations [LKDB10, RR04]. Transformer [Hua23]. Transient [Tým99]. Transit [Zha20a, ZWG^{+}20]. transition [BBLR03, FP95, NDSG07]. Translation [Ano23-38, BQR^{+}20, Pan23a, ARFS05, AJJF14, OPE^{+}95]. translator [OPE^{+}95]. Transliterate [JK22]. Transmission [CWS^{+}24, Li23f, YXW^{+}22, Yan23c, Zha23k]. transparent [KABW11, NPP^{+}00]. Transport [MCZ23, ZWCY16]. Transportation [CCHJ23, HWY^{+}23, HYK^{+}20, HT22, HXPZ24, LWZ24, Liu23g, Zha23l]. Transvaginal [HSS23, LLQ^{+}23, WZ23d]. Transversus [WLLY23]. Traumatic [Ano23i, WMZM23]. Treating [Ano23v].

Treatment [Ano23-29, DM23, FLY^{+}23, HZL^{+}23, HSL^{+}23, LCW22, LCK^{+}23, LTR^{+}23, OWL^{+}23, QYL23, TY23, WZS^{+}22, WLZJ24, XWLL23, YHZG23, ZRX^{+}24]. Tree [ACA19, BH23, CBRRC^{+}16, FXYY17, SW20, SLDL17, UJS^{+}20, yXHtJC23, XSG23, War14]. Trees [PPJ^{+}15, QGS23]. Trend [CW23, GZD23, Xia23a, YY23a]. Trends [LYWS23]. Triadic [KZK^{+}23]. Triangulation [LZ23c]. Tricuspid [CTZC24]. Trigger [JFY23]. Triggered [HXX16]. Trilinos [HW12, HR12a, HR12b, OSLK12, Spo12]. Trios [OSLK12]. TRIZ [Liu23f]. Trochanteric [KYR^{+}23]. Truck [LLGL20]. true [FO96]. Trust [AA23, BYZ^{+}23, BL18, DPMB16, JRS^{+}21, SX23, SFZ^{+}20, TL17, ZLF18]. Trusted [HHD^{+}16, SAE^{+}23, SCD21]. TSP [Ano07a]. Tuberculosis [CLHH20]. Tumor [CH23b, HLL23, HHZ^{+}23, LX23d, SSY^{+}23, WSL23, YHJ^{+}23, ZHHX23]. Tuning [GA18, GRB23, BFGH14, CMM^{+}02, DF14, FML^{+}14, HCJ08, MSCS14, STD^{+}14, SCB02]. Tunnel [MLZ23, Wu23b]. turbulence [RX04]. tutorials [Met99a, Met99b]. Twin [JGDK23, XYCC23, ZZH22]. Twin-to-Twin [ZZH22]. Twitter [KZS^{+}20]. Two [ACLS16, DDB^{+}18, TFZ^{+}20, TC96, WKY23, ZDMY23, KMB09, Kok07, LKDB10, OB96]. Two-Dimensional [DDB^{+}18, Kok07]. Two-Phase [ACLS16, TC96]. two-point [KMB09]. two-sided [LKDB10]. Two-Stage [TFZ^{+}20, ZDMY23]. Two-Step [WKY23]. TXNIP [ZMM^{+}23]. Type [Ano23-34, CZY20, RMX05, Rou08b]. Types [CCB^{+}20, WZL23].

U [FZW23, HFH^{+}23, LZZ^{+}23b, WOC99, Yan23g, ZZDY24]. U-Net [FZW23, HFH^{+}23, LZZ^{+}23b, WOC99, Yan23g, ZZDY24]. U-Net/SLE [WOC99]. UAV [HHCC23]. UAVs [AMT20].

Ubiquitin [SQJG23]. ubiquitous [ASCH^{+}07, ACB^{+}02]. Ultrasonic [JLJ24, WZ23d]. Ultrasonography [Zha23a]. Ultrasound [Ano23-31, Ano23-47, CTD23, CLW^{+}23, CH23b, DHW^{+}23, GSL23, HCC23, HSS23, HHZ^{+}23, LCW22, LLQ^{+}23, LX23b, LCK^{+}23, LBW^{+}24, LLY^{+}23a, LWZC21, LLW23a, LMQ^{+}24, LLGH24, QYL23, SLH23, WCW^{+}21, WHL^{+}21, WZS^{+}22, Wan23f, WZC^{+}23, XWLL23, XLLL23, YPY23, YLLY23, ZLZ^{+}22, ZLHZ23, ZXZZ23, Zho22, ZZH22]. Umbilical [ZZH22]. UML [NCMF15, NL19]. Unbalanced [ZQL23]. Uncertain [DQQ17, LRX^{+}17, LSY20, YG16, ZXHQ17, ZZW22]. Uncertainties [ZJZ^{+}16, ZHZS19]. Uncertainty [SZC16, GW11]. Unconsciousness [YWGL20]. Undergoing [BGZL23, YLLY23]. Underground [VVNHTHNT20]. Underpass [Wu23b]. Undersampling [Ona19]. Understanding [Ano07b, DFT^{+}23, Yan23h]. Underwater [DCF17, WYC20b]. Underwent [ZWLL23]. Unfavorable [Bai95]. UNICORE [Rom02]. unified [SZ09, TF04]. Unique [HCGGSS17]. United [WWYR22]. Units [FSLG17, BB09]. Universities [Ano23-46, Han23, Hon23, Liu22, LGZ23, WZ23b, yZcR23]. University [Ano23e, Ano23-28, Ano23b, CC23, DW23, LZ23b, Son20, Son23b, Zha23g]. Unix [Hil97]. Unmanned [YLZ23, ZDMY23]. Unmeasured [HMQS23]. Unrestricted [Yua20]. Unsteady [MLYH17]. Unstructured [PMEH17, SM17, SO15, SCSJ09]. UPC [LLP^{+}19, Mar05, Zhe10]. Update [FO96]. Update-in-place [FO96]. Updated [LL23a]. Updating [SCZ18]. UPPAAL [SMB17]. UPPAAL-SMC [SMB17]. Urban [CZWY23, FYH^{+}19, GLW^{+}23, HCM23, Li23c, LJZ18, LLZW23, PZH23, WZF^{+}20, YB23, YSSL23, Zha20a, ZJZ23, Zho23]. Urdu [HNMB19, AAS^{+}23]. Urdu-English [HNMB19]. Urinary [LCK^{+}23, ZLZ^{+}22]. URL [LDZQ24]. Usability [AAA^{+}23, GNKuR23]. Usage [AAA^{+}23, CM21]. Use [ARGC19, AWB^{+}23, GW11, SY20, TH23, WJCZ21, BH12, HS03, MPP^{+}04]. Used [LH23a]. User [COB^{+}00, CGA17, DPMB16, FHS^{+}23, HCGGSS17, LNH^{+}20, PLW^{+}20, SJD20, XST^{+}16, YC23, Hav00b, Hem00, WOC99]. User-Controlled [CGA17]. user-customizable [WOC99]. User-Customized [XST^{+}16]. User-Defined [PLW^{+}20]. user-friendly [Hem00]. User-Generated [FHS^{+}23]. user-implemented [Hav00b]. User-level [COB^{+}00]. Users [AAA^{+}23, GKK^{+}20, SFZ^{+}20, ZG15]. Using [AMS17, ASS23, AHM23, Alk19, AAH^{+}23, AFB23, Ano23u, Ano23-27, Ano23-31, AAS^{+}23, BD16, BZKA20, BC23, CIU20, Cer21, CTCG15, CS94, CSZ^{+}19, CLL^{+}20, qCtWfPbZ23, CSM15, CKKK15, CCBPGA15, CPD21, DBWY24, DS97, DRMG^{+}18, DZ18, DCF17, Don23, EBDB24, FLR23, FC22, FCSL23, GAH^{+}23, GMS^{+}23, GL23, GNKuR23, GMM20, HJ15, HBR^{+}20, HWSL22, HAW23, HAA^{+}22, He22, HPGC19, JBLB15, JNK24, JMGH21, KZZ^{+}20, KFU^{+}20, KZS^{+}20, KKD^{+}20, KAA^{+}23, KF20, LPBSSE17, Li23b, Lia23b, LFQA20, LCL^{+}23a, LWZC21, Liu22, Liu23b, LLW23a, LLY^{+}23b, LYZ^{+}23, MLYH17, MYH19, MZLC21, Ma23b, MTK^{+}23, MWQ16, MqWp23, NTNTTK^{+}19, NLMT23, NLPL^{+}20, OPP11, PSW23, PMEH17, PBY^{+}20, QZY^{+}20, QHAN23, RRM^{+}15, RKA^{+}24, RLU^{+}20, RVJ^{+}23, RBJ^{+}19, STM^{+}23, SKB^{+}24, SDM^{+}19, SBK^{+}24, SGK^{+}23, SAJ^{+}20, Sch94, SCD^{+}15, SJ15, SAZ^{+}24, SSSJ20, SMB17, SYW23]. Using [SK21, SZC16, SLDL17, SL23b, Tan23, TKZ^{+}24, TTL^{+}23, TKKZ23, WCLL17, WQLR17, WXL19, WSM^{+}21, WHH^{+}21, WZS^{+}22, Wan23f, WYHS23, WTY23, Wen23, WJZC23, XNL^{+}22, XWG23, XHC^{+}23, tXChL^{+}24, XBG17, XTCC15, XYY^{+}20, YZQ^{+}23, YYY23, Yan23g, YHJ^{+}23, Yi23, ZLLZ20, ZYY^{+}23, Zha23b, ZC20, ZZH22, ZWZ^{+}23b, AMN^{+}12, ADS95, AGIS94, BAN02, CO93, DRL13, FSV14, HMCH07, IJL^{+}01, IMB^{+}13, JKR92, JJY^{+}03, JR10, JDVM10, KK11, LSB12, LLW^{+}14, LKDB10, MJLM07, PDA^{+}08, PKE^{+}10, SBM^{+}10, VvAC^{+}09, WHRH07, WCKD07, WC96]. USLE [FCSL23]. Utilization [QZY^{+}20]. utilized [KJG^{+}08]. Utilizing [ZAW^{+}21].

V [CNW^{+}23]. V2X [XWL^{+}23]. Validation [AFB23, BDV03]. Valuation [ZJZ23]. Value [Ano23i, HMQS23, HH23b, LSS^{+}23, LF23b, LTR^{+}23, LWX23b, SSY^{+}23, Yan23b, YLL^{+}23, YZ16, YPY23, ZJZ23, Zho22, ZLZ^{+}21]. Values [WCF^{+}21, ZH19]. Variability [DGL^{+}17]. Variable [LWZ^{+}23c, RQD^{+}20, Sch03a, Tým99]. Variables [LYL^{+}23]. Variance [HY15, WZC^{+}23, HB12]. Variation [LYWS23]. Variational [ASZF23, GLH23]. Variations [ZH19]. Variety [WQT^{+}20]. Varying [ZCT^{+}20]. Vascular [ZLZ^{+}21]. Vector [BCHL05, CLTX16, DBWY24, EBDB24, FZH23, GL23, Hui23, LCK^{+}23, MCB^{+}20, MZLC21, MqWp23, PWP21, WPDZ19, WZS^{+}23, ZHL^{+}23, ZCLL23]. Vectorization [FHT23, TSP^{+}15, Bra17]. Vectorized [Kok07]. Vectors [JNK24, Hil97]. Vegetation [SCW^{+}23]. VEGFA [Ano23-29]. Vehicle [CCG17, HT22, TT23, WQT^{+}20, XWL^{+}23, YX23, YLZ23, Zha23l, ZXX^{+}23]. Vehicles [Ano23w, JRS^{+}21, PSW23, WLCC18, ZDMY23]. Vehicular [RKA^{+}24]. Vein [He24]. Vena [CZFX23]. Ventilation [SZW^{+}23]. Ventricle [YXL^{+}20]. Ventricular [LCW22, LWZ^{+}23c, ZSZ^{+}23a].

Venture [YLW21, YP24]. Verification [HHD^{+}16, KLK^{+}20a, MKS^{+}24, SSSJ20, tZhYsHlZ19]. Verifying [Ano23-29]. version [Ros00, Met99a, Met99b]. Versions [IPST18]. versus [OB96, YYF^{+}19]. Vertebral [HZL^{+}23]. Vessel [HFH^{+}23, XC23]. VFC [Ben99]. via [BC23, FDY^{+}19, JHS^{+}20, KLK^{+}20b, LLL^{+}23b, LTZS18, LLZW23, LYWS23, LWQJ24, MWZ^{+}23, PLY^{+}23, SKHZ24, TZF^{+}18, WTY23, WWW^{+}17, ZXS^{+}23]. Video [Ano23m, Ano23-41, ATZ^{+}15, BKMS23, DC22, DZL23, HLH23, LWW23a, MqWp23, PZ24, QS23, QH23, TZCZ23, WH20, Xu20, Yan23a, YZ20, ZWS19]. Videos [Li22, Ma23c]. Vienna [Ben99, CMZ92]. View [AHM23, FLS17, Wu16, GKL^{+}96]. Violation [YZ20, ZHL16]. Violations [YY23b]. Virtual [CGC16, Che22, GWZ16, KFFZ05, LWW16, LLS^{+}20, SGK^{+}23, WC23a, XST^{+}16, YB24, ZAW^{+}21, ZHL16, ZZW22, ABB^{+}02, Bry96, WOC99, AGIS94]. Virtualization [NsP16, PGP19]. Virtualized [AK18, HLPY16]. Virus [ZCZ^{+}23, LSB12]. Visible [LLL^{+}23b]. Vision [JWPW24, LW23, LH23b, RVJ^{+}23, Sun23b, XLF^{+}24, Yan23c, ZZSQ23]. Visual [DRQ^{+}18, Ray17, TXD^{+}20, Xia23b, XXM^{+}23, Yua20, Zha23k, ZC23, Zhu22]. Visualization [DDB^{+}18, FYY23, FC22, Hua21, KGV97, Lin23b, XZ22, Zhu22]. Vitality [PZH23]. Vitro [Ano23-29]. VLAM [ABB^{+}02, KVW^{+}07]. VLAM-G [ABB^{+}02, KVW^{+}07]. VLIW [YZZ15]. VM [FZW^{+}18, JFPL16]. Voice [SL23b, Zha23b]. VoIP [CMTAC^{+}16]. Vol [Nag05b]. volatile [JIC^{+}14]. Volleyball [Guo23a, Qin23, Tan23]. Volume [Ano99a, Ano00, Ano08a, Ano09, Ano11a, Ano12, Ano14, Ano23i, JP23, LBW^{+}24, LZ23c, WWH^{+}23, Xu23f, ZCW^{+}19, Ano97a]. volumes [Ano03]. Voronoi [YYL17]. VPP700 [DTV00]. VR [Ano23f, Ano23-41, GWZ16, LH23a]. VR-Cluster [GWZ16]. VRP [SLBZ16]. Vulnerabilities [DWZ^{+}20]. Vulnerability [Ano22a, BRPV24, CZWY23, YJ23]. VVC [ZCS23, ZWZ20].

W [Ste97]. Wait [LGL23]. Walk [LLC^{+}23]. Walker [Nag05b]. Warehouse [FSS^{+}17, Liu23f]. Warning [Li23f, PLJ^{+}23, RAA23, TMG^{+}19, ZCZ21]. Waste [BDAH20, GLW^{+}23, HQL^{+}23]. WASTK [FXYY17]. Water [AMVGC^{+}23, HLHC23, Lia23a, LFY^{+}18, MLAS21, QK21, UAU^{+}23, VVNHTHNT20, ZHZS19, ZL22, KBRS95]. Watermarking [AHM23, XZZY20]. Watts [Ano07a]. Wave [XYL^{+}23, ZCZ^{+}23, ZL22, Mal01]. Wavelet [OWL^{+}23, SLH23, WZML23, XLLL23, RLU^{+}20]. Wavelet-OFDM [RLU^{+}20]. Way [DQQ17, ZH19]. WBIETS [Du24]. WCO [ZJZ^{+}16]. WCO-for-Biodiesel [ZJZ^{+}16]. Weak [MWZ^{+}23, ZLS^{+}17]. Weakly [ZZWL24]. Wealth [YC23]. Wearable [FSLG17]. Weather [ARAM19, MAC^{+}20, O’K00, RAdASGR^{+}19, WHH^{+}21, MGK^{+}13, Mal01, WCG95]. Web [AB18, He20, HZD17, KAA^{+}20, QYZ17, SD23, AAF19, BB08, FGROL^{+}16, GB15, IR02]. Web-Based [AAF19]. Weber [AHM23]. Website [JW23b]. Weight [CZP^{+}24, FYZ^{+}20, HY15, KLL^{+}18, LWCZ21, ZCZ21]. Weighted [DYYW24, FXYY17, GL23, HTZL23, LYL^{+}24, MWQ16, MABU18, WZF24, XY23a, Xu23c, Yan22b, YPC^{+}23, ZFX17, ZGY^{+}20, ZZJ^{+}24]. Weighted-ELM [ZFX17]. Weighting [YWX23]. Well [NO18, YLLL20]. Well-Founded [NO18]. Western [XZLZ24]. Wetsuit [WC23a]. Whale [Ton20]. White [KSYC23]. Whitefly [SKB^{+}24]. Whitening [ZCZ21]. Who [ZWLL23]. Willingness [Ano23-48]. Window [ZCT^{+}20]. Windows [CCG17, GLW^{+}23, SLBZ16, Zha23l, HLM^{+}11]. Wireless [Ano23x, Ano23-36, Don23, HLC^{+}24, LJL23, Li23f, Lou23, WXZW17, Wan23g, WWYR22, WC23b, XSS^{+}23, YYL17, YXZ23, Yan23e, ZM23, ZY24]. Wise [TPSM17]. Withdrawal [PP23]. within [MRLF12, tZhYsHlZ19]. WITMED [LWQ^{+}23]. WNN [LHMC23]. WNN-Based [LHMC23]. Wolf [LGL23]. Word [WTY23, XXM^{+}23, ZHL^{+}23]. Words [LWCZ21]. work [OdSSP13, Sch94]. Workable [GN22]. Workers [HKQ^{+}23, KJA^{+}23]. Workflow [BTV^{+}20, PBY^{+}20, TPSM17, ZWZ16, BCK07, KVW^{+}07, MAG^{+}07, Pro07, SVR^{+}07]. Workflows [MFB^{+}15, SPPH15, ASCH^{+}07, BBŁ08, BBBM12, DSS^{+}05, Gil09, JDVM10, KeKR^{+}11, SZ04, WD07, WHRH07]. Workload [DBVF01]. Works [SC23]. workspaces [KFFZ05]. workstation [Was95]. workstations [Hil97]. World [YSSL23, ZLL^{+}17, JR10]. Worldwide [VCT05]. Wound [OWL^{+}23]. WRF [MAC^{+}20]. Write [ZLL^{+}15]. Write-Combined [ZLL^{+}15]. Writing [Ano07c, Jua23b]. Wrong [LH23b]. WSF2 [FGROL^{+}16].

X [FMG^{+}23, GLD^{+}24, LZWL23, SZW^{+}23, ZWW^{+}23b]. X-Ray [GLD^{+}24, LZWL23, SZW^{+}23, ZWW^{+}23b]. X-Rays [FMG^{+}23]. XCAT [KBG^{+}02]. XCell [KGBB09]. Xeon [SCD^{+}15, SRO^{+}15, TSP^{+}15, ACCTA^{+}15, DGH^{+}15]. Xeon-Phi [ACCTA^{+}15]. xHPF [DS97]. Xiamen [WHH^{+}21]. Xilingol [SCW^{+}23]. XMatch [AMM05]. XMPP [SCKW23]. XPath [CWH23]. Xuan [HSL^{+}23].

Yangtze [HWY^{+}23, Xu23a]. year [Met99a, Met99b]. Years [Szy07]. Yellow [PFQ^{+}23]. YOLO [AMY^{+}21]. YOLOv3 [YY23b]. YOLOv4 [CLL^{+}20, TMY^{+}23, YXW^{+}22]. YOLOv4-Tiny [TMY^{+}23]. Youth [LF23b].

Zhejiang [QHJ23]. Zoltan [BÇCD12]. Zone [HWY^{+}23].

References

Asiri:2023:PEB

[AA23] Ahmad Yahya Asiri and Sultan S. Alshamrani. Performance evaluation of a B2C model based on trust requirements and factors. Scientific Programming, 2023(1): 9935849:1–9935849:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                                         https://onlinelibrary.
                                                         wiley.com/doi/epdf/10.
                                                         1155/2021/9935849.

Awan:2023:UBE

[AAA^{+}23] Mujtaba Awan, Sikandar Ali, Mushtaq Ali, Muhammad Faisal Abrar, Hamid Ullah, and Dawar Khan. Usability barriers for elderly users in smartphone app usage: an analytical hierarchical process-based prioritization. Scientific Programming, 2023(1):2780257:1– 2780257:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                                         https://onlinelibrary.
                                                         wiley.com/doi/epdf/10.
                                                         1155/2021/2780257.

Ayala:2019:EES

[AAF19] Inmaculada Ayala, Mercedes Amor, and Lidia Fuentes. An energy efficiency study of Webbased communication in Android phones. Scientific Programming, 2019 (1):8235458:1–8235458:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/8235458.

Almarashi:2023:NES

[AAH^{+}23] Abdullah M. Almarashi, Ali Algarni, Amal S. Hassan, M. Elgarhy, Farrukh Jamal, Christophe Chesneau, Khudir Alrashidi, Wali Khan Mashwani, and Heba F. Nagy. A new estimation study of the stress-strength reliability for the Topp– Leone distribution using advanced sampling methods. Scientific Programming, 2023(1):2404997:1– 2404997:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2404997.

Awan:2023:SCU

[AAS^{+}23] Malik Daler Ali Awan, Sikandar Ali, Ali Samad, Nadeem Iqbal, Malik Muhammad Saad Missen, and Niamat Ullah. Sentence classification using N-grams in Urdu language text. Scientific Programming, 2023(1): 1296076:1–1296076:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1296076.

Appelbe:1996:STH

[AB96] Bill Appelbe and Donna Bergmark. Software tools for high-performance computing: survey and recommendations. Scientific Programming, 5(3):239– 249, Fall 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ataee:2018:IAS

[AB18] Shahin Mehdipour Ataee and Zeki Bayram. An improved abstract state machine based choreography specification and execution algorithm for semantic web services. Scientific Programming, 2018(1): 4094951:1–4094951:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/4094951.

Afsarmanesh:2002:VGG

[ABB^{+}02] H. Afsarmanesh, R. G.

Belleman, A. S. Z. Belloum, et al. VLAM-G: a Grid-based virtual laboratory. Scientific Programming, 10(2): 173–181, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=1f99bpyvlg7t46lx8ue3%
26referrer=parent%26backto=
issue%2C8%2C8%3Bjournal%
2C1%2C12%3Blinkingpublicationresults%
2C1%2C1.

Abdulla:2002:HBT

[Abd02] Waleed H. Abdulla. HMMbased techniques for speech segments extraction. Scientific Programming, 10(3):221–239, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ashfaq:2023:IAA

[ABK^{+}23] Fariha Ashfaq, Imran Sarwar Bajwa, Rafaqut Kazmi, Akmal Khan, and Muhammad Ilyas. An intelligent analytics approach to minimize complexity in ambiguous software requirements. Scientific Programming, 2023(1): 6616564:1–6616564:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6616564.

Asgari-Bidhendi:2023:PBP

[ABNJMB23] Majid Asgari-Bidhendi, Mehrdad Nasser, Behrooz Janfada, and Behrouz Minaei-Bidgoli. PER-LEX: a bilingual persian-English gold dataset for relation extraction. Scientific Programming, 2023 (1):8893270:1–8893270:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/8893270.

Anwar:2017:DIM

[ABR17] Waheed Anwar, Imran Sarwar Bajwa, and Shabana Ramzan. Design and implementation of a machine learningbased authorship identification model. Scientific Programming, 2017(1): 9431073:1–9431073:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2019/9431073.

Almasri:2019:EEM

[ACA19] Ammar Almasri, Erbug Celebi, and Rami S.

Alkhawaldeh. EMT: Ensemble meta-based tree model for predicting student performance. Scientific Programming, 2019 (1):3610248:1–3610248:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/3610248.

Aloisio:2002:GRB

[ACB^{+}02] Giovanni Aloisio, Massimo Cafaro, Euro Blasi, et al. The Grid Resource Broker, a ubiquitous grid computing framework. Scientific Programming, 10(2):113– 119, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=1f99bpyvlg7t46lx8ue3%
26referrer=parent%26backto=
issue%2C3%2C8%3Bjournal%
2C1%2C12%3Blinkingpublicationresults%
2C1%2C1.

Ardito:2020:TBP

[ACBV20] Luca Ardito, Riccardo Coppola, Luca Barbato, and Diego Verga. A tool-based perspective on software code maintainability metrics: a systematic literature review. Scientific Programming, 2020(1):8840389:1– 8840389:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/8840389.

Arroyo:2015:PSD

[ACCTA^{+}15] Maricela Arroyo, Carlos Couder-Castañeda, Alfredo Trujillo-Alcantara, Israel-Enrique Herrera-Diaz, and Nain Vera-Chavez. A performance study of a dual Xeon-Phi cluster for the forward modelling of gravitational fields. Scientific Programming, 2015(??):316012:1– 316012:14, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://www.hindawi.
                                        com/journals/sp/2015/
                                        316012/.

Ancourt:1997:LAF

[ACIK97] Corinne Ancourt, Fabien Coelho, Francois Irigoin, and Ronan Keryell. Linear algebra framework for static High Performance Fortran code distribution. Scientific Programming, 6(1):3–27, Spring 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ahn:2016:TPA

[ACLS16] Jun-Woo Ahn, Tai-Woo Chang, Sung-Hee Lee, and Yong Won Seo. Twophase algorithm for optimal camera placement. Scientific Programming, 2016(??):4801784:1–4801784:16, ???? 2016. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2016/
4801784/.

Ardito:2019:MGM

[ACMT19] Luca Ardito, Riccardo Coppola, Maurizio Morisio, and Marco Torchiano. Methodological guidelines for measuring energy consumption of software applications. Scientific Programming, 2019(1):5284645:1– 5284645:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/5284645.

Adalsteinsson:2008:CAC

[ADLN08] Helgi Adalsteinsson, Bert J.

Debusschere, Kevin R. Long, and Habib N. Najm. Components for atomistic-to-continuum multiscale modeling of flow in micro- and nanofluidic systems. Scientific Programming, 16(4):297– 313, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anderson:1995:DMS

[ADS95] Charles W. Anderson, Saikumar V. Devulapalli, and Erik A. Stolz. Determining mental state from EEG signals using parallel implementations of neural networks. Scientific Programming, 4(3):171– 183, Fall 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Aslot:2003:QPA

[AE03] Vishal Aslot and Rudolf Eigenmann. Quantitative performance analysis of the SPEC OMPM2001 benchmarks. Scientific Programming, 11(2):105– 124, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ali:2019:MQS

[AE19] Ahmed K. Ali and Ergun Erçelebi. An M-QAM signal modulation recognition algorithm in AWGN channel. Scientific Programming, 2019(1):6752694:1– 6752694:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2019/6752694.

Amara:2023:EAV

[AFB23] Dalila Amara, Ezzeddine Fatnassi, and Latifa Ben Arfa Rabai. An empirical assessment and validation of redundancy metrics using defect density as reliability indicator. Scientific Programming, 2023(1):8325417:1– 8325417:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8325417.

Ahmad:2018:LBP

[AFU18] Najeeb Ahmad, Muhammad Nufail Farooqi, and Didem Unat. Load balancing for parallel multiphase flow simulation. Scientific Programming, 2018(1):6387049:1– 6387049:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/6387049.

Ayguade:1997:DRT

[AGG^{+}97] Eduard Ayguade, Jordi Garcia, Merce Girones, M. Luz Grande, and Jesus Labarta. DDT: a research tool for automatic data distribution in High Performance Fortran. Scientific Programming, 6(1): 73–94, Spring 1997. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Averbuch:1994:PES

[AGIS94] A. Averbuch, E. Gabber, S. Itzikowitz, and B. Shoham. On the parallel elliptic single/ multigrid solutions about aligned and nonaligned bodies using the Virtual Machine for Multiprocessors. Scientific Programming, 3(1):13–32, Spring 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

AlHasan:2012:BKD

[AHCZ12] Mohammad Al Hasan, Jun Huan, Jake Chen, and Mohammed J. Zaki. Biological knowledge discovery and data mining. Scientific Programming, 20(1):1–2, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ajili:2023:CWA

[AHM23] Sondes Ajili, Mohamed Ali Hajjaji, and Abdellatif Mtibaa. Crypto-watermarking algorithm using Weber’s Law and AES: a view to transfer safe medical image. Scientific Programming, 2023(1):5559191:1– 5559191:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5559191.

Alor-Hernandez:2018:SPT

[AHMMÁR18] Giner Alor-Hernández, Jezreel Mejía-Miranda, and José María Álvarez-Rodríguez. Scientific programming techniques and algorithms for dataintensive engineering environments. Scientific Programming, 2018(1): 1351239:1–1351239:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/1351239.

Aftab:2021:ANR

[AIA^{+}21] Zain Aftab, Waheed Iqbal, Khaled Mohamad Almustafa, Faisal Bukhari, and Muhammad Abdullah. Automatic NoSQL to relational database transformation with dynamic schema mapping. Scientific Programming, 2021 (1):8813350:1–8813350:??, 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.

                                        wiley.com/doi/epdf/10.
                                        1155/2020/8813350.

Alaghband:1994:OFS

[AJ94] Gita Alaghband and Harry F. Jordan. Overview of the force scientific parallel language. Scientific Programming, 3(1):33–47, Spring 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Arabas:2014:FTB

[AJJF14] Sylwester Arabas, Dorota Jarecka, Anna Jaruga, and Maciej Fijalkowski. Formula translation in Blitz++, NumPy and modern Fortran: A case study of the language choice tradeoffs. Scientific Programming, 22(3):201– 222, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Adufu:2018:MIB

[AK18] Theodora Adufu and Yoonhee Kim. Mitigating interference between scientific applications in OS-Level virtualized environments. Scientific Programming, 2018(1): 7548527:1–7548527:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2018/7548527.

Amazal:2023:NBD

[AK23] Houda Amazal and Mohamed Kissi. A new big data feature selection approach for text classification. Scientific Programming, 2023(1):6645345:1– 6645345:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6645345.

Alsahli:2016:TAA

[AKA16] Abdulaziz Alsahli, Hameed Khan, and Sultan Alyahya. Toward an agile approach to managing the effect of requirements on software architecture during global software development. Scientific Programming, 2016(??): 8198039:1–8198039:16, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
8198039/.

Aljameel:2023:MLB

[AKA^{+}23] Sumayh S. Aljameel, Irfan Ullah Khan, Nida Aslam, Malak Aljabri, and Eman S. Alsulmi. Machine learning-based model to predict the disease severity and outcome in COVID-19 patients. Scientific Programming, 2023(1):5587188:1–5587188:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5587188.

Alghaili:2020:FFI

[ALA20] Mohammed Alghaili, Zhiyong Li, and Hamdi A. R. Ali. FaceFilter: Face identification with deep learning and filter algorithm. Scientific Programming, 2020(1): 7846264:1–7846264:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/7846264.

Alkhawaldeh:2019:DGR

[Alk19] Rami S. Alkhawaldeh. DGR: Gender recognition of human speech using one-dimensional conventional neural network. Scientific Programming, 2019(1):7213717:1– 7213717:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2019/7213717.

Ai:2016:EMC

[ALL^{+}16] Wei Ai, Kenli Li, Shenglin Lan, Fan Zhang, Jing Mei, Keqin Li, and Rajkumar Buyya. On elasticity measurement in cloud computing. Scientific Programming, 2016 (??):7519507:1–7519507:13, ???? 2016. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2016/
7519507/.

AlShawi:2016:ADM

[AlS16] Amany AlShawi. Applying data mining techniques to improve information security in the cloud: A single cache system approach. Scientific Programming, 2016(??): 2385654:1–2385654:5, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
2385654/.

Alturki:2023:ROS

[Alt23] Ryan Alturki. Research onion for smart IoTenabled mobile applications. Scientific Programming, 2023(1):4270998:1– 4270998:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4270998.

Alla:2023:MPN

[AMB23] Hajar Alla, Lahcen Moumoun, and Youssef Balouki. A multilayer perceptron neural network with selective-data training for flight arrival delay prediction. Scientific Programming, 2023(1):5558918:1– 5558918:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5558918.

Ali:2023:MEC

[AMH23] Daler Ali, Malik Muhammad Saad Missen, and Mujtaba Husnain. Multiclass event classification from text. Scientific Programming, 2023(1): 6660651:1–6660651:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6660651.

Andreozzi:2005:XLS

[AMM05] Sergio Andreozzi, Danilo Montesi, and Rocco Moretti. XMatch: a language for satisfactionbased selection of Grid services. Scientific Programming, 13(4):299–316, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Agrawal:2012:LCS

[AMN^{+}12] Ankit Agrawal, Sanchit Misra, Ramanathan Narayanan, Lalith Polepeddi, and Alok Choudhary. Lung cancer survival prediction using ensemble data mining on SEER data. Scientific Programming, 20(1):29–42, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Afrati:2017:CCM

[AMS17] Foto Afrati, Zaid Momani, and Nikos Stasinopoulos. Cross-checking multiple data sources using multiway join in MapReduce. Scientific Programming, 2017(1):3072813:1– 3072813:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/3072813.

Ashraf:2020:OPG

[AMT20] Adnan Ashraf, Amin Majd, and Elena Troubitsyna. Online path generation and navigation for swarms of UAVs. Scientific Programming, 2020 (1):8530763:1–8530763:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8530763.

Ali:2023:RAA

[AMU^{+}23] Farhad Ali, Wali Khan Mashwani, Hamayat Ullah, Ahmed Hussein Msmali, Ikramullah Ikramullah, and Zabidin Salleh. Reduction of asymptotic approximate expansion of Navier–Stokes equation and solution of inviscid Burgers equation by similarity transformation. Scientific Programming, 2023(1):9054328:1– 9054328:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9054328.

Alcala:2023:SPT

[AMVGC^{+}23] Francisco J. Alcalá, Jaime Martínez-Valderrama, Francisco Gomáriz-Castillo, Carlos G. Hernández, and José M. Cecilia. Scientific programming tools for water management. Scientific Programming, 2023(1):9828596:1– 9828596:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9828596.

Ahmad:2021:ODT

[AMY^{+}21] Tanvir Ahmad, Yinglong Ma, Muhammad Yahya, Belal Ahmad, Shah Nazir, and Amin ul Haq. Object detection through modified YOLO neural network. Scientific Programming, 2021(1):8403262:1– 8403262:??, 2021. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8403262.

Altun:2019:SSB

[AN19] Oguz Altun and Orhan Nooruldeen. SKETRACK: Stroke-based recognition of online hand-drawn sketches of arrow-connected diagrams and digital logic circuit diagrams. Scientific Programming, 2019 (1):6501264:1–6501264:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/6501264.

Allan:2008:MSS

[ANEA08] Benjamin A. Allan, Boyana Norris, Wael R. Elwasif, and Robert C. Armstrong. Managing scientific software complexity with Bocca and CCA. Scientific Programming, 16(4):315–327, ???? 2008.

CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Angus:1993:ADC

[Ang93]  Ian G. Angus.  Ap-

plications demand classspecific optimizations: the C++ compiler can do more. Scientific Programming, 2(4):123–131, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:1993:HPF

[Ano93] Anonymous. High Performance FORTRAN/ journal of development. Scientific Programming, 2(1–2):1–165, Spring– Summer 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:1997:AIV

[Ano97a] Anonymous. Author index volume 6. Scientific Programming, 6(4): 403–??, Winter 1997. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:1997:CP

[Ano97b] Anonymous. Call for papers. Scientific Programming, 6(4):401, 1997. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

http://iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=a5tkxhqqy9eefb7hwkak%
26referrer=parent%26backto=
issue%2C7%2C7%3Bjournal%
2C9%2C9%3Blinkingpublicationresults%
2C1%2C1.

Anonymous:1999:AIV

[Ano99a] Anonymous. Author index volume 7 (1999). Scientific Programming, 7(3– 4):335–336, 1999. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

http://iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=53f7mftrrm4r73yyrqau%
26referrer=parent%26backto=
issue%2C12%2C12%3Bjournal%
2C6%2C9%3Blinkingpublicationresults%
2C1%2C1.

Anonymous:1999:HPJ

[Ano99b] Anonymous. High performance Java compilation and runtime issues. Scientific Programming, 7 (2):85, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2000:AIV

[Ano00] Anonymous. Author index: Volume 7 (1999). Scientific Programming, 8 (1):335–336, 2000. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2001:I

[Ano01] Anonymous. Introduction. Scientific Programming, 9(2–3):69–71, 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2002:I

[Ano02] Anonymous. Introduction. Scientific Programming, 10(3):183–184, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2003:RV

[Ano03] Anonymous. Reviewers for volumes 9–10. Scientific Programming, 11 (1):77–78, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2004:I

[Ano04a] Anonymous. Introduction. Scientific Programming, 12(4):199, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2004:P

[Ano04b] Anonymous. Preface. Scientific Programming, 12 (2):63, ???? 2004. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2005:F

[Ano05] Anonymous. Foreword.

Scientific Programming, 13(2):65–66, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2007:BRT

[Ano07a] Anonymous. Book review: TSP^{SM} Coaching Development Teams (SEI Series in Software Engineering), by Watts S. Humphrey. Scientific Programming, 15(1): 67–69, ???? 2007. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2007:BRU

[Ano07b] Anonymous. Book review: Understanding and Implementing the Finite Element Method, by Mark Gockenbach. Scientific Programming, 15(2):117– 119, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2007:BRW

[Ano07c] Anonymous. Book review: Writing Scientific Software: A Guide to Good Style, by Suely Oliveira and David Stewart. Scientific Programming, 15(3):189–190, ???? 2007. CODEN SCIPEV.

ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2008:AIV

[Ano08a] Anonymous. Author index volume 16 (2008). Scientific Programming, 16(4):343–344, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2008:HPC

[Ano08b] Anonymous. High performance computing on Cell B.E. processors. Scientific Programming, 16(1): 99, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2008:RS

[Ano08c] Anonymous. Regular section. Scientific Programming, 16(1):79, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2009:AIV

[Ano09] Anonymous. Author index volume 17 (2009). Scientific Programming, 17(4):347–348, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2011:AIV

[Ano11a] Anonymous. Author index volume 19 (2011). Scientific Programming, 19(4):265–266, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2011:BR

[Ano11b] Anonymous. Book review. Scientific Programming, 19(2–3):179– 184, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2012:AIV

[Ano12] Anonymous. Author index volume 20 (2012– 2013). Scientific Programming, 20(4):393–394, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2014:AIV

[Ano14] Anonymous. Author index volume 22 (2014). Scientific Programming, 22(4):331–332, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2019:REH

[Ano19] Anonymous. [Retracted] enhancing health risk prediction with deep learning on big data and revised fusion node paradigm. Scientific Programming, 2019(1): 1901876:1–1901876:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/1901876.

Anonymous:2022:RRA

[Ano22a] Anonymous. [Retracted] research on analysis and classification of vulnerability of electromagnetic pulse with a STM32 single-chip microcomputer. Scientific Programming, 2022(1):6836158:1– 6836158:??, 2022. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6836158.

Anonymous:2022:RRC

[Ano22b] Anonymous. [Retracted] research on crude oil trade procurement model based on DEA-Malmquist algorithm. Scientific Programming, 2022(1): 6360439:1–6360439:??, 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6360439.

Anonymous:2023:RCMb

[Ano23a] Anonymous. [Retracted] a college music teaching system designed based on Android platform. Scientific Programming, 2023 (1):7460924:1–7460924:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7460924.

Anonymous:2023:RQE

[Ano23b] Anonymous. [Retracted] a quantitative enhancement mechanism of university students’ employability and entrepreneurship based on deep learning in the context of the digital era. Scientific Programming, 2023(1): 7245465:1–7245465:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7245465.

Anonymous:2023:RET

[Ano23c] Anonymous. [Retracted] An English teaching resource recommendation system based on network behavior analysis. Scientific Programming, 2023(1):6191543:1– 6191543:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6191543.

Anonymous:2023:RIE

[Ano23d] Anonymous. [Retracted] an intelligent evaluation method of information course teaching effect based on image analysis. Scientific Programming, 2023(1):3200865:1– 3200865:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3200865.

Anonymous:2023:RATa

[Ano23e] Anonymous. [Retracted] analysis of tai chi ideological and political course in university based on big data and graph neural networks. Scientific Programming, 2023(1): 9914908:1–9914908:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9914908.

Anonymous:2023:RAAa

[Ano23f] Anonymous. [Retracted] analysis of the 3D application evaluation system of landscape based on hybrid cooperation of VR and AI. Scientific Programming, 2023(1): 4595808:1–4595808:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2021/4595808.

Anonymous:2023:RAD

[Ano23g] Anonymous. [Retracted] apparel design and development based on 3D scanning technology. Scientific Programming, 2023 (1):7933206:1–7933206:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7933206.

Anonymous:2023:RAAb

[Ano23h] Anonymous. [Retracted] application analysis of 3D printing technology in design field: Taking shoe design as an example. Scientific Programming, 2023(1):5662460:1– 5662460:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5662460.

Anonymous:2023:RAV

[Ano23i] Anonymous. [Retracted] application and value of dynamic volume CT multiparameter imaging in the diagnosis of traumatic high-flow penile abnormal erection. Scientific Programming, 2023(1):4234177:1– 4234177:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/4234177.

Anonymous:2023:RAC

[Ano23j] Anonymous. [Retracted] application of computer simulation analysis in green city garden plant landscape design. Scientific Programming, 2023 (1):9422417:1–9422417:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9422417.

Anonymous:2023:RAI

[Ano23k] Anonymous. [Retracted] application of intelligent sensor algorithm in student management information fusion. Scientific Programming, 2023(1): 3053538:1–3053538:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/3053538.

Anonymous:2023:RATb

[Ano23l] Anonymous. [Retracted] application of thematic context-based deep learning in foreign language teaching. Scientific Programming, 2023(1):

8664219:1–8664219:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8664219.

Anonymous:2023:RAR

[Ano23m] Anonymous. [Retracted] application research on roller skater detection, tracking, and trajectory prediction based on video stream. Scientific Programming, 2023(1): 2702272:1–2702272:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2702272.

Anonymous:2023:RCT

[Ano23n] Anonymous. [Retracted] computed tomography imaging based on edge detection algorithm in diagnosis and rehabilitation nursing of stroke patients with motor dysfunction. Scientific Programming, 2023(1):5499351:1– 5499351:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5499351.

Anonymous:2023:RCC

[Ano23o] Anonymous. [Retracted] construction of college English mobile learning model based on distributed terminal. Scientific Programming, 2023 (1):2220096:1–2220096:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2220096.

Anonymous:2023:RCI

[Ano23p] Anonymous. [Retracted] construction of innovation and entrepreneurship platform based on deep learning algorithm. Scientific Programming, 2023 (1):1833979:1–1833979:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1833979.

Anonymous:2023:RCMa

[Ano23q] Anonymous. [Retracted] construction of machine learning model based on text mining and ranking of meituan merchants. Scientific Programming, 2023(1):5165115:1–5165115:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5165115.

Anonymous:2023:RCR

[Ano23r] Anonymous. [Retracted] construction of the 3D reconstruction system of building construction scene based on deep learning. Scientific Programming, 2023(1):5839391:1– 5839391:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5839391.

Anonymous:2023:RCL

[Ano23s] Anonymous. [Retracted] construction of the luxury marketing model based on machine learning classification algorithm. Scientific Programming, 2023 (1):6511552:1–6511552:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6511552.

Anonymous:2023:RCO

[Ano23t] Anonymous. [Retracted] construction of the open oral evaluation model based on the neural network. Scientific Programming, 2023(1):3928246:1– 3928246:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

                                        wiley.com/doi/epdf/10.
                                        1155/2021/3928246.

Anonymous:2023:RDL

[Ano23u] Anonymous. [Retracted] deep learning-based identification of spinal metastasis in lung cancer using spectral CT images. Scientific Programming, 2023(1):2779390:1– 2779390:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/2779390.

Anonymous:2023:RDB

[Ano23v] Anonymous. [Retracted] DenseNet-based classification of MRI images for detecting the difference before and after treating liver cancer. Scientific Programming, 2023(1):4609256:1– 4609256:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/4609256.

Anonymous:2023:RDH

[Ano23w] Anonymous. [Retracted] design of human-machine interaction interface for autonomous vehicles based on multidimensional perceptual context. Scientific Programming, 2023(1):

5859800:1–5859800:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5859800.

Anonymous:2023:RDI

[Ano23x] Anonymous. [Retracted] design of ideological and political multimedia network teaching resources integration system based on wireless network. Scientific Programming, 2023(1):4293771:1– 4293771:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4293771.

Anonymous:2023:RDO

[Ano23y] Anonymous. [Retracted] design of online music teaching system based on B/S architecture. Scientific Programming, 2023 (1):1297019:1–1297019:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1297019.

Anonymous:2023:REP

[Ano23z] Anonymous. [Retracted] enlightenment of physical education teaching experiment based on cloud computing to the current physical education reform. Scientific Programming, 2023(1):6607539:1– 6607539:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6607539.

Anonymous:2023:REU

[Ano23-27] Anonymous. [Retracted] exploration of using modern information technology to realize college English distance teaching. Scientific Programming, 2023(1):3317549:1– 3317549:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3317549.

Anonymous:2023:RIS

[Ano23-28] Anonymous. [Retracted] ice and snow sports education based on 5G cloud computing to improve the social adaptability of southern university students. Scientific Programming, 2023(1):3828624:1– 3828624:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3828624.

Anonymous:2023:RIV

[Ano23-29] Anonymous. [Retracted] identifying and verifying AR, ERBB2, and VEGFA are the targets of qigesan in the treatment of esophageal adenocarcinoma in silico and in vitro. Scientific Programming, 2023(1):5463724:1– 5463724:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5463724.

Anonymous:2023:RIP

[Ano23-30] Anonymous. [Retracted] intelligent prediction algorithm of cross-border e-commerce logistics cost based on cloud computing. Scientific Programming, 2023(1):7038294:1– 7038294:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7038294.

Anonymous:2023:RMC

[Ano23-31] Anonymous. [Retracted] K-means clustering algorithmbased detection of carotid atherosclerotic plaque using contrast-enhanced ultrasound images. Scientific Programming, 2023 (1):2223344:1–2223344:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/2223344.

Anonymous:2023:RMM

[Ano23-32] Anonymous. [Retracted] mind map construction for English grammar teaching based on knowledge map. Scientific Programming, 2023(1): 4921825:1–4921825:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/4921825.

Anonymous:2023:RPC

[Ano23-33] Anonymous. [Retracted] prediction of college students’ employment rate based on gray system. Scientific Programming, 2023(1):4182011:1– 4182011:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/4182011.

Anonymous:2023:RPS

[Ano23-34] Anonymous. [Retracted] preliminary study on the evaluation of multimodal effects of type 2 diabetic nephropathy. Scientific Programming, 2023(1):2563477:1– 2563477:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2563477.

Anonymous:2023:RRDa

[Ano23-35] Anonymous. [Retracted] research on deep learning algorithm in cultural and creative product design. Scientific Programming, 2023(1):3212882:1– 3212882:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3212882.

Anonymous:2023:RRDb

[Ano23-36] Anonymous. [Retracted] research on distance teaching system of English course based on wireless network technology. Scientific Programming, 2023(1):3275340:1– 3275340:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3275340.

Anonymous:2023:RRE

[Ano23-37] Anonymous. [Retracted] research on English education auxiliary teaching system based on MOOC. Scientific Programming, 2023(1):2459228:1–2459228:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/2459228.

Anonymous:2023:RRI

[Ano23-38] Anonymous. [Retracted] research on intelligent English translation method based on the improved attention mechanism model. Scientific Programming, 2023(1):9667255:1–9667255:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9667255.

Anonymous:2023:RRN

[Ano23-39] Anonymous. [Retracted] research on network layer recursive reduction model compression for image recognition. Scientific Programming, 2023(1): 4054435:1–4054435:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/4054435.

Anonymous:2023:RRS

[Ano23-40] Anonymous. [Retracted] research on students’ adaptive learning system based on deep learning model. Scientific Programming, 2023(1): 6593438:1–6593438:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6593438.

Anonymous:2023:RRA

[Ano23-41] Anonymous. [Retracted] research on the application of environmental art design based on the combination of VR and panoramic video technology. Scientific Programming, 2023(1):9820550:1– 9820550:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9820550.

Anonymous:2023:RRG

[Ano23-42] Anonymous. [Retracted] research on the generation and design method of residential community scheme based on improving design. Scientific Programming, 2023(1): 5053494:1–5053494:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5053494.

Anonymous:2023:RROa

[Ano23-43] Anonymous. [Retracted] research on the optimization strategy of innovation behavior and entrepreneurship intention in entrepreneurship teaching. Scientific Programming, 2023(1): 4872108:1–4872108:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/4872108.

Anonymous:2023:RROb

[Ano23-44] Anonymous. [Retracted] retrospection of the optimization model for designing the power train of a formula student race car. Scientific Programming, 2023(1):9465702:1– 9465702:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9465702.

Anonymous:2023:RSR

[Ano23-45] Anonymous. [Retracted] study on resource sharing strategy of e-commerce innovation and entrepreneurship education based on cloud computing. Scientific Programming, 2023 (1):8268000:1–8268000:??, 2023. CODEN SCIPEV.

ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8268000.

Anonymous:2023:RPM

[Ano23-46] Anonymous. [Retracted] the process and model innovation of ideological education network communication in colleges and universities based on cloud computing. Scientific Programming, 2023(1):7302877:1– 7302877:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7302877.

Anonymous:2023:RUI

[Ano23-47] Anonymous. [Retracted] ultrasound image-guided pudendal nerve block on analgesic effect of perineotomy under optimized fast super resolution reconstructed convolutional neural network algorithm. Scientific Programming, 2023(1): 4768673:1–4768673:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4768673.

Anonymous:2023:RWE

[Ano23-48] Anonymous. [Retracted] willingness and evaluation model of college students’ online learning behavior based on distributed cognition. Scientific Programming, 2023(1):6386455:1– 6386455:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6386455.

Anonymous:2024:RSA

[Ano24] Anonymous. [Retracted] sports action recognition based on image processing technology and analysis of the development of sports industry pattern. Scientific Programming, 2024(1):4815097:1– 4815097:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/4815097.

Alvarez-Rodriguez:2018:SSP

[ÁRAHMM18] Jose María Álvarez-Rodríguez, Giner Alor-Hernández, and Jezreel Mejía-Miranda. Survey of scientific programming techniques for the management of dataintensive engineering environments. Scientific Programming, 2018(1): 8467413:1–8467413:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/8467413.

Ashworth:2019:FSP

[ARAM19] Mike Ashworth, Graham D. Riley, Andrew Attwood, and John Mawer. First steps in porting the LFRic weather and climate model to the FPGAs of the EuroExa architecture. Scientific Programming, 2019(1):7807860:1– 7807860:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/7807860.

Alvarez-Rodriguez:2015:STH

[ÁRCPS15] Jose María Álvarez-Rodríguez, Ricardo Colomo-Palacios, and Vladimir Stantchev. Skillrank: Towards a hybrid method to assess quality and confidence of professional skills in social networks. Scientific Programming, 2015 (??):451476:1–451476:13, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://www.hindawi.
                                        com/journals/sp/2015/
                                        451476/.

Agosta:2005:JJT

[ARFS05] Giovanni Agosta, Stefano Crespi Reghizzi, Gerlando Falauto, and Martino Sykora. JIST: Just-In-Time Scheduling translation for parallel processors. Scientific Programming, 13(3):239–253, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Arevalo:2019:PEU

[ARGC19] C. Arevalo, I. Ramos, J. Gutiérrez, and M. Cruz. Practical experiences in the use of patternrecognition strategies to transform software project plans into software business processes of information technology companies. Scientific Programming, 2019(1):7973289:1– 7973289:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2019/7973289.

Addison:2003:OIA

[ARvW03] C. Addison, Y. Ren, and M. van Waveren. OpenMP issues arising in the development of parallel BLAS and LAPACK libraries. Scientific Programming, 11(2):95–104, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Al-Shakarchi:2007:DWU

[ASCH^{+}07] Eddie Al-Shakarchi, Pasquale Cozza, Andrew Harrison, Carlo Mastroianni, Matthew Shields, Domenico Talia, and Ian Taylor. Distributing workflows over a ubiquitous P2P network. Scientific Programming, 15(4):269–281, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ahirwar:2023:CID

[ASS23] Manish Kumar Ahirwar, Piyush Kumar Shukla, and Rakesh Singhai. CBO-IE: a data mining approach for healthcare IoT dataset using chaotic biogeography-based optimization and information entropy. Scientific Programming, 2023(1): 8715668:1–8715668:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8715668.

Alsharif:2023:ASP

[ASZF23] Hussain Zaid H. Alsharif, Tong Shu, Bin Zhu, and Zeyad Farisi. An adaptive smoothness parameter strategy for variational optical flow model. Scientific Programming, 2023(1): 7594636:1–7594636:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7594636.

Asif:2015:HSB

[ATZ^{+}15] Muhammad Asif, Imtiaz A. Taj, S. M. Ziauddin, Maaz Bin Ahmad, and M. Tahir. A hybrid scheme based on pipelining and multitasking in mobile application processors for advanced video coding. Scientific Programming, 2015 (??):197843:1–197843:16, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
197843/.

Ali:2020:FMD

[AUA^{+}20] Sikandar Ali, Niamat Ullah, Muhammad Faisal Abrar, Zhongguo Yang, and Jiwei Huang. Fuzzy multicriteria decisionmaking approach for measuring the possibility of cloud adoption for software testing. Scientific Programming, 2020(1):6597316:1– 6597316:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/6597316.

Awan:2023:EIC

[AUA^{+}23] Mujtaba Awan, Niamat Ullah, Sikandar Ali, Irshad Ahmed Abbasi, Muhammad Shabbir Hassan, Hizbullah Khattak, and Jiwei Huang. An empirical investigation of the challenges of cloud-based ERP adoption in Pakistani SMEs. Scientific Programming, 2023(1): 5547237:1–5547237:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5547237.

Atwan:2023:USA

[AWB^{+}23] Jaffar Atwan, Mohammad Wedyan, Qusay Bsoul, Ahmad Hammadeen, and Ryan Alturki. The use of stemming in the Arabic text and its impact on the accuracy of classification. Scientific Programming, 2023(1):1367210:1– 1367210:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/1367210.

AlBdairi:2020:IEP

[AXA20] Ahmed Jawad A. AlBdairi, Zhu Xiao, and Mohammed Alghaili. Identifying ethnics of people through face recognition: a deep CNN approach. Scientific Programming, 2020(1): 6385281:1–6385281:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/6385281.

An:2023:SCD

[AXWS23] Xin-Dai An, Xiang-Wen Xie, Di Wu, and Ke-Feng Song. Slope collapse detection based on image processing. Scientific Programming, 2023(1): 5565329:1–5565329:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5565329.

An:2024:NCP

[AZGJ24] Tong An, Tao Zhang, Yanzhang Geng, and Haiquan Jiao. Normalized combinations of proportionate affine projection sign subband adaptive filter. Scientific Programming, 2024(1):8826868:1– 8826868:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8826868.

Aziguli:2017:RTC

[AZX^{+}17] Wulamu Aziguli, Yuanyu Zhang, Yonghong Xie, Dezheng Zhang, Xiong Luo, Chunmiao Li, and Yao Zhang. A robust text classifier based on denoising deep neural network in the analysis of big data. Scientific Programming, 2017(1):3610378:1– 3610378:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/3610378.

Bernabe:2018:PSC

[BA18] Gregorio Bernabé and Manuel E. Acacio. On the parallelization of stream compaction on a low-cost SDC cluster. Scientific Programming, 2018(1):2037272:1– 2037272:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

                                        wiley.com/doi/epdf/10.
                                        1155/2018/2037272.

Bailey:1995:USC

[Bai95] David H. Bailey. Unfavorable strides in cache memory systems. Scientific Programming, 4(2):53–58, Summer 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). Also available as RNR technical report RNR-92-015.

Bailey:2008:BR

[Bai08] David H. Bailey. Book review. Scientific Programming, 16(1):97–98, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bai:2023:GTM

[Bai23] Rui Bai. Game teaching method in preschool education based on big data technology. Scientific Programming, 2023(1): 4751263:1–4751263:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/4751263.

Bakir:2023:SLM

[Bak23]  Cigdem Bakir.  A

single-label model to ensure data consistency in information security. Scientific Programming, 2023(1):9913645:1– 9913645:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9913645.

Bryan:2002:AER

[BAN02] Greg L. Bryan, Tom Abel, and Michael L. Norman. Achieving extreme resolution in numerical cosmology using adaptive mesh refinement: resolving primordial star formation. Scientific Programming, 10(4):291–302, 2002. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Beamer:2013:DOB

[BAP13] Scott Beamer, Krste Asanović, and David Patterson. Directionoptimizing breadth-first search. Scientific Programming, 21(3–4):137– 148, ???? 2013. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Baker:2008:SPA

[BB08] Mark Baker and Richard Boakes. Slogger: a profiling and analysis system based on Semantic Web technologies. Scientific Programming, 16(2– 3):183–204, ???? 2008.

CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bodin:2009:HMP

[BB09] Francois Bodin and Stephane Bihan. Heterogeneous multicore parallel programming for graphics processing units. Scientific Programming, 17 (4):325–336, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Baranowski:2012:CWS

[BBBM12] Mikolaj Baranowski, Adam Belloum, Marian Bubak, and Maciej Malawski. Constructing workflows from script applications. Scientific Programming, 20(4):359–377, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Budimlic:2010:CC

[BBC^{+}10] Zoran Budimlić, Michael Burke, Vincent Cavé, Kathleen Knobe, Geoff Lowney, Ryan Newton, Jens Palsberg, David Peixotto, Vivek Sarkar, Frank Schlimbach, and Sagnak Tasirlar. Concurrent collections. Scientific Programming, 18(3– 4):203–217, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Badin:2011:IAM

[BBDN11] Matthew Badin, Lubomir Bic, Michael Dillencourt, and Alexandru Nicolau. Improving accuracy for matrix multiplications on GPUs. Scientific Programming, 19(1):3–11, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bali:2004:GEL

[BBF^{+}04] Bartosz Bali, Marian Bubak, Włodzimierz Funika, Roland Wismüller, Marcin Radecki, Tomasz Szepieniec, Tomasz Arod, and Marcin Kurdziel. Grid environment for online application monitoring and performance analysis. Scientific Programming, 12(4):239–251, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bodin:1993:DPB

[BBG^{+}93] François Bodin, Peter Beckman, Dennis Gannon, Srinivas Narayana, and Shelby X. Yang. Distributed pC++: Basic ideas for an object parallel language. Scientific Programming, 2(3): 7–22, Fall 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL ftp:

//cica.cica.indiana.

                                        edu/pub/sage/pC++SciPro.
                                        ps.gz.

Batar:2023:DRB

[BBI23] Mustafa Batar, Kökten Ulas Birant, and Ali Hakan Isik. Development of rulebased software risk assessment and management method with fuzzy inference system. Scientific Programming, 2023(1): 5532197:1–5532197:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5532197.

Blazewicz:2011:CPA

[BBK^{+}11] Marek Blazewicz, Steven R. Brandt, Michal Kierzynka, Krzysztof Kurowski, Bogdan Ludwiczak, Jian Tao, and Jan Weglarz. CaKernel — a parallel application programming framework for heterogeneous computing architectures. Scientific Programming, 19(4):185–197, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Balis:2008:MGS

[BBŁ08] Bartosz Balis, Marian Bubak, and Bartłomiej Łabno. Monitoring of Grid scientific workflows. Scientific Programming, 16(2–3):205–216, ????

2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bischof:2003:IES

[BBLR03] Christian H. Bischof, H. Martin Bücker, Bruno Lang, and Arno Rasch. An interactive environment for supporting the transition from simulation to optimization. Scientific Programming, 11 (4):263–272, ???? 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Blount:1999:EJN

[BC99] Brian Blount and Siddhartha Chatterjee. An evaluation of Java for numerical computing. Scientific Programming, 7 (2):97–110, 1999. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

http://iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=f277qlrwwjr5m4vxjywv%
26referrer=parent%26backto=
issue%2C3%2C8%3Bjournal%
2C7%2C9%3Blinkingpublicationresults%
2C1%2C1.

Bratu:2023:DAA

[BC23] Andrei Bratu and Gabriela Czibula. DAuGAN: an approach for augmenting time series imbalanced datasets via latent space sampling using adversarial techniques. Scientific Programming, 2023(1): 7877590:1–7877590:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/7877590.

Bischof:1992:AGD

[BCC^{+}92] Christian H. Bischof, Alan Carle, George F. Corliss, Andreas Griewank, and Paul Hovland. ADI-FOR: Generating derivative code from Fortran programs. Scientific Programming, 1(1):11–29, 1992. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bhatt:1993:OOS

[BCC^{+}93] Sandeep Bhatt, Marina Chen, James Cowie, Cheng-Yee Lin, and Pangfeng Liu. Objectoriented support for adaptive methods on parallel machines. Scientific Programming, 2(4):179–192, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bircsak:2000:EON

[BCC^{+}00] John Bircsak, Peter Craig, RaeLyn Crowell, et al. Extending OpenMP for NUMA machines. Scientific Programming, 8 (3):163–181, 2000. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Boman:2012:ZIP

[BÇCD12] Erik G. Boman, Ümit V.

Çatalyürek, Cédric Chevalier, and Karen D. Devine. The Zoltan and Isorropia parallel toolkits for combinatorial scientific computing: Partitioning, ordering and coloring. Scientific Programming, 20 (2):129–150, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Blake:2005:TES

[BCCP05] R. J. Blake, P. V.

Coveney, P. Clarke, and S. M. Pickles. The TeraGyroid experiment — Supercomputing 2003. Scientific Programming, 13(1):1–17, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Banicescu:2005:VNT

[BCHL05] Ioana Banicescu, Ricolindo L. Cariño, Jane L. Harvill, and John Patrick Lestrade. Vector nonlinear time-series analysis of gamma-ray burst datasets on heterogeneous clusters. Scientific Programming, 13(2):67–77, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bahsi:2007:CWM

[BCK07] Emir M. Bahsi, Emrah Ceyhan, and Tevfik Kosar. Conditional workflow management: a survey and analysis. Scientific Programming, 15 (4):283–297, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bala:2001:APC

[BCS01] Piotr Bała, Terry Clark, and L. Ridgway Scott. Application of Pfortran and Co-Array Fortran in the parallelization of the GROMOS96 molecular dynamics module. Scientific Programming, 9(1):61–68, ???? 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

                                        //iospress.metapress.
                                        com/app/home/contribution.
                                        asp%3Fwasp=f2779jvvqg63jq64qwtm%
                                        26referrer=parent%26backto=
                                        issue%2C6%2C6%3Bjournal%
                                        2C2%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Beamonte:2016:DCP

[BD16] Raphaël Beamonte and Michel R. Dagenais. Detection of common problems in real-time and multicore systems using model-based constraints. Scientific Programming, 2016(??):9792462:1–9792462:18, ???? 2016. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2016/
9792462/.

Bano:2020:ABS

[BDAH20] Aniqa Bano, Ikram Ud Din, and Asma A. Al-Huqail. AIoT-based smart bin for real-time monitoring and management of solid waste. Scientific Programming, 2020(1):6613263:1–6613263:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/6613263.

Beguelin:1994:HHN

[BDG^{+}94] Adam Beguelin, Jack J.

Dongarra, George Al Geist, Robert Manchek, and Keith Moore. HeNCE: a heterogeneous network computing environment. Scientific Programming, 3(1):49–60, Spring 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//www.netlib.org/utk/
people/JackDongarra/
PAPERS/HeNCE-A-Heterogeneous-

                                        Network-Computing-Environment.
                                        pdf.

Brodtkorb:2010:SAH

[BDH^{+}10] Andre R. Brodtkorb, Christopher Dyken, Trond R. Hagen, Jon M. Hjelmervik, and Olaf O. Storaasli. State-of-the-art in heterogeneous computing. Scientific Programming, 18 (1):1–33, ???? 2010. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Berry:1994:PPD

[BDLL94] Michael W. Berry, Jack J. Dongarra, Brian H. LaRose, and Todd A. Letsche. PDS: a performance database server. Scientific Programming, 3(2): 147–156, Summer 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Barrett:2004:MTG

[BDM^{+}04] Chris Barrett, Martin Drozda, Madhav V. Marathe, S. S. Ravi, and James P. Smith. A mobility and traffic generation framework for modeling and simulating ad hoc communication networks. Scientific Programming, 12(1):1–23, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Briguglio:2003:PPM

[BDV03] Sergio Briguglio, Beniamino Di Martino, and Gregorio Vlad. A performance-prediction model for PIC applications on clusters of symmetric MultiProcessors: Validation with hierarchical HPF + OpenMP implementation. Scientific Programming, 11(2): 159–176, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bochev:2012:SPI

[BEK^{+}12] P. Bochev, H. C. Edwards, R. C. Kirby, K. Peterson, and D. Ridzal. Solving PDEs with Intrepid. Scientific Programming, 20(2):151–180, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Benkner:1999:VVF

[Ben99] Siegfried Benkner. VFC:

The Vienna Fortran Compiler. Scientific Programming, 7(1):67–81, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=64cr5a4mg33tuhcbdr02%
26referrer=parent%26backto=
issue%2C5%2C7%3Bjournal%

                                        2C8%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Bozkus:2017:AHC

[BF17] Cem Bozkus and Basilio B. Fraguela. Accelerating the HyperLogLog cardinality estimation algorithm. Scientific Programming, 2017(1): 2040865:1–2040865:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2017/2040865.

Benkner:2014:SIA

[BFGH14] Siegfried Benkner, Franz Franchetti, Hans Michael Gerndt, and Jeffrey K. Hollingsworth. Special issue on automatic application tuning for HPC architectures. Scientific Programming, 22(4):259– 260, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Burkhart:1996:SPP

[BFH96] Helmar Burkhart, Robert Frank, and Guido Hachler. Structured parallel programming: how informatics can help overcome the software dilemma. Scientific Programming, 5 (1):33–45, Spring 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bik:1999:IIM

[BGH99] Aart J. C. Bik, Milind Girkar, and Mohammad R. Haghighat. Incorporating Intel(R) MMX^{T M} technology into a Java^{T M} JIT compiler. Scientific Programming, 7(1): 167–184, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=64cr5a4mg33tuhcbdr02%
26referrer=parent%26backto=
issue%2C7%2C7%3Bjournal%
2C8%2C9%3Blinkingpublicationresults%
2C1%2C1.

Ballance:1993:FBE

[BGLR93] Robert A. Ballance, Anthony J. Giancola, George F. Luger, and Timothy J. Ross. Framework-based environment for objectoriented scientific codes. Scientific Programming, 2 (4):111–121, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bai:2023:AIB

[BGZL23] Manyun Bai, Renzhong Guo, Qian Zhao, and Yufang Li. Artificial intelligence-based CT images in analysis of postoperative recovery of patients undergoing laparoscopic cholecystectomy under balanced anesthesia. Scientific Programming, 2023(1):1125573:1– 1125573:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/1125573.

Bohm:1996:FIJ

[BH96] A. P. W. Bohm and R. E. Hiromoto. Functional implementations of the Jacobi Eigensolver. Scientific Programming, 5(2): 111–120, Summer 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Boys:2002:DOM

[BH02] R. J. Boys and D. A. Henderson. On determining the order of Markov dependence of an observed process governed by a hidden Markov model. Scientific Programming, 10(3): 241–251, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Baker:2012:TUG

[BH12] C. G. Baker and M. A. Heroux. Tpetra, and the use of generic programming in scientific computing. Scientific Programming, 20(2):115–128, ???? 2012. CODEN SCIPEV.

ISSN 1058-9244 (print), 1875-919X (electronic).

Bai:2023:RAD

[BH23] Jie Bai and Tian He. Research on audit data analysis and decision tree algorithm for benefit distribution of enterprise financing alliance. Scientific Programming, 2023 (1):1910156:1–1910156:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1910156.

Blazewicz:2013:PMR

[BHK^{+}13] Marek Blazewicz, Ian Hinder, David M. Koppelman, Steven R. Brandt, Milosz Ciznicki, Michal Kierzynka, Frank Löffler, Erik Schnetter, and Jian Tao. From physics model to results: an optimizing framework for crossarchitecture code generation. Scientific Programming, 21(1–2):1–16, ???? 2013. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Barekas:2003:MAO

[BHP^{+}03] Vasileios K. Barekas, Panagiotis E. Hadjidoukas, Eleftherios D. Polychronopoulos, et al. A multiprogramming aware OpenMP implementation. Scientific Programming, 11(2):133–141, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bavier:2012:ABD

[BHRT12] Eric Bavier, Mark Hoemmen, Sivasankaran Rajamanickam, and Heidi Thornquist. Amesos2 and Belos: Direct and iterative solvers for large sparse linear systems. Scientific Programming, 20 (3):241–255, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bian:2023:SMM

[Bia23]  Qing Bian.  Social

media marketing optimization method based on deep neural network and evolutionary algorithm. Scientific Programming, 2023(1): 5626351:1–5626351:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5626351.

Bjorstad:2000:CFN

[Bjø00] Petter E. Bjørstad. Coordinate free numerics. Scientific Programming, 8 (4):209, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=2dyrcfdc1p5ytlerkc3u%
26referrer=parent%26backto=
issue%2C1%2C5%3Bjournal%
2C3%2C9%3Blinkingpublicationresults%
2C1%2C1.

Bak:2011:GTD

[BKK^{+}11] Slawomir Bak, Marcin Krystek, Krzysztof Kurowski, Ariel Oleksiak, Wojciech Piatek, and Jan Waglarz. GSSIM — a tool for distributed computing experiments. Scientific Programming, 19(4):231–251, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bubak:2001:CJN

[BKLS01] Marian Bubak, Dawid Kurzyniec, Piotr Luszczek, and V. Sunderam. Creating Java to Native Code Interfaces with Janet. Scientific Programming, 9(1):39–50, ???? 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=f2779jvvqg63jq64qwtm%
26referrer=parent%26backto=
issue%2C4%2C6%3Bjournal%
2C2%2C9%3Blinkingpublicationresults%
2C1%2C1.

Bouaafia:2023:CLL

[BKMS23] Soulef Bouaafia, Randa Khemiri, Amna Maraoui, and Fatma Elzahra Sayadi. CNN-LSTM learning approachbased complexity reduction for high-efficiency video coding standard. Scientific Programming, 2023(1):6628041:1–6628041:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6628041.

Budimlic:1999:CBO

[BKP99] Zoran Budimlić, Ken Kennedy, and Jeff Piper. The cost of being objectoriented: a preliminary study. Scientific Programming, 7(2): 87–95, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

                                        //iospress.metapress.
                                        com/app/home/contribution.
                                        asp%3Fwasp=f277qlrwwjr5m4vxjywv%
                                        26referrer=parent%26backto=
                                        issue%2C2%2C8%3Bjournal%
                                        2C7%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

BinSiddique:2023:MLB

[BKU^{+}23] Zeeshan Bin Siddique, Mudassar Ali Khan, Ikram Ud Din, Ahmad Almogren, Irfan Mohiuddin, and Shah Nazir. Machine learning-based detection of spam emails. Scientific Programming, 2023(1):6508784:1–6508784:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6508784.

Binmad:2018:PIT

[BL18] Ruchdee Binmad and Mingchu Li. Psychologyinspired trust restoration framework in distributed multiagent systems. Scientific Programming, 2018(1):7515860:1– 7515860:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/7515860.

Bai:2023:ARK

[BL23] XuJing Bai and Jia-Jun Li. Applied research of knowledge in the field of artificial intelligence in the intelligent retrieval of teaching resources. Scientific Programming, 2023(1): 9924435:1–9924435:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9924435.

Bao:2023:CTI

[BLL^{+}23] Kunyang Bao, Chao Liu, Jin Li, Xiang Liu, Wenzhang Luo, and Changren Huang. Computed tomography image under optimized iterative reconstruction algorithm to analyze the characteristics of blood flow field in cerebral aneurysm before and after stent implantation. Scientific Programming, 2023(1):7915729:1– 7915729:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/7915729.

Li:2018:SPP

[bLwHcY18] Jia bin Li, Shi wei He, and Wei chuan Yin. The study of pallet pooling information platform based on cloud computing. Scientific Programming, 2018(1):5106392:1– 5106392:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2018/5106392.

Bai:2023:IFA

[BLZG23] Manyun Bai, Yufang Li, Qian Zhao, and Renzhong Guo. Imaging features of artificial intelligence algorithm in the analysis of cerebral protective effect of craniotomy hematoma removal under propofol anesthesia in patients with chronic subdural hematoma. Scientific Programming, 2023(1): 6435476:1–6435476:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6435476.

Bozkus:1997:POH

[BMN^{+}97] Zeki Bozkus, Larry Meadows, Steven Nakamoto, Vincent Schuster, and Mark Young. PGHPF — an optimizing High Performance Fortran compiler for distributed memory machines. Scientific Programming, 6(1):29–40, Spring 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bachelet:2013:TMT

[BMY13] Bruno Bachelet, Antoine Mahul, and Loïc Yon. Template metaprogramming techniques for concept-based specialization. Scientific Programming, 21(1–2):43–61, ???? 2013. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bui:2020:MMP

[BNH^{+}20] Toan Bui, Tram Nguyen, Huy M. Huynh, Bay Vo, Jerry Chun-Wei Lin, and Tzung-Pei Hong. Multiswarm multiobjective particle swarm optimization with simulated annealing for extracting multiple tests. Scientific Programming, 2020(1): 7081653:1–7081653:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/7081653.

Borchers:2004:BRL

[Bor04] Brian Borchers. Book review: The Lanczos Method: Evolution and Application, by Louis Komzsik. Scientific Programming, 12(3):197–198, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bellens:2009:CST

[BPC^{+}09] Pieter Bellens, Josep M. Perez, Felipe Cabarcas, Alex Ramirez, Rosa M. Badia, and Jesus Labarta. CellSs: Scheduling techniques to better exploit memory hierarchy. Scientific Programming, 17 (1–2):77–95, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Brunner:2000:SMD

[BPK00] Robert K. Brunner, James C.

Phillips, and Laxmikant V. Kalé. Scalable molecular dynamics for large biomolecular systems. Scientific Programming, 8 (3):195–207, 2000. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bin:2020:MTL

[BQR^{+}20] Chenzhong Bin, Saige Qin, Guanjun Rao, Tianlong Gu, and Liang Chang. Multiview translation learning for knowledge graph embedding. Scientific Programming, 2020(1):7084958:1–7084958:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/7084958.

Brandic:2011:SIS

[BR11] Ivona Brandic and Ioan Raicu. Special issue on science-driven cloud computing. Scientific Programming, 19(2–3):71–73, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bramas:2017:INI

[Bra17] Berenger Bramas. Inastemp: a novel intrinsics-as-template library for portable SIMD-Vectorization. Scientific Programming, 2017 (1):5482468:1–5482468:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2017/5482468.

Brownston:2009:BR

[Bro09] Lee S. Brownston. Book review. Scientific Programming, 17(4):339–341, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Brown:2011:BRa

[Bro11a] Adrian Brown. Book review. Scientific Programming, 19(1):63–65, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Brown:2011:BRb

[Bro11b] Adrian Brown. Book review. Scientific Programming, 19(4):259–264, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Brown:2012:BR

[Bro12] Adrian Brown. Book review. Scientific Programming, 20(3):355–358, ????

2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Baena-Ruiz:2024:GST

[BRPV24] Leticia Baena-Ruiz and David Pulido-Velazquez. GIS-SWIAS: Tool to summarize seawater intrusion status and vulnerability at aquifer scale. Scientific Programming, 2024(1): 8818634:1–8818634:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8818634.

Bertolini:2018:NMG

[BRSP18] Vittorio Bertolini, Carlos Rey, Mauricio Sepulveda, and Victor Parada. Novel methods generated by genetic programming for the guillotine-cutting problem. Scientific Programming, 2018(1):6971827:1– 6971827:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/6971827.

Bryson:1996:PCV

[Bry96]  Steve Bryson.  Picto-

rial communication in virtual and real environments, edited by Steve Ellis, Mary Kaiser, and Art Grunwald. Scientific Programming, 5(1):89–??, Spring 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Blikberg:2001:NPA

[BS01] Ragnhild Blikberg and Tor Sørevik. Nested parallelism: Allocation of threads to tasks and OpenMP implementation. Scientific Programming, 9(2–3): 185–194, Spring–Summer 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL http://iospress.
                                        metapress.com/app/home/
                                        contribution.asp%3Fwasp=
                                        7pab6qgbaf8vxg991rwy%
                                        26referrer=parent%26backto=
                                        issue%2C11%2C11%3Bjournal%
                                        2C1%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Bollman:1996:FDI

[BSF96] Dorothy Bollman, Jaime Seguel, and John Feo. Fast digit-index permutations. Scientific Programming, 5(2):137–146, Summer 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Breslow:2014:EFP

[BTS^{+}14] Alex D. Breslow, Ananta Tiwari, Martin Schulz, Laura Carrington, Lingjia Tang, and Jason Mars. Enabling fair pricing on high performance computer systems with node sharing. Scientific Programming, 22(2):59–74, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Broll:2020:MLG

[BTV^{+}20] Brian Broll, Umesh Timalsina, Péter Völgyesi, Tamás Budavári, and Ákos Lédeczi. A machine learning gateway for scientific workflow design. Scientific Programming, 2020(1):8867380:1– 8867380:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8867380.

Bushnell:2009:BR

[Bus09] David H. Bushnell. Book review. Scientific Programming, 17(3):275–277, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bin:2023:ITC

[BYZ^{+}23] Hou Bin, Xue Yu, Yanling Zheng, Yaohui Jiang, and Huanfang Wang. The influence of trust on crowd logistics enterprise’s operational performance: a SEM-PLS model. Scientific Programming, 2023 (1):6403293:1–6403293:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6403293.

Boudjellal:2020:BRE

[BZKA20] Nada Boudjellal, Huaping Zhang, Asif Khan, and Arshad Ahmad. Biomedical relation extraction using distant supervision. Scientific Programming, 2020(1):8893749:1– 8893749:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8893749.

Bubak:2001:PLM

[BŻvA^{+}01] Marian Bubak, Dariusz Żbik, Dick van Albada, et al. Portable library of migratable sockets. Scientific Programming, 9(4): 211–222, 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=64t4wprhwk589ellmv56%
26referrer=parent%26backto=
issue%2C2%2C4%3Bjournal%
2C3%2C12%3Blinkingpublicationresults%
2C1%2C1.

Cao:2023:DIH

[Cao23a] Jun Cao. Design and implementation of humancomputer interaction system in parallel digital library system based on neural network. Scientific Programming, 2023(1): 9921551:1–9921551:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9921551.

Cao:2023:DIM

[Cao23b]  Ming Cao.  Design

and implementation of multidimensional interaction in online English course under the assistance of omnimedia. Scientific Programming, 2023(1):3713161:1– 3713161:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3713161.

Chauveau:1999:MEH

[CB99] Stéphane Chauveau and François Bodin. Menhir: An environment for high performance Matlab. Scientific Programming, 7(3–4):303– 312, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

                                        //iospress.metapress.
                                        com/app/home/contribution.
                                        asp%3Fwasp=53f7mftrrm4r73yyrqau%
                                        26referrer=parent%26backto=
                                        issue%2C9%2C12%3Bjournal%
                                        2C6%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Chang:1993:PIP

[CBCM93] L. Chang, G. Bourianoff, B. Cole, and S. Machida. Parallel implementation of particle tracking with space charge effects on an Intel iPSC/860. Scientific Programming, 2(3): 37–47, Fall 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Chraibi:2023:MOC

[CBE23] Amine Chraibi, Said Ben Alla, and Abdellah Ezzati. Makespan optimisation in cloudlet scheduling with improved DQN algorithm in cloud computing. Scientific Programming, 2023(1):7216795:1– 7216795:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/7216795.

Contreras-Bolton:2016:APA

[CBRRC^{+}16] Carlos Contreras-Bolton, Carlos Rey, Sergio Ramos-Cossio, Claudio Rodríguez, Felipe Gatica, and Victor Parada. Automatically produced algorithms for the generalized minimum spanning tree problem. Scientific Programming, 2016(??): 1682925:1–1682925:11, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
1682925/.

Cao:2023:OPB

[CC23] Yuan Cao and Qing Cao. Optimization of post-based merit salary based on CPI and university characteristic data. Scientific Programming, 2023(1):2287596:1–2287596:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2287596.

Cui:2020:SIS

[CCB^{+}20] Qingyu Cui, Yi Cao, Wenzheng Bao, Bin Yang, and Yuehui Chen. SubRF Seq: Identification of sub-golgi protein types with random forest with partial sequence information. Scientific Programming, 2020(1):8862468:1– 8862468:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/8862468.

Couder-Castaneda:2015:PCM

[CCBPGA15] C. Couder-Castañeda, H. Barrios-Piña, I. Gitler, and M. Arroyo. Performance of a code migration for the simulation of supersonic ejector flow to SMP, MIC, and GPU using OpenMP, OpenMP+LEO, and OpenACC directives. Scientific Programming, 2015 (??):739107:1–739107:20, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://www.hindawi.
                                        com/journals/sp/2015/
                                        739107/.

Cesini:2017:PEC

[CCF^{+}17] Daniele Cesini, Elena Corni, Antonio Falabella, Andrea Ferraro, Lucia Morganti, Enrico Calore, Sebastiano Fabio Schifano, Michele Michelotto, Roberto Alfieri, Roberto De Pietri, Tommaso Boccali, Andrea Biagioni, Francesca Lo Cicero, Alessandro Lonardo, Michele Martinelli, Pier Stanislao Paolucci, Elena Pastorelli, and Piero Vicini. Power-efficient computing: Experiences from the COSA project. Scientific Programming, 2017(1): 7206595:1–7206595:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/7206595.

Chen:2017:MHS

[CCG17] Shifeng Chen, Rong Chen, and Jian Gao. A modified harmony search algorithm for solving the dynamic vehicle routing problem with time windows. Scientific Programming, 2017(1):1021432:1– 1021432:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/1021432.

Chen:2023:CEA

[CCHJ23] Hua Chen, Ming Cai, Ke Huang, and Shuxin Jin. Classification and evolution analysis of key transportation technologies based on bibliometrics. Scientific Programming, 2023(1):2977998:1– 2977998:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2977998.

Chen:2023:SCR

[CCM^{+}23] Ming Chen, Junqiang Cheng, Guanghua Ma, Liang Tian, Xiaohong Li, and Qingmin Shi. Service composition recommendation method based on recurrent neural network and naive Bayes. Scientific Programming, 2023 (1):1013682:1–1013682:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/1013682.

Chandrasekaran:2015:PML

[CCTY15] Sunita Chandrasekaran, Barbara Chapman, Xinmin Tian, and Yonghong Yan. Programming models, languages, and compilers for manycore and heterogeneous architectures. Scientific Programming, 2015(??):376317:1, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://www.hindawi.
                                        com/journals/sp/2015/
                                        376317/.

Chen:2023:GBP

[CD23] Xinfang Chen and Venkata Dinavahi. Group behavior pattern recognition algorithm based on spatiotemporal graph convolutional networks. Scientific Programming, 2023(1): 2934943:1–2934943:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2934943.

Caron:2005:MDP

[CDD^{+}05] Eddy Caron, Bruno DelFabbro, Frédéric Desprez, Emmanuel Jeannot, and Jean-Marc Nicod. Managing data persistence in network enabled servers. Scientific Programming, 13(4):333– 354, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Choi:1996:DIS

[CDO^{+}96] Jaeyoung Choi, Jack J.

Dongarra, L. Susan Ostrouchov, Antoine P. Petitet, David W. Walker, and R. Clint Whaley. Design and implementation of the ScaLAPACK LU, QR, and Cholesky factorization routines. Scientific Programming, 5(3):173– 184, Fall 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

www.netlib.org/netlib/
lapack/lawns/lawn80.
ps;   http://www.netlib.
org/netlib/lapack/lawnspdf/
lawn80.pdf;   http:/

                                        /www.netlib.org/utk/
                                        papers/factor/ftcover.
                                        html.

Cruz:2017:FES

[CEMM17] Henry Cruz, Martina Eckert, Juan M. Meneses, and J. F. Martínez. Fast evaluation of segmentation quality with parallel computing. Scientific Programming, 2017(1): 5767521:1–5767521:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2017/5767521.

Ceraolo:2021:MSP

[Cer21] Massimo Ceraolo. Modelling and simulation of power lines using Modelica language, starting from the line geometry. Scientific Programming, 2021(1):8837884:1– 8837884:??, ???? 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2024/8837884.

Chen:2023:SPF

[CFL^{+}23] Yu Chen, Ruixin Fang, Ting Liang, Zongyu Sha, Shicheng Li, Yugen Yi, Wei Zhou, and Huilin Song. Stock price forecast based on CNN-BiLSTM-ECA model. Scientific Programming, 2023(1): 2446543:1–2446543:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2446543.

Cardellini:2014:DPS

[CFR14] Valeria Cardellini, Salvatore Filippone, and Damian W. I. Rouson. Design patterns for sparse-matrix computations on hybrid CPU/ GPU platforms. Scientific Programming, 22(1): 1–19, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Cui:2023:MDM

[CFY23] Yumei Cui, Xinqun Feng, and Xinxin Yang. A matching degree management model of human body shape and fashion design based on big data analysis. Scientific Programming, 2023(1): 9384404:1–9384404:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9384404.

Chiu:2017:LPU

[CGA17] Alton Chiu, Joseph Garvey, and Tarek S. Abdelrahman. A language and preprocessor for user-controlled generation of synthetic programs. Scientific Programming, 2017(1): 1046161:1–1046161:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/1046161.

Chen:2016:OVM

[CGC16] Tao Chen, Xiaofeng Gao, and Guihai Chen. Optimized virtual machine placement with trafficaware balancing in data center networks. Scientific Programming, 2016 (??):3101658:1–3101658:10, ???? 2016. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2016/
3101658/.

Chmiel:2005:EJA

[CGK^{+}05] Krzysztof Chmiel, Maciej Gawinecki, Pawel Kaczmarek, Michal Szymczak, and Marcin Paprzycki. Efficiency of JADE agent platform. Scientific Programming, 13(2):159–172, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Chan:2008:EFN

[CGL08] Anthony Chan, William Gropp, and Ewing Lusk. An efficient format for nearly constant-time access to arbitrary time intervals in large trace files. Scientific Programming, 16(2–3):155–165, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Cruz:2019:PAP

[CGMRO^{+}19] Bruno Cruz, Silvana Gómez-Meire, David Ruano-Ordás, Helge Janicke, Iryna Yevseyeva, and Jose R. Méndez. A practical approach to protect IoT devices against attacks and compile security incident datasets. Scientific Programming, 2019(1):9067512:1–9067512:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/9067512.

Chen:2023:RAI

[CH23a] Huani Chen and Jian Huang. Research and application of the interactive English online teaching system based on the Internet of Things. Scientific Programming, 2023(1): 3636533:1–3636533:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/3636533.

Chen:2023:RFT

[CH23b] Lifei Chen and Yingying Hu. Risk factors of tumor patients with peripherally inserted central catheter related thrombosis and nursing intervention analysis mediated by algorithmbased ultrasound images. Scientific Programming, 2023(1):7295093:1– 7295093:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/7295093.

Chang:2023:CEF

[Cha23a] Heli Chang. College English flipped classroom teaching model based on big data and deep neural networks. Scientific Programming, 2023(1): 9918433:1–9918433:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9918433.

Chang:2023:AFC

[Cha23b]  Lijun Chang.  Appli-

cation of fuzzy comprehensive evaluation based on genetic algorithm in psychological measurement. Scientific Programming, 2023(1):9607006:1– 9607006:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9607006.

Carver:2015:SEC

[CHC15] Jeffrey C. Carver, Neil Chue Hong, and Selim Ciraci. Software engineering for CSE. Scientific Programming, 2015(??):591562:1– 591562:2, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
591562/.

Cheng:1993:HCN

[Che93a] Harry H. Cheng. Handling of complex numbers in the C^{H} programming language. Scientific Programming, 2(3): 77–106, Fall 1993. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Cheng:1993:SCC

[Che93b] Harry H. Cheng. Scientific computing in the C^{H} programming language. Scientific Programming, 2 (3):49–75, Fall 1993. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Chen:2022:RAO

[Che22]  Hongtao Chen.  RE-

TRACTION: Application of offensive and defensive linkage in a virtual football game in the Internet of Things decisionmaking system. Scientific Programming, 2022(1): 9816218:1–9816218:??, ???? 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2024/9816218.

Chen:2023:RAT

[Che23a] Chen Chen. Research on aerobics training and evaluation method based on artificial intelligenceaided modeling. Scientific Programming, 2023(1): 9545909:1–9545909:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9545909.

Chen:2023:SFM

[Che23b] Qunying Chen. Stepped frequency multiresolution digital signal processing. Scientific Programming, 2023(1):9081988:1– 9081988:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9081988.

Chen:2023:MLB

[Che23c] Ruling Chen. Migration learning-based bridge structure damage detection algorithm. Scientific Programming, 2023(1): 1102521:1–1102521:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1102521.

Chen:2023:RDD

[CHL^{+}23] Wei Chen, ChenYu He, JianRong Lu, Kui Yan, Jin Liu, Feng Zhou, Xin Xu, and Xiao Hao. Research and design of distributed fire alarm system of indoor Internet of Things based on LoRa. Scientific Programming, 2023(1):7462331:1– 7462331:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7462331.

Chapman:1997:OCL

[CHM^{+}97] Barbara Chapman, Matthew Haines, Piyush Mehrotra, Hans Zima, and John Van Rosendale. Opus: a coordination language for multidisciplinary applications. Scientific Programming, 6(4):345–362, Winter 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=a5tkxhqqy9eefb7hwkak%
26referrer=parent%26backto=
issue%2C2%2C7%3Bjournal%
2C9%2C9%3Blinkingpublicationresults%
2C1%2C1.

Carter:1999:IMS

[CHS^{+}99] John B. Carter, Wilson C. Hsieh, Leigh B. Stoller, et al. Impulse: Memory system support for scientific applications. Scientific Programming, 7(3–4):195– 209, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=53f7mftrrm4r73yyrqau%
26referrer=parent%26backto=
issue%2C2%2C12%3Bjournal%
2C6%2C9%3Blinkingpublicationresults%
2C1%2C1.

Cheng:2023:TMS

[CHTZ23] Yanjin Cheng, Hao Hao, Shipeng Tao, and Yanjun Zhou. Traceability management strategy of the EV power battery based on the blockchain. Scientific Programming, 2023(1): 5601833:1–5601833:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5601833.

Chun:2021:SIP

[Chu21] Hyunjin Chun. A study on the impact of 3D printing and artificial intelligence on education and learning process. Scientific Programming, 2021(1):2247346:1– 2247346:??, 2021. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2247346.

Chen:2017:ALD

[CHZY17] Hao Chen, Dengxin Hua, Yikun Zhang, and Qing Yan. Atmospheric lidar data storage model based on ontology. Scientific Programming, 2017(1):3607068:1– 3607068:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/3607068.

Cornu:1996:DIT

[CIN^{+}96] Thierry Cornu, Paolo Ienne, Dagmar Niebur, Patrick Thiran, and Marc A. Viredaz. Design, implementation, and test of a multi-model systolic neural-network accelerator. Scientific Programming, 5(1):47–61, Spring 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Celikmih:2020:FPA

[CIU20] Kadir Celikmih, Onur Inan, and Harun Uguz. Failure prediction of aircraft equipment using machine learning with a hybrid data preparation method. Scientific Programming, 2020(1): 8616039:1–8616039:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/8616039.

Cao:2002:AAB

[CJS^{+}02] Junwei Cao, Stephen A. Jarvis, Subhash Saini, et al. ARMS: An agentbased resource management system for grid computing. Scientific Programming, 10(2):135– 148, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=1f99bpyvlg7t46lx8ue3%
26referrer=parent%26backto=
issue%2C5%2C8%3Bjournal%
2C1%2C12%3Blinkingpublicationresults%
2C1%2C1.

Chapman:2008:F

[CK08] Barbara Chapman and Dieter Kranzlmüller. Foreword. Scientific Programming, 16(2–3):101– 103, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ciznicki:2015:MLB

[CKKK15] Milosz Ciznicki, Michal Kulczewski, Piotr Kopta, and Krzysztof Kurowski. Methods to load balance a GCR pressure solver using a stencil framework on multi- and manycore architectures. Scientific Programming, 2015 (??):648752:1–648752:13, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
648752/.

Choi:2009:EMT

[CKS^{+}09] Wonil Choi, Hyunhee Kim, Wook Song, Jiseok Song, and Jihong Kim. ePRO-MP: a tool for profiling and optimizing energy and performance of mobile multiprocessor applications. Scientific Programming, 17(4):285–294, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Chow:2004:CCM

[CL04] P. Chow and C.-H. Lai. Collaborating components in mesh-based electronic packaging. Scientific Programming, 12 (2):65–70, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Cai:2023:DOC

[CL23a] Lingyi Cai and Wei Liu. Decisions on the orderliness and collaborative operation mechanism of each subsystem under the organizational model of the Internet of Things. Scientific Programming, 2023(1): 2179343:1–2179343:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/2179343.

Chen:2023:CAB

[CL23b] Liang Chen and Qiong Li. Classification algorithmbased CT imaging in diagnosis of acute respiratory distress syndrome and analysis of pathogenic factors. Scientific Programming, 2023(1):4100856:1– 4100856:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4100856.

Cheng:2023:IAM

[CL23c] Jingjing Cheng and Qi Li.

Integrated architecture model of tourism information service in smart scenic spots based on hybrid cloud. Scientific Programming, 2023(1): 3827306:1–3827306:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3827306.

Chapuis:2015:PSB

[CLA^{+}15] Guillaume Chapuis, Mathilde Le Boudic-Jamin, Rumen Andonov, Hristo Djidjev, and Dominique Lavenier. Parallel seed-based approach to multiple protein structure similarities detection. Scientific Programming, 2015 (??):279715:1–279715:12, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://www.hindawi.
                                        com/journals/sp/2015/
                                        279715/.

Cao:2020:ARC

[CLDX20] Kerang Cao, Yuqing Liu, Lini Duan, and Tian Xie. Adaptive residual channel attention network for single image superresolution. Scientific Programming, 2020(1): 8877851:1–8877851:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/8877851.

Chen:2023:ACT

[CLF^{+}23] Yan Chen, Changkun Lin, Qimao Fu, Jinghui Huang, and Chuizhi Huang. Adoption of computed tomography images under iterative reconstruction algorithm in diagnosis of gastric cancer. Scientific Programming, 2023(1):8968288:1– 8968288:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/8968288.

Chen:2023:SDE

[CLG^{+}23] Zhikui Chen, Chaojie Li, Jing Gao, Jianing Zhang, and Peng Li. Semisupervised deep embedded clustering with adaptive labels. Scientific Programming, 2023(1):6613452:1– 6613452:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6613452.

Chen:2020:IFP

[CLHH20] Xiaoping Chen, Jiamin Lin, Han Huang, and Yunpeng Huang. Impact factors on posterior modified transfacet debridement for thoracic spinal tuberculosis basing on regression and classification analysis. Scientific Programming, 2020(1):8892815:1– 8892815:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8892815.

Chang:2017:DMB

[CLL17] Bao Rong Chang, Yun-Da Lee, and Po-Hao Liao. Development of multiple big data analytics platforms with rapid response. Scientific Programming, 2017(1): 6972461:1–6972461:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2017/6972461.

Chu:2018:RFS

[CLL^{+}18] Feng Chu, Ming Liu, Xin Liu, Chengbin Chu, and Juan Jiang. Reentrant flow shop scheduling considering multiresource qualification matching. Scientific Programming, 2018(1): 2615096:1–2615096:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2018/2615096.

Chen:2020:DCO

[CLL^{+}20] Wenkang Chen, Shenglian Lu, Binghao Liu, Guo Li, and Tingting Qian. Detecting citrus in orchard environment by using improved YOLOv4. Scientific Programming, 2020 (1):8859237:1–8859237:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/8859237.

Cao:2023:HIE

[CLLZ23] HuiRu Cao, Xiaomin Li, Songyao Lian, and Choujun Zhan. A hotspot information extraction hybrid solution of online posts’ textual data. Scientific Programming, 2023(1):6619712:1– 6619712:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6619712.

Cai:2005:PPP

[CLM05] Xing Cai, Hans Petter Langtangen, and Halvard Moe. On the performance of the Python programming language for serial and parallel scientific computations. Scientific Programming, 13 (1):31–56, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Cai:2024:EIT

[CLN^{+}24] Tao Cai, Peiyao Liu, Dejiao Niu, Jiancong Shi, and Lei Li. The embedded IoT time series database for hybrid solid-state storage system. Scientific Programming, 2024(1):9948533:1– 9948533:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9948533.

Collados-Lara:2020:STG

[CLPVPI20] Antonio-Juan Collados-Lara, David Pulido-Velazquez, and Eulogio Pardo-Igúzquiza. A statistical tool to generate potential future climate scenarios for hydrology applications. Scientific Programming, 2020(1): 8847571:1–8847571:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8847571.

Chen:2016:PGA

[CLTX16] Zhi Chen, Tao Lin, Ningjiu Tang, and Xin Xia. A parallel genetic algorithm based feature selection and parameter optimization for support vector machine. Scientific Programming, 2016(??): 2739621:1–2739621:10, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
2739621/.

Chen:2023:PCB

[CLW^{+}23] Ping Chen, Wenmin Lu, Jue Wang, Zhanling Guo, and Lianchang Liu. Pattern classification-based analysis on gastrointestinal multifunctional apparatus guided by ultrasound image after resection of esophageal carcinoma. Scientific Programming, 2023(1):4034216:1– 4034216:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4034216.

Cui:2023:DDS

[CLXZ23] Huanqing Cui, Ruixue Liu, Shaohua Xu, and Chuanai Zhou. DMGA: a distributed shortest path algorithm for multistage graph. Scientific Programming, 2023(1): 6639008:1–6639008:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6639008.

Choi:2016:TCB

[CLYL16] HeeSeok Choi, Jong-Beom Lim, Heonchang Yu, and EunYoung Lee. Task classification based energy-aware consolidation in clouds. Scientific Programming, 2016(??): 6208358:1–6208358:13, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
6208358/.

Cappello:2002:CSR

[CM02] Peter Cappello and Dimitrios Mourloukos. CX: a scalable, robust network for parallel computing. Scientific Programming, 10(2):159– 171, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=1f99bpyvlg7t46lx8ue3%
26referrer=parent%26backto=
issue%2C7%2C8%3Bjournal%
2C1%2C12%3Blinkingpublicationresults%
2C1%2C1.

Ceraolo:2021:EML

[CM21] Massimo Ceraolo and Mirko Marracci. An example of Modelica-LabVIEW communication usage to implement hardware-in-the-loop experiments. Scientific Programming, 2021(1): 9648349:1–9648349:??, ???? 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2024/9648349.

Cesar:2002:DPT

[CMM^{+}02] Eduardo César, Anna Morajko, Tomàs Margalef, et al. Dynamic performance tuning supported by program specification. Scientific Programming, 10 (1):35–44, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=9ejnuvwuvby9737jte27%
26referrer=parent%26backto=
issue%2C4%2C9%3Bjournal%
2C2%2C12%3Blinkingpublicationresults%
2C1%2C1.

Cortes-Mendoza:2016:BVS

[CMTAC^{+}16] Jorge M. Cortés-Mendoza, Andrei Tchernykh, Fermin A. Armenta-Cano, Pascal Bouvry, Alexander Yu. Drozdov, and Loic Didelot. Biobjective VoIP service management in cloud infrastructure. Scientific Programming, 2016(??): 5706790:1–5706790:14, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
5706790/.

Chapman:1992:PVF

[CMZ92] Barbara M. Chapman, Piyush Mehrotra, and Hans P. Zima. Programming in Vienna Fortran. Scientific Programming, 1 (1):31–50, Fall 1992. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Chu:2023:BDV

[CNW^{+}23] Xiao Chu, Shah Nazir, Kunhao Wang, Zeqi Leng, and Wajeeha Khalil. Big data and its v’s with IoT to develop sustainability. Scientific Programming, 2023(1):3780594:1– 3780594:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/3780594.

Calhoun:1993:SAD

[CO93] Donna Calhoun and Roy Overstreet. Sensitivity analysis of a dynamical system using C++. Scientific Programming, 2 (4):157–169, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Casanova:2000:APS

[COB^{+}00] Henri Casanova, Graziano Obertelli, Francine Berman, et al. The AppLeS parameter sweep template: User-level middleware for the Grid. Scientific Programming, 8(3):111–126, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Crooks:1995:LCD

[CP95] P. Crooks and R. H. Perrott. Language constructs for data partitioning and distribution. Scientific Programming, 4(2):59–85, Summer 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Cole:2003:DCC

[CP03] Martin J. Cole and Steven G. Parker. Dynamic compilation of C++ template code. Scientific Programming, 11 (4):321–327, ???? 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Czarnul:2021:SMA

[CPD21] Paweł Czarnul, Jerzy Proficz, and Krzysztof Drypczewski. Survey of methodologies, approaches, and challenges in parallel programming using high-performance computing systems. Scientific Programming, 2021 (1):4176794:1–4176794:??, 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/4176794.

Chen:2023:SPE

[CPG^{+}23] Shaobo Chen, Yinzhen Pi, Haiyan Gong, Huaijun Wang, and Shu Liu. Single-photon emission computed tomography image-assisted diagnosis of thyroid diseases under convolutional network neural algorithm. Scientific Programming, 2023(1): 4200271:1–4200271:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/4200271.

Czarnul:2019:EAH

[CPK19] Pawel Czarnul, Jerzy Proficz, and Adam Krzywaniak. Energy-aware high-performance computing: Survey of stateof-the-art tools, techniques, and environments. Scientific Programming, 2019(1):8348791:1–8348791:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2019/8348791.

Clayton:2005:ASC

[CQF05] Brian C. Clayton, Thomas B. Quillinan, and Simon N. Foley. Automating security configuration for the Grid. Scientific Programming, 13(2):113–125, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Craus:2005:PFC

[CR05] Mitică Craus and Laurent.iu Rudeanu. Parallel framework for cooperative processes. Scientific Programming, 13(3):205–217, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Chapin:1994:SIS

[CS94] Steve J. Chapin and Eugene H. Spafford. Support for implementing scheduling algorithms using MESSIAHS. Scientific Programming, 3(4): 325–340, Winter 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Chen:2016:OCR

[CSC16] Zhengyu Chen, Jianhua Sun, and Hao Chen. Optimizing checkpoint restart with data deduplication. Scientific Programming, 2016(??): 9315493:1–9315493:11, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
9315493/.

Ciesielczyk:2015:ELR

[CSM15] Michał Ciesielczyk, Andrzej Szwabe, and Mikołaj Morzy. On efficient link recommendation in social networks using actorfact matrices. Scientific Programming, 2015 (??):450215:1–450215:9, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
450215/.

Crawford:2018:AHF

[CSM^{+}18] Broderick Crawford, Ricardo Soto, Marco Alarcón San Martín, Hanns de la Fuente-Mella, Carlos Castro, and Fernando Paredes. Automatic highfrequency trading: an application to emerging Chilean stock market. Scientific Programming, 2018(1):8721246:1– 8721246:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/8721246.

Cytowski:2017:IAB

[CSU^{+}17] Maciej Cytowski, Zuzanna Szymańska, Piotr Umiński, Grzegorz Andrejczuk, and Krzysztof Raszkowski. Implementation of an agent-based parallel tissue modelling framework for the Intel MIC architecture. Scientific Programming, 2017(1):8721612:1– 8721612:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/8721612.

Chen:2023:DNT

[CSWW23] Yu Chen, Lingyan Sun, Zonghui Wang, and Jinghua Wang. Distribution network topology identification based on IEC 61850 logical nodes. Scientific Programming, 2023(1): 6639432:1–6639432:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6639432.

Chen:2018:NCL

[CSZ^{+}18] Tianfei Chen, Lijun Sun, Qiuwen Zhang, Xiang Wu, and Defeng Wu. Nonmetric correction of lens distortion based on entropy measure. Scientific Programming, 2018(1):4230185:1– 4230185:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/4230185.

Chen:2019:FGC

[CSZ^{+}19] Tianfei Chen, Lijun Sun, Qiuwen Zhang, Xiang Wu, and Defeng Wu. Field geometric calibration method for line structured light sensor using single circular target. Scientific Programming, 2019(1):1526706:1– 1526706:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/1526706.

Chen:2023:RCI

[CT23] Yu Chen and Zhong Tang. Research on the construction of intelligent community emergency service platform based on convolutional neural network. Scientific Programming, 2023(1):5089236:1– 5089236:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5089236.

Chang:2015:ISI

[CTC^{+}15] Bao Rong Chang, Hsiu-Fen Tsai, Chia-Yen Chen, Chien-Feng Huang, and Hung-Ta Hsu. Implementation of secondary index on cloud computing NoSQL database in big data environment. Scientific Programming, 2015 (??):560714:1–560714:10, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
560714/.

Chang:2015:EAH

[CTCG15] Bao Rong Chang, Hsiu-Fen Tsai, Chia-Yen Chen, and Cin-Long Guo. Empirical analysis of high efficient remote cloud data center backup using HBase and Cassandra. Scientific Programming, 2015(??):294614:1– 294614:10, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
294614/.

Chen:2023:NNA

[CTD23] Lei Chen, Yingying Tian, and Yujie Deng. Neural network algorithm-based three-dimensional ultrasound evaluation in the diagnosis of fetal spina bifida. Scientific Programming, 2023(1):3605739:1– 3605739:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3605739.

Chang:2023:ACT

[CTS23] Bao Rong Chang, Hsiu-Fen Tsai, and Po-Wen Su. Applying code transform model to newly generated program for improving execution performance. Scientific Programming, 2023(1): 6691010:1–6691010:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6691010.

Chen:2024:ESE

[CTZC24] Xuliang Chen, Sichao Tai, Pengxiang Zhou, and Xiao Chen. Evaluation of surgical effect of atrial septal defect with tricuspid regurgitation by transesophageal 3D echocardiography based on MC image reconstruction algorithm. Scientific Programming, 2024(1): 9944191:1–9944191:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9944191.

Cigsar:2019:CDM

[ÇÜ19] Begüm Çigsar and Deniz Ünal. Comparison of data mining classification algorithms determining the default risk. Scientific Programming, 2019(1):8706505:1– 8706505:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/8706505.

Cui:2023:AST

[Cui23] Li Cui. Analysis of spatial and temporal distribution characteristics of land desertification based on GIS and remote sensing images. Scientific Programming, 2023(1):7557175:1– 7557175:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7557175.

Clark:1997:EDP

[CvHK97] Terry W. Clark, Reinhard v. Hanxleden, and Ken Kennedy. Experiences in data-parallel programming. Scientific Programming, 6(1):153–158, Spring 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Crutchfield:1993:OOI

[CW93] William Y. Crutchfield and Michael L. Welcome. Object-oriented implementation of adaptive mesh refinement algorithms. Scientific Programming, 2(4):145–156, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Cong:2023:PED

[CW23] Rui Cong and Hailong Wang. Prediction of evolution and development trend in sports industry cluster based on particle swarm optimization. Scientific Programming, 2023(1):7607623:1– 7607623:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7607623.

Chen:2023:BAL

[CWD23] Wenjie Chen, Xiaogang Wu, and Ngabo Desire. Benefit analysis of lowcarbon policy mix innovation based on consumer perspective in smart city. Scientific Programming, 2023(1):3282398:1– 3282398:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3282398.

Chen:2023:PXQ

[CWH23] Rongxin Chen, Zongyue Wang, and Yuling Hong. Pipelined XPath query based on cost optimization. Scientific Programming, 2023(1):5559941:1– 5559941:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2021/5559941.

Cheng:2023:OAS

[CWLW23] Chunhong Cheng, Junyan Wang, Hongbo Li, and Liang Wang. Optimal atlas segmentation on CT images for diagnosis of pediatric mycoplasma pneumonia. Scientific Programming, 2023(1):2586956:1– 2586956:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2586956.

Cha:2023:IAB

[CWMZ23] Yufeng Cha, Zhili Wei, Chi Ma, and Lei Zhang. Intelligent algorithm-based computed tomography image features in diagnosis of the effect of radiofrequency ablation in liver cancer. Scientific Programming, 2023(1): 3422484:1–3422484:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3422484.

Chen:2024:ADD

[CWS^{+}24] Yuquan Chen, Hongxing Wang, Jie Shen, Xingwei Zhang, and Xiaowei Gao. Application of datadriven iterative learning algorithm in transmission line defect detection. Scientific Programming, 2024(1):9976209:1– 9976209:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9976209.

Chen:2023:PGD

[CWSF23] Xieyi Chen, Dongyun Wang, Jinjun Shao, and Jun Fan. Plastic gasket defect detection based on transfer learning. Scientific Programming, 2023(1):5990020:1– 5990020:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5990020.

Chen:2017:EEN

[CYJ^{+}17] Yong Chen, Juncheng Yao, Hai Jin, Chunjiang He, and Hanhua Chen. Exploring the evolution of new mobile services. Scientific Programming, 2017(1):5159690:1– 5159690:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2017/5159690.

Chen:2020:TTP

[CYJ^{+}20] Wenjie Chen, Qiliang Yang, Ziyan Jiang, Jianchun Xing, Qianchuan Zhao, Qizhen Zhou, and Deshuai Han. Touch: a textual programming language for developing APPs of insect intelligent building. Scientific Programming, 2020(1):8887588:1– 8887588:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8887588.

Chen:2023:NDB

[CYLD23] Ya Chen, Hongliang Yuan, Tingting Liu, and Nan Ding. Name disambiguation based on graph convolutional network. Scientific Programming, 2023(1):5577692:1– 5577692:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5577692.

Chen:2023:STP

[CYSZ23] Yingyu Chen, Shenhua Yang, Yongfeng Suo, and Minjie Zheng. Ship track prediction based on DLGWO-SVR. Scientific Programming, 2023(1): 9085617:1–9085617:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9085617.

Chen:2020:GPL

[CYZ^{+}20] Wenjie Chen, Qiliang Yang, Shuo Zhao, Jianchun Xing, and Qizhen Zhou. A graphical programming language and its supporting tool for insect intelligent building. Scientific Programming, 2020(1):9634389:1– 9634389:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/9634389.

Chen:2023:MBD

[CZ23] Donghua Chen and Runtong Zhang. MapReducebased dynamic partition join with Shannon entropy for data skewness. Scientific Programming, 2023(1):1602767:1– 1602767:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1602767.

Cai:2024:NTE

[CZB24] Jiangwei Cai, Lu Zhao, and Anqi Bi. A novel transfer enhanced αexpansion move learning model for EEG signals. Scientific Programming, 2024(1):9957845:1– 9957845:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9957845.

Chen:2023:APS

[CZCL23a] Binbin Chen, Rui Zhang, Long Chen, and Shengjie Long. Adaptive particle swarm optimization with Gaussian perturbation and mutation. Scientific Programming, 2023(1):6676449:1– 6676449:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6676449.

Chen:2023:DCA

[CZCL23b] Lin Chen, Peng Zhan, Luhui Cao, and Xueqing Li. Discrimination and correlation analysis of multiview SAR images with application to target recognition. Scientific Programming, 2023(1):6646388:1– 6646388:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6646388.

Chen:2020:ACI

[CZD^{+}20] Weixiang Chen, Jianfu Zhao, Zhenhui Dai, Mingyue Lv, Zhenhua Yang, and Yuqin Zhang. Application of CT image technology based on nearest neighbor propagation clustering segmentation algorithm in lung cancer radiotherapy. Scientific Programming, 2020(1):7289102:1– 7289102:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7289102.

Chen:2023:AIB

[CZFX23] Yantao Chen, Qinyao Zeng, Biao Feng, and Haixia Xiong. Artificial intelligence-based inferior vena cava images under dezocine anesthesia in detection of bile duct injury after laparoscopic cholecystectomy. Scientific Programming, 2023(1):4661206:1– 4661206:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4661206.

Chi:2024:PIO

[CZH24] Wentao Chi, Xuemei Zhao, and Lufei Huang. The price impact of order book events from a dimension of time. Scientific Programming, 2024 (1):9949565:1–9949565:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9949565.

Cao:2024:NAM

[CZKY24] Le Cao, Wenyan Zhang, Xiu Kan, and Wei Yao. A novel adaptive mutation PSO optimized SVM algorithm for sEMGbased gesture recognition. Scientific Programming, 2024(1):9988823:1– 9988823:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9988823.

Chen:2024:WPF

[CZP^{+}24] Jiashun Chen, Hao Zhang, Dechang Pi, Mehmed Kantardzic, Qi Yin, and Xin Liu. A weight possibilistic fuzzy cmeans clustering algorithm. Scientific Programming, 2024(1):9965813:1– 9965813:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9965813.

Cui:2023:ASO

[CZS^{+}23] Yanqing Cui, Islam Zada, Sara Shahzad, Shah Nazir, Shafi Ullah Khan, Naveed Hussain, and Muhammad Asshad. Analysis of service-oriented architecture and scrum software development approach for IIoT. Scientific Programming, 2023(1): 6611407:1–6611407:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6611407.

Cheng:2023:PDM

[CZW^{+}23] Ruihan Cheng, Longfei Zhang, Shiqi Wu, Sen Xu, Shang Gao, and Hualong Yu. Probability density machine: a new solution of class imbalance learning. Scientific Programming, 2023(1):7555587:1– 7555587:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7555587.

Chu:2023:HIM

[CZWY23] Jinlong Chu, Qiang Zhang, Ai Wang, and Haoran Yu. A hybrid intelligent model for urban seismic risk assessment from the perspective of possibility and vulnerability based on particle swarm optimization. Scientific Programming, 2023(1):2218044:1– 2218044:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2218044.

Chen:2020:QRA

[CZY20] Xiaoping Chen, Lihui Zheng, and Jianqi Yao. Quantile regression analysis of depression and clinical symptom degree in Chinese patients with spinocerebellar ataxia type 3. Scientific Programming, 2020(1): 1394617:1–1394617:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/1394617.

Chen:2023:IQS

[CZY^{+}23] Guangxi Chen, Liang Zeng, Liu Yang, Yixian Yu, Panli Sun, and Tao Yao. Improved quantitative susceptibility mapping under Laplace algorithm in diagnosis of Parkinson’s disease. Scientific Programming, 2023(1):8210526:1– 8210526:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8210526.

Dai:2023:AIT

[Dai23] Dan Dan Dai. Artificial intelligence technology assisted music teaching design. Scientific Programming, 2023(1):9141339:1– 9141339:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9141339.

deBulnes:2020:DMH

[dBMT20] Darian Reyes Fernandez de Bulnes, Yazmin Maldonado, and Leonardo Trujillo. Development of multiobjective highlevel synthesis for FP-GAs. Scientific Programming, 2020(1):7095048:1– 7095048:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/7095048.

DiMartino:2001:WDS

[DBVF01] Beniamino Di Martino, Sergio Briguglio, Gregorio Vlad, and Giuliana Fogaccia. Workload decomposition strategies for shared memory parallel systems with OpenMP. Scientific Programming, 9(2–3):109–122, Spring– Summer 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=7pab6qgbaf8vxg991rwy%
26referrer=parent%26backto=
issue%2C5%2C11%3Bjournal%
2C1%2C9%3Blinkingpublicationresults%
2C1%2C1.

Dai:2024:SVM

[DBWY24] Huajia Dai, Yuhao Bian, Libin Wang, and Junfeng Yang. Support vector machine-based backprojection algorithm for detection of gastric cancer lesions with abdominal endoscope using magnetic resonance imaging images. Scientific Programming, 2024(1):9964203:1– 9964203:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9964203.

Dai:2022:IDM

[DC22] Liqin Dai and ChunHua Chen. Intelligent detection method of English text in natural scenes in video. Scientific Programming, 2022(1):6239112:1– 6239112:??, 2022. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6239112.

Dong:2017:UMC

[DCF17] Mingjie Dong, Wusheng Chou, and Bin Fang. Underwater matching correction navigation based on geometric features using sonar point cloud data. Scientific Programming, 2017(1):7136702:1– 7136702:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2017/7136702.

Dong:2023:IPG

[DCH^{+}23] Yi Dong, Long Chang, Long Hei, Sensen Yang, Wenxin Ma, and Huiqiang Ding. Inhibition of PPAR γ gene particle downregulates the differentiation of rabbit bone marrow mesenchymal stem cells into adipocytes. Scientific Programming, 2023(1):

5335241:1–5335241:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5335241.

Dastani:2023:ICP

[DD23a] Meisam Dastani and Farshid Danesh. Iranian COVID-19 publications in LitCovid: Text mining and topic modeling. Scientific Programming, 2023(1):3315695:1– 3315695:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3315695.

Duan:2023:HSR

[DD23b] Meimei Duan and Lijuan Duan. High spatial resolution remote sensing data classification method based on spectrum sharing. Scientific Programming, 2023(1):4356957:1– 4356957:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4356957.

Dogan:2018:NTD

[DDB^{+}18] Yunus Dogan, Feristah Dalkiliç, Derya Birant, Recep Alp Kut, and Reyat Yilmaz. Novel twodimensional visualization approaches for multivariate centroids of clustering algorithms. Scientific Programming, 2018(1): 9253295:1–9253295:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/9253295.

Dolgopolovas:2020:MGD

[DDJ20] Vladimiras Dolgopolovas, Valentina Dagiene, and Tatjana Jevsikova. Methodological guidelines for the design and integration of software learning objects for scientific programming education. Scientific Programming, 2020(1):6807515:1– 6807515:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/6807515.

Dolgopolovas:2014:PSC

[DDMS14] Vladimiras Dolgopolovas, Valentina Dagiene, Saulius Minkevicius, and Leonidas Sakalauskas. Python for scientific computing education: Modeling of queueing systems. Scientific Programming, 22(1):37–51, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Dolgopolovas:2015:TSC

[DDMS15] Vladimiras Dolgopolovas, Valentina Dagienė, Saulius Minkevičius, and Leonidas Sakalauskas. Teaching scientific computing: A model-centered approach to pipeline and parallel programming with C. Scientific Programming, 2015 (??):820803:1–820803:18, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
820803/.

Doolin:1999:JCL

[DDS99] David M. Doolin, Jack Dongarra, and Keith Seymour. JLAPACK — compiling LAPACK FOR-TRAN to Java. Scientific Programming, 7(2):111– 138, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=f277qlrwwjr5m4vxjywv%
26referrer=parent%26backto=
issue%2C4%2C8%3Bjournal%
2C7%2C9%3Blinkingpublicationresults%
2C1%2C1;   http://www.
netlib.org/utk/people/
JackDongarra/PAPERS/
f2jrep~1.pdf.  The

software is available on the World-Wide Web

                                        at http://www.cs.utk.
                                        edu/f2j/.

Dongarra:2001:RAS

[DEŁ01] Jack Dongarra, Victor Eijkhout, and Piotr Łuszczek. Recursive approach in sparse matrix LU factorization. Scientific Programming, 9(1): 51–60, 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

                                        //iospress.metapress.
                                        com/app/home/contribution.
                                        asp%3Fwasp=f2779jvvqg63jq64qwtm%
                                        26referrer=parent%26backto=
                                        issue%2C5%2C6%3Bjournal%
                                        2C2%2C9%3Blinkingpublicationresults%
                                        2C1%2C1;   http://www.
                                        netlib.org/netlib/utk/
                                        people/JackDongarra/
                                        PAPERS/recur-sparse-
                                        sciprog.pdf; http://
                                        www.netlib.org/utk/people/
                                        JackDongarra/PAPERS/
                                        rlu03.pdf.

Dennis:1996:SMF

[Den96] Jack B. Dennis. Static mapping of functional programs: an example in signal processing. Scientific Programming, 5(2): 121–135, Summer 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Dongarra:2001:ISB

[DEvdV01] Jack Dongarra, Victor Eijkhout, and Henk van der Vorst. An iterative solver benchmark. Scientific Programming, 9(4):223– 231, 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=64t4wprhwk589ellmv56%
26referrer=parent%26backto=
issue%2C3%2C4%3Bjournal%
2C3%2C12%3Blinkingpublicationresults%
2C1%2C1;   http://www.
netlib.org/utk/people/
JackDongarra/PAPERS/
sparse-bench.pdf.

Durillo:2014:SMO

[DF14] Juan Durillo and Thomas Fahringer. From singleto multi-objective autotuning of programs: Advantages and implications. Scientific Programming, 22(4):285–297, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Deboni:1996:HIF

[DFP96] Thomas Deboni, John Feo, and Doug Peters. Hybrid imperative and functional molecular mechanics application. Scientific Programming, 5 (2):97–109, Summer 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Dou:2023:IUA

[DFT^{+}23] Shaosong Dou, Zhiquan Feng, Jinglan Tian, Xue Fan, Ya Hou, and Xin Zhang. An intention understanding algorithm based on multimodal information fusion. Scientific Programming, 2023(1):8354015:1– 8354015:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/8354015.

Dongarra:2015:HPI

[DGH^{+}15] Jack Dongarra, Mark Gates, Azzam Haidar, Yulu Jia, Khairul Kabir, Piotr Luszczek, and Stanimire Tomov. HPC programming on Intel Many-Integrated-Core hardware with MAGMA port to Xeon Phi. Scientific Programming, 2015 (??):502593:1–502593:11, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://www.hindawi.
                                        com/journals/sp/2015/
                                        502593/.

Du:2017:MPV

[DGL^{+}17] Zhihui Du, Rong Ge, Victor W. Lee, Richard Vuduc, David A. Bader, and Ligang He. Modeling the power variability of core speed scaling on homogeneous multicore systems. Scientific Programming, 2017(1):8686971:1– 8686971:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/8686971.

Dinesh:2000:APS

[DHH00] T. B. Dinesh, Magne Haveraaen, and Jan Heering. An algebraic programming style for numerical software and its optimization. Scientific Programming, 8(4):247– 259, ???? 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=2dyrcfdc1p5ytlerkc3u%
26referrer=parent%26backto=
issue%2C4%2C5%3Bjournal%
2C3%2C9%3Blinkingpublicationresults%
2C1%2C1.

Deng:2023:PDF

[DHW^{+}23] Xinglong Deng, Suhui He, Qiumei Wu, Zongjie Weng, Minmin Yang, and Min Liu. Prenatal diagnosis of fetal cleft lip and palate with three-dimensional ultrasound information technology. Scientific Programming, 2023(1): 1531724:1–1531724:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/1531724.

Dinda:1999:SPH

[Din99] Peter A. Dinda. The statistical properties of host load. Scientific Programming, 7(3–4):211– 229, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

                                        //iospress.metapress.
                                        com/app/home/contribution.
                                        asp%3Fwasp=53f7mftrrm4r73yyrqau%
                                        26referrer=parent%26backto=
                                        issue%2C3%2C12%3Bjournal%
                                        2C6%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Deelman:2002:GC

[DK02] Ewa Deelman and Carl Kesselman. Grid computing. Scientific Programming, 10(2):101– 102, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

                                        //iospress.metapress.
                                        com/app/home/contribution.
                                        asp%3Fwasp=1f99bpyvlg7t46lx8ue3%
                                        26referrer=parent%26backto=
                                        issue%2C1%2C8%3Bjournal%
                                        2C1%2C12%3Blinkingpublicationresults%
                                        2C1%2C1.

Ding:2023:AGF

[DLP^{+}23] Yong Ding, Yuebin Liu, Cong Peng, Huanmei Wang, Yuqin Xu, Shengrong Jiao, Huan Xu, Yan Zhao, and Mingyu Liu. Analysis of glomerular filtration rate in ischemic cerebrovascular diseases under the magnetic resonance angiography image segmentation algorithm. Scientific Programming, 2023(1): 8399153:1–8399153:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8399153.

Dong:2023:MLB

[DM23] Jinjiang Dong and Jianhao Mao. Machine learning-based head computerized tomography imaging in diagnosis and surgery treatment of hypertension cerebral hemorrhage. Scientific Programming, 2023(1): 7880477:1–7880477:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7880477.

DAgostino:2023:HSS

[DMAC23] Daniele D’Agostino, Ivan Merelli, Marco Aldinucci, and Daniele Cesini. Hardware and software solutions for energy-efficient computing in scientific programming. Scientific Programming, 2023(1): 5514284:1–5514284:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5514284.

Diaz:2019:FQA

[DMV^{+}19] Gregorio Díaz, Hermenegilda Macià, Valentín Valero, Juan Boubeta-Puig, and Guadalupe Ortiz. Facilitating the quantitative analysis of complex events through a computational intelligence model-driven tool. Scientific Programming, 2019(1):2604148:1– 2604148:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/2604148.

Decyk:2004:SMI

[DN04] Viktor K. Decyk and Charles D. Norton. A simplified method for implementing run-time polymorphism in Fortran95. Scientific Programming, 12(1):45–55, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Decyk:1997:HEC

[DNS97] Viktor K. Decyk, Charles D.

Norton, and Bolesław K. Szymanski. How to express C++ concepts in Fortran90. Scientific Programming, 6(4):363–390, Winter 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=a5tkxhqqy9eefb7hwkak%
26referrer=parent%26backto=
issue%2C3%2C7%3Bjournal%
2C9%2C9%3Blinkingpublicationresults%
2C1%2C1.

Dong:2023:EQG

[Don23] Rui Dong. An effective quantum genetic algorithm based on drama resource mining using wireless sensing technology. Scientific Programming, 2023(1):4122372:1– 4122372:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4122372.

Dong:2016:TSO

[DPMB16] Wang Dong, Wang Peng, Li Meizi, and Zhang Bo. A topic space oriented user group discovering scheme in social network: A trust chain based interest measuring perspective. Scientific Programming, 2016(??): 4091323:1–4091323:25, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://www.hindawi.
                                        com/journals/sp/2016/
                                        4091323/.

Ding:2017:FSS

[DQQ17] Zhe Ding, Zhen Qin, and Zhiguang Qin. Frequent symptom sets identification from uncertain medical data in differentially private way. Scientific Programming, 2017(1):7545347:1– 7545347:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2017/7545347.

Dichev:2013:ERN

[DRL13] Kiril Dichev, Fergal Reid, and Alexey Lastovetsky. Efficient and reliable network tomography in heterogeneous networks using BitTorrent broadcasts and clustering algorithms. Scientific Programming, 21(3–4):79–92, ???? 2013. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Delgado:2018:AMO

[DRMG^{+}18] Elayne Rubio Delgado, Lisbeth Rodríguez-Mazahua, José Antonio Palet Guzmán, Jair Cervantes, José Luis Sánchez Cervantes, Silvestre Gustavo Peláez-Camarena, and Asdrúbal López-Chau. Analysis of medical opinions about the nonrealization of autopsies in a Mexican hospital using association rules and Bayesian networks. Scientific Programming, 2018(1): 4304017:1–4304017:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/4304017.

Deng:2018:LIV

[DRQ^{+}18] Fuhu Deng, Panlong Ren, Zhen Qin, Gu Huang, and Zhiguang Qin. Leveraging image visual features in content-based recommender system. Scientific Programming, 2018(1): 5497070:1–5497070:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/5497070.

Dummler:2012:CSM

[DRR12] Jörg Dümmler, Thomas Rauber, and Gudula Rünger. Combined scheduling and mapping for scalable computing with parallel tasks. Scientific Programming, 20 (1):45–67, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Dart:2014:SDN

[DRT^{+}14] Eli Dart, Lauren Rotman, Brian Tierney, Mary Hester, and Jason Zurawski. The Science DMZ: A network design pattern for data-intensive science. Scientific Programming, 22(2):173–185, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

DeSturler:1997:SPH

[DS97] Eric De Sturler and Volker Strumpen. Scientific programming with High Performance Fortran: a case study using the xHPF compiler. Scientific Programming, 6 (1):127–152, Spring 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Du:2023:APP

[DS23] Zhan Du and Jie Su. Analysis of the practice path of the flipped classroom model assisted by big data in English teaching. Scientific Programming, 2023(1):1831892:1– 1831892:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1831892.

Deelman:2005:PFM

[DSS^{+}05] Ewa Deelman, Gurmeet Singh, Mei-Hui Su, James Blythe, Yolanda Gil, Carl Kesselman, Gaurang Mehta, Karan Vahi, G. Bruce Berriman, John Good, Anastasia Laity, Joseph C. Jacob, and Daniel S. Katz. Pegasus: a framework for mapping complex scientific workflows onto distributed systems. Scientific Programming, 13(3):219–237, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Derby:1996:NLD

[DSZ96] Thomas Derby, Robert Schnabel, and Benjamin Zorn. New language design for prototyping numerical computation. Scientific Programming, 5 (4):279–300, Winter 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Desgagne:2000:PME

[DTV00] Michel Desgagné, Stephen Thomas, and Michel Valin. Performance of MC2 and the ECMWF IFS forecast model on the Fujitsu VPP700 and NEC SX-4M. Scientific Programming, 8(1): 23–30, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=h82chcapth0xynh5tw5w%
26referrer=parent%26backto=
issue%2C3%2C6%3Bjournal%
2C5%2C9%3Blinkingpublicationresults%
2C1%2C1.

Du:2023:CCS

[Du23a]  Xingliang Du.  Con-

struction cost simulation based on artificial intelligence and BIM. Scientific Programming, 2023(1):9744286:1– 9744286:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9744286.

Du:2023:REP

[Du23b] Xingliang Du. Research on engineering project management method based on BIM technology. Scientific Programming, 2023 (1):7230585:1–7230585:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7230585.

Du:2024:ENT

[Du24] Haibao Du. An English network teaching method supported by artificial intelligence technology and WBIETS system. Scientific Programming, 2024(1):8783899:1– 8783899:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8783899.

Duan:2023:RSM

[DW23] Sha Duan and Ziwei Wang. Research on the service mode of the university library based on data mining. Scientific Programming, 2023(1):5564326:1– 5564326:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5564326.

Deng:2020:PBS

[DWZ^{+}20] Fenglei Deng, Jian Wang, Bin Zhang, Chao Feng, Zhiyuan Jiang, and Yunfei Su. A pattern-based software testing framework for exploitability evaluation of metadata corruption vulnerabilities. Scientific Programming, 2020(1):8883746:1– 8883746:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/8883746.

Daihong:2023:FER

[DyLJ23] Jiang Daihong, Hu yuanzheng, Dai Lei, and Peng Jin. Facial expression recognition based on attention mechanism. Scientific Programming, 2023(1): 6624251:1–6624251:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6624251.

Du:2023:ACC

[DYT23] Songbo Du, Fang Yang, and Xuedong Tian. Ancient Chinese character image retrieval based on dual hesitant fuzzy sets. Scientific Programming, 2023(1):6621037:1– 6621037:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6621037.

Dong:2023:AIS

[DYW^{+}23] Bo Dong, Shi Ying, Bingming Wang, Guoli Cheng, and Jiangyi Geng. Analysing the impact of scaling out SaaS software on response time. Scientific Programming, 2023(1):6629867:1– 6629867:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6629867.

Dai:2024:KMC

[DYYW24] Yawen Dai, Guanghui Yuan, Zhaoyuan Yang, and Bin Wang. Kmodes clustering algorithm based on weighted overlap distance and its application in intrusion detection. Scientific Programming, 2024(1): 9972589:1–9972589:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9972589.

Ding:2018:PBR

[DZ18] Xuefeng Ding and Junhui Zhong. Power battery recycling mode selection using an extended MULTI-MOORA method. Scientific Programming, 2018 (1):7675094:1–7675094:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/7675094.

Ding:2013:CDF

[DZKS13] Wei Ding, Yuanrui Zhang, Mahmut Kandemir, and Seung Woo Son. Compilerdirected file layout optimization for hierarchical storage systems. Scientific Programming, 21 (3–4):65–78, ???? 2013. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ding:2020:NLZ

[DZL^{+}20] Xue-Feng Ding, Li-Xia Zhu, Mei-Shun Lu, Qi Wang, and Yi-Qi Feng. A novel linguistic Z-number QUALIFLEX method and its application to large group emergency decision making. Scientific Programming, 2020(1): 1631869:1–1631869:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/1631869.

Du:2023:SEB

[DZL23] Yufeng Du, Quan Zhao, and Xiaochun Lu. Semantic extraction of basketball game video combining domain knowledge and in-depth features. Scientific Programming, 2023(1):9080120:1– 9080120:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9080120.

Dong:2024:NLH

[DZL^{+}24] Yongle Dong, Fan Zhang, Xuan Li, Lifang Zhang, Jia Yu, Yongmei Mao, and Guanglong Jiang. Nonlinear load harmonic prediction method based on power distribution Internet of Things. Scientific Programming, 2024 (1):9978900:1–9978900:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9978900.

Edjlali:1997:RTC

[EAS^{+}97] Guy Edjlali, Gagan Agrawal, Alan Sussman, Jim Humphries, and Joel Saltz. Runtime and compiler support for programming in adaptive parallel environments. Scientific Programming, 6(2):215–227, Summer 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Esseissah:2024:TSI

[EBDB24] Mohamed S. Esseissah, Ashraf Bhery, Sameh S. Daoud, and Hatem M. Bahig. Three strategies for improving shortest vector enumeration using GPUs. Scientific Programming, 2024(1): 8852497:1–8852497:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8852497.

Liu:2020:ICA

[eLA20] Jun e Liu and Feng-Ping An. Image classification algorithm based on deep learning-kernel function. Scientific Programming, 2020(1):7607612:1– 7607612:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/7607612.

Enriquez:2019:RCS

[EMTCA^{+}19] J. G. Enríquez, L. Morales-Trujillo, Fernando Calle-Alonso, F. J. Domínguez-Mayo, and J. M. Lucas-Rodríguez. Recommendation and classification systems: a systematic mapping study. Scientific Programming, 2019(1): 8043905:1–8043905:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/8043905.

Escalada:2017:EPA

[EOS17] Javier Escalada, Francisco Ortin, and Ted Scully. An efficient platform for the automatic extraction of patterns in native code. Scientific Programming, 2017(1): 3273891:1–3273891:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/3273891.

Edwards:2012:MPP

[ESP^{+}12] H. Carter Edwards, Daniel Sunderland, Vicki Porter, Chris Amsler, and Sam Mish. Manycore performanceportability: Kokkos multidimensional array library. Scientific Programming, 20(2):89–114, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Eide:1999:FIC

[ESSL99] Eric Eide, James L. Simister, Tim Stack, and Jay Lepreau. Flexible IDL compilation for complex communication patterns. Scientific Programming, 7(3–4):275– 287, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.

                                        asp%3Fwasp=53f7mftrrm4r73yyrqau%
                                        26referrer=parent%26backto=
                                        issue%2C7%2C12%3Bjournal%
                                        2C6%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Faisal:2024:CBA

[FA24] Mohammad Faisal and Sa’ed Abed. Clusterbased antiphishing (CAP) model for smart phones. Scientific Programming, 2024(1):9957323:1–9957323:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9957323.

Feng:2020:TEE

[FAA^{+}20] Chong Feng, Muhammad Adnan, Arshad Ahmad, Ayaz Ullah, and Habib Ullah Khan. Towards energy-efficient framework for IoT big data healthcare solutions. Scientific Programming, 2020(1):7063681:1– 7063681:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/7063681.

Ferner:1999:ACS

[FB99] Clayton S. Ferner and Robert G. Babb II. Automatic choice of scheduling heuristics for parallel/distributed computing. Scientific Programming, 7(1):47–65, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=64cr5a4mg33tuhcbdr02%
26referrer=parent%26backto=
issue%2C4%2C7%3Bjournal%
2C8%2C9%3Blinkingpublicationresults%
2C1%2C1.

Faure:2001:CGS

[FC01] Christèle Faure and Isabelle Charpentier. Comparing global strategies for coding adjoints. Scientific Programming, 9 (1):1–10, 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=f2779jvvqg63jq64qwtm%
26referrer=parent%26backto=
issue%2C1%2C6%3Bjournal%
2C2%2C9%3Blinkingpublicationresults%
2C1%2C1.

Feng:2022:RND

[FC22] Hui Feng and Guozhen Chen. RETRACTION: a novel data visualization model based on autoencoder using big data analysis and distributed processing technology. Scientific Programming, 2022 (1):9763849:1–9763849:??, ???? 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL https:/

                                        /onlinelibrary.wiley.
                                        com/doi/epdf/10.1155/
                                        2024/9763849.

Fu:2023:EIL

[FCSL23] Anmin Fu, Yulin Cai, Tao Sun, and Feng Li. Estimating the impact of land cover change on soil erosion using remote sensing and GIS data by USLE model and scenario design. Scientific Programming, 2023(1):6633428:1– 6633428:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6633428.

Fan:2017:NHS

[FCZ^{+}17] Xiaoping Fan, Zhijie Chen, Liangkun Zhu, Zhifang Liao, and Bencai Fu. A novel hybrid similarity calculation model. Scientific Programming, 2017(1): 4379141:1–4379141:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2017/4379141.

Fan:2019:SDP

[FDY^{+}19] Guisheng Fan, Xuyang Diao, Huiqun Yu, Kang Yang, and Liqiong Chen. Software defect prediction via attention-based recurrent neural network. Scientific Programming, 2019(1):6230953:1– 6230953:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/6230953.

Feng:2023:RMM

[Fen23a]  Chunmei Feng.  Re-

search on mathematical model of smart service for the elderly in smalland medium-sized cities based on image processing. Scientific Programming, 2023(1):1023187:1– 1023187:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1023187.

Feng:2023:RHE

[Fen23b] Liu Feng. Research on higher education evaluation and decision-making based on data mining. Scientific Programming, 2023(1):6195067:1– 6195067:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

                                        wiley.com/doi/epdf/10.
                                        1155/2021/6195067.

Fdez-Glez:2016:WNF

[FGROL^{+}16] J. Fdez-Glez, D. Ruano-Ordás, R. Laza, J. R. Méndez, R. Pavón, and F. Fdez-Riverola. WSF2: A novel framework for filtering Web spam. Scientific Programming, 2016 (??):6091385:1–6091385:18, ???? 2016. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://www.hindawi.
                                        com/journals/sp/2016/
                                        6091385/.

Fernandez:2017:DSS

[FGSCR^{+}17] Eduardo Fernandez, Claudia Gómez-Santillán, Laura Cruz-Reyes, Nelson Rangel-Valdez, and Shulamith Bastiani. Design and solution of a surrogate model for portfolio optimization based on project ranking. Scientific Programming, 2017(1):1083062:1– 1083062:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2017/1083062.

Feng:2023:HSA

[FH23] Xiaochun Feng and Xiangpei Hu. A heuristic solution approach to order batching and sequencing for manual picking and packing lines considering fatiguing effect. Scientific Programming, 2023(1):8863391:1– 8863391:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8863391.

Fu:2022:MRI

[FHC^{+}22] Qimao Fu, Chuizhi Huang, Yan Chen, Nailong Jia, Jinghui Huang, and Changkun Lin. Magnetic resonance imaging image under low-rank matrix denoising algorithm in the diagnosis and evaluation of tibial plateau fracture combined with meniscus injury. Scientific Programming, 2022(1):6329020:1– 6329020:??, 2022. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6329020.

Fang:2023:IFR

[FHS^{+}23] Bing Fang, Enpeng Hu, Junyang Shen, Jingwen Zhang, and Yang Chen. Implicit feedback recommendation method based on user-generated content. Scientific Programming, 2023(1):3982270:1– 3982270:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/3982270.

Feng:2023:ECC

[FHT23] Jing Ge Feng, Ye Ping He, and Qiu Ming Tao. Evaluation of compilers’ capability of automatic vectorization based on source code analysis. Scientific Programming, 2023(1):3264624:1– 3264624:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/3264624.

Filippone:2014:BR

[Fil14] Salvatore Filippone. Book review. Scientific Programming, 22(1):53–55, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Fan:2020:HCP

[FLC^{+}20] Shixiong Fan, Xingwei Liu, Ying Chen, Zhifang Liao, Yiqi Zhao, Huimin Luo, and Haiwei Fan. How to construct a power knowledge graph with dispatching data? Scientific Programming, 2020(1):8842463:1– 8842463:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8842463.

Fang:2023:CDP

[FLR23] Degen Fang, Chunlei Li, and Yanhong Ren. Classification and diagnosis of pulmonary nodules in thoracic surgery using CT image segmentation algorithm. Scientific Programming, 2023(1): 3367677:1–3367677:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3367677.

Fei:2017:SRR

[FLS17] Hongying Fei, Qian Li, and Dan Sun. A survey of recent research on optimization models and algorithms for operations management from the process view. Scientific Programming, 2017(1):7219656:1– 7219656:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/7219656.

Fu:2023:IIB

[FLY^{+}23] Xiaoli Fu, Minxiang Li, Mantian Yin, Qing Li, and Ying Chen. IMRT images based on swarm intelligence algorithm in the treatment of nasopharyngeal carcinoma and pain. Scientific Programming, 2023(1):7007169:1– 7007169:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/7007169.

Faisal:2016:AEO

[FMA16] Mohammed Faisal, Hassan Mathkour, and Mansour Alsulaiman. AntStar: Enhancing optimization problems by integrating an ant system and algorithm. Scientific Programming, 2016(??): 5136327:1–5136327:12, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://www.hindawi.
                                        com/journals/sp/2016/
                                        5136327/.

Fu:2023:IMA

[FMG^{+}23] Xiaohua Fu, Fei Miao, Jianbin Guo, Weijuan He, Yuying Deng, Yushan Wang, Xiaxiang Zeng, Wenbing Wang, and Yunfeng Liu. Image mosaic algorithm-based analysis of effects of low-dose Xrays on blood immune cells. Scientific Programming, 2023(1):2888567:1– 2888567:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2888567.

Fursin:2014:CMT

[FML^{+}14] Grigori Fursin, Renato Miceli, Anton Lokhmotov, Michael Gerndt, Marc Baboulin, Allen D. Malony, Zbigniew Chamski, Diego Novillo, and Davide Del Vento. Collective mind: Towards practical and collaborative auto-tuning. Scientific Programming, 22 (4):309–329, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Fitzgerald:1996:UPA

[FO96] Steven M. Fitzgerald and Rodney R. Oldehoeft. Update-in-place analysis for true multidimensional arrays. Scientific Programming, 5(2):147–160, Summer 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Foster:1992:PPP

[FOT92] I. Foster, R. Olson, and S. Tuecke. Productive parallel programming: The PCN approach. Scientific Programming, 1 (1):51–66, Fall 1992. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ford:1995:PPA

[FP95] R. W. Ford and D. I. A. Poll. Parallel processing approach to transition prediction for laminar flow control system design. Scientific Programming, 4(3):203–217, Fall 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Fahringer:2000:PPE

[FP00a] T. Fahringer and A. Požgaj. P^{3}T+: a performance estimator for distributed and parallel programs. Scientific Programming, 8 (2):73–93, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

                                        //iospress.metapress.
                                        com/app/home/contribution.
                                        asp%3Fwasp=253x52trrm4r87tkuw1h%
                                        26referrer=parent%26backto=
                                        issue%2C2%2C3%3Bjournal%
                                        2C4%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Francioni:2000:DSH

[FP00b] Joan M. Francioni and Cherri M. Pancake. A debugging standard for high-performance computing. Scientific Programming, 8(2):95–108, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL http://iospress.
metapress.com/app/home/
contribution.asp%3Fwasp=
253x52trrm4r87tkuw1h%
26referrer=parent%26backto=
issue%2C3%2C3%3Bjournal%
2C4%2C9%3Blinkingpublicationresults%
2C1%2C1.

Fatoohi:2001:DID

[FS01] Rod Fatoohi and Lance Smith. Development and implementation of a distributed-object jobexecution environment. Scientific Programming, 9(1):27–37, ???? 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

http://iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=f2779jvvqg63jq64qwtm%
26referrer=parent%26backto=
issue%2C3%2C6%3Bjournal%
2C2%2C9%3Blinkingpublicationresults%
2C1%2C1.

Feng:2024:IDH

[FS24] Jiangfan Feng and Wenzheng Sun. Improved deep hashing with scalable interblock for tourist image retrieval. Scientific Programming, 2024(1): 9937061:1–9937061:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9937061.

Fang:2017:DWD

[FSLG17] Bin Fang, Fuchun Sun, Huaping Liu, and Di Guo. Development of a wearable device for motion capturing based on magnetic and inertial measurement units. Scientific Programming, 2017(1): 7594763:1–7594763:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2017/7594763.

Faisal:2017:TED

[FSS^{+}17] Sidra Faisal, Mansoor Sarwar, Khurram Shahzad, Shahzad Sarwar, Waqar Jaffry, and Muhammad Murtaza Yousaf. Temporal and evolving data warehouse design. Scientific Programming, 2017(1):7392349:1– 7392349:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2017/7392349.

Fang:2014:API

[FSV14] Jianbin Fang, Henk Sips, and Ana Lucia Varbanescu. Aristotle: A performance impact indicator for the OpenCL kernels using local memory. Scientific Programming, 22(3):239–257, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Tu:2023:SMM

[fTPsLZ23] Li fen Tu, Qi Peng, Chun sheng Li, and Aiqun Zhang. 2D in situ method for measuring plant leaf area with camera correction and background color calibration. Scientific Programming, 2023(1):6650099:1– 6650099:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6650099.

Fahad:2020:ACO

[FTS^{+}20] Labiba Gillani Fahad, Syed Fahad Tahir, Waseem Shahzad, Mehdi Hassan, Hani Alquhayz, and Rabia Hassan. Ant colony optimization-based streaming feature selection: an application to the medical image diagnosis. Scientific Programming, 2020(1):1064934:1– 1064934:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/1064934.

Fang:2023:TTC

[FWX23] Lei Fang, Qiang Wei, and Cheng Jian Xu. Technical and tactical command decision algorithm of football matches based on big data and neural network. Scientific Programming, 2023(1):5544071:1– 5544071:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5544071.

Fan:2023:IPM

[FX23] Bingqian Fan and Xuanxuan Xing. Intelligent prediction method of building energy consumption based on deep learning. Scientific Programming, 2023(1):3323316:1– 3323316:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/3323316.

Xu:2017:RAO

[fXhCL17] Xiao feng Xu, Wei hong Chang, and Jing Liu. Resource allocation optimization model of collaborative logistics network based on bilevel programming. Scientific Programming, 2017(1):4587098:1– 4587098:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/4587098.

Fu:2017:WWA

[FXYY17] Deqiang Fu, Yanyan Xu, Haoran Yu, and Boyang Yang. WASTK: a weighted abstract syntax tree kernel method for source code plagiarism detection. Scientific Programming, 2017(1): 7809047:1–7809047:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/7809047.

Fei:2020:ADN

[FYC^{+}20] Jiaxuan Fei, Qigui Yao, Mingliang Chen, Xiangqun Wang, and Jie Fan. The abnormal detection for network traffic of power IoT based on device portrait. Scientific Programming, 2020(1):8872482:1– 8872482:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

                                        wiley.com/doi/epdf/10.
                                        1155/2020/8872482.

Fan:2019:SUS

[FYH^{+}19] Yong Fan, Juhui Yao, Zongyi He, Biao He, and Minmin Li. Study of urban system spatial interaction based on microblog data: a case of Huaihe River Basin, China. Scientific Programming, 2019(1): 2074329:1–2074329:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2019/2074329.

Fu:2023:IMC

[FYH23] Yi Fu, Min Yang, and Di Han. Interactive marketing e-commerce recommendation system driven by big data technology. Scientific Programming, 2023(1):3873059:1– 3873059:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/3873059.

Fan:2023:CAV

[FYY23] Chao Fan, Zhihui Yang, and Yuyi Yuan. Cluster analysis and visualization for the Legend of the Condor Heroes based on social network. Scientific Programming, 2023(1):9439583:1– 9439583:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9439583.

Fei:2020:DLS

[FYZ^{+}20] Rong Fei, Quanzhu Yao, Yuanbo Zhu, Qingzheng Xu, Aimin Li, Haozheng Wu, and Bo Hu. Deep learning structure for cross-domain sentiment classification based on improved cross entropy and weight. Scientific Programming, 2020(1): 3810261:1–3810261:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/3810261.

Fu:2023:PTD

[FZ23] JinFeng Fu and Hongli Zhang. Personality trait detection based on ASM localization and deep learning. Scientific Programming, 2023(1): 5675917:1–5675917:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5675917.

Fan:2023:MCE

[FZH23] Kun Fan, Ting Zhang, and Weihong He. Multilevel clustering-evolutionary random support vector machine cluster algorithmbased functional magnetic resonance imaging in diagnosing cerebral ischemic stroke. Scientific Programming, 2023(1): 3729379:1–3729379:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/3729379.

Fu:2018:EAV

[FZW^{+}18] Xiong Fu, Qing Zhao, Junchang Wang, Lin Zhang, and Lei Qiao. Energy-aware VM initial placement strategy based on BPSO in cloud computing. Scientific Programming, 2018(1): 9471356:1–9471356:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2018/9471356.

Fang:2023:STN

[FZW23] Jie Fang, QingBiao Zhou, and Shuxia Wang. Segmentation technology of nucleus image based on U-Net network. Scientific Programming, 2023(1):

1892497:1–1892497:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1892497.

Gustafson:1996:MPS

[GA96] John Gustafson and Srinivas Aluru. Massively parallel searching for better algorithms or how to do a cross product with five multiplications. Scientific Programming, 5(3): 203–217, Fall 1996. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Garvey:2018:SAP

[GA18] Joseph D. Garvey and Tarek S. Abdelrahman. A strategy for automatic performance tuning of stencil computations on GPUs. Scientific Programming, 2018(1): 6093054:1–6093054:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/6093054.

Gafsi:2020:ICB

[GAH^{+}20] Mohamed Gafsi, Nessrine Abbassi, Mohamed Ali Hajjaji, Jihene Malek, and Abdellatif Mtibaa. Improved chaos-based cryptosystem for medical image encryption and decryption. Scientific Programming, 2020(1): 6612390:1–6612390:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/6612390.

Ghaderzadeh:2023:MLD

[GAH^{+}23] Mustafa Ghaderzadeh, Farkhondeh Asadi, Azamossadat Hosseini, Davood Bashash, Hassan Abolghasemi, and Arash Roshanpour. Machine learning in detection and classification of leukemia using smear blood images: a systematic review. Scientific Programming, 2023(1):9933481:1– 9933481:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9933481.

Gerber:1994:NCS

[GB94] Stefan Gerber and Helmar Burkhart. Numbercrunching software and the input problem: guidelines and a case study. Scientific Programming, 3(1):1–11, Spring 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gundy-Burlet:2010:BR

[GB10] Karen Gundy-Burlet. Book review. Scientific Programming, 18(3–4):219– 220, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ghayekhloo:2015:PSI

[GB15] Samira Ghayekhloo and Zeki Bayram. Prefiltering strategy to improve performance of Semantic Web service discovery. Scientific Programming, 2015(??):576463:1– 576463:15, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
576463/.

Gerstenberger:2014:EHS

[GBH14] Robert Gerstenberger, Maciej Besta, and Torsten Hoefler. Enabling highlyscalable remote memory access programming with MPI-3 One Sided. Scientific Programming, 22 (2):75–91, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Guclu:2020:NSE

[GBH20] Mehmet Guclu, Cigdem Bakir, and Veli Hakkoymaz. A new scalable and expandable access control model for distributed database systems in data security. Scientific Programming, 2020(1):8875069:1– 8875069:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8875069.

George:1994:PCS

[GBJ94] William George, Ralph G. Brickner, and S. Lennart Johnsson. Polyshift communications software for the Connection Machine systems CM-2 and CM-200. Scientific Programming, 3(1):83–99, Spring 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gunzinger:1996:PEH

[GBK^{+}96] A. Gunzinger, B. Baumle, M. Klebl, M. Kocheisen, P. Kohler, R. Morel, U. Muller, and M. Rosenthal. Programming environment for a highperformance parallel supercomputer with intelligent communication. Scientific Programming, 5 (1):25–32, Spring 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gu:2023:NAM

[GCB^{+}23] Tianlong Gu, Hongliang Chen, Chenzhong Bin, Liang Chang, and Wei Chen. Neighborhood attentional memory networks for recommendation systems. Scientific Programming, 2023(1): 8880331:1–8880331:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8880331.

Gao:2023:LRC

[GDA^{+}23] Wenyun Gao, Sheng Dai, Stanley Ebhohimhen Abhadiomhen, Wei He, and Xinghui Yin. Low rank correlation representation and clustering. Scientific Programming, 2023(1):6639582:1– 6639582:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6639582.

Garcia-Diaz:2019:NAD

[GDNVSMM19] Vicente García-Díaz, Edward R. Núñez-Valdez, Vijender K. Solanki, and Carlos E. Montenegro-Marin. Novel advances in the development of machine learning solutions for scientific programming. Scientific Programming, 2019(1):7896462:1– 7896462:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2019/7896462.

Ge:2023:SPC

[Ge23]  Xianyun Ge.  Smart

payment contract mechanism based on blockchain smart contract mechanism. Scientific Programming, 2023(1):3988070:1– 3988070:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/3988070.

Gerndt:2002:GEP

[Ger02] Michael Gerndt. Guesteditorial: PADDA 2001. Scientific Programming, 10(1):1–2, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

                                        //iospress.metapress.
                                        com/app/home/contribution.
                                        asp%3Fwasp=9ejnuvwuvby9737jte27%
                                        26referrer=parent%26backto=
                                        issue%2C1%2C9%3Bjournal%
                                        2C2%2C12%3Blinkingpublicationresults%
                                        2C1%2C1.

GautierdeLahaut:1995:SAC

[GG95] Damien Gautier de Lahaut and Cecile Germain. Static approach for compiling communications in parallel scientific programs. Scientific Programming, 4(4):291– ??, Winter 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gravvanis:2005:PPC

[GG05] George A. Gravvanis and Konstantinos M. Giannoutakis. Parallel preconditioned conjugate gradient square method based on normalized approximate inverses. Scientific Programming, 13(2): 79–91, ???? 2005. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Goldman:2004:IBS

[GGJ04] Jacki P. Goldman, William J.

Gullick, and Colin G. Johnson. Individualbased simulation of the clustering behaviour of epidermal growth factor receptors. Scientific Programming, 12(1):25–43, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Getov:1999:MLP

[GGMS99] Vladimir Getov, Paul Gray, Sava Mintchev, and Vaidy Sunderam. Multilanguage programming environments for high performance Java computing. Scientific Programming, 7(2):139–146, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=f277qlrwwjr5m4vxjywv%
26referrer=parent%26backto=
issue%2C5%2C8%3Bjournal%
2C7%2C9%3Blinkingpublicationresults%
2C1%2C1.

Gschwind:2009:HPC

[GGP09] Michael Gschwind, Fred Gustavson, and Jan F. Prins. High performance computing with the Cell Broadband Engine. Scientific Programming, 17 (1–2):1–2, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Guo:2023:DLD

[GGSL23] Xiangyan Guo, Hui Gao, Xiaofang Sun, and Surong Li. Deep learning for detection of colonic polyps from computed tomography colonoscopy images combined with colonoscopy. Scientific Programming, 2023(1): 1238805:1–1238805:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2021/1238805.

Greer:2002:SCI

[GHH^{+}02] Bruce Greer, John Harrison, Greg Henry, et al. Scientific computing on the Itanium(R) processor. Scientific Programming, 10(4):329–337, 2002. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gaidamour:2012:DCF

[GHST12] Jérémie Gaidamour, Jonathan Hu, Chris Siefert, and Ray Tuminaro. Design considerations for a flexible multigrid preconditioning library. Scientific Programming, 20(3):223– 239, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Grant:2000:CFP

[GHW00] Philip W. Grant, Magne Haveraaen, and Michael F. Webster. Coordinate free programming of computational fluid dynamics problems. Scientific Programming, 8(4):211–230, ???? 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=2dyrcfdc1p5ytlerkc3u%
26referrer=parent%26backto=

                                        issue%2C2%2C5%3Bjournal%
                                        2C3%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Glikman:1995:PAM

[GIKP95] Eli Glikman, Ludmila Ioffe, Itzhak Kelson, and Shlomit S. Pinter. Parallel algorithms for molecular dynamics simulation of irradiation effects in crystals. Scientific Programming, 4(3):185– ??, Fall 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gil:2009:DKD

[Gil09] Yolanda Gil. From data to knowledge to discoveries: Artificial intelligence and scientific workflows. Scientific Programming, 17(3):231–246, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gul:2020:PSO

[GKK^{+}20] Noor Gul, Muhammad Sajjad Khan, Su Min Kim, Marc St-Hilaire, Ihsan Ullah, and Junsu Kim. Particle swarm optimization in the presence of malicious users in cognitive IoT networks with data. Scientific Programming, 2020(1):8844083:1– 8844083:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8844083.

Gennart:1996:GVM

[GKL^{+}96] B. A. Gennart, B. Krummenacher, L. Landron, R. D. Hersch, B. Saugy, J.-C. Hadorn, and D. Muller. Giga view multiprocessor multidisk image server. Scientific Programming, 5(1):3–13, Spring 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gerbessiotis:2004:RMA

[GL04] Alexandros V. Gerbessiotis and Seung-Yeop Lee. Remote memory access: a case for portable, efficient and library independent parallel programming. Scientific Programming, 12(3):169–183, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gu:2023:AMK

[GL23] Yi Gu and Bo Li. An autocontouring method for the kidneys using an adaptive weighted multikernel support vector machines. Scientific Programming, 2023(1): 1885512:1–1885512:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1885512.

Girard:2017:AGH

[GLBB17] Simon R. Girard, Vincent Legault, Guy Bois, and Jean-François Boland. Avionics graphics hardware performance prediction with machine learning. Scientific Programming, 2017(1):9195845:1– 9195845:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/9195845.

Geng:2024:DLB

[GLD^{+}24] Yaqin Geng, Ting Liu, Yu Ding, Wei Liu, Jiayi Ye, Linlin Hu, and Li Ruan. Deep learningbased self-efficacy X-ray images in the evaluation of rheumatoid arthritis combined with osteoporosis nursing. Scientific Programming, 2024(1): 9959617:1–9959617:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9959617.

Gou:2023:MCS

[GLH23] Zhinan Gou, Yan Li, and Zheng Huo. A method for constructing supervised time topic model based on variational autoencoder. Scientific Programming, 2023(1):6623689:1– 6623689:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6623689.

Gao:2023:ROU

[GLW^{+}23] Jiajing Gao, Haolin Li, Jingwen Wu, Junyan Lyu, Zheyi Tan, and Zhufan Jin. Routing optimisation of urban medical waste recycling network considering differentiated collection strategy and time windows. Scientific Programming, 2023(1):5523910:1– 5523910:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5523910.

Gropp:2014:SIS

[GM14] William Gropp and Satoshi Matsuoka. Special issue: SC13 — The International Conference for High Performance Computing, Networking, Storage and Analysis. Scientific Programming, 22 (2):57–58, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gao:2015:LPM

[GMCT15] Fei Gao, Katarzyna Musial, Colin Cooper, and Sophia Tsoka. Link prediction methods and their accuracy for different social networks and network metrics. Scientific Programming, 2015 (??):172879:1–172879:13, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://www.hindawi.
                                        com/journals/sp/2015/
                                        172879/.

Gudakahriz:2020:OTC

[GMM20] Sajjad Jahanbakhsh Gudakahriz, Amir Masoud Eftekhari Moghadam, and Fariborz Mahmoudi. Opinion texts clustering using manifold learning based on sentiment and semantics analysis. Scientific Programming, 2020(1):7842631:1– 7842631:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/7842631.

Gobinathan:2023:NMS

[GMS^{+}23] B. Gobinathan, M. A. Mukunthan, S. Surendran, K. Somasundaram, Syed Abdul Moeed, P. Niranjan, V. Gouthami, G. Ashmitha, Gouse Baig Mohammad, V. K. Shanmuganathan, Yuvaraj Natarajan, K. Srihari, and Venkatesa Prabhu Sundramurthy. A novel method to solve real time security issues in software industry using advanced cryptographic techniques. Scientific Programming, 2023(1): 3611182:1–3611182:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3611182.

Guo:2022:ITB

[GN22] Jiayi Guo and Shah Nazir.

Internet of Things based intelligent techniques in workable computing: an overview. Scientific Programming, 2022(1): 6805104:1–6805104:??, 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6805104.

Guan:2023:SBU

[GNKuR23] Keqing Guan, Shah Nazir, Xianli Kong, and Sadaqat ur Rehman. Software birthmark usability for source code transformation using machine learning algorithms. Scientific Programming, 2023(1):5547766:1–5547766:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5547766.

Gogodze:2023:RMM

[Gog23] Joseph Gogodze. Ranking methods for multicriteria decision-making: Application to benchmarking of solvers and problems. Scientific Programming, 2023(1):5513860:1– 5513860:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5513860.

Guan:2017:DIA

[GPP^{+}17] Fei Guan, Long Peng, Luc Perneel, Hasan Fayyad-Kazan, and Martin Timmerman. A design that incorporates adaptive reservation into mixedcriticality systems. Scientific Programming, 2017 (1):3403685:1–3403685:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2017/3403685.

Guo:2021:PAP

[GPX^{+}21] Haifeng Guo, Ke Peng, Xiaolei Xu, Shuai Tao, and Zhen Wu. The prediction analysis of peer-to-peer lending platforms default risk based on comparative models. Scientific Programming, 2021(1):8816419:1– 8816419:??, 2021. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8816419.

Golumbic:1993:ISA

[GR93] Martin Charles Golumbic and Vladimir Rainish. Instruction scheduling across control flow. Scientific Programming, 2(3): 1–5, Fall 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gupta:2023:IPT

[GRB23] Mansi Gupta, Kumar Rajnish, and Vandana Bhattacharjee. Impact of parameter tuning for optimizing deep neural network models for predicting software faults. Scientific Programming, 2023(1): 6662932:1–6662932:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6662932.

Gomes:2003:POG

[GRC03] Maria Cecília Gomes, Omer F. Rana, and José C. Cunha. Pattern operators for grid environments. Scientific Programming, 11(3):237–261, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Germain-Renaud:1999:JBC

[GRN99] Cécile Germain-Renaud and Vincent Néri. Javabased coupling for parallel predictive-adaptive domain decomposition. Scientific Programming, 7 (2):185–189, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        http://iospress.metapress.
                                        com/app/home/contribution.
                                        asp%3Fwasp=f277qlrwwjr5m4vxjywv%
                                        26referrer=parent%26backto=
                                        issue%2C8%2C8%3Bjournal%
                                        2C7%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Gonzalez:2003:TPM

[GRR^{+}03] J. A. González, C. Rodríguez, G. Rodríguez, et al. A tool for performance modeling of parallel programs. Scientific Programming, 11(3):191–198, 2003. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ge:2023:GJE

[GSL23] Xiujuan Ge, Xiaopeng Song, and Linyan Li. Global joint entropy algorithm-based evaluation of dexmedetomidine combined local anesthetics in brachial plexus block from ultrasound images. Scientific Programming, 2023(1):4327596:1– 4327596:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4327596.

Gregor:2003:DPL

[GSM03] Douglas Gregor, Sibylle Schupp, and David R. Musser. Design patterns for library optimization. Scientific Programming, 11(4):309–320, ???? 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Guo:2023:PEM

[Guo23a] Chengcheng Guo. Prediction and evaluation model of physical training for volleyball players’ effect based on grey Markov theory. Scientific Programming, 2023(1): 6147032:1–6147032:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

                                        wiley.com/doi/epdf/10.
                                        1155/2021/6147032.

Guo:2023:IBI

[Guo23b] Yuheng Guo. Impact on biometric identification systems of COVID-19. Scientific Programming, 2023(1):3225687:1– 3225687:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/3225687.

Gustafsson:1993:OOI

[Gus93] Kjell Gustafsson. Objectoriented implementation of software for solving ordinary differential equations. Scientific Programming, 2(4):217–225, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gaurav:2011:UGC

[GW11] Gaurav and Steven F. Wojtkiewicz. Use of GPU computing for uncertainty quantification in computational mechanics: a case study. Scientific Programming, 19(4):199–212, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gan:2023:HMB

[GWH^{+}23] Y. S. Gan, Shih-Yuan Wang, Chieh-En Huang, Yi-Chen Hsieh, Hsiang-Yu Wang, Wen-Hung Lin, Shing-Nam Chong, and Sze-Teng Liong. How many bedrooms do you need? a realestate recommender system from architectural floor plan images. Scientific Programming, 2023 (1):9914557:1–9914557:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9914557.

Gao:2016:VCD

[GWZ16] Xianming Gao, Baosheng Wang, and Xiaozhe Zhang. VR-Cluster: Dynamic migration for resource fragmentation problem in virtual router platform. Scientific Programming, 2016(??):3976965:1–3976965:14, ???? 2016. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2016/
3976965/.

Gao:2023:SPB

[GWZ23a] Ya Gao, Rong Wang, and Enmin Zhou. Stock prediction based on optimized LSTM and GRU models. Scientific Programming, 2023(1):4055281:1– 4055281:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/4055281.

Gao:2023:NNA

[GWZ23b] Zhuorong Gao, Guangsheng Wang, and Zhi Zhang. Neural network algorithm MRI images for analysis of influencing factors for patellar dislocation in exercise. Scientific Programming, 2023(1):1348922:1– 1348922:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/1348922.

Guo:2020:TCT

[GXY^{+}20] Wenping Guo, Zhuoming Xu, Xijian Ye, Shiqing Zhang, Xiaoming Zhao, and Xue Li. A time-critical topic model for predicting the survival time of sepsis patients. Scientific Programming, 2020(1): 8884539:1–8884539:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/8884539.

Xue:2017:HPS

[gXySR^{+}17] Yun gang Xue, Hua you Su, Ju Ren, Mei Wen, Chun yuan Zhang, and Li quan Xiao. A highly parallel and scalable motion estimation algorithm with GPU for HEVC. Scientific Programming, 2017(1): 1431574:1–1431574:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/1431574.

Guo:2023:AED

[GZD23] Li Guo, Kunlin Zhu, and Ruijun Duan. Analysis of economic development trend in postepidemic era based on improved clustering algorithm. Scientific Programming, 2023(1): 4467001:1–4467001:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4467001.

Hassan:2022:SIF

[HAA^{+}22] Abid Hassan, Md. Iftekhar Ali, Rifat Ahammed, Mohammad Monirujjaman Khan, Nawal Alsufyani, and Abdulmajeed Alsufyani. Secured insurance framework using blockchain and smart contract. Scientific Programming, 2022(1):6787406:1– 6787406:??, 2022. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6787406.

Han:2023:FLM

[Han23] Zhong Han. A fuzzy logic and multilevel analysisbased evaluation algorithm for digital teaching quality in colleges and universities. Scientific Programming, 2023(1): 7026531:1–7026531:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7026531.

Haveraaen:2000:CSA

[Hav00a] Magne Haveraaen. Case study on algebraic software methodologies for scientific computing. Scientific Programming, 8 (4):261–273, ???? 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

http://iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=2dyrcfdc1p5ytlerkc3u%
26referrer=parent%26backto=
issue%2C5%2C5%3Bjournal%
2C3%2C9%3Blinkingpublicationresults%
2C1%2C1.

Haveraaen:2000:MCA

[Hav00b] Magne Haveraaen. Machine and collection abstractions for user-implemented data-parallel programming. Scientific Programming, 8(4):231–246, ???? 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=2dyrcfdc1p5ytlerkc3u%
26referrer=parent%26backto=
issue%2C3%2C5%3Bjournal%
2C3%2C9%3Blinkingpublicationresults%
2C1%2C1.

Hansrajh:2023:DOF

[HAW23] Arvin Hansrajh, Timothy T. Adeliyi, and Jeanette Wing. Detection of online fake news using blending ensemble learning. Scientific Programming, 2023(1):3434458:1– 3434458:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3434458.

Hockney:1994:PRP

[HB94] R. Hockney and M. Berry.

Parkbench report: public international benchmarks for parallel computers. Scientific Programming, 3 (2):iii, Summer 1994. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hu:2012:MLV

[HB12] Zhen Hu and Raj Bhatnagar. Mining low-variance biclusters to discover coregulation modules in sequencing datasets. Scientific Programming, 20 (1):15–27, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hind:1994:ESP

[HBCM94] Michael Hind, Michael Burke, Paul Carini, and Sam Midkiff. An empirical study of precise interprocedural array analysis. Scientific Programming, 3(3):255–271, Fall 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL http://www.mcs.
                                        newpaltz.edu/~hind.

Hameed:2020:IIB

[HBR^{+}20] Kashif Hameed, Imran Sarwar Bajwa, Shabana Ramzan, Waheed Anwar, and Akmal Khan. An intelligent IoT based healthcare system using fuzzy neural networks. Scientific Programming, 2020(1):8836927:1–8836927:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2020/8836927.

He:2020:RMD

[HBSZ20] Qinlu He, Genqing Bian, Bilin Shao, and Weiqi Zhang. Research on multifeature data routing strategy in deduplication. Scientific Programming, 2020(1):8869237:1– 8869237:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8869237.

He:2023:DLB

[HCC23] Yusheng He, Fei Chen, and Qingyun Chen. Deep learning-based analysis of the effect of cardiac color ultrasound on chronic obstructive pulmonary disease under mask region. Scientific Programming, 2023(1):2504848:1– 2504848:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2504848.

Herrera-Cubides:2017:TCU

[HCGGSS17] Jhon Francined Herrera-Cubides, Paulo Alonso Gaona-García, and Geiner Alexis Salcedo-Salgado. Towards the construction of a user unique authentication mechanism on LMS platforms through model-driven engineering (MDE). Scientific Programming, 2017(1): 9313571:1–9313571:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/9313571.

Hernandez:2008:PTM

[HCJ08] Oscar Hernandez, Barbara Chapman, and Haoqiang Jin. A performance tuning methodology with compiler support. Scientific Programming, 16(2– 3):135–153, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hao:2023:IUS

[HCM23] Liangfeng Hao, Xue Chen, and Chengpeng Min. The impact of urban sprawl and smart city construction on regional coordination. Scientific Programming, 2023(1):5589571:1– 5589571:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5589571.

Huang:2017:RGS

[HCQ17] Dong Huang, Y. L. Chen, and Jianping Qiao. Research on geological survey data management and automatic mapping technology. Scientific Programming, 2017(1): 3618139:1–3618139:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/3618139.

He:2020:ERL

[He20] Xiaoying He. An English reading and learning system based on web. Scientific Programming, 2020(1):7281269:1– 7281269:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7281269.

He:2022:EGE

[He22]  Zhenhui He.  English

grammar error detection using recurrent neural networks. Scientific Programming, 2022(1): 7058723:1–7058723:??, 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7058723.

He:2023:ROG

[He23] Wei He. Research on outdoor garden scene reconstruction based on PMVS algorithm. Scientific Programming, 2023(1): 4491382:1–4491382:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/4491382.

He:2024:ACD

[He24] Li He. Application of color Doppler imaging of portal vein based on iterative reconstruction algorithm in the diagnosis of exercise-induced abdominal pain. Scientific Programming, 2024(1): 9975833:1–9975833:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9975833.

Hemler:2000:KEU

[Hem00] Richard S. Hemler. Key elements of the userfriendly, GFDL SKYHI general circulation model. Scientific Programming, 8 (1):39–47, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

                                        //iospress.metapress.
                                        com/app/home/contribution.

asp%3Fwasp=h82chcapth0xynh5tw5w%
26referrer=parent%26backto=
issue%2C5%2C6%3Bjournal%
2C5%2C9%3Blinkingpublicationresults%
2C1%2C1.

Horasan:2019:ALR

[HEVD19] Fahrettin Horasan, Hasan Erbay, Fatih Varçin, and Emre Deniz. Alternate low-rank matrix approximation in latent semantic analysis. Scientific Programming, 2019(1):1095643:1– 1095643:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/1095643.

Huang:2023:ARV

[HFH^{+}23] Zihe Huang, Ying Fang, He Huang, Xiaomei Xu, Jiwei Wang, and Xiaobo Lai. Automatic retinal vessel segmentation based on an improved U-Net approach. Scientific Programming, 2023(1): 5520407:1–5520407:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5520407.

Hafsa:2023:FII

[HGMM23] Amal Hafsa, Mohamed Gafsi, Jihene Malek, and Mohsen Machhout.

FPGA implementation of improved security approach for medical image encryption and decryption. Scientific Programming, 2023(1):6610655:1– 6610655:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6610655.

Hu:2023:ERT

[HH23a] Jin Hu and Li Han. Exchange rate transaction of international trade goods based on fuzzy granulation and deep learning. Scientific Programming, 2023(1):6954732:1– 6954732:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6954732.

Huang:2023:VBI

[HH23b] Lu Huang and Hongfei Han. The value of biological image technology in improving athletes’ motor skills. Scientific Programming, 2023(1):9489505:1– 9489505:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9489505.

Huang:2023:ABS

[HHCC23] Sibo Huang, Jianfeng Huang, Zhaoquan Cai, and Han Cui. Adaptive backstepping sliding mode control for quadrotor UAV. Scientific Programming, 2023(1): 3997648:1–3997648:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3997648.

Huang:2016:RLT

[HHD^{+}16] Chenlin Huang, Chuanwang Hou, Huadong Dai, Yan Ding, Songling Fu, and Mengluo Ji. Research on Linux trusted boot method based on reverse integrity verification. Scientific Programming, 2016(??): 4516596:1–4516596:12, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
4516596/.

Herrera:2005:PSA

[HHML05] J. Herrera, E. Huedo, R. S. Montero, and I. M. Llorente. Porting of scientific applications to Grid Computing on Grid-Way. Scientific Programming, 13(4):317–331, ???? 2005. CODEN SCIPEV.

ISSN 1058-9244 (print), 1875-919X (electronic).

Hu:2023:FCM

[HHZ^{+}23] Jianyun Hu, Pinglin He, Bixin Zhang, Bin Su, Jing Chen, and Haifeng Hu. Fuzzy c-means algorithmbased adoption of obturator nerve block under adaptive ultrasound imaging for bladder tumor. Scientific Programming, 2023(1):9054812:1– 9054812:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9054812.

Hacibeyoglu:2018:NMP

[HI18] Mehmet Hacibeyoglu and Mohammed H. Ibrahim. A novel multimean particle swarm optimization algorithm for nonlinear continuous optimization: Application to feed-forward neural network training. Scientific Programming, 2018(1): 1435810:1–1435810:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2018/1435810.

Forum:1993:HPF

[Hig93] High Performance Fortran Forum. High Performance Fortran language specification. Scientific Programming, 2(1– 2):1–170, Spring–Summer 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hillman:1997:RSE

[Hil97]  L. Hillman.  Review:

Scientist’s and engineer’s guide to workstations and supercomputers: coping with Unix, RISC, vectors, and programming, by Rubin H. Landau and Paul J. Fink, Jr. Scientific Programming, 6(4):391–393, Winter 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hu:1996:IBA

[HJ96] Y. Charlie Hu and S. Lennart Johnsson. Implementing O(N) N-body algorithms efficiently in data-parallel languages. Scientific Programming, 5 (4):337–364, 1996. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ha:2015:EAF

[HJ15] Ok-Kyoon Ha and Yong-Kee Jun. An efficient algorithm for on-the-fly data race detection using an epoch-based technique. Scientific Programming, 2015(??):205827:1– 205827:14, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
205827/.

He:2020:SAF

[HJO^{+}20] Lifei He, Ming Jiang, Ryutarou Ohbuchi, Takahiko Furuya, Min Zhang, and Pengfei Li. Scale adaptive feature pyramid networks for 2D object detection. Scientific Programming, 2020(1):8839979:1– 8839979:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8839979.

Huang:2010:ELA

[HJYC10] Lei Huang, Haoqiang Jin, Liqi Yi, and Barbara Chapman. Enabling locality-aware computations in OpenMP. Scientific Programming, 18 (3–4):169–181, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Han:2017:ADP

[HKH^{+}17] Seunghee Han, Bosung Kim, Jaemin Han, Kyehee Kim, and JooSeok Song. Adaptive data placement for improving performance of online social network services in a multicloud environment. Scientific Programming, 2017(1):2824782:1– 2824782:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/2824782.

haq:2020:DAM

[hKJ^{+}20] Ayesha Kamran Ul haq, Amira Khattak, Noreen Jamil, M. Asif Naeem, and Farhaan Mirza. Data analytics in mental healthcare. Scientific Programming, 2020(1):2024160:1– 2024160:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/2024160.

Howle:2012:PHP

[HKL^{+}12] Victoria E. Howle, Robert C.

Kirby, Kevin Long, Brian Brennan, and Kimberly Kennedy. Playa: Highperformance programmable linear algebra. Scientific Programming, 20(3):257– 273, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Huiqi:2023:MAA

[HKQ^{+}23] Zhao Huiqi, Abdullah Khan, Xu Qiang, Shah Nazir, Yasir Ali, and Farhad Ali. MCDM approach for assigning task to the workers by selected features based on multiple criteria in crowdsourcing. Scientific Programming, 2023(1):4600764:1– 4600764:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4600764.

Ho-Kieu:2018:CPD

[HKVVNT18] D. Ho-Kieu, T. Vo-Van, and T. Nguyen-Trang. Clustering for probability density functions by new k-medoids method. Scientific Programming, 2018 (1):2764016:1–2764016:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/2764016.

Han:2023:EFP

[HL23a] Xianjing Han and Guoxin Liang. Echocardiographic features of patients with coronary heart disease and angina pectoris under deep learning algorithms. Scientific Programming, 2023(1): 8336959:1–8336959:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8336959.

He:2023:EEF

[HL23b] Zhiyong He and Hanguang Luo. An efficient early frame breaking strategy for RFID tag identification in large-scale industrial Internet of Things. Scientific Programming, 2023(1):4603629:1–4603629:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4603629.

Hu:2023:COH

[HL23c] Jingnan Hu and Hui Li.

Composition and optimization of higher education management system based on data mining technology. Scientific Programming, 2023(1): 5631685:1–5631685:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5631685.

Huang:2023:RDG

[HL23d] Qian Huang and Xue Wen Li. Research on the design of government affairs platform in the context of big data. Scientific Programming, 2023(1):

9936217:1–9936217:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9936217.

He:2024:KBN

[HLC^{+}24] Wenxiu He, Fangfang Lu, Jingjing Chen, Yi Ruan, Tingjuan Lu, and Yi Zhang. A kernel-based node localization in anisotropic wireless sensor network. Scientific Programming, 2024(1):9944358:1–9944358:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9944358.

Huang:2023:FPS

[HLH23] Guangliang Huang, Zhuangxu Lan, and Guo Huang. Football players’ shooting posture norm based on deep learning in sports event video. Scientific Programming, 2023(1): 1552096:1–1552096:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/1552096.

He:2023:RDW

[HLHC23] Jian’qiang He, Naian Liu, Mei’lin Han, and Yao Chen. Research on danjiang water quality prediction based on improved artificial bee colony algorithm and optimized BP neural network. Scientific Programming, 2023(1):3688300:1– 3688300:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3688300.

Ha:2019:EPM

[HLKY19] Ok-Kyoon Ha, Keonpyo Lee, Wan-Jin Kim, and Kun Su Yoon. Effective parallelization method for object recognition in 2D sonar images based on task partitioning. Scientific Programming, 2019(1):8908950:1– 8908950:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/8908950.

Han:2023:DLB

[HLL23] Guocan Han, Weifeng Lin, and Wei Lin. Deep learning-based magnetic resonance imaging in the evaluation of tumor, node, and metastasis staging of renal cancer. Scientific Programming, 2023(1):5989870:1– 5989870:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5989870.

Huedo:2005:DEI

[HLM^{+}05] E. Huedo, A. Lepinette, R. S. Montero, I. M. Llorente, and L. Vázquez. Development and execution of an impact cratering application on a computational Grid. Scientific Programming, 13(1): 19–30, ???? 2005. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hill:2011:EOP

[HLM^{+}11] Zach Hill, Jie Li, Ming Mao, Arkaitz Ruiz-Alvarez, and Marty Humphrey. Early observations on the performance of Windows Azure. Scientific Programming, 19(2–3):121– 132, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hong:2016:AAQ

[HLPY16] Cheol-Ho Hong, Kyungwoon Lee, Hyunchan Park, and Chuck Yoo. ANCS: Achieving QoS through dynamic allocation of network resources in virtualized clouds. Scientific Programming, 2016(??):4708195:1–4708195:10, ???? 2016. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2016/
4708195/.

Hou:2024:ECM

[HLS^{+}24] Li Hou, Qi Liu, Kiran Saeed, Saqib Ali Haidery, M. Irfan Uddin, and Hizbullah Khattak. Enhancement of the capability maturity model for improving the quality of software projects in developing countries. Scientific Programming, 2024(1): 9982227:1–9982227:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9982227.

He:2018:SDM

[HLW^{+}18] Yun He, Tong Li, Wei Wang, Wei Lan, and Xiang Li. A structure-driven method for information retrieval-based software change impact analysis. Scientific Programming, 2018(1):5494209:1– 5494209:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/5494209.

Hao:2017:AOP

[HLY^{+}17] Yun Hao, Gaofeng Li, Pingpeng Yuan, Hai Jin, and Xiaofeng Ding. An association-oriented partitioning approach for streaming graph query. Scientific Programming, 2017(1):2573592:1–2573592:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2017/2573592.

Hu:2020:CBM

[HLY20] Bin Hu, Jianlin Lv, and Kun Yang. Costbenefit models on integrating information technology services in automotive production management. Scientific Programming, 2020(1): 8877780:1–8877780:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/8877780.

Hao:2021:CGS

[HLZ^{+}21] Lujie Hao, Qinghua Lv, Xiaosan Zhang, Qingquan Jiang, Songxian Liu, and Lin Ping. Conquering gender stereotype threat in “Digit sports”: Effects of gender swapping on female players’ continuous participation intention in ESports. Scientific Programming, 2021(1):8818588:1– 8818588:??, 2021. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8818588.

Haines:1997:DPP

[HMC97] Matthew Haines, Piyush Mehrotra, and David Cronk. Data-parallel programming in a multithreaded environment. Scientific Programming, 6(2):187–200, Summer 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hill:2007:ISC

[HMCH07] Chris Hill, Dimitris Menemenlis, Bob Ciotti, and Chris Henze. Investigating solution convergence in a global ocean model using a 2048processor cluster of distributed shared memory machines. Scientific Programming, 15(2):107–115, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Huang:2023:AAV

[HMQS23] Lihong Huang, Jianbing Ma, Xiaochun Qiu, and Tao Suo. Assess the application of the e-value in the unmeasured confounder evaluation of observational pharmaceutical studies. Scientific Programming, 2023(1): 9618161:1–9618161:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9618161.

Haveraaen:2015:HPD

[HMR^{+}15] Magne Haveraaen, Karla Morris, Damian Rouson, Hari Radhakrishnan, and Clayton Carson. Highperformance design patterns for modern Fortran. Scientific Programming, 2015(??):942059:1– 942059:14, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
942059/.

Huck:2008:KSA

[HMSM08] Kevin A. Huck, Allen D. Malony, Sameer Shende, and Alan Morris. Knowledge support and automation for performance analysis with PerfExplorer 2.0. Scientific Programming, 16(2–3):123–134, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hansen:1992:OOH

[HMSW92] D. Hansen, D. Maier, J. Stanley, and J. Walpole. Object-oriented heterogeneous database for materials science. Scientific Programming, 1(2): 115–??, 1992. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hao:2023:RMD

[HNG^{+}23] Qinxia Hao, Shah Nazir, Xiaoxu Gao, Li Ma, and Muhammad Ilyas. A review on multicriteria decision support system and industrial Internet of Things for source code transformation. Scientific Programming, 2023(1):6661272:1– 6661272:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6661272.

Haneef:2019:DDL

[HNMB19] Israr Haneef, Rao Muhammad Adeel Nawab, Ehsan Ullah Munir, and Imran Sarwar Bajwa. Design and development of a large cross-lingual plagiarism corpus for urdu-english language pair. Scientific Programming, 2019(1):2962040:1– 2962040:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2019/2962040.

Hong:2023:ACT

[HNSG23] Mu Hong, Shah Nazir, Zhang Shuo, and Wang Guan. Analyzing the classification techniques for bulk of cursive languages data: an overview. Scientific Programming, 2023 (1):6624397:1–6624397:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6624397.

Hong:2023:RDI

[Hon23] Juan Hong. Research on the development of innovation path of ideological and political education in colleges and universities based on cloud computing and kmeans clustering algorithm model. Scientific Programming, 2023(1): 4263981:1–4263981:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/4263981.

Hong:2024:CMH

[Hon24] Jianhua Hong. Care of maintenance hemodialysis patients through intelligent algorithm-based low-dose digital subtraction angiography. Scientific Programming, 2024 (1):8746962:1–8746962:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8746962.

Hey:2002:PEP

[HP02] Tony Hey and Juri Papay.

Performance engineering, PSEs and the GRID. Scientific Programming, 10 (1):3–17, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=9ejnuvwuvby9737jte27%
26referrer=parent%26backto=
issue%2C2%2C9%3Bjournal%
2C2%2C12%3Blinkingpublicationresults%
2C1%2C1.

Hadjidoukas:2009:EFG

[HPD09] P. E. Hadjidoukas, G. Ch.

Philos, and V. V. Dimakopoulos. Exploiting fine-grain thread parallelism on multicore architectures. Scientific Programming, 17(4):309–323, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hernandez:2019:NSA

[HPGC19] Fernando López Hernández, Elena Verdú Pérez, J. Javier Rainer Granados, and Rubén González Crespo. A nondisturbing service to automatically customize notification sending using implicitfeedback. Scientific Programming, 2019(1): 1293194:1–1293194:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2019/1293194.

Hilbrich:2013:MRE

[HPS^{+}13] Tobias Hilbrich, Joachim Protze, Martin Schulz, Bronis R. de Supinski, and Matthias S. Müller. MPI runtime error detection with MUST: Advances in deadlock detection. Scientific Programming, 21(3–4):109– 121, ???? 2013. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

He:2016:DSS

[HPZ16] Hua He, Shanchen Pang, and Zenghua Zhao. Dynamic scalable stochastic Petri net: A novel model for designing and analysis of resource scheduling in cloud computing. Scientific Programming, 2016 (??):9259248:1–9259248:13,

????  2016.  CO-

DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2016/
9259248/.

Hu:2017:SDC

[HPZ17] Yao Hu, Ting Peng, and Lianming Zhang. Software-defined congestion control algorithm for IP networks. Scientific Programming, 2017(1): 3579540:1–3579540:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/3579540.

He:2023:ASO

[HQL^{+}23] Xueting He, Hao Quan, Wanlong Lin, Weiliang Deng, and Zheyi Tan. AGV scheduling optimization for medical waste sorting system. Scientific Programming, 2023(1):4313749:1– 4313749:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4313749.

Heroux:2012:SITa

[HR12a] Michael A. Heroux and Damian Rouson. Special issue on the Trilinos project, Part 1 of 2. Scientific Programming, 20 (2):81, ???? 2012. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Heroux:2012:SITb

[HR12b] Michael A. Heroux and Damian Rouson. Special issue on the Trilinos project, Part 2 of 2. Scientific Programming, 20 (3):221, ???? 2012. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hwang:2003:IDU

[HS03] Yuan-Shin Hwang and Joel Saltz. Interprocedural definition-use chains of dynamic pointer-linked data structures. Scientific Programming, 11(1):3–37, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hu:2020:CPC

[HS20] Qiang Hu and Jiaji Shen. A cluster and process collaboration-aware method to achieve service substitution in cloud service processes. Scientific Programming, 2020(1): 1298513:1–1298513:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2020/1298513.

Hu:2023:DFE

[HS23a] Bin Hu and Weida Shi. Design of a flexible exhibition management information system for exhibition services. Scientific Programming, 2023(1):4789408:1– 4789408:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4789408.

Huang:2023:RMEb

[HS23b] Chun Huang and Diao Shen. Research on music emotion intelligent recognition and classification algorithm in music performance system. Scientific Programming, 2023(1):7886570:1– 7886570:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7886570.

Huang:2023:ISC

[HS23c] Xin Huang and Huilin Song. Investor sentiment combined with multisource information to predict stock prices: an analysis of China’s Ashare market. Scientific Programming, 2023(1): 9094032:1–9094032:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9094032.

Hao:2023:ROP

[HSH23] Gang Hao, Qing Sun, and Ping Han. Research on optimization path of intelligent pension industry based on intelligent fusion algorithm of multisource information. Scientific Programming, 2023(1):6193743:1– 6193743:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6193743.

Hu:2023:QFX

[HSL^{+}23] Po Hu, Suofang Shi, Li Liu, Shuomiao Yin, and Lingfu Wang. Qu Feng Xuan Bi formula alleviates airway inflammation and guides clinical treatment by improving bronchial immune balance in the asthma model. Scientific Programming, 2023(1): 6648834:1–6648834:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2021/6648834.

He:2023:ELM

[HSS23] Man He, Wan Sun, and Naxian Sha. Extreme learning machine denoising algorithm-based transvaginal three-dimensional ultrasound imaging on the diagnosis efficiency of intrauterine adhesions. Scientific Programming, 2023(1):5460333:1– 5460333:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5460333.

Hamilton:1993:FOO

[HST^{+}93] Lisa Hamilton, Mark Stalzer, R. Steven Turley, John Visher, and Stephen Wandzura. Fastscat: an object-oriented program for fast scattering computation. Scientific Programming, 2(4):171–178, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Huang:2023:CSA

[HSY^{+}23] Yiwei Huang, Xiaoyun Sun, Yinping Yao, Yejun Chen, Yan Chen, and Yuping He. Compressed sensing algorithm-based magnetic resonance imaging guided for prognosis of early neurological deterioration after intravenous thrombolysis in ischemic stroke patients. Scientific Programming, 2023(1): 9568856:1–9568856:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9568856.

Hu:2022:RDI

[HT22] Zhihui Hu and Hai Tang. RETRACTION: Design and implementation of intelligent vehicle control system based on Internet of Things and intelligent transportation. Scientific Programming, 2022(1):9829801:1– 9829801:??, ???? 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2024/9829801.

Hu:2023:DWI

[HTZL23] Yong Hu, Jie Tang, Shenghao Zhao, and Ye Li. Diffusion-weighted imaging-magnetic resonance imaging information under class-structured deep convolutional neural network algorithm in the prognostic chemotherapy of osteosarcoma. Scientific Programming, 2023 (1):4989166:1–4989166:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4989166.

Hu:2023:AEF

[Hu23a]  Jingxiao Hu.  Analy-

sis of enterprise financial and economic impact based on background deep learning model under business administration. Scientific Programming, 2023(1):7178893:1– 7178893:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7178893.

Hu:2023:FPP

[Hu23b]  Xin Hu.  Football

player posture detection method combining foreground detection and neural networks. Scientific Programming, 2023(1): 4102294:1–4102294:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4102294.

Huang:2021:IAB

[Hua21] Xianhua Huang. Intelligent algorithms-based CT image segmentation in patients with cardiovascular diseases and realization of visualization algorithms. Scientific Programming, 2021(1): 2285884:1–2285884:??, 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2285884.

Huang:2023:STN

[Hua23]  Yuxin Huang.  Spa-

tial transformer networkbased automatic modulation recognition of blind signals. Scientific Programming, 2023(1): 9450961:1–9450961:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9450961.

Hui:2022:EBO

[Hui22]  Yanli Hui.  Evalua-

tion of blended oral English teaching based on the mixed model of SPOC and deep learning. Scientific Programming, 2022(1):7044779:1– 7044779:??, 2022. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7044779.

Huimin:2023:RPS

[Hui23] Yao Huimin. Research on parallel support vector machine based on spark big data platform. Scientific Programming, 2023(1):7998417:1– 7998417:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7998417.

Hascoet:2008:CAR

[HUN08] L. Hascoët, J. Utke, and U. Naumann. Cheaper adjoints by reversing address computations. Scientific Programming, 16 (1):81–92, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Heroux:2012:NOT

[HW12] Michael A. Heroux and James M. Willenbring. A new overview of the Trilinos project. Scientific Programming, 20(2):83– 88, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Huang:2023:RDL

[HW23a] Boning Huang and Junkang Wei. Research on deep learning-based financial risk prediction. Scientific Programming, 2023(1): 6913427:1–6913427:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6913427.

Huang:2023:CBM

[HW23b] Tianlin Huang and Ning Wang. A clusteringbased method for business hall efficiency analysis. Scientific Programming, 2023(1):7622576:1– 7622576:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7622576.

Huang:2018:LCN

[HWJ18] Han Huang, Hongyu Wang, and Dawei Jin. A low-cost named entity recognition research based on active learning. Scientific Programming, 2018(1):1890683:1– 1890683:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/1890683.

Han:2022:CDE

[HWSL22] Dongfang Han, Dajun Wang, Ba Sangzeren, and Xiaomei Li. Clinical diagnosis of endometrial polyps using multioperator algorithm combined with hysteroscopy. Scientific Programming, 2022(1): 6254873:1–6254873:??, 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6254873.

He:2023:ERB

[HWY^{+}23] Junliang He, Minghui Wei, Hang Yu, Jun Yuan, and Yanbing Chen. Evaluating the rail-based multimodal freight transportation after HSR entry in Yangtze River Delta Economics Zone. Scientific Programming, 2023 (1):5530343:1–5530343:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5530343.

Hu:2023:SAB

[HX23] Yan Hu and Yong Xu.

Study on the automatic basketball shooting system based on the background subtraction method. Scientific Programming, 2023(1): 2634662:1–2634662:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/2634662.

Huang:2024:GTC

[HXPZ24] Lufei Huang, Ying Xu, Xiaohui Pan, and Tao Zhang. Green technology collaboration network analysis of China’s transportation sector: a patent-based analysis. Scientific Programming, 2024(1):9961071:1– 9961071:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9961071.

Huang:2016:ETF

[HXX16] Ling Huang, Xuhuan Xie, and Wenbo Xie. Eventtriggered faults tolerant control for stochastic systems with time delays. Scientific Programming, 2016(??):5606234:1–5606234:13, ???? 2016. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://www.hindawi.
                                        com/journals/sp/2016/
                                        5606234/.

Hongjiu:2015:EMC

[HY15] Liu Hongjiu and Hu Yanrong. An evaluating method with combined assigning-weight based on maximizing variance. Scientific Programming, 2015(??):290379:1– 290379:8, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
290379/.

Hu:2023:SOD

[HY23] Xiaohong Hu and Ziyang Yao. Selection of outline descriptors based on LightGBM with application to infrared image target recognition. Scientific Programming, 2023(1):4940338:1– 4940338:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4940338.

Hu:2020:ABD

[HYK^{+}20] Li Qin Hu, Amit Yadav, Asif Khan, Hong Liu, and Amin Ul Haq. Application of big data fusion based on cloud storage in green transportation: an application of healthcare. Scientific Programming, 2020(1):1593946:1– 1593946:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

                                        wiley.com/doi/epdf/10.
                                        1155/2020/1593946.

Huang:2019:SPD

[HYS^{+}19] Wanrong Huang, Xiaodong Yi, Yichun Sun, Yingwen Liu, Shuai Ye, and Hengzhu Liu. Scalable parallel distributed coprocessor system for graph searching problems with massive data. Scientific Programming, 2019(1):1496104:1– 1496104:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2017/1496104.

Hempel:1999:AMP

[HZ99] Rolf Hempel and Falk Zimmermann. Automatic migration from PARMACS to MPI in parallel Fortran applications. Scientific Programming, 7(1):39–46, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

                                        //iospress.metapress.
                                        com/app/home/contribution.
                                        asp%3Fwasp=64cr5a4mg33tuhcbdr02%
                                        26referrer=parent%26backto=
                                        issue%2C3%2C7%3Bjournal%
                                        2C8%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Huang:2023:RMEa

[HZ23] Cuiqing Huang and Qiang Zhang. Research on music emotion recognition model of deep learning based on musical stage effect. Scientific Programming, 2023(1):3807666:1– 3807666:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3807666.

Huang:2024:RLS

[HZ24] Dong Huang and WeiXin Zhang. Research on learning state based on students’ attitude and emotion in class learning. Scientific Programming, 2024(1):9944176:1– 9944176:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9944176.

Hu:2017:CMI

[HZD17] Qiang Hu, Zhen Zhao, and JunWei Du. A clustering method for isomorphic evolution of web services. Scientific Programming, 2017(1):5725864:1– 5725864:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/5725864.

Hu:2023:IIT

[HZFH23] Hangyu Hu, Xuemeng Zhai, Gaolei Fei, and Guangmin Hu. IT-DPM: an internet topology dynamic propagation model based on generative adversarial learning. Scientific Programming, 2023(1):2390466:1– 2390466:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/2390466.

Huang:2023:CER

[HZHW23] Kai Huang, Tongqing Zhang, Zhiliang He, and Fengxia Wang. Comparison of effects of radiofrequency ablation of liver cancer guided by CT images based on deep learning algorithm. Scientific Programming, 2023(1): 7773473:1–7773473:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/7773473.

Han:2023:FDA

[HZJ23] Haoyang Han, Jundong Zhang, and Ruizheng Jiang. Feature design assessment of the ship fire alarm system. Scientific Programming, 2023(1):3934428:1– 3934428:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3934428.

He:2023:PKG

[HZL^{+}23] Yanming He, Shujun Zhang, Xueguang Liu, Dong Mao, and Zhenzhong Sun. Percutaneous kyphoplasty guided by CT images based on SEPB algorithm in the treatment of elderly osteoporotic thoracolumbar vertebral compression fractures. Scientific Programming, 2023(1): 6044121:1–6044121:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6044121.

Ibrahim:2009:ESD

[IB09] Khaled Z. Ibrahim and François Bodin. Efficient SIMDization and data management of the Lattice QCD computation on the Cell Broadband Engine. Scientific Programming, 17(1– 2):153–172, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ierotheou:2001:SAP

[IJL^{+}01] C. S. Ierotheou, S. P. Johnson, P. F. Leggett, M. Cross, E. W. Evans, H. Jin, M. Frumkin, and J. Yan. The semi-automatic parallelisation of scientific application codes using a computer aided parallelisation toolkit. Scientific Programming, 9 (2–3):163–173, Spring– Summer 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

                                        //iospress.metapress.
                                        com/app/home/contribution.
                                        asp%3Fwasp=7pab6qgbaf8vxg991rwy%
                                        26referrer=parent%26backto=
                                        issue%2C9%2C11%3Bjournal%
                                        2C1%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Islam:2013:MSC

[IMB^{+}13] Tanzima Zerin Islam, Kathryn Mohror, Saurabh Bagchi, Adam Moody, Bronis R. de Supinski, and Rudolf Eigenmann. McrEngine: a scalable checkpointing system using data-aware aggregation and compression. Scientific Programming, 21(3–4):149–163, ???? 2013. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Insa:2018:BPA

[IPST18] David Insa, Sergio Pérez, Josep Silva, and Salvador Tamarit. Behaviour preservation across code versions in Erlang. Scientific Programming, 2018 (1):9251762:1–9251762:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/9251762.

Iyengar:2002:AWS

[IR02] Arun Iyengar and Daniela Rosu. Architecting Web sites for high performance. Scientific Programming, 10(1):75– 89, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=9ejnuvwuvby9737jte27%
26referrer=parent%26backto=
issue%2C8%2C9%3Bjournal%
2C2%2C12%3Blinkingpublicationresults%
2C1%2C1.

Ioannidis:1999:CCR

[IRSD99] Sotiris Ioannidis, Umit Rencuzogullari, Robert Stets, and Sandhya Dwarkadas. CRAUL: Compiler and run-time integration for adaptation under load. Scientific Programming, 7 (3–4):261–273, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

http://iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=53f7mftrrm4r73yyrqau%
26referrer=parent%26backto=
issue%2C6%2C12%3Bjournal%
2C6%2C9%3Blinkingpublicationresults%
2C1%2C1.

Iosup:2008:IOG

[ITF^{+}08] Alexandru Iosup, Todd Tannenbaum, Matthew Farrellee, Dick Epema, and Miron Livny. Interoperating grids through Delegated MatchMaking. Scientific Programming, 16(2–3):233–253, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Jamous:2023:NAP

[JAED23] Razan Jamous, Hosam Alrahhal, and Mohamed El-Darieby. A new ANN-Particle swarm optimization with center of gravity (ANN-PSOCoG) prediction model for the stock market under the effect of COVID-

19. Scientific Programming, 2023(1):6656150:1– 6656150:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6656150.

Janssens:2015:FEA

[JBLB15] Bart Janssens, Támas Bányai, Karim Limam, and Walter Bosschaerts. Finite element assembly using an embedded domain specific language. Scientific Programming, 2015(??):797325:1–797325:22, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
797325/.

Juve:2010:ERP

[JDVM10] Gideon Juve, Ewa Deelman, Karan Vahi, and Gaurang Mehta. Experiences with resource provisioning for scientific workflows using Corral. Scientific Programming, 18 (2):77–92, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Jacob:2007:PMC

[JEM07] Joseph C. Jacob, Peter Eisenhardt, and David Makovoz. Parallel MOPEX: Computing mosaics of large-area Spitzer surveys on a cluster computer. Scientific Programming, 15(2):75–81, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Jespersen:2010:ACC

[Jes10] Dennis C. Jespersen. Acceleration of a CFD code with a GPU. Scientific Programming, 18(3– 4):193–201, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Jiang:2020:GKF

[JFD20] Wei Jiang, Xi Fang, and Jianmei Ding. Gaussian kernel fuzzy Cmeans algorithm for service resource allocation. Scientific Programming, 2020(1):8889480:1– 8889480:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/8889480.

Jiang:2016:FAF

[JFPL16] Jianhua Jiang, Yunzhao Feng, Milan Parmar, and Keqin Li. FP-ABC: Fast and parallel ABC based energy-efficiency live VM allocation policy in data centers. Scientific Programming, 2016(??): 9524379:1–9524379:9, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://www.hindawi.
                                        com/journals/sp/2016/
                                        9524379/.

Jin:2023:CNN

[JFY23] Xin Jin, Lei Fan, and Yongling Yao. Convolutional neural network optimization algorithmbased magnetic resonance imaging in analysis of chronic pain caused by the myofascial trigger point. Scientific Programming, 2023(1):4505147:1– 4505147:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4505147.

Juneja:2023:FCT

[JGDK23] Sapna Juneja, Mamta Gahlan, Gaurav Dhiman, and Sandeep Kautish. Futuristic cyber-twin architecture for 6G technology to support internet of everything. Scientific Programming, 2023(1): 9101782:1–9101782:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9101782.

Ji:2023:NDL

[JGF23] Zhiwei Ji, Jiaheng Gong, and Jiarui Feng. A novel deep learning approach for anomaly detection of time series data. Scientific Programming, 2023(1):6636270:1– 6636270:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6636270.

Jin:2023:ODR

[JHL^{+}23] Qiong Jin, Mian Huang, Jun Lin, Shansan Wu, Zhang Shen, and Haibing Wang. Optimized double-regional filtering algorithm on MRI threedimensional reconstructed images for the evaluation of effects of delivery on the pelvis of primiparas. Scientific Programming, 2023(1):7985624:1– 7985624:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7985624.

Jolivet:2014:SDD

[JHNP14] Pierre Jolivet, Frédéric Hecht, Frédéric Nataf, and Christophe Prud’homme. Scalable domain decomposition preconditioners for heterogeneous elliptic problems. Scientific Programming, 22(2):157– 171, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Jiang:2020:FIC

[JHS^{+}20] Xianwei Jiang, Bo Hu, Suresh Chandra Satapathy, Shui-Hua Wang, and Yu-Dong Zhang. Fingerspelling identification for Chinese sign language via AlexNet-based transfer learning and Adam optimizer. Scientific Programming, 2020(1):3291426:1– 3291426:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/3291426.

Jung:2014:EFC

[JIC^{+}14] Myoungsoo Jung, Ellis H. Wilson III, Wonil Choi, John Shalf, Hasan Metin Aktulga, Chao Yang, Erik Saule, Umit V. Catalyurek, and Mahmut Kandemir. Exploring the future of out-of-core computing with computelocal non-volatile memory. Scientific Programming, 22(2):125–139, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Jin:2003:AMP

[JJY^{+}03] Haoqiang Jin, Gabriele Jost, Jerry Yan, et al. Automatic multilevel parallelization using OpenMP. Scientific Programming, 11(2):177–190, 2003. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Jost:2010:SIE

[JK10] Gabriele Jost and Alice Koniges. Special issue: Exploring languages for expressing medium to massive on-chip parallelism. Scientific Programming, 18(3–4):125– 126, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Jaf:2022:MBT

[JK22] Ashti Afasyaw Jaf and Sema Koç Kayhan. Machinebased transliterate of Ottoman to Latin-based script. Scientific Programming, 2022(1):7152935:1– 7152935:??, 2022. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/7152935.

Jahan:2023:BLM

[JKA^{+}23] Muhammad Shah Jahan, Habib Ullah Khan, Shahzad Akbar, Muhammad Umar Farooq, Sarah Gul, and Anam Amjad. Bidirectional language modeling: a systematic literature review. Scientific Programming, 2023(1):6641832:1– 6641832:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6641832.

Jacobson:1992:ADD

[JKR92] P. Jacobson, B. Kågström, and M. Ränner. Algorithm development for distributed memory multicomputers using CON-LAB. Scientific Programming, 1(??):185–203, 1992. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Jiang:2023:ETR

[JL23a] Yizhang Jiang and Bo Li.

Exploration on the teaching reform measure for machine learning course system of artificial intelligence specialty. Scientific Programming, 2023(1): 8971588:1–8971588:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8971588.

Jiao:2023:TOA

[JL23b] Yipei Jiao and Yu Liu.

The teaching optimization algorithm mode of integrating mobile cloud teaching into ideological and political courses under the internet thinking mode. Scientific Programming, 2023(1):6492009:1– 6492009:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6492009.

Ji:2024:MLA

[JLJ24] Yanni Ji, Wenqian Lou, and Jianwei Ji. Machine learning algorithm-based analysis of efficacy of pulmonary surfactant combined with mucosolvan in meconium aspiration syndrome of newborns through ultrasonic images. Scientific Programming, 2024(1):8469487:1– 8469487:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8469487.

Jing:2020:DSI

[JLX^{+}20] Hongwei Jing, Xiaoming Li, Guangquan Xu, Mengli Zhu, Li Shen, FangYuan Liu, and Haoyang Peng. Dynamic study of intelligent traffic behaviour based on multiple traffic modes. Scientific Programming, 2020(1): 7254907:1–7254907:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7254907.

Jiang:2021:AMD

[JMGH21] Xu Jiang, Baolei Mao, Jun Guan, and Xingli Huang. Android malware detection using finegrained features. Scientific Programming, 2021 (1):5190138:1–5190138:??, 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/5190138.

Jackson:2011:PCA

[JMR^{+}11] Keith R. Jackson, Krishna Muriki, Lavanya Ramakrishnan, Karl J. Runge, and Rollin C. Thomas. Performance and cost analysis of the Supernova factory on the Amazon AWS cloud. Scientific Programming, 19 (2–3):107–119, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Javed:2024:HFF

[JNK24] Muhammad Fasih Javed, Waqas Nawaz, and Kifayat Ullah Khan. HOVA-FPPM: Flexible periodic pattern mining in time series databases using hashed occurrence vectors and apriori approach. Scientific Programming, 2024(1):8841188:1–8841188:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8841188.

Jin:2023:AMS

[JP23] Hongjuan Jin and Weihai Peng. Application of multislice spiral CT volume reconstruction in congenital microtia reconstruction. Scientific Programming, 2023(1):8312296:1– 8312296:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8312296.

Jost:2010:EUH

[JR10] Gabriele Jost and Bob Robins. Experiences using hybrid MPI/OpenMP in the real world: Parallelization of a 3D CFD solver for multi-core node clusters. Scientific Programming, 18(3–4):127– 138, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Junejo:2021:PPA

[JRS^{+}21] Muhammad Haleem Junejo, Ab Al-Hadi Ab Rahman, Riaz Ahmed Shaikh, Kamaludin Mohamad Yusof, Imran Memon, Hadiqua Fazal, and Dileep Kumar. A privacy-preserving attack-resistant trust model for internet of vehicles ad hoc networks. Scientific Programming, 2021(1): 8831611:1–8831611:??, 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8831611.

Jin:2021:DNN

[JSY^{+}21] Yanxia Jin, Zhiru Shi, Jing Yang, Yabian Liu, Xingyu Qiao, and Ling Zhang. Deep neural network-based cloth collision detection algorithm. Scientific Programming, 2021(1): 7889278:1–7889278:??, ???? 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2024/7889278.

Jan:2023:SAA

[JUA^{+}23a] Rooh Ullah Jan, Muhammad Usman, Muhammad Faisal Abrar, Najeeb Ullah, Muhammad Asshad, and Sikandar Ali. Scaling agile adoption motivators from management perspective: an analytical hierarchy process approach. Scientific Programming, 2023(1): 4522273:1–4522273:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/4522273.

Juan:2023:CFA

[Jua23b]  Cao Juan.  A cul-

tural and functional approach to the assessment of logical thinking ability in English writing. Scientific Programming, 2023(1):1783384:1– 1783384:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/1783384.

Juan:2023:RAP

[Jua23c] Hu Juan. Recognition algorithm of popular elements of ethnic minority traditional clothing based on PCA. Scientific Programming, 2023(1): 4834944:1–4834944:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/4834944.

Jiang:2023:FGH

[JW23a] Chunmao Jiang and Peng Wu. A finegrained horizontal scaling method for containerbased cloud. Scientific Programming, 2023(1): 6397786:1–6397786:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6397786.

Jiao:2023:FII

[JW23b] Lei Jiao and Yiping Wang. Factors influencing the information service quality of the online website of hospitals in China. Scientific Programming, 2023(1): 3283481:1–3283481:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3283481.

Jiang:2024:MVB

[JWPW24] Hui Jiang, Ping Wang, Lei Peng, and Xiaofeng Wang. Machine vision and big data-driven sports athletes action training intervention model. Scientific Programming, 2024(1):9956710:1–9956710:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9956710.

Jia:2023:DGR

[JYLL23] Junjie Jia, Yewang Yao, Zhipeng Lei, and Pengtao Liu. Dynamic group recommendation algorithm based on member activity level. Scientific Programming, 2023(1):1969118:1– 1969118:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/1969118.

Jiang:2022:RMH

[JZ22] Sheng Jiang and ChuanZhen Zheng. RETRACTION: Multisensor human resource data fusion and its application in industrial distribution. Scientific Programming, 2022(1): 9875938:1–9875938:??, ???? 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2024/9875938.

Jia:2023:PEP

[JZH^{+}23] Yunfei Jia, Zhaohui Zhang, Zejun He, Panpan Zhu, Yibei Zhang, and Tianhua Sun. Production efficiency prediction of pig breeding industry by optimized LSTM computer algorithm under environmental regulation. Scientific Programming, 2023(1):3074167:1– 3074167:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3074167.

Jiang:2020:IPR

[JZL^{+}20] Qingquan Jiang, Xiaosan Zhang, Qiaozhen Lin, Guofu Chen, Rui Zhang, and Songxian Liu. Internet penetration and regional financial development in China: Empirical evidence based on Chinese provincial panel data. Scientific Programming, 2020(1):8856944:1– 8856944:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8856944.

Jiang:2021:RSA

[JZS21] Zhaoyin Jiang, Fuyou Zhang, and Laishuang Sun. Retracted: Sports action recognition based on image processing technology and analysis of the development of sports industry pattern. Scientific Programming, 2021(1):9764270:1– 9764270:??, ???? 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2024/9764270.

Jin:2023:PFT

[JZY23] Jiulin Jin, Ping Zhao, and Taijie You. Picture fuzzy TOPSIS method based on CPFRS model: an application to risk management problems. Scientific Programming, 2023(1): 6628745:1–6628745:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6628745.

Ji:2020:QEO

[JZZW20] Xuechun Ji, Maoxian Zhao, Mingyu Zhai, and Qingxi Wu. Query execution optimization in spark SQL. Scientific Programming, 2020(1):6364752:1– 6364752:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/6364752.

Khan:2020:QAM

[KAA^{+}20] Rafiullah Khan, Arshad Ahmad, Alhuseen Omar Alsayed, Muhammad Binsawad, Muhammad Arshad Islam, and Mohib Ullah. QuPiD attack: Machine learning-based privacy quantification mechanism for PIR protocols in health-related web search. Scientific Programming, 2020(1): 8868686:1–8868686:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8868686.

Khan:2023:CCD

[KAA^{+}23] Irfan Ullah Khan, Nida Aslam, Rawan Alshehri, Seham Alzahrani, Manal Alghamdi, Atheer Almalki, and Maryam Balabeed. Cervical cancer diagnosis model using extreme gradient boosting and bioinspired firefly optimization. Scientific Programming, 2023(1):5540024:1– 5540024:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5540024.

Kosar:2011:PRE

[KABW11] Tevfik Kosar, Ismail Akturk, Mehmet Balman, and Xinqi Wang. PetaShare: a reliable, efficient and transparent distributed storage management system. Scientific Programming, 19(1):27– 43, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kazienko:2015:CAS

[KAS15] Przemysław Kazienko, Reda Alhajj, and Jaideep Srivastava. Computational aspects of social network analysis. Scientific Programming, 2015 (??):961610:1–961610:2, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://www.hindawi.
                                        com/journals/sp/2015/
                                        961610/.

Kohn:1996:ICG

[KB96] Scott R. Kohn and Scott B. Baden. Irregular coarse-grain data parallelism under LPARX. Scientific Programming, 5(3):185–201, Fall 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kaiser:2001:OCC

[KB01] Timothy H. Kaiser and Scott B. Baden. Overlapping communication and computation with OpenMP and MPI. Scientific Programming, 9(2– 3):73–81, Spring–Summer 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL http://iospress.
                                        metapress.com/app/home/

contribution.asp%3Fwasp=
7pab6qgbaf8vxg991rwy%
26referrer=parent%26backto=
issue%2C2%2C11%3Bjournal%
2C1%2C9%3Blinkingpublicationresults%
2C1%2C1.

Krishnan:2002:XSP

[KBG^{+}02] Sriram Krishnan, Randall Bramley, Dennis Gannon, et al. The XCAT Science Portal. Scientific Programming, 10(4):303–317, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Korn:1995:PTD

[KBRS95] C. Falco Korn, J. M. Bull, G. D. Riley, and P. K. Stansby. Parallelization of a three-dimensional shallow-water estuary model on the KSR-1. Scientific Programming, 4 (3):155–169, Fall 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kenny:2005:CIM

[KCO^{+}05] Stuart Kenny, Brian Coghlan, David O’Callaghan, John Ryan, Rob Byrom, Laurence Field, Steve Hicks, Manish Soni, Antony Wilson, Xiaomei Zhu, Roney Cordenonsi, Ari Datta, Linda Cornwall, Abdeslem Djaoui, and Norbert Podhorszki. The CanonicalProducer: an instrument monitoring component of the Relational Grid Monitoring Architecture (R-GMA). Scientific Programming, 13(2):151–158, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kuzmin:2015:PSS

[KCS15] Konstantin Kuzmin, Mingming Chen, and Bolesław K. Szymanski. Parallelizing SLPA for scalable overlapping community detection. Scientific Programming, 2015(??):461362:1– 461362:18, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://www.hindawi.
                                        com/journals/sp/2015/
                                        461362/.

Kurzak:2009:QFC

[KD09] Jakub Kurzak and Jack Dongarra. QR factorization for the Cell Broadband Engine. Scientific Programming, 17(1– 2):31–42, ???? 2009. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kim:2011:AMA

[KeKR^{+}11] Hyunjoo Kim, Yaakoub el Khamra, Ivan Rodero, Shantenu Jha, and Manish Parashar. Autonomic management of application workflows on hybrid computing infrastructure. Scientific Programming, 19(2–3):75–89, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kessler:1996:PDA

[Kes96] Christoph W. Kessler.

Pattern-driven automatic parallelization. Scientific Programming, 5(3):251– 274, Fall 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Khozium:2020:CBI

[KF20] Mohamed O. Khozium and Norah S. Farooqi. Cooperative business intelligence model using a multiagent platform. Scientific Programming, 2020 (1):8898719:1–8898719:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8898719.

Keahey:2005:VWA

[KFFZ05] K. Keahey, I. Foster, T. Freeman, and X. Zhang. Virtual workspaces: Achieving quality of service and quality of life in the Grid. Scientific Programming, 13(4):265–275, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Khan:2020:OPL

[KFU^{+}20] Zahid Alam Khan, Zhengyong Feng, M. Irfan Uddin, Noor Mast, Syed Atif Ali Shah, Muhammad Imtiaz, Mahmoud Ahmad Al-Khasawneh, and Marwan Mahmoud. Optimal policy learning for disease prevention using reinforcement learning. Scientific Programming, 2020(1):7627290:1– 7627290:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/7627290.

King:2008:SSL

[KG08] Brian R. King and Chittibabu Guda. Semisupervised learning for classification of protein sequence data. Scientific Programming, 16(1):5–29, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kistler:2009:PLB

[KGBB09] Michael Kistler, John Gunnels, Daniel Brokenshire, and Brad Benton. Programming the Linpack benchmark for the IBM Power XCell 8i processor. Scientific Programming, 17(1–2):43–57, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kanan:2019:LCS

[KGIL19] Awos Kanan, Fayez Gebali, Atef Ibrahim, and Kin Fun Li. Lowcomplexity scalable architectures for parallel computation of similarity measures. Scientific Programming, 2019(1): 3185137:1–3185137:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/3185137.

Kiran:2015:GSH

[KGMN15] D. C. Kiran, S. Gurunarayanan, Janardan Prasad Misra, and Abhijeet Nawal. Global scheduling heuristics for multicore architecture. Scientific Programming, 2015 (??):860891:1–860891:12, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
860891/.

Khan:2020:CSO

[KGU^{+}20] Atif Khan, Muhammad Adnan Gul, M. Irfan Uddin, Syed Atif Ali Shah, Shafiq Ahmad, Muhammad Dzulqarnain Al Firdausi, and Mazen Zaindin. Corrigendum to “Summarizing online movie reviews: a machine learning approach to big data analytics”. Scientific Programming, 2020(1):7871490:1– 7871490:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7871490.

Khan:2021:SOM

[KGU^{+}21] Atif Khan, Muhammad Adnan Gul, M. Irfan Uddin, Syed Atif Ali Shah, Shafiq Ahmad, Muhammad Dzulqarnain Al Firdausi, and Mazen Zaindin. Summarizing online movie reviews: a machine learning approach to big data analytics. Scientific Programming, 2021(1):5812715:1– 5812715:??, 2021. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/5812715.

Koppler:1997:VDD

[KGV97] Rainer Koppler, Siegfried Grabner, and Jens Volkert. Visualization of distributed data structures for High Performance Fortran-like languages. Scientific Programming, 6(1):115–126, Spring 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kumar:1995:TPF

[KHSJ95] B. Kumar, C.-H. Huang, P. Sadayappan, and R. W. Johnson. Tensor product formulation of Strassen’s matrix multiplication algorithm with memory reduction. Scientific Programming, 4 (4):275–??, Winter 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Khan:2021:MLA

[KIU^{+}21] Atif Khan, Ibrahim Ibrahim, M. Irfan Uddin, Muhammad Zubair, Shafiq Ahmad, Muhammad Dzulqarnain Al Firdausi, and Mazen Zaindin. Machine learning approach for answer detection in discussion forums: an application of big data analytics. Scientific Programming, 2021(1):4621196:1– 4621196:??, 2021. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

                                        wiley.com/doi/epdf/10.
                                        1155/2020/4621196.

Khan:2023:EBA

[KJA^{+}23] Tehseen Khan, Huma Javed, Mohammad Amin, Omar Usman, Syed Ishtiaq Hussain, Amjad Mehmoood, and Carsten Maple. EEG based aptitude detection system for stress regulation in health care workers. Scientific Programming, 2023(1):4620487:1– 4620487:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/4620487.

Kenny:2008:CAC

[KJG^{+}08] Joseph P. Kenny, Curtis L. Janssen, Mark S. Gordon, Masha Sosonkina, and Theresa L. Windus. A component approach to collaborative scientific software development: Tools and techniques utilized by the Quantum Chemistry Science Application Partnership. Scientific Programming, 16(4):287–296, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Knepley:2009:MAP

[KK09] Matthew G. Knepley and Dmitry A. Karpeev. Mesh algorithms for PDE with Sieve I: Mesh distribution. Scientific Programming, 17(3):215–230, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kunzman:2011:PHC

[KK11] David M. Kunzman and Laxmikant V. Kalé. Programming heterogeneous clusters with accelerators using object-based programming. Scientific Programming, 19(1):47–62, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Khan:2022:ITB

[KKA^{+}22] Zahid Khan, Muhammad Zahid Khan, Sikandar Ali, Irshad Ahmed Abbasi, Haseeb Ur Rahman, Umar Zeb, Hizbullah Khattak, and Jiwei Huang. Internet of Things-based smart farming monitoring system for bolting reduction in onion farms. Scientific Programming, 2022(1): 7101983:1–7101983:??, 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7101983.

Khan:2020:TCS

[KKD^{+}20] Usman Ali Khan, Iftikhar Ahmed Khan, Ahmad Din, Waqas Jadoon, Rab Nawaz Jadoon, Muhammad Amir Khan, Fiaz Gul Khan, and Abdul Nasir Khan. Towards a complete set of gym exercises detection using smartphone sensors. Scientific Programming, 2020(1):6471438:1– 6471438:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/6471438.

Kim:2018:NAT

[KKK^{+}18] Taeuk Kim, Awais Khan, Youngjae Kim, Preethika Kasu, and Scott Atchley. NUMA-Aware thread scheduling for big data transfers over terabits network infrastructure. Scientific Programming, 2018(1):4120561:1– 4120561:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2018/4120561.

Kang:2023:DID

[KL23] Yang Kang and Qiang Li. Design and implementation of data sharing traceability system based on blockchain smart contract. Scientific Programming, 2023(1):1455814:1– 1455814:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1455814.

Khan:2020:DCB

[KLK^{+}20a] Sajid Khan, Dong-Ho Lee, Asif Khan, Ahmad Waqas, Abdul Rehman Gilal, and Zahid Hussain Khand. A digital camera-based rotationinvariant fingerprint verification method. Scientific Programming, 2020 (1):9758049:1–9758049:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/9758049.

Khan:2020:IIG

[KLK^{+}20b] Sajid Khan, Dong-Ho Lee, Muhammad Asif Khan, Muhammad Faisal Siddiqui, Raja Fawad Zafar, Kashif Hussain Memon, and Ghulam Mujtaba. Image interpolation via gradient correlation-based edge direction estimation. Scientific Programming, 2020 (1):5763837:1–5763837:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/5763837.

Ke:2018:IOW

[KLL^{+}18] Hai-Feng Ke, Cheng-Bo Lu, Xiao-Bo Li, Gao-Yan Zhang, Ying Mei, and Xue-Wen Shen. An incremental optimal weight learning machine of single-layer neural networks. Scientific Programming, 2018(1): 3732120:1–3732120:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2018/3732120.

Kurowski:2004:DGS

[KLN^{+}04] K. Kurowski, B. Ludwiczak, J. Nabrzyski, A. Oleksiak, and J. Pukacki. Dynamic Grid scheduling with job migration and rescheduling in the Grid-Lab resource management system. Scientific Programming, 12(4):263–273, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kropf:1996:SPC

[KLS^{+}96] Peter G. Kropf, Edgar F. A. Lederer, Thomas Steffen, Karl Guggisberg, Jean-Guy Schneider, and Peter Schwab. SPINET: a parallel computing approach to spine simulations. Scientific Programming, 5(1):15–??, Spring 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Khan:2020:POD

[KLuH^{+}20] Asif Khan, Jian Ping Li, Amin ul Haq, Shah Nazir, Naeem Ahmad, Naushad Varish, Asad Malik, and Sarosh H. Patel. Partial observer decision process model for crane-robot action. Scientific Programming, 2020(1):6349342:1– 6349342:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/6349342.

Kindratenko:2009:ITP

[KMB09] Volodymyr V. Kindratenko, Adam D. Myers, and Robert J. Brunner. Implementation of the twopoint angular correlation function on a highperformance reconfigurable computer. Scientific Programming, 17 (3):247–259, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kamachi:1997:KPH

[KMR^{+}97] T. Kamachi, A. Muller, R. Ruhl, Y. Seo, K. Suehiro, and M. Tamura. Kemari: a portable High Performance Fortran system for distributed memory parallel processors. Scientific Programming, 6 (1):41–58, Spring 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Koko:2007:VMC

[Kok07]  Jonas Koko.  Vector-

ized Matlab codes for linear two-dimensional elasticity. Scientific Programming, 15(3):157–172, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Knies:1994:HPF

[KOM94] Allan Knies, Matthew O’Keefe, and Tom Mac-Donald. High Performance Fortran: a practical analysis. Scientific Programming, 3(3):187– 199, Fall 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kim:2018:CCR

[KPHY18] Youngpil Kim, Hyunchan Park, Cheol-Ho Hong, and Chuck Yoo. CRIM: Conditional remapping to improve the reliability of solid-state drives with minimizing lifetime loss. Scientific Programming, 2018(1):8171096:1– 8171096:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/8171096.

Kholod:2016:CDD

[KPS16] Ivan Kholod, Ilya Petukhov, and Andrey Shorov. Cloud for distributed data analysis based on the actor model. Scientific Programming, 2016(??): 1050293:1–1050293:11, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
1050293/.

Kim:2018:HSP

[KPW^{+}18] Youngjae Kim, Sungyong Park, Feiyi Wang, Guangyu Sun, and Shangguang Wang. HPC software and programming environments for big data applications. Scientific Programming, 2018(1): 6015685:1–6015685:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/6015685.

Kalaivani:2015:FRB

[KS15] P. Kalaivani and K. L. Shunmuganathan. Feature reduction based on genetic algorithm and hybrid model for opinion mining. Scientific Programming, 2015 (??):961454:1–961454:15, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://www.hindawi.
                                        com/journals/sp/2015/
                                        961454/.

Kumari:2019:DAS

[KS19] Monika Kumari and G. Sahoo. Design and analysis of sustainable and seasonal profit scaling model in cloud environment. Scientific Programming, 2019(1):7457938:1– 7457938:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2019/7457938.

Kang:2023:DKB

[KSYC23] Yongkun Kang, Chuang Sun, Chao Yang, and Honghai Chen. Diffusion kurtosis-based brain image segmentation for the structural remodeling of white matter in patients with hypoxicischemic encephalopathy. Scientific Programming, 2023(1):7692460:1– 7692460:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7692460.

Katsaloulis:2005:SAL

[KTP05] P. Katsaloulis, T. Theoharis, and A. Provata. Statistical algorithms for long DNA sequences: Oligonucleotide distributions and homogeneity maps. Scientific Programming, 13(3):177–188, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kufel:2015:NLS

[Kuf15]  Lukasz Kufel.  Net-

work latency in systems event monitoring for multiple locations. Scientific Programming, 2015 (??):371620:1–371620:6, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
371620/.

Khan:2023:IBS

[KV23] Yunus Khan and Sunita Verma. An intelligent blockchain and softwaredefined networking-based evidence collection architecture for cloud environment. Scientific Programming, 2023(1):7294206:1– 7294206:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7294206.

Korkhov:2007:VGI

[KVW^{+}07] Vladimir Korkhov, Dmitry Vasyunin, Adianto Wibisono, Adam S. Z. Belloum, Márcia A. Inda, Marco Roos, Timo M. Breit, and L. O. Hertzberger. VLAM-G: Interactive data driven workflow engine for Grid-enabled resources. Scientific Programming, 15(3):173–188, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kang:2023:DDA

[KYR^{+}23] Yuxiang Kang, Jie Yu, Zhipeng Ren, Guokai Zhang, Wen Cao, Yinguang Zhang, and Qiang Dong. Dense dilated attentive network for automatic classification of femur trochanteric fracture. Scientific Programming, 2023(1):1929800:1– 1929800:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1929800.

Khan:2023:CBT

[KZK^{+}23] Nauman Ali Khan, Wuyang Zhou, Mudassar Ali Khan, Ahmad Almogren, and Ikram Ud Din. Correlation between triadic closure and homophily formed over location-based social networks. Scientific Programming, 2023(1): 5553566:1–5553566:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5553566.

Kohr:1994:POO

[KZRR94] David R. Kohr, Jr., Xingbin Zhang, Mustafizur Rahman, and Daniel A. Reed. Performance of an object-oriented, parallel operating system. Scientific Programming, 3 (4):301–324, Winter 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Khan:2020:PPS

[KZS^{+}20] Asif Khan, Huaping Zhang, Jianyun Shang, Nada Boudjellal, Arshad Ahmad, Asmat Ali, and Lin Dai. Predicting politician’s supporters’ network on Twitter using social network analysis and semantic analysis. Scientific Programming, 2020(1):9353120:1– 9353120:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/9353120.

Khan:2020:ITS

[KZZ^{+}20] Nauman Ali Khan, Sihai Zhang, Wuyang Zhou, Ahmad Almogren, Ikram Ud Din, and Muhammad Asif. Inferring ties in social IoT using locationbased networks and identification of hidden suspicious ties. Scientific Programming, 2020(1): 6667610:1–6667610:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/6667610.

Kong:2021:PRA

[KZZ^{+}21] Jianguo Kong, Zhiwei Zhao, Xiangwei Zhang, Hanwen Chang, and Haijun Liang. PPA-ResNet: an airport bird recognition model based on improved ResNet. Scientific Programming, 2021(1): 4200980:1–4200980:??, ???? 2021. CODEN SCIPEV.

ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2024/4200980.

Lancaster:2001:PSE

[Lan01] David Lancaster. A problem solving environment based on CORBA. Scientific Programming, 9(4): 233–242, 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=64t4wprhwk589ellmv56%
26referrer=parent%26backto=
issue%2C4%2C4%3Bjournal%
2C3%2C12%3Blinkingpublicationresults%
2C1%2C1.

Landry:2003:IHP

[Lan03] Walter Landry. Implementing a high performance tensor library. Scientific Programming, 11 (4):273–290, ???? 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Larrabee:1993:PPP

[Lar93] Allan R. Larrabee. P4 parallel programming system, the Linda environment, and some experiences with parallel computation. Scientific Programming, 2(3):23– 35, Fall 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Li:2002:BAS

[LB02] Cen Li and Gautam Biswas. A Bayesian approach for structural learning with hidden Markov models. Scientific Programming, 10(3): 201–219, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Long:2012:SHL

[LBvBW12] Kevin Long, Paul T. Boggs, and Bart G. van Bloemen Waanders. Sundance: High-level software for PDE-constrained optimization. Scientific Programming, 20(3):293– 310, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Li:2024:ABC

[LBW^{+}24] Yufang Li, Manyun Bai, Xin Wang, Di Wu, and Qian Zhao. Artificial bee colony algorithmbased ultrasound image features in the analysis of the influence of different anesthesia methods on lung air volume in orthopedic surgery patients. Scientific Programming, 2024(1): 9958392:1–9958392:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9958392.

Li:2023:IEA

[LBY^{+}23] Dan Li, Jinan Bao, Sizhen Yuan, Hongdong Wang, Likai Wang, and Weiwei Liu. Image enhancement algorithm based on depth difference and illumination adjustment. Scientific Programming, 2023(1):6612471:1– 6612471:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6612471.

Li:2023:SBQ

[LCH^{+}23] Ke Li, Hu Chen, Faxiu Huang, Shenggui Ling, and Zhisheng You. Sharpness and brightness quality assessment of face images for recognition. Scientific Programming, 2023(1):4606828:1–4606828:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4606828.

Li:2024:REP

[LCH^{+}24] Xiangqian Li, Caiyun Chen, Li Huang, Huawei Chen, and Cunquan Huang. Research on the evolution path of China’s provincial innovation chain model based on complex network model. Scientific Programming, 2024(1): 8473021:1–8473021:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8473021.

Li:2023:SVM

[LCK^{+}23] Hu Li, Zhijun Chen, Dezhi Kong, Zhiqiang Huang, and Ningning Wang. Support vector machine-based analysis of holmium laser lithotripsy in treatment of urinary calculus from ultrasound images. Scientific Programming, 2023(1): 9880332:1–9880332:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9880332.

Lin:2023:DLB

[LCL^{+}23a] Haijun Lin, Qi Chen, Caijuan Li, Aifen Zheng, Lei Yang, Jiemin Hong, Hanqing Chen, Xuni He, and Wuna Feng. Deep learning-based automatic detection of rectal polyps using abdominal CT images guided by cold snare polypectomy. Scientific Programming, 2023(1):1179016:1– 1179016:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1179016.

Liu:2023:JOF

[LCL^{+}23b] Zhongyu Liu, Xu Chu, Yan Lu, Wanli Yu, Shuguang Miao, and Enjie Ding. A joint optimization framework of the embedding model and classifier for metalearning. Scientific Programming, 2023(1): 1538914:1–1538914:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1538914.

Li:2022:OEU

[LCW22] Zhi Li, Guihe Chen, and Feng Wang. Optimization of esophageal ultrasound under artificial fish swarm algorithm and its adoption in treatment of ventricular septal defect. Scientific Programming, 2022(1):7126251:1– 7126251:??, 2022. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/7126251.

Li:2023:DSS

[LD23] Wei Li and Zhao Deng. Decision support system in a memristor-based mobile CIM architecture applied on big data computation and storage. Scientific Programming, 2023(1):9041150:1– 9041150:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9041150.

Lemeire:2007:CAP

[LDV07] Jan Lemeire, Erik Dirkx, and Frederik Verbist. Causal analysis for performance modeling of computer programs. Scientific Programming, 15(3):121– 136, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Liu:2023:TMD

[LDW23] Yushan Liu, Huijuan Dong, and Liang Wang. Trampoline motion decomposition method based on deep learning image recognition. Scientific Programming, 2023(1): 1215065:1–1215065:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1215065.

Li:2024:EMT

[LDZQ24] Yiqian Li, Tao Du, Lianjiang Zhu, and Shouning Qu. An efficient minimal text segmentation method for URL domain names. Scientific Programming, 2024(1): 9946729:1–9946729:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9946729.

Liu:2023:EPF

[LDZZ23] Huaijin Liu, Jixiang Du, Yong Zhang, and Hongbo Zhang. Enhancing point features with spatial information for pointbased 3D object detection. Scientific Programming, 2023(1):4650660:1– 4650660:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4650660.

Lesk:2010:BR

[Les10] Michael Lesk. Book review. Scientific Programming, 18(3–4):221– 223, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Li:2023:RNT

[LF23a] Meng Li and Ning Fan. Research on night tourism recommendation based on intelligent image processing technology. Scientific Programming, 2023(1): 2624621:1–2624621:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/2624621.

Lian:2023:JVO

[LF23b] Zhihua Lian and Nana Feng. Job value and organizational socialization of the youth of guangdong-hong kongmacao greater bay area: The mediation of career exploration? Scientific Programming, 2023(1): 9168504:1–9168504:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9168504.

Liao:2020:ERB

[LFQA20] Xiaoqun Liao, Mohammad Faisal, Qing QingChang, and Amjad Ali. Evaluating the role of big data in IIOT-Industrial Internet of Things for executing ranks using the analytic network process approach. Scientific Programming, 2020(1): 8859454:1–8859454:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8859454.

Leiva:2017:ASC

[LFS^{+}17] Claudio Leiva, Víctor Flores, Felipe Salgado, Diego Poblete, and Claudio Acuña. Applying softcomputing for copper recovery in leaching process. Scientific Programming, 2017(1):6459582:1– 6459582:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/6459582.

Luo:2020:OCB

[LFS^{+}20] Cheng Luo, Hongying Fei, Dana Sailike, Tingyi Xu, and Fuzhi Huang. Optimization of continuous berth scheduling by taking into account double-line ship mooring. Scientific Programming, 2020(1):8863994:1– 8863994:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/8863994.

Liu:2018:RPM

[LFY^{+}18] Yang Liu, Yaoling Fan, Xinqing Yan, Xuemei Liu, and Bin Yang. Research on the prewarning method for the safety of south-to-north water transfer project driven by monitoring data. Scientific Programming, 2018(1):3287065:1– 3287065:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2018/3287065.

Lu:2003:MLH

[LG03] Quanming Lu and Vladimir Getov. Mixed-language high-performance computing for plasma simulations. Scientific Programming, 11(1):57–66, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Li:2018:CDA

[LG18] Yan Li and Yan Guo. Cultural distance-aware service recommendation approach in mobile edge computing. Scientific Programming, 2018(1): 2181974:1–2181974:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/2181974.

Lagos:2016:MAC

[LGC^{+}16] Carolina Lagos, Guillermo Guerrero, Enrique Cabrera, Stefanie Niklander, Franklin Johnson, Fernando Paredes, and Jorge Vega. A matheuristic approach combining local search and mathematical programming. Scientific Programming, 2016(??): 1506084:1–1506084:7, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
1506084/.

Li:2018:HHC

[LGJ18] Lei Li, Xue Gao, and Lianwen Jin. HCR-CaaS: a handwritten character recognition container as a service based on QoS guarantee algorithm. Scientific Programming, 2018(1): 6509275:1–6509275:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/6509275.

Lai:2023:SNW

[LGL23] Rongshen Lai, Bo Gao, and Wenguang Lin. Solving no-wait flow shop scheduling problem based on discrete wolf pack algorithm. Scientific Programming, 2023(1): 4731012:1–4731012:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4731012.

Liu:2023:RPT

[LGZ23] Feiping Liu, Qiuling Gong, and Junjie Zhou. Reform of the practice teaching system of entrepreneurship education based on 5G fog computing in colleges and universities. Scientific Programming, 2023(1):2466441:1– 2466441:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2466441.

Li:2023:TMD

[LGZG23] Wen-Ting Li, Shang-Bing Gao, Jun-Qiang Zhang, and Shu-Xing Guo. Training method and device of chemical industry Chinese language model based on knowledge distillation. Scientific Programming, 2023(1):5753693:1– 5753693:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5753693.

Lickly:1993:CMP

[LH93] Daniel J. Lickly and Philip J. Hatcher. C++ and massively parallel computers. Scientific Programming, 2(4):193– 202, Winter 1993. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Li:2023:CSC

[LH23a] Xue Li and Xiaobing Hu.

Current status of ceramic industry and VR technology used in ceramic display and dissemination. Scientific Programming, 2023(1):7555550:1– 7555550:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7555550.

Lu:2023:IRM

[LH23b] Wang Lu and JiangYuan Hou. Intelligent recognition method of athlete wrong movement based on image vision. Scientific Programming, 2023(1):8467906:1– 8467906:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8467906.

Lin:2023:CTI

[LHC^{+}23] Changkun Lin, Chuizhi Huang, Yan Chen, Nailong Jia, Jinghui Huang, and Qimao Fu. Computed tomography images under optimized iterative reconstruction algorithm for blood flow field characteristics in cerebral aneurysm before and after stent implantation. Scientific Programming, 2023(1):8982101:1– 8982101:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8982101.

Li:2023:AWB

[LHMC23] Guihua Li, Chenyu Han, Hong Mei, and Shuai Chen. Application of the WNN-based SCG optimization algorithm for predicting soft soil foundation engineering settlement. Scientific Programming, 2023(1):9936285:1– 9936285:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2021/9936285.

Liu:2023:SDM

[LHZ^{+}23] Dianting Liu, Kangzheng Huang, Chenguang Zhang, Danling Wu, and Shan Wu. Study on discovery method of cooperative research team based on improved Louvain algorithm. Scientific Programming, 2023(1): 3234280:1–3234280:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3234280.

Li:2022:RDA

[Li22] Lina Li. RETRACTION:

Design and application of the piano teaching system integrating videos and images. Scientific Programming, 2022(1):9789404:1– 9789404:??, ???? 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2024/9789404.

Li:2023:RME

[Li23a] Chunlan Li. Research on methods of English text detection and recognition based on neural network detection model. Scientific Programming, 2023(1):

6406856:1–6406856:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6406856.

Li:2023:DFE

[Li23b] Hui Li. Dynamic facial expression recognition using sparse reserved projection algorithm for low illumination images. Scientific Programming, 2023(1):2658471:1– 2658471:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/2658471.

Li:2023:AMH

[Li23c] Juan Li. Analysis of the mental health of urban migrant children based on cloud computing and data mining algorithm models. Scientific Programming, 2023(1):7615227:1– 7615227:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/7615227.

Li:2023:MLB

[Li23d]  Linyan Li.  Machine

learning based image feature recognition and clinical nursing of children’s rheumatoid arthritis-related lung injury. Scientific Programming, 2023(1): 8362318:1–8362318:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8362318.

Li:2023:RIM

[Li23e]  Qiuxia Li.  Research

on the impact mechanism and application of financial digitization and optimization on smalland medium-sized enterprises. Scientific Programming, 2023(1): 9534976:1–9534976:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9534976.

Li:2023:PSD

[Li23f]  Yaling Li.  Psycho-

logical stress detection and early warning system based on wireless network transmission. Scientific Programming, 2023(1):3739045:1– 3739045:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3739045.

Li:2024:RCP

[Li24]  Yunfei Li.  Research

on e-commerce purchasing model in crude oil trade. Scientific Programming, 2024(1):9994965:1– 9994965:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9994965.

Liang:2023:WPP

[Lia23a] Li Liang. Water pollution prediction based on deep belief network in big data of water environment monitoring. Scientific Programming, 2023(1):8271950:1– 8271950:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/8271950.

Liang:2023:IDE

[Lia23b]  Wei Liang.  Innova-

tive development of egg carving cultural and creative products using 3D printing technology based on Internet of Things. Scientific Programming, 2023(1): 3267155:1–3267155:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3267155.

Lin:2004:GTD

[Lin04] H. X. Lin. Graph transformation and designing parallel sparse matrix algorithms beyond data dependence analysis. Scientific Programming, 12 (2):91–100, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Lin:2023:ABD

[Lin23a] Li Lin. Application of big data model in financial taxation management. Scientific Programming, 2023(1):7001456:1– 7001456:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7001456.

Lin:2023:IVP

[Lin23b]  Xiang Lin.  Informa-

tion visualization from the perspective of big data analysis and fusion. Scientific Programming, 2023(1):8934632:1– 8934632:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8934632.

Liu:2020:RIE

[Liu20] Lili Liu. Research on IT English flipped classroom teaching model based on SPOC. Scientific Programming, 2020(1): 7273981:1–7273981:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/7273981.

Liu:2022:RAS

[Liu22] Bingduan Liu. RETRAC-TION: The acceptance status of traditional moral culture in colleges and universities using convolutional neural network. Scientific Programming, 2022(1):9851785:1– 9851785:??, ???? 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2024/9851785.

Liu:2023:PPD

[Liu23a] Chuyi Liu. Process product design based on multimode interaction design. Scientific Programming, 2023(1):1817746:1– 1817746:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2021/1817746.

Liu:2023:DII

[Liu23b] Fang Liu. Design and implementation of intelligent educational administration system using fuzzy clustering algorithm. Scientific Programming, 2023(1): 9485654:1–9485654:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9485654.

Liu:2023:SIF

[Liu23c] Guoying Liu. Scale invariant feature transformation algorithm-based magnetic resonance imagingassisted diagnosis of patients with stroke and rehabilitation nursing. Scientific Programming, 2023(1):4677210:1–4677210:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4677210.

Liu:2023:OKM

[Liu23d] Luosha Liu. Optimizing kenmi manipulation courses of high school sports based on CDIO model under the background of cloud computing. Scientific Programming, 2023(1):9031150:1– 9031150:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9031150.

Liu:2023:AIT

[Liu23e]  Na Liu.  Application

of 3D image technology in motor and sensory nerve classification. Scientific Programming, 2023(1):2219110:1– 2219110:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/2219110.

Liu:2023:CBI

[Liu23f]  Tingting Liu.  Cross-

border Internet of Things e-commerce warehouse control system based on TRIZ theory. Scientific Programming, 2023(1): 2656508:1–2656508:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/2656508.

Liu:2023:SPS

[Liu23g]  Zhihuan Liu.  Short-

est path selection algorithm for cold chain logistics transportation based on improved artificial bee colony. Scientific Programming, 2023(1): 9007162:1–9007162:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9007162.

Liu:2023:MSC

[LJ23] Jun Liu and Xiaolong Jiang. Multislice spiral computed tomography under hybrid iterative reconstruction algorithm in staging diagnosis of bladder cancer and ceramide glycosylation. Scientific Programming, 2023(1):3031282:1– 3031282:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3031282.

Li:2020:BDS

[LJF20] Ju Li, Muhammad Nazir Jan, and Mohammad Faisal. Big data, scientific programming, and its role in internet of industrial things: a decision support system. Scientific Programming, 2020(1):8850096:1– 8850096:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/8850096.

Lu:2023:STR

[LJG^{+}23] Zhengwu Lu, Guosong Jiang, Yurong Guan, Qingdong Wang, and Jianbo Wu. A SAR target recognition method based on decision fusion of multiple features and classifiers. Scientific Programming, 2023(1):1258219:1– 1258219:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/1258219.

Li:2023:ORA

[LJL23] Liqing Li, Linli Jiang, and Zixuan Liu. Optimization research of artificial intelligence and wireless sensor networks in smart pension. Scientific Programming, 2023(1): 5421668:1–5421668:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5421668.

Liu:2023:ESL

[LJSJ23] Shanwei Liu, Yue Jiao, Qinting Sun, and Jinghui Jiang. Estimation of sea level change in the South China Sea from satellite altimetry data. Scientific Programming, 2023(1): 6618135:1–6618135:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6618135.

Liu:2018:DSC

[LJZ18] Zhihan Liu, Yi Jia, and Xiaolu Zhu. Deployment strategy for carsharing depots by clustering urban traffic big data based on affinity propagation. Scientific Programming, 2018(1):3907513:1– 3907513:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/3907513.

Ltaief:2010:STS

[LKDB10] Hatem Ltaief, Jakub Kurzak, Jack Dongarra, and Rosa M. Badia. Scheduling two-sided transformations using tile algorithms on multicore architectures. Scientific Programming, 18(1):35–50, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Lee:2018:IIE

[LKS18] Kyong-Ha Lee, Woo Lam Kang, and Young-Kyoon Suh. Improving i/O efficiency in hadoop-based massive data analysis programs. Scientific Programming, 2018(1): 2682085:1–2682085:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/2682085.

Li:2023:TRS

[LL23a] Jingyu Li and Cungen Liu. Target recognition of synthetic aperture radar images by updated classifiers. Scientific Programming, 2023(1):7181221:1– 7181221:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7181221.

Li:2023:ROM

[LL23b] Linna Li and Renjun Liu. Research on optimal matching scheme of public resource management based on the computational intelligence model. Scientific Programming, 2023(1): 7960972:1–7960972:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7960972.

Li:2023:RMT

[LL23c] Yunpeng Li and Xi Li.

Research on multi-target network security assessment with attack graph expert system model. Scientific Programming, 2023(1):9921731:1–9921731:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9921731.

Liu:2023:DAE

[LL23d] Xiaowei Liu and Hongjin Liu. Design and application of English grammar intelligent question bank system. Scientific Programming, 2023(1): 9483719:1–9483719:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9483719.

Lu:2024:CMA

[LL24] Yu Lu and Wang Lizhi.

Construction of multimedia assisted legal classroom teaching model based on data mining algorithm. Scientific Programming, 2024(1): 9948800:1–9948800:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

                                        wiley.com/doi/epdf/10.
                                        1155/2021/9948800.

Liu:2020:HPM

[LLC^{+}20] Yang Liu, Xiang Li, Xianbang Chen, Xi Wang, and Huaqiang Li. Highperformance machine learning for large-scale data classification considering class imbalance. Scientific Programming, 2020(1): 1953461:1–1953461:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/1953461.

Liu:2023:SNF

[LLC^{+}23] Jian Liu, Wenfeng Liu, Yuhu Cheng, Shuguang Ge, and Xuesong Wang. Similarity network fusion based on random walk and relative entropy for cancer subtype prediction of multigenomic data. Scientific Programming, 2023(1):2292703:1– 2292703:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/2292703.

Lou:2024:BBU

[LLGH24] Yanfang Lou, Bing Liang, Qing Gu, and Zhiwei He. Bispectral-based ultrasound image restoration algorithm for neurological disorders analysis in patients anesthetized with sevoflurane. Scientific Programming, 2024(1):9975089:1– 9975089:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9975089.

Lin:2020:MHA

[LLGL20] Min Lin, Jun-Yan Lyu, Jia-Jing Gao, and Ling-Yu Li. Model and hybrid algorithm of collaborative distribution system with multiple drones and a truck. Scientific Programming, 2020(1):8887057:1– 8887057:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8887057.

Li:2023:CAT

[LLL^{+}23a] Dong Li, Zhanxin Li, Libin Li, Jiangmin Zheng, Qingwei Wang, Li Zhang, and Zhigang Guo. Clinical application of threedimensional reconstruction technique in thoracoscopic pneumonectomy. Scientific Programming, 2023(1):7510437:1– 7510437:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/7510437.

Li:2023:IVI

[LLL^{+}23b] Zongping Li, Wenxin Lei, Xudong Li, Tingting Liao, and Jianming Zhang. Infrared and visible image fusion via fast approximate bilateral filter and local energy characteristics. Scientific Programming, 2023(1):3500116:1– 3500116:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/3500116.

Liu:2023:RBP

[LLL23c] Rui Liu, Ziqi Liu, and Shuyong Liu. Recognition of basketball player’s shooting action based on the convolutional neural network. Scientific Programming, 2023(1): 3045418:1–3045418:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/3045418.

Lian:2023:RSH

[LLLZ23] Zhiqiang Lian, Shanyun Liu, Jieming Lu, and Luxing Zhou. Research on sports and health intelligent diagnosis based on cluster analysis. Scientific Programming, 2023(1):4872947:1– 4872947:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4872947.

Lagraviere:2019:POM

[LLP^{+}19] Jérémie Lagravière, Johannes Langguth, Martina Prugger, Lukas Einkemmer, Phuong Hoai Ha, and Xing Cai. Performance optimization and modeling of fine-grained irregular communication in UPC. Scientific Programming, 2019(1): 6825728:1–6825728:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/6825728.

Li:2023:NMC

[LLP23] Tongyan Li, Yingxiang Li, and Chen Yi-Ping Phoebe. Novel multidimensional collaborative filtering algorithm based on improved item rating prediction. Scientific Programming, 2023(1):2592604:1– 2592604:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/2592604.

Li:2023:DLB

[LLQ^{+}23] Ji Li, Dan Liu, Xiaofeng Qing, Lanlan Yu, and Huizhen Xiang. Deep learningbased three-dimensional transvaginal ultrasound in diagnosis of intrauterine adhesion. Scientific Programming, 2023(1): 2947021:1–2947021:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/2947021.

Liu:2020:LCV

[LLS^{+}20] Zhenpeng Liu, Jiahuan Lu, Nan Su, Bin Zhang, and Xiaofei Li. Locationconstrained virtual machine placement (LCVP) algorithm. Scientific Programming, 2020(1): 8846087:1–8846087:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/8846087.

Lv:2023:KFE

[LLT23] Changhai Lv, Junfeng Li, and Jian Tian. Key frame extraction for sports training based on improved deep learning. Scientific Programming, 2023(1):1016574:1– 1016574:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1016574.

Laney:2014:AED

[LLW^{+}14] Daniel Laney, Steven Langer, Christopher Weber, Peter Lindstrom, and Al Wegener. Assessing the effects of data compression in simulations using physically motivated metrics. Scientific Programming, 22(2):141–155, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Li:2022:CNN

[LLW^{+}22] Guang Li, Fangfang Liu, Yuping Wang, Yongde Guo, Liang Xiao, and Linkai Zhu. A convolutional neural network (CNN) based approach for the recognition and evaluation of classroom teaching behavior. Scientific Programming, 2022(1):6336773:1– 6336773:??, 2022. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6336773.

Liu:2023:ESI

[LLW23a] Jieqiong Liu, Tingting Lei, and Fengyun Wu. Evaluation of severity of infectious pneumonia for newborn using ultrasound image under adaptive image denoising algorithm. Scientific Programming, 2023(1): 6191448:1–6191448:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6191448.

Liu:2023:RRA

[LLW23b] Wang Liu, Xiao Li, and Fengjiao Wu. Research on restoration algorithm of tomb murals based on sequential similarity detection. Scientific Programming, 2023(1):6842353:1– 6842353:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6842353.

Liu:2020:MPC

[LLY^{+}20] Xiazhen Liu, Zhonghua Lu, Wu Yuan, Wenpeng Ma, and Jian Zhang. Massively parallel CFD simulation software: CCFD development and optimization based on Sunway Taihu-Light. Scientific Programming, 2020(1):8847481:1– 8847481:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8847481.

Liao:2023:STA

[LLY^{+}23a] Chunfeng Liao, Hui Luo, Jianqing Yang, Xianliang Wu, and Min Zhao. Speckle tracking algorithm-based cardiac color ultrasound in diagnosis of patients with atrial fibrillation combined with heart failure. Scientific Programming, 2023(1):8457008:1– 8457008:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8457008.

Luo:2023:EPO

[LLY^{+}23b] Jiawei Luo, Lan Lan, Dujiang Yang, Shixin Huang, Mengjiao Li, Jin Yin, Juan Xiao, and Xiaobo Zhou. Early prediction of organ failures in patients with acute pancreatitis using text mining. Scientific Programming, 2023(1):6683942:1– 6683942:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6683942.

Luo:2023:AAI

[LLY23c] Tingting Luo, Guangyao Li, and Naijiang Yu. Application of artificial intelligence and collaborative knowledge for manufacturing design. Scientific Programming, 2023(1): 5846952:1–5846952:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5846952.

Li:2024:MAT

[LLY24] Shisen Li, Chao Liu, and Guoliang Yuan. Martial arts training prediction model based on big data and MEMS sensors. Scientific Programming, 2024(1):9993916:1– 9993916:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9993916.

Liu:2023:CAD

[LLZW23] Hui Liu, Yang Liu, Ran Zhang, and Xia Wu. A clustering algorithm via density perception and hierarchical aggregation based on urban multimodal big data for identifying and analyzing categories of poverty-stricken households in China. Scientific Programming, 2023(1): 6692975:1–6692975:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6692975.

Luo:2023:IRT

[LML^{+}23] Long Luo, Rukuo Ma, Yuan Li, Fangnan Yang, and Zhanfei Qiu. Image recognition technology with its application in defect detection and diagnosis analysis of substation equipment. Scientific Programming, 2023(1):2021344:1– 2021344:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2021344.

Liu:2024:MIE

[LMQ^{+}24] Weina Liu, Jing Ma, Yanxia Qiao, Haiyan Ge, Cuncun Shen, Junran Li, Yaya Qin, and Jingjing Qiu. Medical image enhancement algorithm for b-mode ultrasound image analysis of neonatal respiratory distress syndrome. Scientific Programming, 2024(1): 8552537:1–8552537:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8552537.

Li:2023:RCS

[LMZP23] Jinhai Li, Yunlei Ma, Xiang Zhan, and Jiaming Pei. Research of contextual semantic reasoning model based on domain ontology. Scientific Programming, 2023(1): 4011190:1–4011190:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4011190.

Liao:2020:SPA

[LNH^{+}20] Zitian Liao, Shah Nazir, Anwar Hussain, Habib Ullah Khan, and Muhammad Shafiq. Software piracy awareness, policy, and user perspective in educational institutions. Scientific Programming, 2020(1):6647819:1– 6647819:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/6647819.

Lee:2015:OPE

[LNK^{+}15] Joo Hwan Lee, Nimit Nigania, Hyesoon Kim, Kaushik Patel, and Hyojong Kim. OpenCL performance evaluation on modern multicore CPUs. Scientific Programming, 2015(??):859491:1–859491:20, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
859491/.

Labarta:2001:NOD

[LOHA01] J. Labarta, J. Oliver, D. S. Henty, and Eduard Ayguadé. New OpenMP directives for irregular data access loops. Scientific Programming, 9(2–3):175–183, Spring– Summer 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=7pab6qgbaf8vxg991rwy%
26referrer=parent%26backto=
issue%2C10%2C11%3Bjournal%
2C1%2C9%3Blinkingpublicationresults%
2C1%2C1.

Lou:2023:QAI

[Lou23] Qinqin Lou. Quantitative analysis of the impact of wireless internet technology on college students’ innovation and entrepreneurship under the background of “Internet plus”. Scientific Programming, 2023(1):9282092:1– 9282092:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9282092.

Lin:1999:APS

[LP99] Yuan Lin and David Padua. On the automatic parallelization of sparse and irregular Fortran programs. Scientific Programming, 7(3–4):231– 246, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=53f7mftrrm4r73yyrqau%
26referrer=parent%26backto=
issue%2C4%2C12%3Bjournal%
2C6%2C9%3Blinkingpublicationresults%
2C1%2C1.

Leon-Paredes:2017:HSB

[LPBSSE17] Gabriel A. León-Paredes, Liliana I. Barbosa-Santillán, and Juan J. Sánchez-Escobar. A heterogeneous system based on latent semantic analysis using GPU and multi-CPU. Scientific Programming, 2017(1):8131390:1– 8131390:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/8131390.

Leon-Paredes:2019:SSF

[LPBSSEPL19] Gabriel Alejandro León-Paredes, Liliana Ibeth Barbosa-Santillán, Juan Jaime Sánchez-Escobar, and Antonio Pareja-Lora. Ship-SIBISCaS: a first step towards the identification of potential maritime law infringements by means of LSA-based image. Scientific Programming, 2019(1):1371328:1– 1371328:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/1371328.

Li:2020:SOD

[LPD20] Sibo Li, Roberto Paoli, and Michael D’Mello. Scalability of OpenFOAM density-based solver with Runge–Kutta temporal discretization scheme. Scientific Programming, 2020(1):9083620:1–9083620:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.

                                        wiley.com/doi/epdf/10.
                                        1155/2020/9083620.

Liddell:2004:PCA

[LPHD04] Heather M. Liddell, D. Parkinson, G. S. Hodgson, and P. Dzwig. Parallel computing applications and financial modelling. Scientific Programming, 12 (2):81–90, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Li:2018:ABE

[LQLX18] Deng Li, Liying Qiu, Jiaqi Liu, and Congwen Xiao. Analysis of behavioral economics in crowdsensing: a loss aversion cooperation model. Scientific Programming, 2018(1): 4350183:1–4350183:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2018/4350183.

Lu:2023:RSP

[LQW23] Fei Lu, Wei Qin, and Yu-Xuan Wang. Research on spatial pattern dynamic evolution algorithm and optimization model construction and driving mechanism of provincial tourism eco-efficiency in China under the background of cloud computing. Scientific Programming, 2023(1):1951264:1– 1951264:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1951264.

Liao:2021:HEP

[LQZ^{+}21] Zhifang Liao, Xiaofei Qi, Yan Zhang, Xiaoping Fan, and Yun Zhou. How to evaluate the productivity of software ecosystem: a case study in GitHub. Scientific Programming, 2021(1): 8814247:1–8814247:??, 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8814247.

Lastovetsky:2005:DPM

[LR05] Alexey Lastovetsky and Ravi Reddy. Data partitioning for multiprocessors with memory heterogeneity and memory constraints. Scientific Programming, 13(2):93–112, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Li:2017:UDM

[LRX^{+}17] Xiang Li, Dan Ralescu, Xiaofeng Xu, Lei Li, and Irene Diaz. Uncertain decision-making: a mathematical programming perspective. Scientific Programming, 2017(1): 6372752:1–6372752:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/6372752.

Lam:2012:AIV

[LSB12] Ham Ching Lam, Srinand Sreevatsan, and Daniel Boley. Analyzing influenza virus sequences using binary encoding approach. Scientific Programming, 20(1): 3–13, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Li:2023:DVC

[LSS^{+}23] Xia Li, Zhanqiang Song, Guiping Shen, Ying Huang, and Junyu Chen. Diagnostic value of chest CT images based on full model iterative reconstruction algorithm for lung cancer patients. Scientific Programming, 2023(1):5257682:1–5257682:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5257682.

Liu:2020:SCO

[LSY20] Xinquan Liu, Xiaojing Shen, and Ming You. Study on coordination and optimization of contract farming supply chain based on uncertain conditions. Scientific Programming, 2020(1): 8858812:1–8858812:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8858812.

Lan:2002:DLB

[LTB02] Zhiling Lan, Valerie E.

Taylor, and Greg Bryan. Dynamic load balancing of SAMR applications on distributed systems. Scientific Programming, 10 (4):319–328, 2002. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Li:2023:DCS

[LTF23] Xin Li, Zhang Tao, and Nana Feng. A dualchannel sale system in financially constrained supply chain. Scientific Programming, 2023(1): 8872728:1–8872728:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8872728.

Liu:2023:TEC

[LTL^{+}23] Yongjuan Liu, Yongpan Tan, Rongxia Liu, Xuekui Ye, Lina Wang, Wei Cui, Dan Lu, Jing Guo, and Dongmei Chu. Therapeutic effect of cervical cerclage on cervical insufficiency during pregnancy analyzed by magnetic resonance image under neural network algorithm. Scientific Programming, 2023(1): 5879255:1–5879255:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5879255.

Liu:2023:VMR

[LTR^{+}23] Shuang Liu, Min Tang, Shuqin Ruan, Feng Wei, and Jiaxi Lu. Value of magnetic resonance imaging features in diagnosis and treatment of breast cancer under intelligent algorithms. Scientific Programming, 2023(1): 5289128:1–5289128:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5289128.

Liao:2018:IPR

[LTZS18] Jinzhi Liao, Jiuyang Tang, Xiang Zhao, and Haichuan Shang. Improving POI recommendation via dynamic tensor completion. Scientific Programming, 2018(1): 3907804:1–3907804:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/3907804.

Luo:2023:AMT

[Luo23] Haifeng Luo. Automatic manipulator tracking control based on moving target trajectory prediction. Scientific Programming, 2023(1):7944300:1– 7944300:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7944300.

Lv:2023:DMT

[Lv23] Zongming Lv. The design of mathematics teaching optimization scheme based on data-driven decision-making technology. Scientific Programming, 2023(1):5377784:1– 5377784:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5377784.

Lvping:2023:BTM

[Lvp23] Shen Lvping. Blockchain technology for management of intangible cultural heritage. Scientific Programming, 2023(1): 2613656:1–2613656:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/2613656.

Lin:2023:AIM

[LW23] Yujia Lin and Xin Wang. Analysis and influence of media degradation image propagation path based on image vision. Scientific Programming, 2023(1):6903255:1– 6903255:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6903255.

Lv:2023:RDE

[LWCJ23] Ding Lv, Qunqi Wu, Bo Chen, and Yahong Jiang. Research on decision evaluation model of HOV lane setting. Scientific Programming, 2023 (1):1688824:1–1688824:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2021/1688824.

Liu:2021:RHW

[LWCZ21] Qingxue Liu, Lifang Wang, Yuan Chang, and Jixuan Zhang. Recognition of hotspot words for disease symptoms incorporating contextual weight and co-occurrence degree. Scientific Programming, 2021(1):7863381:1– 7863381:??, ???? 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2024/7863381.

Lei:2023:SMM

[LWL^{+}23a] En Lei, ChaoBo Wang, Wen Xue Li, Yue Dong Wang, Yong Bing Yang, Hua Bin Zheng, and Qi Yuan Tang. Straw mulching with minimum tillage is the best method suitable for straw application under mechanical grain harvesting. Scientific Programming, 2023(1):6878176:1– 6878176:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6878176.

Li:2023:PCA

[LWL23b] Biqiu Li, Jiabin Wang, and Xueli Liu. Parallel cleaning algorithm for similar duplicate Chinese data based on BERT. Scientific Programming, 2023(1):5916748:1–5916748:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5916748.

Li:2023:CCT

[LWL^{+}23c] Jie Li, Wei Wang, Shizhi Long, Xin Liu, Long Huang, and Yuanxun Li. Chest computerized tomography images under iterative model reconstruction algorithm in patients with lung cancer. Scientific Programming, 2023(1):7897784:1– 7897784:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7897784.

Liu:2023:ADM

[LWQ^{+}23] Gang Liu, Wei Wang, He Qin, Qingguo Zhou, Jianbing Ma, Xiaokang Zhou, and Haiyan Zhen. Application of data mining in WITMED: Identification of prognostic genes in oral cancer. Scientific Programming, 2023(1):4907687:1– 4907687:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4907687.

Liu:2024:SAN

[LWQJ24] Yu Liu, Yong Wang, Haojin Qi, and Xiaoming Ju. SuperPruner: Automatic neural network pruning via super network. Scientific Programming, 2024(1):9971669:1– 9971669:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9971669.

Li:2018:OBL

[LWT18] Zhihong Li, Lanteng Wu, and Hongting Tang. Optimizing the borrowing limit and interest rate in P2P system: From borrowers’ perspective. Scientific Programming, 2018(1):2613739:1– 2613739:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/2613739.

Lin:2016:HTS

[LWW16] Weiwei Lin, Wentai Wu, and James Z. Wang. A heuristic task scheduling algorithm for heterogeneous virtual clusters. Scientific Programming, 2016(??):7040276:1–7040276:10, ???? 2016. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2016/
7040276/.

Liu:2023:COH

[LWW23a] Jinhua Liu, Caiping Wang, and Yanhua Wu. Construction and optimization of higher education management system based on internet video online technology. Scientific Programming, 2023(1):5520662:1– 5520662:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5520662.

Liu:2023:DLS

[LWW23b] Wenjian Liu, Baoping Wang, and Wennan Wang. Deep learning software defect prediction methods for cloud environments research. Scientific Programming, 2023 (1):2323100:1–2323100:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2021/2323100.

Wang:2023:OAM

[lWwFAH23] Li li Wang, Xian wen Fang, Esther Asare, and Fang Huan. An optimization approach for mining of process models with infrequent behaviors integrating data flow and control flow. Scientific Programming, 2023(1): 8874316:1–8874316:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8874316.

Liu:2023:ITI

[LWX23a] Jinhua Liu, Caiping Wang, and Xianchun Xiao. Internet of Things (IoT) technology for the development of intelligent decision support education platform. Scientific Programming, 2023(1):6482088:1– 6482088:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6482088.

Lv:2023:III

[LWX23b] Jingjing Lv, Nan Wang, and Shaoxin Xiang. The influence of information interaction behavior on value co-creation business model of online education enterprises performances from the perspective of supply chain. Scientific Programming, 2023(1):4648814:1– 4648814:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4648814.

Li:2016:NMO

[LWYS16] Yongfei Li, Shicheng Wang, Dongfang Yang, and Dawei Sun. A novel metric online monocular SLAM approach for indoor applications. Scientific Programming, 2016 (??):5369780:1–5369780:8, ???? 2016. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2016/
5369780/.

Lei:2023:RPE

[LWZ^{+}23a] Lei Lei, Jian Wu, Shuhai Zheng, Xinyi Zhang, Liang Wang, Yanfei Wang, and Hao Wan. Recognition of power equipment based on multitask sparse representation. Scientific Programming, 2023(1):8322361:1– 8322361:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8322361.

Li:2023:BNN

[LWZ^{+}23b] Yufang Li, Xin Wang, Qian Zhao, Xiaoqing Zhang, and Manyun Bai. Backpropagation neural network artificial intelligence algorithmbased magnetic resonance imaging image feature analysis in the general anesthesia hip arthroplasty. Scientific Programming, 2023(1): 6892979:1–6892979:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6892979.

Liu:2023:EVP

[LWZ^{+}23c] Yanhua Liu, Xiaojie Wu, Yu Zhang, Na Li, and Hua Luan. Evaluation of variable parameter active demons algorithmbased diffusion tensor imaging on ventricular remodeling after myocardial infarction and nursing. Scientific Programming, 2023(1):6419240:1– 6419240:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6419240.

Li:2024:AIT

[LWZ24] Na Li, Ze Wu, and Zhongbiao Zhao. Application of Internet of Things compressed sensing and information interaction technology in intelligent transportation layout. Scientific Programming, 2024(1):9979198:1– 9979198:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9979198.

Liu:2021:DGD

[LWZC21] Yuhui Liu, Yu Wang, Yang Zhang, and Rulei Cheng. Detection of gestational diabetes mellitus and influence on perinatal outcomes from bmode ultrasound images using deep neural network. Scientific Programming, 2021(1):2283768:1– 2283768:??, 2021. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/2283768.

Li:2023:DAB

[LX23a] Tong Li and Yingzhe Xiao. Domain adaptationbased automatic modulation recognition. Scientific Programming, 2023 (1):4277061:1–4277061:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4277061.

Li:2023:MAB

[LX23b] Yousheng Li and Guitao Xia. Multioperator algorithm-based ultrasound imaging combined with hysteroscopy for the diagnosis of endometrial polyps. Scientific Programming, 2023(1): 1808665:1–1808665:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1808665.

Li:2023:SHS

[LX23c] Zijian Li and Chuqiao Xiao. ER-Store: a hybrid storage mechanism with erasure coding and replication in distributed database systems. Scientific Programming, 2023(1):9910942:1– 9910942:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9910942.

Liu:2023:ADT

[LX23d] Yasong Liu and Wentao Xu. Adoption of diffusion tensor imaging under optimized fuzzy C-means cluster algorithm in intracerebral benign and malignant tumor resection. Scientific Programming, 2023(1):3215087:1– 3215087:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/3215087.

Li:2021:RLK

[LXAC21] Chunhua Li, Xuefeng Xian, Xusheng Ai, and Zhiming Cui. Representation learning of knowledge graphs with embedding subspaces. Scientific Programming, 2021(1): 4741963:1–4741963:??, 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/4741963.

Li:2023:MAA

[LXL^{+}23] Tao Li, Yuanyuan Xu, Jiliang Luo, Jianan He, and Shiming Lin. A method of amino acid terahertz spectrum recognition based on the convolutional neural network and bidirectional gated recurrent network model. Scientific Programming, 2023(1): 2097257:1–2097257:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2097257.

Li:2021:PAI

[LXRC21] Yibing Li, Fei Xie, Xiaoye Ren, and Fenyun Cao. The prediction of atherosclerosis index based on photoplethysmograph. Scientific Programming, 2021(1): 2234514:1–2234514:??, 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2234514.

Liu:2024:RKD

[LXW24] Cuiyin Liu, Jishang Xu, and Feng Wang. A review of keypoints’ detection and feature description in image registration. Scientific Programming, 2024(1):8509164:1– 8509164:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8509164.

Li:2018:RSB

[LXWD18] Guangshun Li, Shuzhen Xu, Junhua Wu, and Heng Ding. Resource scheduling based on improved spectral clustering algorithm in edge computing. Scientific Programming, 2018(1): 6860359:1–6860359:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2018/6860359.

Li:2020:SDB

[LY20] Bin Li and Jianmin Yao. Selection of in-domain bilingual sentence pairs based on topic information. Scientific Programming, 2020(1):8879570:1– 8879570:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/8879570.

Li:2023:PNE

[LY23a] Zonghua Li and Zhengwei Ye. A petri nets evolution method that supports BPMN model changes. Scientific Programming, 2023(1):6610795:1–6610795:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2021/6610795.

Liang:2023:RDO

[LY23b] Nan Liang and Mengxuan Yu. Research on design optimization of prefabricated residential houses based on BIM technology. Scientific Programming, 2023(1):1422680:1– 1422680:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1422680.

Liu:2023:IRM

[LYL^{+}23] Zhenpeng Liu, Xianwei Yang, Yi Liu, Yonggang Zhao, and Xiaofei Li. ImReMuDF: Redundant mutants identification method based on definition and reference of variables. Scientific Programming, 2023(1):7543896:1– 7543896:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7543896.

Liu:2024:SCB

[LYL^{+}24] Xiaoyu Liu, Zining Yao, Nan Li, Kai Gao, and Sheng Zhu. Study on the correlation between nonalcoholic fatty liver and metabolic syndrome based on magnetic resonance diffusion-weighted imaging. Scientific Programming, 2024(1): 8756116:1–8756116:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8756116.

Lai:2020:HMC

[LYTL20] Jianqi Lai, Hang Yu, Zhengyu Tian, and Hua Li. Hybrid MPI and CUDA parallelization for CFD applications on multi-GPU HPC clusters. Scientific Programming, 2020(1):8862123:1– 8862123:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8862123.

Liu:2023:FVT

[LYWS23] Yezhen Liu, Xilong Yu, Yanhua Wu, and Shuhong Song. Forecasting variation trends of stocks via multiscale feature fusion and long shortterm memory learning. Scientific Programming, 2023(1):5113151:1– 5113151:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5113151.

Luo:2023:IFM

[LYZ^{+}23] Meimei Luo, Run Yang, Haijie Zhang, Yi Wang, and Shengmei Gao. Image fusion of multislice spiral CT with magnetic resonance imaging (MRI) in the diagnosis and nursing of malignant bone diseases using ANOVA. Scientific Programming, 2023(1): 4751845:1–4751845:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4751845.

Liu:2024:NPO

[LYZY24] He Liu, Zekun Yu, Xiangzhi Zhong, and Helong Yu. Network public opinion monitoring system for agriculture products based on big data. Scientific Programming, 2024(1):9976001:1– 9976001:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9976001.

Li:2023:ASC

[LZ23a] Honger Li and Lixia Zhao.

Algorithmic study of the characteristics of electrocardiograph signals in patients with coronary heart disease. Scientific Programming, 2023(1): 2304072:1–2304072:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2304072.

Lin:2023:REO

[LZ23b] Qingna Lin and Lizheng Zhuo. Research on the evaluation and optimization method of the impact of chorus education on university culture based on coevolution model in the background of artificial intelligence. Scientific Programming, 2023(1):9261934:1– 9261934:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9261934.

Liu:2023:AVC

[LZ23c] Yong Liu and Yanwei Zheng. Accurate volume calculation driven by Delaunay triangulation for coal measurement. Scientific Programming, 2023(1):6613264:1– 6613264:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6613264.

Lu:2023:SAM

[LZ23d] Xinxin Lu and Hong Zhang. Sentiment analysis method of network text based on improved AT-BiGRU model. Scientific Programming, 2023 (1):6669664:1–6669664:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6669664.

Li:2024:TPL

[LZC^{+}24] Lei Li, Yuquan Zhu, Tao Cai, Dejiao Niu, Huaji Shi, and Tingting Zou. A temporal pool learning algorithm based on location awareness. Scientific Programming, 2024(1):9956244:1– 9956244:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9956244.

Li:2016:ROP

[LZF16] Gang Li, Qingpu Zhang, and Zhengqian Feng. Research on optimal path of data migration among multisupercomputer centers. Scientific Programming, 2016(??): 5018213:1–5018213:8, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
5018213/.

Liu:2023:SAD

[LZK^{+}23] Desheng Liu, Hang Zhen, Dequan Kong, Xiaowei Chen, Lei Zhang, Mingrun Yuan, and Hui Wang. Sensors anomaly detection of industrial Internet of Things based on isolated forest algorithm and data compression. Scientific Programming, 2023(1):6699313:1– 6699313:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6699313.

Li:2023:CAL

[LZL^{+}23] Aijun Li, Yuehua Zheng, Yan Li, Tao Zhou, Peicheng Cao, Shaobo Qiu, and Jinpeng Wang. Comparative analysis of lowrank matrix denoising algorithm-based MRI and CT images in diagnosis of cerebral aneurysms. Scientific Programming, 2023(1):2480037:1–2480037:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2480037.

Liu:2023:PDS

[LZLJ23] JiMin Liu, HuiQi Zhao, Chen Liu, and Quan-Qiu Jia. Privacy data security policy of medical cloud platform based on lightweight algorithm model. Scientific Programming, 2023(1):5543714:1– 5543714:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5543714.

Liu:2019:TFB

[LZLZ19] Xiang Liu, Pei Ru Zhu, Ye Liu, and Jing Wen Zhao. Tracking full-body motion of multiple fish with midline subspace constrained multicue optimization. Scientific Programming, 2019(1): 1868797:1–1868797:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/1868797.

Li:2023:SDD

[LZSC23] Xiaoguang Li, Juan Zhu, Haoran Shi, and Zijian Cong. Surface defect detection of seals based on k-means clustering algorithm and particle swarm optimization. Scientific Programming, 2023(1):3965247:1– 3965247:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3965247.

Li:2023:MSG

[LZWJ23] Pengfei Li, Min Zhang, Jian Wan, and Ming Jiang. Multi-scale guided attention network for crowd counting. Scientific Programming, 2023(1): 5596488:1–5596488:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5596488.

Liu:2023:ITD

[LZWL23] Zhengyu Liu, Qingming Zhang, Juyong Wang, and Shibao Lu. Intelligent three-dimensional reconstruction algorithm-based X-ray film for analysis of the changes in sagittal parameters and curative effect of anterior cervical surgery. Scientific Programming, 2023(1): 5194613:1–5194613:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5194613.

Lin:2020:DCB

[LZYJ20] Ping Lin, Xiaosan Zhang, Shuming Yan, and Qingquan Jiang. Dynamic capabilities and business model innovation of platform enterprise: a case study of DiDi taxi. Scientific Programming, 2020(1): 8841368:1–8841368:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8841368.

Luo:2019:SFS

[LZZ19] Qifang Luo, Sen Zhang, and Yongquan Zhou. Stochastic fractal search algorithm for template matching with lateral inhibition. Scientific Programming, 2019(1): 1803934:1–1803934:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/1803934.

Li:2023:HAB

[LZZ23a] Lianhuan Li, Zheng Zhang, and Shaoda Zhang. Hybrid algorithm based on content and collaborative filtering in recommendation system optimization and simulation. Scientific Programming, 2023(1):7427409:1– 7427409:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7427409.

Liu:2023:IUN

[LZZ^{+}23b] Bingshuai Liu, Jiawei Zheng, Hongwei Zhang, Peijie Chen, Shipeng Li, and Yuexian Wen. An improved 2D U-Net model integrated squeezeand-excitation layer for prostate cancer segmentation. Scientific Programming, 2023(1):8666693:1– 8666693:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8666693.

Ma:2022:RRP

[Ma22]  Yanjun Ma.  rmvPF-

BAM: Removing primers from BAM files based on amplicon-based nextgeneration sequencing and cloud computing when analyzing personal genome data. Scientific Programming, 2022(1): 6536470:1–6536470:??, 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6536470.

Ma:2023:BCM

[Ma23a] Jue Ma. Blockchain consensus mechanism based on improved distributed consistency and hash entropy. Scientific Programming, 2023(1):2030810:1– 2030810:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2030810.

Ma:2023:STR

[Ma23b]  Li Ma.  SAR target

recognition using improved sparse representation with local reconstruction. Scientific Programming, 2023(1):2446848:1– 2446848:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2446848.

Ma:2023:AED

[Ma23c]  Qinmin Ma.  Abnor-

mal event detection in videos based on deep neural networks. Scientific Programming, 2023(1): 6412608:1–6412608:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6412608.

Ma:2023:IAH

[Ma23d]  Zongpei Ma.  IoT-

Assisted hybrid intelligent learning architecture based on digital education in a diverse society. Scientific Programming, 2023(1):7203873:1– 7203873:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/7203873.

Majeed:2021:ISS

[MAA^{+}21] Rizwan Majeed, Nurul Azma Abdullah, Imran Ashraf, Yousaf Bin Zikria, Muhammad Faheem Mushtaq, and Muhammad Umer. An intelligent, secure, and smart home automation system. Scientific Programming, 2021(1): 4579291:1–4579291:??, 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/4579291.

Mohammed:2018:FRB

[MABU18] Tareq Abed Mohammed, Shaymaa Alhayali, Oguz Bayat, and Osman N.

Uçan. Feature reduction based on hybrid efficient weighted gene genetic algorithms with artificial neural network for machine learning problems in the big data. Scientific Programming, 2018(1):2691759:1– 2691759:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/2691759.

Moreno:2020:ANM

[MAC^{+}20] R. Moreno, E. Arias, D. Cazorla, J. J. Pardo, A. Navarro, T. Rojo, and F. J. Tapiador. Analysis of a new MPI process distribution for the weather research and forecasting (WRF) model. Scientific Programming, 2020(1): 8148373:1–8148373:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8148373.

McGough:2007:GRT

[MAG^{+}07] A. Stephen McGough, Asif Akram, Li Guo, Marko Krznaric, Luke Dickens, David Colling, Janusz Martyniak, Roger Powell, Paul Kyberd, Chenxi Huang, Constantinos Kotsokalis, and Panayiotis Tsanakas. GRIDCC: a real-time Grid workflow system with QoS. Scientific Programming, 15 (4):213–234, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Malfetti:2001:AOW

[Mal01]  Paolo Malfetti.  Ap-

plication of OpenMP to weather, wave and ocean codes. Scientific Programming, 9 (2–3):99–107, Spring– Summer 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

                                        //iospress.metapress.
                                        com/app/home/contribution.
                                        asp%3Fwasp=7pab6qgbaf8vxg991rwy%
                                        26referrer=parent%26backto=
                                        issue%2C4%2C11%3Bjournal%
                                        2C1%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Mannarswamy:2008:BR

[Man08] Sandya S. Mannarswamy. Book review. Scientific Programming, 16(4):341– 342, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Marowka:2005:EMT

[Mar05] Ami Marowka. Execution model of three parallel languages: OpenMP, UPC and CAF. Scientific Programming, 13(2):127– 135, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Memon:2023:FED

[MAS^{+}23] Zojan Memon, Hamideh Aghian, Muhammad Shahzad Sarfraz, Akhtar Hussain Jalbani, Rozita Jamili Oskouei, Khuda Bux Jalbani, and Ghulam Hussain Jalbani. Framework for educational domain-based multichatbot communication system. Scientific Programming, 2023(1):5518309:1– 5518309:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5518309.

Mattson:1994:ELG

[Mat94] Timothy G. Mattson. Efficiency of Linda for general purpose scientific programming. Scientific Programming, 3(1):61–71, Spring 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Mattson:2003:HGO

[Mat03] Timothy G. Mattson.

How good is OpenMP. Scientific Programming, 11(2):81–93, 2003. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Mu:2017:CKF

[MBBDP17] Wenxin Mu, Frederick Benaben, Nicolas Boissel-Dallier, and Herve Pingaud. Collaborative knowledge framework for mediation information system engineering. Scientific Programming, 2017(1):9026387:1– 9026387:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2017/9026387.

Ma:2020:ASP

[MCB^{+}20] Wenzheng Ma, Yi Cao, Wenzheng Bao, Bin Yang, and Yuehui Chen. ACT-SVM: Prediction of protein-protein interactions based on support vector basis model. Scientific Programming, 2020 (1):8866557:1–8866557:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/8866557.

McCalpin:1996:CSS

[McC96] John D. McCalpin. A case study of some issues in the optimization of Fortran 90 array notation. Scientific Programming, 5(3):219–237, Fall 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL http://reality.
sgi.com/employees/mccalpin/
papers/f90.ps.

Min:2023:CIG

[MCF^{+}23] Huasong Min, Ziming Chen, Bin Fang, Ziwei Xia, Yixu Song, Zongtao Wang, Quan Zhou, Fuchun Sun, and Chunfang Liu. Cross-individual gesture recognition based on long short-term memory networks. Scientific Programming, 2023(1): 6680417:1–6680417:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6680417.

Markus:2010:CFF

[MCvM10] A. A. Markus, W. M. G.

Courage, and M. C. L. M. van Mierlo. A computational framework for flood risk assessment in The Netherlands. Scientific Programming, 18 (2):93–105, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ma:2023:SAE

[MCZ23] Chongsen Ma, Yun Chen, and Yinghui Zhang. Simulation analysis of the evolution of sustainable operation of transport infrastructure projects under government regulation based on prospect theory and BP neural network. Scientific Programming, 2023(1):6868487:1– 6868487:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6868487.

Meng:2023:CCD

[Men23] Xiuying Meng. Concrete crack detection algorithm based on deep residual neural networks. Scientific Programming, 2023 (1):3137083:1–3137083:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3137083.

Metcalf:1999:IFCa

[Met99a]  Mike Metcalf.  Infor-

mation file on compilers, tools, books, courses, tutorials, and the standard for the Fortran language and its derivatives: Version of 20 May 1999 (the penultimate year of the millennium). Scientific Programming, 7(1): 327–333, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Metcalf:1999:IFCb

[Met99b] Mike Metcalf. Information file on compilers, tools, books, courses, tutorials, and the standard for the Fortran language and its derivatives: Version of 20 May 1999 (the penultimate year of the millennium). Scientific Programming, 7(3–4): 327–333, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=53f7mftrrm4r73yyrqau%
26referrer=parent%26backto=
issue%2C11%2C12%3Bjournal%
2C6%2C9%3Blinkingpublicationresults%
2C1%2C1.

Malawski:2015:SMD

[MFB^{+}15] Maciej Malawski, Kamil Figiela, Marian Bubak, Ewa Deelman, and Jarek Nabrzyski. Scheduling multilevel deadlineconstrained scientific workflows on clouds based on cost optimization. Scientific Programming, 2015 (??):680271:1–680271:13, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.

                                        com/journals/sp/2015/
                                        680271/.

Malakar:2013:DCS

[MGK^{+}13] Preeti Malakar, Thomas George, Sameer Kumar, Rashmi Mittal, Vijay Natarajan, Yogish Sabharwal, Vaibhav Saxena, and Sathish S. Vadhiyar. A divide and conquer strategy for scaling weather simulations with multiple regions of interest. Scientific Programming, 21(3–4):93– 107, ???? 2013. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Merlin:1995:IHP

[MH95] John Merlin and Anthony Hey. An introduction to High Performance Fortran. Scientific Programming, 4(2):87–113, Summer 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Michael:1997:RET

[Mic97]  G. A. Michael.  Re-

view: Enabling technologies for petaflops computing, by Thomas Sterling, Paul Messina, Paul H. Smith. Scientific Programming, 6(4):395–397, Winter 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Michalakes:2000:SSP

[Mic00] John Michalakes. The same-source parallel MM5. Scientific Programming, 8 (1):5–12, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=h82chcapth0xynh5tw5w%
26referrer=parent%26backto=
issue%2C1%2C6%3Bjournal%
2C5%2C9%3Blinkingpublicationresults%
2C1%2C1.

Mathur:1995:AAC

[MJ95] Kapil K. Mathur and S. Lennart Johnsson. Allto-all communication on the Connection Machine CM-200. Scientific Programming, 4(4):251–273, Winter 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Muddukrishna:2015:LAT

[MJB15] Ananya Muddukrishna, Peter A. Jonsson, and Mats Brorsson. Localityaware task scheduling and data distribution for OpenMP programs on NUMA systems and manycore processors. Scientific Programming, 2015 (??):981759:1–981759:16,

????  2015.  CO-

DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://www.hindawi.
                                        com/journals/sp/2015/
                                        981759/.

Misko:2023:EEP

[MJK23] Joshua Misko, Shrikant S. Jadhav, and Youngsoo Kim. Extensible embedded processor for convolutional neural networks. Scientific Programming, 2023(1): 6630552:1–6630552:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6630552.

Madison:2007:PCR

[MJLM07] Richard Madison, Abhinandan Jain, Christopher Lim, and Mark Maimone. Performance characterization of a rover navigation algorithm using large-scale simulation. Scientific Programming, 15(2):95–105, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Muhammad:2019:MCM

[MKH^{+}19] Yousaf Shad Muhammad, Saima Khan, Ijaz Hussain, Alaa Mohamd Shoukry, Sadaf Shamsuddin, and Showkat Gani. Minimum cost multiobjective programming model for target efficiency in sample selection. Scientific Programming, 2019(1):7193726:1– 7193726:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/7193726.

Mahalakshmi:2024:PAS

[MKS^{+}24] K. Mahalakshmi, K. Kousalya, Himanshu Shekhar, Aby K. Thomas, L. Bhagyalakshmi, Sanjay Kumar Suman, S. Chandragandhi, Prashant Bachanna, K. Srihari, and Venkatesa Prabhu Sundramurthy. Public auditing scheme for integrity verification in distributed cloud storage system. Scientific Programming, 2024(1): 8533995:1–8533995:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8533995.

Ma:2023:CEA

[ML23a] Ruixin Ma and Junying Lou. CPGAN: an efficient architecture designing for text-to-image generative adversarial networks based on canonical polyadic decomposition. Scientific Programming, 2023(1):5573751:1– 5573751:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5573751.

Miao:2023:TRS

[ML23b] Xinying Miao and Yunlong Liu. Target recognition of SAR images based on complex bidimensional empirical mode decomposition. Scientific Programming, 2023(1):6642316:1– 6642316:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6642316.

Miao:2024:TRS

[ML24] Xinying Miao and Yunlong Liu. Target recognition of SAR images based on azimuthal constraint reconstruction. Scientific Programming, 2024(1): 9974723:1–9974723:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9974723.

Mondaca:2021:FEM

[MLAS21] S. L. Mondaca, C. A. Leiva, C. A. Acuña, and E. A. Serey. Flow enhancement of mineral pastes to increase water recovery in tailings: a Matlab-based imaging processing tool. Scientific Programming, 2021(1): 5607242:1–5607242:??, 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/5607242.

Miao:2020:MBB

[MLHC20] Kehua Miao, Jie Li, Wenxing Hong, and Mingtao Chen. A microservice-based big data analysis platform for online educational applications. Scientific Programming, 2020(1): 6929750:1–6929750:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/6929750.

Meng:2023:DIO

[MLL23] Siming Meng, Ge Lin, and Xiaoyan Liang. Detection, integration, and optimization of acoustic field simulation in the closed space. Scientific Programming, 2023(1): 9301571:1–9301571:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9301571.

Meng:2023:DLB

[MLY23] Yang Meng, Guoxin Liang, and Mei Yue. Deep learning-based arrhythmia detection in electrocardiograph. Scientific Programming, 2023(1): 9926769:1–9926769:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9926769.

Ma:2017:PUA

[MLYH17] Wenpeng Ma, Zhonghua Lu, Wu Yuan, and Xiaodong Hu. Parallelization of an unsteady ALE solver with deforming mesh using OpenACC. Scientific Programming, 2017(1): 4610138:1–4610138:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2017/4610138.

Ma:2023:STC

[MLZ23] Qiangqiang Ma, Wentao Li, and Yongjun Zhang. Subway tunnel construction settlement analysis based on the combination of numerical simulation and neural network. Scientific Programming, 2023(1):4678744:1– 4678744:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4678744.

Muralidharan:2016:BBS

[MM16] D. Muralidharan and R. Muthaiah. Bus based synchronization method for CHIPPER based NoC. Scientific Programming, 2016(??): 1907521:1–1907521:11, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
1907521/.

Moreira:2002:NJH

[MMG^{+}02] José E. Moreira, Samuel P.

Midkiff, Manish Gupta, et al. NINJA: Java for high performance numerical computing. Scientific Programming, 10(1): 19–33, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=9ejnuvwuvby9737jte27%
26referrer=parent%26backto=
issue%2C3%2C9%3Bjournal%
2C2%2C12%3Blinkingpublicationresults%
2C1%2C1.

Morgenstern:1994:MPS

[Mor94] Craig Morgenstern. Methods for precise submesh allocation. Scientific Programming, 3(4):353–364, Winter 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Morris:2015:EMI

[Mor15]  Karla Morris.  Emu-

lating multiple inheritance in Fortran 2003/ 2008. Scientific Programming, 2015(??):126069:1– 126069:7, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://www.hindawi.
                                        com/journals/sp/2015/
                                        126069/.

Mhadhbi:2016:ETH

[MOS16] Imene Mhadhbi, Slim Ben Othman, and Slim Ben Saoud. An efficient technique for hardware/ software partitioning process in codesign. Scientific Programming, 2016(??): 6382765:1–6382765:11, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://www.hindawi.
                                        com/journals/sp/2016/
                                        6382765/.

Mossberg:1997:OOS

[MOT97] Eva Mossberg, Kurt Otto, and Michael Thune.

Object-oriented software tools for the construction of preconditioners. Scientific Programming, 6(3): 285–295, Fall 1997. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Memishi:2016:FBR

[MPA16] Bunjamin Memishi, María S.

Pérez, and Gabriel Antoniu. Feedback-based resource allocation in MapReduce-based systems. Scientific Programming, 2016(??):7241928:1– 7241928:13, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2016/
7241928/.

Miles:2004:TPA

[MPP^{+}04] Simon Miles, Juri Papay, Terry Payne, Michael Luck, and Luc Moreau. Towards a protocol for the attachment of metadata to grid service descriptions and its use in semantic discovery. Scientific Programming, 12(4):201– 211, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Min-qing:2023:ACM

[MqWp23] Zhang Min-qing and Li Wen-ping. An automatic classification method of sports teaching video using support vector machine. Scientific Programming, 2023(1):4728584:1– 4728584:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4728584.

Moreau:2002:MOJ

[MR02] Luc Moreau and Daniel Ribbens. Mobile objects in Java. Scientific Programming, 10(1): 91–100, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=9ejnuvwuvby9737jte27%
26referrer=parent%26backto=
issue%2C9%2C9%3Bjournal%
2C2%2C12%3Blinkingpublicationresults%
2C1%2C1.

Majeed:2018:DNF

[MR18] Shakaiba Majeed and Minsoo Ryu. Debugging nondeterministic failures in Linux programs through replay analysis. Scientific Programming, 2018(1):8939027:1– 8939027:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2018/8939027.

Morris:2012:ECW

[MRLF12] Karla Morris, Damian W. I. Rouson, M. Nicole Lemaster, and Salvatore Filippone. Exploring capabilities within ForTrilinos by solving the 3D Burgers equation. Scientific Programming, 20 (3):275–292, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Maley:2000:CCS

[MS00] David Maley and Ivor Spence. Config: a case study in combining software engineering techniques. Scientific Programming, 8(2): 59–71, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=253x52trrm4r87tkuw1h%
26referrer=parent%26backto=
issue%2C1%2C3%3Bjournal%
2C4%2C9%3Blinkingpublicationresults%
2C1%2C1.

Malik:2020:ESL

[MSA^{+}20] Saud Malik, Ahthasham Sajid, Arshad Ahmad, Ahmad Almogren, Bashir Hayat, Muhammad Awais, and Kyong Hoon Kim. An efficient skewed line segmentation technique for cursive script OCR. Scientific Programming, 2020(1):8866041:1–8866041:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8866041.

Martinez:2014:ELS

[MSCS14] Andrea Martínez, Anna Sikora, Eduardo César, and Joan Sorribes. ELAS-TIC: A large scale dynamic tuning environment. Scientific Programming, 22(4):261–271, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Moon:1999:CCT

[MSH99] Sungdo Moon, Byoungro So, and Mary W. Hall. Combining compiletime and run-time parallelization. Scientific Programming, 7(3–4):247– 260, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=53f7mftrrm4r73yyrqau%
26referrer=parent%26backto=
issue%2C5%2C12%3Bjournal%
2C6%2C9%3Blinkingpublicationresults%
2C1%2C1.

Mubarak:2013:PGA

[MSLS13] Misbah Mubarak, Seegyoung Seol, Qiukai Lu, and Mark S. Shephard. A parallel ghosting algorithm for the flexible distributed mesh database. Scientific Programming, 21(1– 2):17–42, ???? 2013. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Manguoglu:2011:PMS

[MSSG11] Murat Manguoglu, Faisal Saied, Ahmed Sameh, and Ananth Grama. Performance models for the Spike banded linear system solver. Scientific Programming, 19(1):13– 25, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Meng:2020:AII

[MSZZ20] Tian Meng, Mengnan Sun, Yixuan Zhao, and Bo Zhu. Analysis of the impact of interest rate liberalization on financial services management in Chinese commercial banks. Scientific Programming, 2020(1): 8860076:1–8860076:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8860076.

Manoharan:2023:PTA

[MTK^{+}23] Hariprasath Manoharan, Yuvaraja Teekaraman, Ramya Kuppusamy, Arun Radhakrishnan, and Mohamed Yaseen Jabarulla. A prognostic three-axis coordination model for supply chain regulation using machine learning algorithm. Scientific Programming, 2023(1): 1894768:1–1894768:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/1894768.

Mansoor:2020:DLA

[MUJN20] Asif Mansoor, Muhammad Waleed Usman, Noreen Jamil, and M. Asif Naeem. Deep learning algorithm for braincomputer interface. Scientific Programming, 2020 (1):5762149:1–5762149:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/5762149.

Muller:2003:OCB

[Mül03] Matthias S. Müller. An OpenMP compiler benchmark. Scientific Programming, 11(2):125–131, 2003. CODEN SCIPEV.

ISSN 1058-9244 (print), 1875-919X (electronic).

Martinez-Valderrama:2020:AST

[MVIA20] Jaime Martínez-Valderrama, Javier Ibáñez, and Francisco J. Alcalá. AQUA-COAST: a simulation tool to explore coastal groundwater and irrigation farming interactions. Scientific Programming, 2020(1):9092829:1– 9092829:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/9092829.

Martinez-Valderrama:2020:SSA

[MVIAM20] Jaime Martínez-Valderrama, Javier Ibáñez, Francisco J. Alcalá, and Silvio Martínez. SAT: a software for assessing the risk of desertification in Spain. Scientific Programming, 2020(1): 7563928:1–7563928:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/7563928.

Mi:2016:MCF

[MWQ16] Aizhong Mi, Lei Wang, and Junyan Qi. A multiple classifier fusion algorithm using weighted decision templates. Scientific Programming, 2016 (??):3943859:1–3943859:10, ???? 2016. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2016/
3943859/.

Mu:2023:SDW

[MWZ^{+}23] Nan Mu, Hongyu Wang, Yu Zhang, Hongyu Han, and Jun Yang. Saliency detection in weak light images via optimal feature selection-guided seed propagation. Scientific Programming, 2023(1): 9921831:1–9921831:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9921831.

Ma:2019:IOC

[MYH19] Wenpeng Ma, Wu Yuan, and Xiaodong Hu. Implementation and optimization of a CFD solver using overlapped meshes on multiple MIC coprocessors. Scientific Programming, 2019(1):4254676:1– 4254676:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/4254676.

Ma:2021:DDD

[MZLC21] He Ma, Yi Zuo, Tieshan Li, and C. L. Philip Chen. Data-driven decision-support system for speaker identification using e-vector system. Scientific Programming, 2021(1):4748606:1– 4748606:??, 2021. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/4748606.

Nikolopoulos:2003:SNR

[NAA^{+}03] Dimitrios S. Nikolopoulos, Ernest Artiaga, Eduard Ayguadé, et al. Scaling non-regular sharedmemory codes by reusing custom loop schedules. Scientific Programming, 11(2):143–158, 2003. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nagel:2004:BRI

[Nag04] Dan Nagel. Book review: Industrial Strength Parallel Computing: Programming Massively Parallel Processors, by Alice E. Koniges. Scientific Programming, 12(1): 57–62, ???? 2004. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nagle:2005:BRH

[Nag05a] Dan Nagle. Book review: High Performance Linux Clusters, by A. Joseph and D. Sloan. Scientific Programming, 13(2):173– 175, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nagle:2005:BRM

[Nag05b] Dan Nagle. Book review: MPI — The Complete Reference, Vol. 1, The MPI Core, 2nd ed., Scientific and Engineering Computation Series, by Marc Snir, Steve Otto, Steven Huss–Lederman, David Walker and Jack Dongarra. Scientific Programming, 13(1):57–63, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nagle:2009:BRa

[Nag09a] Dan Nagle. Book review. Scientific Programming, 17(3):279–282, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nagle:2009:BRb

[Nag09b] Dan Nagle. Book review. Scientific Programming, 17(4):343–345, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nagle:2011:BRa

[Nag11a] Dan Nagle. Book review. Scientific Programming, 19(1):67–70, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nagle:2011:BRb

[Nag11b] Dan Nagle. Book review. Scientific Programming, 19(4):253–258, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nagle:2012:BR

[Nag12] Dan Nagle. Book review. Scientific Programming, 20(3):349–353, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nanthaamornphong:2015:EUC

[NCMF15] Aziz Nanthaamornphong, Jeffrey Carver, Karla Morris, and Salvatore Filippone. Extracting UML class diagrams from object-oriented Fortran: ForUML. Scientific Programming, 2015 (??):421816:1–421816:15, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
421816/.

Norton:2007:TAM

[NDSG07] Charles D. Norton, Viktor K. Decyk, Bolesław K. Szymanski, and Henry Gardner. The transition and adoption to modern programming concepts for scientific computing in Fortran. Scientific Programming, 15(1):27–44, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nazir:2023:HBD

[NGMAA23] Shah Nazir, Iván García-Magariño, Rodziah Binti Atan, and Shaukat Ali. Healthcare big data management and analytics in scientific programming. Scientific Programming, 2023(1):9780175:1– 9780175:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9780175.

Niu:2023:GAB

[NHHW23] Yuan Niu, Xuejie He, Guijuan Hao, and Liang Wang. Genetic algorithmbased computed tomography image analysis for the diagnosis and mental health of COVID-19 patients in early lowincidence areas. Scientific Programming, 2023(1):

2297206:1–2297206:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2297206.

Nielsen:2008:MCB

[NJ08] Ida M. B. Nielsen and Curtis L. Janssen. Multicore challenges and benefits for high performance scientific computing. Scientific Programming, 16 (4):277–285, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nakano:2002:SAS

[NKV^{+}02] Aiichiro Nakano, Rajiv K.

Kalia, Priya Vashishta, et al. Scalable atomistic simulation algorithms for materials research. Scientific Programming, 10(4): 263–270, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nanthaamornphong:2019:EFA

[NL19] Aziz Nanthaamornphong and Anawat Leatongkam. Extended ForUML for automatic generation of UML sequence diagrams from object-oriented Fortran. Scientific Programming, 2019(1):2542686:1– 2542686:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/2542686.

Ng:2023:IDL

[NL23] Anfernee Joan B. Ng and Kun-Hong Liu. The investigation of different loss functions with capsule networks for speech emotion recognition. Scientific Programming, 2023(1):9916915:1– 9916915:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9916915.

Nguyen:2023:UAE

[NLMT23] Thuy-Anh Nguyen, Hai-Bang Ly, Hai-Van Thi Mai, and Van Quan Tran. Using ANN to estimate the critical buckling load of Y shaped cross-section steel columns. Scientific Programming, 2023(1): 5530702:1–5530702:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5530702.

Novo-Loures:2020:UNL

[NLPL^{+}20] María Novo-Lourés, Reyes Pavón, Rosalía Laza, David Ruano-Ordas, and Jose R. Méndez. Using natural language preprocessing architecture (NLPA) for big data text sources. Scientific Programming, 2020(1): 2390941:1–2390941:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/2390941.

Narayanamoorthy:2016:IOF

[NM16] S. Narayanamoorthy and S. Maheswari. The intelligence of octagonal fuzzy number to determine the fuzzy critical path: A new ranking method. Scientific Programming, 2016(??): 6158208:1–6158208:8, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
6158208/.

Nieves:2018:EWF

[NO18] Juan Carlos Nieves and Mauricio Osorio. Extending well-founded semantics with clark’s completion for disjunctive logic programs. Scientific Programming, 2018(1): 4157030:1–4157030:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

                                        wiley.com/doi/epdf/10.
                                        1155/2018/4157030.

Norton:2007:HPC

[Nor07] Charles D. Norton. High performance computing for mission-enabling space applications. Scientific Programming, 15(2):71– 73, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nikolopoulos:2000:TRD

[NPP^{+}00] Dimitrios S. Nikolopoulos, Theodore S. Papatheodorou, Constantine D. Polychronopoulos, et al. A transparent runtime data distribution engine for OpenMP. Scientific Programming, 8(3): 143–162, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

No:2016:MMC

[NsP16] Jaechun No and Sung soon Park. MultiCache: Multilayered cache implementation for I/O virtualization. Scientific Programming, 2016(??): 3780163:1–3780163:13, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://www.hindawi.
                                        com/journals/sp/2016/
                                        3780163/.

Nguyen-Trang:2019:EHO

[NTNTTK^{+}19] T. Nguyen-Trang, T. Nguyen-Thoi, T. Truong-Khac, A. T. Pham-Chau, and HungLinh Ao. An efficient hybrid optimization approach using adaptive elitist differential evolution and spherical quadratic steepest descent and its application for clustering. Scientific Programming, 2019(1):7151574:1– 7151574:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/7151574.

Ning:2017:PPA

[NYML17] Jianguo Ning, Xinpeng Yuan, Tianbao Ma, and Jian Li. Parallel pseudo arc-length moving mesh schemes for multidimensional detonation. Scientific Programming, 2017 (1):5896940:1–5896940:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/5896940.

OBoyle:1996:EPV

[OB96] M. F. P. O’Boyle and J. M. Bull. Expert programmer versus parallelizing compiler: a comparative study of two approaches for distributed shared memory. Scientific Programming, 5(1):63–88, Spring 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Orellana:2016:FAS

[OCC^{+}16] Julio Proaño Orellana, Blanca Caminero, Carmen Carrión, Luis Tomas, Selome Kostentinos Tesfatsion, and Johan Tordsson. FPGA-aware scheduling strategies at hypervisor level in cloud environments. Scientific Programming, 2016(??): 4670271:1–4670271:12, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://www.hindawi.
                                        com/journals/sp/2016/
                                        4670271/.

Olivier:2013:CMW

[OdSSP13] Stephen L. Olivier, Bronis R. de Supinski, Martin Schulz, and Jan F. Prins. Characterizing and mitigating work time inflation in task parallel programs. Scientific Programming, 21(3–4):123– 136, ???? 2013. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Oehmke:2000:SAA

[OHS00] Robert Oehmke, Janis Hardwick, and Quentin F. Stout. Scalable algorithms for adaptive statistical designs. Scientific Programming, 8(3): 183–193, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

OKeefe:2000:GEP

[O’K00] Matthew O’Keefe. Guesteditorial parallel software design for weather simulation codes. Scientific Programming, 8(1):1–3, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ok:2023:SII

[OKHS23] Jin-Sung Ok, Soon-Do Kwon, Cheol-Eun Heo, and Young-Kyoon Suh. A survey of industrial Internet of Things platforms for establishing centralized data-acquisition middleware: Categorization, experiment, and challenges. Scientific Programming, 2023(1): 6641562:1–6641562:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6641562.

Ouyang:2023:CAR

[OL23] Tianhao Ouyang and Xiaoyong Lu. Clustering analysis of risk divergence of China government’s debts. Scientific Programming, 2023(1):7033597:1– 7033597:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/7033597.

Oldehoeft:2000:BPS

[Old00] Rod Oldehoeft. Best papers from SC2000. Scientific Programming, 8(3): 109–110, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Onan:2019:CCB

[Ona19]  Aytug Onan.  Con-

sensus clustering-based undersampling approach to imbalanced learning. Scientific Programming, 2019(1):5901087:1– 5901087:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2019/5901087.

Ott:2020:FKD

[OPB^{+}20] Jordan Ott, Mike Pritchard, Natalie Best, Erik Linstead, Milan Curcic, and Pierre Baldi. A Fortran– Keras deep learning bridge for scientific computing. Scientific Programming, 2020(1):8888811:1– 8888811:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8888811.

OKeefe:1995:FPT

[OPE^{+}95] Matthew O’Keefe, Terence Parr, B. Kevin Edgar, Steve Anderson, Paul Woodward, and Hank Dietz. Fortran-P translator: towards automatic translation of Fortran 77 programs for massively parallel processors. Scientific Programming, 4(1):1–21, Spring 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ostermann:2011:UNE

[OPP11] Simon Ostermann, Kassian Plankensteiner, and Radu Prodan. Using a new event-based simulation framework for investigating resource provisioning in Clouds. Scientific Programming, 19 (2–3):161–178, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ouyang:2023:ASF

[OQZ23] Chengtian Ouyang, Yaxian Qiu, and Donglin Zhu. Adaptive spiral flying sparrow search algorithm. Scientific Programming, 2023(1): 6505253:1–6505253:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6505253.

OHallaron:1999:SSS

[OS99] David R. O’Hallaron and Bolesław K. Szymanski. Software systems for scalable computers. Scientific Programming, 7(3–4): 191–193, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

                                        //iospress.metapress.
                                        com/app/home/contribution.
                                        asp%3Fwasp=53f7mftrrm4r73yyrqau%
                                        26referrer=parent%26backto=
                                        issue%2C1%2C12%3Bjournal%
                                        2C6%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Oldfield:2012:TST

[OSLK12] Ron A. Oldfield, Gregory D. Sjaardema, Gerald F. Lofstead II, and Todd Kordenbrock. Trilinos I/O Support (Trios). Scientific Programming, 20(2):181–196, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ozturan:1994:CTP

[OSS94] Can Ozturan, Balaram Sinharoy, and Bolesław K. Szymanski. Compiler technology for parallel scientific computation. Scientific Programming, 3 (3):201–225, Fall 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Otto:1993:PAC

[Ott93] Steve W. Otto. Parallel array classes and lightweight sharing mechanisms. Scientific Programming, 2(4):203–216, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ou:2023:MRI

[OWL^{+}23] Xiaoran Ou, Qi Wang, Chunxiao Li, Hongjin Zhao, and Lei Guo. Magnetic resonance imaging based on wavelet algorithm in the diagnosis and treatment of tibial osteomyelitis wound infection. Scientific Programming, 2023(1):2130089:1– 2130089:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2021/2130089.

Oralbekova:2023:NAS

[OZI^{+}23] Zhanar Oralbekova, Tamara Zhukabayeva, Kazizat Iskakov, Makpal Zhartybayeva, Nargiz Yessimova, Alma Zakirova, and Ainur Kussainova. A new approach to solving the problem of atmospheric air pollution in the industrial city. Scientific Programming, 2023(1):8970949:1– 8970949:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8970949.

Ozturan:2004:RBD

[Özt04] Can Özturan. Resource bartering in data grids. Scientific Programming, 12(3):155–168, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Pan:2023:EMT

[Pan23a] Wenxia Pan. English machine translation model based on an improved self-attention technology. Scientific Programming, 2023(1):2601480:1– 2601480:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2601480.

Pan:2023:QAP

[Pan23b] Xiaohui Pan. Quantitative analysis and prediction of global terrorist attacks based on machine learning. Scientific Programming, 2023(1): 7890923:1–7890923:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7890923.

PARKBENCH:1994:PRP

[PAR94] PARKBENCH Committee/Assembled by R.Hockney (Chairman) and M. Berry (Secretary). PARK-BENCH report: Public international benchmarks for parallel computers. Scientific Programming, 3(2):101–146, Summer 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Plazek:2001:CMP

[PBK01] Joanna Płażek, Krzysztof Banaś, and Jacek Kitowski. Comparison of message-passing and shared memory implementations of the GM-RES method on MIMD computers. Scientific Programming, 9(4):195– 209, 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=64t4wprhwk589ellmv56%
26referrer=parent%26backto=
issue%2C1%2C4%3Bjournal%
2C3%2C12%3Blinkingpublicationresults%
2C1%2C1.

Peng:2020:EAS

[PBY^{+}20] Zhihao Peng, Behnam Barzegar, Maryam Yarahmadi, Homayun Motameni, and Poria Pirouzmand. Energy-aware scheduling of workflow using a heuristic method on green cloud. Scientific Programming, 2020(1): 8898059:1–8898059:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8898059.

Perez-Carro:2014:CJM

[PCGLO14] Pascual Pérez-Carro, Francisco Grimaldo, Miguel Lozano, and Juan M. Orduña. Characterization of the Jason multiagent platform on multicore processors. Scientific Programming, 22(1):21– 35, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Prins:1999:ICF

[PCS99] Jan F. Prins, Siddhartha Chatterjee, and Martin Simons. Irregular computations in Fortran — expression and implementation strategies. Scientific Programming, 7(3–4): 313–326, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=53f7mftrrm4r73yyrqau%
26referrer=parent%26backto=
issue%2C10%2C12%3Bjournal%
2C6%2C9%3Blinkingpublicationresults%
2C1%2C1.

Pu:2018:DNH

[PCYZ18] Xun Pu, ShanXiong Chen, XianPing Yu, and Le Zhang. Developing a novel hybrid biogeography-based optimization algorithm for multilayer perceptron training under big data challenge. Scientific Programming, 2018(1): 2943290:1–2943290:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/2943290.

Peachey:2008:FFD

[PDA^{+}08] T. C. Peachey, N. T. Diamond, D. A. Abramson, W. Sudholt, A. Michailova, and S. Amirriazi. Fractional factorial design for parameter sweep experiments using Nimrod/E. Scientific Programming, 16(2–3):217– 230, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Pike:2005:IDP

[PDGQ05] Rob Pike, Sean Dorward, Robert Griesemer, and Sean Quinlan. Interpreting the data: Parallel analysis with Sawzall. Scientific Programming, 13(4):277–298, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Peng:2018:LAC

[PDRC18] Han Peng, Chenglie Du, Lei Rao, and Fu Chen. A LTS approach to control in Event-B. Scientific Programming, 2018(1): 8765186:1–8765186:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2018/8765186.

Perez-Diaz:2016:BAC

[PDROFRM16] N. Pérez-Díaz, D. Ruano-Ordás, F. Fdez-Riverola, and J. R. Méndez. Boosting accuracy of classical machine learning antispam classifiers in real scenarios by applying rough set theory. Scientific Programming, 2016 (??):5945192:1–5945192:10, ???? 2016. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2016/
5945192/.

Peng:2023:RCE

[Pen23] Hua Peng. Research on credit evaluation of financial enterprises based on the genetic backpropagation neural network. Scientific Programming, 2023(1):7745920:1– 7745920:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7745920.

Perrott:2008:MMR

[Per08] Ron Perrott. Memorial:

In memory of Robert Gordon Babb II. Scientific Programming, 16(1): 1, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ping:2023:DSS

[PFQ^{+}23] Wang Ping, Jie Fu, Wenyu Qiao, Muhammad Yasir, Sheng Hui, Md Sakaouth Hossain, and Shah Nazir. Decision support system for hyperspectral remote-sensing data of Yellow River Estuary, China. Scientific Programming, 2023(1): 1376167:1–1376167:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1376167.

Park:2019:BGM

[PGP19] Younghun Park, Minwoo Gu, and Sungyong Park. Ballooning graphics memory space in full GPU virtualization environments. Scientific Programming, 2019(1): 5240956:1–5240956:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/5240956.

Papiani:1995:GBI

[PHH95] Mark Papiani, Anthony J. G. Hey, and Roger W. Hockney. Graphical benchmark information service. Scientific Programming, 4(4):219–227, Winter 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Preissl:2010:OCC

[PKE^{+}10] Robert Preissl, Alice Koniges, Stephan Ethier, Weixing Wang, and Nathan Wichmann. Overlapping communication with computation using OpenMP tasks on the GTS magnetic fusion code. Scientific Programming, 18(3– 4):139–151, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Plale:2004:FBD

[Pla04] Beth Plale. Framework for bringing data streams to the grid. Scientific Programming, 12(4):213– 223, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Pham:2023:CGP

[PLAAH23] Binh Thai Pham, Hai-Bang Ly, Nadhir Al-Ansari, and Lanh Si Ho. A comparison of Gaussian process and M5P for prediction of soil permeability coefficient. Scientific Programming, 2023(1):3625289:1– 3625289:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3625289.

Pan:2023:MCE

[PLJ^{+}23] Wen-Tsao Pan, Yi Liu, Huan Jiang, Ya-Ting Chen, Ting Liu, Yan Qing, Guo-Hui Huang, and Rong Li. Model construction of enterprise financial early warning based on quantum FOA-SVR. Scientific Programming, 2023(1):5018917:1– 5018917:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5018917.

Peng:2020:UDL

[PLW^{+}20] Tao Peng, Jierong Liu, Guojun Wang, Qin Liu, Jianer Chen, and Jiawei Zhu. A userdefined location-sharing scheme with efficiency and privacy in mobile social networks. Scientific Programming, 2020(1): 7832875:1–7832875:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/7832875.

Pan:2023:FRN

[PLY^{+}23] Jie Pan, Li-Ping Li, Chang-Qing Yu, Zhu-Hong You, Zhong-Hao Ren, and Jing-Yu Tang. FWHT-RF: a novel computational approach to predict plant proteinprotein interactions via an ensemble learning method. Scientific Programming, 2023(1):1607946:1– 1607946:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1607946.

Ponnusamy:1994:GCM

[PMCF94] R. Ponnusamy, N. Mansour, A. Choudhary, and G. C. Fox. Graph contraction for mapping data on parallel computers: a quality-cost tradeoff. Scientific Programming, 3 (1):73–82, Spring 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Pech:2017:SAU

[PMEH17] Fernando Pech, Alicia Martinez, Hugo Estrada, and Yasmin Hernandez. Semantic annotation of unstructured documents using concepts similarity. Scientific Programming, 2017(1):7831897:1– 7831897:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/7831897.

Pase:1994:CFP

[PMM94] Douglas M. Pase, Tom MacDonald, and Andrew Meltzer. CRAFT Fortran programming model. Scientific Programming, 3 (3):227–253, Fall 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Papadimitriou:2012:SAP

[PMTL12] Stergios Papadimitriou, Seferina Mavroudi, Kostas Theofilatos, and Spiridon Likothanasis. The software architecture for performing scientific computation with the JLA-PACK libraries in ScalaLab. Scientific Programming, 20(4):379–391, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Papadimitriou:2014:MLS

[PMTL14] Stergios Papadimitriou, Seferina Mavroudi, Kostas Theofilatos, and Spiridon Likothanasis. MATLABlike scripting of Java scientific libraries in ScalaLab. Scientific Programming, 22(3):187–199, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Pang:2023:AMS

[PP23] Yaowen Pang and Xiang Peng. Analysis of mental state of patients after drug addiction and withdrawal guided by PETCT image based on optimized image fusion algorithm. Scientific Programming, 2023(1): 5943410:1–5943410:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5943410.

Preciado:2019:HFD

[PPB^{+}19] Juan Carlos Preciado, Álvaro E. Prieto, Rafael Benitez, Roberto Rodríguez-Echeverría, and José María Conejero. A highfrequency data-driven machine learning approach for demand forecasting in smart cities. Scientific Programming, 2019(1):8319549:1–8319549:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/8319549.

Petcu:2005:DIG

[PPD05] Dana Petcu, Marcin Paprzycki, and Diana Dubu. Design and implementation of a Grid extension for Maple. Scientific Programming, 13(2):137–149, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Paszynska:2015:QOE

[PPJ^{+}15] A. Paszyńska, M. Paszyński, K. Jopek, M. Woźniak, D. Goik, P. Gurgul, H. AbouEisha, M. Moshkov, V. M. Calo, A. Lenharth, D. Nguyen, and K. Pingali. Quasi-optimal elimination trees for 2D grids with singularities. Scientific Programming, 2015 (??):303024:1–303024:18, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://www.hindawi.
                                        com/journals/sp/2015/
                                        303024/.

Pawlowski:2012:AEAa

[PPS12a] Roger P. Pawlowski, Eric T. Phipps, and Andrew G. Salinger. Automating embedded analysis capabilities and managing software complexity in multiphysics simulation, Part I: Templatebased generic programming. Scientific Programming, 20(2):197–219, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Pawlowski:2012:AEAb

[PPS^{+}12b] Roger P. Pawlowski, Eric T. Phipps, Andrew G. Salinger, Steven J.

Owen, Christopher M. Siefert, and Matthew L. Staten. Automating embedded analysis capabilities and managing software complexity in multiphysics simulation, Part II: Application to partial differential equations. Scientific Programming, 20(3):327–345, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Prentice:1999:LE

[Pre99] J. K. Prentice. Letter to the editors. Scientific Programming, 7(1):83–84, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Paul:2014:CEM

[PRM^{+}14] Indrani Paul, Vignesh Ravi, Srilatha Manne, Manish Arora, and Sudhakar Yalamanchili. Coordinated energy management in heterogeneous processors. Scientific Programming, 22(2):93–108, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Prodan:2007:SRW

[Pro07]  Radu Prodan.  Spec-

ification and runtime workflow support in the ASKALON Grid environment. Scientific Programming, 15(4):193–211, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Prabhakar:2023:PAH

[PRW23] Sunil Kumar Prabhakar, Harikumar Rajaguru, and Dong-Ok Won. Performance analysis of hybrid deep learning models with attention mechanism positioning and focal loss for text classification. Scientific Programming, 2023(1):2420254:1– 2420254:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2420254.

Papadimitriou:2015:SGC

[PSM^{+}15] Stergios Papadimitriou, Kirsten Schwark, Seferina Mavroudi, Kostas Theofilatos, and Spiridon Likothanasis. ScalaLab and GroovyLab: Comparing Scala and Groovy for scientific computing. Scientific Programming, 2015(??):498618:1– 498618:13, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
498618/.

Perez:2019:ATP

[PST19] Sergio Pérez, Josep Silva, and Salvador Tamarit. Automatic testing of program slicers. Scientific Programming, 2019(1): 4108652:1–4108652:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/4108652.

Patnaik:2008:INN

[PSU08] Debprakash Patnaik, P. S.

Sastry, and K. P. Unnikrishnan. Inferring neuronal network connectivity from spike data: a temporal data mining approach. Scientific Programming, 16(1):49–77, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Pang:2023:DCC

[PSW23] Meiyu Pang, Jianing Shen, and Lixiu Wu. A distributed congestion control strategy using harmonic search algorithm in internet of vehicles. Scientific Programming, 2023(1):5519492:1– 5519492:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

                                        wiley.com/doi/epdf/10.
                                        1155/2021/5519492.

Pllana:2009:ISP

[PT09a] Sabri Pllana and Jesper Larsson Träff. Introduction to the scientific programming special issue: Software development for multi-core computing systems. Scientific Programming, 17(4):283– 284, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Pllana:2009:RSP

[PT09b] Sabri Pllana and Jesper Larsson Träff. Reviewers for scientific programming special issue on software development for multi-core computing systems. Scientific Programming, 17(4):337, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Park:2013:EBB

[PTS^{+}13] Jongsoo Park, Ping Tak Peter Tang, Mikhail Smelyanskiy, Daehyun Kim, and Thomas Benson. Efficient backprojectionbased synthetic aperture radar computation with many-core processors. Scientific Programming, 21(3–4):165–179, ???? 2013. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Pan:2023:DLL

[PTZL23] Shuguang Pan, Wei Tang, Tiejun Zhou, and Wei Luo. Deep learning in laparoscopic colorectal carcinoma surgery under magnetic resonance imaging. Scientific Programming, 2023(1):1911381:1– 1911381:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1911381.

Paredes-Valverde:2017:SAS

[PVCPdPSZVG17] Mario Andrés Paredes-Valverde, Ricardo Colomo-Palacios, María del Pilar Salas-Zárate, and Rafael Valencia-García. Sentiment analysis in Spanish for improvement of products and services: a deep learning approach. Scientific Programming, 2017(1): 1329281:1–1329281:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/1329281.

Park:2001:PPE

[PVKE01] Insung Park, Michael J.

Voss, Seon Wook Kim, and Rudolf Eigenmann. Parallel programming environment for OpenMP. Scientific Programming, 9(2–3):143–161, Spring– Summer 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=7pab6qgbaf8vxg991rwy%
26referrer=parent%26backto=
issue%2C8%2C11%3Bjournal%
2C1%2C9%3Blinkingpublicationresults%
2C1%2C1.

Parent:2004:ALB

[PVL^{+}04] Johan Parent, Katja Verbeeck, Jan Lemeire, Ann Nowe, Kris Steenhaut, and Erik Dirkx. Adaptive load balancing of parallel applications with multi-agent reinforcement learning on heterogeneous systems. Scientific Programming, 12(2):71–79, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Pan:2021:SEM

[PWP21] Shuang Pan, Jianguo Wei, and Hao Pan. Study on evaluation model of Chinese P2P online lending platform based on hybrid kernel support vector machine. Scientific Programming, 2021(1):4561834:1– 4561834:??, 2021. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/4561834.

Pu:2023:OBF

[PWY23] Aihua Pu, Hua Wang, and Jichong Ying. Optimized backprojection filtration algorithm for postoperative reduction and analysis of respiratory infection-related factors of pelvic fractures by CT imaging. Scientific Programming, 2023(1):3554718:1– 3554718:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3554718.

Pei:2017:IBA

[PY17] Jisheng Pei and Xiaojun Ye. Informationbalance-aware approximated summarization of data provenance. Scientific Programming, 2017 (1):4504589:1–4504589:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/4504589.

Peng:2024:RAR

[PZ24] FenTian Peng and Hongkai Zhang. Research on action recognition method of dance video image based on human-computer interaction. Scientific Programming, 2024(1): 8763133:1–8763133:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8763133.

Peng:2023:TRF

[PZC^{+}23] Han Peng, Xiaoli Zhang, Guozhen Cao, Zhouzhou Liu, Yuejuan Jing, and Lei Rao. A time refinement framework based on iUML-B state machine. Scientific Programming, 2023(1):6672717:1– 6672717:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6672717.

Pan:2023:ICD

[PZH23] Chen Pan, Junling Zhou, and Xiaohua Huang. Impact of check-in data on urban vitality in the Macao Peninsula. Scientific Programming, 2023 (1):7179965:1–7179965:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2021/7179965.

Peng:2019:IBD

[PZL19] Wei Peng, Xinlei Zhang, and Xin Li. Intelligent behavior data analysis for internet addiction. Scientific Programming, 2019(1):2753152:1– 2753152:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/2753152.

Chen:2023:GPB

[qCtWfPbZ23] Ling qing Chen, Mei ting Wu, Li fang Pan, and Ru bin Zheng. Grade prediction in blended learning using multisource data. Scientific Programming, 2023(1):4513610:1– 4513610:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4513610.

Qiu:2023:RIF

[QFX^{+}23] Xiaoyu Qiu, Zhiquan Feng, Tao Xu, Xiaohui Yang, Ya Hou, and Xin Zhang. Research on intention flexible mapping algorithm for elderly escort robot. Scientific Programming, 2023(1):5541269:1– 5541269:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5541269.

Qiu:2020:CMS

[QFYY20] Hong Qiu, Chongdi Fan, Jie Yao, and Xiaohan Ye. Chinese microblog sentiment detection based on CNN-BiGRU and multihead attention mechanism. Scientific Programming, 2020(1):8865983:1– 8865983:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/8865983.

Qiu:2023:CFC

[QGS23] Jing Qiu, Xiaoxu Geng, and Guanglu Sun. Compressed firmware classification based on extra trees and Doc2Vec. Scientific Programming, 2023 (1):2666153:1–2666153:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/2666153.

Qing:2023:SJM

[QH23] Guo Qing and HuBao Hui. Standardized judgment method of shooting training action based on digital video technology. Scientific Programming, 2023(1):4725875:1– 4725875:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4725875.

Qiang:2023:CBO

[QHAN23] Xu Qiang, Zhao Huiqi, Farhad Ali, and Shah Nazir. Criterial based opinion leader’s selection for decision-making using ant colony optimization. Scientific Programming, 2023(1):4624334:1– 4624334:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4624334.

Qi:2023:SDB

[QHJ23] Shanzhong Qi, Zhilei Huang, and Lina Ji. Sustainable development based on green GDP accounting and cloud computing: a case study of Zhejiang Province. Scientific Programming, 2023 (1):7953164:1–7953164:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7953164.

Qin:2023:TTM

[Qin23]  Hao Qin.  Trajectory

tracking method of volleyball player’s arm hitting image based on dp algorithm. Scientific Programming, 2023(1): 4848036:1–4848036:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/4848036.

Qiu:2023:STR

[Qiu23]  Youchun Qiu.  SAR

target recognition based on joint representation of multimode representations. Scientific Programming, 2023(1):6153831:1– 6153831:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6153831.

Qi:2021:FPP

[QK21] Yongqiang Qi and Yi Ke. Fast path planning for onwater automatic rescue intelligent robot based on constant thrust artificial fluid method. Scientific Programming, 2021(1): 8828039:1–8828039:??, 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2020/8828039.

Qu:2023:PSR

[QL23] Guangfu Qu and Won Hyung Lee. Point set registration based on improved KL divergence. Scientific Programming, 2023(1): 1207569:1–1207569:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1207569.

Qi:2023:RTM

[QM23] Min Qi and Hongying Meng. Research on teaching method and class evaluation for international online teaching. Scientific Programming, 2023(1):4120921:1– 4120921:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4120921.

Qiu-Qiang:2023:CSS

[QQWJWYQZ23] Zheng Qiu-Qiang, Liang Wen-Jun, Fan Wan-Yi, and Zhang Qi-Zhe. College students’ selfacceptance: a paint therapy group counseling intervention. Scientific Programming, 2023(1): 5180607:1–5180607:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5180607.

Qiao:2023:HAR

[QS23] Lei Qiao and QiuHao Shen. Human action recognition technology in dance video image. Scientific Programming, 2023(1):6144762:1– 6144762:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6144762.

Quoc:2020:IDE

[QTD^{+}20] Huu Dang Quoc, Loc Nguyen The, Cuong Nguyen Doan, Toan Phan Thanh, and Neal N. Xiong. Intelligent differential evolution scheme for network resources in IoT. Scientific Programming, 2020(1):8860384:1– 8860384:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/8860384.

Qiu:2023:AAO

[QWP23] Zemin Qiu, Feng Wang, and Zhihong Pan. Adaptive adjustment object detection algorithm under multiple mechanisms based on GAN. Scientific Programming, 2023(1): 5875320:1–5875320:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5875320.

Qiao:2020:DCE

[QWZP20] Shicheng Qiao, Qinghu Wang, Jun Zhang, and Zhili Pei. Detection and classification of early decay on blueberry based on improved deep residual 3D convolutional neural network in hyperspectral images. Scientific Programming, 2020(1): 8895875:1–8895875:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8895875.

Qi:2023:DLB

[QYL23] Ruifang Qi, Kun Yang, and Rongmin Li. Deep learning-based ultrasound imaging diagnosis for gonadotropin-releasing hormone agonists treatment of central precocious puberty. Scientific Programming, 2023(1):4512506:1– 4512506:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/4512506.

Qu:2017:MIL

[QYW^{+}17] Xiaobo Qu, Wen Yi, Tingsong Wang, Shuaian Wang, Lin Xiao, and Zhiyuan Liu. Mixedinteger linear programming models for teaching assistant assignment and extensions. Scientific Programming, 2017(1): 9057947:1–9057947:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2017/9057947.

Qi:2017:ICO

[QYZ17] Lianyong Qi, Jiguo Yu, and Zhili Zhou. An invocation cost optimization method for web services in cloud environment. Scientific Programming, 2017(1):4358536:1– 4358536:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2017/4358536.

Qian:2020:OUL

[QZY^{+}20] Tingting Qian, Xiuguo Zheng, Juan Yang, Yeying Xu, Yan Wang, Qiang Zhou, Shenglian Lu, and Xiaotao Ding. Optimal utilization of light energy in semiclosed greenhouse using three-dimensional cucumber model. Scientific Programming, 2020(1): 8855063:1–8855063:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8855063.

Rehman:2023:FMP

[RAA23] Abdul Rehman, Nadeem Akhtar, and Omar H. Alhazmi. Formal modeling, proving, and model checking of a flood warning, monitoring, and rescue system-of-systems. Scientific Programming, 2023 (1):6685978:1–6685978:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6685978.

Rio-Alvarez:2019:ECW

[RAdASGR^{+}19] A. Rio-Alvarez, J. de Andres-Suarez, M. Gonzalez-Rodriguez, D. Fernandez-Lanvin, and B. López Pérez. Effects of challenging weather and illumination on learningbased license plate detection in noncontrolled environments. Scientific Programming, 2019(1): 6897345:1–6897345:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2019/6897345.

Ray:2017:SVP

[Ray17] Partha Pratim Ray. A survey on visual programming languages in Internet of Things. Scientific Programming, 2017(1): 1231430:1–1231430:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2017/1231430.

Ren:2023:FDF

[RAZ^{+}23] Honge Ren, Walid Atwa, Haosu Zhang, Shafiq Muhammad, and Mahmoud Emam. Frame duplication forgery detection and localization algorithm based on the improved Levenshtein distance. Scientific Programming, 2023(1):5595850:1– 5595850:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5595850.

Ramzan:2019:IDA

[RBJ^{+}19] Bushra Ramzan, Imran Sarwar Bajwa, Noreen Jamil, Riaz Ul Amin, Shabana Ramzan, Farhan Mirza, and Nadeem Sarwar. An intelligent data analysis for recommendation systems using machine learning. Scientific Programming, 2019(1):5941096:1– 5941096:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/5941096.

Rodriguez-Echeverria:2019:PBD

[REPRL^{+}19] Roberto Rodriguez-Echeverria, Juan C. Preciado, Álvaro Rubio-Largo, José M. Conejero, and Álvaro E. Prieto. A pattern-based development approach for interaction flow modeling language. Scientific Programming, 2019(1): 7904353:1–7904353:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/7904353.

Rahman:2023:ESL

[RJA^{+}23] Syed Ijaz Ur Rahman, Misbah Jadoon, Sikandar Ali, Hizbullah Khattak, and Jiwei Huang. Efficient segmentation of lymphoblast in acute lymphocytic leukemia. Scientific Programming, 2023 (1):7488025:1–7488025:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/7488025.

Raees:2024:CAS

[RKA^{+}24] Muhammad Raees, Tamim Ahmed Khan, Khurrum Mustafa Abbasi, Afzal Ahmed, Samina Fazilat, and Inaam Ahmed. Contextaware services using MANETs for long-distance vehicular systems: a cognitive agent-based model. Scientific Programming, 2024(1):8835859:1–8835859:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/8835859.

Rehg:1999:ITD

[RKR^{+}99] James M. Rehg, Kathleen Knobe, Umakishore Ramachandran, Rishiyur S. Nikhil, and Arun Chauhan. Integrated task and data parallel support for dynamic applications. Scientific Programming, 7(3– 4):289–302, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        http://iospress.metapress.
                                        com/app/home/contribution.
                                        asp%3Fwasp=53f7mftrrm4r73yyrqau%

26referrer=parent%26backto=
issue%2C8%2C12%3Bjournal%
2C6%2C9%3Blinkingpublicationresults%
2C1%2C1.

Rommelse:2004:EAC

[RLC04] J. R. Rommelse, H. X.

Lin, and T. F. Chan. Efficient active contour and K-means algorithms in image segmentation. Scientific Programming, 12 (2):101–120, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Rackl:2002:EBL

[RLL^{+}02] Günther Rackl, Thomas Ludwig, Markus Lindermeier, et al. Efficiently building on-line tools for distributed heterogeneous environments. Scientific Programming, 10(1): 67–74, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=9ejnuvwuvby9737jte27%
26referrer=parent%26backto=
issue%2C7%2C9%3Bjournal%
2C2%2C12%3Blinkingpublicationresults%
2C1%2C1.

Rubio-Largo:2019:DDC

[RLPI19] Álvaro Rubio-Largo, Juan Carlos Preciado, and Luis Iribarne. Data-driven computational intelligence for scientific programming. Scientific Programming, 2019(1): 5235706:1–5235706:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/5235706.

Raji:2020:NAE

[RLU^{+}20] Mordecai F. Raji, Jian-Ping Li, Amin Ul Haq, Victor Ejianya, Jalaluddin Khan, Asif Khan, Mudassir Khalil, Amjad Ali, Ghufran A. Khan, Mohammad Shahid, Bilal Ahamad, Amit Yadav, and Imran Memon. A new approach for enhancing the services of the 5G mobile network and IOT-Related communication devices using Wavelet-OFDM and its applications in healthcare. Scientific Programming, 2020(1):3204695:1– 3204695:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/3204695.

Raja:2021:SRH

[RMS21] Rakesh Raja, Indrajit Mukherjee, and Bikash Kanti Sarkar. A systematic review of healthcare big data. Scientific Programming, 2021(1):5471849:1– 5471849:??, 2021. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/5471849.

Rouson:2005:DMA

[RMX05] Damian W. I. Rouson, Karla Morris, and Xiaofeng Xu. Dynamic memory de-allocation in Fortran 95/2003 derived type calculus. Scientific Programming, 13(3):189– 203, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Reid:2007:CAN

[RN07] John Reid and Robert W.

Numrich. Co-arrays in the next Fortran Standard. Scientific Programming, 15(1):9–26, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Romberg:2002:UGI

[Rom02] Mathilde Romberg. The UNICORE Grid infrastructure. Scientific Programming, 10(2):149– 157, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=1f99bpyvlg7t46lx8ue3%

26referrer=parent%26backto=
issue%2C6%2C8%3Bjournal%
2C1%2C12%3Blinkingpublicationresults%
2C1%2C1.

Ruano-Ordas:2023:NTM

[ROMFST23] David Ruano-Ordás, Jose R. Méndez, Vítor Basto Fernandes, and Guillermo Suárez-Tangil. Novel tools for the management, representation, and exploitation of textual information. Scientific Programming, 2023(1): 9781923:1–9781923:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9781923.

Rosmond:2000:SVN

[Ros00] Thomas E. Rosmond. A scalable version of the Navy Operational Global Atmospheric Prediction System spectral forecast model. Scientific Programming, 8(1): 31–38, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=h82chcapth0xynh5tw5w%
26referrer=parent%26backto=
issue%2C4%2C6%3Bjournal%
2C5%2C9%3Blinkingpublicationresults%
2C1%2C1.

Rouson:2008:CSC

[Rou08a] Damian W. I. Rouson.

Complexity in scalable computing. Scientific Programming, 16(4):275– 276, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Rouson:2008:TAD

[Rou08b] Damian W. I. Rouson.

Towards analysis-driven scientific software architecture: The case for abstract data type calculus. Scientific Programming, 16(4):329–339, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ren:2020:AAL

[RQD^{+}20] Yanduo Ren, Jiangbo Qian, Yihong Dong, Yu Xin, and Huahui Chen. AVBH: Asymmetric learning to hash with variable bit encoding. Scientific Programming, 2020(1):2424381:1– 2424381:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/2424381.

Rauber:2004:ILO

[RR04] Thomas Rauber and Gudula Rünger. Improving locality for ODE solvers by program transformations. Scientific Programming, 12(3):133– 154, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Rauber:2007:MTD

[RR07] Thomas Rauber and Gudula Rünger. Mixed task and data parallel executions in general linear methods. Scientific Programming, 15(3):137– 155, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Radhakrishnan:2015:UCP

[RRM^{+}15] Hari Radhakrishnan, Damian W. I. Rouson, Karla Morris, Sameer Shende, and Stavros C. Kassinos. Using coarrays to parallelize legacy Fortran applications: Strategy and case study. Scientific Programming, 2015(??):904983:1– 904983:12, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
904983/.

Rico:2009:ATL

[RRV09] Alejandro Rico, Alex Ramirez, and Mateo Valero. Available tasklevel parallelism on the Cell BE. Scientific Programming, 17(1–2):59–76, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Rosing:1994:FLC

[RS94] Matt Rosing and Robert Schnabel. Flexible language constructs for large parallel programs. Scientific Programming, 3(3): 169–186, Fall 1994. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Rosing:1995:LLM

[RS95] Matt Rosing and Joel Saltz. Low latency messages on distributed memory multiprocessors. Scientific Programming, 4 (1):35–43, Spring 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Reussner:2002:SCB

[RST02] Ralf Reussner, Peter Sanders, and Jesper Larsson Träff. SKaMPI: a comprehensive benchmark for public benchmarking of MPI. Scientific Programming, 10 (1):55–65, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.

asp%3Fwasp=9ejnuvwuvby9737jte27%
26referrer=parent%26backto=
issue%2C6%2C9%3Bjournal%
2C2%2C12%3Blinkingpublicationresults%
2C1%2C1.

Raza:2024:EEM

[RuHMS24] Zeeshan Raza, Irfan ul Haq, Muhammad Muneeb, and Omair Shafiq. Energy efficient multiprocessing solo mining algorithms for public blockchain systems. Scientific Programming, 2024(1):9996132:1– 9996132:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9996132.

Ramadoss:2023:CVH

[RVJ^{+}23] Janarthanan Ramadoss, J. Venkatesh, Shubham Joshi, Piyush Kumar Shukla, Sajjad Shaukat Jamal, Majid Altuwairiqi, and Basant Tiwari. Computer vision for humancomputer interaction using noninvasive technology. Scientific Programming, 2023(1):3902030:1– 3902030:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3902030.

Rahman:2020:NNB

[RWN20] Md. Mostafizer Rahman, Yutaka Watanobe, and Keita Nakamura. A neural network based intelligent support model for program code completion. Scientific Programming, 2020(1):7426461:1– 7426461:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/7426461.

Ruan:2017:OMA

[RWY^{+}17] Junhu Ruan, Xuping Wang, Chengyan Yue, Guo Chen, and Minsoo Kim. Optimization models and algorithms for operation and control with advanced information technologies. Scientific Programming, 2017(1): 4706714:1–4706714:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/4706714.

Rouson:2004:DMQ

[RX04] Damian W. I. Rouson and Yi Xiong. Design metrics in quantum turbulence simulations: How physics influences software architecture. Scientific Programming, 12(3):185–196, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Shah:2023:TTA

[SAE^{+}23] Syed Luqman Shah, Irshad Ahmed Abbasi, Alwalid Bashier Gism Elseed, Sikandar Ali, Zahid Anwar, Qasim Rajpoot, and Maria Riaz. TAME^{C}: Trusted augmented mobile execution on cloud. Scientific Programming, 2023(1): 5542852:1–5542852:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5542852.

Sahin:2020:CPI

[Sah20] Ismail Sahin. A comparative performance investigation of swarm optimisers on the design of hydrostatic thrust bearing. Scientific Programming, 2020(1):8856770:1– 8856770:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/8856770.

Sarrafpour:2020:EEA

[SAJ^{+}20] Bahman A. Sassani (Sarrafpour), Mohammed Alkorbi, Noreen Jamil, M. Asif Naeem, and Farhaan Mirza. Evaluating encryption algorithms for sensitive data using different storage devices. Scientific Programming, 2020(1):6132312:1– 6132312:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/6132312.

Shah:2019:NPS

[SAN^{+}19] Syed Ahsin Ali Shah, Wajid Aziz, Malik Sajjad Ahmed Nadeem, Majid Almaraashi, Seong-O. Shim, and Turki M. Habeebullah. A novel phase space reconstruction- (PSR-)Based predictive algorithm to forecast atmospheric particulate matter concentration. Scientific Programming, 2019(1):6780379:1–6780379:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/6780379.

Shang:2024:ASK

[SAZ^{+}24] Rui Shang, Balqees Ara, Islam Zada, Shah Nazir, Zaid Ullah, and Shafi Ullah Khan. Analysis of simple k-mean and parallel k-mean clustering for software products and organizational performance using education sector dataset. Scientific Programming, 2024(1): 9988318:1–9988318:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9988318.

Smith:2001:DMM

[SB01] Lorna Smith and Mark Bull. Development of mixed mode MPI/ OpenMP applications. Scientific Programming, 9(2–3):83–98, Spring– Summer 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

                                        //iospress.metapress.
                                        com/app/home/contribution.
                                        asp%3Fwasp=7pab6qgbaf8vxg991rwy%
                                        26referrer=parent%26backto=
                                        issue%2C3%2C11%3Bjournal%
                                        2C1%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Srirama:2011:SPS

[SBJV11] Satish Narayana Srirama, Oleg Batrashev, Pelle Jakovits, and Eero Vainikko. Scalability of parallel scientific applications on the cloud. Scientific Programming, 19 (2–3):91–105, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Samad:2024:MTC

[SBK^{+}24] Ali Samad, Hairulnizam Bin Mahdin, Rafaqat Kazmi, Rosziati Ibrahim, and Zirawani Baharum. Multiobjective test case prioritization using test case effectiveness: Multicriteria scoring method. Scientific Programming, 2024(1):9988987:1–9988987:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9988987.

Shan:2010:PMP

[SBM^{+}10] Hongzhang Shan, Filip Blagojević, Seung-Jai Min, Paul Hargrove, Haoqiang Jin, Karl Fuerlinger, Alice Koniges, and Nicholas J. Wright. A programming model performance study using the NAS parallel benchmarks. Scientific Programming, 18(3–4):153– 167, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Shen:2016:POS

[SC16] Zhe-Ping Shen and Walter W. Chen. Profile orientation and slope stability analysis. Scientific Programming, 2016(??): 7029786:1–7029786:10, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
7029786/.

Shan:2023:DAI

[SC23] Xiaowen Shan and Hao Chen. Design of automatic integration algorithm for popular science microanimation works in the context of new media. Scientific Programming, 2023(1):8269439:1– 8269439:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8269439.

Soto:2016:EPS

[SCAP16] Ricardo Soto, Broderick Crawford, Boris Almonacid, and Fernando Paredes. Efficient parallel sorting for migrating birds optimization when solving machine-part cell formation problems. Scientific Programming, 2016(??): 9402503:1–9402503:39, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
9402503/.

Smallen:2002:AST

[SCB02] Shava Smallen, Henri Casanova, and Francine Berman. Applying scheduling and tuning to on-line parallel tomography. Scientific Programming, 10(4):271–289, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Schweitzer:2015:PEM

[SCD^{+}15] P. Schweitzer, S. Cipière, A. Dufaure, H. Payno, Y. Perrot, D. R. C. Hill, and L. Maigne. Performance evaluation of multithreaded Geant4 simulations using an Intel Xeon Phi cluster. Scientific Programming, 2015 (??):980752:1–980752:10, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
980752/.

Sun:2021:TEC

[SCD21] Shuang Sun, Shudong Chen, and Rong Du. Trusted and efficient cross-domain access control system based on blockchain. Scientific Programming, 2021(1): 8832568:1–8832568:??, 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/8832568.

Schmidt:1994:UNS

[Sch94] Henning Schmidt. Using naming strategies to make massively parallel systems work. Scientific Programming, 3(4):289–300, Winter 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Schonfelder:2003:VPA

[Sch03a] J. L. Schonfelder. Variable precision arithmetic: a Fortran 95 module. Scientific Programming, 11 (1):67–76, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

                                        pcwww.liv.ac.uk/~jls/
                                        vpa20.f95;   http://
                                        pcwww.liv.ac.uk/~jls/
                                        vpa20.htm.

Schupp:2003:LBC

[Sch03b] Sibylle Schupp. Lifting a butterfly — a componentbased FFT. Scientific Programming, 11(4):291– 307, ???? 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Sun:2023:RDF

[SCKW23] Lingyan Sun, Yu Chen, Chuiyue Kong, and Jinghua Wang. Research on distributed feeder automation communication based on XMPP and GOOSE. Scientific Programming, 2023(1): 6650725:1–6650725:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6650725.

Strecansky:2020:CMI

[SCR20] Peter Strecanský, Stanislav Chren, and Bruno Rossi. Comparing maintainability index, SIG method, and SQALE for technical debt identification. Scientific Programming, 2020(1):2976564:1– 2976564:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/2976564.

Sahni:2009:SSA

[SCSJ09] Onkar Sahni, Christopher D. Carothers, Mark S. Shephard, and Kenneth E. Jansen. Strong scaling analysis of a parallel, unstructured, implicit solver and the influence of the operating system interference. Scientific Programming, 17 (3):261–274, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Song:2023:RXF

[SCW^{+}23] Xiangyang Song, Xiang Chen, Xiaodong Wang, Nitu Wu, Aijun Liu, Ritu Su, Yong Yang, Shujuan Chang, and Zhijun Wei. Responses of Xilingol fractional vegetation cover to climate change and human activities. Scientific Programming, 2023(1):5169913:1– 5169913:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5169913.

Sun:2018:ABC

[SCZ18] Lijun Sun, Tianfei Chen, and Qiuwen Zhang. An artificial bee colony algorithm with random location updating. Scientific Programming, 2018(1): 2767546:1–2767546:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2018/2767546.

Shan:2023:WSC

[SD23] Chuang Shan and Yugen Du. A web service clustering method based on semantic similarity and multidimensional scaling analysis. Scientific Programming, 2023(1):6661035:1– 6661035:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6661035.

Salguero:2019:OBF

[SDM^{+}19] Alberto G. Salguero, Pablo Delatorre, Javier Medina, Macarena Espinilla, and Antonio J. Tomeu. Ontology-based framework for the automatic recognition of activities of daily living using class expression learning techniques. Scientific Programming, 2019(1): 2917294:1–2917294:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/2917294.

Schattler:2000:RPP

[SDS00] Ulrich Schättler, Günther Doms, and Jürgen Steppeler. Requirements and problems in parallel model development at DWD. Scientific Programming, 8(1):13– 22, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

                                        //iospress.metapress.
                                        com/app/home/contribution.
                                        asp%3Fwasp=h82chcapth0xynh5tw5w%
                                        26referrer=parent%26backto=
                                        issue%2C2%2C6%3Bjournal%
                                        2C5%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Skinner:1995:PPC

[SE95] Gregg Skinner and Rudolf Eigenmann. Parallel performance of a combustion chemistry simulation. Scientific Programming, 4 (3):127–139, Fall 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Serafini:2018:ISC

[Ser18] Paolo Serafini. An iterative scheme to compute size probabilities in random graphs and branching processes. Scientific Programming, 2018(1): 3791075:1–3791075:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2018/3791075.

Sheshadri:2003:GSPb

[SF03a] K. Sheshadri and Peter Fritzson. A general symbolic PDE solver generator: Beyond explicit schemes. Scientific Programming, 11(3):225–235, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Sheshadri:2003:GSPa

[SF03b] K. Sheshadri and Peter Fritzson. A general symbolic PDE solver generator: Explicit schemes. Scientific Programming, 11(1):39–55, 2003. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Su:2023:FCD

[SF23] Fei Su and Zhe Fan.

Flipped classroom design of college ideological and political courses based on long short-term memory networks. Scientific Programming, 2023(1): 6971906:1–6971906:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6971906.

Shi:2020:CTE

[SFZ^{+}20] Congcong Shi, Jiaxuan Fei, Xiaojian Zhang, Qigui Yao, and Jie Fan. Continuous trust evaluation of power equipment and users based on risk measurement. Scientific Programming, 2020(1): 8895804:1–8895804:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8895804.

Smith:1996:DDS

[SG96] Barry F. Smith and William D. Gropp. Design of data-structure-neutral libraries for the iterative solution of sparse linear systems. Scientific Programming, 5(4):329–336, Winter 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Saravanakumar:2023:ETV

[SGK^{+}23] C. Saravanakumar, M. Geetha, S. Manoj Kumar, S. Manikandan, C. Arun, and K. Srivatsan. An efficient technique for virtual machine clustering and communications using task-based scheduling in cloud computing. Scientific Programming, 2023(1):5586521:1– 5586521:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5586521.

Schulz:2008:OSO

[SGM^{+}08] Martin Schulz, Jim Galarowicz, Don Maghrak, William Hachfeld, David Montoya, and Scott Cranford. Open — SpeedShop: an open source infrastructure for parallel performance analysis. Scientific Programming, 16(2– 3):105–121, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Soto:2019:DSA

[SGPCLG19] Ricardo Soto, Juan A.

Gómez-Pulido, Stéphane Caro, and José M. Lanza-Gutiérrez. Data science and AI-based optimization in scientific programming. Scientific Programming, 2019(1):7154765:1– 7154765:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/7154765.

Schonauer:1994:EGB

[SH94] W. Schonauer and H. Hafner.

Explaining the gap between theoretical peak performance and real performance for supercomputer architectures. Scientific Programming, 3(2):157–168, Summer 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Stupar:2019:MBE

[SH19] Ivana Stupar and Darko Huljenić. Model-based extraction of knowledge about the effect of cloud application context on application service cost and quality of service. Scientific Programming, 2019(1):5075412:1– 5075412:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/5075412.

Shi:2023:AOD

[SHCZ23] Peng Shi, Bin Hou, Jing Chen, and Yunxiao Zu. An algorithm of occlusion detection for the surveillance camera. Scientific Programming, 2023(1):6698160:1– 6698160:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6698160.

Shen:2023:ATL

[She23] Hexue Shen. Application of transfer learning algorithm and real time speech detection in music education platform. Scientific Programming, 2023(1):1093698:1– 1093698:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1093698.

Sato:2001:CEO

[SHHI01] Mitsuhisa Sato, Hiroshi Harada, Atsushi Hasegawa, and Yutaka Ishikawa. Cluster-enabled OpenMP: An OpenMP compiler for the SCASH software distributed shared memory system. Scientific Programming, 9 (2–3):123–130, Spring– Summer 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=7pab6qgbaf8vxg991rwy%
26referrer=parent%26backto=
issue%2C6%2C11%3Bjournal%
2C1%2C9%3Blinkingpublicationresults%
2C1%2C1.

Skillicorn:1997:QAA

[SHM97] David B. Skillicorn, Jonathan M. D. Hill, and W. F. Mc-Coll. Questions and answers about BSP. Scientific Programming, 6(3): 249–274, Fall 1997. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

ftp://ftp.comlab.ox.
ac.uk/pub/Documents/
techpapers/Jonathan.
Hill/SkillHillMcColl_
QA.ps.gz.

Shu:1994:ADP

[Shu94] Wei Shu. Adaptive dynamic process scheduling on distributed memory parallel computers. Scientific Programming, 3 (4):341–352, Winter 1994.

CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Sun:2023:ACS

[SHW23] Jian Sun, Guobin Hu, and Chenghua Wang. Analog circuit soft fault diagnosis based on sparse random projections and k-nearest neighbor. Scientific Programming, 2023(1): 8040140:1–8040140:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/8040140.

Sinharoy:1999:COI

[Sin99] Balaram Sinharoy. Compiler optimization to improve data locality for processor multithreading. Scientific Programming, 7(1):21–37, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

                                        //iospress.metapress.
                                        com/app/home/contribution.
                                        asp%3Fwasp=64cr5a4mg33tuhcbdr02%
                                        26referrer=parent%26backto=
                                        issue%2C2%2C7%3Bjournal%
                                        2C8%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Shahrivari:2015:FPA

[SJ15] Saeed Shahrivari and Saeed Jalili. Fast parallel all-subgraph enumeration using multicore machines. Scientific Programming, 2015 (??):901321:1–901321:11, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
901321/.

Setia:2020:HHM

[SJD20] Sonia Setia, Verma Jyoti, and Neelam Duhan. HPM: a hybrid model for user’s behavior prediction based on Ngram parsing and access logs. Scientific Programming, 2020(1):8897244:1– 8897244:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8897244.

Schonherr:2014:TSS

[SJR14] Jan H. Schönherr, Ben Juurlink, and Jan Richling. TACO: A scheduling scheme for parallel applications on multicore architectures. Scientific Programming, 22(3):223– 237, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Somu:2021:ALE

[SK21] Renugadevi Somu and Preethi Ashok Kumar. Analysis of learner’s emotional engagement in online learning using machine learning Adam robust optimization algorithm. Scientific Programming, 2021(1): 8886197:1–8886197:??, ???? 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2024/8886197.

Shah:2020:CSM

[SKA^{+}20a] Sahar Shah, Mahnoor Khan, Ahmad Almogren, Ihsan Ali, Lianwen Deng, Heng Luo, and Muazzam A. Khan. Corrigendum to “Security measurement in industrial IoT with cloud computing perspective: Taxonomy, issues, and future directions”. Scientific Programming, 2020(1): 3671835:1–3671835:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/3671835.

Shah:2020:SMI

[SKA^{+}20b] Sahar Shah, Mahnoor Khan, Ahmad Almogren, Ihsan Ali, Lianwen Deng, Heng Luo, and Muazzam A. Khan. Security measurement in industrial IoT with cloud computing perspective: Taxonomy, issues, and future directions. Scientific Programming, 2020(1):8871315:1– 8871315:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8871315.

Saleem:2024:IBC

[SKB^{+}24] Rana Muhammad Saleem, Rafaqat Kazmi, Imran Sarwar Bajwa, Amna Ashraf, Shabana Ramzan, and Waheed Anwar. IOTbased cotton whitefly prediction using deep learning. Scientific Programming, 2024(1):8824601:1– 8824601:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8824601.

Song:2024:FTL

[SKHZ24] Yunsheng Song, Xiaohan Kong, Shuoping Huang, and Chao Zhang. Fast training logistic regression via adaptive sampling. Scientific Programming, 2024(1):9991859:1– 9991859:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

                                        wiley.com/doi/epdf/10.
                                        1155/2021/9991859.

Satoh:2001:COT

[SKS01] Shigehisa Satoh, Kazuhiro Kusano, and Mitsuhisa Sato. Compiler optimization techniques for OpenMP programs. Scientific Programming, 9 (2–3):131–142, Spring– Summer 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

                                        //iospress.metapress.
                                        com/app/home/contribution.
                                        asp%3Fwasp=7pab6qgbaf8vxg991rwy%
                                        26referrer=parent%26backto=
                                        issue%2C7%2C11%3Bjournal%
                                        2C1%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Shi:2024:SAI

[SKT^{+}24] Wenjing Shi, Xiangyu Kong, Wei Tian, Yujin Yan, and Yusi Chen. Segmentation algorithm for intracranial magnetic resonance images for cerebral stroke identification and nursing evaluation. Scientific Programming, 2024(1):9996977:1– 9996977:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9996977.

Shi:2009:ISC

[SKU^{+}09] Guochun Shi, Volodymyr V.

Kindratenko, Ivan S. Ufimtsev, Todd J. Martinez, James C. Phillips, and Steven A. Gottlieb. Implementation of scientific computing applications on the Cell Broadband Engine. Scientific Programming, 17(1– 2):135–151, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Shui:2021:IPD

[SL21] Wenbing Shui and Mengxia Li. Integrated pricing and distribution planning for community group purchase of fresh agricultural products. Scientific Programming, 2021(1): 8839398:1–8839398:??, 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8839398.

Shi:2023:HRB

[SL23a] Weiwei Shi and Qiuzuo Li. Human resources balanced allocation method based on deep learning algorithm. Scientific Programming, 2023(1): 4681959:1–4681959:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4681959.

Sun:2023:SHM

[SL23b] Lina Sun and Mingzhi Li. Sports and health management using big data based on voice feature processing and Internet of Things. Scientific Programming, 2023(1): 3271863:1–3271863:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/3271863.

Sun:2023:DPP

[SL23c] Wei Sun and Lijun Li. Design of process products based on image processing multimode interaction. Scientific Programming, 2023(1):9925764:1– 9925764:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9925764.

Shi:2016:HBP

[SLBZ16] Chenghua Shi, Tonglei Li, Yu Bai, and Fei Zhao. A heuristicsbased parthenogenetic algorithm for the VRP with potential demands and time windows. Scientific Programming, 2016(??): 8461857:1–8461857:12, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
8461857/.

Shi:2023:MCA

[SLC^{+}23] Lei Shi, Jia Luo, Gang Cheng, Xia Liu, and Gang Xie. A multifeature complementary attention mechanism for image topic representation in social networks. Scientific Programming, 2023(1): 5304321:1–5304321:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5304321.

Sun:2016:MSR

[SLD^{+}16] Xiaobing Sun, Bin Li, Yucong Duan, Wei Shi, and Xiangyue Liu. Mining software repositories for automatic interface recommendation. Scientific Programming, 2016(??): 5475964:1–5475964:11, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
5475964/.

Sun:2017:UHL

[SLDL17] Xiaobing Sun, Xiangyue Liu, Yucong Duan, and Bin Li. Using hierarchical latent Dirichlet allocation to construct feature tree for program comprehension. Scientific Programming, 2017(1):4382348:1– 4382348:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/4382348.

Shen:2023:DLB

[SLH23] Zhuxiang Shen, Wei Li, and Hui Han. Deep learning-based wavelet threshold function optimization on noise reduction in ultrasound images. Scientific Programming, 2023(1):3471327:1– 3471327:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3471327.

Song:2000:MST

[SLJ^{+}00] H. J. Song, X. Liu, D. Jakobsen, et al. The MicroGrid: a scientific tool for modeling Computational Grids. Scientific Programming, 8(3): 127–141, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Sun:2017:CCP

[SLL^{+}17] Xiaobing Sun, Xiangyue Liu, Bin Li, Bixin Li, David Lo, and Lingzhi Liao. Clustering classes in packages for program comprehension. Scientific Programming, 2017(1):3787053:1– 3787053:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/3787053.

Slota:2012:SQP

[Slo12]  Renata Slota.  Stor-

age QoS provisioning for execution programming of data-intensive applications. Scientific Programming, 20(1):69–80, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Siddiqui:2018:SBH

[SLP18] Zahid Ali Siddiqui, Jeong-A Lee, and Unsang Park. SEDC-based hardware-level fault tolerance and fault secure checker design for big data and cloud computing. Scientific Programming, 2018(1):7306837:1– 7306837:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/7306837.

Shan:2017:RAL

[SLS17] Xian Shan, Kang Liu, and Pei-Liang Sun. Risk analysis on leakage failure of natural gas pipelines by fuzzy Bayesian network with a bow-tie model. Scientific Programming, 2017(1):3639524:1–3639524:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2017/3639524.

Sharma:2017:MCF

[SM17] Anuj Sharma and Irene Moulitsas. MPI to Coarray Fortran: Experiences with a CFD solver for unstructured meshes. Scientific Programming, 2017 (1):3409647:1–3409647:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2017/3409647.

Shang:2023:SCC

[SMA^{+}23] Rui Shang, YongMei Ma, Farhad Ali, ChuanShuang Hu, Shah Nazir, Huafei Wei, and Abdullah Khan. Selection of crowd in crowdsourcing for smart intelligent applications: a systematic mapping study. Scientific Programming, 2023(1):9368128:1– 9368128:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9368128.

Shengxin:2017:CSR

[SMB17] Dai Shengxin, Hong Mei, and Guo Bing. A comparative study of reliability-ignorant and reliability-aware energy management schemes using UPPAAL-SMC. Scientific Programming, 2017 (1):2621089:1–2621089:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/2621089.

Soheili:2020:SAP

[SMD20] Majid Soheili, Amir-Masoud Eftekhari Moghadam, and Mehdi Dehghan. Statistical analysis of the performance of rank fusion methods applied to a homogeneous ensemble feature ranking. Scientific Programming, 2020(1):8860044:1– 8860044:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8860044.

Schopf:2002:GTT

[SN02] Jennifer M. Schopf and Bill Nitzberg. Grids: The top ten questions. Scientific Programming, 10(2): 103–111, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

                                        //iospress.metapress.
                                        com/app/home/contribution.
                                        asp%3Fwasp=1f99bpyvlg7t46lx8ue3%
                                        26referrer=parent%26backto=
                                        issue%2C2%2C8%3Bjournal%
                                        2C1%2C12%3Blinkingpublicationresults%
                                        2C1%2C1.

Snelling:1995:AAG

[Sne95] David F. Snelling. Applications analysis: Guest Editor’s introduction. Scientific Programming, 4(3):123–??, Fall 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Sun:2024:MDM

[SNH24] Lei Sun, Shah Nazir, and Anwar Hussain. Multicriteria decision making to continuous software improvement based on quality management, assurance, and metrics. Scientific Programming, 2024(1):9953618:1– 9953618:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2021/9953618.

Snyder:2007:SPF

[Sny07] W. Van Snyder. Scientific programming in Fortran. Scientific Programming, 15(1):3–8, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Sharp:1993:CSO

[SO93] Michael D. Sharp and Steve W. Otto. Classspecific optimizing compiler. Scientific Programming, 2(4):235–238, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Sen:2011:SII

[SO11] Alper Sen and Can Ozturan. Special issue on the 9th International Symposium on Parallel and Distributed Computing. Scientific Programming, 19 (1):1, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Soner:2015:GME

[SO15] Seren Soner and Can Ozturan. Generating multibillion element unstructured meshes on distributed memory parallel machines. Scientific Programming, 2015 (??):437480:1–437480:10, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
437480/.

Sati:2018:APC

[SO18] Nur Uylas Sati and Burak Ordin. Application of the polyhedral conic functions method in the text classification and comparative analysis. Scientific Programming, 2018(1):5349284:1– 5349284:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/5349284.

Song:2020:CUO

[Son20] Shixian Song. Construction of university online examination system based on cloud computing technology. Scientific Programming, 2020(1): 7849255:1–7849255:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7849255.

Song:2023:SRB

[Son23a] Huilin Song. A study on the relationship between public derivative big data and industrial policymaking: Taking bike sharing as an example. Scientific Programming, 2023(1):9347432:1– 9347432:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9347432.

Song:2023:UEQ

[Son23b] Longhao Song. University employment quality evaluation system based on multicriteria decision and data analysis. Scientific Programming, 2023(1):3838140:1– 3838140:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3838140.

Song:2020:IRR

[SPH20] Huilin Song, Diyun Peng, and Xin Huang. Incorporating research reports and market sentiment for stock excess return prediction: a case of mainland China. Scientific Programming, 2020(1): 8894757:1–8894757:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

                                        wiley.com/doi/epdf/10.
                                        1155/2020/8894757.

Spotz:2012:PRA

[Spo12] William F. Spotz. PyTrilinos: Recent advances in the Python interface to Trilinos. Scientific Programming, 20(3):311–325, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Santana-Perez:2015:TRS

[SPPH15] Idafen Santana-Perez and María S. Pérez-Hernández. Towards reproducibility in scientific workflows: An infrastructurebased approach. Scientific Programming, 2015 (??):243180:1–243180:11, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://www.hindawi.
                                        com/journals/sp/2015/
                                        243180/.

Sun:2021:LMA

[SQ21] Yingjia Sun and Xin Qi. A DE-LS metaheuristic algorithm for hybrid flow-shop scheduling problem considering multiple requirements of customers. Scientific Programming, 2021(1): 8811391:1–8811391:??, 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8811391.

Shan:2023:PAP

[SQJG23] Shujun Shan, Yue Qi, Jihong Jiang, and Song Guo. Prediction and analysis of protein ubiquitin sites in the model plant A. thaliana. Scientific Programming, 2023(1): 6694846:1–6694846:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6694846.

Shan:2021:PRK

[SQW^{+}21] Xin Shan, Jingyi Qiu, Bo Wang, Yongcheng Dang, Tingxiang Lu, and Yiming Zheng. Place retrieval in knowledge graph. Scientific Programming, 2021(1):5060635:1– 5060635:??, 2021. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/5060635.

Szustak:2015:AMH

[SRO^{+}15] Lukasz Szustak, Krzysztof Rojek, Tomasz Olas, Lukasz Kuczynski, Kamil Halbiniak, and Pawel Gepner. Adaptation of MPDATA heterogeneous stencil computation to Intel Xeon Phi coprocessor. Scientific Programming, 2015(??):642705:1– 642705:14, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
642705/.

Soto:2016:SPC

[SRTCL16] Ricardo Soto, Eduardo Rodriguez-Tello, Stéphane Caro, and Frédéric Lardeux. Scientific programming in computational intelligence. Scientific Programming, 2016(??): 5039793:1–5039793:2, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
5039793/.

Schaffer:2000:DPA

[SS00] Daniel S. Schaffer and Max J. Suárez. Design and performance analysis of a massively parallel atmospheric general circulation model. Scientific Programming, 8(1): 49–57, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=h82chcapth0xynh5tw5w%
26referrer=parent%26backto=

issue%2C6%2C6%3Bjournal%
2C5%2C9%3Blinkingpublicationresults%
2C1%2C1.

Silva:1997:IPD

[SSC97] Luis M. Silva, Joao Gabriel Silva, and Simon Chapple. Implementation and performance of DSMPI. Scientific Programming, 6(2):201–214, Summer 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Stockinger:2002:GDM

[SSM^{+}02] Heinz Stockinger, Asad Samar, Shahzad Muzaffar, et al. Grid Data Mirroring Package (GDMP). Scientific Programming, 10(2):121–133, 2002. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

http://iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=1f99bpyvlg7t46lx8ue3%
26referrer=parent%26backto=
issue%2C4%2C8%3Bjournal%
2C1%2C12%3Blinkingpublicationresults%
2C1%2C1.

Shahriary:2015:CBA

[SSN15] Saeed Reza Shahriary, Mohsen Shahriari, and Rafidah MD Noor. A community-based approach for link prediction in signed social networks. Scientific Programming, 2015(??):602690:1–602690:10, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://www.hindawi.
                                        com/journals/sp/2015/
                                        602690/.

Shaukat:2020:IEO

[SSSJ20] N. Shaukat, S. Shuja, S. K. Srinivasan, and S. Jabeen. Improved efficiency of object code verification using statically abstracted object code. Scientific Programming, 2020(1):6791891:1– 6791891:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/6791891.

Sun:2023:VCS

[SSY^{+}23] Wei-Li Sun, Yong Shen, Yuan Yuan, Xiao-Jing Zhou, and Wei-Peng Li. The value and clinical significance of tumor marker detection in cervical cancer. Scientific Programming, 2023(1):6643782:1– 6643782:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6643782.

Schone:2014:TMM

[STD^{+}14] Robert Schöne, Jan Treibig, Manuel F. Dolz, Carla Guillen, Carmen Navarrete, Michael Knobloch, and Barry Rountree. Tools and methods for measuring and tuning the energy efficiency of HPC systems. Scientific Programming, 22(4):273–283, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Stevenson:1997:RGI

[Ste97] D. E. Stevenson. Review: Great ideas in computer science: a gentle introduction, by Alan W. Biermann. Scientific Programming, 6(4):399–400, Winter 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

S:2023:HAR

[STM^{+}23] Madhusudhan H. S, Satish Kumar T, S. M. F. D. Syed Mustapha, Punit Gupta, and Rajan Prasad Tripathi. Hybrid approach for resource allocation in cloud infrastructure using random forest and genetic algorithm. Scientific Programming, 2023(1): 4924708:1–4924708:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4924708.

Su:2023:LDL

[Su23]  Kaituo Su.  Logistics

distribution location algorithm based on improved imperial competition algorithm. Scientific Programming, 2023(1): 5382868:1–5382868:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5382868.

Sun:2023:DAS

[Sun23a]  Sihua Sun.  Digital

audio scene recognition method based on machine learning technology. Scientific Programming, 2023(1):2388697:1– 2388697:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/2388697.

Sun:2023:DPI

[Sun23b] Xiaolifei Sun. Design of the poster image system based on human vision. Scientific Programming, 2023(1):1411145:1– 1411145:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/1411145.

Sun:2023:ESR

[Sun23c] Zhenzi Sun. An empirical study on the relationship between education and economic development based on PVAR model. Scientific Programming, 2023(1): 6052182:1–6052182:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6052182.

Spafford:2015:ADS

[SV15] Kyle L. Spafford and Jeffrey S. Vetter. Automated design space exploration with Aspen. Scientific Programming, 2015 (??):157305:1–157305:10, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
157305/.

Singh:2007:OWD

[SVR^{+}07] Gurmeet Singh, Karan Vahi, Arun Ramakrishnan, Gaurang Mehta, Ewa Deelman, Henan Zhao, Rizos Sakellariou, Kent Blackburn, Duncan Brown, Stephen Fairhurst, David Meyers, G. Bruce Berriman, John Good, and Daniel S. Katz. Optimizing workflow data footprint. Scientific Programming, 15 (4):249–268, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Sitek:2016:HPF

[SW16] Paweł Sitek and Jarosław Wikarek. A hybrid programming framework for modeling and solving constraint satisfaction and optimization problems. Scientific Programming, 2016(??): 5102616:1–5102616:13, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://www.hindawi.
                                        com/journals/sp/2016/
                                        5102616/.

Song:2020:CCS

[SW20] Chuandong Song and Haifeng Wang. Cir-BiTree: Citrullination site inference based on a fuzzy neural network and flexible neural tree. Scientific Programming, 2020(1):8847694:1– 8847694:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/8847694.

Song:2023:IRD

[SW23] Su Song and Fangzheng Wang. Innovative research on the development of online education mode of internet thinking based on the discrimination of learning attention under the analysis of head posture. Scientific Programming, 2023(1):7183278:1– 7183278:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7183278.

Shu:2024:DLB

[SW24] Yiwen Shu and Xiwen Wu. Deep learning based coronary angiography in diagnosis of myocardial ischemia. Scientific Programming, 2024(1): 8491976:1–8491976:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8491976.

Sun:2024:ETT

[SWDZ24] Jie Sun, Zhiruo Wang, Xiaoyi Dang, and Yang Zhang. Eye-tracking technology in online real estate rental. Scientific Programming, 2024(1): 8851657:1–8851657:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

                                        wiley.com/doi/epdf/10.
                                        1155/2021/8851657.

Sun:2023:LRA

[SWH23] Long Sun, Haitao Wang, and Jian Huang. Logistic regression analysis of relationship between changes of cerebrospinal fluid and communicating hydrocephalus after decompressive craniectomy in craniocerebral injury under computed tomography images. Scientific Programming, 2023(1): 1746412:1–1746412:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/1746412.

Shi:2024:DMK

[SWW^{+}24] Hengyue Shi, Dong Wang, Peng Wu, Yi Cao, and Yuehui Chen. Deep multiple kernel learning for prediction of MicroRNA precursors. Scientific Programming, 2024(1): 9969282:1–9969282:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9969282.

Sun:2023:CID

[SWY23] Wenwen Sun, Dongliang Wang, and Xianliang Yan. CT imaging in the diagnosis of lung injury of organophosphorus poisoning and analysis of its correlation with procalcitonin and c-reactive protein levels. Scientific Programming, 2023(1):9011630:1– 9011630:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9011630.

Su:2024:FCA

[SWZL24] Haoliang Su, Fang Wang, Leying Zhang, and Guiyang Li. Fuzzy clustering algorithm-segmented MRI images in analysis of effects of mental imagery on neurorehabilitation of stroke patients. Scientific Programming, 2024(1): 9945153:1–9945153:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9945153.

Sun:2017:NIP

[SWZS17] Limei Sun, Siqin Wu, Zili Zhu, and Alec Stephenson. Noninterest income and performance of commercial banking in China. Scientific Programming, 2017(1): 4803840:1–4803840:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2017/4803840.

Shang:2023:IST

[SX23] Wenli Shang and Xiangyu Xing. ICS software trust measurement method based on dynamic length trust chain. Scientific Programming, 2023 (1):6691696:1–6691696:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6691696.

Shi:2023:ADL

[SXP23] Wenlei Shi, Lei Xu, and Dongli Peng. Application of deep learning in financial management evaluation. Scientific Programming, 2023(1):2475885:1– 2475885:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/2475885.

Satratzemi:2018:ESP

[SXTK18] Maya Satratzemi, Stelios Xinogalos, Despina Tsompanoudi, and Leonidas Karamitopoulos. Examining student performance and attitudes on distributed pair programming. Scientific Programming, 2018(1):6523538:1– 6523538:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/6523538.

Song:2020:UCS

[SY20] Chuandong Song and Bin Yang. Use chou’s 5-step rule to classify protein modification sites with neural network. Scientific Programming, 2020(1):8894633:1– 8894633:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8894633.

Shen:2023:IAI

[SY23] Yan Shen and Fang Yu.

The influence of artificial intelligence on art design in the digital age. Scientific Programming, 2023(1):4838957:1– 4838957:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4838957.

Shi:2023:ORA

[SYW23] Ying Shi, Yongsheng Yang, and Fengjun Wang. Optimized reconstruction algorithm based analysis of effects of exercise therapy combined with celecoxib capsules on the rehabilitation of knee osteoarthritis using computed tomography images. Scientific Programming, 2023(1):9590266:1– 9590266:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9590266.

Sakellariou:2004:LCR

[SZ04] Rizos Sakellariou and Henan Zhao. A low-cost rescheduling policy for efficient mapping of workflows on Grid systems. Scientific Programming, 12(4):253–262, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Savage:2009:EMA

[SZ09] John E. Savage and Mohammad Zubair. Evaluating multicore algorithms on the unified memory model. Scientific Programming, 17(4):295–308, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Song:2020:PCC

[SZ20] Hui Song and Siyu Zhang.

Perceptual characteristics of Chinese speech intelligibility in noise environment. Scientific Programming, 2020(1):8859152:1– 8859152:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8859152.

Samudrala:2015:PFD

[SZAG15] Sai Kiranmayee Samudrala, Jarosław Zola, Srinivas Aluru, and Baskar Ganapathysubramanian. Parallel framework for dimensionality reduction of large-scale datasets. Scientific Programming, 2015(??):180214:1–180214:12, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
180214/.

Song:2016:SOU

[SZC16] Jiekun Song, Kaixin Zhang, and Zijian Cao. 3Es system optimization under uncertainty using hybrid intelligent algorithm: A fuzzy chance-constrained programming model. Scientific Programming, 2016 (??):2675759:1–2675759:13, ???? 2016. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2016/
2675759/.

Shi:2023:DCD

[SZL^{+}23] Zhigang Shi, Yunlong Zhao, Zhanshuang Liu, Yanan Zhang, and Le Ma. Diagnosis and classification decision analysis of overheating defects of substation equipment based on infrared detection technology. Scientific Programming, 2023(1): 3356044:1–3356044:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3356044.

Sang:2023:NHF

[SZW^{+}23] Xu Sang, Zhen Zhang, Yumeng Wu, Wansheng Peng, and Xin Chen. Noninvasive high-frequency shock ventilation based on chest X-ray reconstruction algorithm for neonatal respiratory distress syndrome. Scientific Programming, 2023(1): 4535136:1–4535136:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4535136.

Shang:2023:TCG

[SZWZ23] Wenli Shang, Guanyu Zhang, Tianyu Wang, and Rui Zhang. A test cases generation method for industrial control protocol test. Scientific Programming, 2023(1):6611732:1– 6611732:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6611732.

Shi:2020:ASS

[SZX^{+}20] Chenjun Shi, Ji Zhu, Mingqian Xu, Xu Wu, and Yan Peng. An approach of spectra standardization and qualitative identification for biomedical materials based on terahertz spectroscopy. Scientific Programming, 2020(1):8841565:1–8841565:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8841565.

Song:2016:MMB

[SZXL16] Aibo Song, Maoxian Zhao, Yingying Xue, and Junzhou Luo. MHDFS: A memory-based Hadoop framework for large data storage. Scientific Programming, 2016(??):1808396:1– 1808396:12, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://www.hindawi.
                                        com/journals/sp/2016/
                                        1808396/.

Szymanski:2007:FPL

[Szy07] Bolesław K. Szymanski. Fortran programming language and scientific programming: 50 years of mutual growth. Scientific Programming, 15(1): 1–2, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Tang:2023:ARM

[Tan23] JinGen Tang. An action recognition method for volleyball players using deep learning. Scientific Programming, 2023(1): 3934443:1–3934443:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/3934443.

Thakur:1996:ETP

[TC96] Rajeev Thakur and Alok Choudhary. An Extended Two-Phase Method for Accessing Sections of Out-of-Core Arrays. Scientific Programming, 5 (4):301–317, Winter 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

http://www.mcs.anl.gov/
~thakur/papers/ext2ph.
ps.

Truong:2004:SGU

[TF04] Hong-Linh Truong and Thomas Fahringer. SCALEA-G: a unified monitoring and performance analysis system for the Grid. Scientific Programming, 12 (4):225–237, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Tian:2023:LMG

[TF23] Yanjia Tian and Xiang Feng. Large margin graph embedding-based discriminant dimensionality reduction. Scientific Programming, 2023(1): 2934362:1–2934362:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2934362.

Turcu:2011:RTD

[TFN11] Gabriela Turcu, Ian Foster, and Svetlozar Nestorov. Reshaping text data for efficient processing on Amazon EC2. Scientific Programming, 19 (2–3):133–145, ???? 2011.

CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Tang:2020:ABR

[TFZ^{+}20] Qiong Tang, Zhuo Fu, Dezhi Zhang, Hao Guo, and Minyi Li. Addressing the bike repositioning problem in bike sharing system: a twostage stochastic programming model. Scientific Programming, 2020(1): 8868892:1–8868892:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8868892.

Thalij:2018:MGS

[TH18] Saadi Hamad Thalij and Veli Hakkoymaz. Multiobjective glowworm swarm optimization-based dynamic replication algorithm for real-time distributed databases. Scientific Programming, 2018(1):2724692:1–2724692:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/2724692.

Tian:2023:SAC

[TH23] Ye Tian and Xiaobing Hu. SWOT analysis of China’s ceramic industry and the use of computers for scientific and technological innovation research. Scientific Programming, 2023(1): 5395988:1–5395988:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5395988.

Takizawa:2015:ODT

[THS^{+}15] Hiroyuki Takizawa, Shoichi Hirasawa, Makoto Sugawara, Isaac Gelado, Hiroaki Kobayashi, and Wen mei W. Hwu. Optimized data transfers based on the OpenCL event management mechanism. Scientific Programming, 2015(??):576498:1– 576498:16, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
576498/.

Tian:2023:OHM

[Tia23] Hua Tian. Optimization of hybrid multimedia art and design teaching mode in the era of big data. Scientific Programming, 2023(1):8266436:1– 8266436:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

                                        wiley.com/doi/epdf/10.
                                        1155/2021/8266436.

Turjo:2023:SSC

[TKKZ23] Manoshi Das Turjo, Mohammad Monirujjaman Khan, Manjit Kaur, and Atef Zaguia. Smart supply chain management using the blockchain and smart contract. Scientific Programming, 2023(1): 6092792:1–6092792:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6092792.

Tao:2002:MAB

[TKS02] Jie Tao, Wolfgang Karl, and Martin Schulz. Memory access behavior analysis of NUMA-based shared memory programs. Scientific Programming, 10(1):45–53, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL http://iospress.
                                        metapress.com/app/home/
                                        contribution.asp%3Fwasp=
                                        9ejnuvwuvby9737jte27%
                                        26referrer=parent%26backto=
                                        issue%2C5%2C9%3Bjournal%
                                        2C2%2C12%3Blinkingpublicationresults%
                                        2C1%2C1.

Tang:2024:CFR

[TKZ^{+}24] Bing Tang, Linyao Kang, Li Zhang, Feiyan Guo, and Haiwu He. Collaborative filtering recommendation using nonnegative matrix factorization in GPU-Accelerated spark platform. Scientific Programming, 2024(1): 8841133:1–8841133:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8841133.

Tian:2017:PSR

[TL17] Hao Tian and Peifeng Liang. Personalized service recommendation based on trust relationship. Scientific Programming, 2017(1):4106134:1– 4106134:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/4106134.

Tran:2015:CLC

[TLC15] Nhat-Phuong Tran, Myungho Lee, and Dong Hoon Choi. Cache localitycentric parallel string matching on many-core accelerator chips. Scientific Programming, 2015 (??):937694:1–937694:20, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.

                                        com/journals/sp/2015/
                                        937694/.

Tran:2017:POL

[TLH17] Nhat-Phuong Tran, Myungho Lee, and Sugwon Hong. Performance optimization of 3D lattice Boltzmann flow solver on a GPU. Scientific Programming, 2017(1):1205892:1– 1205892:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2017/1205892.

Tang:2023:SKE

[TLW^{+}23] Mingjing Tang, Tong Li, Wei Wang, Rui Zhu, Zifei Ma, and Yahui Tang. Software knowledge entity relation extraction with entity-aware and syntactic dependency structure information. Scientific Programming, 2023(1):7466114:1– 7466114:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/7466114.

Tang:2023:DMH

[TLZ^{+}23] Yahui Tang, Tong Li, Rui Zhu, Fei Du, Jishu Wang, and Zifei Ma. A discovery method for hierarchical software execution behavior models based on components. Scientific Programming, 2023(1):4788357:1– 4788357:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4788357.

Torti:2019:DRN

[TMG^{+}19] Emanuele Torti, Mirto Musci, Federico Guareschi, Francesco Leporati, and Marco Piastra. Deep recurrent neural networks for edge monitoring of personal risk and warning situations. Scientific Programming, 2019(1): 9135196:1–9135196:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/9135196.

Teboulbi:2023:RTI

[TMHM23] Safa Teboulbi, Seifeddine Messaoud, Mohamed Ali Hajjaji, and Abdellatif Mtibaa. Realtime implementation of AI-based face mask detection and social distancing measuring system for COVID-19 prevention. Scientific Programming, 2023(1):8340779:1– 8340779:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/8340779.

Tian:2023:DSS

[TMY^{+}23] Yangyang Tian, Wandeng Mao, Shaoguang Yuan, Diming Wan, and Yuanhui Chen. A decision support system for power components based on improved YOLOv4-Tiny. Scientific Programming, 2023(1):4447271:1– 4447271:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/4447271.

Tanaka:2014:ISB

[TOF^{+}14] Teruo Tanaka, Ryo Otsuka, Akihiro Fujii, Takahiro Katagiri, and Toshiyuki Imamura. Implementation of d-spline-based incremental performance parameter estimation method with ppOpen-AT. Scientific Programming, 22 (4):299–307, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Tong:2020:HAF

[Ton20] Wangyu Tong. A hybrid algorithm framework with learning and complementary fusion features for whale optimization algorithm. Scientific Programming, 2020(1): 5684939:1–5684939:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/5684939.

Tang:2013:FLC

[TPKP13] Ping Tak Peter Tang, Jongsoo Park, Daehyun Kim, and Vladimir Petrov. A framework for lowcommunication 1-D FFT. Scientific Programming, 21(3–4):181–195, ???? 2013. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Thant:2017:MLW

[TPSM17] Phyo Thandar Thant, Courtney Powell, Martin Schlueter, and Masaharu Munetomo. Multiobjective level-wise scientific workflow optimization in IaaS public cloud environment. Scientific Programming, 2017(1):5342727:1– 5342727:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/5342727.

Tang:2023:EMP

[TS23a] Yushou Tang and Jianhuan Su. Eye movement prediction based on adaptive BP neural network. Scientific Programming, 2023(1):4977620:1– 4977620:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/4977620.

Song:2023:CCI

[tS23b] Jian tao Song. Construction of corporate investment decision support model based on deep learning. Scientific Programming, 2023(1): 2665041:1–2665041:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/2665041.

Thakar:2011:LSD

[TSCT11] Ani Thakar, Alex Szalay, Ken Church, and Andreas Terzis. Large science databases — are cloud services ready for them? Scientific Programming, 19(2–3):147– 159, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Tian:2015:ESV

[TSP^{+}15] Xinmin Tian, Hideki Saito, Serguei V. Preis, Eric N. Garcia, Sergey S. Kozhukhov, Matt Masten, Aleksei G. Cherkasov, and Nikolay Panchenko. Effective SIMD vectorization for Intel Xeon Phi coprocessors. Scientific Programming, 2015 (??):269764:1–269764:14, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
269764/.

Tang:2016:EER

[TT16] Xiaoyong Tang and Weizhen Tan. Energy-efficient reliability-aware scheduling algorithm on heterogeneous systems. Scientific Programming, 2016 (??):9823213:1–9823213:13, ???? 2016. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2016/
9823213/.

Tian:2023:VRB

[TT23] Shuai Tian and Xuedong Tian. Vehicle reidentification based on MAPANet and k-reciprocal encoding. Scientific Programming, 2023(1):8711288:1– 8711288:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8711288.

Tian:2023:MCA

[TTL^{+}23] Chuangeng Tian, Lu Tang, Xiao Li, Kaili Liu, and Jian Wang. Morphological component analysisbased perceptual medical image fusion using convolutional sparsitymotivated PCNN. Scientific Programming, 2023 (1):6647200:1–6647200:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6647200.

Tunay:2023:NDM

[Tun23] Mustafa Tunay. A new design of metaheuristic search called improved monkey algorithm based on random perturbation for optimization problems. Scientific Programming, 2023(1):5557259:1– 5557259:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5557259.

Tellez-Velazquez:2023:FFF

[TVCB23] Arturo Téllez-Velázquez and Raúl Cruz-Barbosa. On the feasibility of Fast Fourier Transform separability property for distributed image processing. Scientific Programming, 2023(1):1780931:1– 1780931:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1780931.

Xie:2024:LBA

[tXChL^{+}24] Shu tong Xie, Qiong Chen, Kun hong Liu, Qing zhao Kong, and Xiu juan Cao. Learning behavior analysis using clustering and evolutionary error correcting output code algorithms in small private online courses. Scientific Programming, 2024(1): 9977977:1–9977977:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9977977.

Tiankai:2020:RAA

[TXD^{+}20] Sun Tiankai, Wang Xingyuan, Jiang Daihong, Lin Da, Ding Bin, and Li Dan. A robust authentication algorithm for medical images based on fractal Brownian model and visual cryptography. Scientific Programming, 2020(1):6642586:1– 6642586:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/6642586.

Tang:2023:AIB

[TY23] Xiaobin Tang and Jiang Yuan. Artificial intelligencebased computed tomography imaging characteristics for the diagnosis effect of high flow nasal cannula in the treatment of patients with advanced lung cancer and chronic obstructive pulmonary disease. Scientific Programming, 2023(1):5204017:1– 5204017:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5204017.

Tuerxun:2023:DLB

[TYCL23] Mayila Tuerxun, Lixin Yin, Huiqun Chen, and Jingqian Lin. Deep learning-based positron emission tomography molecular imaging in the assessment of cognitive dysfunction in patients with epilepsy. Scientific Programming, 2023(1): 2714222:1–2714222:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2714222.

Tyma:1999:TVC

[Tým99] Paul Týma. Transient variable caching in Java’s stack-based intermediate representation. Scientific Programming, 7(2): 157–166, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=f277qlrwwjr5m4vxjywv%
26referrer=parent%26backto=
issue%2C7%2C8%3Bjournal%
2C7%2C9%3Blinkingpublicationresults%
2C1%2C1.

Tian:2023:SSR

[TZC23] Liang Tian, Xiaorou Zhong, and Ming Chen. Semantic segmentation of remote sensing image based on GAN and FCN network model. Scientific Programming, 2023(1): 9491376:1–9491376:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9491376.

Tan:2023:PBR

[TZCZ23] Yonghong Tan, Xuebin Zhou, Aiwu Chen, and Songqing Zhou. Pedestrian behavior recognition based on improved dual-stream network with differential feature in surveillance video. Scientific Programming, 2023(1):3279957:1– 3279957:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/3279957.

Tang:2020:EFD

[TZDZ20] Qian Tang, Huan Zhang, Jun Dong, and Lianming Zhang. Elephant flow detection mechanism in SDN-based data center networks. Scientific Programming, 2020(1): 8888375:1–8888375:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/8888375.

Tan:2018:KGR

[TZF^{+}18] Zhen Tan, Xiang Zhao, Yang Fang, Bin Ge, and Weidong Xiao. Knowledge graph representation via similarity-based embedding. Scientific Programming, 2018(1):

6325635:1–6325635:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/6325635.

Tan:2020:TRS

[TZH20] Xiaojing Tan, Ming Zou, and Xiqin He. Target recognition in SAR images based on multiresolution representations with 2D canonical correlation analysis. Scientific Programming, 2020(1):7380790:1– 7380790:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/7380790.

Zeng:2019:VIA

[tZhYsHlZ19] Le tian Zeng, Chun hui Yang, Mao sheng Huang, and Yue long Zhao. Verification of imaging algorithm for signal processing software within synthetic aperture radar (SAR) system. Scientific Programming, 2019(1):7105281:1– 7105281:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/7105281.

Tang:2024:CCB

[TZL^{+}24] Jingfan Tang, Meijia Zhou, Pengfei Li, Min Zhang, and Ming Jiang. Crowd counting based on multiresolution density map and parallel dilated convolution. Scientific Programming, 2024(1): 8831458:1–8831458:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/8831458.

Tan:2023:HMM

[TZZ23] Shijie Tan, Hongjun Zhou, and Jinjin Zheng. A hybrid model method for accurate surface deformation and incision based on FEM and PBD. Scientific Programming, 2023(1):8343312:1– 8343312:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/8343312.

Ullah:2018:BDC

[UAK18] Saeed Ullah, M. Daud Awan, and M. Sikander Hayat Khiyal. Big data in cloud computing: a resource management perspective. Scientific Programming, 2018(1): 5418679:1–5418679:??, 2018.

CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/5418679.

Ullah:2023:EIB

[UAU^{+}23] Rafi Ullah, Arbab Waseem Abbas, Mohib Ullah, Rafi Ullah Khan, Irfan Ullah Khan, Nida Aslam, and Sumayh S. Aljameel. EEWMP: an IoT-based energy-efficient water management platform for smart irrigation. Scientific Programming, 2023(1):5536884:1– 5536884:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5536884.

Ullah:2020:MDT

[UJS^{+}20] Faizan Ullah, Qaisar Javaid, Abdu Salam, Masood Ahmad, Nadeem Sarwar, Dilawar Shah, and Muhammad Abrar. Modified decision tree technique for ransomware detection at runtime through API calls. Scientific Programming, 2020 (1):8845833:1–8845833:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.

                                        wiley.com/doi/epdf/10.
                                        1155/2020/8845833.

Vermeulen:1993:OSI

[VC93] Al Vermeulen and Margaret Chapman. OON-SKI: an introduction. Scientific Programming, 2 (4):109–110, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Varela:2005:WCA

[VCT05] Carlos A. Varela, Paolo Ciancarini, and Kenjiro Taura. Worldwide computing: Adaptive middleware and programming technology for dynamic Grid environments. Scientific Programming, 13 (4):255–263, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

vanDam-Kleese:2001:ETI

[vDKH01] Kerstin van Dam-Kleese and Michael Hopewell. Enabling technologies for improved data management: Hardware. Scientific Programming, 9(1): 11–25, 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

                                        //iospress.metapress.
                                        com/app/home/contribution.
                                        asp%3Fwasp=f2779jvvqg63jq64qwtm%
                                        26referrer=parent%26backto=
                                        issue%2C2%2C6%3Bjournal%

2C2%2C9%3Blinkingpublicationresults%
2C1%2C1.

Viklund:1995:OOO

[VF95] Lars Viklund and Peter Fritzson. ObjectMath

— an object-oriented language and environment for symbolic and numerical processing in scientific computing. Scientific Programming, 4(4):229– 250, Winter 1995. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Vishwas:2009:IPM

[VGC09] B. C. Vishwas, Abhishek Gadia, and Mainak Chaudhuri. Implementing a parallel matrix factorization library on the Cell Broadband Engine. Scientific Programming, 17(1–2):3– 29, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Verner:1993:DGF

[VHBR93] Diane A. Verner, Gregory L. Heileman, Kent G. Budge, and Allen C. Robinson. Development of generic field classes for finite element and finite difference problems. Scientific Programming, 2 (4):227–234, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Velu:2016:ETB

[VMRM16] Vengadeshwaran Velu, Norman Mariun, Mohd Amran Mohd Radzi, and Nashiren Farzilah Mailah. Equalization technique for balancing the modulation ratio characteristics of the singlephase-to-three-phase matrix converter. Scientific Programming, 2016(??): 6187926:1–6187926:10, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
6187926/.

Vollebregt:1997:ALP

[Vol97] Edwin Vollebregt. Abstract level parallelization of finite difference methods. Scientific Programming, 6(4):331–344, Winter 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=a5tkxhqqy9eefb7hwkak%
26referrer=parent%26backto=
issue%2C1%2C7%3Bjournal%
2C9%2C9%3Blinkingpublicationresults%
2C1%2C1.

Vetter:2013:SIS

[VR13] Jeffrey S. Vetter and Padma Raghavan. Special issue: Selected papers from Super Computing 2012. Scientific Programming, 21(3–4):63–64, ???? 2013. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Visser:2002:FHM

[VRM02] Ingmar Visser, Maartje E. J. Raijmakers, and Peter C. M. Molenaar. Fitting hidden Markov models to psychological data. Scientific Programming, 10(3):185–199, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Verstak:2003:BBS

[VRW^{+}03] Alex Verstak, Naren Ramakrishnan, Layne T. Watson, et al. BSML: a binding schema markup language for data interchange in problem solving environments. Scientific Programming, 11(3): 199–224, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Varbanescu:2009:BHR

[VvAC^{+}09] Ana Lucia Varbanescu, Alexander S. van Amesfoort, Tim Cornwell, Ger van Diepen, Rob van Nieuwpoort, Bruce G. Elmegreen, and Henk Sips. Building highresolution sky images using the Cell/B.E. Scientific Programming, 17(1– 2):113–134, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Vaderna:2018:GDA

[VVDM18] Renata Vaderna, Zeljko Vuković, Igor Dejanović, and Gordana Milosavljević. Graph drawing and analysis library and its domainspecific language for graphs’ layout specifications. Scientific Programming, 2018(1):7264060:1– 7264060:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/7264060.

Vo-Van:2020:NCA

[VVNHTHNT20] T. Vo-Van, A. Nguyen-Hai, M. V. Tat-Hong, and T. Nguyen-Trang. A new clustering algorithm and its application in assessing the quality of underground water. Scientific Programming, 2020(1):6458576:1– 6458576:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/6458576.

Wang:2002:OPG

[Wan02] Ping Wang. OpenMP programming for a global inverse model. Scientific Programming, 10(3): 253–261, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wang:2021:DDO

[Wan21]  Liangliang Wang.  A

data-driven optimization model of important multidimensional factors affecting college students’ cognitive engagement in ideological and political theory course. Scientific Programming, 2021(1): 2257415:1–2257415:??, 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2257415.

Wang:2022:RSF

[Wan22] Ruijun Wang. RETRAC-TION: Spring festival holiday tourism data mining based on the deep learning model. Scientific Programming, 2022(1): 9819746:1–9819746:??, ???? 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2024/9819746.

Wang:2023:MER

[Wan23a] Jian Wang. Music education to rescue psychological stress in social crisis based on fuzzy prediction algorithm. Scientific Programming, 2023(1): 2039235:1–2039235:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/2039235.

Wang:2023:IFE

[Wan23b] Junhuan Wang. Improved facial expression recognition method based on GAN. Scientific Programming, 2023(1):2689029:1– 2689029:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/2689029.

Wang:2023:MMS

[Wan23c] Junshu Wang. A management model of small- and medium-sized enterprises based on deep learning algorithm. Scientific Programming, 2023(1): 5996597:1–5996597:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5996597.

Wang:2023:IPE

[Wan23d] Nan Wang. Ideological and political education recommendation system based on AHP and improved collaborative filtering algorithm. Scientific Programming, 2023 (1):2648352:1–2648352:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2648352.

Wang:2023:RST

[Wan23e] Peng Wang. Research on sports training action recognition based on deep learning. Scientific Programming, 2023(1): 3396878:1–3396878:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3396878.

Wang:2023:DLBa

[Wan23f] Shu Wang. Deep learningbased assessment of adverse cardiovascular events in elderly patients with coronary heart disease after percutaneous coronary intervention using intravascular ultrasound images. Scientific Programming, 2023(1): 3314457:1–3314457:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

                                        wiley.com/doi/epdf/10.
                                        1155/2021/3314457.

Wang:2023:AIA

[Wan23g]  Shuaiqi Wang.  Ar-

tificial intelligence applications in the new model of logistics development based on wireless communication technology. Scientific Programming, 2023(1):5166993:1– 5166993:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5166993.

Wang:2023:GAB

[Wan23h] Shuxin Wang. Genetic algorithm and BP neural network for college physical education teaching evaluation. Scientific Programming, 2023(1): 9921960:1–9921960:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9921960.

Wang:2023:RSI

[Wan23i] Yiqin Wang. Remote sensing image semantic segmentation algorithm based on improved ENet network. Scientific Programming, 2023(1): 5078731:1–5078731:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5078731.

Warren:1996:PDP

[War96] Karen H. Warren. PDDP, a data parallel programming model. Scientific Programming, 5(4):319– 327, Winter 1996. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Warren:2014:IPH

[War14] Michael S. Warren. 2HOT:

an improved parallel hashed oct-tree N-body algorithm for cosmological simulation. Scientific Programming, 22(2):109– 124, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wasserman:1995:BTN

[Was95] Harvey J. Wasserman.

Benchmark tests on the new IBM RISC System/ 6000 590 workstation. Scientific Programming, 4 (1):23–24, Spring 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Williams:1995:SCF

[WB95] Dan Williams and Luc Bauwens. Simulation of compressible flow on a massively parallel architecture. Scientific Programming, 4(3):193– 201, Fall 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wang:2016:GOS

[WB16] Chun-Feng Wang and Yan-Qin Bai. Global optimization for solving linear multiplicative programming based on a new linearization method. Scientific Programming, 2016(??): 3204368:1–3204368:9, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
3204368/.

Wang:2017:RBS

[WBZL17] Xiaokun Wang, Xiaojuan Ban, Yalan Zhang, and Xu Liu. Rigid body sampling and individual time stepping for rigid-fluid coupling of fluid simulation. Scientific Programming, 2017(1):8502691:1– 8502691:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/8502691.

Wolski:1996:CEC

[WC96] Rich Wolski and David Cann. Compiler-enforced cache coherence using a functional language. Scientific Programming, 5(2):161–171, Summer 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wang:2023:RVD

[WC23a] Haifeng Wang and Joung Hyung Cho. Research on virtual display of wetsuit based on CLO3D. Scientific Programming, 2023(1): 1052510:1–1052510:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1052510.

Wu:2023:RME

[WC23b] Yongli Wu and Jingliang Chen. Realization of mobile education resource sharing method based on wireless broadband connection. Scientific Programming, 2023(1): 4018500:1–4018500:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4018500.

Wang:2018:SPH

[WCF18] Junchang Wang, Shaojin Cheng, and Xiong Fu. SDN programming for heterogeneous switches with flow table pipelining. Scientific Programming, 2018(1):2848232:1– 2848232:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/2848232.

Wang:2021:VSA

[WCF^{+}21] Yifei Wang, Ying Chen, Jincai Fang, Yunnong Song, Jianguo Shen, and Jianchao Wang. Values of sinogram affirmed iterative reconstruction algorithm-based low-dose computed tomography imaging in clinical diagnosis of cerebral hemorrhage. Scientific Programming, 2021(1):2228062:1– 2228062:??, 2021. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2228062.

Wolters:1995:DPN

[WCG95] Lex Wolters, Gerard Cats, and Nils Gustafsson. Data-parallel numerical weather forecasting. Scientific Programming, 4(3):141–??, Fall 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wang:2007:EPP

[WCKD07] Joseph Wang, Yong Cao, Raed Kafafy, and Viktor Decyk. Electric propulsion plume simulations using parallel computer. Scientific Programming, 15(2):83–94, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wang:2017:FJS

[WCLL17] Lei Wang, Jingcao Cai, Ming Li, and Zhihu Liu. Flexible job shop scheduling problem using an improved ant colony optimization. Scientific Programming, 2017(1): 9016303:1–9016303:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/9016303.

Wang:2021:IDR

[WCR^{+}21] Xin Wang, Yuanyi Chen, Wei Ruan, Qiang Gao, Guode Ying, and Li Dong. Intelligent detection and recovery of missing electric load data based on cascaded convolutional autoencoders. Scientific Programming, 2021(1): 8828745:1–8828745:??, 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2020/8828745.

Wang:2021:MLA

[WCW^{+}21] Nana Wang, Lingyang Chen, Hongzhu Wang, Yibing Wang, and Binhao Ruan. MATrix LABoratory algorithm in ultrasound image-guided general drug anesthesia along with lumbar and sacral plexus block in hip replacement under. Scientific Programming, 2021(1):2273244:1– 2273244:??, 2021. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2273244.

Wang:2016:IEH

[WCZ^{+}16] Wenfeng Wang, Xi Chen, Hongwei Zheng, Zhihan Lv, Zhengjia Liu, Jing Qian, and Ping Hu. Intelligence in ecology: How Internet of Things expands insights into the missing CO_{2} sink. Scientific Programming, 2016(??): 4589723:1–4589723:8, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
4589723/.

Wang:2023:MCA

[WCZW23] Luyi Wang, Ying Chen, Jian Zhang, and Congxiao Wang. Monte Carlo algorithm-based multimodal magnetic resonance imaging prognosis prediction in analysis of rehabilitation effect of exercise learning on stroke patients and influencing factors of memory function. Scientific Programming, 2023(1):9289494:1– 9289494:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9289494.

Walker:2007:DCW

[WD07] David Walker and Ewa Deelman. Dynamic computational workflows: Discovery, optimisation and scheduling. Scientific Programming, 15(4):191– 192, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Waqar:2023:ESB

[WDD^{+}23] Muhammad Waqar, Hassan Dawood, Hussain Dawood, Nadeem Majeed, Ameen Banjar, and Riad Alharbey. An efficient SMOTE-based deep learning model for heart attack prediction. Scientific Programming, 2023(1):6621622:1– 6621622:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6621622.

Wu:2023:EAR

[WDY23] Shijun Wu, Jianghong Dai, and Jiujiu Yang. An experimental and algorithm research on the influence of OTO teaching mode on college students’ PE learning interest based on cloud computing. Scientific Programming, 2023(1):2042158:1– 2042158:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2042158.

Wang:2020:MTR

[WDZY20] Yingying Wang, Jixiang Du, Hongbo Zhang, and Xiuhong Yang. Mushroom toxicity recognition based on multigrained cascade forest. Scientific Programming, 2020(1): 8849011:1–8849011:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8849011.

Wei:2023:REE

[Wei23] Chunyan Wei. Research on the effect of English talents gathering based on big data hotspot collection technology. Scientific Programming, 2023(1): 2846621:1–2846621:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2846621.

Wen:2023:DDC

[Wen23] Lin Wen. Determining the degree of characteristics for internet of healthcare devices using fuzzy ANP. Scientific Programming, 2023(1):9292496:1– 9292496:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9292496.

Wester:2008:SR

[Wes08] Michael Wester. Software review. Scientific Programming, 16(1):93– 96, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wang:2020:IBT

[WFHC20] Cheng Wang, Weihua Fu, Haiyang Huang, and Jianwei Chen. Isomapbased three-dimensional operational modal analysis. Scientific Programming, 2020(1):6348372:1– 6348372:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/6348372.

Wu:2023:RAL

[WFZ^{+}23] Haoqiang Wu, Yiran Fu, Quanxing Zha, Aidong Chen, and Hongyuan Jing. Regional atmospheric light optimisation algorithm for heterogeneous image dehazing. Scientific Programming, 2023(1):3377905:1– 3377905:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3377905.

Wu:2021:CAD

[WG21] Donglei Wu and Minwei Guo. Corrigendum to “Application of data mining in traditional benchmark evaluation model for buildings energy consumption”. Scientific Programming, 2021(1):9786563:1– 9786563:??, ???? 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2024/9786563.

Wu:2022:EGE

[WG22] Erhan Wu and Zhiyi Gai. Evaluation of government ecological environment governance effect from the public’s perspective based on the entropy method: Take the Kubuqi Desert in Inner Mongolia as an example. Scientific Programming, 2022(1):6319358:1– 6319358:??, 2022. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6319358.

Wang:2023:SMI

[WG23] Jin Wang and Yanfei Gao.

Suspect multifocus image fusion based on sparse denoising autoencoder neural network for police multimodal big data analysis. Scientific Programming, 2023(1):6614873:1– 6614873:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6614873.

Wu:2024:ADM

[WG24] Donglei Wu and Weimin Guo. Application of data mining in traditional benchmark evaluation model for buildings energy consumption. Scientific Programming, 2024(1):8610050:1– 8610050:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/8610050.

Weissman:1993:POO

[WGF93] Jon B. Weissman, Andrew S. Grimshaw, and R. D. Ferraro. Parallel object-oriented computation applied to a finite element problem. Scientific Programming, 2 (4):133–144, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wylie:2008:PMA

[WGW08] Brian J. N. Wylie, Markus Geimer, and Felix Wolf. Performance measurement and analysis of large-scale parallel applications on leadership computing systems. Scientific Programming, 16 (2–3):167–181, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wang:2020:TTA

[WH20] You Jun Wang and Guo Huang. Target tracking algorithm of basketball video based on improved grey neural network. Scientific Programming, 2020(1):7808456:1– 7808456:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7808456.

Wang:2023:ACC

[WH23] Huafeng Wang and Rong Huang. Application of cloud computing in the optimization of college calisthenics teaching mode. Scientific Programming, 2023(1):9265238:1– 9265238:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9265238.

Willis:1993:ADA

[WHG93] A. G. Willis, M. P.

Healey, and B. E. Glendenning. AIPS++ ndimensional array classes. Scientific Programming, 2 (4):239–246, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wang:2021:STI

[WHH^{+}21] Zhijin Wang, Yaohui Huang, Bingyan He, Ting Luo, Yongming Wang, and Yonggang Fu. Shortterm infectious diarrhea prediction using weather and search data in Xiamen, China. Scientific Programming, 2021(1): 8814222:1–8814222:??, 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8814222.

Wang:2018:IMI

[WHL^{+}18] Wei Wang, Yun He, Tong Li, Jiajun Zhu, and Jinzhuo Liu. An integrated model for information retrieval based change impact analysis. Scientific Programming, 2018(1):5913634:1– 5913634:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/5913634.

Wang:2021:NEM

[WHL^{+}21] Weili Wang, Mingwei Huang, Tingting Lin, Chengzhi Lu, and Jiandong Liu. Neuroprotective effect of monosialotetrahexosylganglioside (GM1) on patients with Parkinson’s disease anesthetized by ketamine under denoising algorithm-based ultrasound image diagnosis. Scientific Programming, 2021(1):2253431:1– 2253431:??, 2021. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2253431.

Walker:2007:DSS

[WHRH07] David W. Walker, Lican Huang, Omer F. Rana, and Yan Huang. Dynamic service selection in workflows using performance data. Scientific Programming, 15(4):235–247, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wang:2021:UBD

[WJCZ21] Yingzi Wang, Muhammad Nazir Jan, Sisi Chu, and Yue Zhu. Use of big data tools and industrial Internet of Things: an overview. Scientific Programming, 2021(1): 8810634:1–8810634:??, 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8810634.

Wei:2023:TMC

[WJS^{+}23] Hong Wei, Fang Jiang, Fang Shao, Denghui Zhang, Fang Gu, Ying Yang, Qiuxia Chen, and Zheng Ai. Temperature measurement and control application in a laser plastic surgery real temperature detection system. Scientific Programming, 2023(1):8382482:1– 8382482:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8382482.

Wu:2023:CTI

[WJZC23] Lijuan Wu, Jianwei Ji, Shiyong Zhao, and Jiaolei Chen. Computed tomography image segmentation using edge correction algorithm for refractory mycoplasma pneumonia in children. Scientific Programming, 2023(1): 3578971:1–3578971:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3578971.

Wang:2023:DBC

[WKW^{+}23] Shujuan Wang, Yan Kang, Xiaoqing Wang, Lanling Wei, Yanan Zhu, and Huilin Yuan. Design of breakdown and checklist for continuous renal replacement therapy. Scientific Programming, 2023(1):3485058:1– 3485058:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3485058.

Wang:2023:AES

[WKY23] Haotian Wang, Chen Ke, and Xiaojun Yang. Algorithm of ecocompensation in sloping land conversion program based on heckman’s twostep model. Scientific Programming, 2023(1): 7880461:1–7880461:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7880461.

Wang:2023:PMC

[WL23a] Bing Wang and Sitong Liu. Prediction method of college students’ psychological pressure based on deep neural network. Scientific Programming, 2023(1):2943678:1– 2943678:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2943678.

Wang:2023:CDR

[WL23b] Jing Wang and Wei Li. The construction of a digital resource library of English for higher education based on a cloud platform. Scientific Programming, 2023(1):4591780:1– 4591780:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/4591780.

Wang:2023:MIT

[WL23c] Pengyuan Wang and Jie Li. Medical image three-dimensional simulation technology in hospital clinical practice. Scientific Programming, 2023(1):1169676:1– 1169676:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/1169676.

Wang:2018:LLB

[WLCC18] Hai Wang, Xinyu Lou, Yingfeng Cai, and Long Chen. A 64-line lidarbased road obstacle sensing algorithm for intelligent vehicles. Scientific Programming, 2018(1): 6385104:1–6385104:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/6385104.

Wang:2022:REE

[WLJ22] Lingtuo Wang, Xiaobin Li, and Yufang Jiao. RETRACTION: Effect of English reading on the cultivation of College students’ humanistic spirit based on embedded sensor system. Scientific Programming, 2022(1):9852168:1– 9852168:??, ???? 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2024/9852168.

Wang:2023:NEE

[WLL23a] Xuezhi Wang, Wenhui Li, and Qingliang Li. A new embedded estimation model for soil temperature prediction. Scientific Programming, 2023(1):5881018:1– 5881018:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5881018.

Wang:2023:SIP

[WLL^{+}23b] Zheng Wang, Yang Li, Li-Ping Li, Zhu-Hong You, and Wen-Zhun Huang.

Self-interacting proteins prediction from PSSM based on evolutionary information. Scientific Programming, 2023(1): 6677758:1–6677758:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6677758.

Wei:2023:CAE

[WLLY23] Meiyu Wei, Ming Liu, Jie Liu, and Haitao Yang. A comparison of analgesic effect between preoperative and postoperative transversus abdominis plane (TAP) blocks for different durations of laparoscopic gynecological surgery. Scientific Programming, 2023(1): 6668496:1–6668496:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6668496.

Wang:2021:TSS

[WLP^{+}21] Shudong Wang, Yanqing Li, Shanchen Pang, Qinghua Lu, Shuyu Wang, and Jianli Zhao. A task scheduling strategy in edge-cloud collaborative scenario based on deadline. Scientific Programming, 2021(1):3967847:1– 3967847:??, 2021. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/3967847.

Wang:2023:RMF

[WLQ23] Lei Wang, Qian Li, and Jin Qin. Rotating machinery fault diagnosis method based on improved semisupervised generative confrontation network. Scientific Programming, 2023(1): 1761446:1–1761446:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1761446.

Wu:2021:BDM

[WLZ^{+}21] Hanqian Wu, Mumu Liu, Shangbin Zhang, Zhike Wang, and Siliang Cheng. Big data management and analytics in scientific programming: a deep learningbased method for aspect category classification of question-answering-style reviews. Scientific Programming, 2021(1): 4690974:1–4690974:??, 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/4690974.

Wang:2024:TDR

[WLZJ24] Lijun Wang, Haiyan Lin, Hong Zheng, and Yuying Jiang. Threedimensional reconstruction algorithm of computed tomography imaging for surgical treatment and rapid rehabilitation nursing of renal cell carcinoma. Scientific Programming, 2024(1):9948656:1– 9948656:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9948656.

Wang:2023:LDA

[WM23a] Chuxin Wang and Haoran Mo. Learning deep attention network from incremental and decremental features for evolving features. Scientific Programming, 2023(1):1492828:1– 1492828:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/1492828.

Wang:2023:RGR

[WM23b] Xiaowen Wang and Yijun Mu. Research on grey relational clustering model of multiobjective human resources based on time constraint. Scientific Programming, 2023(1): 5551255:1–5551255:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5551255.

Wu:2021:FPC

[WmG21] Hui Wu and Xiao min Gu. Fuzzy principal component analysis model on evaluating innovation service capability. Scientific Programming, 2021(1):8834901:1– 8834901:??, 2021. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8834901.

Wheat:1994:POS

[WMR^{+}94] Stephen R. Wheat, Arthur B.

Maccabe, Rolf Riesen, David W. van Dresser, and T. Mack Stallcup. PUMA: an operating system for massively parallel systems. Scientific Programming, 3(4):275–288, Winter 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wu:2023:SAP

[WMZM23] Zongtao Wu, Jingku Ma, Xiaolong Zhou, and Xiumei Ma. Selfadaptive particle swarm optimization of CT images for diagnosis of severe traumatic brain injury. Scientific Programming, 2023(1):9891663:1– 9891663:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9891663.

Wang:2020:BSM

[WNL^{+}20] Yingqiang Wang, Zhaohua Nian, Chang Liu, Wei Han, and Maowei Lin. A bandwidth statistical multiplexing and control method for satellite broadcasting. Scientific Programming, 2020(1): 8841006:1–8841006:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8841006.

Welsh:1999:UNS

[WOC99] Matt Welsh, David Oppenheimer, and David Culler. U-Net/SLE: a Java-based user-customizable virtual network interface. Scientific Programming, 7(2):147–156, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.

asp%3Fwasp=f277qlrwwjr5m4vxjywv%
26referrer=parent%26backto=
issue%2C6%2C8%3Bjournal%
2C7%2C9%3Blinkingpublicationresults%
2C1%2C1.

Wang:2019:RPS

[WPDZ19] Kuaini Wang, Huimin Pei, Xiaoshuai Ding, and Ping Zhong. Robust proximal support vector regression based on maximum correntropy criterion. Scientific Programming, 2019(1):7102946:1– 7102946:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/7102946.

Wang:2017:NOR

[WQLR17] Xuping Wang, Jian Qiu, Tong Li, and Junhu Ruan. A network optimization research for product returns using modified plant growth simulation algorithm. Scientific Programming, 2017 (1):1080468:1–1080468:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/1080468.

Wang:2020:MHV

[WQT^{+}20] Bochen Wang, Qiyuan Qian, Zheyi Tan, Peng Zhang, Aizhi Wu, and Yi Zhou. Multidepot heterogeneous vehicle routing problem for a variety of hazardous materials with risk analysis. Scientific Programming, 2020(1):8839628:1– 8839628:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8839628.

Wang:2015:HMS

[WS15] Jie-Sheng Wang and Na-Na Shen. Hybrid multiple soft-sensor models of grinding granularity based on cuckoo searching algorithm and hysteresis switching strategy. Scientific Programming, 2015(??):146410:1– 146410:11, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
146410/.

Wilcox:2011:TSS

[WSB11] Chris Wilcox, Michelle Mills Strout, and James M. Bieman. Tool support for software lookup table optimization. Scientific Programming, 19(4):213–229, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wen:2023:RGA

[WSL23] Yufang Wen, Dongfang Su, and Qing Lin. Regiongrowing algorithm on CT angiography images for detection of gynecological malignant tumor. Scientific Programming, 2023(1):9875886:1– 9875886:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9875886.

Wang:2021:NER

[WSM^{+}21] Yu Wang, Yining Sun, Zuchang Ma, Lisheng Gao, and Yang Xu. Named entity recognition in Chinese medical literature using pretraining models. Scientific Programming, 2021(1): 8812754:1–8812754:??, 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8812754.

Wu:1994:OSS

[WSP94] Min-You Wu and Wolfgang Schroder-Preikschat. Operating system support for massively parallel computer architectures: an introduction. Scientific Programming, 3 (4):273–??, Winter 1994.

CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wirawan:2009:HPP

[WSZK09] Adrianto Wirawan, Bertil Schmidt, Huiliang Zhang, and Chee Keong Kwoh. High performance protein sequence database scanning on the Cell Broadband Engine. Scientific Programming, 17(1– 2):97–111, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wang:2023:MDU

[WTY23] Rong Wang, Cong Tian, and Lin Yan. Malware detection using CNN via word embedding in cloud computing infrastructure. Scientific Programming, 2023(1):8381550:1– 8381550:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/8381550.

Wu:2016:DSR

[Wu16] Xiuguo Wu. Data sets replicas placements strategy from cost-effective view in the cloud. Scientific Programming, 2016 (??):1496714:1–1496714:13, ???? 2016. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2016/
1496714/.

Wu:2023:SIA

[Wu23a] Boyin Wu. Sports intelligent assistance system based on deep learning. Scientific Programming, 2023(1):3481469:1– 3481469:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3481469.

Wu:2023:NSA

[Wu23b] Huajun Wu. Numerical simulation algorithm design of influence on existing tunnel by underpass construction of new tunnel. Scientific Programming, 2023(1):1734308:1– 1734308:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1734308.

Wu:2023:AIB

[Wu23c] Yue Wu. Application of improved boosting algorithm for art image classification. Scientific Programming, 2023(1): 3480414:1–3480414:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/3480414.

Wang:2022:RMM

[WW22] Yifei Wang and Yongsheng Wang. Research on multiperson motion capture system combining target positioning and inertial attitude sensing technology. Scientific Programming, 2022(1):6808521:1– 6808521:??, 2022. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6808521.

Wang:2023:OEO

[WW23a] Huiling Wang and Jiasheng Wang. Online education optimization based on edge computing under the COVID-19 pandemic. Scientific Programming, 2023(1): 3667252:1–3667252:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/3667252.

Wang:2023:AAR

[WW23b] Qian Wang and Mingzhe Wang. Aerobics action recognition algorithm based on threedimensional convolutional neural network and multilabel classification. Scientific Programming, 2023 (1):3058141:1–3058141:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3058141.

Wang:2023:AML

[WW23c] Zhihui Wang and Jinyu Wang. Applications of machine learning in public security information and resource management. Scientific Programming, 2023(1):4734187:1– 4734187:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4734187.

Wang:2023:DLBc

[WWH^{+}23] Han Wang, Hui Wang, Zhonglve Huang, Huajun Su, Xiang Gao, and Feifei Huang. Deep learning-based computed tomography images for quantitative measurement of the correlation between epicardial adipose tissue volume and coronary heart disease. Scientific Programming, 2023(1):9866114:1– 9866114:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9866114.

Wang:2023:SPH

[WWL^{+}23] Hui Wang, Ning Wang, MeiJie Li, Simeng Mi, and YaYa Shi. Student physical health information management model under big data environment. Scientific Programming, 2023(1):5795884:1– 5795884:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5795884.

Wang:2016:HNT

[WWLG16] Guan Wang, Yuxin Wang, Hui Liu, and He Guo. HSIP: A novel task scheduling algorithm for heterogeneous computing. Scientific Programming, 2016(??):3676149:1– 3676149:11, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://www.hindawi.
                                        com/journals/sp/2016/
                                        3676149/.

Wang:2023:DNN

[WWQ23] Haifeng Wang, Chengche Wang, and Hongchun Qu. Deep neural network for somatic mutation classification. Scientific Programming, 2023(1): 5529202:1–5529202:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5529202.

Wu:2017:TCD

[WWW^{+}17] Jibing Wu, Zhifei Wang, Yahui Wu, Lihua Liu, Su Deng, and Hongbin Huang. A tensor CP decomposition method for clustering heterogeneous information networks via stochastic gradient descent algorithms. Scientific Programming, 2017(1): 2803091:1–2803091:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/2803091.

Wang:2023:APR

[WWW23] Zhuo Wang, Zixuan Wang, and Likai Wang. Automatic 3D pollen recognition based on convolutional neural network. Scientific Programming, 2023(1):5577307:1– 5577307:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

                                        wiley.com/doi/epdf/10.
                                        1155/2021/5577307.

Wu:2022:RCD

[WWYR22] Ke Wu, Jiwei Wang, Weisong Yang, and Xuebing Ruan. RETRAC-TION: Cultural differences of basketball between China and the United States and its concept in basketball teaching under the background of wireless network intelligence technology. Scientific Programming, 2022(1):9876879:1– 9876879:??, ???? 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2024/9876879.

Wang:2017:KQE

[WWZ17] Yingqi Wang, Nianbin Wang, and Lianke Zhou. Keyword query expansion paradigm based on recommendation and interpretation in relational databases. Scientific Programming, 2017(1): 7613026:1–7613026:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2017/7613026.

Wan:2023:IRM

[WX23] Jixin Wan and Yu Xiaobo. Intelligent retrieval method of approximate painting in digital art field. Scientific Programming, 2023(1):5796600:1– 5796600:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5796600.

Wang:2019:BIM

[WXL19] Guo-Zheng Wang, Li Xiong, and Hu-Chen Liu. A Bayesian inference method using Monte Carlo sampling for estimating the number of communities in bipartite networks. Scientific Programming, 2019 (1):9471201:1–9471201:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/9471201.

Wang:2017:HRS

[WXZW17] Xiaoding Wang, Li Xu, Shuming Zhou, and Wei Wu. Hybrid recovery strategy based on random terrain in wireless sensor networks. Scientific Programming, 2017(1): 5807289:1–5807289:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

                                        wiley.com/doi/epdf/10.
                                        1155/2017/5807289.

Wang:2023:APB

[WY23] Anzhi Wang and Xiuling Yi. Attitude perception of badminton players based on mobile edge computing. Scientific Programming, 2023(1):2436635:1– 2436635:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/2436635.

Wang:2020:ASO

[WYC^{+}20a] Xin Wang, Yue Yang, Mingsong Chen, Qin Wang, Qin Qin, Hua Jiang, and Huijiao Wang. AGNES-SMOTE: an oversampling algorithm based on hierarchical clustering and improved SMOTE. Scientific Programming, 2020(1): 8837357:1–8837357:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/8837357.

Wu:2020:UNR

[WYC20b] Di Wu, Fei Yuan, and En Cheng. Underwater no-reference image quality assessment for display module of ROV. Scientific Programming, 2020(1):8856640:1– 8856640:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8856640.

Wang:2023:DLBd

[WYHS23] Chao Wang, Jiabin Yuan, Zebin Huang, and Zhicai Shi. Deep learning-based correlation analysis between spine surgery lumbar facet joint and lumbar disc herniation using magnetic resonance images. Scientific Programming, 2023(1):9623991:1– 9623991:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9623991.

Wang:2021:LBA

[WYY21] Bingming Wang, Shi Ying, and Zhe Yang. A log-based anomaly detection method with efficient neighbor searching and automatic k neighbor selection. Scientific Programming, 2021(1): 4365356:1–4365356:??, 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

                                        wiley.com/doi/epdf/10.
                                        1155/2020/4365356.

Wang:2022:MMD

[WZ22] Yingyan Wang and Rui Zeng. The model of makerspace development element and performance analysis based on NVivo classification. Scientific Programming, 2022(1): 7123961:1–7123961:??, 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/7123961.

Wang:2023:DLBb

[WZ23a] Qiaoli Wang and Jinfu Zhu. Deep learning-based CT imaging in perioperative period and nursing of esophageal carcinoma patients. Scientific Programming, 2023(1): 4453317:1–4453317:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/4453317.

Wang:2023:OCT

[WZ23b] Yongqiang Wang and Na Zhang. The optimization of classroom teaching in colleges and universities based on network topology. Scientific Programming, 2023(1): 1271438:1–1271438:??, 2023.

CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1271438.

Wei:2023:IDB

[WZ23c] Hui Wei and Wei Zheng.

Image denoising based on improved Gaussian mixture model. Scientific Programming, 2023(1): 7982645:1–7982645:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7982645.

Wu:2023:ELM

[WZ23d] Jing Wu and Zhikun Zhang. Extreme learning machine denoising algorithm based analysis of transvaginal 3dimensional ultrasonic image for the diagnostic effect of intrauterine adhesion. Scientific Programming, 2023(1): 9629884:1–9629884:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9629884.

Wei:2023:MVA

[WZC^{+}23] Yong Wei, Hongru Zhu, Peng Chen, Wenren Zuo, Wenhui Qian, and Qingyi Zhu. Minimum variance algorithm-based correlation analysis between body mass index and the malignant degree of prostate cancer mediated under ultrasound images. Scientific Programming, 2023(1):4990942:1– 4990942:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4990942.

Wu:2020:UTS

[WZF^{+}20] Yongrong Wu, Yijie Zhou, Yanming Feng, Yutian Xiao, Shaojie He, Junsheng Zhou, Tianhe Ren, Jinling Chen, Mingsong Chen, Jianbing Xiahou, and Fan Lin. Urban traffic signal control based on multiobjective joint optimization. Scientific Programming, 2020(1):8839720:1– 8839720:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8839720.

Wang:2024:ASB

[WZF24] Jia-Bao Wang, Chun-An Zou, and Guang-Hui Fu. AWSMOTE: an SVM-based adaptive weighted SMOTE for class-imbalance learning. Scientific Programming, 2024(1):9947621:1– 9947621:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9947621.

Wang:2017:AEM

[WZG^{+}17] Lei Wang, Lei Zhao, Guan Gui, Baoyu Zheng, and Ruochen Huang. Adaptive ensemble method based on spatial characteristics for classifying imbalanced data. Scientific Programming, 2017(1): 3704525:1–3704525:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/3704525.

Wang:2023:AIF

[WZL23] Shuang Wang, Yi Zhang, and Yongkun Liu. Analysis of image features and TCM syndrome types of lobar pneumonia in children based on mean square deviation lung CT image registration algorithm. Scientific Programming, 2023(1): 4016914:1–4016914:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

                                        wiley.com/doi/epdf/10.
                                        1155/2021/4016914.

Wang:2023:DSB

[WZML23] Li Wang, Weiguang Zheng, Xiaojun Ma, and Shiming Lin. Denoising speech based on deep learning and wavelet decomposition. Scientific Programming, 2023(1): 8677043:1–8677043:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/8677043.

Wang:2022:ESA

[WZS^{+}22] Jiyuan Wang, Hanqing Zou, Shaokun Sun, Wenqian Xu, and Jie Jin. Edge segmentation algorithmbased prevention of recurrent laryngeal nerve injury in treatment of thyroid nodules using ultrasound images. Scientific Programming, 2022(1): 6583110:1–6583110:??, 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6583110.

Wang:2023:SVM

[WZS^{+}23] Jiyuan Wang, Hanqing Zou, Shaokun Sun, Wenqian Xu, and Jie Jin. Support vector machine parameter optimization for positron emission tomography images for estimation of recurrent laryngeal nerve injury with thyroid nodules. Scientific Programming, 2023(1): 2553244:1–2553244:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2553244.

Wang:2019:FMS

[WZW19] Yongzhi Wang, Jianwen Zheng, and Hui Wang. Fast mesh simplification method for three-dimensional geometric models with feature-preserving efficiency. Scientific Programming, 2019(1):4926190:1– 4926190:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/4926190.

Wang:2020:ADP

[WZW^{+}20] Pengwei Wang, Caihui Zhao, Yi Wei, Dong Wang, and Zhaohui Zhang. An adaptive data placement architecture in multicloud environments. Scientific Programming, 2020(1):1704258:1–1704258:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/1704258.

Wang:2023:MAM

[WZY^{+}23] Zhiyong Wang, Jingzhao Zhang, Yaran Yu, Jian Liu, Wei Liu, Na Jiang, and Donge Guo. Monitoring, analyzing, and modeling for single subsidence basin in coal mining areas based on SAR interferometry with l-band data. Scientific Programming, 2023(1):6662097:1– 6662097:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6662097.

Wang:2023:NNS

[WZZ23] Yan Wang, Yanmin Zhang, and Chunyue Zhang. Neural network segmentation algorithm-based magnetic resonance imaging to explore the relationship between cerebrospinal fluid flow with communicating hydrocephalus after decompressive craniectomy for craniocerebral injury. Scientific Programming, 2023 (1):4723329:1–4723329:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2021/4723329.

Xing:2017:PFT

[XBG17] Yiming Xing, Xiaojuan Ban, and Chong Guo. Probabilistic forecasting of traffic flow using multikernel based extreme learning machine. Scientific Programming, 2017(1):2073680:1– 2073680:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/2073680.

Xue:2023:VTP

[XC23] Han Xue and Tian Chai.

Vessel track prediction based on fractional gradient recurrent neural network with maneuvering behavior identification. Scientific Programming, 2023(1):5526082:1– 5526082:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5526082.

Xia:2018:RMP

[XCW18] Xu Xia, Zhigang Chen, and Wei Wei. Research on monitoring and prewarning system of accident in the coal mine based on big data. Scientific Programming, 2018(1):9308742:1– 9308742:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/9308742.

Xiao:2020:SPC

[XH20] Wen Xiao and Juan Hu. A survey of parallel clustering algorithms based on spark. Scientific Programming, 2020(1):8884926:1– 8884926:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8884926.

Xie:2023:PLB

[XHC^{+}23] Shu-Tong Xie, Zong-Bao He, Qiong Chen, Rong-Xin Chen, Qing-Zhao Kong, and Cun-Ying Song. Predicting learning behavior using log data in blended teaching. Scientific Programming, 2023(1):4327896:1– 4327896:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4327896.

Xiaolong:2023:SCD

[XHL^{+}23] He Xiaolong, Zhao Huiqi, Zhong Lunchao, Shah Nazir, Deng Jun, and Adnan Shahid Khan. Soft computing and decision support system for software process improvement: a systematic literature review. Scientific Programming, 2023(1): 7295627:1–7295627:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7295627.

Xu:2017:PGE

[XHSL17] Lixiong Xu, Yuan Huang, Xiaodong Shen, and Yang Liu. Parallelizing gene expression programming algorithm in enabling large-scale classification. Scientific Programming, 2017(1): 5081526:1–5081526:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/5081526.

Xiong:2023:BDD

[XHX23] Wenying Xiong, Dongqin Huang, and Wei Xu. Big data and deep learning model for FMS score prediction of aerobics athletes. Scientific Programming, 2023(1):3370580:1– 3370580:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/3370580.

Xiao:2023:RAG

[Xia23a] Daiyou Xiao. Research on the application of generative adversarial networks in the generation of stock market forecast trend images. Scientific Programming, 2023(1):7321671:1– 7321671:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/7321671.

Xiao:2023:RVI

[Xia23b] Yahui Xiao. Research on visual image texture rendering for artistic aided design. Scientific Programming, 2023(1): 1190912:1–1190912:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/1190912.

Xie:2020:MDL

[Xie20] Xiaoling Xie. MQHD: Dynamic load balancing method for energy-aware storage systems. Scientific Programming, 2020 (1):6843572:1–6843572:??, 2020. CODEN SCIPEV.

ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/6843572.

Xiao:2023:RHF

[XJH23] Wen Xiao, Ping Ji, and Juan Hu. RnkHEU: a hybrid feature selection method for predicting students’ performance. Scientific Programming, 2023(1): 1670593:1–1670593:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1670593.

Xie:2023:DPA

[XL23] Xingyu Xie and Bin Lv. Design of painting art style rendering system based on convolutional neural network. Scientific Programming, 2023(1):4708758:1– 4708758:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4708758.

Xie:2023:DLB

[XLB^{+}23] Xiaoxiao Xie, Zhen Li, Lu Bai, Ri Zhou, Canfeng Li, Xiaocheng Jiang, Jianwei Zuo, and Yulong Qi. Deep learning-based MRI in diagnosis of fracture of tibial plateau combined with meniscus injury. Scientific Programming, 2023(1):9935910:1– 9935910:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9935910.

Xue:2024:MVR

[XLF^{+}24] Lixia Xue, Meian Li, Liang Fan, Aixia Sun, and Tian Gao. Monocular vision ranging and camera focal length calibration. Scientific Programming, 2024(1):9979111:1– 9979111:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9979111.

Xuan:2023:WBF

[XLLL23] Zhidong Xuan, Haixia Liu, Chao Li, and Yongrong Liu. Wavelet bilateral filter algorithmbased high-frequency ultrasound image analysis on effects of skin scar repair. Scientific Programming, 2023(1):9573474:1– 9573474:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2021/9573474.

Xu:2018:RSD

[XLY18] Xinjun Xu, Yang Lei, and Feng Yang. Railway subgrade defect automatic recognition method based on improved faster R-CNN. Scientific Programming, 2018(1):4832972:1– 4832972:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/4832972.

Xu:2016:IBH

[XML^{+}16] Honglong Xu, Rui Mao, Hao Liao, He Zhang, Minhua Lu, and Guoliang Chen. Index based hidden outlier detection in metric space. Scientific Programming, 2016(??): 8048246:1–8048246:14, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
8048246/.

Xiao:2022:DSS

[XNL^{+}22] Hui Xiao, Shah Nazir, Hanmin Li, Habib Ullah Khan, and Chengwei Li. Decision support system to risk stratification in the acute coronary syndrome using fuzzy logic. Scientific Programming, 2022(1):6571905:1– 6571905:??, 2022. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6571905.

Xiaolong:2023:LBO

[XNLJ23] He Xiaolong, Shah Nazir, Zhong Lunchao, and Deng Jun. Librarybased overview of multicriteria decision making for continuous software improvement for internet of software industry. Scientific Programming, 2023(1):5519900:1– 5519900:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5519900.

Xue:2004:PPD

[XNQF04] Caijun Xue, Hong Nie, Qingying Qiu, and Peien Feng. A peer-to-peer distributed collaborative optimization system. Scientific Programming, 12 (2):121–131, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Xie:2023:RMC

[XSG23] Wenhao Xie, Yanhong She, and Qiao Guo. Research on multiple classification based on improved SVM algorithm for balanced binary decision tree. Scientific Programming, 2023(1):5560465:1– 5560465:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5560465.

Xilong:2023:DIW

[XSS^{+}23] Qu Xilong, Liu Shengzong, Fu Sha, He Hong, Hu Ying, and Xiao Leyi. Design and implementation of wireless environment monitoring system based on STM32. Scientific Programming, 2023(1): 6070664:1–6070664:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6070664.

Xiao:2016:UCV

[XST^{+}16] Lei Xiao, Yu Sheng, Guanlan Tan, Jianxin Wang, and Yi Pan. A user-customized virtual network platform for NaaS cloud. Scientific Programming, 2016(??): 9315672:1–9315672:6, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://www.hindawi.
                                        com/journals/sp/2016/
                                        9315672/.

Xu:2018:SPA

[XSZ18] Lei Xu, Anping Song, and Wu Zhang. Scalable parallel algorithm of multiple-relaxation-time lattice Boltzmann method with large eddy simulation on multi-GPUs. Scientific Programming, 2018(1):1298313:1– 1298313:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2018/1298313.

Xu:2015:MGS

[XTCC15] Rengan Xu, Xiaonan Tian, Sunita Chandrasekaran, and Barbara Chapman. Multi-GPU support on single node using directivebased programming model. Scientific Programming, 2015(??):621730:1–621730:15, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://www.hindawi.
                                        com/journals/sp/2015/
                                        621730/.

Xue:2020:LRP

[XTZ20] Chen Xue, Wuxu Tian, and Xiaotao Zhao. The literature review of platform economy. Scientific Programming, 2020(1): 8877128:1–8877128:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8877128.

Xu:2020:STV

[Xu20] Yunjun Xu. A sports training video classification model based on deep learning. Scientific Programming, 2020(1): 7252896:1–7252896:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7252896.

Xu:2023:DCT

[Xu23a] Jian Xu. Design of a cultural tourism passenger flow prediction model in the Yangtze River Delta based on regression analysis. Scientific Programming, 2023(1):9913468:1– 9913468:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9913468.

Xu:2023:RCM

[Xu23b] Ke Xu. Recognition and classification model of music genres and Chinese traditional musical instruments based on deep neural networks. Scientific Programming, 2023(1): 2348494:1–2348494:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2348494.

Xu:2023:ERT

[Xu23c]  Qiang Xu.  Evalua-

tion of rural tourism spatial pattern based on multifactor-weighted neural network algorithm model in big data era. Scientific Programming, 2023(1):8108287:1– 8108287:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8108287.

Xu:2023:RIE

[Xu23d] Yong Xu. Research on investment environment performance evaluation of blockchain industry with intuitionistic fuzzy CO-DAS method. Scientific Programming, 2023(1): 1387062:1–1387062:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1387062.

Xu:2023:RHJ

[Xu23e] Yu-Bin Xu. The research on huanglian jiedu decoction against atopic dermatitis. Scientific Programming, 2023(1): 5557908:1–5557908:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5557908.

Xu:2023:RVM

[Xu23f] Yuequn Xu. Recognition and volume measurement of intracranial hematoma through CT images under intelligent recognition algorithm. Scientific Programming, 2023(1): 8045719:1–8045719:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8045719.

Xue:2023:MFA

[Xue23] Li Xue. MRI findings of acute sports injury of the gastrocnemius muscle. Scientific Programming, 2023(1):9899036:1– 9899036:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9899036.

Xiao:2023:MBN

[XWG23] Ying Xiao, Deyan Wang, and Ya Gao. A mobile Bayesian network structure learning method using genetic incremental K2 algorithm and random attribute order technology. Scientific Programming, 2023(1):4743752:1– 4743752:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/4743752.

Xiong:2019:NCE

[XWL19] Li Xiong, Guo-Zheng Wang, and Hu-Chen Liu. New community estimation method in bipartite networks based on quality of filtering coefficient. Scientific Programming, 2019(1):4310561:1– 4310561:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2019/4310561.

Xu:2023:ISA

[XWL^{+}23] Cheng Xu, Hongjun Wu, Hongzhe Liu, Xuewei Li, Li Liu, and Pengfei Wang. An intelligent scheduling access privacy protection model of electric vehicle based on 5G-V2X. Scientific Programming, 2023(1):1198794:1– 1198794:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1198794.

Xuan:2023:LGD

[XWLL23] Zhidong Xuan, Na Wu, Chao Li, and Yongrong Liu. Local Gaussian distribution fitting boundary image segmentation algorithm for ultrasound images in avoiding recurrent laryngeal nerve injury during thyroid nodules treatment. Scientific Programming, 2023(1):6115040:1– 6115040:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6115040.

Xue:2023:RSI

[XXM^{+}23] Jiaming Xue, Shun Xiong, Chaoguang Men, Zhiming Liu, and Yongmei Liu. Remote-sensing image tile-level annotation based on discriminative features and expressive visual word descriptors. Scientific Programming, 2023(1):3490254:1– 3490254:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/3490254.

Xu:2023:AEP

[XXSL23] Wei Xu, Wenying Xiong, Zhe Shao, and Yun Li. Analysis of effectiveness and performance prediction of sports flipped classroom teaching based on neural networks. Scientific Programming, 2023 (1):5284457:1–5284457:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5284457.

Xiao:2021:RET

[XY21] Meng Xiao and Haibo Yi. RETRACTION: an efficient teaching model of international cooperation based on artificial intelligence. Scientific Programming, 2021(1):9856181:1– 9856181:??, ???? 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2024/9856181.

Xiao:2022:ETM

[XY22] Meng Xiao and Haibo Yi. An efficient teaching model of international cooperation based on artificial intelligence. Scientific Programming, 2022(1): 7049857:1–7049857:??, 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7049857.

Xia:2023:CMT

[XY23a] Xiongjun Xia and Jin Yan. Construction of music teaching evaluation model based on weighted naïve Bayes. Scientific Programming, 2023(1): 7196197:1–7196197:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7196197.

Xu:2023:ARF

[XY23b] Qingxiang Xu and Jiesen Yin. Application of random forest algorithm in physical education. Scientific Programming, 2023 (1):1996904:1–1996904:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1996904.

Xu:2023:DMI

[XYCC23] Lijun Xu, Shengzan Yan, Zhe Chen, and Xin Chen. Design of the museum interactive lighting system based on the digital twin technology. Scientific Programming, 2023(1):4824417:1– 4824417:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4824417.

Xie:2023:IAB

[XYL^{+}23] Xinjing Xie, Jun Yang, Guoxin Liang, Chunxiang Li, and Zhongyuan Li. Imaging algorithmbased real-time shear wave elastography combined with thyroglobulin antibodies in diagnosis of thyroid cancer. Scientific Programming, 2023(1):7189334:1– 7189334:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7189334.

Xu:2019:DDS

[XYY^{+}19] Jianzhong Xu, Fu Yan, Kumchol Yun, Sakaya Ronald, Fengshu Li, and Jun Guan. Dynamically dimensioned search embedded with piecewise opposition-based learning for global optimization. Scientific Programming, 2019(1):2401818:1– 2401818:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/2401818.

Xu:2020:CJF

[XYY^{+}20] Yuanyuan Xu, Wan Yan, Genke Yang, Jiliang Luo, Tao Li, and Jianan He. CenterFace: Joint face detection and alignment using face as point. Scientific Programming, 2020(1):7845384:1– 7845384:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/7845384.

Xia:2022:KMG

[XZ22] Shiliang Xia and Kaiyang Zhong. Knowledge mapping of green technology visualization with bibliometric tools. Scientific Programming, 2022(1): 6298813:1–6298813:??, 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6298813.

Xiao:2023:CMI

[XZ23a] Yue Xiao and Zhiqing Zeng. A construction method of intelligent manufacturing system under industry 4.0 model. Scientific Programming, 2023(1):4775237:1–4775237:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4775237.

Xu:2023:PME

[XZ23b] Yuquan Xu and Renfeng Zhao. A prediction model of endometrial cancer lesion metastasis under region of interest target detection algorithm. Scientific Programming, 2023(1): 9928842:1–9928842:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9928842.

Xu:2023:DTO

[XZ23c] Ziyan Xu and Wenxue Zhou. A data technology oriented to information fusion to build an intelligent accounting computerized model. Scientific Programming, 2023(1): 6031324:1–6031324:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6031324.

Xue:2024:CTI

[XZLZ24] Jun Xue, Jun Zhang, Li Li, and Ping Zang. Computed tomography images under iterative reconstruction algorithm in analysis of the efficacy of honey-fried herba ephedrae along with Western medicine on acute exacerbation chronic obstructive pulmonary disease. Scientific Programming, 2024(1): 8734187:1–8734187:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8734187.

Xue:2023:CPP

[XZNL23] Han Xue, Weicheng Zhang, Chao Ni, and Xiping Lu. Cross product and partitioned filtering-based Graham convex hull for buoy drifting area demarcating. Scientific Programming, 2023(1):7713884:1– 7713884:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7713884.

Xiao:2020:MWA

[XZZY20] Shaozhang Xiao, Zhengwei Zhang, Yue Zhang, and Changhui Yu. Multipurpose watermarking algorithm for medical images. Scientific Programming, 2020(1):8848885:1– 8848885:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8848885.

Yang:2022:RAM

[Yan22a]  Hongling Yang.  RE-

TRACTION: Application of multilayer neural network in sports psychology. Scientific Programming, 2022(1):9854125:1– 9854125:??, ???? 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2024/9854125.

Yangyudongnanxin:2022:FCR

[Yan22b] Guo Yangyudongnanxin. Financial credit risk control strategy based on weighted random forest algorithm. Scientific Programming, 2022(1): 6276155:1–6276155:??, 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6276155.

Yan:2023:VPM

[Yan23a] Qingbo Yan. A video production method of microclass combined with MOOC. Scientific Programming, 2023(1):9925165:1– 9925165:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9925165.

Yan:2023:VCN

[Yan23b] Rong Yan. The value of convolutional-neural-network-algorithm-based magnetic resonance imaging in the diagnosis of sports knee osteoarthropathy. Scientific Programming, 2023(1):2803857:1– 2803857:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2803857.

Yang:2023:RTD

[Yan23c]  Donghui Yang.  Re-

search on traffic detection method of secure transmission industrial Internet of Things based on computer vision. Scientific Programming, 2023(1):3406612:1– 3406612:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/3406612.

Yang:2023:RBG

[Yan23d] Ke Yang. Research on the basketball goal recognition method based on improved MobileNet. Scientific Programming, 2023 (1):5862037:1–5862037:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5862037.

Yang:2023:RRP

[Yan23e] Lijuan Yang. Research on the realization path of college English education based on the SVM algorithm model under the background of cloud computing and wireless communication. Scientific Programming, 2023(1): 6182824:1–6182824:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6182824.

Yang:2023:QEM

[Yan23f]  Yun Yang.  Quality

evaluation method of a mathematics teaching model reform based on an improved genetic algorithm. Scientific Programming, 2023(1): 6395349:1–6395349:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6395349.

Yang:2023:NBI

[Yan23g]  Zhuqing Yang.  A

novel brain image segmentation method using an improved 3D U-Net model. Scientific Programming, 2023(1): 4801077:1–4801077:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4801077.

Yanli:2023:RSL

[Yan23h]  Hui Yanli.  Research

on spoken language understanding based on deep learning. Scientific Programming, 2023(1): 8900304:1–8900304:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8900304.

Yang:2023:TDS

[YB23] Xiaozhou Yang and Fan Bai. Three-dimensional structure analysis of urban landscape based on big data technology and digital technology. Scientific Programming, 2023(1):7970870:1– 7970870:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7970870.

Ying:2024:BVS

[YB24] Fan Ying and Zhou Bo. Building virtual scene construction and environmental impact analysis based on image processing. Scientific Programming, 2024(1):9979862:1– 9979862:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9979862.

Ye:2023:PRM

[YC23] Xiangyu Ye and Mengmeng Chen. Personalized recommendation for mobile internet wealth management based on user behavior data analysis. Scientific Programming, 2023(1):9326932:1– 9326932:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9326932.

Yu:2023:MTF

[YCLK23] Fangping Yu, Hang Chen, Jiaqi Luo, and Haibo Kuang. Measuring total factor productivity of China provincial nonlife insurance market: a DEA-Malmquist index method. Scientific Programming, 2023(1): 3022658:1–3022658:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3022658.

Yang:2017:ROA

[YCWL17] Lifeng Yang, Liangming Chen, Ningwei Wang, and Zhifang Liao. Routing optimization algorithms based on node compression in big data environment. Scientific Programming, 2017(1):2056501:1– 2056501:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/2056501.

Yang:2016:NPM

[YDC^{+}16] Quan Yang, Zhihui Du, Zhoujian Cao, Jian Tao, and David A. Bader. A new parallel method for binary black hole simulations. Scientific Programming, 2016(??):

2360492:1–2360492:14, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://www.hindawi.
                                        com/journals/sp/2016/
                                        2360492/.

Yang:2024:DCI

[YDC^{+}24] Feng Yang, Guixin Dong, Chaoran Cui, Xiaojie Li, Yaxi Su, and Yilong Yin. Digital currency illegal behavior detection based on mutual information prior loss. Scientific Programming, 2024(1):9954204:1– 9954204:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9954204.

Ye:2023:ACP

[Ye23] Dan Ye. An algorithm for construction project cost forecast based on particle swarm optimizationguided BP neural network. Scientific Programming, 2023(1):4309495:1– 4309495:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/4309495.

Yang:2016:URA

[YG16] Guibin Yang and Hongyu Gao. Uncertain risk assessment of knowledge management: Based on set pair analysis. Scientific Programming, 2016 (??):2025892:1–2025892:8, ???? 2016. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2016/
2025892/.

Yang:2024:BEA

[YGZ^{+}24] Wenbing Yang, Xiaoqi Gao, Chunlei Zhang, Feng Tong, Guantian Chen, and Zhijian Xiao. Bridge extraction algorithm based on deep learning and high-resolution satellite image. Scientific Programming, 2024(1): 9961963:1–9961963:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9961963.

Yasir:2023:SCD

[YHH^{+}23] Muhammad Yasir, Sheng Hui, Zheng Hongxia, Md Sakaouth Hossain, Hong Fan, Li Zhang, and Zhao Jixiang. A spatiotemporal change detection analysis of coastline data in Qingdao, East China. Scientific Programming, 2023(1): 6632450:1–6632450:??, 2023.

CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6632450.

Ye:2023:TGO

[YHJ^{+}23] Jianming Ye, He Huang, Weiwei Jiang, Xiaomei Xu, Chun Xie, Bo Lu, Xiangcai Wang, and Xiaobo Lai. Tumor grade and overall survival prediction of gliomas using radiomics. Scientific Programming, 2023(1): 9913466:1–9913466:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9913466.

Yang:2023:SPB

[YHZ^{+}23] Junfeng Yang, Yuwen Huang, Ruili Zhang, Fuxian Huang, Qinggang Meng, and Shixin Feng. Study on PPG biometric recognition based on multifeature extraction and naive Bayes classifier. Scientific Programming, 2023(1):5597624:1– 5597624:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5597624.

Yuan:2023:ABM

[YHZG23] Feng Yuan, Ling He, Yancui Zhu, and Honglei Guo. ANOVA-based magnetic resonance imaging in the diagnosis of prostate cancer with bone metastasis and rehabilitation treatment. Scientific Programming, 2023(1):9926652:1– 9926652:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9926652.

Yi:2023:NTT

[Yi23] Pan Yi. A novel trade transaction agreement algorithm using blockchain consensus mechanism. Scientific Programming, 2023(1):5343337:1–5343337:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5343337.

Yin:2023:IDB

[Yin23]  Zhanneng Yin.  Im-

proved DenseNet-based MRI in pulmonary nodules diagnosis and benign and malignant differentiation. Scientific Programming, 2023(1):8689068:1– 8689068:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/8689068.

Yang:2023:ARM

[YJ23] BaoPing Yang and Kun Jiang. Automatic repair method for D2D communication routing buffer overflow vulnerability in cellular network. Scientific Programming, 2023(1):3963574:1– 3963574:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/3963574.

Yu:2020:BBS

[YL20] Liang Yu and Da Lin. Bayesian-based search decision framework and search strategy analysis in probabilistic search. Scientific Programming, 2020(1):8865381:1–8865381:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/8865381.

Yang:2023:ADM

[YL23] Jinpeng Yang and Ying Liu. Application of data mining in the evaluation of enterprise lean management effect. Scientific Programming, 2023(1):4774140:1– 4774140:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4774140.

Yan:2023:VTI

[YLL^{+}23] Caoxin Yan, Zhiyan Luo, Zimei Lin, Shouxing Xu, Yunkai Luo, and Jian Chen. Value of thyroid imaging reporting and data system in initial Bethesda category III thyroid nodules. Scientific Programming, 2023(1):3482111:1– 3482111:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3482111.

Yin:2020:PIM

[YLLL20] Rongwang Yin, Qingyu Li, Peichao Li, and Detang Lu. Parameter identification of multistage fracturing horizontal well based on PSO-RBF neural network. Scientific Programming, 2020(1):6810903:1– 6810903:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

                                        wiley.com/doi/epdf/10.
                                        1155/2020/6810903.

Yu:2023:OAR

[YLLY23] Miao Yu, Junhui Li, Yue Li, and Lejing Yu. Optimized algebraic reconstruction algorithm based hemodynamic changes in color Doppler ultrasound in monitoring the patients with colorectal cancer undergoing peripherally inserted central catheter and catheterization nursing. Scientific Programming, 2023(1):8375290:1– 8375290:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/8375290.

Yin:2021:MDP

[YLW21] L. Yin, Y. Liu, and Z. Wang. Model for design of portfolio venture investment contract when taking moral hazards into account. Scientific Programming, 2021(1): 8821371:1–8821371:??, 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/8821371.

Yang:2018:CEE

[YLWW18] Yang Yang, Ming Li, Chen Wang, and QingYue Wei. Cylindricity error evaluation based on an improved harmony search algorithm. Scientific Programming, 2018(1): 2483781:1–2483781:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/2483781.

Yu:2023:NSP

[YLZ23] Wangwang. Yu, Jun.

Liu, and Jie. Zhou. A novel sparrow particle swarm algorithm (SPSA) for unmanned aerial vehicle path planning. Scientific Programming, 2023(1):5158304:1– 5158304:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5158304.

Yuan:2023:OCT

[YN23] Fang Yuan and Yong Nie. Online classroom teaching quality evaluation system based on facial feature recognition. Scientific Programming, 2023(1):7374846:1– 7374846:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

                                        wiley.com/doi/epdf/10.
                                        1155/2021/7374846.

Yu:2023:MRF

[YNK23] Ning Yu, Lin Nan, and Tao Ku. Multipolicy robot-following model based on reinforcement learning. Scientific Programming, 2023(1): 5692105:1–5692105:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5692105.

Yin:2024:RMC

[YP24] Linsen Yin and Ane Pan. Replacing management or not: Contract renegotiation to prevent double moral hazards of venture capital investments. Scientific Programming, 2024(1): 9974235:1–9974235:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9974235.

Yuan:2023:MID

[YPC^{+}23] Quan Yuan, Zhenyun Peng, Zhencheng Chen, Yanke Guo, Bin Yang, and Xiangyan Zeng. Medical image denoising algorithm based on sparse nonlocal regularized weighted coding and low rank constraint. Scientific Programming, 2023(1):7008406:1–7008406:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7008406.

Yannibelli:2021:CAN

[YPM^{+}21] Virginia Yannibelli, Elina Pacini, David Monge, Cristian Mateos, and Guillermo Rodriguez. A comparative analysis of NSGA-II and NSGA-III for autoscaling parameter sweep experiments in the cloud. Scientific Programming, 2021(1):4653204:1– 4653204:??, 2021. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/4653204.

Yang:2023:CVH

[YPY23] Jing Yang, Hanfei Peng, and Guangyan Yan. The clinical value of highfrequency ultrasound in the diagnosis of psoriatic arthritis. Scientific Programming, 2023(1): 8292597:1–8292597:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8292597.

Yang:2018:HHP

[YSL^{+}18] Haoduo Yang, Huayou Su, Qiang Lan, Mei Wen, and Chunyuan Zhang. HPGraph: Highperformance graph analytics with productivity on the GPU. Scientific Programming, 2018(1): 9340697:1–9340697:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2018/9340697.

Yu:2023:SRW

[YSSL23] Hao Yu, Xu Sun, Wei Deng Solvang, and Gilbert Laporte. Solving a realworld urban postal service system redesign problem. Scientific Programming, 2023(1):3058472:1– 3058472:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/3058472.

Yao:2023:MLA

[YSZ^{+}23] Chunhua Yao, Xinyu Song, Xuelei Zhang, Weicheng Zhao, and Ao Feng. Multitask learning for aspect-based sentiment classification. Scientific Programming, 2023 (1):2055555:1–2055555:??, 2023. CODEN SCIPEV.

ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2055555.

Yao:2022:DLB

[YT22] Wei Yao and Stefanie Thomas. Deep learningbased magnetic resonance imaging image feature analysis for pathological classification of brain glioma. Scientific Programming, 2022(1): 6778009:1–6778009:??, 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6778009.

Yin:2023:PMS

[YT23] Xiaoting Yin and Xiaosha Tao. Prediction of merchandise sales on e-commerce platforms based on data mining and deep learning. Scientific Programming, 2023(1):2179692:1– 2179692:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2179692.

Yamazaki:2015:CLR

[YTD15] Ichitaro Yamazaki, Stanimire Tomov, and Jack Dongarra. Computing low-rank approximation of a dense matrix on multicore CPUs with a GPU and its application to solving a hierarchically semiseparable linear system of equations. Scientific Programming, 2015(??):246019:1– 246019:17, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
246019/.

Yu:2022:RMM

[Yu22] Zhao Yu. Research on multimodal music emotion recognition method based on image sequence. Scientific Programming, 2022(1):7087588:1–7087588:??, 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7087588.

Yuan:2020:FDR

[Yua20] Zhenguo Yuan. Face detection and recognition based on visual attention mechanism guidance model in unrestricted posture. Scientific Programming, 2020(1):8861987:1– 8861987:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2020/8861987.

Yu:2024:DAL

[YW24] Dian Yu and Tongyao Wang. A distributed algorithm for large-scale linearly coupled resource allocation problems with selfish agents. Scientific Programming, 2024(1): 9939805:1–9939805:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9939805.

Yuan:2018:RMP

[YWC^{+}18] Zejian Yuan, Xin Wang, Jiuwen Cao, Haiquan Zhao, and Badong Chen. Robust matching pursuit extreme learning machines. Scientific Programming, 2018(1): 4563040:1–4563040:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/4563040.

Yang:2020:CCS

[YWGL20] Haitao Yang, Guan Wang, Jinxia Gao, and Jie Liu. A crossover comparison of the sensitivity and the specificity between BIS and AEP in predicting unconsciousness in general anesthesia. Scientific Programming, 2020(1):8899957:1– 8899957:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8899957.

Yao:2020:MAL

[YWGZ20] Tuozhong Yao, Wenfeng Wang, Yuhong Gu, and Qiuguo Zhu. Multiview active learning for scene classification with high-level semanticbased hypothesis generation. Scientific Programming, 2020(1):3878153:1– 3878153:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/3878153.

Yang:2020:EGM

[YWH20] Kun Yang, Wan Wang, and Bin Hu. Evolutionary game models on multiagent collaborative mechanism in responsible innovation. Scientific Programming, 2020(1):8875099:1– 8875099:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8875099.

Yu:2023:CJR

[YWX23] Liqun Yu, Lu Wang, and Yongxing Xu. Combination of joint representation and adaptive weighting for multiple features with application to SAR target recognition. Scientific Programming, 2023(1):9063419:1– 9063419:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9063419.

Yan:2023:ECS

[YX23] Haitao Yan and Yongzhi Xu. Energy control strategy for parallel hybrid electric vehicle based on terminal neural network. Scientific Programming, 2023(1):7328008:1– 7328008:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7328008.

Xie:2023:IAT

[yXHtJC23] Zhi ying Xie, Yuan-Rong He, Yuan tong Jiang, and Chih-Cheng Chen. Improved AND/ OR tree search algorithm in analysis of stochastic and time-dependent shortest path problem. Scientific Programming, 2023(1): 9922466:1–9922466:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9922466.

Yang:2020:DRD

[YXL^{+}20] Meijun Yang, Xiaoyan Xiao, Zhi Liu, Longkun Sun, Wei Guo, Lizhen Cui, Dianmin Sun, Pengfei Zhang, and Guang Yang. Deep RetinaNet for dynamic left ventricle detection in multiview echocardiography classification. Scientific Programming, 2020(1): 7025403:1–7025403:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/7025403.

Yang:2022:MDT

[YXW^{+}22] Zhen Yang, Xuefei Xu, Keke Wang, Xin Li, and Chi Ma. Multitarget detection of transmission lines based on DANet and YOLOv4. Scientific Programming, 2022(1): 6235452:1–6235452:??, 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6235452.

Yang:2023:CCP

[YXZ23] Rui Yang, Mengying Xu, and Jie Zhou. Clone chaotic parallel evolutionary algorithm for low-energy clustering in high-density wireless sensor networks. Scientific Programming, 2023(1): 6630322:1–6630322:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6630322.

Yan:2023:STF

[YY23a] Yingying Yan and Daguang Yang. A stock trend forecast algorithm based on deep neural networks. Scientific Programming, 2023(1):7510641:1–7510641:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7510641.

Yang:2023:RTV

[YY23b] Qiong Yang and Lifeng Yu. Recognition of taxi violations based on semantic segmentation of PSPNet and improved YOLOv3. Scientific Programming, 2023(1): 4520190:1–4520190:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/4520190.

Yu:2023:SPA

[YY23c] Jia Yu and Rui Yang. Study on the predictive algorithm of plant restoration under heavy metals. Scientific Programming, 2023(1):6193182:1– 6193182:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6193182.

Yang:2019:LVG

[YYF^{+}19] Xingguang Yang, Huiqun Yu, Guisheng Fan, Kai Shi, and Liqiong Chen. Local versus global models for just-in-time software defect prediction. Scientific Programming, 2019(1):2384706:1–2384706:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2019/2384706.

Yang:2017:LAB

[YYL17] Xi Yang, Fang Yan, and Jun Liu. 3D localization algorithm based on Voronoi diagram and rank sequence in wireless sensor network. Scientific Programming, 2017(1):4769710:1– 4769710:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/4769710.

Yang:2023:SFD

[YYW^{+}23] Zhe Yang, Shi Ying, Bingming Wang, Yiyao Li, Bo Dong, Jiangyi Geng, and Ting Zhang. A system fault diagnosis method with a reclustering algorithm. Scientific Programming, 2023(1): 6617882:1–6617882:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6617882.

Yang:2023:HAF

[YYY23] Tao Yang, Guoliang Yuan, and Jing Yan. Health analysis of footballer using big data and deep learning. Scientific Programming, 2023(1): 9608147:1–9608147:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9608147.

Yin:2017:ADI

[YYZ^{+}17] Hong-Jian Yin, Hai Yu, Yu-Li Zhao, Zhi-Liang Zhu, and Wei Zhang. Analysis of the dynamic influence of social network nodes. Scientific Programming, 2017(1): 5046905:1–5046905:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/5046905.

Yang:2016:RSB

[YZ16] Kai Yang and Zhuhong Zhang. Racing sampling based microimmune optimization approach solving constrained expected value programming. Scientific Programming, 2016(??):2148362:1– 2148362:9, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2016/
2148362/.

Yuan:2020:VDL

[YZ20] Chao Yuan and Jie Zhang. Violation detection of live video based on deep learning. Scientific Programming, 2020(1):1895341:1– 1895341:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/1895341.

Zheng:2023:AQC

[yZcR23] Hai yan Zheng and Xing cheng Ran. Application of QR code online testing technology in nursing teaching in colleges and universities. Scientific Programming, 2023(1):6380501:1– 6380501:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6380501.

Yan:2023:DCC

[YZL23] Wei Yan, Huijun Zhou, and Hui Li. Decision and coordination of crossborder e-commerce supply chain: Based on four modes of cooperation. Scientific Programming, 2023(1):5561357:1– 5561357:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5561357.

Yan:2023:SAS

[YZQ^{+}23] Wei Yan, Lifan Zhou, Zhengjiang Qian, Le Xiao, and Haixia Zhu. Sentiment analysis of student texts using the CNN-BiGRU-AT model. Scientific Programming, 2023 (1):8405623:1–8405623:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8405623.

Yu:2017:EFJ

[YZSZ17] Lianfei Yu, Cheng Zhu, Jianmai Shi, and Weiming Zhang. An extended flexible job shop scheduling model for flight deck scheduling with priority, parallel operations, and sequence flexibility. Scientific Programming, 2017(1):2463252:1– 2463252:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/2463252.

Yuan:2023:CEC

[YZWF23] Yongjie Yuan, Yongjun Zhang, Junyuan Wang, and Ping Fang. Classification of electrocardiogram of congenital heart disease patients by neural network algorithms. Scientific Programming, 2023(1): 3801675:1–3801675:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3801675.

Yu:2017:CBA

[YZX17] Z. Yu, Y. Zuo, and W. C. Xiong. Concurrency bug avoiding based on optimized software transactional memory. Scientific Programming, 2017(1):9404323:1– 9404323:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/9404323.

Yang:2015:IOE

[YZZ15] Xu Yang, Mingbin Zeng, and Yanjun Zhang. Implementing and optimizing of entire system toolkit of VLIW DSP processors for embedded sensor-based systems. Scientific Programming, 2015(??):507896:1– 507896:7, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
507896/.

Zhu:2010:SFT

[ZA10] Qian Zhu and Gagan Agrawal. Supporting fault-tolerance for timecritical events in distributed environments. Scientific Programming, 18(1):51–76, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Zhang:2021:UVC

[ZAW^{+}21] Ying Zhang, Farhad Ali, Kunhao Wang, Shah Nazir, and Zeqi Leng. Utilizing virtual crowd for global software development. Scientific Programming, 2021(1):2259594:1– 2259594:??, 2021. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/2259594.

Zhao:2020:FER

[ZC20] Yifeng Zhao and Deyun Chen. A facial expression recognition method using improved capsule network model. Scientific Programming, 2020(1): 8845176:1–8845176:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/8845176.

Zhou:2023:MIU

[ZC23] Bin Zhou and Min Chen. MRI images under the optimized registration algorithm for primary open angle glaucoma visual path damage. Scientific Programming, 2023(1):4921276:1– 4921276:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4921276.

ZhuanSun:2023:MEB

[ZCCS23] Fengqin ZhuanSun, Jiaojiao Chen, Wenlong Chen, and Yan Sun. The mechanism of evolution and balance for ecommerce ecosystem under blockchain. Scientific Programming, 2023(1): 5984306:1–5984306:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5984306.

Zhang:2023:ARP

[ZCG23] Zheng Zhang, Juan Chen, and Qing Guo. AGVs route planning based on region-segmentation dynamic programming in smart road network systems. Scientific Programming, 2023(1):9589476:1– 9589476:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9589476.

Zhang:2020:OEE

[ZCL^{+}20] Jun Zhang, Jianpeng Chang, Ping Lin, Minzi Song, and Yanqiu Gong. Operation efficiency evaluation of the China– Europe freight train based on grey cross-efficiency DEA. Scientific Programming, 2020(1):8843733:1– 8843733:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8843733.

Zeng:2024:OCA

[ZCL^{+}24a] Chao Zeng, Jing Chen, Wenbing Liu, Kang Liang, Hui Li, Jing Wang, Jingge Li, and Haibo Xu. Optimized CSMRI algorithmbased MRI image analysis in the active rehabilitation method for patients with acute cerebral infarction. Scientific Programming, 2024(1):8494784:1– 8494784:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8494784.

Zheng:2024:AIB

[ZCL^{+}24b] Yanling Zheng, Yongqing Cao, Jiheng Liu, Hui Chen, and Liming Peng. Artificial intelligence-based dynamic contrast-enhanced magnetic resonance imaging for quantitative evaluation of the efficacy of targeted therapy for liver metastasis from colon cancer and nursing. Scientific Programming, 2024(1):8734108:1– 8734108:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8734108.

Zhang:2023:MCE

[ZCLL23] Zhili Zhang, Guo Cheng, Guifang Liu, and Gaixia Li. Multilevel clusteringevolutionary random support vector machine cluster algorithm-based blood oxygenation leveldependent functional magnetic resonance imaging images in analysis of therapeutic effects on cerebral ischemic stroke. Scientific Programming, 2023(1): 7706782:1–7706782:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7706782.

Zhang:2023:FCP

[ZCS23] Qiuwen Zhang, Tengyao Cui, and Rijian Su. Fast CU partition decision based on texture for h.266/VVC. Scientific Programming, 2023(1): 5527713:1–5527713:??, 2023.

CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5527713.

Zhang:2017:FAR

[ZCSL17] Xiaomei Zhang, Qiang Chen, Zhicai Shi, and Jianru Liang. Faultaware resource allocation for heterogeneous data sources with multipath routing. Scientific Programming, 2017(1): 9749581:1–9749581:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2017/9749581.

Zong:2020:MWD

[ZCT^{+}20] Yongshuo Zong, Jinling Chen, Siyi Tao, Cheng Wang, and Jianbing Xiahou. Moving window differential evolution independent component analysisbased operational modal analysis for slow linear time-varying structures. Scientific Programming, 2020(1):8879086:1– 8879086:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/8879086.

Zhang:2019:GBS

[ZCW^{+}19] Xianzhe Zhang, Gang Chen, Jiechen Wang, Manchun Li, and Liang Cheng. A GIS-based spatial-temporal autoregressive model for forecasting marine traffic volume of a shipping network. Scientific Programming, 2019(1):2345450:1– 2345450:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/2345450.

Zhang:2016:IAS

[ZCXQ16] Zhiying Zhang, Xiaozhen Chen, Wenwen Xiao, and Guijie Qi. Identifying and analyzing strong components of an industrial network based on cycle degree. Scientific Programming, 2016(??): 7340859:1–7340859:11, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
7340859/.

Zhang:2021:AEF

[ZCZ21] Tianjiao Zhang, Qiangxing Chen, and Xiaolong Zhu. Analysis of enterprise financial risk early warning model based on the evidence theory and whitening weight function. Scientific Programming, 2021(1):9207782:1– 9207782:??, ???? 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2024/9207782.

Zhao:2023:ROC

[ZCZ^{+}23] Zhexia Zhao, Xiaoyan Chen, Yanxia Zhang, Qian Song, Jiping Xue, and Chunsong Kang. Receiver operator characteristics (ROC) analysis of real-time shear wave elastography information health monitoring in liver fibrosis with chronic hepatitis B virus infection. Scientific Programming, 2023(1):8393697:1– 8393697:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8393697.

Zhao:2023:INT

[ZD23a] Guotao Zhao and Jie Ding. Image network teaching resource retrieval algorithm based on deep hash algorithm. Scientific Programming, 2023(1):9683908:1–9683908:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9683908.

Zhu:2023:RKT

[ZD23b] Yikun Zhu and Zhiling Du. Research on the key technologies of network security-oriented situation prediction. Scientific Programming, 2023(1): 5527746:1–5527746:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5527746.

Zeng:2023:TSC

[ZDMY23] Chenxi Zeng, Zhongliang Deng, Jiyang Ma, and Shengsong Yang. Twostage channel adaptive algorithm for unmanned aerial vehicles localization with cellular networks. Scientific Programming, 2023(1):8946885:1–8946885:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8946885.

Zeng:2022:RST

[Zen22]  Yan Zeng.  Research

on social talent governance based on genetic algorithm. Scientific Programming, 2022(1): 6288679:1–6288679:??, 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6288679.

Zhang:2024:AFA

[ZFFP24] Liyi Zhang, Mingyue Fu, Teng Fei, and Xuhua Pan. Application of FWA-Artificial fish swarm algorithm in the location of low-carbon cold chain logistics distribution center in beijingtianjin-hebei metropolitan area. Scientific Programming, 2024(1): 9945583:1–9945583:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9945583.

Zhang:2017:ACT

[ZFX17] Sen Zhang, Qiang Fu, and Wendong Xiao. Advertisement click-through rate prediction based on the weighted-ELM and Adaboost algorithm. Scientific Programming, 2017(1):2938369:1–2938369:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2017/2938369.

Zygmunt:2015:CUA

[ZG15] Anna Zygmunt and Bogdan Gliwa. The comparison of users activity on the example of Polish and American blogosphere. Scientific Programming, 2015 (??):907547:1–907547:11, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
907547/.

Zhang:2018:IGP

[ZG18] Lixia Zhang and Jianliang Gao. Incremental graph pattern matching algorithm for big graph data. Scientific Programming, 2018(1):6749561:1– 6749561:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/6749561.

Zhou:2023:RIC

[ZG23] Acheng Zhou and Chao Gao. Research on the involvement of computer graphics algorithms in systems for the creation of public sculpture. Scientific Programming, 2023(1):4520642:1– 4520642:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

                                        wiley.com/doi/epdf/10.
                                        1155/2021/4520642.

Zhang:2023:LFM

[ZGFL23] Kai Zhang, Wei Guo, Jian Feng, and Mei Liu. Load forecasting method based on improved deep learning in cloud computing environment. Scientific Programming, 2023(1):3250732:1– 3250732:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/3250732.

Zhao:2018:PFS

[ZGL18] Wenbing Zhao, Longxiang Gao, and Anfeng Liu. Programming foundations for scientific big data analytics. Scientific Programming, 2018(1):2707604:1– 2707604:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2018/2707604.

Zhao:2023:EKC

[ZGL23] Xian Zhao, Zhenjie Guo, and Shanqin Liu. Exploring key competencies and professional development of music teachers in primary schools in the era of artificial intelligence. Scientific Programming, 2023(1):5097003:1– 5097003:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5097003.

Zhou:2016:API

[ZGS16] Yancong Zhou, Xudong Guo, and Xiaochen Sun. Acquisition pricing and inventory decisions on dual-source spare-part system with final production and remanufacturing. Scientific Programming, 2016(??): 8038045:1–8038045:10, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
8038045/.

Zou:2023:MDI

[ZGT^{+}23] Fumin Zou, Feng Guo, Junshan Tian, Sijie Luo, Xiang Yu, Qing Gu, and Lyuchao Liao. The method of dynamic identification of the maximum speed limit of expressway based on electronic toll collection data. Scientific Programming, 2023(1):4702669:1– 4702669:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/4702669.

Zhang:2008:DGS

[ZGW08] Ji Zhang, Qigang Gao, and Hai Wang. Discover gene specific local co-regulations from time-course gene expression data. Scientific Programming, 16(1):31– 47, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Zheng:2020:SDP

[ZGY^{+}20] Shang Zheng, Jinjing Gai, Hualong Yu, Haitao Zou, and Shang Gao. Software defect prediction based on fuzzy weighted extreme learning machine with relative density information. Scientific Programming, 2020(1):8852705:1– 8852705:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/8852705.

Zhi:2019:ITW

[ZH19] Huilai Zhi and Shulin Hu. Influences of three-way concept lattice caused by variations of attribute values. Scientific Programming, 2019(1):7865197:1– 7865197:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/7865197.

Zhang:2023:MIC

[ZH23] Bing Zhang and Xu Hu.

A medical image classification model based on adversarial lesion enhancement. Scientific Programming, 2023(1):4265650:1– 4265650:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4265650.

Zhang:2020:MRB

[Zha20a] Lin Zhang. The model of the relationship between urban rail transit and residential location. Scientific Programming, 2020(1):8851637:1– 8851637:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8851637.

Zhang:2020:SMR

[Zha20b] Song Zhang. Selection of multimode resourceconstrained project scheduling scheme based on DEA method. Scientific Programming, 2020(1): 2742437:1–2742437:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/2742437.

Zhang:2022:RDD

[Zha22] Jie Zhang. RETRAC-TION: Data-driven learning teaching model of College English based on mega data analysis. Scientific Programming, 2022(1):9871025:1– 9871025:??, ???? 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2024/9871025.

Zhang:2023:EEB

[Zha23a] Chuanlong Zhang. Evaluation of the efficacy of budesonide combined with pulmonary surfactants on the neonatal respiratory distress syndrome by pulmonary ultrasonography. Scientific Programming, 2023(1): 2329524:1–2329524:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2329524.

Zhang:2023:VKR

[Zha23b] Hongli Zhang. Voice keyword retrieval method using attention mechanism and multimodal information fusion. Scientific Programming, 2023(1): 6662841:1–6662841:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6662841.

Zhang:2023:MFE

[Zha23c]  Jingwen Zhang.  Mu-

sic feature extraction and classification algorithm based on deep learning. Scientific Programming, 2023(1):1651560:1– 1651560:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1651560.

Zhang:2023:ADL

[Zha23d] Menghan Zhang. Applications of deep learning in news text classification. Scientific Programming, 2023(1):6095354:1– 6095354:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6095354.

Zhang:2023:HPP

[Zha23e] Qingqi Zhang. Housing price prediction based on multiple linear regression. Scientific Programming, 2023(1):7678931:1– 7678931:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/7678931.

Zhang:2023:DLB

[Zha23f] Wenbo Zhang. Deeplearning-based MRI images for analysis of sportinduced ankle joint injury. Scientific Programming, 2023(1):5544160:1– 5544160:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5544160.

Zhang:2023:PSU

[Zha23g] Xiaoqian Zhang. Planning the structure of university teaching staff based on multiobjective optimization method. Scientific Programming, 2023(1):1773561:1–1773561:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/1773561.

Zhang:2023:ECM

[Zha23h] Yangqianhui Zhang. Exploration of cross-modal text generation methods in smart justice. Scientific Programming, 2023(1): 3225933:1–3225933:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3225933.

Zhang:2023:NMA

[Zha23i] Zhe Zhang. New media advertising communication analysis model based on extension neural network. Scientific Programming, 2023(1):5969446:1– 5969446:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5969446.

Zhang:2023:MRP

[Zha23j] Zhihao Zhang. A method of recommending physical education network course resources based on collaborative filtering technology. Scientific Programming, 2023(1):9531111:1– 9531111:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2021/9531111.

Zhao:2023:AVI

[Zha23k]  Jing Zhao.  Art vi-

sual image transmission method based on cartesian genetic programming. Scientific Programming, 2023(1):4628563:1– 4628563:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4628563.

Zhao:2023:OFR

[Zha23l] Wanying Zhao. Optimal fixed route for multimodal transportation of vehicle logistics in context of soft time windows. Scientific Programming, 2023(1):2657918:1– 2657918:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2657918.

Zhao:2023:RDA

[Zha23m]  Yu Zhao.  Research

and design of automatic scoring algorithm for English composition based on machine learning. Scientific Programming, 2023(1):3429463:1– 3429463:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3429463.

Zheng:2010:OUP

[Zhe10] Yili Zheng. Optimizing UPC programs for multi-core systems. Scientific Programming, 18 (3–4):183–191, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Zhu:2023:CTI

[ZHHX23] Ling Zhu, Yucheng He, Nan He, and Lanhua Xiao. Computed tomography image based on intelligent segmentation algorithm in the diagnosis of ovarian tumor. Scientific Programming, 2023(1):7323654:1– 7323654:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7323654.

Zhou:2016:VMP

[ZHL16] Zhou Zhou, Zhigang Hu, and Keqin Li. Virtual machine placement algorithm for both energyawareness and SLA violation reduction in cloud data centers. Scientific Programming, 2016(??):

5612039:1–5612039:11, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
5612039/.

Zhang:2023:CME

[ZHL^{+}23] Qinghui Zhang, Lei Hou, Pengtao Lv, Mengya Zhang, and Hongwei Yang. Chinese medical entity recognition model based on character and word vector fusion. Scientific Programming, 2023(1):5933652:1– 5933652:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5933652.

Zhou:2020:HMC

[Zho20] Guanghong Zhou. Humanmachine cooperation and path planning for complex road conditions. Scientific Programming, 2020(1): 7262281:1–7262281:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7262281.

Zhou:2022:PVU

[Zho22] Yuehong Zhou. Prediction and value of ultrasound image in diagnosis of fetal central nervous system malformation under deep learning algorithm. Scientific Programming, 2022(1): 6246274:1–6246274:??, 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6246274.

Zhou:2023:PUR

[Zho23] Wusheng Zhou. Prediction of urban and rural tourism economic forecast based on machine learning. Scientific Programming, 2023(1):4072499:1– 4072499:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4072499.

Zhao:2023:SFE

[ZHP^{+}23] Zheng-Yang Zhao, Wen-Zhun Huang, Jie Pan, Yu-An Huang, Shan-Wen Zhang, and Chang-Qing Yu. A sparse feature extraction method with elastic net for drugtarget interaction identification. Scientific Programming, 2023(1): 6686409:1–6686409:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

                                        wiley.com/doi/epdf/10.
                                        1155/2021/6686409.

Zhu:2022:SBD

[Zhu22] Weiming Zhu. A study of big-data-driven data visualization and visual communication design patterns. Scientific Programming, 2022(1):6704937:1– 6704937:??, 2022. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6704937.

Zhu:2023:RHR

[Zhu23a]  Hong Zhu.  Research

on human resource recommendation algorithm based on machine learning. Scientific Programming, 2023(1):8387277:1– 8387277:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/8387277.

Zhu:2023:NCR

[Zhu23b] Qianyao Zhu. Network course recommendation system based on doublelayer attention mechanism. Scientific Programming, 2023(1):7613511:1– 7613511:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7613511.

Zhou:2021:MDA

[ZHW^{+}21] Yilian Zhou, Ligang He, Bin Wang, Yi Su, and Hao Chen. MCAF: Developing an annotationbased offloading framework for mobile cloud computing. Scientific Programming, 2021(1): 5304612:1–5304612:??, 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/5304612.

Zhang:2022:RSI

[ZHY22] Kai Zhang, Chengquan Hu, and Hang Yu. Remote sensing image land classification based on deep learning. Scientific Programming, 2022(1): 6203444:1–6203444:??, 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6203444.

Zhai:2020:REB

[ZHZF20] Pengjun Zhai, Xin Huang, Beibei Zhang, and Yu Fang. Relation extraction based on fusion dependency parsing from Chinese EMRs. Scientific Programming, 2020(1):8658040:1– 8658040:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2020/8658040.

Zhang:2019:IMS

[ZHZS19] Hong Zhang, Minghu Ha, Hongyu Zhao, and Jianwei Song. Inexact multistage stochastic chance constrained programming model for water resources management under uncertainties. Scientific Programming, 2019(1):1680813:1– 1680813:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2017/1680813.

Zima:2007:FLA

[Zim07] Hans P. Zima. From FORTRAN 77 to localityaware high productivity languages for peta-scale computing. Scientific Programming, 15(1):45– 65, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Zhang:2023:MAI

[ZJ23] Weihua Zhang and Yufeng Jia. Modern art interactive design based on artificial intelligence technology. Scientific Programming, 2023(1):5223034:1– 5223034:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5223034.

Zhang:2016:ROR

[ZJZ^{+}16] Yong Zhang, Yunjian Jiang, Miner Zhong, Nana Geng, and Dandan Chen. Robust optimization on regional WCOfor-biodiesel supply chain under supply and demand uncertainties. Scientific Programming, 2016(??): 1087845:1–1087845:15, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
1087845/.

Zhang:2023:RVA

[ZJZ23] Jian-Guo Zhang, Jian Jiang, and Rui Zhang. Recreational value assessment of urban productive landscape of baguatian in hangzhou based on contingent valuation method (CVM) and cloud computing. Scientific Programming, 2023(1): 9276457:1–9276457:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9276457.

Zhu:2023:OCM

[ZKA23] Meng-Xi Zhu, In-Jae Kim, and Zhi-Quan An. Optimizing the construction of multidimensional system of entrepreneurship education from the perspective of the second classroom. Scientific Programming, 2023(1):2344527:1– 2344527:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/2344527.

Zhou:2022:PEA

[ZL22] Zhangkai Zhou and Yihan Li. Parameter estimation of attribute scattering center based on water wave optimization algorithm. Scientific Programming, 2022(1): 6733510:1–6733510:??, 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6733510.

Zhang:2023:TGM

[ZL23a] Lin Zhang and Xinquan Liu. Traffic game model with the contract model. Scientific Programming, 2023(1): 6189075:1–6189075:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6189075.

Zhang:2023:AST

[ZL23b] Yingxue Zhang and Zhe Li. Automatic synthesis technology of music teaching melodies based on recurrent neural network. Scientific Programming, 2023(1):1704995:1– 1704995:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1704995.

Zhao:2023:ICM

[ZL23c] Haiyan Zhao and Shuangxi Li. Intelligent classification method of low occupancy big data based on grid index. Scientific Programming, 2023(1): 7965297:1–7965297:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7965297.

Zhang:2020:PPM

[ZLCW20] Zhaohui Zhang, Qiuwen Liu, Ligong Chen, and Pengwei Wang. A peak prediction method for subflow in hybrid data flow. Scientific Programming, 2020(1):2548351:1– 2548351:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/2548351.

Zhou:2021:DBR

[ZLDX21] Shuyue Zhou, Xiaobo Li, Yihong Dong, and Hao Xu. A decoupling and bidirectional resampling method for multilabel classification of imbalanced data with label concurrence. Scientific Programming, 2021(1): 8829432:1–8829432:??, 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8829432.

Zhang:2018:CCT

[ZLF18] Junwei Zhang, Deyu Li, and Xiaoqin Fan. A customer-centric trust evaluation model for personalized service selection. Scientific Programming, 2018(1):4819195:1– 4819195:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2018/4819195.

Zhu:2022:RAI

[ZLF^{+}22] Jiaming Zhu, Zhi Li, Jing Fu, Fangcheng He, Xiaoling Mou, and Pengjv Wu. RETRACTION: Application of Internet of Things technology in student management evaluation system. Scientific Programming, 2022(1):9806369:1– 9806369:??, ???? 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2024/9806369.

Zhang:2023:BNN

[ZLHZ23] Xiaoqing Zhang, Yinsu Lou, Sunhao Hu, and Dan Zhu. Backpropagation neural network algorithm-based color Doppler ultrasound detection of gestational diabetes mellitus and perinatal outcomes. Scientific Programming, 2023(1): 4258360:1–4258360:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4258360.

Zhang:2015:WCL

[ZLL^{+}15] Wenzhe Zhang, Kai Lu, Mikel Luján, Xiaoping Wang, and Xu Zhou. Write-combined logging: An optimized logging for consistency in NVRAM. Scientific Programming, 2015(??):398369:1–398369:13, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
398369/.

Zhao:2017:SPT

[ZLL^{+}17] Wenbing Zhao, Xiong Luo, HuaPing Liu, Kun Hua, and Chaomin Luo. Scientific programming towards a smart world. Scientific Programming, 2017(1):3706232:1–3706232:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/3706232.

Zhou:2020:GTG

[ZLL20] Yufeng Zhou, Yufeng Li, and Zhi Li. A grey target group decision method with dual hesitant fuzzy information considering decisionmaker’s loss aversion. Scientific Programming, 2020(1):8930387:1–8930387:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2020/8930387.

Zhang:2020:FBC

[ZLLS20] Feng Zhang, Guofan Li, Cong Liu, and Qian Song. Flowchart-based crosslanguage source code similarity detection. Scientific Programming, 2020 (1):8835310:1–8835310:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8835310.

Zhang:2020:FCG

[ZLLZ20] Feng Zhang, Lulu Li, Cong Liu, and Qingtian Zeng. Flow chart generation-based source code similarity detection using process mining. Scientific Programming, 2020(1):8865413:1– 8865413:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8865413.

Zhang:2017:FEM

[ZLS^{+}17] Yan Zhang, Xiang Liu, Yunyu Shi, Yunqi Guo, Chaoqun Xu, Erwen Zhang, Jiaxun Tang, and Zhijun Fang. Fashion evaluation method for clothing recommendation based on weak appearance feature. Scientific Programming, 2017(1): 8093057:1–8093057:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/8093057.

Zhang:2023:RSC

[ZLSW23] Fujun Zhang, Aichuan Li, Jianfei Shi, and Dongxin Wang. Research on school classroom teaching model based on clustering algorithm and fuzzy control. Scientific Programming, 2023(1):7435202:1– 7435202:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7435202.

Zhu:2021:MLP

[ZLT^{+}21] Fubao Zhu, Xiaonan Li, Haipeng Tang, Zhuo He, Chaoyang Zhang, Guang-Uei Hung, Pai-Yi Chiu, and Weihua Zhou. Machine learning for the preliminary diagnosis of dementia. Scientific Programming, 2021(1):5629090:1– 5629090:??, 2021. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2020/5629090.

Zhang:2023:ODC

[ZLWW23] Diandian Zhang, Yan Liu, Zhuowei Wang, and Depei Wang. OCR with the deep CNN model for ligature scriptbased languages like Manchu. Scientific Programming, 2023(1): 5520338:1–5520338:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5520338.

Zhu:2023:AEB

[ZLX^{+}23] Xinming Zhu, Haiyan Liu, Qing Xu, Junńan Liu, and Xiaoyang Lihua. Advances in an event-based spatiotemporal data modeling. Scientific Programming, 2023 (1):3532845:1–3532845:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3532845.

Zou:2021:DVC

[ZLZ^{+}21] Mingyang Zou, Junjie Liao, Yurong Zeng, Qianwen Guan, and Bowen Lan. Diagnostic value of CT angiography combined with high-resolution magnetic resonance angiography in vascular lesions in acute stroke. Scientific Programming, 2021(1): 2274443:1–2274443:??, 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2274443.

Zhang:2022:PFU

[ZLZ^{+}22] Hongbin Zhang, Hezhou Li, Xin Zhao, Juan Wu, Xiao Liang, and Haiyan Lu. Pelvic floor ultrasound under particle swarm intelligent optimization algorithm in preoperative and postoperative evaluation of female stress urinary incontinence. Scientific Programming, 2022(1): 6517725:1–6517725:??, 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6517725.

Zhao:2023:SMF

[ZLZ^{+}23] Haihong Zhao, Bo Li, Xiaonian Zhang, Xiaolin Guo, and Lixia Qian. Study on memory function of stroke patients under exercise relearning based on DWI image analysis based on optimized registration algorithm. Scientific Programming, 2023(1): 1536515:1–1536515:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1536515.

Zhang:2023:EEA

[ZM23] Bingtao Zhang and Lingyan Meng. Energy efficiency analysis of wireless sensor networks in precision agriculture economy. Scientific Programming, 2023(1):8346708:1– 8346708:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8346708.

Zhang:2018:SPM

[ZMLZ18] E. Zhang, Qihuang Mei, Ming Liu, and Feifeng Zheng. Stowage planning in multiple ports with shifting fee minimization. Scientific Programming, 2018(1):3450726:1– 3450726:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/3450726.

Zhao:2023:LNP

[ZMM^{+}23] Yan Zhao, Wenjuan Ming, Xiaohuan Ma, Yuchan Zheng, and Yanli Yan. LncRNA NKILA promotes cardiomyocytes apoptosis by targeting miR22-3p-TXNIP signal axis to inhibit proliferation, migration, and invasion of cardiomyocytes under high glucose-induced condition. Scientific Programming, 2023(1):6626845:1– 6626845:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6626845.

Zhu:2019:DAA

[ZPA19] Yuxiang Zhu, Yanjun Peng, and Arsineh Boodaghian Asl. Dual adaptive adjustment for customized garment pattern. Scientific Programming, 2019 (1):8069373:1–8069373:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/8069373.

Zhang:2022:RAE

[ZQ22] Lei Zhang and Liefeng Qiu. RETRACTION: Aerobic exercise fatigue detection based on spatiotemporal entropy and label technology. Scientific Programming, 2022 (1):9815392:1–9815392:??, ???? 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL https:/

/onlinelibrary.wiley.
com/doi/epdf/10.1155/
2024/9815392.

Zheng:2023:APS

[ZQL23] Dawei Zheng, Chao Qin, and Peipei Liu. Adaptive particle swarm optimization algorithm ensemble model applied to classification of unbalanced data. Scientific Programming, 2023(1):7589756:1– 7589756:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7589756.

Zaki:2008:BDM

[ZRP08] Mohammed J. Zaki, Naren Ramakrishnan, and Srinivasan Parthasarathy. Biological data mining. Scientific Programming, 16(1):3, ???? 2008. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Zhang:2024:CIA

[ZRX^{+}24] Ying Zhang, Mo Ruan, Yongqing Xu, Jianfei Jiang, Xinneng Li, and Hongbo Tan. Cartilage injury after patellar dislocation surgery and evaluation of the effect of platelet-rich plasma treatment by algebraic reconstruction techniques based MRI. Scientific Programming, 2024(1): 8808310:1–8808310:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8808310.

Zhang:1999:CEC

[ZS99] Xiaodong Zhang and Lin Sun. Comparative evaluation and case studies of shared-memory and data-parallel execution patterns. Scientific Programming, 7(1):1–19, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=64cr5a4mg33tuhcbdr02%
26referrer=parent%26backto=
issue%2C1%2C7%3Bjournal%
2C8%2C9%3Blinkingpublicationresults%
2C1%2C1.

Zhen:2023:RCN

[ZS23] Lihua Zhen and Xiaoqi Sun. The research of convolutional neural network based on integrated classification in question classification. Scientific Programming, 2023(1):4176059:1– 4176059:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4176059.

Zhang:2022:AQP

[ZSD22] Lina Zhang, Yu Sang, and Donghai Dai. Accurate quaternion polar harmonic transform for color image analysis. Scientific Programming, 2022(1):7162779:1– 7162779:??, 2022. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/7162779.

Zhou:2010:ABD

[ZSS^{+}10] Min Zhou, Onkar Sahni, Mark S. Shephard, Christopher D. Carothers, and Kenneth E. Jansen. Adjacencybased data reordering algorithm for acceleration of finite element computations. Scientific Programming, 18(2):107–123, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Zhang:2023:PVC

[ZSZ^{+}23a] Xianrong Zhang, Muhammad Shafiq, Guijun Zheng, Junping Wan, and Zhe Sun. Premature ventricular contractions’ detection based on active learning. Scientific Programming, 2023(1):5556011:1– 5556011:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5556011.

Zheng:2023:AEC

[ZSZ^{+}23b] Tao Zheng, Guofeng Shao, Qingyun Zhou, Qinning Wang, and Mengmeng Ye. Abdominal enhanced computed tomography image by artificial intelligence algorithm in the diagnosis of abdominal aortic aneurysm. Scientific Programming, 2023(1): 8721464:1–8721464:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8721464.

Zhu:2023:MLB

[ZSZ23c] Jianbin Zhu, Xiaojun Shi, and Shuanghua Zhang. Machine learning-based grammar error detection method in English composition. Scientific Programming, 2023(1): 4213791:1–4213791:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4213791.

Zhou:2023:CTI

[ZTL^{+}23] Chunfang Zhou, Shufang Tian, Fei Lv, Rui Shang, and Xuejiao Zheng. Computerized tomography imaging omics under iterative reconstruction algorithm in diagnosis of gastric cancer. Scientific Programming, 2023(1):2987080:1– 2987080:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2987080.

Zhu:2018:BDM

[ZTLH18] Yueqin Zhu, Yongjie Tan, Xiong Luo, and Zhijie He. Big data management for cloud-enabled geological information services. Scientific Programming, 2018(1):1327214:1– 1327214:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/1327214.

Zheng:2018:PRB

[ZW18] Jianxing Zheng and Yanjie Wang. Personalized recommendations based on sentimental interest community detection. Scientific Programming, 2018(1):8503452:1– 8503452:??, 2018. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2018/8503452.

Zhang:2023:SIC

[ZW23a] Dawei Zhang and Daling Wang. Study on the intentional choice mechanism of course selection based on swarm intelligence algorithm. Scientific Programming, 2023 (1):5510039:1–5510039:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5510039.

Zhou:2023:RRP

[ZW23b] Lina Zhou and Chunxia Wang. Research on recommendation of personalized exercises in English learning based on data mining. Scientific Programming, 2023(1): 5042286:1–5042286:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5042286.

Zhou:2023:ASP

[ZW23c] Lu Zhou and Huiling Wang. An approach to study the poverty reduction effect of digital inclusive finance from a multidimensional perspective based on clustering algorithms. Scientific Programming, 2023(1): 4645596:1–4645596:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/4645596.

Zhao:2016:BRS

[ZWCY16] Yang Zhao, Lei Wang, Xue Chen, and Futao Yang. Bulk restoration for SDN-based transport network. Scientific Programming, 2016(??): 1938312:1–1938312:8, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
1938312/.

Zhao:2023:RIE

[ZWF23] Hua Zhao, Aibo Wang, and Ying Fan. Research on the identification and evaluation of aerobics movements based on deep learning. Scientific Programming, 2023(1):6433260:1– 6433260:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6433260.

Zhang:2020:STP

[ZWG^{+}20] Zhe Zhang, Cheng Wang, Yueer Gao, Jianwei Chen, and Yiwen Zhang. Shortterm passenger flow forecast of rail transit station based on MIC feature selection and ST-LightGBM considering transfer passenger flow. Scientific Programming, 2020(1):3180628:1– 3180628:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/3180628.

Zhou:2023:MDA

[ZWJN23] Jian Zhou, Zhuping Wang, Yingjie Jiao, and Cong Nie. Material discrimination algorithm based on hyperspectral image. Scientific Programming, 2023(1): 8329974:1–8329974:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8329974.

Zhang:2019:ESA

[ZWL^{+}19] Chi Zhang, Yuxin Wang, Yuanchen Lv, Hao Wu, and He Guo. An energy and SLA-Aware resource management strategy in cloud data centers. Scientific Programming, 2019(1):3204346:1– 3204346:??, 2019. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/3204346.

Zhu:2023:DLB

[ZWLL23] Chen Zhu, Bin Wang, Lei Li, and Tianzuo Li. Deep learning-based CT images in pulmonary function assessment of patients who underwent laparoscopic surgery under guidance of electrical impedance tomography. Scientific Programming, 2023(1):9889488:1– 9889488:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9889488.

Zhang:2019:LCT

[ZWS19] Qiuwen Zhang, Shuaichao Wei, and Rijian Su. Low-complexity texture video coding based on motion homogeneity for 3D-HEVC. Scientific Programming, 2019(1): 1574081:1–1574081:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/1574081.

Zhang:2016:ODR

[ZWW16] Jianhua Zhang, Shuliang Wang, and Yixing Wang. Optimal design on robustness of scale-free networks based on degree distribution. Scientific Programming, 2016(??): 5659687:1–5659687:7, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL https://www.hindawi.
com/journals/sp/2016/
5659687/.

Zhang:2023:ARI

[ZWW^{+}23a] Ziwen Zhang, Xuelian Wang, Yongdong Wu, Zengpeng Zhao, and Yang E. Applied research on In-SAR and GPS data fusion in deformation monitoring. Scientific Programming, 2023(1):3888975:1– 3888975:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3888975.

Zong:2023:DXR

[ZWW^{+}23b] Yinghong Zong, Limei Wu, Dongping Wu, Jing Jiang, and Shimei Yang. Dual X-ray image under neural network adopted in comparison of efficacy and safety of deep hydrolyzed protein milk powder and parenteral nutrition in intervention of neonatal noninfectious abdominal distension. Scientific Programming, 2023(1):6872291:1– 6872291:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6872291.

Zheng:2021:HAM

[ZWXL21] Feifeng Zheng, Zhaojie Wang, Yinfeng Xu, and Ming Liu. Heuristic algorithms for MapReduce scheduling problem with open-map task and seriesreduce tasks. Scientific Programming, 2021(1): 8810215:1–8810215:??, 2021. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8810215.

Zheng:2016:RAS

[ZWZ16] Wei Zheng, Chen Wang, and Dongzhan Zhang. A randomization approach for stochastic workflow scheduling in clouds. Scientific Programming, 2016(??):9136107:1–9136107:13, ???? 2016. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2016/
9136107/.

Zhao:2020:ACS

[ZWZ20] Jinchao Zhao, Yihan Wang, and Qiuwen Zhang. Adaptive CU split decision based on deep learning and multifeature fusion for h.266/ VVC. Scientific Programming, 2020(1):8883214:1– 8883214:??, 2020. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2020/8883214.

Zhao:2023:FCS

[ZWZ23a] Jinchao Zhao, Yihan Wang, and Qiuwen Zhang. Fast CU size decision method based on just noticeable distortion and deep learning. Scientific Programming, 2023(1): 3813116:1–3813116:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

wiley.com/doi/epdf/10.
1155/2021/3813116.

Zhu:2023:MMP

[ZWZ^{+}23b] Yingying Zhu, Haiyan Wu, Dawei Zhou, Qiushi Wang, Zhendi Sun, Chuanzhen Niu, Liping Ye, Xinghan Tian, and Chunting Wang. A multivariate model for predicting the progress of COVID-19 using clinical data besides chest CT scan. Scientific Programming, 2023(1):3432010:1– 3432010:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3432010.

Zhong:2017:REH

[ZX17] Hongye Zhong and Jitian Xiao. Retracted: Enhancing health risk prediction with deep learning on big data and revised fusion node paradigm. Scientific Programming, 2017(1): 9757658:1–9757658:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/9757658.

Zhang:2019:MSA

[ZXDY19] Gongjie Zhang, Chunli Xie, Yongquan Dong, and Qiao Yu. Mutant selecting according to the nondominated original statements. Scientific Programming, 2019(1): 4702345:1–4702345:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2019/4702345.

Zhang:2017:LMD

[ZXHQ17] Hao Zhang, Ying Xiong, Mingke He, and Chongchong Qu. Location model for distribution centers for fulfilling electronic orders of fresh foods under uncertain demand. Scientific Programming, 2017(1): 3423562:1–3423562:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/3423562.

Zhang:2023:MSK

[ZXS^{+}23] Hongbin Zhang, Qipeng Xiong, Haowei Shi, Weinan Liang, Xiang Zhong, Zhiliang Zhu, Xiong Li, and Donghong Ji. Mining sufficient knowledge via progressive feature fusion for efficient material recognition. Scientific Programming, 2023(1):8971349:1– 8971349:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/8971349.

Zheng:2023:SPS

[ZXX^{+}23] Weiguang Zheng, Weiwei Xin, Enyong Xu, Shuilong He, Jirong Qin, and Heng Wang. Sizing of the propulsion system for a heavy-duty fuel cell commercial vehicle. Scientific Programming, 2023(1):1497178:1– 1497178:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/1497178.

Zhang:2023:EAE

[ZXZZ23] Yiqun Zhang, Lu Xue, Chunlian Zhang, and Jianying Zhou. Evaluation of the application effect of contrast-enhanced ultrasound image technology based on threedimensional image fusion algorithm in the diagnosis of adenomyosis. Scientific Programming, 2023(1): 2896752:1–2896752:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/2896752.

Zhao:2023:MRI

[ZY23] Qingyan Zhao and Shuai Yuan. Magnetic resonance imaging (MRI) based on machine learning algorithms for the diagnosis in efficacy of dexmedetomidine along with modified electroconvulsive therapy nursing on first episode schizophrenia. Scientific Programming, 2023(1): 1897571:1–1897571:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/1897571.

Zheng:2024:MMD

[ZY24] Bing Zheng and Jing Yang. Measurement method of distributed nodes in wireless sensor networks based on multiple attributes. Scientific Programming, 2024(1): 9936337:1–9936337:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9936337.

Zhao:2023:RBM

[ZYL23] Jingyuan Zhao, Liyan Yu, and Zhuo Liu. Research based on multimodal deep feature fusion for the auxiliary diagnosis model of infectious respiratory diseases. Scientific Programming, 2023(1):5576978:1– 5576978:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5576978.

Zeng:2015:RIF

[ZYLZ15] Yi Zeng, Shiqun Yin, Jiangyue Liu, and Miao Zhang. Research of improved FP-growth algorithm in association rules mining. Scientific Programming, 2015 (??):910281:1–910281:6, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.
com/journals/sp/2015/
910281/.

Zhao:2021:IPM

[ZYX^{+}21] Shuo Zhao, Qiliang Yang, Jianchun Xing, Qizhen Zhou, Guangtong Xue, and Wenjie Chen. INR: a programming model for developing APPs of insect intelligent building. Scientific Programming, 2021(1):3659849:1– 3659849:??, 2021. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.

                                        wiley.com/doi/epdf/10.
                                        1155/2020/3659849.

Zhang:2023:DDL

[ZYY^{+}23] Yongzhao Zhang, Jianshi Yin, Han Yan, Jun Liu, and Junsheng Wang. Denoising of degenerative lumbar spine lesions MRI images using blockmatching and 3D filtering. Scientific Programming, 2023(1):2430380:1– 2430380:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/2430380.

Zhang:2023:ASA

[ZYZ23] Jinling Zhang, Jun Yang, and Min Zhao. Automatic segmentation algorithm of magnetic resonance image in diagnosis of liver cancer patients under deep convolutional neural network. Scientific Programming, 2023(1):4614234:1– 4614234:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/4614234.

Zhang:2023:SSC

[ZYZL23] Bo Zhang, Lan Yang, Meng Zhang, and Yanhui Li. Sales strategy considering advertising in advance-selling and spot-selling integration mode for fresh product. Scientific Programming, 2023(1):9181912:1– 9181912:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9181912.

Zhong:2023:REE

[ZZ23] Chongxia Zhong and Jie Zhu. Rehabilitation effect of exercise therapy on knee osteoarthritis evaluated by computed tomography image under optimized reconstruction algorithm. Scientific Programming, 2023(1): 3580360:1–3580360:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3580360.

Zhang:2023:SMA

[ZZD23a] Lifen Zhang, Qiang Zhou, and Huifang Ding. Simulation and modeling algorithm for terminal container handling intelligent management based on Internet of Things and big data technology. Scientific Programming, 2023(1):8667945:1– 8667945:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/8667945.

Zhong:2023:REM

[ZZD^{+}23b] Xin Zhong, Gang Zheng, Min Ding, Jing-Bing Li, and Qi-Hong Qian. The role of endothelial mesenchymal transformation on infantile hemangioma. Scientific Programming, 2023(1): 5511095:1–5511095:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5511095.

Zheng:2024:IUN

[ZZDY24] Ruiyong Zheng, Yongguo Zheng, and Changlei Dong-Ye. Improved 3D U-Net for COVID-19 chest CT image segmentation. Scientific Programming, 2024(1):9999368:1– 9999368:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/9999368.

Zhu:2022:RBT

[ZZH22] Gaibian Zhu, Qiuyan Zhang, and Sujuan He.

Relationship between twinto-twin selective intrauterine growth restriction with sFas/sFasL level of umbilical cord blood using Doppler ultrasound fetal heart rhythm detection algorithm. Scientific Programming, 2022(1): 6256562:1–6256562:??, 2022. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6256562.

Zou:2024:EWL

[ZZJ^{+}24] Yi Zou, Jijuan Zhong, Zhihao Jiang, Hong Zhang, and Xuyu Pu. Experience weighted learning in multiagent systems. Scientific Programming, 2024(1):9948156:1– 9948156:??, 2024. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9948156.

Zheng:2023:KPI

[ZZS^{+}23] Mengze Zheng, Islam Zada, Sara Shahzad, Javed Iqbal, Muhammad Shafiq, Muhammad Zeeshan, and Amjad Ali. Key performance indicators for the integration of the serviceoriented architecture and scrum process model for IOT. Scientific Programming, 2023(1):6613579:1– 6613579:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/6613579.

Zhu:2023:IRP

[ZZSQ23] Dongdong Zhu, Honglei Zhang, Yulong Sun, and Haijie Qi. Injury risk prediction of aerobics athletes based on big data and computer vision. Scientific Programming, 2023(1):5526971:1– 5526971:??, 2023. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2021/5526971.

Zhen:2017:OMA

[ZZW17] Lu Zhen, Si Zhang, and Xinchang Wang. Optimization models and algorithms for services and operations management. Scientific Programming, 2017(1):4216785:1– 4216785:??, 2017. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1155/2017/4216785.

Zhou:2022:SRE

[ZZW22] Hang Zhou, Xinying Zhu, and Jian Wang. A specific risk evaluation system for live virtual machine migration based on the uncertain theory. Scientific Programming, 2022(1):6784419:1– 6784419:??, 2022. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6784419.

Zhang:2023:TDR

[ZZW^{+}23] Zhongxing Zhang, Yan Zhang, Xiaohui Wang, Xiaoli Han, Xin Zhang, and Hong Chen. Threedimensional reconstruction algorithm for CT pulmonary angiography in patients with pulmonary embolism combined with syncope. Scientific Programming, 2023(1): 5268650:1–5268650:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/5268650.

Zhao:2024:MAD

[ZZWL24] Hongwei Zhao, Danyang Zhang, Jiaxin Wu, and Pingping Liu. Multiproxies adaptive distribution loss with weakly supervised feature aggregation for fine-grained retrieval. Scientific Programming, 2024(1): 9945840:1–9945840:??, 2024. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/9945840.

Zhang:2023:AMB

[ZZWZ23] Xinchen Zhang, Qincheng Zhou, Shijie Weng, and Hui Zhang. ARIMA model-based fire rescue prediction. Scientific Programming, 2023(1): 3212138:1–3212138:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/3212138.

Zhu:2017:ILK

[ZZX^{+}17] Yueqin Zhu, Wenwen Zhou, Yang Xu, Ji Liu, and Yongjie Tan. Intelligent learning for knowledge graph towards geological data. Scientific Programming, 2017(1): 5072427:1–5072427:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2017/5072427.

Zhang:2023:LRM

[ZZZ^{+}23] Daigui Zhang, Lihua Zhou, Tingdi Zhang, Shuai Wang, and Yue Li. Low-rank matrix denoising algorithm-based magnetic resonance imaging combined with computed tomography images in the diagnosis of cerebral aneurysm. Scientific Programming, 2023(1): 6191230:1–6191230:??, 2023. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://onlinelibrary.
wiley.com/doi/epdf/10.
1155/2021/6191230.